A Complete Bibliography of Publications in the Journal of Theoretical Biology: 1980–1989

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

01 June 2019 Version 1.00

Title word cross-reference

−3/2 [Pah85a]. 1 [Aki85]. 1/F [SS85]. 10 [Sto89b]. 2 [Aki85, BSR^{+}88, KSH87]. 2μ [Fut86, WB88]. 5 [BVDA87]. 57 [UN82]. 6 [KSH87]. 8 [SRO85]. ^{+} [Fri85, HF88, NA84, Ped88]. ^{−} [NK88]. ^{14} [Law84]. ^{2+} [AH86, BRW89, BLR83, BH83, FK83, Fri85, JS80, KK83c, KT81, Szu83, TY82]. 31 [DDA85]. _{0} [BW87]. _{0}.5 [Yos86]. _{1} [EP86, HK80, KS81, KK81, Was86a]. _{1β} [KS81, KK81]. _{2} [BT85, PRH81]. _{2α} [KS81, KK81]. _{a} [UN82]. _{M1} [WR86]. _{M2} [WR86]. _{o}^{+} [Ped88]. A_{L} [PWH^{+}88]. α [AL83, BD80, CH82, Jea85, Kot83, LSFT85, MATGR87, SKK84, SK84, SLFT86, VM82, WMR86, Yok83]. β [BRM82, HSd^{+}88, Mur85, Mur87, NSKS85, Syv84, UN82, WMR86]. d [Mor86]. D = 0 [Has81]. D ≠ 0 [Has81]. F [BM87, BMR89]. g_{1}(τ) [LF86]. g_{2}(τ) [LF86]. γ [All85b, BLR83, Osu84]. h [Cha80b, LP82]. K [LVN87]. K_{m} [Bar80a, Fel80]. l [Ven84]. λ [WS85b]. m [Cha80b, LP82]. N [FGDS87, Cha80b, LP82, Szu83]. ϕ [Dug81, OK82]. × [Rei84]. → [VM82, WMR86]. V_{m} [Bar80a].

-amino [Mor86, Ven84]. -aminobutiric [BLR83]. -Asp [UN82]. -ATPase [FK83, Fri85]. -azaguanine [SRO85]. -chloro- [Aki85]. -chymotrypsin [LSFT85, SLFT86]. -compartment [FGDS87]. -crystallin [CH82]. -D-dihydrosphingosine [WMR86]. -D-galactose [WMR86]. -D-glucosyl-N- [WMR86]. -D-N0acetylneuraminyl- [WMR86]. -furylethylenes [BSR^{+}88]. -gangliosides [WR86]. -globin [Yok83]. -globins [Syv84]. -glutamyl [All85b, Osu84]. -helices [SK84]. -helix [AL83, VM82]. -heltical [BD80]. -Isoxazolol [BVDA87]. -kinase [KSH87]. -kinase/fructose [KSH87]. -lactalbumins [MATGR87]. -lactam [HSd^{+}88]. -methyl [Kot83]. -phenanthroline [Sto89b]. -phosphate [KSH87]. -phosphofructo- [KSH87]. -polypeptide [Jea85]. -Pump [Ped88]. -sheet [BRM82, NSKS85]. -test [BMR89]. -trypsin [UN82]. -turn [Mur85]. -turns [Mur87].

1-277 [Ano81-36]. 12th [Mur85]. 15- [Kot83]. 179 [Ano80j]. 18s [MDPF86].

2 [KK83c]. 2-D-hydroxyoctadecanoyl [WMR86].

3 [Iid88, MDPF86]. 3- [BVDA87]. 3’-splice [Iid88]. 3’-terminus [MDPF86].

5 [Fel80, MDPF86]. 5’-cyclic [Fel80]. 5’-Leader [MDPF86]. 546 [Phi82]. 5S [JC80].

6-bisphosphatase [KSH87]. 61st [Mur85].

76 [Ano80j].

93 [Phi82].

A.C. [Bla80]. Ab-initio [PR82]. abasic [Kyp88]. aberrations [Cra80a, KM81]. ability [FF83, Reg88, Sim85, TDA86]. abnormal [KN83]. abnormality [Was86a]. abortificient [KS81, KK81, Kot82, KK83c, Kot83]. absence [GW88]. Absolute [Kra82, SSM81, MB85]. absorbance [ZG82]. Absorbent [DC87a, SBPF83]. absorption [AHFH81, FGDS87, HF85, MHII83, SC84b, YMIH86]. absorptive [Win89b]. abundance [And80a, GP82c, Jak88, Rou80, Tat88]. Acanthaster [McC87]. acceleration [Kri80]. acceptor [Gho82, MJN88, NSW82, NW85]. acceptors [SNW81]. access [TCR85]. accessible [GVP^{+}80, Sta86, SPG87]. Accession [MB85]. accident [TIK89b]. accidental [Ale84]. according [BBS88]. account [Bel80]. accretion [Car83]. accumulating [Oht87]. accumulation [CC83b, Swa82, dBH87]. Accuracy [Sav81, TC80b, FS80, LB87, SR87b]. accurate [RMS81]. Acetabularia [BG88, GT85a, HSV84, HH85a]. acetyl [YBBL82b]. acetylation [Spe84]. acetylcarnitine [MRR80]. acetylcholine [DSR89, RTM82, TS85]. acetylglycine [RTM81b]. acid [Arg81, BS88a, BS85, BLR83, CC82b, CC83a, Cla82d, CF82, CH82, För88, Goo84, HT87, HR85, Kum82, LVN87, LTZ88, MLB88, McG80, McG81, McG84, MT84, Mor86, MF85, NU82, NMM84, PP80, RB82b, RB83a, Sco82, See82, SK84, Tay86, TDA86, WFL86, dW81b]. acids [Bre88, CV81, Frö84, HSd^{+}88, Mik81, Mik83, Mur85, Mur87, MRR80, Osu84, Ven84]. acoustic [Fle88]. acquired [BKMS86, BKMS87, BR83, JL89]. Acrididae [HLS85]. across [AWP80, HGG82, Ipi87, OK87b, Osu84, TMGB80]. actin [Gre85, MB88a]. actin-myosin [MB88a]. actinomycin [Mil87]. action [Bar81d, BHM^{+}85, CY84, Gol83b, Gre80, HBRN81, Hal80b, KS81, KK81, Kot82, KK83c, Kot83, Kum82, Mar80a, RT85, Ros84c, SRO84, She80, She81, Shi81, WK81, ZgQy84]. activated [IHK80, ML84, Poz87]. activating [VCC^{+}88]. Activation [Gar87, And88, CVNO87, KK85, Kri85a, Kri85b, Mad88, MUI85, NDB83, NMB85, Rao82, TS85, VCC^{+}88]. activator [Tys83b, WS82]. Active [ER84, Jan82, AH86, Fle80, Hig85, MdMF83, SH80a, Sab80, SBW83, UN82]. activities [BSR^{+}88, RP84]. activity [Bar88, CVNO87, CC83b, DSW83, GK81, GB87, JL88, KLK^{+}88, KHU^{+}83, KM85, KK86a, ML86, MHB83, Oon82, PTD81, Rak80, RSSK87, Sco84a, She81, WT87, WS81, Woo80]. Acyl [GBS89]. acylenzyme [GBS89]. acylenzyme-nucleophile [GBS89]. ad [FBP87]. adair [Whi80b, Whi81b]. Adaptation [BS88b, Pin84, Bro81b, Dus89a, Mad88, MT81, SGDK86, Stu85b]. adapting [MM80]. Adaptive [CK82, PC83, HT84, Jam89c, KL87, Lib88, SCN85]. added [Law83]. Addendum [Ano80a, Ano81a, Ano84a]. additional [Gie88, SA80]. Additive [SGd84, Dom83]. Additivity [HR85]. address [Ste80, Ste82]. adenine [Kyp88, Nus80a]. adeno [RMFM85]. adeno-associated [RMFM85]. adenoma [All85a]. Adenosine [Szu83, Fel80, ZBL80a, ZBL80b]. adenylate [Blu84, Kre82]. adenylates [TKB80]. adherence [PJK^{+}83]. Adhesion [Toz89, Bir81, BP87, HW86, HW87, Hsu87a, Hsu87b, Hsu87d, Hsu87c, Hsu87e, MM83d, SG82, Vro83]. adhesive [MCB^{+}86]. adjacent [VDH82]. adjustable [HT89c]. adjustment [HSV84, TC87]. administering [MSBB84]. admitting [KK86b]. ADP [ZBL80b, Sto83, Wil84]. ADP-ribose [Sto83]. adrenergic [KK86a]. adsorbed [WP81]. Adsorption [Kur84a, VA88a, AWV^{+}88, Fed80, Kol82, TJP83, WWN82]. adsorptive [Rak86]. adult [Hof83]. advantage [BP86, Cha86, Gen88, Gri83, Lay87, Man82a, Man84]. advantageous [CB81]. Advantages [Sir86, vHS82, Gro82, TH81]. adversity [SG86]. aerobic [CF87b]. Aerobiosis [YC89]. Aerodynamic [Hum83, WS84]. affect [Ber87, FDC85]. affected [HM80, Yok84]. affecting [Mur85]. affects [BT85]. affinities [Jak88, KK83c]. affinity [BGG84, SH85b]. after [AK86, BC81b, BC83, CG85a, Gal84a, Gal84b, HSV84, Har87b, MH81, OWN89]. after-effects [CG85a]. against [Col85, For85, MH80d, MC89, vdHN85, Orl82, SM88b, Vic87]. age [Abu86, AA83, Aug85, Cas82, Cas84, CC81, CT81a, CCDF81, Cus80, DO83, Ebe87, Ebe88, HCW80, Hir80, Hir82, KF84, Lib88, Man85, McN89, SVN89, ZL84]. age-class [Aug85]. age-dependent [Hir80, Hir82]. age-specific [Cas82, Cas84]. age-structured [CC81, Ebe87, McN89, SVN89, ZL84]. ageing [Bar81e, Dil86, HK81, HK83, Mor82, Rei83b, Uma82]. agent [VTL85]. agents [Ber85, Cru80, Hua80a, KK86a, RW82, YHJ89]. aggregates [TCR85]. Aggregation [Bri89, COG83, BE82, HU85, JK87, Kil80, PO81, TCR85]. aggression [Tre81a]. aging [Ada80, DCF80, Ric80, Wag89, vZ87]. agonism [BVDA87]. agonist [Lef86]. agonists [LH89]. aided [CMM86]. AIDS [LG89b, San88]. al. [AK86]. alamethicin [Bru85b]. alanine [YHJ89]. alarm [Tam87]. alarmone [Var83]. alga [PL85c]. algae [GF81b]. algal [Gor87, MVV89]. Algebraic [Rik85, Whi81b]. algorithm [LVN87]. algorithms [Wat84]. Alignment [Got86, Alt89, Wat84]. alkaline [NBZ^{+}83]. alkoxy [Ano81-37]. alkylated [DD89]. alkylating [BSR^{+}88, RW82, SBW83]. alkylation [MT84]. all-or-none [Wri80]. Allele [Gre84a, SFR86, KK86b, Orl81a]. alleles [Yok85]. allelic [VC88a]. Allocation [Cha89c, Ryt87, CC82a, CB85, CB86b, GC86, Pug87a, Tay85b]. allocations [Llo83]. allogeneic [OWN89]. allometric [Cal83a, Don86, PK82, Per84, RL84, Whi81a]. allometry [Cal87, Cha89b, Har82d, Jol89, PK85, PS81b, Rei85b, SS83b, Smi80, Smi84, TSRK88]. allosteric [KN82, KK82a, Kur83, SM88a, Str85c, WIN80, Whi80c, Yos86]. allosterism [Dun82]. allow [Boy89]. allowed [Whi80b, Whi81b]. allyl [GPM89]. almost [AM87]. along [ANS83, LP88b]. alpha [Cla83a, FY81]. alpha-helical [FY81]. alpha-helix [Cla83a]. also [MM86a]. Alsomitra [AO87]. altered [Sto89a]. alternans [AWM85]. alternation [Cha80e]. alternative [BR87, BJ82, Bel87, Cha81, HAC82, Loc82, Mic88, RL84, Sac88]. alternatives [Col85]. Altman [Ola86]. Altruism [Tor82, Eld83, Nee89b, Orl82, Rot80, Tam87, Tre82]. alu [BJ84]. alu-type [BJ84]. alveolar [Tro83]. alveoli [Col85]. alveolus [DGST82]. always [Cha86, Iwa88, Met82, SC86a, SR89]. ambient [DY82, Sea80]. ambiguities [AOB85]. ambush [LG85]. amide [RTM81b]. amiloride [MTM86b]. amines [TN80]. Amino [RB82b, Spe84, BS85, Bre88, CC82b, CC83a, Cla82d, CV81, CH82, Frö84, HT87, HR85, HSd^{+}88, LVN87, MLB88, Mor86, Mur85, Mur87, NU82, Osu84, PP80, RB83a, SK84, Tay86, TDA86, Ven84]. aminoacyl [Jak88, MKM80]. aminoacyl-tRNA [Jak88]. aminoacyl-tRNAs [MKM80]. aminobutiric [BLR83]. Ammonium [EG87a]. among [Bro81b, CW84, EPH86, FE85, Gri81a, Gri81b, Gri81c, Gri81d, Gri82b, Gri82c, Gri82d, Gri82e, Gri82a, Gri82f, Hai86, MM83b, MGL81, NDB83, PH87a, SS89a, Sve82, Vro83, WL87]. amount [AWV^{+}88]. AMP [ZBL80a]. amphetamine [Aki86]. amphetamines [Ano81-37]. amphibia [Bar81c, CPLA85]. amphibian [Pat84]. Amplification [FC81, Gra84, CA89, Fut86, MS89a, MAC^{+}83, OO84a, Tor80, WB88, RN85]. amplifiers [ER84]. amplitude [AZE83, RMS81, Vau86]. anaerobic [SSF80]. anaerobiosis [FQ81, YC89]. analgesics [GPM89]. analog [McC80a]. analogous [MG84a]. analogs [FK83, YHJ89]. analogues [ABS82].

Analogy [EPH86, Hof86]. analyse [MB88a]. Analyses [CZJY80, BW81, FW81, Fel80]. Analysis [AMM80, BN80, BLC86, BGG84, CCG^{+}82, CS84, CB86c, CA89, DG87, DDJ83, FBF85, FS85, FS87, GS81, Gol83a, HS81, HF88, KDAT84, LL88, MRG84, MK88, PWH^{+}88, SPS83, SK85b, SH82b, TO80, ATC81, ASK89, ADG80, AK82, Alm83, ASTT80, Ano80-35, AOB85, BM87, BJ82, BG86, BBS88, BB88b, Bro80b, Bro83, Bur83, BT81, CH80, CW88, CG81a, Cla82e, CLF84, Coo81, Dam86, DL85, DW82, Dug82, DRP^{+}87, FM88, FEC87b, FDC^{+}89, FS83, FK83, Fri85, Gar80, Get82, Gho82, Gie88, GBS89, GFA89, GM84, HU85, HW86, HW87, Hsu87e, HT88, HM85b, Iid88, Jaf80, JK89, JP85, Jøn85c, JR80, KHB83, KUT85, KT87a, KYU^{+}86, KDG85, KA89, KK87a, KK87b, Kil80, KHU^{+}83, KMSR84, KTG86, KT87b, KK82a, KC82, KL83, Kri85a, LD83b, Leo84]. analysis [LL87a, Mae84, Mah84b, Mar80b, MAC89, MHB83, MH81, MKM80, NFP88, PL85a, Pen82, Per82, Phi81b, PDCD88, RTM88, Rei83b, Ros84b, Sch80, SSW81, SM81b, Sim81a, SF89, SK87, SC89, SG82, SF83, SW84, TATS81, UKNS87, WMS81, Whi89b, WF82, YA89, YHJ89, ZL84, Zac81, Zam86, FP89, PF89]. Analytical [SDM^{+}82, Has81, KL80]. analyzed [LF86]. Analyzing [Koo85]. anatomical [CF87b]. ancestor [Cla82b]. ancestors [TH81]. anchor [Kur84a]. anchorage [Rap84]. and/or [Eco83]. anemotropic [PSG83]. angiosperms [RB82a, Sch80]. angles [WH86, ZC89]. animal [BSS83, BB88b, Cou80, Cze87, DD85b, DW87, Fel80, GR88, GM89, GJ88, Goy82, Hok86b, Kum82, Luf82, McA87, MH80b, Mur81, Sel80, SR81, Wel83, Wel86, WB83, Zee81, vRV80]. animal-pollinated [GR88]. animals [Ale87, Bro82b, CHT80, CPPR87, CPPRS87, Dob86, FCBK88, Har83, JR80, MB80, Niv83, Nor82, Smi82a, SR87b, SV89, Süh89, Tat88, Zam86]. animals-exploration [Süh89]. anion [Mon83]. anisogamous [KR86, Ros82]. anisogamy [Hoe80, Hoe84, Wie81]. anisotropic [YMS84]. anisotropy [DBB80]. Announcement [Ano80b, Ano81b, Ano82a, Ano86a, Ano87a, Ano80c]. Announcements [Ano88a]. annual [Bul81, MW81]. annuals [CM87]. anomalous [PCTF83, Rob87, Wol80]. anserine [Bro81b]. ant [IAHA80]. Antagonism [RN84c, BVDA87]. Antagonistic [Nee89a, SJ81]. Antarctic [LS81]. antedating [Lib85, Mas86, Wol89]. antenna [FBF85]. anterior [Kwo84]. Anti [Fra88]. Anti-sense [Fra88]. Antibacterial [BSR^{+}88, SSF87, Cru80]. antibiotics [HSd^{+}88, LH86, SRO84]. antibodies [Koo85, Wu88]. antibody [BJ89, BGG84, BF81, CSH^{+}81, CU88, GS81, GAD80, MW83a, SW87, SW81b]. anticancer [SBW83]. antico [MGL81]. antico-operative [MGL81]. anticodon [BS83, Mik81]. antigelling [YHJ89]. antigen [BJ89, BBH84, CU88, Fis80, GS81, Tho85, WP81]. antigen-antibody [BJ89]. antigen-heterogeneous [GS81]. antigen-presenting [Tho85]. antigenic [Frö88]. antigens [GAD80, SM88b]. antimetabolite [SRO85]. antiparallelism [Rei84]. antipredation [Kil80]. ants [DPV83, Her81a, Jaf80, Jaf84, LH83, Orl81b, Tay81]. anuran [CPLA85]. any [Eva80a]. aorta [Li88]. aortic [TLS83]. apical [MTM86a, MTM86b]. Apis [OL86, Omh87a, Omh87b]. APLYSIA [TG83]. apolar [Frö84]. apoproteins [PP85]. apoptosis [Suc83]. apparatus [Bou84, OCMP89]. Apparent [BHM^{+}85, BWM80, BT85, IHTN88, SH85b, SPG87]. appearance [IHTN88, KSH87]. appendages [CK87, Koe83]. Applicability [Fuk84, HAG82, VKV83, Amt86, Wil81b]. applicable [Lev80a]. Application [BB84, Bro80b, BKJT80, Dug82, HSd^{+}88, Kil80, MH81, RP84, Rob87, RSM80, SCH84b, SBPF83, BRRK87, Bar82a, BW81, BK88, Els83, FE85, FCBK88, HSV84, HR80b, HBP89, HH81, HM85b, KW89, KT87b, KS88, Kwo84, MH80b, Nag81, PS81a, PSL^{+}85, RTM81a, RW82, SSF80, Tan80, WB89, WD87, Zaj86, KK86a, Pap82, Sim80b]. Applications [AL81, BAB84, Cha89b, Uma82, BSB88b, BS84b]. applied [Asp88, Fer85b, Fri85, HK88, MB84b, Mul87, PWH^{+}88, Ras87, Rik82, TLN85]. applies [MG84a]. apply [Par85]. Applying [San88]. appraisal [BG87, Sib81]. Approach [HdL84, Abu86, Alm85, BJ82, Bor82, BKB80a, BKB80b, dCB84, Cal83a, CFLF88, CV89, CN84, DV83, Dug86, FT85, GTSNO^{+}88, GM89, Get84, GS85, HKC83, HS81, HC89, Hof83, HSd^{+}88, Ipi87, IP83, JD82, JA88, Leo84, LQC89, MSA^{+}81, MP85a, Mau87, Mic86, MV88a, Nun81, PGKE87, PGKE88, Peu83, PP80, PPC82, Red88, RMB81, SF81, ST86a, ST86b, SC83, SM87, TIK89a, TLT80, VM82, VDH82, VLCK86, WB85, Wax81, YC89, dVR85]. approaches [FFM82, Mor86, RS83]. approaching [Boo84b, LGK84]. approximate [Rat87]. approximation [Aal89, Bar80a, PSL^{+}85, SK84, TC80b]. approximations [Lán84b, SM80b, Tuc87]. April [Ano80x, Ano80-29, Ano81y, Ano81-32, Ano82y, Ano82-30, Ano83x, Ano83-32, Ano84-28, Ano84-34, Ano85-30, Ano85-36, Ano86-29, Ano86-35, Ano87-28, Ano87-35, Ano88-34, Ano88-39, Ano89y, Ano89-31]. aquatic [AS86]. aqueous [JA88, PP85]. arborescences [Per82]. arborescent [CHT80, Gol88]. archaebacterial [Cla82e]. architecture [Fuj82]. area [BK88, Frö84, KO82, KK83a, Kwo84, Kwo89, Sve82]. areas [Cor83b, DB88, GVP^{+}80]. arenarius [IAHA80]. argument [MJ80, SA80]. Arguments [Col85]. argyrostoma [PS80b]. arise [Syv84]. Arms [Par83a]. aromatic [IO83, MHB83, SH80a, Sýk84, TN80]. arrangement [GPLC84, WS85b]. arrangements [AL81]. arrays [HE80a, NSJW84]. Arrhenius [Kub85, SM81b]. arterial [Bol80, Lef83, Sau80, Zam81, ZB84, Zam86]. arteries [AA84, Bol82a]. artery [MH81, TGP^{+}86]. arthropod [MM83d]. arthropods [Tay85a]. articles [MDW80a, MDW80b, MDW80c]. artieries [Bol82b]. Artificial [AZE83, KH80]. aryl [YHJ89]. aryl-substituted [YHJ89]. ascites [Dug82]. Ascorbate [LH80]. ascorbic [NMM84, Sco82]. Asexual [MD86, SC82, dSTSM83, Cha80e, Mog86]. Asp [UN82]. aspartate [MH80a]. aspects [Aki84, Bar81b, Ber82, Blo85, BvdHB84, FT87, FM88, GG84a, HT89a, HN84, Hoe80, Hum83, Kre82, Mot89, Pet82, PHTCS86, Pro86, Res82, RR83a, RR83b, Sie81, Sil83, SP84, Ten80, WD86, Zam81]. aspiration [SME86]. assay [AWV^{+}88]. assays [AL87, Lat83, Smi82c]. assembled [Yod81, Yod84]. assemblies [Wag89]. Assembly [Kat81, BB81b, BH88, Dun82, KsC85, PHTCS86, TG88a]. assembly/disassembly [PHTCS86]. assess [SM87]. assessing [LM89]. Assessment [PDCD88, BB81a, EL83, Jea87, NP81]. assessments [JGW83]. assignment [PP80]. assists [EGS80]. associated [KD80, Maw85, RMFM85, Top88]. associating [IHTN88, Kur84b, MJN88]. association [Bla84, Ger83b, HR85, KD80, Kur83, McG81, NSJW84, NW85, Sch82a]. Associations [CW84]. Assortment [TS86]. assumed [Wal89a]. assuming [YM89]. assumptions [CB89b]. astogenetic [Gol88]. astray [OK87c]. Asymmetric [EN86, Lap81b, NR85, HP82, MF81, Sel80]. asymmetrical [Bia81, Lam84a, LP88a]. asymmetrically [Lap84]. asymmetries [LE84a]. asymmetry [BT84, DT80, FC81, Faj82, Gra85a, Hoe82, Por81]. Asymptotic [BAB84, Sai82, KH80]. Asymptotics [VP86]. asynchronous [TH80]. AT- [Cla82f, Cla83b]. ATG [Shc89b]. atherosclerosis [Bol80, VTL85]. athlete [Cor82, Kue81]. ation [Dou83]. Atkinson [PH89]. Atkinson-Shorrocks [PH89]. atmospheric [Wal89a]. ATP [ZBL80b, CG81b, FK83, Kas80, Nit86, PC86, Wil84]. ATP-analogs [FK83]. ATPase [FK83, Fri85, Wig82a, Wig82b]. atrioventricular [UKNS87]. attaching [TDA86]. attachment [GK80]. attain [HB84]. attempt [Mal84a]. Attenuation [VG80, FB82b, Man81b, RN85, Von82]. attract [HM87]. attractant [Lap80a, Lap80b]. attraction [CB86b, HJS84]. attrition [HP82, NR85, Ril80]. auctions [HR80a]. August [Ano80y, Ano80-31, Ano81-27, Ano81-31, Ano82v, Ano82-28, Ano83y, Ano83-31, Ano84y, Ano84-31, Ano85z, Ano85-32, Ano86-30, Ano86-37, Ano87-30, Ano87-37, Ano88-33, Ano88-40, Ano89-29, Ano89-35]. Aulacantha [Hun81]. aureus [FB82a]. Autocatalytic [Kau86]. Autogenesis [CK85]. autolysins [Koc88]. Automata [GT88b, TG88a, ABEFS^{+}86, CPPRS87, Ing89, MA83, Wei86]. Automated [KB89b, KM85, KK86a]. autonomous [HC81b]. autoregressive [KTG86]. autorepressor [MRG84]. autosomal [Wer87]. avenae [HL86]. average [Lei81, Mad88, MVV89, MM82c, SPS83]. averaged [Frö88]. avian [Ano80-35, BK88, CG85b, Ham85, Mar80b]. avoidance [Bul84, PSG83, Sto81, WW84]. axes [Yat86]. Axial [SM87, Her83a, LL82, LL84, MB86, Sto89a]. Axial-view [SM87]. Axiomatization [RZ86, ZR86, Jon85b]. Axisymmetric [Wax81, MK88]. Axon [Eil88, dANG83, Sto89a]. axonal [BH88]. axons [AM82, Fra87, Mat84, MUI85]. axoplasmic [AT85, TG84]. aza [ABS82]. azaguanine [SRO85].

B [Hop83, Nus85, WP81, Bec80, BC81a, CM82, EPH86, GWK87, dBH89]. B-cell [WP81]. B-DNA [Hop83, Nus85]. B/r [BC81a, GWK87]. Bacillus [BB81b, BK81, FLNF86]. Back [Gin86]. backbone [CSH^{+}81]. backbones [SW81b]. bacteria [BT85, CDC88, CBY82, Cla83b, CSC^{+}80, GZWR80, GBFH88, Hed85, Hun85b, Koc82, Len84, LH86, Lum84, MRG84, NSCH88, RR83a, RR83b, SK88b, WGRZ80, vdH84]. Bacterial [CS87, Fue80, JPGL86, Bah80, Ble88, Ble89, Bre81, BP87, BPP88, Cla82d, FH89, Gla80, Gro88, HW86, HW87, Hsu87a, Hsu87b, Hsu87d, Hsu87c, Hsu87e, LA82, Len84, LBBL84, Men85, MYS89, vdHN85, OH85, SSF80, SAS85, TD81, YLBB81, YBBL82a, YBBL82b, YBBL82c, vdH84]. Bacteriophage [TG88a, CSC^{+}80]. bacteriophages [Ser87]. bacterium [BK83, RR84]. Balaenoptera [KB88b]. balance [Bor88, BW85, FQ81, FDC^{+}89, MB80, MJ80]. balanced [FS87]. balloon [DGST82]. band [GBFH88]. bands [Mei82, SP88]. barrier [GS80, KB83, LKC^{+}87, Loc82, Ros82]. barriers [Dam86, LL82, LL84]. basal [ATC81, Aok89, Bar88, Gra81b, KK83a]. base [ANOM82, BS83, BHM^{+}85, Cla82d, Cla82c, Cla83b, GPPZ80, Hua80a, Mik81, NP84b, OK82, Row85b, SM81c]. base-pair [Cla83b]. base-ratios [Cla82d]. based [BW86b, Bog87, BSB88a, BSB88b, DW82, DCR82, DT80, Fer83, FBS85, Get81, Hof86, HL89, IYS83, IT80, Ji85, KDAT84, Kin85, Lav84, Law84, LH86, LS87, LPS89, MLB88, Men85, MdMF83, MAGR87, MHB83, Mor85a, MB88a, MS86a, Nev88, SSM81, SC84a, SGDK86, SJ81, Tak89b, Wal89b, Wal89a, Was82b, Wei80]. bases [DD89, KT87a, MT84, MH80c, NC89, Pin83a]. Basic [Cal87, Sch83a, PA86, RB83b, Wes83a]. Basis [Whe84, Ahl81, ACB81, AS86, Bar81e, Bod88, BJE89, Col82, DKW89, Eil88, For85, KP85, Koc86a, LO81, Lil82, MUC81, Pen82, SRO84, WR83, Wil84]. be [Bot82, Boy83, Bro82a, CSL^{+}86, CN85, HRG82, Har82c, JS81, Kor89, Mik84b, Nor82, Pic89, Rob89, SSW81, Smi88, SME86]. beads [SR87a]. bearing [McG84]. beat [Cor83b]. beating [Hao82]. bed [Bas82, HW80b, Lef83]. bee [Ros80a, Ros80b, Ros84b]. begging [Mot89]. begins [Bre86]. Behavior [FT87, Lap81a, And88, AM81b, BL85, BC86b, BSM82, Cha89a, GEL80, HT88, IHK80, KH80, Mat81, MC89, Mot89, New83, OKT^{+}80, OH83, PPC82, Rap84, Ros80a, Ste88, Tro84, VV82, Wol80, YMS84]. Behavior-dependent [FT87]. Behavioral [FHL89, MM83b, SK88a]. behaviors [HC89, Wad82]. Behaviour [BM89a, AK87b, BS83, BSS83, Cra83a, Cre88a, DG85, DG87, DPV83, EL83, EL87, Har87a, HM85a, KN82, KK82a, MH80b, Met82, PCTF83, Pra84, RG87, SR88, Spa80, SF83, Tay81, Whi80c]. behavioural [Leo84]. Being [vRV80]. Bekoff [BB81a]. Belovsky [Rei86a]. bend [MHB89]. Bending [Ale87, SOLG82]. benefits [Abu81]. bent [Cha89b]. benzene [ITK88]. benzo [ITK88]. Bergmann [Sea80]. beta [KK86a]. beta-adrenergic [KK86a]. between [Aki85, Ano81-37, Aok84, BB85, BWM80, BRW89, BR87, BRR87, BGB89, BP81b, BC80, BHM^{+}85, Bri85, BBH84, Cal82, CB86a, Cha80b, Cla82a, CL82, Cra82a, CM87, DD85a, DG85, Dur82, Ebe87, Ebe88, EB87, Fle80, Fra84, Ger83b, GK80, HK88, Her83b, Hig85, HJS84, Hop83, Hum81, IYS83, INMM83, IM85, Jak88, Jam82, JK88, Kar88b, Kat81, KK83b, KHU^{+}83, Koc88, KK81, Lav84, McA87, McC80b, McG80, McG81, MW83b, MATGR87, Mic86, MF81, Mog86, Mot88, Moy82a, NU82, NSW82, Ola82, Omh88b, PK85, PDCD88, PP85, Por82, RN84c, RW82, Rob88, RH85, RSM80, Sab80, SYM86, SCP83, SA86, SNW81, SS88, SA82, Shi87, SHBS89, Ste87, SW83, Sýk84, TS80, Tre81a, VM89, Wag89, WWG^{+}88, Wil86a, WWN82, Yam87, YI88, dBH86]. between [dBH87]. bhopalator [Ji85]. Bi [GVCC82]. biallelic [Gre82b, Gre84a]. bias [BC86b, Orl82, RT83b]. biased [Lap81b, Sil84]. Biases [Pea88a, Pea88b, Pea88c]. bicarbonate [Izu81]. bicolor [IAHA80]. bifurcating [Sto89a]. Bifurcation [Spa80, GMM89]. bifurcations [AS84, CLFL86, MLB^{+}89, PL85b, PG88]. big [BBH84]. big-T [BBH84]. bighorn [LC82]. bilateral [DK82a, DK82b]. bilayer [BP81b, Bru85b, SOLG82]. bilayers [MSS81, TL88]. biligand [Rog84]. bimodality [WR83]. binary [JLS83, MB85, Per82]. Binding [DO83, Inc83a, Kur84b, BA87, Bab88, BWM80, BWW80, BM87, BJ89, BC80, BSB88a, BSB88b, Bri84, Bri85, Bri88, BF81, Di 88, Di 89, EC81, GS81, Gro80, HM80, HL89, Hum81, Inc83b, IHTN88, KA89, KN82, KD80, LE84b, Man81a, MJN88, Mik84b, Mon83, NSKS85, NW85, RC81, RH80, Rei83a, Rei85a, RN84c, RB83b, Sar87, SNW81, Shi81, Sta86, SPG87, Sto83, Sto89b, Whi80b, Whi80c, Whi81b, WWN82]. binds [Gho82]. bineme [Luc82]. binocular [LD83a]. binomial [LPS86]. Bio [Stu85a, AC89, MF89, SB86]. bio- [SB86]. bio-electromechanochemical [AC89]. Bio-informational [Stu85a]. bio-theme [MF89]. bioactivation [RF82]. Bioassembly [Fuj82]. Biochemical [HG87, dVR85, BMR89, Bro81b, CT81a, CB89b, DG85, DG87, DG80, Gar89, Gie88, HS81, Hyv84, LG89a, Mae84, PL85a, PL85b, PG88, Pap82, RMS81, Sco82, Shu84]. biochemistry [BT81, Wel85]. biocides [Aki84]. biocultural [FHL89]. Biodigestion [Hsu88]. bioelectronic [Lac84]. Bioelectrorheological [FP89, PF89]. bioenergetic [MB80]. bioenergetics [MB86]. biogenesis [FT85, RTM88]. Biogenic [CSL^{+}86]. Biol [Ano80j, Phi82]. Biological [ALST81, Bre88, Els83, FSF89, HE83, RT88, Ury81, WdlT84, AK80, Bec85, Blu85, Boo84a, BPH87, CMN81, Cla81, Cla83a, CS84, Coh84, CL82, DDBB83, DR89, DRP^{+}87, DRL88, Eco82, Els81, Fer85a, Fer85b, Fis85, FS80, GEL80, Gla82, GE82, GT88b, Got86, GM82b, Hor89a, Hyv80, Ish81, KH80, Lam84b, Mau87, McC80a, Mea86, Mor85a, MF85, NFP88, Ond80, PCTF83, Pok82, Por81, RR88a, RR88b, Rya80, SSM81, SR80b, ST86a, SCH84b, SM81b, SGJ86, Sou80, Stu85b, Tan80, Tan84, TG88a, Tod85, TLT80, Wei86, WK81, Wil86b, ZG89, vGvGL88, RZ86]. biologists [OK87c]. biology [CF87a, Dra81, Els82, Ewa80, Gar82, GW83, Hol88, Kat81, LH88, Lau82, MW83b, Nus87, PLJ84, Pro86, Ros80a, Tom82]. biomagnetic [Bar87]. biomass [KL81, VdV87, vS85]. Biomechanical [Giv86, Men82]. biomembranes [KB83, MR89, Mul85]. biometric [BGB89]. biomolecular [EG87a, Tra86]. biomolecule [SPST83]. biomolecules [PRH81, PR82, RTM81a, RTM81b, RTM82]. Biophysical [KS81, KK81, Kot82, KK83c, Kot83, VD89, BB85, KP85, Ler83]. biopolymers [Gal83, Kri89, KP81]. bioreactor [SGH85]. biosynthesis [BE81b, Blo87, GT88b, Inf85, SE80, SA80, VdV87]. biosynthetic [RTM88, Tai86]. biosystems [Yag85]. biotransformation [Aki85]. biotransformations [Aki86]. Biphasic [RG87, Bro80b, FK83]. biphenyls [MP87]. Bird [Fle88, LA86, Hou86, Ipi87, Lun87, TS84]. birds [BL88, Fra84, Hou87, Hum83, KLR89, Moy82a, Ray85, TC87, WS84, War85]. birth [Jam88, Jam89b]. bisphosphatase [KSH87]. Bistability [HC87]. bitter [Ven84]. bivalent [NSW82]. bivariate [Jol89]. bleekeri [MUI85]. blockade [Sta86]. blocking [MTM86b]. blood [AA84, Boy88, DK83, Eng89, FW81, JK87, KR85a, KS89, KTG86, KT87b, KIO88, KM87, Lan84a, LR85b, Loc82, MP87, MHLK87, PTD81, SOLG82, TLS83, Wax81, Win81]. blood-brain [Loc82]. blue [KB88b]. bluffing [GM89]. Board [Ano80d, Ano80e, Ano80f, Ano80g, Ano80h, Ano81c, Ano81d, Ano81e, Ano81f, Ano81g, Ano81h, Ano82b, Ano82c, Ano82d, Ano82e, Ano82f, Ano82g, Ano83a, Ano83b, Ano83c, Ano83d, Ano83e, Ano83f, Ano84b, Ano84c, Ano84d, Ano84e, Ano84f, Ano84g, Ano85a, Ano85b, Ano85c, Ano85d, Ano85e, Ano85f, Ano86b, Ano86c, Ano86d, Ano86e, Ano86f, Ano86g, Ano87b, Ano87c, Ano87d, Ano87e, Ano87f, Ano87g, Ano88b, Ano88c, Ano88d, Ano88e, Ano88f, Ano88g, Ano89a, Ano89b, Ano89c, Ano89d, Ano89e, Ano89f]. bodies [Her83b]. Body [Cal83b, Don86, GM87, LIWK86, Rei86b, Aok89, Bel87, BIS^{+}80, BW85, Con83, KK83a, Koo86, KS82, Pec88, Rei86c, Sea80, SR87b, Tat88]. body-weight [Pec88]. Boltzmann [LP83]. bonds [Coh84]. Bone [BB88a, Bon82, OWN89, RL84, SYM86, STN^{+}84, Whi89b]. bones [Ale84]. bony [Mul87]. Boolean [LD83b]. botanometry [Jea80]. both [Ash85, Bel80]. bottom [Mul87]. bound [BO82, DH85, DBB80, EB87, GK80, Inc83b, MN89b, Ten80]. boundaries [TLN85]. boundary [Fer81b, HE80b, HSd^{+}88, JA88]. bounded [Eld80]. Bounding [Ray85, WS84]. bounds [Law83]. bovine [NSKS85, UN82]. bowel [LH80]. Br [WGRZ80]. Bradbury [AK86]. bradycardia [Cor82, Kue81]. Brain [SC86b, SC87b, Ano81-37, AA85, Bar88, Coh85, Hof83, Hon84, Hon86, Lib85, Loc82, Mas86, Nin83, Tou85]. brains [Cal83c, Tod82]. branch [Par85, Pug87b]. branched [Gie88, NSJW84, Wag87]. branches [GMM89, WH83]. Branching [CHT80, Gol88, Gun89, AY85, Hor80, Par85, WH86, Zam81, ZB84, Zam86]. break [Kub85]. Breakage [CCG87]. breakdown [Bre82]. breaking [DfP87, FEC87b, FEC87a, LO82]. breaks [HT89b]. breast [KF84]. breathing [AFHM86, RSSK87, SBC80, Sau80]. Breed [BB81a]. breeders [CB86a, SCS88]. breeding [KK82b, SC86a]. bridges [MS86a, PC86, Toz89]. bristled [CK87]. broadcast [CJ82]. Brood [TC87, Omh87b]. Brooks [HT84]. Brooks-Dyar [HT84]. Brownian [Eld80]. brushborder [WMS81]. bryozoan [GMM89]. bryozoans [Gol88]. bubble [SG82]. budgets [Koo86]. Buffer [KS80a, Izu81]. buffered [NK88]. building [BD80]. bulk [AHFH81].

Bulmer [Cha82b]. bundles [BD80, ST80]. Burk [FLR83, GC83]. Burkitt [Mor85b]. bursal [SRPS89]. bursting [DG87]. butterflies [Ano80j]. BVP [RL88]. bypass [MH81].

c [Coo88, CH82, GDHB87, MATGR87, MAGR87, Mer87, BSB88a, BSB88b, Law84, Mik81, PRH81, RH80, Rei85a, SM81c]. c-fos [Mer87]. c-myc [Coo88, GDHB87]. C1 [LBBL84]. Ca [AH86, BRW89, BLR83, BH83, FK83, Fri85, KK83c, KT81, Szu83, TY82, Wig82a]. Ca-ATPase [Wig82a]. Cable [Kaw84, Ash85, Bla80, Poz87, Sch89]. cables [Bro81a]. cactus [Her83b]. cage [RS86]. calcification [HB89]. calcifies [Bol82b]. Calcium [DLN85, TG83, AWMK85, BA87, Bab88, BSB88a, BSB88b, GT85a, HH85a, HTG89, HF85, LPS86, NSCH88, PS80a, PS81a, PS84, RH80, Rei85a, SJ81, Ten80]. calcium-binding [BSB88a, BSB88b]. calcium-regulated [GT85a, HTG89]. calcium/magnesium [RH80]. Calculated [Rei86b, RCA89, TC80b]. Calculating [SBM89, Gie88, KB81]. Calculation [HR86, Kar88a, Kwo84, RTM81a, RTM81b, RTM82, SH83b, Spe87, Vil83, WMR86, WR86, DML80, HKC83, HCKB83, Kot83, Mon83, MV88a, RC81, SK84, War86, YBBL82c, GVP^{+}80]. calculations [ABS82, BS85, GPLC84, Kot82, Kri89, LBBL84, YLBB81, YBBL82a, YBBL82b, ZG89]. calculus [LRS82, RLS82]. calibration [PW88]. calling [NC89]. calls [Tam87]. Calmodulin [DLN85, Rei83a, Rei85a]. cAMP [BS88b]. Can [Bot82, HB84, Hin87, JS81, Kle81, Nor82, Rob89, Smi88, Bar83a, Boy83, Fel82, FDC85, Gar82, Ibe83, Ker89, Koo86, Kor89, Mor85a, Orl81a, SSW81, You81]. canal [Ber82]. canalization [Ker89]. Cancer [CL84, Gra83b, Gra84, Wes83a, BW86b, Bre82, Con83, DCF80, HG80, Har82c, KF84, KKJ81, LH80, Mel82, SW84, WD86, dGV80]. cancerigenesis [Mel82]. cannibalism [Ste85]. canopy [JT84]. capacitance [HF88]. capacity [CF87b, LE84b, Tak89c]. capillary [And81, BM85, CD84, HW80b]. capita [Get84]. capitula [Rld82]. Capping [Rao82]. capture [BBH86, SG86]. carbaminoprotein [Fox81]. Carbinolamines [SBW83]. carbohydrate [Wri84]. carbon [Fox81, Law84, NL86]. carboxylic [KK83c]. carboxypeptidase [NU82]. Carcinogenesis [Har82b, Alf83, Gar87, Hje86, Joh84, OO84a, Uma82]. carcinogenic [ITK88, JGW83, TN80]. carcinogens [BG87, RF82, SL80b, Sýk84]. carcinoma [SS88]. cardiac [IYS83, JCK80, Li86, Win89a, Won81]. cardiovascular [Zam86]. care [WGS80]. Carlo [EG87a, EG87b, Hsu87a]. carnitine [MRR80]. carnosine [Bro81b]. carrier [Bar82a, Gai80, YMIH86]. carrier-mediated [YMIH86]. carryover [GR88]. Cartesian [Pap89]. cartilage [HB89]. cascade [Cou80, TS88a]. CASE [KK86a, KB89b, AK87b, BF83, CB84, FS83, Hin87, HK83, Mat87, PS83, SL81, VCC^{+}88, IHK80]. cases [PR82]. caste [Gri82a, Gri82f, Lum82]. cat [RS86, RSSK87]. catabolism [MG81]. cataglyphis [IAHA80]. catalase [Dou83]. catalatic [Dou83]. catalysed [Cho81]. catalyses [PA86]. catalysis [AK87a, AK87b, GG82, Hav89, LM85, Ola82, Org86, Res82, SGH85, Shi82, Vol81]. Catalytic [Kri80, RN86, ANS83, Gho82, HAG82, KLK^{+}88, LSFT85, SLYF84, SLFT86]. catalyzed [Dug86, For80, Pen80]. Catastrophe [KRB81, VD89, ST88]. catastrophic [Bre83]. Categorical [Iid88]. cathodoluminescence [Bas80c]. cation [Sve82]. cationic [Cha89a]. cations [KK83c, ST88]. caudal [SJK88]. causal [Bar86, VTL85]. cause [LS89, MH81, Mor85b, Was86a]. caused [Lef86]. causing [BPP88]. cavities [BHM^{+}85]. CCC [Voi85]. Celestial [PW88]. Cell [Bir81, Boy83, BC81b, BC83, Coo87, CM86, EP86, Gal84a, Gal84b, HCVM88, MB82, Mar84, Sza83, Voi85, ATC81, AL81, Ada80, AMM80, ABB89, AWV^{+}88, AC89, BB86, BGB89, Bec81, BG84, Ber83, BO82, BS84a, Bod88, BKB80a, BKB80b, BC81a, Bre86, BWM82, BK83, CZJY80, CCG^{+}82, CP85, CT81a, CPPR87, CPPRS87, CJ83, COR83a, De87, Der87, DG80, DV83, DDJ83, Eng89, FP89, Fri83, Fuj82, GPKBG89a, GPKBG89b, GPAL88, GMM89, GF81b, Goy82, GAD80, HB86a, Hay80a, Hay80b, HR86, Hir83, Hir84a, HK83, HE80b, HO84, HYDS84, Inc83b, IP83, JJ82a, Ji85, Joh84, JP89b, JP89c, KG82, KDAT84, KB88a, Kle81, Koc89, KWG87, KM87, Lan84a, Lat83, LD83b, LBBL84, MK88, MCB^{+}86, Men82, MBJK89, MN89b, MM83d, Mor84, Nes84, Nev88, Nor89]. cell [NL86, OWN89, Ond80, Osu84, PWH^{+}88, PH87a, PO86, PF89, Pet83, Por83, PL87, PTD81, Pro80, Pys80, Rab82, Rei81, Ric87, Rot82b, Rot83, San84, Sch83b, SS88, SCH84b, SGd84, SG82, SCP84, Sve81, SOLG82, Sve82, SH82b, TL86, Ten80, Toz89, Tys83a, Tys83b, TH85, TD86, Tys87, Uma82, Ury81, VA88b, WD86, WKV88, WS82, Whi89a, WP81, Wig81, Wil82, Wol85, Wri84, iWSS88, Zaj86, dBH86, dBH89, dMP85, vZvdSdL81, vHU87]. cell-bound [Inc83b]. cell-cell [CPPR87, CPPRS87, MM83d]. cell-density [Ada80]. cell-killing [WD86]. cell-mediated [GPKBG89a]. cell-specific [Ond80]. cell-types [BGB89]. cells [BF83, Coo84b, Coo87, CM88, Dou84, Dug82, EP86, Fel80, Fra82, GWK87, HRG82, Hol83b, Hun85a, Inc83a, KBR88, KDAT84, Kle81, KIO88, Kor82, Kre84, Ler83, LP86, MdMF83, MG84a, MHLK87, Rap84, Rat86, RN84a, RN84b, RSM80, Sai82, SR87a, SPP84, Sch85, SH83b, Ste88, Suc85a, Suc85b, Tai86, Tak87, Tho85, TCY85, VDH82, Wes83a, WGHY82, WB83, Wol80, Wri80, YKS80, YVIG80]. Cellular [Bun80, Mah84a, Mah84b, Poh81, Suc83, Alf83, Arm84, BS88b, Bas80b, Cha89a, CPPRS87, CCDF81, FM86, GS86, GM84, HK81, Ing89, JLS80, JS82, Kur86, MF80, MA83, MB88b, NDB83, Rap84, RG87, San84, Sco84b, Szu83, Wes83a, Wil84]. cellulose [RQ81]. centered [MSA^{+}81]. centers [MH80c]. Central [GN86, Wee84a]. Central-place [GN86]. centred [PRH81, PR82]. centres [DB88]. centrifugal [Mor84]. centrifuge [IPS83]. centripetal [Sib83]. centromere [RB82a]. centrosomal [Gra88]. ceramide [WMR86]. cereal [LP88a]. cerebellar [Mon88]. cerebellum [AA85, Ito82, KMSR84]. cerebral [Der84, Ipi87, TO80, Zha84]. cerebrospinal [Spe81]. cerevisiae [BW87, Fut86, WB88]. certain [LN85, Man81a, Rot87]. cervix [Asp88, GTH88]. cesium [CS84]. chain [CC83a, Cur88, DW81a, Dou81, Dou85, FJ84, Gli80, HdL84, MVV89, NBZ^{+}83, WJT87]. chain-like [MVV89]. chains [BP81a, DC87a, Jea85, Kuh82, NP84a, SH87, SBPF83]. challenge [Mik84b, MG85]. chamber [MR86, RM83]. chambers [FBB89]. chance [GW83, ZG82]. change [Agm85, Cha86, CH85, SK88a, VA88a]. changes [Asp88, BW89, BW86b, BC83, Cra80b, Der87, DDJ83, HG80, Hav89, KD80, MB80, Mat83, Mul85, Rob87, Sch89, Sir86, SJ81, Tor80]. changing [Kon84a, RT85, Tay89a]. channel [Ash86, Bab88, BE82, Cha88c, CS84, Col84, KK87a, KK87b, KSW^{+}85, Leu87b, Lev80b, LL87b, MTM86b, Mon83, NK88, Pic86, Sta86, Ste87, SJ89, TIK89a, Tuc87]. channelling [KVW88]. channels [BA87, BW86b, Cha80c, Cla83a, Fri83, Gre89, KFPB85, KK87a, Leu87a, Lin80, NA84, Ste87, Ste88]. chaos [RL88, Tom82]. chaotic [Spa80, Wag82]. Characea [TCY85]. characean [TCY85]. character [Oli85, Whi81b]. Characteristic [OK82, Fra84]. characteristics [Bar81c, Coo81, HT87, KHB83, PA82, Zac81]. Characterization [GB87, GH87, BE81a, LP83, RSSK87, Sta86, Tan84]. characterizing [RR88a]. characters [TO89]. Charge [BE82, DV83, DT80, Kri85a, PA86, RR88b]. charged [Kol82, MOK87, OK87a, Rab82, RN84a, ST88, VDH82, Wat85]. charges [DT80, EM81, ZG89]. charging [Fer85b]. cheat [Arm84, Vic87]. chelation [BRW89]. Chemical [Hua80a, MH80c, OC81, ST86a, Alf83, AS86, BSR^{+}88, BG86, Bol82a, Bol82b, Boy83, Coh84, Dur82, FW81, Gal84b, GPLC84, HM80, Jaf80, Jaf84, Kin82, MB84a, Mas84, McC80a, Min82, MP85b, NBZ^{+}83, Pap82, PO81, RW82, Rob84, Rob87, RF82, SB86, TS88a, Wag87, Wil86a, YLBB81, YBBL82a, YBBL82b, YBBL82c]. chemical-physical [Bol82a]. chemically [Bar83a, FS85, KBR88, KKNM80, KK80]. Chemicomechanical [ML84]. Chemiosmotic [Nit86, Cha82a, Mal81b, Wag87]. chemistry [HT83]. chemokinesis [VV82]. chemoklinotactic [BC86b]. chemoreception [KKKK80]. Chemosensory [Lap81b]. chemostat [LH86, Pow85]. chemostatic [FBP87]. chemotactic [Ros85]. chemotaxis [Lap80a]. chiral [DfP87]. chirality [Blu85, Fer82]. chirility [RB82e]. Chitin [BS81]. chitinase [BS81]. Chitty [Ste81]. chlorfenvinphos [Aki85]. chloro [Aki85]. chlorophyll [Law84]. chlorophyllous [FSF89]. Chlorophyta [PL85c]. Chloroplast [RTM88]. Choice [Koc82, Cha88b, GR88, LSH88, MM85, Pom87, RC89, Tay88b, WD81]. choline [IHK80]. chondrogenesis [Oth86]. choosing [BMR89]. Choosy [Man85]. Chou [CC83a]. Chou-Fasman [CC83a]. chromatic [YM89]. chromatid [JJ82a, KM81, WS85a]. chromatin [BN81, BF81, JS80, JS81, SR80a, Sch82a, Str83, Suc85b, Tak89b]. chromatin-dependent [Tak89b]. chromatin-DNA [BN81]. chromatographic [BF83, MKM80, NM81]. chromosomal [NMB85, RSR88, Tak89a, ZgQy84]. Chromosome [CG80, Mat80, Bia81, Bod88, BC81b, CG84, Cra80a, DS87, FBS85, Hun81, Luc82, Mar80a, NSCH88, SP88, Sto84b, TD81, WEHM82, Wil81a, Wri80]. chromosomes [Bas80c, Blu83, JL88, SCP83, Tak89a]. Chronic [RM86, dGV80]. chronological [Lib88]. chymotrypsin [LSFT85, SLFT86]. cilia [MHB89]. Ciliary [Rik82, FB86, SM87]. ciliated [BLA82]. Circadian [CB86c, ATC81, CG85a, KS80b, KFSW82, LS87, Nun81, Pai82, PS80b, Pet80, Pro84, SL87a, SL87b, Sin83, WF82, SL87b]. circle [Fut86, MK83]. circuit [Mad88]. circular [EAL80, Kuh82, MHII83]. circulation [LTM85]. Citric [För88]. citrus [LH80]. clamp [Cha88c]. class [Aug85, BS82, Ham82]. classes [Ebe87, Ebe88]. Classical [Gla82, BPM82, Lim88]. Classification [TL86, ABEFS^{+}86, CB83, GFA89, HAG82, KS88, Pen82, SFR86, SP84, Tay86]. Clathrin [LFP84, McK83]. cleaner [Nev88]. clearance [BAB84, Nor82]. clearly [Rot80]. cleavage [GT80, GL84, Kar88a, Sto89b]. cleft [WGP^{+}88]. clefts [CC83b]. clinical [LBW86, She83]. clinically [SBW83]. clock [Bec85, Bri80, LS87, Men82, Nun81, SL87a, SL87b]. clonal [MKM80, CPPRS87]. clones [AL81]. cloning [GP82a]. close [MG84b, Tys83a]. closed [Kol82]. clotting [Vro83]. clue [Suk89]. clues [DD89]. Cluster [FE85, MKM80, OL86, Omh87a, Omh87b]. clustering [HL86, Nus80a]. clusters [BG84, GH83, GMM89, Sar87]. CNDO [KK83c, YBBL82b]. CNDO/2 [KK83c]. CNS [Pel84]. Co [CBC87, KKNM80, MN89a, Nam84, NC89, RH80, SH83a, Tor80, BWW80, Bas83, BC80, CN80, Gho81, IHTN88, Kar80, KD80, Mor82, NU82, Nee89a, Pap89, PPS86, RN85, RN86, Shc88, Shc89a, Shc89c, Sib81, TO89, TSRK88, Whi80b, Whi80c, Whi81b, Zol83, dBH89, vdH84, HMF88, Sau80, Spe87, SR89, URC83]. co-dominant [Mor82]. co-evolution [Nee89a, TO89]. Co-evolutionarily [MN89a]. Co-existence [Nam84, vdH84]. co-operation [dBH89]. Co-operative [KKNM80, NC89, SH83a, BC80, Shc88, Shc89a, Shc89c, Whi80c]. Co-operativity [CBC87, RH80, BWW80, Bas83, CN80, Gho81, IHTN88, Kar80, KD80, PPS86, RN85, RN86, Whi80b, Whi81b, Zol83]. Co-ordinate [Tor80, Sib81]. co-ordinated [NU82, TSRK88]. co-ordinates [Pap89]. CO2 [Ano81-36]. Coagulation [HB86a, Eld80, KS89]. coalitions [Hug86]. coarse [BL85]. coat [Bar81b, Ler83, Mur81]. coated [Kat83, McK83, SEN82]. coats [DV83]. coccid [ST80]. code [AOB85, Cha81, Cla82b, Duf83, Duf85, Duf86, Gra83a, LJ88, Nus83, Rei84, RB82c, RW83, Shc88, Shc89a, Shc89b, Shc89c, Suk89]. coding [AS87, Alm85, Eld84, FFM82, LM85, Mic86, Row85b, SSW81]. Codon [AS87, Bro83, Bul88, Bro80a, Cla82a, Fil88, LJ88, LvH87, Mik81, RT83b, Sir86, TRS84]. codon-anticodon [Mik81]. Codon-level [Bro83]. codons [VL86]. coefficient [And81, Don86, WR83]. coefficients [BWW80, Gie88, HT87, KTNO81, Nor82]. Coevolution [Len84, RR84, Lew81a, Lew81b, Lum88, RR83a, RR83b, Wer87]. coevolutionary [Bar84]. Coexistence [LH86, Arm88b, Arm89, Hol85, Ive88, SG89b]. cognitive [JPL83]. coherence [MHLK87, PCTF83]. coherent [AM82, NS86, Pok82]. coil [NMMO82, VM82]. coiled [NMMO82]. coiled-coil [NMMO82]. coin [CK82]. coin-flipping [CK82]. cold [Sto81]. ColE1 [BLC86, KP89a, YI88]. ColE1-type [BLC86]. coli [BJL85, BC81a, BC81b, BC83, Bre86, CM82, Cla82a, DS84, GWK87, Jak88, MZ89, Man81a, Man81b, NHT88, Nag81, Nor89, Ros85, WP84, WGRZ80, YI88]. Collagen [CH82, MLB88]. collagens [Fin81]. collective [CF85, CF87a]. colloid [Mor85a, Rap84]. colonially [BL88]. colonies [Boo84b, BPP88, Ede82, MVV89, Omh88a]. colonization [SRPS89]. colony [HCW89]. color [CPPR87, CPPRS87]. combat [LH83]. combination [Ber85]. combinatorial [HKC83, HCKB83, NP84b]. combine [Emi82]. combined [Lam88]. combining [CSH^{+}81, Rog83, SW81b]. combustion [New80, New83]. come [FBB89]. comets [HHP89]. comma [Cla82b]. comma-free [Cla82b]. Comment [GT88a, Met82, GG84a, KP85, McC80b]. Comments [AK86, Bha86, HLS89, Hou83, JLS83, MB88a, Sco89, Tho82, dCJM86]. commitment [Pro80, Sve81]. common [Esh84a, MDBRM87, TDA86]. communal [CB86a]. Communication [RH85, Ibe83, Jaf80, LH83, LL87b, SC86b, Tan84]. communities [AW83, BHMR85, RKL83, Ula80a, Ula80b, Yod81, Yod84]. Community [HE84, AK84b, CM87, GAK88, IH80, MSA^{+}81, MN89a, Mau87, Phi81b, Tay88b, Tay88c]. comparability [Bor82]. Comparative [Li86, Hol89, Kol82, Rid89, SPP84]. comparing [Pea88d, SP88]. Comparision [GWK87]. Comparison [ANS83, BC80, Jøn85c, RMFM85, SK88b, ITK88, LBBL84, MAGR87, Mur82, Blo85, MTM86a]. comparisons [Boo84a, DE82]. compartment [FGDS87, MRB83]. Compartmental [HM85b, FCBK88, SA82, She83]. compartmentalization [Dea80]. compartmentalized [Bun80]. compartments [EGS80, GF81b]. compass [Bri80, PW88]. compensation [Mon88, Nun80, SC87a, YE82]. compensatory [PGKE88]. compete [PC86]. competing [NM80]. Competition [Arm89, CM87, Ebe88, Gat82, GAK88, IM85, Asm83, BP86, BGB89, BB81c, Bul80, CW84, Ebe87, Kor89, LD83a, LS88, LL87b, MG84b, MG85, Nun80, Par82, PH89, Pow88, RKL83, Rog86, Ryt87, Shi87, Sil84, TB80, Van84, Ven82, You81]. Competitive [Nam84, Bri89, Emi82, Hal80a, Hol85, ICL85, Jol88, Kar88b, KS80a, Kol82, Man81a, MMGL87]. competitors [Ive88, Kor89, LH86]. Compilation [Nus86, BT87, Ola86]. complement [DW83]. complementarities [Was82a]. complementarity [BHM^{+}85, FT85, MDPF86]. Complementary [BBH84, Sho85, Bre88, GP82a, ZR84]. complementation [Kat81]. Complete [TD81, Sir86]. completion [Bel80]. Complex [Cre88a, DG85, MLB^{+}89, NMM87, SR88, Bec88, CN85, DG87, FM88, GBS89, HE84, Jol89, KS80a, LH89, Lev80a, NMMO82, Rak86, RTM88, RLS82, SF89, Tay88c, Tre81b, WBKW82]. complexes [Inc83b, Kur86, Mik81, vGvGL88]. complexity [AW83, CJ82, DS87, Gal83, Gar82, HE83, KC85, KC87, Mau87, Nev88, Pap80, Pap82, SH81, Yag85]. complication [SC87a]. component [Bol80, Hof82, MSS81]. components [Joh84, NA84, Rab82]. composed [MKL81]. composite [Asp88]. composition [Bro80a, CBY82, Cla82d, Hol83b, MLB88, NSW82, Row85b]. composition-dependence [NSW82]. compositions [CH82, GNG80, HT87]. compound [HFO83]. compounds [Min82]. comprehensive [JP89b, McA87, PL87]. compressed [BM89a]. compression [Whi82]. Computation [DG80, GJ88, AWV^{+}88, BK88, FW81, GRS82]. Computational [DS87, HCKB83, WC88]. Computationally [Day83a]. computed [GRS82, TLS83]. Computer [CMM86, DDA85, FEC87a, KB89b, KD80, MF80, MRB83, PLJ84, Pok82, Rld82, Ste88, VV82, WB89, WGKG83, AK82, Bor84, BF88, BF81, Fia88, HO84, HYDS84, Hum81, KM85, KK86a, LF86, Ola86, PO81, PSL^{+}85, Sal84, SC84a]. computer-simulated [Bor84]. Computing [Whi89a, CL82, JLS83]. concealed [Gun89]. Concentration [MHII83, CG81b, DFS89, ER84, Fel80, Fer83, FDS^{+}88, GS80, HG80, KSH87, KL80, Mal88, RT85, RdL83, SK88b, SR87b, SJ81, WD81]. concentration-dependent [Mal88]. concentrations [AL87, BT85, BF88, Lan82a, Law83, SAS85, VCC^{+}88]. concept [Coo84c, HAG82, JGW83, Kno82, LL87a, Mas86, Mel82, Nes84, RT85, SJ85, SOLG82, Sza83, dL83, MB86]. conception [Jam85, SCB81]. conceptive [SCB81]. concepts [Coh84, Els83, Ji85, LH88, MSA^{+}81, Mor85a]. conceptual [BG87, Wic85b]. Concerning [All85b, BWW80, Bou84, Hok86a, Wag82]. conclusions [Har82b, DV83]. Concordance [JS82]. condensation [JS81, Mat80]. condenstion [JS80]. condenstion-decondensation [JS80]. condition [Fer81b, Hsu88]. conditional [FS88]. conditioning [Top88]. Conditions [LA86, Whi81b, Aok89, BM88, DY82, NM81, Omh87b, Pop82, RM86, SM88a, Whi80b, Yok85, dBH89, dMP85]. conductance [Bru85b, Col84, Str85a, Str85b, Str85c, YVIG80]. conducting [SS85]. conduction [KYU^{+}86, RN84b, UKNS87]. cone [Sim80a, Sim80b, Sim81a, YM89]. confer [Gri83]. Conference [Ano80c]. Configuration [ZG82, ZG89]. Configurational [And81]. Configurations [WP81, HYDS84]. Conflict [MG84b, Bul81, CT81b, Ell86, Esh84b, Har80b, Lat83, Que83, Rei87, Rob80, TC80a]. conflicting [TJP83, CPPR87]. Conflicts [Hai86, BC86a, Esh84a, LE84a, Sel80, Zee81, vRV80]. Conformation [Kot82, Kot83, MN89b, Aid88, Flo84, FB82a, HKC83, HCKB83, LBBL84, RTM81a, RTM81b, RTM82, WMR86, WR86, RTM81b, RTM82]. Conformational [BS85, Bur82a, Kri85a, Agm85, BS83, BW89, Hav89, JC80, MSS81, MAL87, Nus85, Ola82, PA82, Rak86, YLBB81, YBBL82a, YBBL82b, YBBL82c, LBBL84]. Conformations [GPM89, MRR80, GVP^{+}80, ZBL80a, ZBL80b]. conformons [Ji85]. confrontation [NDB83]. Congenital [BGB89]. congruence [GFA89]. conjugate [GPKBG89a]. conjugated [Toz89]. conjugates [GPKBG89b]. conjugation [TD81]. Conjugational [TD81]. conjugative [vdH84]. Conjugatophyceae [PL85c]. connectance [RKL83]. connected [AK80]. connection [KO82, KK83a]. connections [Wil86a, YM89]. conscious [Hon84, Sny88]. consequence [Der84, GKB86, Har80a]. Consequences [CM82, Coh84, Cla82d, Cla83b, JCK80, MJ88, McG81, Pin83a, SG88, dSTSM83, WWN82, KD80]. conservation [LR85a, Tay86]. Conserved [Nus88, Syv84]. Consideration [MJN88, MZM89, NA84, TCR85, DKW89, Ond80, Van82]. Considerations [GRJ81, RJ80, Bee80, BW89, BIS^{+}80, BE81b, Blu84, Dix83, FQ81, Gra80, JLS80, LFP84, Miy86, Pap80, Tho83, VFS80, WS84, WWN82, Woj83]. considered [KO82, KK83a]. considering [WG81]. Consistency [Cha80b]. Consistent [Tay88b]. consistently [Coh84]. consisting [Gri82d, Gri82e, Gri82a, Gri82f]. conspecifics [Tre81a]. Constancy [Hin86, KF87]. constant [Bar83a, BS82, Hir82, HM80, Kri89, Lef86, RdL83]. constants [BWM80, BSB88a, BSB88b, Cal82, DY82, Dur82, GC83, HL89, KC83a, KD80, LR80, WIN80]. constituents [KDAT84, SE80]. constrained [ES88, HM85a]. Constraint [Gar89, LSH88, SL81]. Constraints [SS89a, Wri84, BJE89, BGG84, Che84b, Fer88, Giv86, Hol83a, JL88, NP84b, Pra84, Ras87, RL82]. constructed [LdJ86]. constructing [GP82a]. Construction [McG84, FFM82, Tay88c]. Consumer [Sch81]. consumption [LG85]. contact [Ber83, Rld82, Rob84, Rob87, SA80]. contacting [RR84]. contagion [Ros83]. contain [Ber89]. containing [BK81, CH80, GMTG80, SR87a, Sir86, ZG89]. content [Gar82, Kon84b, Mat83, RGR82, RT83a, RH85, SV84a, Sve82]. contest [MG85, Sch88]. contests [EN86, GM89, MH89, Moy82b]. context [Kha86, Shc89a]. contexts [FT87]. continuity [Ere85]. Continuous [Agm85, KR82, Ble89, CT81a, CSC^{+}80, Esh83, Fai89, FS86, HN85, HE80a, PM80, SA80, VA85]. Continuum [Coo82, War86, Coo88, Cou80, GW83, Lan84a, Lan85, MB84b, Pla82, Wax81, SKF83]. contract [HE80b]. contractile [Rik82]. contractility [Reg88]. contraction [BM89b, Bar82b, BvdHB84, CG81a, CG81b, GG84a, HE80a, Kas80, Lan85, MB88b, Obe81, See86, SS87, UK82, Won81]. contrary [HK81]. Contrasting [LL87c]. contribution [Cze87, Ibe83, MS86b]. Control [ASK89, Gie88, Hut89, Kho84, OH85, Pro80, AFHM86, BJ84, Bar88, Bar80b, Ble88, Ble89, BLC86, BS89, Bur82b, Can87, Car81b, CH86, CS87, CMN81, CLF84, CLFL86, COR83a, Col84, CB89b, DCR82, Hal80b, Han86, HH85a, JJ82b, Jam86b, Jon85a, KAM85, Kuh82, LS87, Mah84a, Mah84b, MRG84, MPPF80, Nag88, PL84a, PJL84, PPL85, PL85a, PL85b, PG88, Phi81a, Phi82, RMS81, Red88, Rei87, Sau80, Sek85, SF89, Stu86, Tor80, Tys83b, TH85, TD86, Tys87, Von82, WKV88, WS82]. Controlled [Omh88a, PH89, Bar82a, FEC87a, SBC80]. controlling [KFSW82, YI88]. controls [Bre82, Coo87, EP86, Nor89]. controversy [Ros84b]. convection [BRR87, TGP^{+}86]. convection-dispersion [BRR87]. convective [EAL80]. conversion [HN85, OH84, Peu83, Peu85, Peu86, WOT85]. converted [Cal87]. convex [Wei85]. convey [Smi82a]. cooperation [FT87, Hal80a]. Cooperative [KK80, FEC87b, Lan82a, MS86a, TCR85]. Cooperativity [Bri84, Bri85, GG84b]. cooption [Car85]. coordinate [Lew81c, Pap84, Sib83, War85]. coordinates [Hun85a, Lew81c, Pel84]. Coordination [Pel84, LCDC84]. copper [Bro81b, Sto89b]. coprophagous [Har85]. copulation [Yam87]. Copy [Fut86]. Coral [BHMR85, AK84b, AK86]. cord [FW81]. core [KFSW82, Wes83b]. cornea [Kwo84]. corneal [Hor89b]. coronary [LTM85, MH81]. corpuscle [Eng89]. Corrected [RC81, Kap85]. correcting [RW83]. correction [Blo85, HT80, Tay81]. corrects [Kat85]. correlate [WK81]. correlated [BB88b, LN85]. correlates [Hof85, Man85, MKM80]. Correlation [SW83, Aki85, BB85, Frö84, HCVM88, Jak88, KTNO81, LF86, Yam87]. Correlations [MM83b, Cal87, FBF85, RW82]. correspondences [DD85a]. cortex [GT85b, Ipi87, TO80, TG88b, Zha84]. Cortical [CVNO87, GT89, Tod82]. corticalization [Hof85]. Corticosteroid [Web83]. cosexual [CB86b]. Cost [ZB84, Zam86, Abu85b, Bul82, Cha80e, Cha80d, Cha82b, HvL86, Ols87, Par83b, Sav81, SPKJ89, WH86]. costing [SC84b]. costs [Alt89, Cas82, Cas84, Par83a, Pom87, RS83]. Could [CSL^{+}86, Mik84b, Syv84]. Coulombic [Mal84b, Mal84a, Mal85]. counter [EG87b, Lay87]. counter-current [Lay87]. counter-ions [EG87b]. counterdefenses [Len84]. Counterexamples [Eva80a]. couple [UN82]. Coupled [Hav89, BESD84, Cum89, FM88, Ger85, KS80b, LG89a, Lin80, LK83, Tai86, dSTSM83]. couples [SOLG82]. Coupling [AG89, HK88, MGL81, Blu84, Lac84, LM80, Nit86, Ola82, Omo89, RN85, Won81, Woo80]. course [MMGM80, Sen88]. courses [FK83]. courtship [Mot82]. covalent [Fri85]. Covariance [Ive88, PH89, Seg81, Sto84b]. covary [SR89]. cranio [SJK88]. cranio-caudal [SJK88]. crassulacean [CF82]. Crawling [KF83]. creaons [BH81c]. creatine [DDA85]. creating [Ven82]. creatinine [Bor82]. creep [FP89]. crenated [Lan84a]. Crenation [BKB80a, BKB80b, ABB89, BKB80a]. crest [Bar89]. crick [Ji88]. crime [HBH83]. criteria [SGJ86, BMR89]. criterion [Bar80b, IT80, Smi84]. Critical [Mur87, BG87, FT85, GTH88, KB86, KSH87, NP81, PTD81, Sib81, Zac81]. critique [AA85, Hof87, Wic85a]. Cross [Fel83, NSJT81, Syv85, Hua80a, Hua80b, MS86a, PC86, SW87, Syv84, Toz89, VFS80]. cross-bridges [MS86a, PC86, Toz89]. cross-link [Hua80b]. Cross-linkage [Fel83]. Cross-linking [NSJT81, Hua80a]. cross-reactions [SW87]. Cross-species [Syv85, Syv84]. cross-walls [VFS80]. crossbridges [BM89a]. crossed [TS80]. Crossover [Mag88]. crown [BHMR85, ICL85]. crown-of-thorns [BHMR85]. crowns [McM81]. cruciform [MM82b]. cruciforms [McG89b]. cryoprotectants [TL86]. crypt [Pys80]. crypts [PL87]. crystal [EG87b, Goo84]. crystallin [CH82]. crystallized [Bol82a]. crystals [CH80, Sim86]. cubic [WBKW82]. cuckoos [Kel87]. cues [Was86b]. cultivated [Bre83]. cultural [BR83, FHL89, YE82]. culture [Ble88, Ble89, CSC^{+}80, Gro88, Hir83, Hir84a, Koc86a, Lum88, Pow85, Ree80]. cultured [SPP84]. cultures [EH83, GF81b]. cupping [BKB80a, BKB80b, BKB80b]. cupula [Ber82]. current [BBS88, BB84, Cha80c, KK87a, KK87b, KSW^{+}85, Lay87, Rat87, TMGB80, VG80, VM89]. current-distance [Rat87]. current-voltage [Cha80c]. currents [Bar87, CS84, Leu87b, Lev80b, Mat84, RR88b, Str85c]. curvature [BB88a, BP81b, Eng89, Tod85]. curve [Bar80a, IYS83, IHTN88, JP89a, Per84]. curves [BM87, BAB84, Cha80a, DFS89, DCR82, Dug86, FDS^{+}88, GS81, Nor82, PWH^{+}88, Pea88b, SL87a, SNW81, VA85, WBKW82, Wei83, Wei85, WG81, Zno83]. cuticle [Ale87, Cri81, IAHA80, SI81]. cyanobacteria [Nit86]. cybrids [Har82b]. cyclase [Blu84]. cycle [AMM80, All85b, And80a, Boy83, BC81a, BK83, CCG^{+}82, Coo87, Coo88, De87, Der84, DS84, EP86, För88, Hao82, HCVM88, HC87, HC89, Jon85a, KFSW82, KDAT84, KB88a, KSH87, MS80, MHI85, MHT88, Men82, Nor89, NL86, PS80b, Pet80, RP84, RS82, Rot82b, SS88, SC83, SM87, Tys83a, Tys83b, TH85, TD86, VA85, VA88b, WS82, vZvdSdL81]. cycles [AK86, BHMR85, Cal83a, HE84, Hyv84, KK86b, Len84, SN80, SC84b, Ste81, Voi85]. Cyclic [Rei84, Cla81, Fel80, GVP^{+}80, Gol83a, LH80, Mul85, OH83, Sch83a, VGP^{+}80]. Cyclical [Bul82, Cha82b]. cycling [Fin82, Kir86, San84]. cyclostat [GF81b]. cylinder [TW80b]. cylinders [LL82, LL84]. cylindrical [Ale87, CH80, Hun85a]. cysteine [PA86]. cytochrome [CH82]. cytochromes [PSL^{+}85]. cytokinesis [Hun85a, KDAT84, WB83]. cytometry [Whi80a]. cytophotometry [Sch83b]. Cytoplasmic [CT81b, BS84a, Bre80b, SE80, Wer87]. cytoskeletal [Sco84b, Wag89]. cytoskeleton [Luf82]. cytosociology [WK81]. cytosolic [FQ81]. cytotactic [Hje86]. cytotoxicity [DH85, GPKBG89a].

D [MDBRM87, Por83, Mil87, Was86a, WMR86]. D-cell [Por83]. D-loop [MDBRM87]. daily [JT84, PJ86]. dam [Gri82d]. damage [Cra80a]. damaging [FB82a]. damped [LS87, SL87a, SL87b]. damping [SB86]. dams [Gri82e]. Dance [Ros84b, Ros80a, Ros80b]. dane [Bec80]. dangerous [Col82]. Daniel [Ano80j]. Darwinian [Abu85a, Abu86]. darwinism [Wic85a]. dasyclads [HK88]. data [BM87, BBM87, Bel87, BG86, BW81, BB84, Boy83, BT81, Cha89b, Cha88c, DS87, Di 88, DG80, Dug88, HM85b, JS82, KB81, Mae84, MUC81, Oli85, RC81, RM86, RW82, Sar87]. daughter [HCVM88, KDAT84]. day [Cla82b, Pic89]. dealth [MH81]. death [Hir80, Hir82, KG82, Suc83, Uma82, Wol85]. debris [Sch83b]. Decay [Car81a, AA83, Hir84a]. December [Ano80-28, Ano81-30, Ano81-34, Ano82x, Ano82-31, Ano83-28, Ano83-33, Ano84-29, Ano84-35, Ano85-27, Ano85-34, Ano86-27, Ano86-34, Ano87-33, Ano87-34, Ano88-35, Ano88-42, Ano89-28, Ano89-34]. Decision [EL83, Coo81, CK82, MH80b]. decision-theoretic [CK82]. decline [Ful81, JP89a]. decomposition [Mei82]. decondensation [JS80, Tak89b]. decreased [Mur85]. deduced [MDBRM87]. defect [CH80, LP88b]. defective [FM86]. defects [DW82]. defend [For85]. defense [FM86, MH80d]. defenses [Len84]. deficiency [NMM84]. deficient [BRRK87]. deficits [SD83]. defining [Wag87]. definition [Dix83, HG87, Jøn85c, VC87]. deflection [Ber82]. deformation [Dob86, MB84b, TL88]. deformations [MK88, PF89]. degeneracy [Sib83, TRS84]. degradation [Mad88, MG84a, WGHY82, Whe84]. degree [Bas81, SM88a]. degrees [KH80]. Dehydration [Sto81]. dehydrogenases [FLNF86, Gho82]. delay [ADG80]. delayed [KM81]. delays [Mah84a, Mah84b, SCB81, SVN89]. Deleterious [Kon88, Yok85]. deletions [SK84]. delivery [Pop82, SHBS89]. delta [FB82a]. delta-lysin [FB82a]. demand [Rob89]. demands [MR86]. deme [Tre83]. Demiarcs [BH81c]. demonstrate [McG84]. demonstrates [GTH88]. Denaturation [BRM82, MM83c, VM82]. dendritic [KT83, Sch89]. Density [Asm83, CD87, Hol85, Nam80, Nun80, Wat80, Ada80, CD84, DW81a, Dra81, HINT85, Inc83a, IT80, JJ80, Kno89, Lev80b, Lew82, PP85, Pro80, Res82, Ste81, Sve82, TO80, Van84]. Density-dependence [Wat80]. Density-dependent [Nam80, Pro80]. Density-independent [Hol85]. departure [SPKJ89, ZB84]. Dependence [AZE83, AG89, Has88, MR86, AK80, Bas80a, CC82b, CC83a, CDC88, CCDF81, CD87, FEC87b, FEC87a, Ger85, HS87, KF84, Lan82a, LPS86, NSW82, RM83, Wat80]. dependency [Aid88, KM81, Rap84]. dependent [ASK89, Ada80, AWP80, Asm83, Bru85a, Cre88a, Cre88b, FT87, Fer81a, Fer85b, FS83, Gre84a, HKK82, HINT85, HR80b, Hir80, Hir82, Hou87, IT80, KTS89, Les86, LB87, Mal88, MH89, Mil87, Nam80, Pro80, RMS81, Ros84a, SSM81, SH80b, SMGD80, SGd84, SF83, Tak89b, Wee84b, WGP^{+}88, WWG^{+}88]. depletion [MTM86a, MTM86b, dBH87]. depolymerization [RMB81]. deposition [Gro82]. Depression [And80a, Mul87]. Deregulation [Mor85b]. Derivation [VA85, Cru80, HT80, Kat85, Peu85, RW82, SL81, Wei88]. Derivations [Val85]. derivatives [Bar89, JJ80]. derived [BGB89, Cou80, SJ85]. Deriving [SJ85]. dermis [SA81]. describe [AS86, Nee89b]. described [PP84a]. describing [Cas81, DFS89, DRB86, Top88]. Description [SB86, AY85, Ash86, CTS86, FE85, GT80, JF85, KP89a, Pel84, SDM^{+}82, LS87]. Desensitization [Ger83a, Ger85, Lef86]. Design [BE86, KA89, BMR89, Eco83, EC81, HAG82, Koo85, Lau82, MHT88, Rog84]. designed [McG84]. designing [Gar89]. desmids [HK88]. Destabilization [RN84b]. desynchronize [Nan86]. detected [AWV^{+}88, Der87]. detecting [DB88, For81]. Detection [AVR82, Hua80b, BW89, CA89, RB82a]. deteriorating [HC81b]. determinants [BGB89, Coh84, Frö88, GPM89, Reg88]. Determination [DW81a, Fai89, FS86, JA88, Leg87, MLS80, Smi84, BS88a, Bog87, CB89a, Gai80, GS80, Mag88, Mit88, Miy86, Sto89c, Top88, WIN80, VGP^{+}80]. determinations [HM80]. determine [CF87b, Nus80b]. determined [Kor89]. determines [Bre86]. Determining [Zuc86, AL87, EAL80, KM85, PPS86, Rap84, SW81b, VdV87, Wei83]. Determinism [PDC84]. deterministic [GW83, RR83b, Rot82a, HS81]. detritus [Car81a, Cou80]. deuterium [Tat88]. Development [Goo82, LPS89, Abu88, Bel84, BJE89, Bru85a, CM86, Cum85a, Dea80, Fer85a, HC81a, HLS85, HLS89, JJN86, JPL83, KDW88, LJ88, LR83, Nag89, Ren82, RSE81, Sco89, Sla85, Ste80, SW84, UK82, Ula80a, Ula80b, Wag89, Wee84b, ZR86]. Developmental [Hol83a, KG82, Ker89, Che84b, GF81b, LD83a, Ras87, Stu86]. developments [LJTH85]. developmentt [SH80b]. Deviation [IHTN88]. deviations [Nus85]. Devolution [Mer82]. di-saccharide [YBBL82c]. diabetes [BRRK87, Wee84b]. diagnostic [MGL81]. diagram [Sim80a, Süh89]. diagrams [AK87b]. Diallelic [Rog83]. diameter [Hor80]. Diapause [Tay89a, HLS85, HLS89, Sco89, Tay85a]. diatom [Gor87]. dichlorophenyl [Aki85]. Dictyostelium [HC81a, Bel84, PO81, PO86, SK86, Wri84]. Dielectric [FH89, Kri89, Kri85a, War86]. Diet [LSH88, Lim88]. diethylphosphate [Aki85]. Difference [IYS83, PA86, RT88, She80, She81, WdlT84]. differences [And86, Cla82a, Ebe87, EB87, HJS84, Hoe84, LS89, Moy82a, PDCD88, Tra86]. Different [CB89b, Wol85, BESD84, DCR82, Hoe80, Ing89, KYU^{+}86, KK81, MGL81, MTM86a, MB88b, Niv83, Rya80, Sea80, Vil83]. Differential [LS82, Shu84, Yok84, Blu83, Duf86, NBZ^{+}83]. differentiated [Rap84, Wag89]. Differentiation [CY84, PO86, Rei83b, Bar83b, Bod88, Hua80b, KS84, LC80, Mit89, Nes84, San84, Sve81]. differentiative [CM86]. difficult [Day83a]. difficulties [BBM87]. Diffusion [BH83, HC81a, HTG89, Tuc87, Aal89, Agm85, AWV^{+}88, AL87, BKJT80, DfP87, DSR89, Eva80b, Fle80, GM84, Har87b, HK88, HF88, Hun85a, KSS81, Lap81b, LR80, LL82, LL84, ME82, MB84a, MN85, Mur82, MS86b, NM82, NM85, Nag86, New80, New83, NFP88, Nor82, PV83, Shi80, SCH84b, SHBS89, SEN82, SN85, SSN86, SKF83, Tro83, TC80b, WGP^{+}88, WWG^{+}88]. diffusion-adsorption [AWV^{+}88]. Diffusion-induced [HC81a]. Diffusional [GC83, SA86]. Diffusive [Eva80a, Rot82a, GA80, Pop82]. diffusivities [Mal88]. diffusivity [SR81]. digesta [Fai89, FTDS85, FS86, Mat87]. digestion [Kar88b]. digests [Kar88a]. digit [Hol83a]. dihydrosphingosine [WMR86]. Dilemma [FT87, Boy89, Hop83]. dilution [Koc86a, SR89]. dimension [SM87]. Dimensional [Cal82, GZWR80, GM84, WGRZ80, AL83, BB81b, CSH^{+}81, DG85, FDR83, FS85, Hun85a, KO83, MHII83, NHT88, Nus80b, RdL83, Row85b, SK85a, SP87, SW81b, Wes83b, YMIH86, Zam81]. dimensionality [Mor80, PS81b]. dimensions [Com82, MR86, RM83]. dimer [Whi82]. dimeric [PJL84]. dimers [SJ89]. dimorphic [Rei86c]. dimorphism [LA85, Rei86c, SS82]. dimorphisms [Ell86]. dinucleotide [AS87]. dioecious [Lew82]. diol [IO83, ITK88]. diol-epoxides [IO83]. dioxide [Fox81, NL86]. dipeptide [GVP^{+}80, RTM81b, VGP^{+}80]. diphosphate [Tor80, ZBL80b, Szu83]. diploid [Gri82a, Gri82f, Hin80, HB84, Hin86, Hin87, Row87, Row88, Tre81a, Tre82, Wad82]. diploids [Har85]. diploids-who [Har85]. Diploidy [TH81]. dipolar [KB83, Pen80, Sch84a]. dipole [LP82]. dipoles [Wes83b]. Direct [SC84b, And80a, BHP82, Her83b, MG84b]. directed [Bre80b]. Direction [IPS83]. directional [Bar81c, DRL88, DD85b]. directionality [PP84a]. directly [RR84]. disassembly [PHTCS86]. disc [AL81, MRB83]. discern [PS82a]. discerned [SSW81]. discoideum [Bel84, PO81, PO86, SK86, Wri84]. discontinuous [Ble88, HE80a, MZM89]. discrepancy [Sýk84]. Discrete [CPPR87, CPPRS87, EM81, KR82, Lew81a, Rég84c, SM80a, SKF83, ZG89]. discrete-time [SM80a]. discriminant [Iid88]. discriminate [MGL81, Mic86]. discrimination [AS87, BM87, BMR89, BT84, CPLA85, For85, FS80, Gen88, KS88, dBH87]. Discussion [Mor85a, HS81]. disease [Bar86, Dil86, Ewa80, Joh86, TATS81, YE82]. diseases [Bre83, Rég84c]. disintegrating [Kle81]. Dislocations [ST80, HE80a]. dismutase [BK81, OO84a]. disorder [BB85]. disparity [SCP83, Sho85]. Dispersal [Fra86, CB86b, CHM80, Com82, Kra82, Nam80, Tay88a]. dispersion [BRR87, FSM80, Tho82]. displacement [DD85b, Gro80, Smi82c]. displacements [Res82]. display [LBW86, Maw85]. displaying [Hir82]. displays [And80b]. disposition [BIS^{+}80, Wei83, Wei85]. disregarded [CB83]. disrupted [Mer87]. dissimilar [DD85a]. dissipation [Peu86]. Dissipative [Mal88, TCY85, Wat85, Ji85, Ric80]. dissociating [KK82a, Kur83]. dissociation [HM80, KD80, Lef86, WSMP83, Whi83]. distance [HKC83, HCKB83, MLS80, Pen82, Rat87, Sou80]. distances [Fai85, GRS82, Ree80]. distension [FBP87, Win89b]. distinct [CMN81, KB88a]. distinction [McC80b]. distinctive [Sib83]. distinctness [GH83]. distinguishing [All85a]. distributed [BW85, HK80, Rot82b, TW80b]. Distribution [SN80, Alm85, AS86, AZE83, BJ89, BGG84, CDC88, CT81a, DV83, GKB86, Gra80, GP82c, Ham85, Har84b, HW80b, KM81, KF84, Kor89, KS81, LP83, Mil89, Nam80, NM80, OK87b, Pec88, Rab82, Rei81, SA81, Sax83, Tra85, VL86, Vla87, Wei83, ZG89]. Distributions [Sta81, Aug85, EH83, GPKBG89b, GWK87, KWG87, Ros85, Rou80, TH85, Vau86, VA85, WR83]. disturbance [Arm88b, PGKE87]. disturbances [CLF84]. divalent [Kol82, ST88]. Diversity [SKK84, BW81]. dividing [Lap84, OWN89]. diving [YC89]. Division [CM88, Ber83, BS84a, BC81b, BC83, Bre86, CZJY80, Coo88, Dou84, HR86, Hir83, HYDS84, IBJ^{+}89, KDAT84, Lap84, TD86, Tys87]. divisional [Rei81]. Dixon [FLR83]. DNA [Aid88, Alm83, Alm85, BJ84, BB85, BPP^{+}86, Bec80, Bel80, BN81, BGGV83, Blu83, Bod88, BJE89, BHM^{+}85, BC83, Bro82a, CS87, Cla82d, Cla82c, Cla82f, COR83a, CCG87, Cra80a, DML80, DD89, Dur82, For80, Fuk83, GF81a, Gar80, Gat86, GP82a, GMTG80, GPPZ80, Gra81a, HRG82, Hop83, HT89b, Hua80a, Hua80b, Kar88b, Kle81, Kon84b, Kyp88, LN85, LP86, MRG84, McC84, MM82b, McG89b, Men85, MdMF83, Mic86, MM83c, ML86, MH80c, MM82c, MMG84, ND82, NBZ^{+}83, Nus82a, Nus83, Nus85, PR82, PY88, RSR88, RGR82, RMFM85, Rei81, Rei84, Rob85a, RT83a, Row85b, SCP83, SP88, SBM89, SL86, SV84a, SM81c, Sto83, Sto89b, Tak87, Tak89a, Tak89b, TS80, Whi80a, WS85b, ZC89].

DNA-fluorescence [BGGV83]. DNA-protein [DML80]. DNA/RNA [Gar80]. DNAs [MDBRM87, OK82, SW83]. Do [LH89, LS89, MP87, Smi82a, Var83, Cla86, EGS80, Gul82, Har83, Hoe80, KC85, LBW86, Nan86, SR89, War85, Esh84a]. doable [MAC^{+}83]. doable-minate-mediated [MAC^{+}83]. docking [SLFT86]. Does [Ber87, Luf82, Nee89b, Ola82, Par85, SS81, Bar88, BHP82, BK83, Coh85, Gri83, HE80b, NS86, PS84, Was86b, YI88]. dog [Aki85, KS82, LTM85]. Dollo [Wag82]. domain [Bod88, BJE89, KvDVH87, LL84, RR88a]. domains [AK84a, AKS84]. domestication [LP88a]. Dominance [CB87b, Kac87, Bri82, Bur83, CS80, Hug86, Jam85, McM81, MB85, WB85]. dominant [Mor82]. Donnan [OK87a]. donor [HFO83]. don’t [Arm84]. dopamine [KRB81, Was86a]. dopaminergic [KRB81]. dormancy [FEC87b, FEC87a]. Dörner [Jam89a]. Doryteuthis [MUI85]. dosage [CB86c]. dose [BM87, BE86, dGV80]. dose-response [BM87, BE86, dGV80]. Dosimetry [OC81]. Double [MM86a, BBM87, CCG87, DFS89, Dug88, GDHB87, HT89c, HT89b, McG81, SLYF84, SW83, VA88b]. double-exponential [DFS89]. double-strand [HT89b]. double-stranded [CCG87, GDHB87, SW83]. doublet [CBLM81]. doubling [AS84, RL88]. Doubly [SR89, Spe87, SR87b]. doubly-labelled [Spe87, SR87b]. doubt [Nor83]. down [Leo84]. dragonfly [KSL89, RH85]. drift [Has88]. Driven [MS80, Dou84, GGPM80, Gra85b, Hao82, Sch83a]. Drone [Omh88a]. drones [Omh88a]. droplet [TY82]. Drosophila [AL81, BKJT80, Dea80, DRB86, MM82a, MRB83, Ras87, CB86c, Nag88]. Drug [Rei83a, BSR^{+}88, BIS^{+}80, Ger83b, Ger85, GPM89, MHII83, PGKE87, PGKE88, RT85, SRO84, SHBS89, Wei83, Wei85]. drug-receptor [Ger83b, GPM89]. drugs [Aki86, PS82a, SBW83, SHBS89]. dry [FBP87, Mäk86]. DSB [PL84b]. DSB-repair [PL84b]. Dual [Bec85, Lum84, MDPF86, Ses82]. duct [JD82]. ducts [BLA82, MV88a, MV88b]. due [CCG87, CL84, HRG82, Orl82, SR89]. duplexes [GPLC84, Ste82]. duplication [Nus82a, Nus83]. durations [Pea88d, VA85]. during [AAAT87, Asp88, BK83, BH88, Coo88, CM86, De87, FQ81, Gho82, HYDS84, Hun85a, Jac83, JC80, JS80, Kar88b, KDAT84, Kin83, Koc89, KDW88, MG80, ML86, Moy82b, PO81, RSSK87, SRO85, SCP83, SM81a, SA80, WT87, WFL86, WD86, YC89]. dwell [KSW^{+}85]. dwell-time [KSW^{+}85]. Dyar [HT84]. dye [Gho82]. dye-linked [Gho82]. Dynamic [Bri80, Gal84a, Gal84b, HC89, OKT^{+}80, OH83, PG88, RL84, ST86b, AS86, DG85, Han85, HM87, Jam80, KLR89, KKKK80, MK81, Row87, Row88, SK88a, Shi87, ST86a, Sim81b]. Dynamical [FT87, Ada80, Cre88a, Fer88, FE85, Hay80a, LS82, YMS84, KN82, KN83]. Dynamics [DGST82, Har84a, INMM83, Ric87, Swa82, Zee81, ASTT80, AG89, BKMS86, BKMS87, BG88, CPPR87, CTS86, CM86, CM87, CN80, Ebe88, Get84, Gin86, GGPM80, GAD80, GGC86, GH82, HH89, Har84b, HBP89, Ive88, JP89b, JP89c, KK82b, LTM85, MR86, McN87, MLB^{+}89, MS86a, Mur83, Omh88b, PL84a, PJL84, PPL85, PL85a, PL85b, PG88, PHTCS86, RM83, Rég84c, RS82, Rog86, Rog89, SK85a, SS83a, SH83a, SM80a, SD81, SS89b, Tay88c, TG83, URC83, Vla87, Web81, dBH86]. dynein [MS86a, MHB89]. dynein-tubulin [MS86a]. dystrophies [Inf85].

E. [BJL85, DS84, Ros85, YI88]. each [Row88]. Eadie [FLR83]. ear [LM82, PP84a]. earlier [SC86a]. early [BBH84, EGS80, KDW88, Ste84b, SD87, WFL86]. easily [Pea88d]. easy [Nun81]. eccrine [JK89]. echinocyte [ABB89]. echinoderm [Van82]. eclosion [PS80b]. Ecological [LK88, PDCD88, SC87a, COG83, CN84, DE82, Her81b, HU85, Lil82, Sch81, SG88]. Ecology [PS81b, Bot82, Cal82, Gla87, HD84, MW83b, MMGL87, Pah85b, Pah85a, Tay88b]. economic [Her81b]. Economical [TC87]. Economy [MHT88, Blo87, CF82]. Ecosystem [Fin82, Han86, MSA^{+}81, RR84, RL82, RR83a, RR83b]. ecosystems [BP80, Han85, HBP89, Kra82, Tan84]. ecto [TKB80]. ecto-enzymes [TKB80]. ectotherm [Par83b]. edges [Pin83b]. editing [Lam84b]. editor [McC80c]. Editorial [Ano80d, Ano80e, Ano80f, Ano80g, Ano80h, Ano81c, Ano81d, Ano81e, Ano81f, Ano81g, Ano81h, Ano82b, Ano82c, Ano82d, Ano82e, Ano82f, Ano82g, Ano83a, Ano83b, Ano83c, Ano83d, Ano83e, Ano83f, Ano84b, Ano84c, Ano84d, Ano84e, Ano84f, Ano84g, Ano85a, Ano85b, Ano85c, Ano85d, Ano85e, Ano85f, Ano86b, Ano86c, Ano86d, Ano86e, Ano86f, Ano86g, Ano87b, Ano87c, Ano87d, Ano87e, Ano87f, Ano87g, Ano88b, Ano88c, Ano88d, Ano88e, Ano88f, Ano88g, Ano89a, Ano89b, Ano89c, Ano89d, Ano89e, Ano89f]. editors [KW86a, KW86b, KW86c, KW86d]. EEG [AZE83, KHU^{+}83, Tou85]. EF [BA87]. Effect [Bar80a, Car81b, CC83b, COA85, GPKBG89a, HM80, Hsu88, Inc83a, Rot87, VL86, WSMP83, Win89b, YE82, ADdACP87, AdMM86, And81, AA84, Bar87, Ber85, BS84a, BF88, BK81, Cha89a, CG81b, Cra82a, DLN85, EM81, EL87, Fle80, Fra88, Fri83, Ger83b, HINT85, Hol89, KM83, LSFT85, Man81c, MKS^{+}86, MTM86a, MTM86b, Mot82, ND82, NSKS85, Orl81a, Orl82, PS82a, PGKE87, PGKE88, PH87b, Rab82, SI81, Ste88, Tam87, TGP^{+}86, Wad82, Whi82, Kri80]. Effective [Kri85b, Dem81]. effectively [Gar82, Gre82b]. effectiveness [Reg88]. effector [SM88a]. Effects [BCL83, BC86b, BR83, Hal80a, HdL84, LE84a, Ryt87, SR80b, TL88, Arm88b, BSR^{+}88, Bec81, BW89, Bor84, CG85a, CS84, CU88, Cri81, FJ85, Fer85b, FE85, Gre89, HTG89, HFL89, HLS85, JSJS88, Jol88, KM81, KFPB85, KS80a, Koo85, Lam88, LR80, LL82, LL84, LvH87, Maw85, McN87, McN89, NU82, NMB85, NSW82, PYN85, PS80a, RC89, Rei81, RN85, RR88b, Sch82b, SK84, Str85c, URC83, Wig81, WdlT84, ELM81]. efficiencies [Sch81]. Efficiency [Dus89a, Jur84, KL81, Sar82, FJ85, HN85, RN86, VdV87]. Efficient [TJP83, Wat84, Got86, LVN87, MH87, Nus83]. effort [BS82, HD84]. egg [HLS85, HLS89, Sco89]. eggs [CVNO87, TS84]. eioenergetic [RS83]. Elaboration [Gre80]. elastase [CMM86]. Elastic [Eco83, Eco84, Hok86a, Pug87b, RL82, TLS83, Wax81, iWSS88, Pys80]. elastic/geometric [Eco84]. elasticity [PYN85, iWSS88]. electric [Bla80, Coo84b, HGG82, Rab82, RSM80, Ste88, Tou85, Wes83b, ZG82, ZG89, RN84a]. Electrical [KM87, Win89a, CC83b, EM81, Fer85b, GPAL88, KT83, KP85, LO82, LM82, MR89, Wil86a]. electrically [Rat87]. electrochemical [DG80, She80, She81]. electrodiffusion [Mal88]. electrodynamic [See86]. electroencephalogram [Sie81]. electrogenic [Cha82a, HFL89, JCK80]. electrolytes [Cha89a, Web83]. electromagnetic [PJK^{+}83, Ten80, TG88b]. electromechanochemical [AC89]. electron [BK81, DT80, Gho82, Kam80, Lac84, Res82, Sar82, Sch83a, SR80b]. Electronic [Cri81, Res82, MLS80, VI80, YVIG80]. electrophoresis [Can87, LO81, MN89b]. electrophoretic [Can87, DV83, Gra86b, SFR86]. electrophoretical [DG80]. electrophysiological [JCK80, Pin83a]. Electrorheological [Gre89]. Electrostatic [NSKS85, OK87a, Pen80, ANS83, BRW89, Cha89a, DV83, GPM89, KB83, KS81, LM82, Mon83, PYN85, PSL^{+}85, VDH82, War86, Ler83]. electrostatics [Kri89]. electrostimulation [Rat86]. element [Pic89]. elementary [Dur82]. elements [Hic89, Kre84, ND82]. elevator [Orl81a, Orl82]. eliminate [KC85]. elimination [BRR87, DH85, GB80, JK81]. Ellipsoid [BK88, Kwo89]. Elongation [GWK87, BH88, Jak88]. elucidation [NSJT81]. elution [BPP^{+}86, NBZ^{+}83]. embolized [Pic89]. embryo [Nag88]. embryogenesis [GT80, MP85a]. embryological [Ste80]. embryonal [SS88]. embryonic [BS84b, Sla85, SG82, SJK88, Wis80]. embryos [Pat84]. emendation [Mal85]. Emendations [GRJ81]. Emergence [ABEFS^{+}86, BBHM85, CH85]. emission [DBB80]. emit [LBW86]. emmetropization [Car82a]. empirical [Frö84, SK85b]. employing [Val85]. empty [Bel86]. enabling [MV88a]. enantiomers [Gra85b]. encapsidated [AVR82]. encephalization [Hof82]. encoded [RMS81, Stu86]. encodes [Mer87]. encore [Eco84]. encounter [RC89]. encounters [JK88, TS86]. end [EKSS88, Koc89, MKL81, Top88]. end-plate [MKL81]. end-point [Top88]. endless [Len84]. endogenous [MSBB84, SL80b, Whe84]. endolymph [MV88a, MV88b]. endonuclease [Kar88a, Kar88b]. endoplasmic [ACB81]. endosomes [LL88]. endothelium [AWP80]. endotherms [PJ86]. ends [Ash85, Bel80, Tay85a]. energetic [PJ86, WS84, Wat85]. Energetics [AK87a, AK87b, CRB89, HB88, KVW88, Ano80j, KB88b, Min82, NU82, Sch81, Sea80]. energies [HR85]. energized [Mal84a, Mal84c]. Energy [BH81a, Bar81d, Bar82b, BE81b, BS84b, CN84, Her81b, HT83, Koo86, KKJ81, Lay87, ML86, Mul85, Aid88, BS85, Blo85, Blo87, DV83, EB87, Fox81, FDC^{+}89, Frö84, GPLC84, Gla80, HN85, Hig85, Jur84, KDG85, KK85, KvDVH87, Kot82, Kot83, Kri85a, Kri85b, LBBL84, MB80, Mal84a, Mal85, MLS80, MYS89, Peu83, Peu85, Peu86, Pro84, RTM81a, RTM81b, RTM82, Ryt87, SRO85, Sav81, Sch84a, SK84, SOLG82, Tra86, TG80, VDH82, VI80, Wel85, Wil84, ZG89]. engineered [HL89, Koo85]. engines [Dun82]. engulfment [Wis80]. enhance [Ker89]. enhances [Pin83b]. enhancing [BBH84, RS83]. Enlarged [Ano81i, KRT81a, KRT81b]. entities [Hau82]. entity [CMN81]. entrainment [Hao82, LK83, YVIG80]. entrainment-beating [Hao82]. entrapped [KLK^{+}88]. Entropy [Aok89, Cla81, Phi81b, BW86a, Bor88, BS84b, BLC84, DSD81, Dou84, Jur84, Mar84, PDCD88, Rya80]. entropy-driven [Dou84]. entry [McN87, PS81a]. entry-exit [McN87]. environment [Cha88b, CB85, GN86, HC81b, HH81, IM85, JP89c, Kon84a, Lei81, LSFT85, MN89a, Nam84, Niv83, Reg86, Tak89c, TC87]. environmental [ÅF80, CB89a, HG80, Str80, Was86b, HdL84]. environments [AS86, BR86, CB81, MGH81, Rot82a]. enzymatic [Agm85, Bar82b, KK85, Kri85a, Kri85b, Met82, OKT^{+}80, OH84, RF82, SH87, Vol81]. enzymatically [SG89b]. Enzyme [CB84, Pet82, RN84a, RN84b, AK82, Aki84, AK87a, AK87b, Bar83a, Bas83, BW89, CP85, Cho81, DSW83, EB87, Fri85, GVC81b, GVC81a, GV81, Gar89, Gol83a, Gra85b, GG82, Gre80, Hav89, Kar80, KVW88, KK82a, KB89b, Kri85b, KC83b, KL80, KCM81, Kur83, Kur84b, LR80, LQC89, MM86a, MM82c, OH83, Ola82, Ola86, Pen80, PTD81, Rak80, Res82, RN84c, Ric87, Rog84, SM88a, SF81, Shi82, SC89, TCR85, Top88, Tys83a, VCC^{+}88, WT87, WBKW82, War86, Web81, WD81, WK81, WSMP83, WKV88, Whi83]. enzyme-catalysed [Cho81]. enzyme-ligand [WSMP83, Whi83]. Enzymes [KLK^{+}88, Bun80, Can87, CBC87, Dug86, Gho82, HFL89, HL89, IHK80, KK82a, Kri80, Kur84a, KSM85, Lan82a, ML82, ML84, MAL87, PJL84, RP84, RN82, RN85, RN86, Sch84a, SPS83, TKB80, WT87, Yos86]. Enzymic [Lam84b, ACB81, ME82]. enzymology [Bun80]. epidemic [AS84, SK85a]. epidermal [ATC81, AL81, Sai82]. epidermis [HO84, HM85b, PM80, SA81, WF82]. epigenesis [MA83]. Epigenetic [Sac88, Eld84, JL89, Ste80, Ste82]. epileptic [KA85]. epimorphic [Lew81c, Pap84, Sib81]. epiphyseal [HB89]. epipubic [Whi89b]. Epithelia [SG82, AL81, FM88, Sla85, VG80]. epithelial [BBZ87, MF81, MTM86a, MTM86b]. epithelium [MM83d]. epoxide [ITK88]. epoxides [IO83]. equality [Smi87]. equation [BBM87, BE86, BB84, Don86, Dug88, Eng89, Ful81, IYS83, JK87, KSS81, RT88, SM88a, Shi80, Whi80b, Whi81b]. equations [Cha80b, Cla86, DDA85, GA80, HBRN81, HC81b, HS89, HS81, PK82, Peu85, Peu86, QL82, RC81, Reg84b, RCA89, RSM80, SPS83, Shi80, Shi87, Str85a, Str85b, Str85c, WBKW82, WdlT84, Zaj86]. equilibria [BR87, Cra89, DLN85, Fox81, NR85, Shi82, VC88a, Wel83]. Equilibrium [AH86, KDG85, PP85, Yok83, Hin87, Ish81, KK87a, KC83b, Orl81a, Orl81b, Ril80, SR87b, Tra85, WB85, Whi82]. equivalence [Stu85a]. equivalent [LP82, Sch89, Vis87, WB89]. Errata [Ano80i, Ano87h, Ano87i, Ano88h, Ano88i]. Erratum [Ano81j, Ano81k, Ano81l, Ano82h, Ano83g, Ano83h, Ano83i, Ano83j, Ano85g, Ano88j, Ano88k, Ano88l, Ano88m]. Error [BJL85, Kir80, NS89a, Bar80a, Blo85, COG83, Dur82, For81, GP80, GTMM87, GTSNO^{+}88, Rea84, RW83, Zno83]. error-correcting [RW83]. error-correction [Blo85]. error-detecting [For81]. error-screening [Rea84]. Errors [SR89, Tre83, DPV83, Lef86]. eruption [Hor89a]. erythrocyte [DG80, HGG82, LP88b, PYN85, SME86, Zno83]. erythrocytes [BK88, PCTF83, SME86, Tus88, Van82]. escape [DH85, GB80, KM83]. escaped [Kle81]. escaping [Bec80]. Escherichia [BC81a, BC81b, BC83, Bre86, CM82, Cla82a, GWK87, Jak88, MZ89, Man81a, Man81b, NHT88, Nag81, Nor89, WP84, WGRZ80]. especially [LM82]. ESS [CV88, Cla86, GM89, Hin87, KLR89, Orl81b, PK80, Par83a, VC88b, Vic89, Vic87]. essential [HM80, RHH85, SH80a]. establishing [Jam89c]. establishment [KS84]. estimate [Cor82, SK84, SR89]. estimated [Nor82]. Estimates [Bas83, Bas80c, Law84]. Estimating [FDS^{+}88, HT89c, KB88a, OWN89, Tay85a, CF87b, JA88, SK85b, SCB81]. Estimation [DBB80, Lev80b, SR81, VA88b, Bar80a, BE86, BSB88a, BSB88b, CB87a, Dra81, EC81, GRJ81, HBRN81, KK83c, LR85a, Lef86, MS83, RJ80, Whi80a, BSB88a]. estrogens [BHM^{+}85]. ethanol [Der84]. etiology [Inf85]. eucaryote [Kin85]. eucaryotes [HR80b]. eucaryotic [Hua80b, Wri80]. Euclidean [Peu86]. Eukaryote [Luc82, Bur82b, Iid88, NDB80]. eukaryotes [AMM80, Hir84b, MM82b, ND82, Spe84]. Eukaryotic [LN85, MDPF86, Wil81a, Woj83, Arm87a, Bod88, BJE89, BT87, Coo87, CBLM81, De87, EP86, GMTG80, HB86b, KvDVH87, LP86, Nus80b, Nus82b, Nus86, Omo89]. eurkaryotes [Bee80]. Eusocial [Har85, WL87]. eusociality [Abu81, Cra80b, Cra82a, Cra83b, Gra86a, Iwa81]. eutectic [SA86]. Evaluating [Jol88]. Evaluation [KB89b, RM86, Aid88, CMM86, HW80a, Kap85, KM85, KK86a, Mic86, Mur87, SR89]. evaporative [Dam86]. evasion [WW84]. even [CG84, HO84]. even-numbered [CG84]. event [Wri80]. events [Ger83b, IT82, SM81a, Vro83, WFL86, You85]. Evidence [AS87, BHMR85, Bro80a, Bro83, SS88, SG86, VTL85, MHLK87, NBZ^{+}83, TDA86]. evoked [MKL81, PS84]. Evolution [Amt86, BSM82, Cha88a, CF85, Cra82a, Cra82b, EL83, EL87, FM86, GAK88, HG80, Hir80, Hir82, Hoe82, HJS84, Hoe84, Hof83, LA85, vdHN85, RN82, Abu81, Abu84, ADdACP87, Arm87a, Bel86, Bod88, BJE89, BB81c, BPH87, BR88, BLC84, Bur83, Bur82a, Cha88b, Che84b, Cla82b, CV81, Cra80b, CJ82, Cra83a, Cra83b, CK85, Cur88, Dav87, Dem83, DC87a, Duf86, ER88, Eld83, Emi82, Esh84b, Esh84a, FT87, FHL89, Fis80, GM89, GK81, Gla82, Gol83b, Goo82, Gra85b, Gra83b, Gra84, GK80, Gra86b, Gre80, HBA89, Hoe80, Hol83a, IT80, Iwa81, Iwa88, Jaf84, Joh86, KVW88, Kin85, Kno89, Kon84a, KL80, Lam84a, LG89b, LL87b, LTZ88, Mag86, MJ80, Man81c, McC80a, McC80b, Mik83, Moo88, Nee89a, Nee89b, Nee89c, Nod82, NS89b, Oht87, Pai82, PK80]. evolution [Par83a, RGR82, RN86, Rut83, SH81, Sch83a, SS82, SF81, SG88, SC83, SC84b, Sil84, Sla86, Ste85, Suk89, Syv85, Tam87, Tan80, TO89, Too82, Tro84, Ven82, Wad82, Wad85, WB85, WRW83, Was82a, Was82b, WK81, WGS80, Wic80, Wie81, Wil81b, WS85a, Yok81, Zee81, Kri85b]. Evolution-imposed [Kri85b]. Evolutionarily [CHM80, Com82, Llo83, Mat85, Mot88, Sch88, Ste84a, Tre82, GNG80, Har81, Har83, Hin80, Hou83, Ish84, Kon84b, Lun87, MN89a, RS84, Sel80, TC80a, VC87]. Evolutionary [Cla82a, Esh83, Ewa80, HR80a, Les86, Lil82, MI81, Mat83, O’G84, Rei87, Tre81a, AS87, Boy89, Car80, CS80, Cla82e, Cra89, CD87, Cri82, Dav84, DE82, Duf85, ES85, Gra86a, Gra88, Gri83, Har87a, HHP89, HB86b, HB84, Jon85b, KB81, Kon88, PP84b, Ril80, RHH85, Sza83, Wag81, Wic85b, Wil81c, Wil84, vRV80]. evolutiont [BJ82]. evolve [Hin87]. evolved [Coo81, Mil89]. Exact [TACM82, GPPZ80, New80, SGDK86, Shi80]. exactly [Shc89b]. examination [BH81a, CK82, LRKC83, LJTH85]. examined [Cra83a]. example [Cre88a, FM88, KBR88, LP88a]. examples [Ipi87]. exchange [Fle80, Hig85, KP81, KC83b, MF80, NSKS85, Sch84a, Syv84, TS80, WS85a]. exchanges [Bia81, Fox81]. excision [COR83a]. excitable [KKNM80, KK80, Leu87a, LN81, MS89a, MHB89, Rat86, RSM80, Ste88, TY82]. excitable-membrane [Leu87a]. excitation [IT82, KYU^{+}86, KVB83, MLS80, SB86, UKNS87, Won81, Zac81]. excitation-contraction [Won81]. excitatory [KK83b]. exciton [FY81]. excitonic [NS86]. exclusion [Bri89, vdHN85, Sto83]. exclusively [Gre84b]. excretion [Ond80]. exercise [KB89a, Sau80]. Exhaustion [PS84]. exhibiting [AKS84, KA85]. exist [SS81]. existence [AOB85, GA80, Hir83, Hir84a, Leu87a, Nam84, vdH84]. exit [McN87, WGP^{+}88]. Exocytosis [Bre80b]. exogenous [Coo84b]. exons [BBH84]. expansion [QL82, RL82]. Expected [Coo84c, Ber85, Has88, Orl81b, dW81b]. expedience [FSF89]. experience [Hon84, Lib85]. experiment [LS81]. Experimental [BKMS87, TLN85, Aki85, BMR89, FP89, Gar87, GWK87, HK81, JS82, Kat85, LRKC83, LJTH85, MHLK87, NSJT81, NBZ^{+}83, RW82, SSF87, SH83b]. experimentally [SV89, Süh89]. experiments [AK87a, BE86, EC81, Gar89, IP83, JA88, KBR88, KA89, Lan82b, MGL81, PM80, Ros84a, Sýk84, VA88b, WD81, Wis80, WGRZ80]. explain [Koo86, Mor85a]. explaining [Dou85, IHK80, SH82a]. explanation [Bar81c, Eva80a, Fra84, Hux86, LPS86, PS80a, Sea80, Sýk84]. explicit [Cor82]. exploitation [LS88, MG85, Rog86]. exploiting [Hou86]. Exploration [MTM86a, MTM86b, SL81, Süh89]. explore [Kel87]. Exponent [Don86]. Exponential [PK82, Cha80a, DFS89, QL82, SG89b]. exponentially [IP83, MG84a]. exponents [Rei86b]. expose [Mot89]. exposure [dGV80]. expressed [HK88, KTNO81]. expression [BGGV83, BHP82, BK88, CFLF88, CV89, Con83, CL84, FS88, Fuk83, Fuk84, GDHB87, KDW88, Man81a, Mor85b, SMGD80, TMGB80, VL86, Wil81c]. expressions [MTM86a, Rik85]. extended [BRR87, vDvD86]. Extending [LL87a]. Extension [Hyv80, GS86, GWK87, HBP89, PS82b, Pug87b]. Extensions [HB86a, Cas84]. extent [NSW82, Rak80]. exteriors [Duf86]. External [LR80, SSN86, Ped88, RL88, RN84a, Ric87, Spe81]. extinction [CW84, Coo84c, Gad81, Pea88a, Pea88c, Rot87, SM80b]. Extinctive [Emi82]. Extracellular [Fer80, GT85b, BHP82, Col85, Fer81b, Fer81a, HH85a]. extracts [RQ81]. extragenic [ND82]. extramicrosomal [RF82]. extrapolations [BAB84]. extrema [Bri84]. extreme [ANOM82, Dan86]. extrusion [Dal83].

F [Gre85]. F-actin [Gre85]. facilitate [Hav89, Ibe83]. facilitated [MS86b, MB84a]. facilitation [CV81, PS80a]. fact [Mor85a]. factor [AA84, Blu84, Esh84a, Jak88, Kon88, OKT^{+}80, Sab80, Wes83a]. factor-adenylate [Blu84]. factorability [BWM80]. factors [Bri85, FS87, Per84, SK85b, Wer87]. faecal [DFS89, FDS^{+}88]. faeces [FDS^{+}88]. failed [Har82c]. falsification [GC83]. Familial [Cra83a]. families [BP86, Jak88, Pec88, SH82a]. Family [Yok81, AM81a, JK81, VKV83]. Fasman [CC83a]. fast [Shi82]. fatigue [Ale84]. fatty [Kum82, MRR80]. favor [Cha86]. favour [SA80, You81]. favourable [Man82b]. favouring [Ven82]. fd [OK82]. Feasibility [RL82, KD88]. feature [Sim85]. features [KA85]. February [Ano80w, Ano80-30, Ano81-28, Ano81-33, Ano82-27, Ano82-32, Ano83-27, Ano83-34, Ano84x, Ano84-30, Ano85-31, Ano85-37, Ano86-28, Ano86-33, Ano87-29, Ano87-36, Ano88-32, Ano88-37, Ano89z, Ano89-32]. fecundability [MS83]. feedback [Bar80b, Ble88, Ble89, BPM82, GS85, LS87, Mad88, Mah84a, Mah84b, OH84, Spa80]. feedbacks [FBP87]. Feeding [DK83, Ano80j, Hou87, KR85a, Koe83, Lim87, MOV82, Sil83, Spe86]. female [CS80, Cha89c, GC86, LL87c, TO89]. females [Man85]. fern [Wil81c]. ferredoxin [CH82]. ferroelectric [Leu87a, Leu87b]. Ferroelectricity [Bol80]. fertilities [Mot88]. fertility [Abu85a]. fertilization [Man81c]. fetal [FW81]. Fewer [Wel86]. fibers [CC83b, Rat87]. Fiberstresses [Reg84b]. fibre [Asp88]. fibre-reinforced [Asp88]. fibres [Ale87, BM89a, NM82]. fibrin [WFL86]. fibroblasts [Ada80]. fibrous [TS88b]. fiction [AT85, RR84]. field [Bor84, GT80, GL84, Leg87, MBJK89, Mor84, Pin83a, Pin84, PRH81, PR82, Rab82, RSM80, See86, Sib83, Ten80, TT85, TT87, Tou85, War86, Yor81, ZG82, ZG89]. fields [Bar87, Coo84b, Fer85b, GT85a, GT85b, HTG89, Hej82, IBJ^{+}89, JR80, Sib81, YM89]. fighting [EL83, EL87]. filament [CG81a, CG81b, Dal83, Obe81, SS87]. filaments [Ere85]. file [KSS81, KH80, SF83, SKF83]. filling [BW89]. films [DZJ85]. filters [BPP^{+}86]. Finding [WH83, BL88, IPS83, ZR84]. Finite [GT88b, TG88a, NS89a, RH85, Sch88, Smi88, ZL84, dCJM86]. fintess [Smi85b]. fireflies [SC87b]. firefly [SC86b]. firing [CJ83, Fra87]. First [Gad81, AVR82, BS83, Bas80a, Gin86, Mag86, Nag86, SM88a, SCB81]. first-order [Bas80a]. fish [And88, Gul82, Lay87, MOV82]. fishes [Moy82a, Mul87]. fission [CZJY80, JCZ84]. fit [Tho85]. fitness [AM81a, Abu85a, Abu86, BR83, Coo84c, Iwa88, KW89, Ste84a, Tay88a, TC80a, Wei88, Yok84]. fitnesses [Gre84b]. fitter [Sil83]. Fitting [Cha89b, Dug88, BBM87, Bro80b]. five [MDBRM87]. fixation [DW83, Jon85a, Man82b, Smi87]. flagella [BH83, CBLM81, Fue80, FH89, MHB89, Omo89, Rik85]. flagellar [JPGL86, MS86a, MYS89, Rik82]. flapping [Yat86]. flat [KSL89, VA88a]. flattened [Hun85a]. flexibility [NBZ^{+}83]. flexible [Hum81]. Flexoelectric [BP81b]. flies [HT88]. Flight [AO87, AY89, BH81a, Hou86, Hum83, Ray85, WS84]. flip [Nor89]. flip-out [Nor89]. flipping [CK82]. flocks [War85]. Flow [Bas80a, BLA82, AHFH81, Aok89, Bas83, Boy88, CK87, Cla81, EAL80, Fai89, FM88, FTDS85, FS86, Hsu87d, Ibe83, Kin82, Li88, Luk84, Mas84, Mat87, MHII83, MV88a, MV88b, TLS83, TG80, Whi80a, Wil86a, Win81, YMIH86]. flowers [Bel86]. flowing [Lap80a]. fluctuating [CB81, Cha88b, Nam84, Ste81]. fluctuation [DSW83, Ola82, Peu86]. Fluctuations [Ish86, BP81b, SS85, Spe81]. Fluid [CK87, Hay80a, AHFH81, INMM83, KS82, Spe81, Swa82]. Fluid-dynamical [Hay80a]. fluidity [DW83, VD89]. fluorescence [BGGV83, RC81]. fluorescent [DBB80]. fluoroscopic [Kap85]. Flux [BBH86, BM88, BT85, BN80, HM85b, Jam82, Kat85, KD88, Lan82a, NSCH88, NK88, TCY85, Win89b]. fluxes [KH80, Wag87, Win81]. flying [Cor83b, Orl81b]. Foetal [Jam89a]. folded [SR80a, Tod82]. Folding [NP84a, Cha81, CSH^{+}81, LM89, SA80, Tod82]. foldings [KvDVH87, SW81b]. folds [VM89]. follicles [MN85, NM85]. following [GZWR80, WGRZ80]. Food [BL88, HBP89, BP81a, CB86a, HdL84, Hou86, KL81, Leg87, Lev80a, MH89, NC89, Ste84a]. food-calling [NC89]. food-sharing [CB86a]. food-webs [KL81]. forage [Reg86]. foragers [Jan82]. Foraging [Bel87, Hel80, KLR89, Tay81, Gre82a, GN86, Her81a, Hou87, MH85, Reg84a, YC89]. Force [Kas80, BH88, Dun82, Fer84, Hay80b, Koc86b, LL87b, Mal81b, PRH81, PR82, RL88, Tai86, YA89]. forced [AMI84]. forces [Bir81, GBFH88]. forest [WJS89]. forests [FSM80]. Forgetting [Lin80]. fork [MdMF83]. forks [Var83]. form [Bel84, Cum85a, ITK88, KvDVH87, Kor89, MJ88, NST80, Pet83, Rou80, WJT87]. Formal [TS80, McG81, ZR86]. formalism [Coo81, Kri85a]. formalized [Shc89c]. formation [BH81a, BBZ87, BHH83, BG84, BW86b, Bor88, BPP88, BKJT80, Chi86, CPPR87, Cra80a, Dil86, FE85, FF83, GK80, GBFH88, Hor89a, Hua80b, Hum83, KS80a, KIO88, Las87, Mal88, Mea86, MK87, Mur81, NM82, Nag88, NFP88, OCMP89, Pap89, Rak86, Sac88, SH85a, SK86, ST88, SEN82, TY82, Tus88, Wei80]. formative [GB87]. Formicidae [Jaf84]. forming [GBS89]. Forms [SNW81, BAB84, BRR87, Ji88, KSW^{+}85, WB89]. formulation [Kat85, MB86, Zno83]. forward [FM83]. fos [Mer87]. fossil [Mic88, Pea88a, Pea88b, Pea88c]. fossilized [Pea88d]. found [dW81b]. Foundation [Luc82, TLN85]. foundations [Ros80a]. founder [RB87]. foundresses [Wad85]. Four [McG89a, McG89b, Har87b, Hop86, JP89a, WH86]. four-parameter [JP89a]. Four-strand [McG89b]. four-stranded [Hop86]. fractal [Lef83, LQC89, SGH85, Tak89a]. fraction [ADdACP87, DBB80]. Fractional [Gre84b, Rak80]. fractionated [HMF88]. fracture [Ale84, KDG85, Wei80]. fragment [KK86a, PRH81, RTM81a]. fragments [BN80, GMTG80, KK81, RTM82]. frames [KM89]. frameshift [CM82, Cla82a, Cla82d, Cla82f]. framework [BG87, RSE81]. Fraser [Sta81].

Free [Blo87, EB87, Iwa88, MYS89, VDH82, Blo85, Cla82b, DV83, Fox81, Frö84, GKB86, HR85, Hun81, Jur84, Kor89, Orl81a, SR87b, SJ81, Wel85]. free-living [SR87b]. freeze [KDG85, Wei80]. freeze-fracture [KDG85]. frequencies [Cor83b, Fel82, GRJ81, Hin86, LN85, NS89a, RJ80, Tou85, Yok83]. Frequency [Cre88b, OH84, RMS81, BR87, Cre88a, Fer83, Gre84a, HS87, KM81, Les86, Mil87, MMG84, PWH^{+}88, Rat86, RR88a, Tre81a, VA85]. Frequency-dependent [Cre88b, Cre88a, Les86]. Fretwell [GKB86]. Frisch [Oth86]. Fritzsch [Met82]. frog [HM85b, PP84a]. Fröhlich [Pok82, PJK^{+}83]. fructose [KSH87, Tor80]. frugal [Sza83]. fruit [CB86b]. Full [Ber89]. fully [FGDS87, MH84]. fully-interchanging [FGDS87]. function [BB85, Bar88, Cha80a, Cha89c, CB86c, DKW89, FBF85, Fis80, Fuj82, HT80, Kin83, KP85, Kre82, Mad88, Mon88, SSW81, TLT80, Whi89b]. Functional [Bas81, ABEFS^{+}86, DD85a, DB88, Eil88, HvL86, KB89a, Mad88, MATGR87, Rei85a, RN82, Sch83a, Sin82, Smi82b]. Functionalism [HL84]. functionalist [LH88]. functionally [Vis87]. functioning [Fel80]. functions [Bas80b, CB88a, FDC^{+}89, GT89, LF86, NS82, Vla87]. fundamental [Coo84c, Ibe83, Ond80]. fungal [Ede82]. fungi [ES83, dSTSM83]. fungicide [HL86]. Further [MHT88, Mur87, NDB80, NMB85, Ros84b, GL84]. furylethylenes [BSR^{+}88]. fusion [LKC^{+}87]. futile [RS82].

G [Ano80j, BW87, EP86, HK80, MF85, SM81c, WR86]. G-T [MF85]. G-T/G-U [MF85]. G-U [MF85]. G. [Ano81i]. GA [GPLC84]. GABA [ABS82, BVDA87]. Gabor [TLN85]. Gabor-Nelson [TLN85]. Gain [Ish84, Mul85, Sar89]. galactose [VKV83, WMR86]. galactosides [WP84]. Galton [Gad81]. gambling [LC82]. game [Boy89, BP83, CB86a, HR80a, HM87, KLR89, MHI85, MHT88, Par83a, PPC82, Row85a, Row87, Row88, Shi87, TS86, Wil81b]. game-theoretical [BP83, PPC82]. Games [DC87b, Aok84, Cra80b, Cra89, FHL89, ICL85]. gamete [HJS84, Hoe84]. gametes [Hoe80]. gametophytes [Wil81c]. ganglia [Bar88]. gangliosides [WR86]. Gap [Alt89, Coo84b, BW86b]. Gas [NM81]. gases [Wal89a]. gaster [IAHA80]. gastropod [Hut89]. gate [BS89]. gated [BN80]. gating [BE81a, BE82, Cha88c, Cla83a, Gre89, Lev80b, LP83, MUI85, Str85c]. Gaussian [Tod85]. GC [Cla82f, Cla83b]. GC-rich [Cla82f, Cla83b]. gel [SR87a, SME86]. gels [McC84]. Gene [JJN86, Lum88, RB82a, WOT85, AM87, BW81, BHP82, Bur82b, BBH84, CF87a, Cla82d, Cla82e, CV89, CB88b, CB86c, Fel82, Fuk83, Fuk84, GDHB87, Hin86, JCZ84, KDW88, MAC^{+}83, OO84a, Sek85, Syv84, Syv85, TO89, VL86, Was86a]. Gene-culture [Lum88]. gene-for-gene [TO89]. genera [Pec88]. General [Ash85, Ber85, BGGV83, CG84, Car81a, Coo81, CB84, Fuj82, HBA89, JC80, KL87, Kon88, Las87, Les86, MM83c, Row88, SL81, SM88a, Sim86, Tay85b, dVR85, Lew81a, Pap80]. generalization [JP89a]. generalizations [ES88]. Generalized [Tay88b, Tay88c, Abu86, AOB85, Pel84]. generating [Bar81b, Bar82a, Gol88, Koc86b, Sir86]. Generation [Lan82b, MG80, SR80a, CBLM81, GTH88, HKK82, IBJ^{+}89, Kas80, Mat84, MM83d, MB88b, Sie81, Ste81, TH85]. generations [Cha80e, Gre84a, Rég84c]. genes [Bec80, CM82, Cla82a, Cra81b, Hau82, Iid88, KM89, Mac87, ND82, NDB83, NMB85, Por82, RT83b, San84, SJ85, SW81a, TH81, Yoc81a, Yok83, ZgQy84, dW81b]. Genet [Ing89]. Genetic [Abu84, Duf85, Duf86, FFM82, LM85, LJ88, Sou80, TS80, Abu88, AOB85, Aok84, BR83, BA80, Cha81, CFLF88, Cla82b, Dom83, Duf83, ER88, EGS80, Gra83a, GZ86, HT80, Hic89, Ji88, Kat81, Kuh82, Oht87, Rei84, RB82c, Row87, RW83, Shc88, Shc89a, Shc89b, Shc89c, Ste81, Suk89, SW83, Uma82, WB85, ZG85, dVR85]. Genetically [Get81]. Genetics [Abu85a, Abu86, Aal89, BSS83, Can87, Che84b, Nag82, New80, New83, RZ86, SS89a, Smi87, ZR86, dL83]. genic [Hje86, Rob85a]. genome [Bec80, BS87, Fil88, Mik83, SJ85, Stu86, TH81, Woj83]. genomic [KF87]. genons [BH81c]. Genotype [KK82b, Row88]. genotypes [GNG80]. genotypic [Nee89a]. genotyping [RB82a]. Geographical [Nag82, LA86]. geography [Fel82]. geologic [Pea88d]. geometer [Nin83]. Geometric [Hor89b, Bon85, Eco83, Eco84, Jea85, Pet83, Tod82]. Geometrical [HT89a, TH80, GPLC84, GB87, HYDS84]. Geometry [BG88, HE80a, BF83, CRB89, Car83, Cum89, HKC83, HCKB83, LS82, MZM89, Men85, Peu86]. Germination [Ell86, LA82, Tho86]. Geschwind [Jam89a]. gestation [CY84]. giant [AM82, Mat84, MUI85, dANG83]. Gibbs [Kre82, Wel85]. giemsa [Bas80c]. Gierer [Mac82]. gills [Lay87]. gingivitis [NMM84]. girase [ML86]. giucosamine [YBBL82b]. give [Bar83a]. given [Vis87]. gland [She80]. glands [JK89]. glauca [Hor89b]. gliding [Gor87]. Global [Fra82, Bot82, Hyv84, LCDC84, MF81, RTM81a, RTM81b, RTM82]. globin [Yok83]. globins [Lav84, SK84, Syv84]. globular [CH80, GRS82, PP80, Pra84]. globule [Kri80]. Glomerular [Luk84]. GLPC [NM81]. glu [NU82]. glucagon [KO83, She81]. gluconeogenesis [RP84]. gluconeogenic [Jam80]. glucose [FLNF86]. glucose-dehydrogenases [FLNF86]. glucosyl [WMR86]. glucosyl-ceramide [WMR86]. glutamyl [All85b, Osu84]. glutathione [MB82]. glycan [LBBL84, YLBB81, YBBL82c]. Glycocalyx [Ler83]. glycogen [KD88]. glycolysis [RP84, Tai86, Tor80, Woo80]. glycolytic [KSM85]. glycoprotein [VDH82]. glycosphingolipids [WMR86, WR86]. goes [FBB89]. golden [MK83]. Goldman [BB84]. Golgi [Bou84, Hol83b, OCMP89]. Gompertz [BP81a, GP82c, Hir82, Hol89, SV84b]. Gompertz- [Hir82]. Gompertzian [BGB89]. governing [DDA85, Hun81]. gradient [Bar82a, Lap80b]. gradient-controlled [Bar82a]. gradients [MBJK89, Wal89a]. graft [MH81]. gram [KD85, Koc88, Hun85b, Koc89]. gram-positive [KD85, Koc88, Koc89]. gramicidin [Mon83, SJ89]. grammar [CV89]. granulocyte [Vro83]. graph [Gol83a, GS85, Kat81, KL83, Per82]. graph-theoretic [Gol83a]. graph-theoretical [KL83, Per82]. Graphic [WIN80, BF81]. Graphical [KCM81, HC89, HM80]. graphs [Hyv80, Ola86]. Grasping [BA87]. grass [Bru85a, JP85]. grasshopper [HLS85]. gravity [Mor84]. grazing [CW88, JP85]. great [Mik84b]. grebes [YC89]. groove [Sto89b]. Group [SL80a, SD87, AOB85, Bro82b, GBS89, KK83c, LG89b, Pra84, Rob81, Rob85b]. groups [AM87, BR88, Cha89a, CB88b, Gri81b, Gri81c, Gri81d, Gri82b, Gri82c, Gri82d, Gri82e, Gri82a, Gri82f, KK83b, OK87b, Oli85, WJT87]. grow [BK83]. growing [AK84b, Gro88, IP83, Koc86a, MG84a, Rot83, SR87a, SE80]. Growth [ES83, HR84, KO82, KK83a, Lap84, Reg88, Sil89, AL81, Ada80, AC89, BB86, BM85, BRW89, BW86b, Ble88, Ble89, Bre81, BS84b, Bru85a, Cas81, CHT80, CK80, CN80, Ede82, FW81, Fer87, FDR83, FS85, GAD80, Gro88, Gun89, Har84b, Hej82, HR86, HG87, HCW89, Jea80, JP85, JP89a, KC81, KB88a, KF84, Koc82, KD85, Koc89, KWG87, LRKC83, LJTH85, Mac84, MB80, Mäk86, MB86, MW83b, Mit89, Mur83, Par83b, PK82, PK85, Pla82, Pok86, Pys80, RdL83, Ric87, Ros84a, Ros85, Rot82a, Sab80, SSF80, SAS85, SH82a, Sch81, STN^{+}84, SCN85, SDM^{+}82, SM80a, Str80, SG89b, Tod85, TW80a, TS88b, Val85, Voi85, Wax81, Wes83a, Ble88, Ble89, RdL83]. GT [GPLC84]. GTP [Jak88]. guarded [SH85b]. guarding [GR83, Yam87]. gull [DC87b]. gynodioecy [vDvD86]. gyrase [For80]. gyration [Fis85].

H [NK88]. H1 [Van82, JS81]. H5 [Van82]. Haber [OC81]. Habitat [BR86]. Habituation [Jac82]. haemoglobins [SK85b]. hair [MN85, NM82, NM85]. hairpins [NST80]. Halecostomi [Mul87]. Half [Sch80]. hallucinogenic [KM85]. Hamilton [Kir86, Nee89b]. hand [BA87]. handed [LL87c]. Handedness [Jam86a, Jam88, Jam89a]. handicap [Dom83, Mot82, NH84]. handicap-phenotypic [Mot82]. Handicaps [Smi85b]. handle [Gar82]. Hanes [FLR83]. haplo [Wad82]. haplo-diploid [Wad82]. haplodiploid [Iwa81]. haploid [Asm83, Gri82f, Lew81b, Lum84, Tre82]. Haploids [Moo88]. hard [Wes83b]. hardiness [Sto81]. harmful [Hel80]. Harpending [GT88a]. harvested [Rot87]. harvesting [BS82, FBF85]. having [BESD84, MF81, Mot88, Tay88c]. Hayflick [Bre82]. haystack [BT80, vTdJ86]. head [BHH83, Gra80, Mac82, Ser87]. head-tail [Ser87]. head-theoretical [Gra80]. healed [Pic89]. healing [WFL86]. hearing [Bar81c]. heart [AFHM86, BF88, DKW89, GK85, HBRN81, KYU^{+}86, MR86]. Heat [BW85, For85, OL86, dCB84, CFLF88, Gra80, dANG83, SA81]. heat-shock [CFLF88]. heavy [Wat85]. height [CF87b, ICL85, Mäk86]. helical [Ale87, BHM^{+}85, FY81, LN85, Men85, Nus85, Sil89]. helices [PR82, SK84]. helicoidal [Nev88]. helix [AL83, BD80, Cla83a, Men82, VM82]. helix-helix [BD80]. helminth [BKMS86, BKMS87]. help [Mot88]. helper [dBH86]. helping [BP85]. heltical [BD80]. hematocrit [Boy88]. hematogenous [WD86]. hematologic [CCDF81]. heme [DO83, GG84b]. hemicellulose [Nev88]. Hemodynamic [Li83]. hemoglobin [Bla84, CBF87, Dun82, Gro80, GG84b, Sve82]. hemoglobin-ligand [Gro80]. hemoglobins [FJ84, KN83]. hemolytic [AWV^{+}88]. hemoproteins [HT87, KN82, KN83]. Henderson [Shi80]. hepatic [BRR87]. hepatitis [Bec80, EPH86]. hepatomas [Pet83]. hepatotoxic [WS81]. herbicide [Kno82]. herbicides [Shi81]. herbivores [MH80d]. herbivory [Tro84]. herbs [Pug87a]. hermaphrodite [HvL86]. hermaphrodites [Cha89c, GC86]. hermaphroditism [BB81c]. herpes [RMFM85]. hessians [BWM80, BWW80]. heterogeneity [ACB81]. heterogeneous [Bas80a, BC80, Car81a, CB81, GS81, ME82, Pop82, Ros85, SGH85, WKV88]. heterologous [VL86]. heterospory [HW88]. heterostylous [CC82a]. heterothallism [dSTSM83]. heterozygosity [Has88, dCJM86]. heuristic [TACM82]. hexagonal [AL81, HL86]. hexaploids [CG80]. hidden [DR89]. hierarchical [AW83, Eil88, HU85, Pap84]. Hierarchies [Peu83, Peu85, Peu86, TLT80]. hierarchy [DDJ83, PL87, SH87]. High [Rat86, Fel80, KH80, PP85, Tre81b, VCC^{+}88]. high- [Fel80]. higher [BS87, EP86, Hir84b, Iid88, SBM89]. highly [Inf85]. hiking [Man82a]. Hill [BWM80, BE86, Bri84]. histidine [Dug82]. histocompatibility [Fis80, SM88b, Tre81b]. histograms [BGGV83, KSW^{+}85]. histone [BN81, Bro83, JS81, KDW88]. histones [Van82]. Historical [Lau82, TIK89b]. histories [Cas82, Cas84, ES88]. history [Abu85b, Abu86, Abu88, Fel82, IT80, PC83, SC87a]. Hitch [Man82a]. Hitch-hiking [Man82a]. hitchhiking [Orl82]. HIV [KM89, LG89b]. Hl [BN81]. HLA [WD87]. HnRNA [Sek85, JJ82b]. hoc [Har82b]. Hodgkin [BB84, AMI84, BBS88, Cha80b, HBRN81, HS89, LP82, MS89a, MUI85, RSM80, Str85a, Str85b, Str85c]. Hofstee [FLR83]. holoblastic [GL84]. home [Ben89]. homeostasis [Sco84a]. homeotherms [KB86]. homing [PW88, Wal89b, Wal89a]. hominid [Cal83c]. Homo [AS87]. homochirality [Tra86]. Homoeotic [Sla85]. homogeneous [Web81, WKV88, FDR83]. homologies [FLNF86]. homologous [Hop86, SCP83, TS80, VZR85]. homologue [Mag84]. Homology [KTNO81, vHU87, Fin81, Lav84, Nus80b, Nus83, YI88]. Homomorphism [Str85a]. homopolypeptide [VA88a]. homoptera [Ton86]. homosexuality [Jam89a]. homothallic [JCZ84]. Honderich [Lib85]. honest [vRV80]. honey [Ros80a, Ros80b, Ros84b]. honey-bee [Ros80a, Ros80b]. honeybee [OL86, Omh87a, Omh87b, Omh88a, Omh88b]. honeybees [Omh87b]. Hopf [CLFL86]. horizontal [FHL89]. Hormonal [Jam86b, Kha86, Sch82b, Was86b]. Hormone [Blu84, Jam87, AWV^{+}88, BRW89, Bec88, DO83, Jam85, Jam89b, Jam89c, Ros84c, She80, She81, SH83b]. hormone-receptor [Bec88]. hormones [SH80a]. hornets [IPS83]. Horton [Par85]. hospitalized [MSBB84]. Host [Kno89, ADG80, And80a, Bre80a, BF85, Cze87, JK88, KM83, Lew81a, Lew81b]. host-parasite [KM83]. host-parasitoid [ADG80]. host-pathogen [Bre80a, BF85]. hosts [Kel87, Lew81b]. hot [Cla82a]. hot-spots [Cla82a]. hourglass [SL87b]. household [ASTT80]. hovering [KSL89]. hr [DCR82]. HTdR [WF82]. HTLV [PH87a]. HTLV-III [PH87a]. HTLV-III/LAV [PH87a]. huddling [CRB89]. Human [Joh86, Mik83, Aok89, BRW89, BIS^{+}80, BW85, GDHB87, GTH88, Gra80, Jac83, JP89b, JP89c, KK83a, KFSW82, KF84, Kwo84, Lav84, MP87, NMM84, PH87a, PCTF83, PJK^{+}83, SPP84, Sax83, SN80, SK85b, SME86, Sve82, Tus88, Wu88, Yok84, Yok85, dGV80]. humans [MH81]. hummingbirds [Cor83b]. humoral [Bec81, KT87a]. Huxley [AMI84, BBS88, Cha80b, HBRN81, HS89, LP82, MS89a, MUI85, RSM80, Str85a, Str85b, Str85c]. Huxley-Type [HBRN81]. hyaluronic [WFL86]. hybrid [PWH^{+}88]. hybridizing [Can87]. hydra [Mac82]. hydrated [BLR83]. hydrates [EG87b, Goo84]. Hydration [GVP^{+}80, VGP^{+}80, LKC^{+}87]. hydrocarbon [IO83, Sýk84]. hydrocarbons [MHB83]. Hydrodynamic [GBFH88, Sil83]. hydrodynamical [MOV82]. hydrogen [HG80, HFO83, Jam82, NSKS85]. hydrolysis [GBS89, Kas80]. hydropathic [MATGR87, MAGR87]. hydropathy [Bre88]. hydrophobic [Frö84, KB83]. Hydrophobicity [Ven84, CC82b, LdJ86]. hydroponic [Ros84a]. hydrostatic [WB89]. hydroxyoctadecanoyl [WMR86]. hymenoptera [Abu81, AM81b, Bul81, Cra80b, Cra82b, Get82, Orl82, Sta85, WL87, Jaf84]. hyperbolic [EC81, IHK80, IHTN88]. hypercycle [GTSNO^{+}88, MH83, MH84]. Hypercycles [BNH80, EGS80]. Hypercyclic [EGS80, GTMM87, Ish84, Ish86, Sza88]. hyperplasia [All85a]. hypothalamic [Wee84b]. Hypotheses [BW86b, Fra88, Jam89a, Mal84c, Tre81b]. hypothesis [Ber89, Bog87, Bol82a, Bou84, Dou85, FH89, Gra81a, HLS85, Inf85, KG82, KRB81, Kir86, Klä83, LGK84, Loc82, Lum88, LPS86, Mal81b, Mal84b, Mal84a, Mal85, Mat81, Mon88, NMM84, NP81, NP82, Nic84, OH85, Ond80, Rap84, RH80, Ros80a, Ros80b, Ros84b, Ser87, Sla86, SMGD80, Tak89b, Tho85, TO89, Ula80a, Ula80b, Ury81, Wee84a, WGHY82, Zaj82, ZgQy84, vZ87, Uma82]. Hypothetical [Cla82b, Dou84, Hir80, Kel87, Sýk84]. hypothyroidism [MP87]. hypoxia [LR85b]. hysteretic [JSJS88, KK82a].

ice [CSL^{+}86]. IDDM [Wee84b]. Ideal [Kor89, GKB86, Nus85]. ideality [Kat85]. identical [AM87]. identification [AK82]. identifying [Ols87]. identity [HK83]. Idiotypic [dBH89]. if [Sta85]. IgG [GH82]. IgM [GH82]. ignored [Zno83]. II [Abu85a, AK87b, Bar82b, Ble89, BSB88b, BKB80b, CV88, CMM86, Che83, CPPRS87, CTS86, DD85b, EM81, FT87, Gal84b, GVC81a, GT88b, Gre82b, GZ86, Gri81b, HCKB83, Hoe84, HFO83, JS82, JP89c, KK87b, KN82, KH80, KK81, KK80, LL84, Lew81b, LSFT85, Mah84b, MW83a, Mas84, MTM86b, PJL84, Pap82, PGKE88, Peu85, RdL83, RTM81b, RN84b, RR83b, Sar87, SH81, SL87a, She81, Shi81, Sim80b, SM80b, SF83, VGP^{+}80, WB85, WSMP83, WGRZ80, iWSS88, YBBL82a, Yod84, ZBL80b]. II-Interactions [JP89c]. III [Abu86, BJ84, Che84a, CG81a, DW87, GVC81b, GRS82, Gri81c, HdL84, KKNM80, KN83, Kon88, PPL85, PH87a, Peu86, RTM82, SL87b, SLFT86, SKF83, SME86, YBBL82b]. III/LAV [PH87a]. Ill [Top88, Maw85]. Ill-conditioning [Top88]. ill-effects [Maw85]. illusion [Eco84, Sza88]. illustrative [NMM87]. Image [DL85, HT83, RM86]. image/pressure [RM86]. imaginal [MRB83]. imitate [Rak86]. Immigration [CW84]. Immigration-extinction [CW84]. immobilization [BE82]. immobilized [CP85, LR80, SR87a]. immune [Bec81, Dan86, DH85, FE85, GB80, GH82, Hof86, Hof87, KW89, KUT85, KT87a, MAC89, Nes84, SLN84, Vro83, WC88]. immunity [BKMS86, BKMS87, SSF87, Tre81b]. immunization [FE85]. immunoassay [SN85]. immunoassays [SSN86]. immunogenetics [MW83a, WM83, WD87]. immunogenicity [SLN84]. immunoglobulin [CSH^{+}81, Mor85b, SW81b]. Immunological [dBH87, Fai85, For85]. immunology [GGC86]. impact [WD87]. Impaired [Inf85]. impedance [TLS83]. Implementation [GS86]. implicating [Rao82]. Implication [ND82, Fuj82]. Implications [BHM^{+}85, Can87, Car83, GR88, HB86b, Kue81, Mäk86, MM82b, NSW82, PJK^{+}83, RT85, SS89a, Sla85, SR87b, AS87, Bas83, Bla80, Blu83, Bre88, Cal83c, Cla82a, Coo82, Cor82, EKSS88, Gor87, Gre80, LS81, NMB85, PGKE88, Sau80, SK85b, Syv85, TS80, GK80, MN89b]. Importance [Wei83, And86, Hyv80, Ste81]. important [Esh84a, VTL85, Wis80]. imposed [Kri85b]. imprecision [DS84]. Improved [Law84, BAB84, FS80, ZR85]. Improving [Bus82]. impulse [SM81a]. in-series [For81]. inactivating [Ste88]. inactivation [BE82, Cas81, FK83, Fri85, GVC81b, Top88]. inappropriate [Con83]. inbred [Gri81c, Gri81d]. inbreeding [Cra82a]. incidence [CCDF81, DCF80, HCW80, KF84, MW83a, PC83, YE82]. incipient [Gen88]. include [McC84]. including [DLN85, GBS89, PR82, Rap84, Sch82b]. inclusion [Bro83]. inclusive [AM81a, Tay88a]. incompatibility [Pen82]. incompatible [vdHN85, vdH84]. incomplete [BBS88, Rog83]. incompleteness [Pen82]. incorporated [KBR88]. incorporating [Asm83, Gri81a, Gri81b, Gri81c, Gri81d, Gri82b, Gri82c, Gri82d, Gri82e, Gri82a, Gri82f, SHBS89, TH85, dBH89]. incorporation [Law84]. Increase [RKL83, Coo84b, Dem81, Hir80, Hir82, Rob89, SH81]. increases [Iwa88]. increasing [Lib88]. Indefinite [NSJW84, NW85]. indefinitely [MJN88]. Independent [Reg88, Cal87, GAD80, Hol85, Par83a, Rei86b, SH83b, SBPF83]. Index [Ano84h, Ano84i, Ano84j, Ano84k, Ano84l, Ano85h, Ano85i, Ano85j, Ano85k, Ano85l, Ano85m, Ano86h, Ano86i, Ano86j, Ano86k, Ano86l, Ano86m, Ano87j, Ano87k, Ano87l, Ano87m, Ano87n, Ano87o, Ano88n, Ano88o, Ano88p, Ano88q, Ano88r, Ano89g, Ano89h, Ano89i, Ano89j, Ano89k, Ano89l]. indicates [MKL81]. Indications [Leu87a, Nus82a]. indicators [Bas82, Sta81]. Individual [BB81d, CK82, Di 88, Di 89, GWK87, HH89, Kar88a, LS89, Sar87, vdH84]. individual-site [Di 88]. Individuals [BB81a, Gri81a, Gri81b, Gri81c, Gri81d, Gri82b, Gri82c, Gri82d, Gri82e, Gri82a, Hau82, LG89b]. individuate [Gri82f]. indo [SLFT86, YBBL82b]. indo-ISCRF [SLFT86]. induced [BM85, Eva80b, FS85, HC81a, HL86, HT89b, KM81, Kur83, Rab82, SK87, Tay85a, Tho85, Toz89, VA88a, ZG82]. induction [Aki84, JR80, SM88b, Sve81, Tay89a, dGV80]. inert [BW89]. Inevitable [CB87b, Kac87, Orl82]. inevitably [TIK89b]. Infanticide [HAC82]. infected [LG89b, SV89, Süh89]. infection [BKMS86, BKMS87, GW88, Pok86]. infections [SSF80]. infections-Mathematical [SSF80]. infectious [Ewa80]. infer [Fel82]. inferring [DS87]. infinite [Gre82b]. inflammatory [WFL86]. inflow [Rob88, Rob89]. Influence [BW85, CH82, GTMM87, GTSNO^{+}88, Hir80, Hir83, Hir84a, ME82, WF82, ASTT80, Cla82c, Duf86, Her83a, KB83, Par83b, TATS81, Wad85]. influenced [BSR^{+}88]. Influences [Cla82d, Hua80a]. Influenza [HHP89]. Information [BLC84, HW82, HU85, Mag82, RH85, Rya80, SV84a, TG80, TG88b, AK82, Alf83, Alm85, ALST81, CH85, Coo84a, CB83, EGS80, Gar82, HT80, Ibe83, Ish84, Ji88, Kir80, Kon84b, Lam84b, Mat83, MH87, Oht87, PP80, RGR82, Sar89, SBM89, SW81a, Smi82a, SM88b, Stu85b, Tan80, Tan84, Wal89b, WB82, Wil86b, Yoc81a, Yoc81b]. Information-entropy [Rya80]. information-theory [Alf83]. Informational [GPPZ80, Kre82, LTZ88, Kuh82, RT83a, Stu85a]. Informativeness [Sto84a]. infusion [Fai89, FS86]. inherent [PV83]. Inheritance [Hed85, Col82, CT81b, Hje86, JL89, Lum84, Tre81a, YE82]. inhibited [Ber83]. inhibiting [Ond80]. inhibition [Aki84, BPM82, BC81b, Bro80b, Kar88b, Mil87, Ped88, Pin83a, Pin83b, Pin84, PJK^{+}83, Rak80, SPS83, Shu84]. inhibitor [LH86, NSKS85, SC84a, Top88]. inhibitors [HAG82, JF85, KB89b, Rog84]. inhibitory [AL87, Omh88a]. inhomogeneous [Bro81a]. initial [Hsu88]. initiated [CSL^{+}86]. Initiation [MRG84, UK82, Bah80, BT87, MZ89, Man81a, NM85, Nag86, Nus86, SCH84b]. initiator [Eps86]. Initio [ANOM82, Aid88, BLR83, BVDA87, NA84, NU82, PRH81, PR82, TIK89a]. injection [PM80]. inner [Esh84b]. input [Fin82, KSH87, Poz87]. inputs [GC86, TW80b]. insect [COA85, DK82a, DK82b, KM83, LS87, MM83b, SL87a, SL87b, ST80]. insects [Bul83, Cor83b, Cri81, DK83, Iwa81, KR85a, Rég84c]. inseminations [Wad82, Wad85]. insensible [JK89]. insertion [Kyp88]. insertions [SK84]. Inside-to-outside [KD85]. insight [BG86, PL87]. insights [GP82b]. instabilities [AM82, BKJT80]. Instability [Eva80b, Bar80b, Cus80, GZ86, KF87, LO82, MM82b, Mur82, Pow88, ST86a, Win89a]. Instantaneous [TLS83, FW81, JT84]. instructive [Was82b]. insulating [TLN85]. insulation [KB86]. insulin [BRRK87, Sab80, Wee84b]. insulin-dependent [Wee84b]. intact [Bas83, Nor82]. intake [FBP87]. intakes [LS89]. Integer [LK83]. Integer-ratio [LK83]. integrated [SPS83, SC83, SP84]. integration [BH88, Gra85b, Kur86]. integrity [Sco84b]. intensities [CN84, Her81b]. intensity [DBB80, MM83a]. intentions [Car82b, Moy82b, Smi82a, vRV80]. inter [Aid88, GRS82]. Inter-residue [GRS82]. inter-strand [Aid88]. interacting [Aug85, DV83, GNG80, KVW88, LF86, SV84b, ZR84, ZR85]. Interaction [BESD84, IHTN88, SRO85, TS88b, Aid88, BRW89, BLR83, Blu84, DML80, FBS85, GPLC84, Gri81a, Gri81b, Gri81c, Gri81d, Gri82b, Gri82c, Gri82d, Gri82e, Gri82a, Gri82f, IYS83, KK81, Ler83, McA87, NA84, NSW82, PC86, Rei83a, RSM80, Sch84a, SD81, SS89b, VDH82, WWG^{+}88, Hoe84]. Interactions [Bro81b, JP89c, SNW81, dBH86, Ano81-37, Arg81, BP81b, Bre80a, Bry80, BD80, CG85a, CPLA85, CPPR87, CPPRS87, Fel83, FS83, Goo84, GPM89, HB88, Hol85, HR85, MMGL87, Mon83, NMM87, NSJT81, RN84c, Sar87, SC89, Ten80, VGP^{+}80, Vro83, WW84, Wil82]. Interactive [BF81, Cum85b]. intercalating [Sch82a, Sto83]. intercalative [Sto89b]. intercellular [BPP88, SG82, Wei80]. interception [COA85, Her83a, Her83b]. interchange [Day83b]. interchanging [FGDS87]. intercolony [LH83]. interconnected [Ren82]. interconversion [AK87a, DLN85, SHBS89]. intercostal [RSSK87]. interdependence [CN84]. interfaces [Rya80]. interfacial [Fer85b, KO82]. interference [FW85, Jol88, Mag88]. interferon [Sco82, Mar80a]. intergroup [GT88a, HR87]. interiors [Duf86]. intermediate [GBS89, Hop86]. intermediates [Shi82]. intermitotic [vZvdSdL81]. intermittent [HN85]. intermolecular [Fer84, SR80b]. internal [DSD81, Koc88, Lap81a]. internalization [Bec88]. interphase [Suc85a]. interplay [WWN82]. Interpretation [SS85, BG84, Bol82b, Cha82a, GM82a, IP83, KM81, KBR88, LP82, MUC81, Pet80, Rak80, Sib83, VD89, Whi81a, SGH85]. Intersexual [Car85, Jac82]. interspecific [Moy82a]. Interspersed [Bro82a, GMTG80]. interspike [TC80b, WR83]. interstitial [INMM83]. intertidal [FDC85]. Interval [AWMK85, WR83]. Interval-strength [AWMK85]. intervals [Vau86]. Intestinal [YMIH86, AHFH81, EAL80, HF85, MHII83, NP81, NP82, Pys80, Swa82, Zaj82]. intestine [BWM82, JA88, PL87, Win89b]. Intra [LL87b, Aid88, GTH88]. intra- [Aid88]. Intra-specific [LL87b]. intra-uterine [GTH88]. Intracellular [Bol82a, Whe84, For85, Kir80, KT81, LL88, SJ81, WI80]. intraclonal [SS88]. intracolonial [Omh88b]. intragametic [Esh84a]. Intragenic [LvH87, Bul88, Cla82c]. intragenomic [CT81b]. Intramembrane [SM81a, HL86]. intramembranous [Wei80]. intramuscular [GM82a]. intraspecific [Asm83, Ebe87, TC80a]. intravascular [INMM83]. Intrinsic [MHB83, BESD84, GP82c, JA88, dBH86]. intron [Nus88]. Introns [HB86b, Con83, CL84, For81, HBA89, Hic89, Lav84]. intruder [LE84a]. Invalidity [Shi80]. invariance [Nag82, PV83, vS83]. invariants [Bas83]. invasion [HE84, Has86, SW84]. inverse [Nus82a, Pel84]. inversion [DS87, RN85, WEHM82, Wis80]. invertebrates [FQ81, SC82]. invest [Rei85b]. Investigation [Sim85, BD80, HG87]. investigations [DE82]. investment [Blu83]. involve [Ola82]. involved [Flo84, Hje86]. Involvement [Wri80, Bee80, HW80a, Wee84b]. involves [Inc83b]. involving [BRW89, EN86, FEC87b, GDHB87, Lac84, ND82, WT87]. ion [BF88, Cla83a, DLN85, DT80, EG87a, Fer80, Fer81b, Fer81a, Fer83, FS83, Fri83, HG80, KS80a, KSS81, KK87a, Leu87b, MOK87, MH80c, NA84, NU82, OK87a, SS85, Sta86, SF83, SKF83, Str85a, TMGB80]. ion-penetrable [MOK87]. ion-protein [KS80a]. Ionic [Mon83, BN81, Cha80c, KFPB85, Lev80b, NU82, Ric87, RR88b, Ste87, SME86, TIK89a, VA88a, You85]. ionizable [Cha89a, OK87b]. ionization [Sar87]. ionizing [Lam88]. ionotropic [ST86b]. ions [CS84, EG87b, Hal80b, JD82, KB83, Rab82, SJ81, Ten80, ZG89]. irradiated [Cra80a, LP86, MdMF83]. irradiation [Szu83]. irreversible [AK87b, HC87, LR88, Peu83, Peu85, Rak80, Top88, VI80, Wag82, WT87]. ISCRF [SLFT86]. Ising [TLT80, YMS84]. islands [CW84]. Isobolic [Sto89c]. isocline [Nun80]. isogamous [Hoe82, Wie81]. isolate [Gla87]. Isolated [HS89, SR81]. isolating [Rob85a]. isolation [Mac87]. isomerism [Rak86]. isomers [WS81]. isomorphic [Nag89]. Isomorphism [Str85b, Str85c]. isosmotic [Mas84]. isothermal [Web81]. isotonic [FM88]. isotope [KBR88, SR80b, SR89]. isotopes [McC80a]. Isotopic [AK87a]. isotropic [Sar87]. isotropy [Fer82]. Isoxazolol [BVDA87]. isozyme [GFA89]. isozymes [LG89a]. itself [Mot89]. IV [GVCC82]. IV. [BD80, CG81b, Gri81d, Kot82, PL85a, YBBL82c]. IX. [Gri82a].

J [Ano80j]. J. [Ano80j, Phi82]. January [Ano80k, Ano80q, Ano81o, Ano81u, Ano82n, Ano82t, Ano83p, Ano83u, Ano84r, Ano84-36, Ano85r, Ano85x, Ano86p, Ano86v, Ano87q, Ano87w, Ano88s, Ano88z, Ano89m, Ano89t]. jet [MB88a]. jog [DW82]. joining [WJT87]. joint [Koo85]. jointly [Jam89a]. July [Ano80l, Ano80r, Ano81p, Ano81v, Ano82j, Ano82p, Ano83m, Ano83s, Ano84n, Ano84t, Ano85o, Ano85t, Ano86r, Ano86x, Ano87r, Ano87x, Ano88u, Ano88-27, Ano89p, Ano89w]. juncos [PPC82]. junction [BBH84, RMFM85, RCA89, Ser87, Vau86, VM89, WH83]. junctions [BW86b, Coo84b, DSR89, Nus88, See82, TGP^{+}86, Wei80, WGP^{+}88, WWG^{+}88]. June [Ano80z, Ano80-32, Ano81x, Ano81-29, Ano82u, Ano82z, Ano83z, Ano83-29, Ano84-27, Ano84-32, Ano85-29, Ano85-35, Ano86-31, Ano86-36, Ano87-31, Ano87-38, Ano88-31, Ano88-38, Ano89-27, Ano89-33]. junk [Gra81a]. juvenile [Kha86, McN89].

K-ATPase [Wig82b]. Kananyan [Ano81i]. karyotype [MI81]. Katz [BB84]. Kauffman [Sta88]. Kaufman [Hof87]. Kaufman-Urbain-Thomas [Hof87]. Kawato [KP85]. kernel [AFHM86]. keyed [Oli85]. Keywords [LH88]. Kh [Ano81i]. killing [WD86]. Kin [CC81, Hug86, Que83, SH82a, AM81a, Cra82a, Cra83b, Get81, Get82, Rot80, TACM82, Wal87]. kinase [DDA85, GFA89, IHK80, Kre82, Tor80]. kinase/fructose [KSH87]. kindling [AdMM86]. kinds [FJ84, Hao82, Moy82b]. kinematic [Zaj86]. kineses [DD85b]. kinesis [DW87]. Kinetic [BBS88, Cha88c, Der87, Gho82, GBS89, HFL89, HT88, Ped88, Pet82, RTM88, TIK89a, Yos86, AK82, BS81, Bar83a, BBM87, BE81b, Boy83, Bro80b, Can87, Cha82a, CFLF88, CTS86, CA89, Dam86, Dug88, EC81, FGDS87, FCBK88, Fri85, Gai80, GC83, HH85a, HCW89, HT89b, Inc83b, IHK80, JP89b, KP89a, KK82a, KCM81, LR80, RN85, RN86, SL81, SC89, Top88, WBKW82, WD81, Whi89a, YMS84]. Kinetics [BSR^{+}88, HW87, KDW88, Rak80, Rak86, SPS83, SEN82, VCC^{+}88, VKV83, WT87, Agm85, BRRK87, Bas83, BW89, BC80, BWM82, BK81, Bun80, Cas81, Cha80b, DLN85, DRB86, FTDS85, GV81, GPKBG89a, Gre80, HBH83, Kar88a, KP89b, KS89, KK82a, Koc82, KKJ81, KD80, KC83b, Lap84, LA82, LQC89, LPS89, ME82, MH80a, MUI85, MB88a, PL84a, PL84b, SH80b, TCR85, WGKG83, Wei85, ZG82, vZvdSdL81, GVC81b, TD81]. King [Ola86]. Kingsolver [Ano80j]. kink [DW82]. kinks [McG89b]. Kinship [Seg81]. kittiwake [DC87b]. klinokinesis [Dus89a]. klinotactic [Bor84]. know [Was86b]. Knowledge [HL89]. Knowledge-based [HL89]. known [Wu88]. Koshland [Whi80c]. Krebs [Der84]. Krebs-cycle [Der84]. Krogh [SA82].

L [Ano80j, Por83, RMFM85]. L-cell [Por83]. L/S [RMFM85]. L1 [Ber89]. label [KBR88]. labeled [LS81]. labeling [VA85, VA88b]. labelled [HMF88, Rei81, Smi82c, Spe87, SR87b, SR89]. labelled-ligand [Smi82c]. labelling [KBR88, PM80]. laboratory [SV89, Süh89]. lac [CM82]. lack [Der84]. lactalbumins [MATGR87]. lactam [HSd^{+}88]. lactate [Coh85, Jam80, Jam82]. lactose [Man81a]. Ladder [Bro81a]. laevis [KDW88]. lag [CN80]. Lake [Car83, Har82a]. lake-water [Har82a]. lambda [Ano81i, KRT81a, KRT81b]. Laminar [Li88, EAL80, MHII83, YMIH86]. land [Eco81, Hok86b]. landscapes [KL87, KW89, SC83, Wei88]. language [Cal83c, Hux86, Ibe83, Ros80a, Ros80b, Ros84b, SC87b, WM83]. langurs [HAC82]. Laplace [Li86]. Large [Nus85, CN85, Gar82, LH80, NST80, NP84a, NP84b, TCR85, WR83]. larger [Rei85b, Rei86c]. largest [Eco81, Hok86b]. larvae [RH85]. latent [HT83]. later [Bro83]. lateral [FBS85, Pin83a, Pin83b, Pin84]. latter [KR82]. lattice [LFP84, McK83, Mor85a, Sta88]. lattices [See82]. launch [Cal83c]. LAV [PH87a]. Law [Cri82, Els83, GKB86, Gra81b, Hir82, Li86, MLS80, Oon82, Pah85a, Par85, SGH85, Wag82]. laws [Cho81, Voi85]. layer [Col85, Gai80, GS80, HF88, HSd^{+}88, JA88, OK87b, VDH82]. layers [TG88b]. Leader [MDPF86, KvDVH87]. leading [Aki84]. leaf [Tak89c]. leakage [WWG^{+}88]. leaky [FM88, TGP^{+}86]. learn [Har83]. learned [Get82]. Learning [Cra89, Har81, Har87a, HS87, Reg86, Hou83, MH85, Sal84]. least [Bro80b]. leaves [Her83a, Her83b]. lecithin [SG89a]. led [OK87c]. left [LL87c]. left-handed [LL87c]. leg [MM83d]. legume [SS89b]. length [Ber89, CB83, Fra84, Hor80, NBZ^{+}83, Par85, Rob88, WS85a]. lengths [WH83]. lens [Hor89b]. lesions [Kyp88]. less [KR82, Rei85b]. Lessons [HT83]. lethal [YE82]. lethality [Wol85]. Letter [McC80c]. leukocyte [MSBB84, PJK^{+}83]. leukocytes [PJK^{+}83]. levation [Mul87]. level [Ber87, Bro83, GPKBG89a, Gla80, GM84, Nun80, SP88, Sch82b]. levels [BESD84, CA89, Jam85, Jam87, Jam88, Jam89a, Jam89b, Jam89c, MKS^{+}86, MP87, Rya80, Tre82]. Lewi [vDvD86]. lexotypes [Zuc86]. Li [NA84]. liable [Ale84]. Libet [Hon86, Mas86]. libitum [FBP87]. Lichen [CK80, KC81]. lidear [ML82]. life [Abu85b, Abu86, Abu88, And80a, BS85, Bar81a, Cas82, Cas84, Dix83, Dou81, ES88, HE84, IT80, Nic84, Nod82, Org86, PC83, RB82e, SC83, SC84b, SC87a, SD87, Yoc81b, ZL84]. life-cycle [SC83]. life-history [SC87a]. life-proposed [Dou81]. lifespan [Cor82, Kue81]. lifetime [Lei81]. lift [Orl81a]. ligament [Fer83]. Ligand [Sta86, SPG87, BM87, Bri84, Bri85, Bri88, Gro80, HM80, IHTN88, KN82, KD80, Kur83, Kur84b, LL88, MJN88, MM86a, NSW82, NW85, SNW81, Smi82c, WSMP83, Whi80b, Whi80c, Whi83, WWN82]. ligand-binding [Whi80c]. ligand-induced [Kur83]. ligands [BO82, DBB80, Hum81, Inc83a, NU82, PC86, RN84a, Sch82a, SEN82]. light [Cha80a, FBF85, Fis85, HN85, MdMF83, MT81, MPPF80, PW88, Sch83a, Tak89c, Whi82, ZG82, vGvGL88]. light-driven [Sch83a]. light-harvesting [FBF85]. lignification [Hwa85]. like [Ber89, DG80, MVV89, WB89]. likelihood [LR85a]. likely [Kac87]. limb [Oth86, RL84, TT85, TT87]. limbs [Jac83]. limit [Bre82, Hao82, Hyv84, KK86b, Lap81b, MS80, NL86, PS80b, Pet80, Pug87b]. limit-cycle [Hao82, MS80]. limitations [Hol88, KK80, SR87a]. limited [GF81b, Gro88, PTD81, Ros85, SN85]. limiting [WP84]. Limits [Ågr85, Dus88, Mei82, ÅF80, Phi81a, Phi82, Sch81, Tre81a]. line [BT81, SS88]. Linear [Han85, SC89, Abu81, BM87, BMR89, Bel80, Bro81a, Bro80b, CN84, DKW89, Dug82, Hol85, Hsu87b, Jam82, Kuh82, LK83, Mae84, Mal88, McC84, MH84, New80, New83, NSJW84, OC81, PS83, Peu85, Pin83a, SSM81, SM81b, Sto89c, Tay88b, TT87, Mah84b, PPL85, Peu83]. linearly [DY82, Kur84b]. linearly-varying [DY82]. lines [Cha89b, Fuj82, MKM80, Pet83]. Lineweaver [FLR83, GC83]. lining [Tro83]. link [CL82, Hua80b, MW83b]. Linkage [Di 88, Di 89, Fel83, Hin87]. linked [BE82, Gho82, Mah84a, Mah84b, YE82]. linking [Dea80, Hua80a, NSJT81, Was86b, WI80]. lipid [BP81b, Hal80b, KM84, LC80, SA86]. lipid-protein [Hal80b]. lipophilic [SA86]. lipophilic-eutectic [SA86]. lipoprotein [JJ80, PP85]. liposome [TL88]. Liquid [RCA89]. Liver [Jam80, ACB81, Bas81, Bas83, IHK80, JK81, KD88, She80, Tor80, Ury81]. livers [VKV83]. living [BS84b, FDR83, Hau82, Ji85, Kha86, Kre84, RdL83, RN84a, RN84b, SR87b, Tho82]. load [BB88a, Mad88, MM85]. Local [Hyv84, Mal84c, PL85a, SE80, Sil84, SHBS89, TB80, Tay89b, FJ85, LKC^{+}87, Mah84b, MF81, Rob89]. localisation [PRH81]. localized [Com82]. localizes [NK88]. locally [BW85, Sto89a]. located [Mur87]. location [Bas80b, DB88, Frö88]. locations [Sho85]. loci [Has81, Tre81b]. locomotion [Dob86]. locus [AM81b, CB86c, Cre88b, Has88, Hin87, KK86b, LP88a, Orl81a, SFR86, VC88a]. locus-two [KK86b]. log [Smi82c, Wei85]. log-convex [Wei85]. logi [SBPF83]. Logic [AK82, MH81]. Logical [Kre84, KUT85, Wag82]. Logistic [Str80, Ful81, HC81b, JP89a, QL82, RT88, Rot82a, TW80a]. logit [Smi82c]. logit-log [Smi82c]. Lomnicki [Rog89]. Lon [GP82c]. long [KK81, MM82b, NS86, New83, PCTF83, STN^{+}84, War86, WWG^{+}88]. long-range [PCTF83, War86]. long-time [New83]. longevity [BGB89, Cal83b, Omh88b]. longitudinal [STN^{+}84]. longlasting [Omh88a]. look [Gat86]. loop [MDBRM87, Wil81a]. loops [PL85a, PL85b, PG88]. Loss [HRG82, Dam86, HMF88, JK89, Pys80]. Lotka [BP81a, Shi87, Tay88b, Tay88c]. Lottery [Fia88]. Low [MH80d, CA89, Fel80, Gal84a, JJ80, SK84]. low- [Fel80]. lower [KB86]. lowest [Ols87]. lowest-cost [Ols87]. LPS [Hun85b]. LSD [RB83b]. LTR [BJ84]. lumen [HF85]. lung [Col85, FB86, Hor80, KF84, Tro83]. lungs [WD86]. lying [Moy82b]. lymph [Jøn85c]. lymphocyte [Bon82, GPKBG89b, GGC86, Rao82]. lymphocyte-target [GPKBG89b]. lymphocytes [FE85, Ses82, Tho85, dBH86]. lymphoma [Mor85b]. lymphotropic [PH87a]. lysin [FB82a]. lysine [Van82]. lysine-rich [Van82]. lysis [KM84]. lysolecithin [VTL85]. lysozymes [MATGR87, MAGR87].

m [Fut86]. MacDope [BIS^{+}80]. machine [OK87c]. macrocarpa [AO87]. macromodel [ABB89]. Macromolecular [CBY82, Dem83, Kat85, WWN82]. macromolecule [KD80, TGP^{+}86]. macromolecules [Ish81, PSL^{+}85, Was82a, Was82b, WGP^{+}88, WWG^{+}88]. macroparasites [And80a]. macrophages [dBH86]. Macroscopic [Oon82, Gre80, IM85, Tom82]. MacWilliams [GS86]. magic [Rei84]. magnesium [Bol82a, DLN85, RH80, SJ81]. Magnetic [JR80, MBJK89, TL88, Yor81]. magnetocardiogram [WB82]. magnetosome [Hed85]. magnetotactic [GBFH88, Lum84]. magnetotropic [Hed85]. magnitude [BWM80]. main [Mor85b]. maintain [Sil89]. maintaining [Fuj82]. maintenance [Bul80, Gra88, GZ86, KR86, MRB83, Par82, Rap84, Ros82, ZG85, vDvD86]. major [Tre81b]. majority [BH81b]. make [Coh85]. male [BCL83, Cha89c, Cra82b, GC86, HAC82, LL87c, Man85, Sil84, TO89, WGS80]. male-biased [Sil84]. male/female [GC86]. Males [KR86, Man84, Sta85]. malignant [Gal84a, Gal84b]. malnourished [MSBB84]. malnutrition [BRRK87]. malnutrition-related [BRRK87]. Malthusian [Dem81]. mammal [Ben89, Eco81, Ipi87, Van82]. mammalian [AWMK85, AWM85, Bar81b, Cal82, Col85, COR83a, Eco83, HO84, Jam89b, Jam89c, Li88, MI81, MdMF83, Pec88, Sai82, Suc83, Suc85a, Suc85b, Syv84, Tak87, WGHY82, Won81]. mammals [Cal83a, CRB89, Fra84, Hof82, Hof83, Hof85, LIWK86, Mit89, Pro84, Whi89b]. mammary [Hol83b, SH83b]. man [Bor82, Sla85, TLS83, WG81, Zam86]. manganese [BK81]. manganese-containing [BK81]. manifestation [LG89b]. manual [Ola86]. many [And80b, AK80, Hir84b, HT87, Par82, Ros84a]. many-parameter [Ros84a]. mapping [RB82a]. maps [DG85, SW83]. March [Ano80n, Ano80t, Ano81n, Ano81t, Ano82m, Ano82s, Ano83k, Ano83q, Ano84p, Ano84v, Ano85s, Ano85y, Ano86n, Ano86t, Ano87p, Ano87v, Ano88w, Ano88-29, Ano89n, Ano89s]. marginal [CB85]. marijuana [BCL83]. marine [FQ81, LIWK86]. marker [DFS89, Fai89, FDS^{+}88]. markers [Can87, Sto84b]. market [FS86]. markings [Mur81]. Markov [Alm83, BE81a, Cur88, DC87a, Gar80, MAC89, SBPF83]. marrow [Bon82, OWN89]. Martindale [Ano80-35]. Mass [Hal80b, PSG83, SR87a, EAL80, FS85, Fra84, Jaf80, Mad88]. masses [Cor83b]. masticatory [VLCK86]. Matching [Wu88, Gar89, HS87, KG82, LVN87, MMG84]. Mate [MM85, Cha88b, GR88, GR83, HM87, Lam84a, TB80]. materials [Asp88]. maternal [HCW80, Jam85]. maternity [Hir82]. Mathematical [Ber83, BW87, BPP88, CD84, FEC87b, HR80b, KS82, Man81a, MW83a, OO84b, PL84a, PJL84, PPL85, PL85a, PL85b, PG88, SSF80, SK87, VD84, WM83, ZL84, AA83, AL87, BRRK87, BM85, BWW80, Bec81, BJ82, Bor82, BWM82, BS89, CP82, Che83, Che84a, Chi86, Fer85a, GG84b, HAC82, HB89, JD82, Jea87, KB89a, KS89, KR85b, LE84b, Mil87, MG84a, MV88a, RSE81, SMGD80, SS89b, Tay85b, URC83, WMS81, WB88, YA89, Zno83, vdH84, BIS^{+}80]. mathematics [FTDS85, WB89, WM83]. Mating [JCZ84, Abu84, Cha80d, DRB86, GZ86, GH87, Gri82a, Gri82f, Hin80, Hoe82, HT88, LA85, Mac87, TO89, WL87, Kon88]. Mating-type [JCZ84]. matrix [BHP82, DE82, Pel84, SF89]. matter [FBP87, Mäk86, Wat85]. Maturation [WC88, KW89, Que83]. maturing [Kha86]. maximal [MK83]. Maximum [Fou84, LR85a, VL86]. May [Ano80m, Ano80s, Ano81m, Ano81r, Ano82i, Ano82o, Ano83n, Ano83t, Ano84o, Ano84u, Ano85q, Ano85w, Ano86s, Ano86y, Ano87t, Ano87y, Ano88t, Ano88y, Ano89o, Ano89u, Cha86, DO83, HRG82, Has82, Niv83, Rak86, SME86, vTdJ86]. may-wigner [Has82]. Maynard [Vic88]. mean [Hin86, MK83, TW80a, YI88]. meaning [CB84, RZ86, Yos86]. meaningful [Hor80, Wol85]. means [Gar89, PDCD88]. measurable [Har80a]. measure [FF83, SL86, Sou80]. Measurement [HMF88, HW80a, Izu81, Kha86]. measurements [Bla80, DV83, Kap85, KTG86, Ols87, Tra85]. Measures [Kar80, Car81b, Lev80a, MM83b, Pen82, SBM89]. Measuring [DY82, SR87b]. Mechanical [AWM85, MK88, WD86, Asp88, DKW89, HN84, KM87, LTM85, MR89, MCB^{+}86, MHLK87, PP84a, STN^{+}84, SG89a, Smi82b, Wil86a]. Mechanics [Eng89, Ano80j, BvdHB84, DK83, GG84a, KR85a, Li86, MB84b, SV84b]. mechanism [Agm85, All85b, Bar82a, BHH83, Ber89, BHM^{+}85, Bur82b, Cla83a, COR83a, Col84, Dal83, Dav84, Dob86, DS84, Dou84, Dou83, Dou85, Fer83, FM86, Fue80, Fuj82, HAG82, HR87, Hsu87b, Hsu87c, Hun81, Hwa85, IHK80, Jac82, KS81, KK81, Kot82, Kot83, Kum82, Lac84, LH89, LSFT85, Mag84, Man81b, MB84a, Mas84, MF81, Mul87, Mur83, Mur81, Nag88, NDB80, ND82, OH84, OH85, Osu84, Ped88, Pen80, PGKE87, PGKE88, PS82b, Rea84, Ros83, Sac88, SGDK86, SM88a, Ser87, Shu84, Sim81b, Sla85, Sla86, SLYF84, SLFT86, Tor80, Von82, Wee84b, Whe84, Wig82a, Wig82b]. Mechanisms [SH85b, SPG87, Wal87, AK82, Bar83a, Bia81, Cru80, Dil86, Dou81, FDC85, GVCC82, HB86b, Kha86, KK83c, Lam84b, LCDC84, MTM86a, MTM86b, MB88b, Mur82, NFP88, PL87, Tus88, WBKW82, WB83]. Mechanistic [Aki86, DD89, Bru85a]. mechano [LM82]. mechano-electrical [LM82]. Mechanochemical [Omo89]. mechanochemistry [Gor87]. Medawar [Gle84]. media [Ree80]. mediated [BW89, Gai80, GPKBG89a, MB84a, Mas84, MAC^{+}83, Osu84, Ton86, YMIH86, Zol83]. mediator [MSBB84]. Medium [KK85, Hig85, Koc88, Kri80]. megaspores [Hai86]. megaterium [FLNF86]. Meinhardt [Mac82]. meiosis [Abu85b, Cha80e, Dav84, Dav87, Mag86]. Meiotic [JL88, HCW80, Mag84, RB82a]. melanogaster [CB86c, Nag88]. melanogenesis [MPPF80]. Melanoplus [HLS85]. Mellifera [Omh87a, OL86, Omh87b]. mellitus [BRRK87, Wee84b]. members [Gri82a, Gri82f]. Membrane [BW86b, BD80, DW83, HSd^{+}88, LKC^{+}87, LF86, SOLG82, ABB89, Bar82a, BBH86, BM88, BE81a, BN80, Bla80, BB84, Bre80b, Cla83a, CC83b, DSD81, DW81a, DW82, DZJ85, DG80, ER84, FB82a, Fri83, GK80, Gre89, HBRN81, HGG82, HFL89, IT82, JA88, KHB83, Kaw84, KKNM80, KK87a, KK87b, Kol82, KM87, KC83b, KK80, Kur84a, Leu87a, LP83, MOK87, MK88, Mar84, Mat84, MF81, Nit86, OK87b, Osu84, Rab82, Sch89, Sco84a, SME86, Sve82, TDA86, TY82, VD89, WP84, Whi82, WP81, Wig81, Zol83, vZ87]. membrane-bound [GK80]. membrane-carrier [Bar82a]. membrane-damaging [FB82a]. membranes [AH86, BP81b, Bru85b, CS84, DGST82, DT80, EM81, Gal84a, Gal84b, GK80, Hal80b, KDG85, Koc86b, KH80, LP88b, LN81, MS89a, OK87a, RAR83, RB85, RN84b, RR88b, SS85, SA86, ST88, SM81b, SG89a, TMGB80, ZG89, vGvGL88]. membranes-the [DGST82]. Memory [MH87, AA83, AK84a, AKS84, BH81b, Bou84, JSJS88, KS84, Rob81, Rob85b, SS83c, dBH87]. Mendelian [BSS83, Tre81a]. Menten [Agm85, BBM87, Dug88, Fer87, HT89c, PL84a, Sen88]. meristems [CCG^{+}82, HR86]. mesh [Sil83]. Meso [Gra85a]. mesosomes [Dou84]. message [Yca82]. messor [IAHA80]. Metabolic [Cal87, CB89b, CA89, JP89b, JP89c, KC82, LIWK86, Red88, ASK89, Don86, Fra84, GJ88, Gra81b, GM87, ITK88, KDAT84, KL83, KC83a, Li83, LL87a, MHT88, PL84a, PJL84, PPL85, PL85a, PL85b, PG88, Rei86b, Sch82b, SF89, SHBS89]. metabolically [CSL^{+}86]. metabolism [CF82, ES83, Jam80, JJ80, KK83a, Kur86, KD88, LC80, McC80b, Min82, Par83b, PP85, Pro84, SGH85, WKV88, Wil84, Wri84]. metabolite [Gra85b, SHBS89]. metabolites [Lil82, SBW83]. Metal [KS80a, Gro82, IHK80]. metamorphose [Was86b]. metastasis [WD86]. methionine [Mik83]. Method [Bas83, ZR84, AFHM86, AK82, All85a, BS88a, BSB88a, BSB88b, Boo84a, DB88, Gal83, Gol83a, HKC83, HCKB83, Jea87, KB81, LM89, MUC81, MCB^{+}86, Ola86, PPS86, SP88, SL81, SK85b, SF89, SK84, SR87b, SW81b, SG82, SJK88, TIK89a, Top88, VdV87, Wag87, WIN80, ZR85]. Methodological [Pro86, Cri82]. methodology [HH85b, LQC89, RTM81a]. methods [CB83, KS88, LH89, Per82, TJP83]. methyl [Kot83, RTM81b]. methylacetate [SLFT86]. methylation [Cla82a]. methyldiazonium [MH80c]. metric [Day83b, LR88]. metrical [Ric80]. MFA [GT88b, TG88a]. Mg [BLR83, JS80, KK83c, Szu83]. MHC [Tre81b]. micellar [SNW81]. micelles [KLK^{+}88]. Michaelis [Agm85, BBM87, Dug88, Fer87, HT89c, PL84a, Sen88]. MICMAC [HH81]. Micrasterias [PL85c]. micro [KB88a, Lap80b, MMGL87]. micro-ecology [MMGL87]. micro-organisms [KB88a, Lap80b]. microaerobically [BT85]. microalgae [CN80]. Microbes [Tro84, Gro82]. Microbial [Lap80a, Bor88, Cas81, Cla82f, Cze87, Lil82, dL83]. microbiological [BW86a]. Microcirculatory [Boy88]. microclimate [McA87]. microdosimetric [Cra81a]. microenvironment [WSMP83]. microevolution [RR83b]. microfibrils [iWSS88]. microfilaments [Rao82]. micrographs [Kam80]. microorganisms [BLA82, Lap81a, Lap81b]. microparticles [Fis85]. Micropipette [SME86]. microscopic [Gre80, IM85]. Microscopical [Zac81]. microspora [Dal83]. microtubular [ST80]. Microtubule [CTS86, PHTCS86, BH88, RMB81]. microtubules [CBLM81, FBS85, HW82, KP81]. microvascular [CB87a, SPKJ89]. microvessels [CB87a]. Microviscosity [Whi83]. microwave [Tou85]. MICs [AL87]. Might [Har82c, Pic89]. Migration [ML82, KB88b, LA86, Lun87, McC84, Moo88, PL87, Rab82]. migratory [HLS85]. mild [MP87]. milieu [RN84a, Ric87]. milk [Hol83b]. millet [LP88a]. mimetics [ABS82]. mimicry [Kel87, OO84b]. minate [MAC^{+}83]. Mind [Hon86, Hon84, Lib85]. mind-brain [Hon84, Lib85]. mineralization [Har82a]. miniature [MKL81]. Minimal [DRP^{+}87, KvDVH87, AL87, DRL88, KD88]. minimum [CB83, DV83, RTM81a, RTM81b, RTM82, SH81, Wag81]. minimum-length [CB83]. minor [ASTT80, SM88b, Sto89b, TATS81]. MIRROR [HH85b]. mispairing [MF85]. missing [Bri88]. mistaken [HK83]. Mistakes [Boy89]. Mitchell [Har80a]. mitochondria [Har82c, Woo80]. Mitochondrial [Joh84, BS87, Kho84, LM80, Mal81b, Mal84b, Mal84a, Mal84c, Mal85, MDBRM87, Mik83, MRR80, She80, She81, TH81, vZ87]. mitogenesis [Rao82]. mitosis [CT81a, Hun81, Hun85a, JS80, JS81, Kin83, Mat80, Ond80, Por83, SSF80, SJ81]. mitosis-inhibiting [Ond80]. mitosis-stimulated [SSF80]. mitotic [PE82]. Mixed [Whi80b, BP86, Cra89, Hsu87d, Row85a, Smi88, Vic87]. mixed-strategy [Cra89]. mixtures [GW88, Lam88, MB85]. MNDO [YLBB81]. MO [TIK89a]. mobile [Toz89]. mobility [DV83, Gra86b, Kri85a, Reg88]. mode [De87, Shi81]. Model [ATC81, AWMK85, AWM85, BBZ87, Ble88, Ble89, Boo84b, Coo82, Cus80, DML80, For80, GPKBG89b, KT87a, KM87, McK83, SJ81, Woj83, ZG89, Aal89, Ada80, ADG80, AMM80, Alf83, AWV^{+}88, AW83, AdMM86, Amt86, And88, AA83, AK84a, AAAT87, Ano81i, Aok84, AC89, Arm87a, Arm87b, AS84, Ash86, Asm83, ABEFS^{+}86, AL87, BS81, BRRK87, BPP^{+}86, BM85, BS88b, Bar81b, BM87, BMR89, Bar84, BRW89, Bar83b, BB81b, BE82, Bec81, Ben89, Ber83, BP80, Bha86, BVW83, Bod88, BJE89, BBS88, Bon82, BB88b, Boy83, BP83, BWM82, Bri82, BS89, BLC84, Bru85a, Bru85b, Bul80, BT80, Bur82b, BD80, BH88, Car81a, CH86, Cas81, CPLA85, CP85, CS80, Cha81, CG81a, CG81b, Chi86, CS82, CCDF81, CBF87, CJ83, CF82, Con83, Coo88, Cre88b]. model [CSC^{+}80, CF87b, CY84, Cum85a, Cum85b, Dea80, DFS89, DC87a, Dil86, DGST82, DG80, DS84, DW87, DRB86, DCR82, Ede82, EH83, Eil88, EAL80, Emi82, Fai85, FY81, Fel83, Fer85a, Fer87, Fia88, FP89, FW85, FEC87b, Fle88, FJFN89, FGDS87, Fue80, FBS85, Fuk83, Fuk84, GDHB87, Gar87, GTH88, Ger85, GT88a, Gli80, GBS89, GF81b, GR83, GT89, GS86, Gro80, Gro88, GP82c, GH82, HW88, Har84b, HB86a, HSV84, HAC82, HCW80, HF85, HS81, HFL89, HBA89, HLS89, Hin80, HL86, Hje86, Hof86, Hol89, Hop86, HBH83, HT88, HT89c, HT89b, IYS83, Inc83b, Ing89, IH80, JC80, Jam80, Jan82, JK88, Jea80, Jea87, JJ80, Ji85, JT84, Jol89, JLS80, JS82, JP89b, KAM85, KRT81a, KRT81b, KW89, KS80b, KFSW82, Kaw84, KP89a, Kel87]. model [KS84, KB89a, KS89, KKNM80, KDAT84, Kin83, Kin85, KL81, KF84, KSL89, KR85b, KIO88, KC83a, KO83, KK86b, KB88b, Kum82, Kur83, KD88, Kwo84, Lan84a, Lan85, Lán84b, LCDC84, LP88a, Las87, Lav84, Lef83, LH86, Ler83, Leu87b, LP82, LTM85, Lew81c, LS87, LG89a, Lim88, Luc82, Lum84, LN81, LM82, LH80, MK81, Mac82, Mac87, MM82a, MZ89, Mäk86, Man81b, MM83a, MRG84, MG81, MS89a, MF80, Mat84, MKL81, McA87, MB86, MAC89, Mea86, MK87, Men85, MBJK89, MdMF83, MA83, Mil87, MTM86a, MTM86b, MM83d, MHII83, MAC^{+}83, MG84a, MH84, MOV82, MV88b, MS86a, MHB89, MYS89, MG85, MS89b, NM81, Nag81, NMM87, Nev88, NBZ^{+}83, NSCH88, NL86, Nus82b, Oht87, OKT^{+}80, Ola86, Orl81b, Oth86, PP84a, Pap84, PP84b, PC86]. model [PF89, Per84, PH89, PS82b, Pla82, Pok86, PL87, Poz87, PH87b, PL84b, Pro80, PJ86, RTM81a, RTM81b, Rao82, Ras87, RAR83, RB85, Rég84c, Rei86a, RKL83, Rik82, Rob84, Rob87, Rob80, RR83b, Rog83, Row87, Row88, RSM80, RS86, SRPS89, SBC80, Sau80, SL87a, SL87b, Sch89, SSF80, Sch83b, Sch82a, SC84a, Sco89, SS83b, SK86, SP87, Ses82, SN80, She83, SG88, STN^{+}84, Sib81, Sib83, SV84b, SS83c, SG89a, Smi82c, SLN84, SS87, SS89b, SBPF83, Spe86, Sta88, SHBS89, Ste80, SF83, Ste84b, Sto83, Sto89b, SCP84, SH82b, Tak87, Tak89a, Tak89b, Tan84, Tat88, Tay85b, Tay88a, Tay89b, TT85, Tho86, Tod82, TS88a, TLT80, TT87, Tou85, TH85, TGP^{+}86, URC83, Val85, VCC^{+}88, VD84, Voi85, WFL86, WC88, WGP^{+}88, WWG^{+}88, WS82]. model [WF82, Wil81a, WS85b, WB88, Won81, YMS84, Yam87, YA89, YVIG80, YMIH86, dGV80, vTdJ86, vdH84]. modeled [DZJ85]. Modeling [Cha80a, FBP87, Har82a, Sim80a, Sim80b, SV89, Süh89, Cha88c, CD84, Gri81a, Hej82, Hsu87b, Hsu87d, Hsu87c, IBJ^{+}89, Ito82, PSG83, PGKE87, PGKE88, SSF87, SH82a, WGKG83]. Modelling [AFHM86, HD84, LP86, PP85, Sim81a, BPP88, HR80b, HH81, HH85b, HB89, KTG86, PL84a, PJL84, PPL85, PL85a, PL85b, PG88, PS83, PR82, Sto84a, Sto84b, War86, vHBL87]. Models [AL81, Bel80, CHT80, LP88a, MW81, MG84b, AM81a, Abu84, AH86, ASTT80, AM81b, Arg81, BMR89, BRR87, Bel87, BP81a, BSS83, BP87, BW87, COG83, CG85b, Cur88, DH85, Der87, FCBK88, FS83, GP80, GT88b, Gre84a, HC81b, HCVM88, IM85, JK81, KC85, KM83, KR82, Koc82, KWG87, KK80, Lan82b, LRKC83, LJTH85, Leg87, Les86, LR83, LE84b, Mac84, Mah84a, Mah84b, MB80, MJ88, Mat87, MGL81, McG80, McG84, McG89a, MAL87, MM80, Mur82, Nun80, Oli85, Pah85b, PK80, QL82, RL82, Sal84, SSM81, SA82, SCH84b, SM80a, SM80b, SS87, SGd84, Ste84a, SF83, SKF83, Sto89c, Stu85b, TH80, Tay88b, Tay88c, TG88a, TW80a, TACM82, Tys83b, Val85, VP86, VKV83, Wal89b, Wal89a, WOT85, Whi80c, Wri84, TH80]. modes [BS87, KR85b, Wol85]. modification [Cra81a, Fri85, HM80, LvH87, Rak80, Rak86, Rob80, SGDK86, WT87, Yok84]. Modifications [LJTH85]. Modified [MUI85, PA82]. modifier [SM88a]. modifiers [Bur83, Lan82a, WB85]. modulation [Ros84c]. modulus [Eng89]. moiety [YLBB81, YBBL82c]. mold [GS86, RSE81]. molds [Arm84, BS88b]. Molecular [Bar81e, Bia81, Bla84, CL82, ELM81, Fra88, GPM89, HT87, ITK88, IT82, SRO84, Sar89, Sar85, Ahl81, ANOM82, Aid88, ALST81, ABS82, Bar81d, Bar82b, Bel80, Bod88, BJE89, BVDA87, DD85a, Els82, Gli80, HKC83, HCKB83, Hol88, HFO83, IO83, Inf85, Ji85, Kat81, KK87a, KK86a, KS81, KK81, Kot82, KK83c, Kot83, LTZ88, Mat81, MHB83, MB88a, MB85, NA84, NU82, Nev88, Nus87, RTM81a, RTM82, SS89a, Sch82b, SR81, Sib83, Suk89, UN82, Was82a, Wic85a, YHJ89, GPM89]. molecularly [Koo85]. molecule [BRW89, FY81, NU82, Rob81, Rob85b]. molecule-group [Rob81, Rob85b]. molecules [DD85a, HFO83, KB83, Kuh82, McG81, MM86b, MG84a, Por81, Sch85, TS80, Wil82]. molecules-Properties [McG81]. molluscs [FDC85, MK87]. moment [Har84b, LP82]. momologies [Sab80]. monitored [Kuh82]. monitoring [BG87]. monkeys [KMSR84, SK87]. Mono [Kol82, YBBL82a]. Mono- [Kol82]. mono-saccharides [YBBL82a]. monoculture [WJS89]. monocultures [FW85, Har84a]. monoecy [Wie81]. monogamy [MM85, Rut83]. monolayered [HE80b]. monomer [KP81, RdL83, Whi82]. monomer-dimer [Whi82]. monomer-polymer [KP81]. monomeric [CBC87]. monomers [Blo87]. mononucleosomes [Sch82a]. monophosphate [Fel80, ZBL80a]. monopole [PSL^{+}85]. monopolization [LS89]. monovalent [MJN88]. Monte [EG87a, EG87b, Hsu87a]. Moore [Pel84]. moose [Rei86a]. MOPC167 [CSH^{+}81]. Moran [Aal89]. Morphogen [Hun85a, Cum89]. morphogenesis [BBZ87, BG88, BKJT80, Cum85b, FJFN89, GT85a, HC81a, HH85a, HK88, HT89a, Klä83, MG80, Nev88, PV83, PL85c, Ste84b]. morphogenetic [RSE81]. Morphological [DfP87, IP83]. morphology [FDC85, Koe83, RS87, Smi82b]. Morphometric [Suc85a, SW84]. mortality [BGB89, Cal83b, Hol85, Ing89, Lib88, SH82a]. mortalization [HK81]. mosaic [Dun82, IBJ^{+}89, Sim81a]. mosquito [Pet80]. most [CB88b]. moth [LL87b]. mother [HCVM88]. mother-daughter [HCVM88]. motif [AM87, CB88b]. motile [SK88b]. motility [Fue80, Gor87, HJS84, Hoe84, MB88a]. motion [Fri83, MM86b, MYS89, NS86, Rik82]. motional [SM81b]. motivational [CG85a]. motive [Dun82, Hay80b, Tai86]. motor [AA85, Fue80, MYS89]. Motors [FH89, CS87]. moulding [GJG83]. mouse [ATC81, MKM80, PM80, PL87, SSF87, Vau86, WF82]. mouth [Mul87, NK88]. Movable [GT88b, TG88a]. move [Jac83]. movement [BO82, Bre80b, FBS85, Hun81, INMM83, MJ88, MdMF83, MB88b, MS86a, Wag82]. movements [AWMK85, BM89b, Ben89, BB88b, MB88a, TSRK88]. moving [HE80a, See86]. mRNA [Fel83, HR80b, HB86b, HBA89, Iid88, MDPF86, NDB80, Nus88, VZR85]. mRNAs [KvDVH87, Nus82b]. MS2 [NST80]. much [Gri83, HE80b, KR82]. Muco [FB86]. Muco-ciliary [FB86]. mucosal [MTM86a, MTM86b]. Muller [NS89a]. multi [Cre88b, Gar89, Hin87, KK87b, Ric87, VC88a, WJT87]. multi-allelic [VC88a]. multi-channel [KK87b]. multi-enzyme [Gar89, Ric87]. multi-locus [Hin87]. multi-peptide [WJT87]. multi-phenotype [Cre88b]. multicompartment [JB80]. multicompartments [IP83]. multienzyme [GS85, Kur86]. multilayered [DZJ85]. multinode [Eil88]. Multiple [All85a, Pok82, WL87, Alt89, AK84a, DG85, DLN85, FS80, Gar87, MAL87, TO89, Ven82, Wad82, Wad85, Whi89a]. multiplicative [Has81, Pin84]. multiplicity [Kra82]. multiply [DG87]. multisite [Whi80b]. multispecies [KR82, PH89]. multistationarity [KT87a]. Multistep [Koc82]. multistranded [McG81, McG84]. multivalent [GS81, KD80, MJN88, NSW82]. multivariable [HS81]. multivariate [Ano80-35, BG86, Ipi87, Mar80b]. münch [SMGD80]. murein [LBBL84]. Muscimol [ABS82]. Muscle [BM89b, MB88b, BM89a, BvdHB84, CG81a, CG81b, GG84a, HBRN81, JCK80, Kas80, PC86, RH80, SS87, WGKG83, Win89a, Won81, YA89, Bar82b]. muscular [Inf85, Obe81, See86]. musculus [KB88b]. must [Bel80, FBB89]. mutagenesis [BG86, Cla82f, Sir86]. mutagenicity [DD89, MHB83]. mutagens [BG87, KM81, OC81, SL80b]. mutagens/carcinogens [BG87, SL80b]. mutant [PWH^{+}88, SK85b]. mutase [War86]. mutation [Alf83, Frö88, HK81, Ish86, MD86, Mur85, NS89a]. Mutational [Cla82e, Cla82b, Cla82c, MM85]. Mutations [Mur85, Row85a, BPH87, CM82, Cla82d, Cla83b, Gen88, Kon88, Man82b, Mor82, RGR82, SV84a, Wag81]. Mutual [Kar88b, YVIG80, Blo87, Mot88]. mutualism [Arm87b, Ton86]. mutually [KK83b, LK83]. mutually-coupled [LK83]. myc [Coo88, GDHB87]. mycelial [ES83, EH83]. mycoplasma [Sla86]. myelin [RB83b]. myelination [Fra87]. myeloma [MKM80]. Myocardial [Reg84b, MR86, RM86, YKS80]. myocardium [AWMK85, AWM85, Bol82a]. myofilament [Mor85a]. myosin [MB88a, Obe81]. myxobacteria [LGK84]. myxomycete [UK82].

N [Ano81i, YBBL82b, GPM89, RTM81b, WMR86]. N-acetyl-giucosamine [YBBL82b]. N-acetylglycine [RTM81b]. N-allyl-Normetazocine [GPM89]. N-methyl [RTM81b]. N0acetylneuraminyl [WMR86]. nalorphine [GPM89]. narcosis [Fox81]. native [Sim85]. Natural [Coo81, Hir82, MLB88, BKMS87, Coo84a, CK82, Dil86, DB88, FSF89, GJG83, Gri81a, Gri81b, Gri81c, Gri81d, Gri82b, Gri82c, Gri82d, Gri82e, Gri82a, Gri82f, HBP89, Hir80, Mac87, Per82, Tak89c, Tat88, Ula80a, Ula80b]. nature [Bân83, Dou84, Esh84a, Hux86, Mal81b, MKM80, Pet83, RSR88, Rei85a, Sve81, TN80]. navigation [Wal89b, Wal89a]. near [Ber89, Nus88, Rab82]. nearest [Day83b]. Nearly [SK85a]. necessary [Eps86, McC80b, Rob80]. nectar [Ano80j]. Negative [Jak88, Gho81, HF88, Hun85b, KD80, Mah84a, Mah84b, ST88]. Negentropy [Jaf84]. Neglect [Pro84]. neglected [Siz87, SP84]. Neighbor [Ree80, Day83b]. neighborhood [Gra81a, Koc86b]. neighbour [Sto83]. neighbour-exclusion [Sto83]. neighbourhood [MG84b]. neighbouring [ND82]. Nelson [TLN85]. Neonatal [GK85, SM88b, Hof83]. neoplastic [LC80, MKM80]. Neosteingelia [ST80]. Nernst [RCA89, Shi80]. nerve [Bar81d, CJ83, HBRN81, Miy86, PS81a, Sab80, SM81a, Str85a, Ten80, TW80b]. nerves [AP80]. nervous [TG80]. nested [CB88a]. nesting [BL88, Bul84]. nestling [BS85]. net [AK84a, AAAT87]. nets [AZE83, AKS84, KA85, Sil83]. Network [FM88, ST88, Bro81a, CP82, Che84a, FE85, GH82, HBP89, Hof86, Hof87, KC82, MTM86a, MTM86b, PP84b, Peu83, Wag81, Peu85]. Networks [Wei86, ABEFS^{+}86, Che83, CB87a, HS81, HU85, JJN86, Kin82, LD83b, LL87a, MA83, PL84a, PJL84, PPL85, PL85a, PL85b, PG88, SPKJ89, dBH89]. neural [And88, AZE83, AK84a, AKS84, AAAT87, Bar89, CG85a, CPLA85, Eil88, Hof86, JPL83, KS80b, KHU^{+}83, KA85, LCDC84, Oon82, PP84b, Sal84]. neurochemical [KC83a]. neurocranium [Mul87]. neuroeffector [DSR89]. neuroendocrine [Wee81, Wee84a]. neurogenic [Wee81]. neurogenic-neuroendocrine [Wee81]. neurology [MLB^{+}89]. Neuromagnetic [Lib80]. neuromelanin [Lac84]. neuromorphology [Mik84a]. neuromuscular [Vau86, VM89]. neuron [Kaw84, KK83b]. Neuronal [Hof85, Ipi87, KG82, KMSR84, Lán84b, TC80b, WR83]. neurone [KT81]. neurons [McG85]. neurophysiological [MM83b]. neurotransmission [MP85b]. Neurotransmitter [LPS89, LPS86, PS80a, PS81a, PS82a, PS84, PPS86]. neutral [IH80, Orl81a]. newly [WGHY82]. Newman [Oth86]. Newman-Frisch [Oth86]. NH [NSKS85]. Niche [Ebe87, GP82b, Mor80]. nicking [CCG87]. nicotinic [TS85]. night [Pic89]. nitella [TY82]. Nitrogen [Jon85a, Bru85a]. NK [KW89]. nm [ZG82]. NMR [DDA85, GPLC84, KC83b]. No [MHLK87, LS81]. node [DSR89, Jøn85c, UKNS87]. noise [GGPM80, Str80, TW80b, Tuc87, VG80]. Noisy [AKS84, KA85]. Nomenclature [JJ80]. Non [Abu81, Bar81e, FJ85, HC81b, Ish81, JJ82a, RF82, SSM81, SM81b, AMI84, AS87, BM87, BMR89, Bro80b, Con83, DKW89, Dug82, FE85, For85, FLR83, Gad81, GA80, Gri82c, Hje86, Hol85, IHK80, Kat85, Kaw84, KCM81, LK83, Mal88, Man81a, Mic86, MG84a, MH84, New80, New83, NC89, PS83, Sch89, SK88b, SM88b, Ste88, Str80, SG89b, TT87, Tra85, TLN85, WB85, dBH87]. Non-autonomous [HC81b]. non-body [Con83]. non-co-operative [NC89]. non-coding [AS87, Mic86]. non-competitive [Man81a]. Non-enzymatic [RF82]. non-enzymatically [SG89b]. Non-equilibrium [Ish81, Tra85, WB85]. non-existence [GA80]. non-extinction [Gad81]. non-genic [Hje86]. non-growing [MG84a]. non-hyperbolic [IHK80]. non-ideality [Kat85]. non-immune [FE85]. non-immunological [For85]. non-inactivating [Ste88]. non-insulating [TLN85]. Non-linear [Abu81, SSM81, SM81b, BM87, BMR89, Bro80b, DKW89, Dug82, Hol85, LK83, Mal88, MH84, New80, New83, PS83, TT87]. Non-local [FJ85]. non-motile [SK88b]. non-periodic [AMI84]. Non-random [JJ82a, Gri82c]. non-self [For85, dBH87]. Non-specific [Bar81e]. non-steady [KCM81]. non-steady-state [FLR83]. non-thymic [SM88b]. non-uniform [Kaw84, Sch89]. non-white [Str80]. none [Wri80]. nonenzymatic [Kri85b]. nonequilibrium [JF85, JPGL86]. Nonlinear [KS88, PY88, BESD84, BM88, Bur83, COG83, Hsu87c, Mur83, Pin83a, Pin83b, Tam87]. nonlinearities [RR88a]. nonstationary [BBH86, BM88]. nontransformed [Kle81]. Normal [RTM81a, AS86, GP82c, Kle81]. normally [HK80]. Normetazocine [GPM89]. Note [KW86a, KW86b, KW86c, KW86d, Was86a, Arm88a, BP87, Jam82, LR83, MC89, SS82, Sel80, WD81]. Notes [Hok86a, KC87]. Notonecta [Hor89b]. novel [BS88a, Col84, Gie88, IHK80, Mon88, She80, WK81]. November [Ano80p, Ano80v, Ano81w, Ano82l, Ano82r, Ano83o, Ano84q, Ano84w, Ano85p, Ano85u, Ano86q, Ano86w, Ano87u, Ano87-27, Ano88x, Ano88-30, Ano89q, Ano89v, Ano81-36]. novo [Fer88]. Nuclear [Suc85b, Bon85, HBA89, Nus88, TH81]. nucleation [CSL^{+}86, FDR83, RdL83]. nuclei [Ipi87]. Nucleic [See82, Arg81, BS88a, Goo84, LTZ88, McG80, McG81, McG84, MT84, MF85, dW81b]. nucleocytoplasmic [CS82]. nucleophile [GBS89]. nucleophilic [MH80c]. Nucleotide [Alm85, EG87b, Fel80, Kre82, LVN87, NDB80, Nus80a, NP84a, NP84b, SSW81, Wu88]. nucleotides [LH80, PRH81, PA82]. nucleus [RB82a, Suc85a]. Number [WJT87, BMR89, Fut86, Has88, HM80, Li88, MdMF83, NST80, VC88a, Wag81, dW81b]. numbered [CG84]. numbers [Gul82]. Numerical [Mac82, PSG83, KG82, RCA89, SM80b]. Nussinov [BT87]. nutrient [Fin82, GF81b, Rob88, Rob89]. nutrition [KKJ81, SH80b]. nutritional [GZWR80, WGRZ80]. nutritive [MH80d].

O [PRH81, BT85]. obeying [Cha82a, RT88, WdlT84]. objectives [CB89b]. objects [Pin83b]. Oblique [FHL89]. observations [HK81, Leg87, SCB81]. observed [Ham85, Wei80]. obtained [DCR82]. obtaining [WBKW82]. occlusion [Fra88]. occur [vTdJ86]. occurrence [Mur85, Tre81b, Vil83, Vla87]. October [Ano80-27, Ano80-33, Ano81s, Ano81z, Ano82w, Ano82-29, Ano83w, Ano83-30, Ano84z, Ano84-33, Ano85-28, Ano85-33, Ano86z, Ano86-32, Ano87-32, Ano87-39, Ano88-36, Ano88-41, Ano89-30, Ano89-36]. ocular [Mon88]. odorants [MM83a]. odour [ELM81]. off [ND82]. offspring [Ell86, Har80b, Jam87, Rei85b, Rei87, Rob80, Tay88a]. OH [NK88]. ohmic [Bar87, Pic86]. old [GP82b, HKC83]. olfaction [MM83b]. olfactory [Ben89, NM81]. oligomer [GK80, LN85]. oligomeric [GD81, RN86, dVR85]. oligomerization [FSF89]. Oligopeptides [Klä83]. on/off [ND82]. oncogene [Gar87, Mer87, NDB83, NMB85]. oncogenesis [Ahl81]. Oncopeltus [Bar82a]. One [AM81b, DG85, FDR83, FS85, För88, Gri82d, HH81, Koc89, KK86b, LP88a, LH86, MM86a, MB85, RdL83, SK85a, SM88a, SV84a, VC88a, vRV80]. One- [AM81b]. one-dimensional [DG85, FDR83, FS85, RdL83]. one-locus [VC88a]. one-step [För88]. ones [KR82]. only [Hoe80, Law83]. Onsager [MF89, Peu86]. onset [NS89a]. ontogenesis [Ano81i, KRT81a, KRT81b]. Ontogenetic [Dil86]. opaque [Fis85]. open [HC89, Hsu87d, Par85, RS82, Web81, WGP^{+}88, WWG^{+}88]. operates [LH89]. operation [TG88a, dBH89]. operative [BC80, KKNM80, MGL81, NC89, Shc88, Shc89a, Shc89c, SH83a, Whi80c]. operativity [BWW80, Bas83, CBC87, CN80, Gho81, IHTN88, Kar80, KD80, PPS86, RH80, RN85, RN86, Whi80b, Whi81b, Zol83]. operator [Was86a]. operators [Oli85]. operon [BPM82, Man81a, Man81b, SR88]. ophidine [Bro81b]. opioid [GPM89]. opponent [Par83a]. opponent-independent [Par83a]. opposed [Kno82]. opposite [Kyp88]. optic [Tho83]. Optical [GK81, Faj82, Hor89b]. optically [Der87]. optically-detected [Der87]. Optimal [BMR89, Bro82b, Cas82, Cas84, EC81, Gre82a, Hou87, KB86, MH85, Pug87a, Rob88, SCS88, Tak89c, WW84, Bel87, BT85, BPH87, Har87a, Hou86, MD86, Rei86a, SPKJ89, WH83, YC89]. optimality [BL85, SH87, WH86, ZB84]. Optimalization [ES88]. Optimization [Blo85, Mul87, Ros84a, SL81, CF82, FBF85, Hor89b, HM85a, Lef83, LB87, RTM81a, RTM81b, RTM82]. Optimum [Ale84, Sea80, Yat86, Spe86]. orb [COA85, FB82b]. orb-web [COA85]. orbital [ANOM82, Aid88, ABS82, BVDA87, GPM89, HFO83, IO83, ITK88, NA84, NU82, PRH81, PR82, UN82]. orbital-centred [PRH81, PR82]. orchidaceae [DL85]. Orchids [DK82a, DK82b]. Order [BB85, Bas80a, KC87, MSS81, Pap80, Pap82, SKK84, SR80a, SBM89]. ordered [Bro80a, vGvGL88]. orderly [Hor89a]. orders [Pec88]. Ordinal [BC86a]. ordinate [Sib81, Tor80]. ordinated [NU82, TSRK88]. ordinates [Pap89]. organ [Nag86, SM88b]. organ-specific [SM88b]. organelle [BS84a]. organelles [BS84a, MB88b, SH83a]. organic [KLK^{+}88, Min82, SAS85]. organism [MSA^{+}81, SGH85, VL86]. organism-centered [MSA^{+}81]. organismal [WRW83]. organismic [Wic85a]. organisms [CK87, Cla82a, Gra85a, Hoe80, Kha86, KB88a, Lap80b, Lum84, Wad82]. organization [AM82, ABEFS^{+}86, Bel84, Bod88, BJE89, EGS80, FDC85, FM83, GTMM87, IH80, Ish84, Ish86, KSM85, Leo84, MYS89, Nag88, Pin84, RS82, Rya80, Sin82, Sla85, Suc85b, Tak87, Wei86, Yoc81b]. organized [TS88a]. organizer [Pat84]. organizing [Sin82]. Organotaxism [Mic88]. organs [Bas80a, Bas80b, dCB84, Hej82, HR84, Was82b, WD86]. orientation [Ben89, Bor84, DD85b, DW87, FDC85]. Orientational [vGvGL88]. orientations [AL83]. oriented [Kam80]. Origin [BBHM84, Dou81, Faj82, AM81b, Bân83, Bar81a, Blu85, Bro83, BP85, Dou84, Eco82, Fer82, HW88, Ham85, HB86b, Koc89, Lam84b, Maw85, MKM80, Nic84, RB82c, Ros83, Sin82, Smi87, SD87, TRS84, Tra86, Wie81, Yoc81b]. origination [Pea88a]. Origins [BS85, Gra86a, Gra83a, Org86, RB82e]. Orthogonal [Ros84a]. orthophosphate [KBR88]. Orthoptera [HLS85]. Oscillating [TS83]. oscillation [KT81, NL86, Tys83a]. Oscillations [NS89b, Sau80, BS88b, COA85, GAD80, MS80]. oscillator [AMI84, DDBB83, DCR82, GGPM80, Hao82, KFSW82, LS87, MS80, Pet80, RL88, SL87a, SL87b]. oscillators [BESD84, KS80b, LG89a, LK83, Nan86, Sin83]. Oscillatory [LG89a, All85b, DG87, HN82, MUK88, Osu84, PB81, SF83, Tys83a]. Osmosis [Fer84]. osmotic [KM84, Rap84, Wag87, Wol80, Mas84]. Osuji [All85b]. other [Bar81b, Der87, Els82, HT87, Koc89, KP81, Mor85a, MG84a, SG82, Wu88]. outcrossing [BA80]. outer [WP84]. output [Fin82, Kre82]. output-input [Fin82]. over- [GZ86, ZG85]. Overcoming [LKC^{+}87]. overcomplete [Pel84]. Overdominance [GZ86]. overlap [Bas81, Gal83, GC86]. Overlapping [SW81a, McM81, Por82, Yoc81a]. oversaturation [AK87a, AK87b]. overview [Bel87]. ovis [LC82]. ovulatory [SN80]. ovule [Lew82]. ovum [Bha86, BVW83]. own [Bar88, San84, SJ85]. owner [LE84a]. owner-intruder [LE84a]. oxidase [Mor86]. oxidation [HFO83]. oxidative [FJ85, JF85, Kho84, Mal84b, Mal84c]. Oxygen [Pop82, BJ82, CBF87, KB89a, Lay87, Ros85, SR87b, SR89, Tat88]. oxygen-18 [SR87b, Tat88].

P [PRH81, Bar86, DDA85]. p21 [Mur85, Mur87]. paced [CJ83]. pacemaker [KS80b, PS80b, YVIG80]. pacemakers [IYS83]. packages [BNH80]. packings [SK84]. Paddles [CK87]. Pages [Ano81-36, Ano80-34, Ano80x, Ano80-29, Ano80y, Ano80-31, Ano80-28, Ano80w, Ano80-30, Ano80k, Ano80q, Ano80z, Ano80-32, Ano80l, Ano80r, Ano80n, Ano80t, Ano80m, Ano80s, Ano80p, Ano80v, Ano80-27, Ano80-33, Ano80o, Ano80u, Ano81y, Ano81-32, Ano81-27, Ano81-31, Ano81-30, Ano81-34, Ano81-28, Ano81-33, Ano81o, Ano81u, Ano81x, Ano81-29, Ano81p, Ano81v, Ano81n, Ano81t, Ano81m, Ano81r, Ano81w, Ano81s, Ano81z, Ano81q, Ano81-35, Ano82y, Ano82-30, Ano82v, Ano82-28, Ano82x, Ano82-31, Ano82-27, Ano82-32, Ano82n, Ano82t, Ano82u, Ano82z, Ano82j, Ano82p, Ano82m, Ano82s, Ano82i, Ano82o, Ano82l, Ano82r, Ano82w, Ano82-29, Ano82k, Ano82q, Ano83v, Ano83x, Ano83-32, Ano83y, Ano83-31, Ano83-28, Ano83-33, Ano83-27, Ano83-34, Ano83p, Ano83u, Ano83z, Ano83-29, Ano83m, Ano83s, Ano83k]. Pages [Ano83q, Ano83n, Ano83t, Ano83o, Ano83w, Ano83-30, Ano83l, Ano83r, Ano84-28, Ano84-34, Ano84y, Ano84-31, Ano84-29, Ano84-35, Ano84x, Ano84-30, Ano84r, Ano84-36, Ano84-27, Ano84-32, Ano84n, Ano84t, Ano84p, Ano84v, Ano84o, Ano84u, Ano84q, Ano84w, Ano84z, Ano84-33, Ano84m, Ano84s, Ano85-30, Ano85-36, Ano85z, Ano85-32, Ano85-27, Ano85-34, Ano85-31, Ano85-37, Ano85r, Ano85x, Ano85-29, Ano85-35, Ano85o, Ano85t, Ano85s, Ano85y, Ano85q, Ano85w, Ano85p, Ano85u, Ano85-28, Ano85-33, Ano85n, Ano85v, Ano86-29, Ano86-35, Ano86-30, Ano86-37, Ano86-27, Ano86-34, Ano86-28, Ano86-33, Ano86p, Ano86v, Ano86-31, Ano86-36, Ano86r, Ano86x, Ano86n, Ano86t, Ano86s, Ano86y, Ano86q, Ano86w, Ano86z, Ano86-32, Ano86o, Ano86u, Ano87-28, Ano87-35, Ano87-30, Ano87-37, Ano87-33, Ano87-34, Ano87-29, Ano87-36]. Pages [Ano87q, Ano87w, Ano87-31, Ano87-38, Ano87r, Ano87x, Ano87p, Ano87v, Ano87t, Ano87y, Ano87u, Ano87-27, Ano87-32, Ano87-39, Ano87s, Ano87z, Ano88-34, Ano88-39, Ano88-33, Ano88-40, Ano88-35, Ano88-42, Ano88-32, Ano88-37, Ano88s, Ano88z, Ano88-31, Ano88-38, Ano88u, Ano88-27, Ano88w, Ano88-29, Ano88t, Ano88y, Ano88x, Ano88-30, Ano88-36, Ano88-41, Ano88v, Ano88-28, Ano89y, Ano89-31, Ano89-29, Ano89-35, Ano89-28, Ano89-34, Ano89z, Ano89-32, Ano89m, Ano89t, Ano89-27, Ano89-33, Ano89p, Ano89w, Ano89n, Ano89s, Ano89o, Ano89u, Ano89q, Ano89v, Ano89-30, Ano89-36, Ano89r, Ano89x]. PAH [Sýk84]. pain [BS89]. pair [Cla83b, Gri82a, Gri82f, SM81c]. pairing [JL88, Mag84, NP84b, RB82b, RB83a, Shc89c]. pairs [BHM^{+}85, GPLC84, KK82b, Mik81, Toz89]. Pairwise [Sar87]. palindromes [Dug81, Vil83]. Palindromic [Blu85, GF81a, MM82b]. pancreas [Wee81]. pancreatic [NSKS85, UN82]. paper [GRJ81, Met82]. papovavirus [BJ84]. paradigm [GP82b, Was82a]. paradigms [OO84b]. paradox [Eco84]. Paradoxes [Ros80b]. parallel [Lin80, YVIG80, vHS82]. Paramecium [VV82]. Parameter [HBRN81, Mur82, BGG84, Dem81, JP89a, Ros84a, SA86, Tay88b, Tod85]. parameterisation [PRH81]. parameters [AZE83, AK80, BE86, Bla80, BW85, CC82b, CC83a, DG80, EC81, FBF85, Gai80, Ger83a, GPPZ80, HT87, HT89c, HSd^{+}88, JA88, Leg87, LN85, LP82, LTZ88, LPS86, MB80, MF81, Moy82a, RW82, Rot87, Sto89c, Top88, Ven84, VA88b, YA89]. Parametric [KHB83, MS89a, Mik84a]. Parasite [GGC86, BP83, JK88, KM83, Kno89, Mot89, SJ85]. parasite-derived [SJ85]. parasites [BNH80, Car81b, Kir86, MC89]. parasitic [Bro82a, Gre82a]. parasitoid [ADG80]. parasystole [IYS83]. parathyroid [All85a]. Parent [Har80b, Ell86, Hou87, Rob80, Yok84]. Parent-offspring [Har80b, Ell86, Rob80]. Parental [Jam89b, Jam89c, Blu83, WGS80]. parents [Jam86a, Jam87]. Parity [Tra86]. Parity-violating [Tra86]. Parkinson [Bar86]. Parsimony [Cri82, Day83a]. Part [JP89c, Was82a, Was82b, Gal84a, Gal84b, Koc88, Wal89a, YBBL82c, BKB80a, BKB80b, JLS80, JP89b, Wal89b]. parthenogenesis [Bul82, Cha82b, KR86, Man81c]. Partial [LG85, Lun87, Bec80, GC86, Kar88a, Ped88, TD81, Was86a]. Particle [BF83, Wei80, Wil86a, WS85b]. particles [Bec80, BF83, CS82, HL86, Kam80, LP83, Wat85]. Partition [Koc88]. partitioning [BS84a, HD84, LS88, Mäk86, Rog89]. parturient [GTH88]. parvalbumin [RH80, Rei85a]. passage [Gad81]. passerines [CJ82]. passive [Sch89]. patch [Arm87b, Arm88b, Cha88c, COR83a, KK87a, KK87b]. patches [Hou86, LSH88]. patchy [CB85, GN86, HH81, IM85, Ive88, MN89a]. Paternity [WGS80]. pathogen [Bre80a, BF85, Joh86, Lew81a, Lew81b]. pathogenicity [Flo84]. Pathogens [Too82, Car81b, For85, Lew81b]. pathway [Bon82, ITK88, McK83, RP84, RF82, SF89, Tho83]. Pathways [LP88b, Gar87, Gie88, HdL84, MHT88, RTM88]. patients [KKJ81, MSBB84, MH81]. Pattern [BKJT80, MM82a, AS87, BBZ87, BHH83, Chi86, CPPR87, CPPRS87, CJ83, Cum85a, Cum85b, HSV84, KS84, KB88a, KIO88, Mal88, Mea86, MK87, Mur81, Nag88, NFP88, PV83, Pap89, Pat84, Phi81b, Sac88, Sai82, SK88a, SK86, Sib81, Tod82, TCY85]. pattern-surface [Cum85b]. patterning [Fra82, SCH84b]. Patterns [CV88, Hun85b, VC88b, AVR82, Bar81b, BHH83, BLA82, Bri80, Gol88, GB87, Gun89, Hol83a, Jea87, LS82, LA86, MN85, NMM87, OK82, PW88, SCH84b, SCN85, Sil89, Vic89, Wei80]. paucity [Dug81]. PCBs [MP87]. PCILO [YBBL82a, YBBL82c]. PDDM [BRRK87]. PDR [San88]. peak [RB87, Vau86]. pearl [LP88a]. peeling [Toz89]. pellet [EH83]. penetrable [MOK87, OK87a]. penetrance [Rog83]. penetration [JD82]. Penrose [Pel84]. pentane [RTM81a]. pentose [MHI85, MHT88]. peptide [Dou81, Dou85, FB82a, Gli80, Mag82, Pet82, SH80a, WJT87]. peptides [GVP^{+}80, PR82, VGP^{+}80]. peptidoglycan [LBBL84, SP87, YLBB81, YBBL82a, YBBL82b, YBBL82c]. per-taxon [Pea88a]. perceive [Nor83]. perceived [MM83a]. percentage [SH83b]. perception [DK82a, DK82b, Nor83, Ric87, VD84]. Percolation [Sta88, FJFN89, SYM86, dBH89]. perennial [Pug87a]. perfection [KVW88]. performance [AY89, Jol88, Koe83]. performed [ML84]. perfused [Bas80a, Bas80b, JA88, VKV83]. perfusion [BRR87, WI80]. Period [RL88, AS84, CB86c, EP86, McN89, Tay85a, Tys83a]. period-doubling [AS84]. Periodic [AMI84, BH81b, Tys83a, AVR82, DG85, GA80, Gro88, HS89, RL88]. periodically [AMI84, GGPM80, Sta86]. periodicities [DR89]. periodicity [SL86]. periods [War85]. Peripheral [CBLM81, Kur84a, Sny88]. permanent [CZJY80]. permeabilities [EAL80, JA88]. permeability [HSd^{+}88, LL82, LL84, MUC81, MTM86a, TL88]. permeation [TMGB80]. permselection [Luk84]. permutation [Hua80a]. permutations [Rei84]. peroxidase [HFO83]. perrenials [CM87]. persistence [Agu85, LS89]. persistent [Vla87]. persons [LL87c]. perspectives [Bur82a, Nus87]. pertaining [NSJT81]. perturbation [Rab82, SF81, WdlT84, YA89]. perturbations [CG81a]. perturbed [RT88]. perturbing [DCR82]. pest [Car81b]. PGA [KS81, KK81, Kot82]. PGB [Kot82]. PGE [Kot82, Kot83]. PGF [KS81, KK81, Kot83]. pH [Koc86b, OK87b, VA88a]. phage [Ano81i, KRT81a, KRT81b, Len84, OK82, RR84, RR83a, RR83b]. phage-bacteria [RR83a, RR83b]. phage-bacterium [RR84]. phagocytes [RF82]. phagocytosis [You85]. pharmacokinetics [Bor82]. pharmacological [Ger83a, LH89, Mae84, Sto89c]. pharmacology [BT81, NS82]. Phase [DCR82, Leu87b, Win89a, Wis80, BW87, CM86, DW82, FF83, GVC81a, GV81, GVCC82, Hor89a, IYS83, JPGL86, MUK86, Ola82, PO81, PP85, SPP84, ST86b, SN85, SSN86, Wri80]. phenanthroline [Sto89b]. Phenetic [CB83, Ipi87, Pen82]. phenomena [Di 89, Ond80, RL83, Ros84a, Rot80, TT85, Wol85, YMS84]. phenomenological [CFLF88, LRS82, MM83a, RLS82]. phenomenon [Bou84, CF85, FT85, SSF80]. phenotype [Cre88b, Les86]. Phenotypic [NH84, Dom83, Mot82, Pet83]. phenylalkylamines [KM85]. pheromonal [HJS84]. Pheromone [FSM80, Omh88a]. pheromones [LL87b]. phloem [SMGD80]. Phonon [Lac84]. Phonon-electron [Lac84]. phosphate [KSH87, MHI85, MHT88]. phosphates [ZBL80a, ZBL80b]. phosphatidylcholines [Inf85]. phosphodiesterases [Fel80]. phosphofructo [KSH87]. phosphofructokinase [Tor80, WGKG83]. phosphoglycerate [War86]. Phospholipase [Sco84a]. Phospholipid [Nor89, RAR83, TL88]. phosphorylation [FJ85, Har80a, JS81, JF85, Kho84, Mal84b, Mal84c]. photographic [HT83]. Photometric [Lat83]. Photooxidation [Pai82]. photoperiodic [LS87, SL87a, SL87b]. photoperiodically [Tay85a]. photoperiodism [Fuk81]. photoperiods [Tay89a]. photophosphorylation [HT83]. Photoresistivity [IAHA80]. photosynthesis [Bru85a, Cha80a, JT84]. Photosynthetic [vGvGL88, BJ82, DT80, FBF85, FSF89, HN85, HT83, Jur84, KVB83, NS86, Tak89c, VI80]. photosystem [Shi81]. Phototautomerism [SM81c]. phototropism [FS87]. Phycomyces [FS87, PSG83]. Phyllotactic [GB87, Jea87, MZM89]. phyllotaxic [MZM89]. Phyllotaxis [SC84a, MK83, Rob84, Rob87]. Phylogenetic [LdJ86, Day83a, Jøn85c]. phylogenies [DS87, Lam84a]. Phylogeny [GFA89, Arm88a, BT84, Fou84]. Physarum [MUK86, MUK88, UK82]. Physical [GG82, Stu85b, Bol80, Bol82a, CL82, Faj82, HL86, HT83, Lum82, MDBRM87, Str85a, Str85b, Str85c, Tan84, Vol81, WB89, Wol89]. Physico [Min82, NBZ^{+}83, BSR^{+}88, MB84a, Mas84]. Physico-chemical [Min82, NBZ^{+}83, BSR^{+}88, MB84a, Mas84]. physio [RW82]. physio-chemical [RW82]. Physiological [DKW89, EKSS88, vS83, BLC86, Gho81, MW83b, Per84, Ros82, SS89a, SG88]. physiology [Goy82, NS82, PS81b]. phytochrome [HW80a]. phytoplankton [LRKC83, Law84, LJTH85]. Picking [Ols87]. picosecond [Mik84b]. pieces [Frö88]. pigeon [Wal89b, Wal89a]. pigeons [PW88, Yor81]. pigment [vGvGL88]. pigment-protein [vGvGL88]. pigmented [Cri81, McG85]. pigments [FSF89, vGvGL88]. pipe [Mäk86, Nev88, Val85]. pipe-cleaner [Nev88]. pipe-model [Val85]. piscivorous [HB88]. pit [dCB84]. pitching [Yat86]. pK [UN82]. place [GN86]. placental [Hof83]. planar [Eld80]. planci [McC87]. Planck [RCA89, Shi80]. planes [KDG85]. planktonic [MVV89]. Plant [Pah85b, Van84, Wel86, iWSS88, Ågr85, AL81, AW83, Amt86, AC89, BS87, Cou80, CM87, DDJ83, FW85, Fuk81, GW88, GJ88, Giv86, Gol83b, Gre82b, Har84a, Har84b, Hay80a, Hay80b, Hej82, HR84, Her83a, Her83b, HL82, Jol88, Klä83, Kno82, Kor82, MW81, MB84b, MB86, MW83b, Nev88, Pah85a, RS83, RQ81, Rob88, Rob89, Ros84a, SD81, SD83, Sto81, Tai86, TL86, VdV87, Wel83]. Plant-animal [Wel86, GJ88, Wel83]. plant-pollinator [SD81]. plantations [Gat82]. plants [Bre83, CC82a, CB86b, CF82, FEC87b, FEC87a, GR88, HW80a, Kha86, MW83b, PDC84, Rob85a, SG86, Wat80]. plants-Evidence [SG86]. plaque [AWV^{+}88]. plasma [Ber87, Bre80b, FF83, MG81, Nor82, Sco84a]. plasmacytomas [MKM80]. plasmalemma [HL86]. plasmid [Fut86, KP89a, KP89b, WB88, YI88]. plasmids [BLC86, vdHN85, vdH84]. plasmodesmatal [GRJ81, RJ80]. plasmodia [UK82]. plasmodium [MUK86, MUK88]. plastic [TG83]. plasticity [LD83a, NH84]. plastid [TH81]. plate [KSL89, MKL81]. Platelet [WG81, JK87, Vro83]. plausible [Pen80, Sib83]. play [Luf82]. plays [FT87]. Pleiotropic [Abu85a]. plot [BWM80, IHTN88, Rak80]. Plots [Law83, FLR83, GC83, Kub85, SM81b]. poikilotherm [LR83, SH80b]. Point [BPH87, CMN81, She81, SV84a, Top88, Whi89a]. points [BMR89]. Poisson [RCA89]. polar [Dal83, Kri80, Lew81c, Pap84, Pap89, RB82a, Sib81, Sib83]. polar-coordinate [Lew81c]. polarity [Hed85]. polarizability [RAR83]. polarization [Bri80, RC81]. polarized [PW88, vGvGL88]. polarly [SC84a]. polarly-translocated [SC84a]. pole [BB81b]. policy [Aok84]. pollen [GR88, Lew82, PS82b]. pollen-ovule [Lew82]. pollinated [BA80, GR88, Lew82]. pollination [Wel83, Wel86]. pollinator [CB86b, SD81, Wel86]. pollution [Phi81a, Phi82]. Poly [Car80, Sto83, TRS84]. poly-codon [TRS84]. polyamines [Goy82]. polycephalum [MUK86, MUK88, UK82]. polychlorinated [MP87]. polycyclic [IO83, MHB83, Sýk84]. Polyelectrolyte [BN81]. polyelectrolytes [MN89b]. polygonal [CM88]. Polymer [iWSS88, KP81]. polymerase [Arm88a, BJ84, Bee80, San84, TIK89b]. polymerases [Arm87a]. polymeric [Hum81, RN85, WP81]. polymerization [FC81]. polymers [Fer87, FDR83, ML82, ML84, RdL83, Ven82]. polymorphic [BF85, Cha89c]. Polymorphism [vTdJ86, Bre80a, Gre82b, Ste81, Too82, Tre81b, Yok81]. polymorphisms [Fra86, GZ86, ZG85]. polynomials [BWM80, BWW80, Bri84, Bri85, Whi81b]. polynucleotide [Dou81]. polynucleotides [Blu85]. polypeptide [Jea85]. polypeptides [Mat83]. polyploids [CG84]. polyploidy [MD86, Mog86, Ury81]. polyvalent [CU88]. pombe [CZJY80, Ham82]. pool [OWN89, WG81]. pools [Kre82, WI80]. poorly [Pea88d]. Population [Ada80, ASTT80, CM87, Get84, JP89a, KTS89, Rei81, Rog86, SS89b, Wel83, Aal89, Agu85, ADdACP87, And80a, BKMS86, BKMS87, BWM82, Cal83a, Can87, CM86, Ebe88, Ful81, GJG83, GPKBG89a, GPKBG89b, Gin86, GGC86, HN82, HKK82, HS81, HB84, Hin86, Hin87, Hir80, Hua80b, Ive88, Lei81, LS88, LS89, LA86, MW81, MW83b, Mur83, Nag82, Nam80, New80, New83, Omh88b, Rég84c, Rog89, Ros85, RB87, Ryt87, SFR86, Sch88, SM80a, SM80b, Smi88, TW80a, Vla87, Yok85, ZL84, dGV80, vdH84]. Population-dependent [KTS89]. Population-dynamical [Ada80]. Population-dynamics [ASTT80]. populations [Aug85, BF85, Car81b, CC81, Cha89c, Cus80, Ebe87, Fra86, GNG80, Giv86, Gre82b, Gre84b, HC81b, HdL84, Ham82, HINT85, Har84b, HS81, Hin80, Hir82, Hoe82, HvL86, IP83, JK88, Kel87, KK82b, Les86, Lib88, McN89, Moo88, NM80, NS89a, QL82, RT88, Rot83, Rot87, Sir86, SGd84, Ste81, SVN89, Wat80, WdlT84, Yok84, Yok85, dMP85, dCJM86, vS85]. pore [NA84]. Pores [SKF83, And81, FS83, KH80, LF86, SK88b, SF83]. porous [MHII83]. position [Bul88, Fra88, LvH87, SJK88]. position-effect [Fra88]. Positional [Kor82, CS87, KS84]. positions [Mur85]. Positive [Zol83, Bar80b, Gho81, KD85, Koc88, Koc89, Mah84a, Mah84b]. possibilities [Wil81c]. Possibility [RR84, Cra81b, Hir84b, Jam89c]. Possible [Ano81-37, CSH^{+}81, FB82a, GF81a, HG80, Hun81, KSW^{+}85, Kyp88, Sch85, Sim81b, Vro83, WB82, Bar83a, HT84, Hux86, Lac84, LN81, McG84, NDB80, Rob85a, Ros84c, Shu84, SW81b, SJ89, dSTSM83, TRS84, Tus88, Was86a, Wig82a, Wig82b, ZG82]. post [GDHB87, Har82b, Mac87, MG84a, RB82a]. post-mating [Mac87]. post-meiotic [RB82a]. post-transcriptional [GDHB87]. post-translational [MG84a]. potassium [CS84, CC83b, DLN85, JCK80, Str85b]. potency [ABS82]. Potential [DB88, JSJS88, Lef86, OK87b, SL80b, SBW83, ANS83, AP80, Bar81d, BP85, CW88, Cla82a, Col82, DS84, HGG82, HFL89, Jam82, KS81, MOK87, Men82, MBJK89, Mer87, NST80, OK87a, She80, She81, SR87b]. potentially [Hel80]. potentials [BW86b, Gol83b, GPM89, HBRN81, MKL81, PSL^{+}85, RCA89]. Power [NS82, AK80, GKB86, Hir82, Pah85a]. power-law [Hir82]. practical [Jea87, KB81]. Pre [Mur81, BHH83, BBH84, EPH86, GS86, HBA89, Nus88, SG89a]. pre- [SG89a]. pre-mRNA [HBA89, Nus88]. Pre-pattern [Mur81, BHH83]. pre-promoter [BBH84]. pre-S [EPH86]. pre-spore [GS86]. pre-stalk [GS86]. Prebiotic [Bur82a, Kin86, Ven82]. preceding [Hut89]. precipitin [NSW82]. precise [BE86, EC81]. precision [Cal83c]. precodons [Por82]. precopulatory [Yam87]. precursor [DH85, Nus80b, RTM88, dBH87]. precursor-product [RTM88]. precursors [Iid88]. predation [AK84b, Arm89, GW88, KLR89, Lim87]. Predator [McC87, Ton86, AG89, BS82, CPLA85, CSC^{+}80, HB88, Mat85, MKS^{+}86, PH87b, Ren82, WW84]. Predator-mediated [Ton86]. Predators [KR85b, HH81, Jan82, LG85, Spe86]. predict [Frö88, Nun81]. predictability [BB88a]. Predicted [AHFH81, KWG87, Whi80c]. Predicting [Pec88, SAS85, DB88, KD88, SM80b, SJK88, YA89]. Prediction [NMMO82, NHT88, BSB88a, BSB88b, Bro81a, CMM86, JD82, Man81a, Mur85, Vis87]. Predictions [HK81, YM89, HL89, Kat85, LRKC83, LJTH85, VZR85]. predictive [Gla87, Spe86, SW81b]. predominantly [BA80]. preference [CPLA85, Kon88, LA85]. preferences [TO89]. preferential [Kyp88, LvH87]. Preferred [Cla82f, TO89, VGP^{+}80]. pregnancy [Asp88]. Preliminary [BD80, MAC^{+}83]. preparation [CY84]. prepattern [HK88, SH85a]. prepatterns [Hun81, Hun85a, Nag89]. prephase [TL88]. prescription [Cri82]. presence [AM87, Gai80, GGPM80, Str80]. present [Cla82b, CB88b]. present-day [Cla82b]. presentation [Vol81]. presenting [Tho85]. pressure [GTH88, LR85b, Nic84, Rap84, RM86, Rld82, Rob84, Rob87, Spe81]. presteady [KL80]. Presumed [Fox81, Ros84b]. presynaptic [MP85b, PS82a]. preventing [LH80]. previous [GRJ81]. prey [AG89, BS82, CPLA85, COA85, CSC^{+}80, HB88, Hel80, HH81, Kil80, LG85, MKS^{+}86, McN87, PH87b, RC89, Ren82, Spe86, WW84]. prey-predator [CPLA85, MKS^{+}86, PH87b]. prigoginian [Ji88]. primarily [Blo87]. Primary [SH85a, VI80, Ber87, Bro83, CB89a, MN85, NM85, SB86, Wee84b]. primate [Nus88, Oli85]. primates [CS80, Rut83]. primeval [Dou81]. primitive [BS85, Dou84]. primitives [CH85]. primordia [Nag86]. principle [Bri89, Hav89, MF89, NH84, San88, SH81, SSF80]. Principles [Els81, GE82, Per82, Cha82a, Gin86, GEL80, Les86, MB84b, Mul87, SH87, WH86]. prion [RS87]. Prior [TC80a]. prisoner [Boy89, FT87]. Probabilistic [DPV83, IBJ^{+}89, MA83]. Probabilities [Mit89, Vil83]. Probability [Miy86, Nod82, AZE83, DW83, Hir83, Mur85, RB87, Sch83b, WBKW82]. probable [SW87]. probe [Car80, Kap85]. probes [BW89]. problem [Fou84, HKC83, Lau82, New80, New83, RR83a, RB82d, Sch83b, Smi84, VI80, WRW83, WEHM82]. Problems [Mae84, Dan82, Day83a, HCKB83, MDW80a, MDW80b, MDW80c, RB82d, Wel85]. procaryotic [Wri80]. procedure [BSB88a, Gie88, Got86]. procedures [ABEFS^{+}86]. process [BE81a, BPH87, Bre86, Cha88c, FM83, Gad81, Gai80, GT80, Gri81a, KO82, KK83a, MB88a, Orl82, Ren82, Sve81, VLCK86]. processes [Blo85, Cla81, CH82, FEC87a, Har82a, HR80b, HH85b, Lac84, LRS82, MB84b, MG84a, Peu86, Sar82, SB86, VI80, Wat85, Zac81, Zha84]. processing [CH85, GDHB87, HW82, MG84a, vHS82]. produce [Hoe80, MP87]. produced [MN85]. producing [Ser87]. Product [Pin83b, Bor88, Cla82d, KL80, ME82, Pin83a, RTM88, SPS83]. Production [vS85, Ågr85, Aok89, BW86a, Bia81, Bon82, BS84b, Car83, DSD81, FDS^{+}88, GAD80, HMF88, Hir84b, JJ82b, Jur84, dANG83, OL86, Omh88a, Spe87, SR89, TT87]. productivity [CW88, RS83]. profile [Eng89, HGG82, Ler83, MHII83, SLN84]. profile-curvature [Eng89]. profiles [Bar83a, KL80, LP86, MAGR87, MKM80]. profiling [Zuc86]. progeny [Jam86a]. progesterone [CY84]. prognostic [She83]. program [KM85, PSL^{+}85, Stu86, Uma82]. Programmed [Kuh82]. programmes [SP88]. programming [CN84]. programs [PLJ84]. Progress [Dug86]. progression [Bon85, KF87]. projectiles [Gor87]. projection [Ros84a]. Projections [RL83]. prokaryotes [Cla82c, MM82b]. prokaryotic [Arm87a, Nus85]. prolactin [AAAT87]. prolate [Reg84b]. proliferating [HRG82]. proliferation [ATC81, Bre82, PL87, Pro80, Pys80, Sve81]. proliferon [NP81, NP82, Zaj82]. proline [WMS81]. prolonged [EP86]. prolylhydroxylase [HAG82]. prominent [Wil84]. promoter [BBH84, Fra88, Lav84, OK82]. promotion [KF84]. proof [HT80, Siz87]. Proofreading [FS80, BE81b, Sav81, SL81]. proofs [Ros84b]. Propagation [KYU^{+}86, MUK86, BJL85, Ede82, GP80, Gli80, Kir80, MHB89]. properly [Cal87]. Properties [Day83b, MK83, McG81, MHB89, Asp88, BSR^{+}88, DRL88, EM81, ELM81, KT83, Kaw84, KK87a, KP85, KM87, MR89, MK88, MR86, McG84, MCB^{+}86, MATGR87, MDBRM87, Min82, RM86, RHH85, SS89a, Ste87, SJ89, TLS83, Wei88, Whi89a, iWSS88, dVR85, vGvGL88]. proportion [Mac82, PO86]. proportion-regulating [Mac82]. proposal [Jam89a, TKB80, Wol85]. proposals [She81]. proposed [BGG84, Con83, Cra81a, Dou81, Dou83, KF84, Mat84, RF82]. propulsion [Yat86]. Prospectives [Nus87]. prostaglandins [KS81, KK81, Kot82, KK83c]. prostaglandins-IV. [Kot82]. prostaglands [Kot83]. proteases [GBS89, LSFT85, SLYF84, SLFT86]. protectedness [Gre84a, Gre82b]. protection [Yok85]. Protein [Arg81, Bri88, Cha81, DSW83, GT88b, HR85, RB83a, Sve81, Whi80c, ZR84, ANS83, BRRK87, BJL85, Blo87, Bri84, Bri85, Bry80, Cze87, DO83, DML80, Duf85, Duf86, DB88, FY81, FCBK88, Frö88, Hal80b, IHTN88, JC80, KFPB85, KS80a, KA89, KS88, Kri80, KTNO81, Lan82b, LdJ86, LB87, LSFT85, Mad88, MG81, Mic86, MSBB84, MM86b, MG84a, Mur85, Mur87, MKM80, Ond80, Rak80, Rak86, RS87, RR83a, RR83b, RB83b, Row85b, Sar85, Shu84, Sim81b, Sim85, Spe84, VZR85, War86, Wes83a, WI80, Whe84, Whi82, Zuc86, vGvGL88, vHU87, vHBL87]. protein-coding [Row85b]. protein-deficient [BRRK87]. Protein-ligand [Bri88, Bri84, Bri85]. Protein-protein [HR85, ZR84]. protein-solvent [War86]. proteinases [ANS83, PA86]. proteinligand [RC81]. Proteins [Sim86, Bab88, BSB88a, BSB88b, Bro80a, Bus82, CH80, CFLF88, CH82, Cra81b, Cur88, Dan86, DO83, DC87a, EPH86, For85, GRS82, GK80, Gra86b, GD81, Hir84b, HM80, Kau86, KN82, Kri85a, Kur84a, LS82, LE84b, MG84a, PP80, Pra84, RR84, SA80, TDA86, Vis87, WJT87, WGHY82, Whe84, Whi80b, Wil86b, ZR84, ZR85, vHU87]. proteins-the [ZR85]. proteolysis [Whe84]. proto [Mer87]. proto-oncogene [Mer87]. protocell [FM83]. Proton [LM80, DT80, Dun82, LSFT85, Mal84c, NU82, Res82, SLYF84, Tai86, TCY85, GPLC84]. proton-electron [Res82]. protonation [MT84]. protonic [She80, She81]. protonmotive [Koc86b, Mal81b]. protoplasm [Fox81]. protoplasmic [Hay80a, Hay80b, TY82]. protozoa [Cze87, KTS89, SC82]. provide [FM88, Law83, Orl81a]. provides [PL87, SM88b]. provisioning [Tay89b]. Pseudo [GK80, WBKW82]. Pseudo-oligomer [GK80]. pseudo-steady [WBKW82]. Pseudochemotaxis [Lap80b]. pseudomurein [LBBL84]. pseudopotential [BLR83]. pseudopregnancy [AAAT87]. pulmonary [Lef83]. pulse [IP83, Sch83b, TLS83]. pulse-cytophotometry [Sch83b]. Pump [Ped88, Cha82a]. pumps [Dun82]. purine [AM87, CB88b]. purine-pyrimidine [AM87, CB88b]. purity [Faj82]. Purkinje [CC83b]. purposes [SPP84]. Putting [MG85]. pyrene [ITK88]. pyridine [ELM81]. pyrimidine [AM87, CB88b]. pyruvate [GFA89, Tor80].

QSAR [YHJ89]. Qualitative [GH82, AK82, FM88, HdL84]. qualitatively [BHMR85]. quality [Mad88, MH80d]. quantal [LPS86, Miy86, RG87, Vau86]. Quantification [MCB^{+}86, LH89]. quantify [Mal84a]. quantitation [SLN84]. Quantitative [Bee80, Che84b, DE82, HH85a, KK86a, WWN82, Zuc86, Abu88, Aok84, Cha81, CB87a, Fle88, Fuk81, Hol83b, KM85, KP81, Mal85, MHB83, MOV82, MV88b, RW82, Rob85a, Sou80]. Quantization [DRL88]. Quantum [MHLK87, YLBB81, YBBL82a, YBBL82b, YBBL82c, BvdHB84, GG84a, Wol89]. quartets [Nus88]. quasi [Bar80a, LL87a]. quasi-steady [LL87a]. quasi-steady-state [Bar80a]. quaternary [BN81]. queen [Bul81]. queens [WL87]. question [Has86]. Questions [Ito82]. quiescent [BH83, Rot82b]. quotients [GJ88].

R [Ano81i, BC81a, GWK87]. R. [AK86]. rabbit [Bor82]. rabies [KAM85]. racemic [Gra85a]. races [Par83a]. radar [Dra81]. radial [SEN82, Wil81a]. radiation [Her83a, Her83b, HT89b, LBW86, PJK^{+}83]. radiation-induced [HT89b]. radiations [Lam88]. Radii [Fis85, SPKJ89]. radioactive [FCBK88]. radioresistant [Suc83, Suc85a, Suc85b]. radiosensitive [Suc83, Suc85a, Suc85b]. rafts [SG82]. rakes [CK87]. ramet [Ing89]. Random [BS84a, Fed80, BC86a, BB88b, CK82, GZ86, Gri81c, Gri82b, Gri82c, Hsu88, JJ82a, Lap81b, MJ88, MGH81, SS88, TJP83, TH85, VM82, Vil83, YM89, dW81b]. randomization [Nee89a]. randomly [Hin80, Kam80]. Randomness [Agu85]. range [KK81, PCTF83, War86]. ranges [Ben89]. Ranging [Dus89b]. Rapid [Fis80, Sch84a, VdV87, VI80, WGHY82]. rapidly [Nan86]. rare [Hir84b, Moy82b]. Rareness [Nod82]. rat [Aki85, AAAT87, Bas81, CMM86, IHK80, VKV83]. ratchet [NS89a]. Rate [FS83, Kon84a, LR85b, MdMF83, SF83, SKF83, AFHM86, Bar83a, BF88, Cho81, Dem81, Don86, Dur82, Fil88, Fra84, Frö88, GK85, Gra81b, GM87, Hsu87b, Hsu87c, JD82, JK87, Kel87, KF84, KC83a, KD80, LIWK86, Li83, LvH87, Man81a, MI81, NSKS85, RP84, Rei86b, RN84c, Rob88, Rob89, SPS83, SSM81, Shi82, Sil89, VL86, Wag89, WBKW82, Wee84a, WP84]. rate-limiting [WP84]. rates [AHFH81, Car83, Dur82, Hal80b, Hej82, HR86, Hir80, Hir82, JK89, Kar88a, KD88, MD86, Pea88a, Pea88c, SE80, SM81b, SPG87, Whi80a]. Ratio [AG89, BBH86, BM88, Ber87, BT80, Bul83, CB89a, Cra80b, FS88, Gre82a, HM85b, Iwa81, Jam85, Jam86b, Jam88, Jam89c, Lew82, LK83, Orl81b, Rei87, SL80a, Sto89c, TB80, Tor82, Wer87, vTdJ86]. Ratio-Dependence [AG89]. ratio-linear [Sto89c]. rationale [BS83, GH82, MSBB84]. ratios [Bog87, Cla82d, Cla82c, Fin82, Gra86a, Hor80, Jam87, Jam89b, Llo83, LL87c, Mit88, Sil84]. Ratner [Ano81i]. RBE [Cra81a]. rDNA [WS85a]. Re [Kap85, BH81a, Cra83a, HW80a]. Re-evaluation [Kap85, HW80a]. re-examination [BH81a]. re-examined [Cra83a]. reactants [NSJT81]. Reaction [SH80b, BKJT80, DfP87, DDA85, För88, For80, Gol83a, Hal80b, Har87b, HK88, Hua80a, Hun85a, KK85, Kin82, Kre82, Mal88, MM86a, Met82, MN85, Mur82, NM82, NM85, Nag86, NFP88, NSW82, PPL85, PV83, RN84c, SSM81, SH87, Sen88, SCH84b, SM81b, SEN82, SN85, TS88a, WT87]. reaction-electrodiffusion [Mal88]. reaction-rate [SSM81]. reactions [Bar82b, Bar81e, Cho81, CU88, Dug86, Gal84b, GVC81b, GV81, Kri80, Kri85a, Kri85b, KP81, ME82, Ped88, PHTCS86, Rak86, RN84a, RN84b, Ric87, SF81, SW87, Tai86, Vro83, Wag87]. reactive [GA80]. reactivity [Hua80a, IO83, MH80c]. reactors [CT81a]. Reading [KM89]. readthrough [EKSS88]. reagent [Sto89b]. real [KC85]. realization [Nod82]. realizing [OKT^{+}80]. really [Eps86, Hau82]. rear [Omh87b]. rearrangement [GZWR80, WGRZ80]. reason [Kyp88]. reasonable [Bar83a]. Reassessment [Cze87]. Reassociation [GMTG80]. Rebuttal [Yoc81a]. Reception [Gla80]. receptive [Pin84, YM89]. receptivity [Yam87]. receptor [BRW89, Bec88, Blu84, Ger83b, GPM89, Hum81, Inc83a, Inc83b, LL88, Mar84, Mik84b, RT85, Ros84c, SGDK86, SEN82, TS85, Was86a, Zol83]. receptor-coated [SEN82]. receptor-coupling [Blu84]. receptor/ligand [LL88]. receptors [Ano81-37, Ger83a, Ger85, GK80, LN81, MM80, MB85, SH85b, Sta86]. Reciprocal [Rot80, Nee89b, Peu86]. reciprocation [BSM82]. reciprocity [BR88, NS89b]. recoding [BH81b]. recognition [Alm85, BB81d, Get81, Get82, KK86a, Lam84a, LRS82, MM82c, RN84a, Rob85b, Sar89, Ses82, Tho85, Wal87, ZR84]. recognizable [MC89]. recognized [PDCD88]. recombination [ADdACP87, BS87, Bul80, CB81, HRG82, Hop86, McG84, McG89a, TS80]. recombinational [WB88]. Reconsideration [PB81]. reconsidered [Lim88]. reconstruction [BB81b, Kam80]. record [Mic88]. recorded [KMSR84, VG80]. recording [SM87]. recordings [Ste87]. records [CZJY80, KSW^{+}85]. recruitment [Jaf80, Jaf84]. rectangular [Sil83]. rectangular-mesh [Sil83]. Recurrent [OO84b]. recycling [Bec88, Ros84c].

Red [ABB89, BKB80a, BKB80b, Eng89, JP89b, JP89c, KIO88, KM87, Lan84a, MK88, MHLK87, PTD81, SOLG82, Sve82, Wol80]. redistribution [VI80]. redox [FQ81, Jam80, Jam82]. reducing [RS83]. reductase [MB82]. reduction [SM88a, SGH85, Smi82b]. redundancy [Duf83, Duf85, Duf86]. Redundant [SM88a]. reef [AK84b, BHMR85]. reference [FGDS87, HR86, PW88, SF83]. refinement [PRH81, PR82]. reflection [And81]. reflex [Mon88]. refolding [Sim85]. refractoriness [AP80]. refuges [McN87]. regard [LB87]. regenerating [BR86]. Regeneration [Bar89, Bar82a, Dea80, Lew81c, Mac82, MG80, Pap84, Sla80, TT85]. regenerativity [AP80]. regimes [DG85, Ing89]. region [AVR82, BBH84, Eld80, GB87, MDBRM87, NMMO82, Pat84, WGP^{+}88, vHU87]. regions [Alm85, BJ84, EPH86, Ipi87, Lav84, MDPF86, Mic86, Syv84]. regression [BM87, Dug82, Mae84, SSM81, SS83b, SC89]. regular [Sai82, Ste81]. regulated [DG87, FBP87, GT85a, HTG89, NL86]. regulating [KS82, Mac82, Whe84]. Regulation [GAD80, Ada80, AC89, Bee80, BW86b, BPP88, Car81b, CV89, CF82, Dou84, Fel80, Fer83, Fuk83, Fuk84, GDHB87, GS85, GS86, GH82, JK89, Joh84, KLK^{+}88, KL83, Lav84, LD83b, Lum82, MM82a, Man81b, McC87, Nes84, Pat84, PO86, PGKE88, PY88, RMS81, Sch82b, Sib81, Tor80, Web83, Wee81, Wee84a]. regulations [PGKE87]. regulator [Mar84]. regulatory [Kin83, Kin85, MM86a, Mer87, Nus85, RN82]. reinforced [Asp88]. rejection [Kel87]. Rejoinder [Hon86]. related [BRRK87, Gho82, GK80, Gri81b, Gri81d, Kat81, Mad88, MK83, SBW83, Tre81a, Wol85]. relatedness [Sou80]. relating [ABS82, BVDA87, HCW80, WS84]. Relation [WH86, AM81a, BRR87, Cha80d, GK85, Hol83b, Jam82, Joh84, Lim87, MUK86, MF81, SYM86, SA82, SOLG82, Whi81b, Wol85, vZvdSdL81, HJS84]. Relations [IM85, Sve82, CS80, Cha80c, Di 88, Dur82, Koo86, LIWK86, Moy82a, Por82, Rat87, Rob88, SL81]. Relationship [MATGR87, Ste87, AWMK85, Bar86, Bri85, BBH84, Fle80, Her83b, KHU^{+}83, Mog86, MRR80, PK85, SA86, Shi87, Van84]. Relationships [BWM80, McG80, Omh88b, BB81d, Fra84, HG80, KM85, KK86a, Mal84a, MM83b, RTM88, Sch81, Smi82c, Vro83, Wag89, WS81]. relative [AL83, CHT80, EL83, Koc89, KK83c, SK88b, Smi84, YA89]. relatively [Rei85b]. relatives [Aok84, Mot88, MG84b]. relativity [Nor83, SH81]. Relaxation [Pow85, MGL81, KK82a]. relaxational [Lev80b]. relay [BS88b, LSFT85]. release [AAAT87, Inc83b, LS81, LPS86, LPS89, Miy86, PS80a, PS81a, PS82a, PS84, PPS86]. relevance [Alf83, LBW86]. Reliability [Bre82, Her81a, Mil89]. relict [HB86b]. remain [NS86]. remark [Reg84a]. remarks [Kon88, MHT88]. remnants [Bro80a]. renal [WMS81]. renegade [Har82c]. renewing [IP83]. renin [Wee84a]. reorganization [KK85, LFP84]. repair [COR83a, HT89b, PL84b]. repeat [Bro83, RMFM85]. Repeated [PM80, Boy89, DLN85, FT87]. repeats [GMTG80, dW81b]. repellent [Lap80a]. repetitious [HRG82]. repetitive [BJ84, Bro82a, BBH84, Koc86a]. replace [EGS80]. Replaceability [CV81]. Replicating [Ste82, Fer87, SG89b]. replication [Bel80, Ber89, BC81b, BC83, BLC86, BS87, Dur82, Fer88, KP89a, KP89b, Luf82, MK81, MZ89, MRG84, MM82b, MdMF83, NSCH88, NS89a, RS87, RB83a, SCP83, SP88, Tak87, Tak89a, Tak89b, Var83, Wil81a, YI88]. replicative [CK85]. Replicator [SS83a]. replicators [SD87]. replicons [MdMF83]. Reply [Ano80-35, Cha82b, Lib85, Vic88, Wil83, And86, BB81a, Coo87, HLS89, Rog89]. represent [Esh84a]. representable [Hyv80]. representation [AL83, ALST81, Bus82, CP82, Che83, Che84a, EG87a, Jea85, LS82, PS80b, Row85b, TRS84]. representations [RL83]. repressible [SR88]. repression [PB81, Tys83a]. reproducing [Gre84b, Har85, HS81]. reproduction [Cas82, Cas84, CT81a, Gra88, MD86, Mog86, Rei85b, Ryt87, SC82, dSTSM83]. Reproductive [ER88, LC82, BP85, HAC82, Mac87, YE82]. reptation [McC84]. reptile [Ipi87]. reptiles [DY82, SG88]. repulsion [OK87a, PYN85, RN84a]. required [SM88b]. requirements [GM82a, Kri85b]. research [HD84, Jam89a]. resemblances [Ipi87]. resembling [Bec80]. resides [PA86]. residual [Bus82]. residue [GRS82, MP87]. residues [AL83, HM80, HR85, NU82, PP80, SH80a, Vis87]. resistance [Aki84, Cru80, Jam80, SJ85]. resistivity [Kaw84, Sto89a]. resolution [Lat83, TJP83]. resolved [DG85]. resonance [Fer83, MS89a, Poh81]. Resource [LS88, Rog89, EL87, GC86, Get84, Ive88, LH86, LS89, MG85, Pug87a, RC89, SG86, Sil84]. resources [EN86]. respect [Blo85]. respiration [Amt86, GK85]. respiratory [CLFL86, DDBB83, SBC80, Sau80]. respiring [BT85]. respirometry [FBB89]. response [BM87, Bec81, BE86, BH83, CLF84, Dan86, Dav87, DCR82, FDC^{+}89, GAD80, Har82c, IYS83, KW89, KUT85, KT87a, KKNM80, KK80, Lan82a, LL82, LL84, MM83b, MAC89, Poz87, SL87a, SLN84, SD83, Sto89a, Sto84a, Sto89c, Szu83, TW80b, WFL86, WC88, Wil81b, Zol83, dGV80]. Responses [Tho86, AMI84, Der87, FS87, KKKK80, KM83, Lap81b, MUK86, PSG83, RG87, SL87b]. responsible [JS81, LCDC84, Sim85]. Resting [EP86, BESD84, Hir83, Hir84a]. restricted [Bro80a, Gad81]. restriction [Kar88a, Kar88b, MM82c, Tre81b]. restructuring [LKC^{+}87]. result [Kle81, Mor82, Mur85, vDvD86]. resulting [BP81b, SH87, Smi82c]. results [BWW80, CB89b, ICL85, Kat85, RC81, TJP83, Wee84b]. retaining [TH81]. retaliatory [Gor87]. retention [Tho85]. Rethinking [Smi80, Har82d, SS83b]. reticulocytest [HR80b]. reticulum [AH86, ACB81, Wig82a]. retina [Ham85, SCH84b, Sim80a, Sim80b]. retinoids [Bar89]. retraction [Har80b]. retrieval [Rob85b, SPST83]. retroposons [Ber89]. retrotransposons [HB86b]. retroviral [EPH86]. retrovirus [BJ84]. return [Bar87]. revealed [Zac81]. reverberation [KK83b]. reverberatory [BH81b]. Reversal [MUK88, Aki84, AP80, Mas86, SL80a]. reverse [AWV^{+}88, Cra81b]. reversed [KLK^{+}88]. reversibility [Gre80]. reversible [AK87b, Fox81, KP81, KD80, ME82, Sen88, WWN82]. revised [KRB81]. Revision [LQC89]. revisited [DR89, GM82b, Jea85, Kwo89]. revived [Lum82]. Reviving [WS89]. rewards [BC86a]. Reynolds [Li88]. Rh [Yok81]. Rhizobium [SS89b]. Rhizobium-legume [SS89b]. Rhizopus [HCW89]. rhythm [KFSW82, PS80b, UK82, WF82]. Rhythmic [Zha84, TSRK88]. rhythmicity [Pai82]. rhythms [ATC81, CG85a, CB86c, Nun81, Pro84]. ribbing [Gun89]. ribonucleic [Mik81, Mik83]. ribonucleoprotein [CS82]. Ribose [Szu83, Sto83]. ribosomal [NHT88, Sek85, Wil86b]. ribosome [Fel83]. ribosomes [Bah80]. ribotype [Bar81a]. rich [Cla82f, Cla83b, Van82]. ride [Orl81a]. rigid [And81, FS83, SF83, SKF83]. rigorous [Wei88]. Risk [Bul84, JGW83, KLR89, Lim87, Reg84a, She83, SCB81]. risk-sensitive [Reg84a]. Rissa [DC87b]. Ritualized [LH83]. RNA [Arm87a, Arm88a, AVR82, BJ84, Bee80, Cha88a, Fer88, Fuj82, Gar80, GDHB87, JC80, LM89, Mal81a, Mil87, NHT88, NDB80, Nee89c, NST80, Nus80b, Org86, Rea84, San84, Ste82, vHS82]. RNAs [YI88]. RNY [Cla82b]. rod [CDC88, GZWR80, WGRZ80]. rod-shaped [CDC88, GZWR80, WGRZ80]. rods [KD85, Koc88, Koc89]. Roger [GT88a]. Role [ANOM82, BN81, BPM82, Iwa81, OK87a, You85, Bir81, BSS83, Bou84, BP85, BvdHB84, DCF80, Don86, Dun82, Dus89a, Faj82, FT85, GTH88, GW83, Gor87, Goy82, Gra85b, GG84a, GM87, HE80a, HF88, Hut89, KF84, KKJ81, Kur86, LC80, Lil82, Luf82, MPPF80, NM82, NM85, Nag86, NMM84, Nic84, NBZ^{+}83, OO84a, Rap84, Ros84c, San84, Sco84b, Sek85, Tho85, TKB80, Ven82, WFL86, Wil84, ZG82, WSMP83]. root [BB86, Bru85a, CCG^{+}82, Pla82, Rob88]. rotary [AY89, Fue80]. rotating [IPS83, Obe81]. rotation [Her83a, JPGL86, Koc89]. rotational [MYS89, MS86b]. Rote [TY82]. rotors [Win89a]. roughness [MCB^{+}86]. rouleau [Tus88]. Rouleaux [FF83, KIO88]. route [RL88]. Routes [LP88b]. RQ [SR87b]. rRNA [MDPF86]. rugged [KL87, KW89]. rule [HT84, Nee89b, OC81, Pic84, Sea80, Shc89a, Whi81a]. rules [BL85, Cho81, CM88, EL83, Gol88, Har87a, HS87, KCM81, Lew81c]. rumen [FGDS87]. Rumer [Shc89a]. ruminant [Cze87]. ruminants [FBP87, FDS^{+}88]. ryegrass [CW88].

S [Kot83, SPP84, EPH86, RMFM85, Wri80]. S-phase [SPP84]. S37 [Dug82]. saccharide [YBBL82c]. saccharides [YBBL82a]. Saccharomyces [BW87, Fut86, WB88]. Sacrificing [BB88a]. safe [Gra81a]. safe-neighborhood [Gra81a]. salamanders [TT87]. saliva [Izu81]. Salmonella [Cla82a, CM82]. salmonid [SH82a]. samara [AO87]. same [Niv83]. Sample [HCKB83]. samples [HG87]. sampling [OO84b]. sanguinipes [HLS85]. sapiens [AS87]. sarcomere [Ere85]. Sarcophaga [PS80b]. sarcoplasmic [AH86, Wig82a]. saturable [Pic86]. saturated [KH80]. Saturation [CB88a, Cha80a, DBB80, KH80]. savings [BH81a, Fou84]. scalar [Hej82]. Scale [CH85, Met82, PV83]. Scaling [Mau87, Pro84, Pro86, RL84, Tay88b, Whi89b]. scanning [PPC82]. scar [Ham82]. scar-class [Ham82]. scars [CZJY80]. Scatchard [IHTN88]. scattering [Fis85, Lan82b, Whi82]. scenarios [Yoc81b]. schema [Rob81, Rob85b]. schematic [Cho81]. schemes [Gol83a]. schizophrenia [KRB81, Was86a]. schizosaccharomyces [CZJY80, Ham82]. schmoluchowski [KSS81]. science [HHP89]. scientists [Ros84b]. scolymantha [Hun81]. Scott [HLS89]. Scrapie [Wil86b]. screening [Kri80, Rea84]. screwing [McC84]. sea [BH83, Rik85]. sealed [Ash85, BN80]. search [BPH87, DD85a, DR89, Kil80, Spe86]. Searching [ZR85, KR85b, Nus83]. Seasonal [PC83, SCS88]. Seasonality [AS84]. seasonally [Nam84]. second [Els83]. Secondary [Sab80, BSB88a, BSB88b, Duf85, Lav84, Lil82, MF85, RMFM85, SH85a, SJK88, Wee84b]. secreting [AWV^{+}88, GAD80]. secretion [BS88b, CH86, Wee84a]. secretory [Hol83b, Win89b]. Section [Eng89]. sedentary [Spe86]. sediment [Car83]. sedimentation [HT87, LP86]. Sedziwy [Rog89]. seebeck [SI81]. seed [Arm89, Ell86, Que83, Tay89b]. seedbank [MW81]. seedlings [Rob87]. seeds [AY89]. seeing [HL84, Mic88]. seen [Mik84b]. segment [MM83d]. segmentation [Dea80]. segments [IAHA80, Swa82]. segregation [BS87, CS87, JJ82a, LA86, Men85, Nag81, NSCH88, RB82a, Wri80]. select [Bah80]. selected [Bro80a, TIK89b, YA89]. Selection [DG85, Dem83, Gre82b, AM81a, Abu85a, Abu86, And86, Asm83, Blo87, BR86, Car85, CC81, Che84b, Col82, CK82, Cra82a, Cra83a, Cra83b, Cre88a, Cre88b, Dil86, Dom83, DB88, ES85, GJG83, GT88a, Gla87, Gre84a, Gri81a, Gri81b, Gri81c, Gri81d, Gri82b, Gri82c, Gri82d, Gri82e, Gri82a, Gri82f, HR87, Has88, HS81, Ish81, IT80, Jac82, Kir86, KK86b, LK88, Lim88, Mac87, Pom87, Que83, Rog83, Rot80, Sac88, SH82a, SG86, Smi85b, Ste84a, SD87, TO89, TACM82, Yok84, Yok85]. Selective [RB82e, CH82, Gro82, LL87b, MG81, SC83, TMGB80, TH81]. selectivity [NA84, Pin83b, TIK89a]. Self [FM83, Hic89, RS82, San84, Yoc81b, ABEFS^{+}86, Bel84, BC86b, BA80, CV81, Dou84, Fer87, For85, Giv86, HS81, IHTN88, KsC85, LO81, Man81c, MF80, MYS89, NSW82, NSJW84, NW85, Per84, Pic84, Sin82, TS88a, Whe84, Whi81a, dBH87]. self-assembly [KsC85]. self-associating [IHTN88]. self-association [NSJW84, NW85]. self-electrophoresis [LO81]. self-facilitation [CV81]. self-fertilization [Man81c]. self-interaction [NSW82]. Self-organization [FM83, RS82, ABEFS^{+}86, Bel84, MYS89]. self-organized [TS88a]. self-organizing [Sin82]. self-pollinated [BA80]. self-regulating [Whe84]. self-regulation [Dou84]. self-replicating [Fer87]. self-reproducing [HS81]. self-sorting [MF80]. Self-splicing [Hic89]. self-steered [BC86b]. self-thinning [Giv86, Per84, Pic84, Whi81a]. Self-transcription [San84]. self/non [For85]. self/non-self [For85]. Selfing [Lew81b]. Selfish [Blu83]. selling [SBPF83]. Semantic [Dix83, Rei83b]. Semi [Dav84, Dav87, Mag86]. Semi-meiosis [Dav84, Dav87, Mag86]. semicircular [Ber82, MV88a, MV88b]. seminal [Ber87]. semirufus [IAHA80]. semisynthesis [Pet82]. senescence [Gle84, Hir80, Hir82, JLS80, JS82, PDC84]. sensation [Dus88, Sny88]. sense [Fra88, NM81]. sensing [JR80]. sensitisation [RW82]. sensitive [dCB84, Cla83a, Reg84a, Tay85a]. Sensitivity [KC83a, MB80, Mil87, RP84, Yor81, Coo84b, KC82, Lan82a, LR85b, NM81]. Sensori [AA85]. Sensori-motor [AA85]. Sensory [Gen88, Hon84, Lib85, MT81, SGDK86]. separable [Rot80]. separate [KvDVH87, Row88]. separated [Gre84a]. separating [FF83]. separation [BF83, DT80, SK87]. separation-induced [SK87]. septation [Nag81, NSCH88]. September [Ano80o, Ano80u, Ano81q, Ano81-35, Ano82k, Ano82q, Ano83l, Ano83r, Ano84m, Ano84s, Ano85n, Ano85v, Ano86o, Ano86u, Ano87s, Ano87z, Ano88v, Ano88-28, Ano89r, Ano89x]. sequel [Orl82]. Sequence [FLNF86, MDBRM87, Aid88, Alt89, BS88a, Bro80a, Bro83, Gho82, GPPZ80, KB81, MG80, Nus80b, Nus83, OK82, RMFM85, SBM89, TDA86, VZR85, Wat84]. sequence-dependency [Aid88]. sequences [Alm83, Alm85, AVR82, Ber89, Bro82a, BBH84, For81, Frö88, Gar80, Got86, KTNO81, LVN87, LdJ86, MDPF86, MM82b, Mic86, MM83c, MMG84, NDB80, ND82, Nus80a, NP84b, Nus86, PPL85, RH85, Row85b, Sar85, SPST83, SSW81, Tak89b, TRS84, VZR85, Wu88, ZC89, dW81b]. sequencing [Gal83]. Sequential [RC89, TATS81, ASTT80, BGGV83, Fed80, SF81, TJP83]. Sequential-encounter [RC89]. sequentially [BC80]. serial [Sla80]. series [BK88, DR89, For81, KMSR84, Pet83]. serine [ANS83, LSFT85, PA86, SLYF84, SLFT86]. serotonin [Ano81-37, RB83b]. seshadri [Met82]. Set [CMN81, CB88a, Gad81]. Set-point [CMN81]. sets [Kau86, Ste84a]. setting [ASTT80]. settlement [TH80]. Several [Lev80a, GVP^{+}80, Gri82e]. Sex [BBHM85, Bog87, Bre80a, BT80, Bul83, CC82a, CB85, CB86b, Col82, Cra80b, FS88, GC86, Jam85, Man82b, Mit88, Mit89, MGH81, Wil81c, vTdJ86, Abu88, BBHM84, Ber87, Blu83, Bul80, Bul82, Cha88a, Cha80d, Cha82b, Cha86, CB89a, Gra86a, Gre82a, Gri83, Hoe82, HvL86, Iwa81, JL88, Jam86b, Jam86a, Jam87, Jam88, Jam89b, Jam89c, KR86, Lew82, Llo83, LL87b, LL87c, MJ80, Man82a, Man84, Nee89c, Orl81b, Rei87, Ros82, Ros83, SL80a, Sil84, TB80, Tay85b, TH81, Too82, Tor82, TATS81, Wer87, Wie81, YE82, Yok84, You81]. sex-linked [YE82]. sex-specific [Abu88]. sexes [Par82]. Sexual [And86, BB81c, Dom83, Kir86, Rei86c, Smi85b, Cha80e, ES85, Gra88, HB84, Hin86, Hin87, LA85, LK88, Les86, Pom87, SS82, TO89]. sexually [Cha89c, Gre84b]. SH [WJT87]. Shape [Luk84, PYN85, Boo84a, Der87, Eil88, Eng89, Hut89, ICL85, MM83d, MHB83, Nun80, SG82, SOLG82, TL88, TS84, VFS80, WJS89, YHJ89]. shaped [CDC88, GZWR80, WGRZ80]. shapes [SL87a]. sharing [CB86a]. shear [Eva80b]. sheathed [Fue80]. sheep [LC82]. sheet [BRM82, HE80b, NSKS85]. shell [Hut89]. shells [MK87]. shift [GZWR80, SS88, WGRZ80]. shift-up [GZWR80, WGRZ80]. shifts [BR87, GPLC84, RB87]. shock [CFLF88, For85, HSV84, TKB80]. shoot [HR86]. shore [FDC85]. Shorrocks [PH89]. Short [MDW80a, MDW80b, MDW80c, AA83, GPLC84, Gra88, KK81, Man82a, Met82, Yam87, WGP^{+}88]. short-term [AA83, Man82a]. Short-time [WGP^{+}88]. Should [McC84, Mot89]. shown [Shi80]. shrinkage [Lat83]. shrinkage-the [Lat83]. shunt [Poz87]. shunts [Pop82]. shuttle [Jam82]. Sib [Bul80, You81]. siblicide [DC87b]. Sibling [BP86]. siblings [Cra82a, LL87c]. side [CC83a, Els82]. signal [DO83, Kap85, Mag82]. signaling [PO81]. signals [GM82a, Gun89, Iid88, KMSR84, RN84a, RN84b, Ric87, RG87, VG80]. Significance [CB88b, GPAL88, RGR82, WGHY82, Duf83, ES85, Gho81, GZ86, Gun89, HT84, LM89, Li83, MF85, Tre81b, Ury81, YMS84, Zac81, ZG85]. significant [Kon84b, Luf82]. signs [Ewa80]. silent [Fil88]. Similarities [EPH86, Sco82]. similarity [ÅF80, Eco82, Eco83, Eco84, GM82b, GM84, Hok86a, JLS83, MKM80, NDB80, RL84, SGJ86]. Simple [Bar83a, HT80, KC85, LE84b, Vol81, AH86, BPP^{+}86, Boy83, BLC84, DG87, Gat86, GF81b, HC87, HL86, HT89b, Jan82, JF85, KAM85, KP89a, Nun81, Pow88, SC84a, SM80b, SK84, Sou80, Spa80, Ste84b, Tod82, Tre81a, TG80, Vic87, Yag85, PJL84]. Simpler [Lew81c]. simplest [BF83, LG89a]. simplex [RMFM85]. simplicity [RN82]. Simplified [DH85, Jol89, SPS83]. Simulated [Bre81, Wal89b, Wal89a, Bor84, BC86b, Boy83]. simulating [AH86]. Simulation [ADdACP87, CC83b, Cra80b, DSR89, DDBB83, KYU^{+}86, KT81, PL85c, TG84, TJP83, UKNS87, BIS^{+}80, BF88, DDA85, FM88, FEC87a, HO84, HYDS84, Hsu87a, Hum81, Kir86, KS82, LC82, LF86, MF80, MRB83, MTM86a, MTM86b, PO81, Pok82, Rld82, VV82, VLCK86, WB89, Wol80]. simulations [EG87a, EG87b, KD80, Mac82, Sal84, Ste88, SL87a]. Simultaneous [Has81, MT81, AHFH81, SHBS89]. Singing [HM87]. Single [KSS81, SLYF84, Bar80a, BE81a, Cha88c, CCG87, Cra80a, Cre88b, DW81a, DV83, Fra82, IP83, KSW^{+}85, KH80, MM83a, Man82b, Mei82, MSS81, NSCH88, Nun80, PL85a, PL85b, PG88, QL82, Sir86, SHBS89, Ste87, SF83, SKF83, Wat80, Web81, Whi89a, Hsu87d]. single-channel [KSW^{+}85]. single-file [KH80, SF83]. single-level [Nun80]. single-locus [Cre88b]. single-species [Wat80]. single-strand [Cra80a]. single-stranded [CCG87]. singular [SF81]. singularities [Win89a]. singularity [Nun81]. sinks [Bar82a]. sinus [DSR89]. sinusoidal [BRR87, CG81a]. sire [Gri82d, Gri82e]. Sister [WS85a, JJ82a]. sit [Jan82]. sit-and-wait [Jan82]. Site [LB87, BE81a, CSH^{+}81, Di 88, Di 89, Hig85, Iid88, Mag88, MM86a, Mer87, Rei83a, RB83b, Sab80, Sar87, Shi81, UN82]. sites [BA87, Bah80, BT87, Cla82a, Cla82f, Hum81, Kar88b, LM80, Mik84b, MM82c, Nus80b, Nus85, Nus86, OK82, Ren82, SH80a, SW81b, SPG87, Vis87, VGP^{+}80, WWG^{+}88]. situ [Law83]. situations [FLR83]. sizable [BR88]. Size [Cha86, Tys87, ADdACP87, Arm88b, Bel87, Bor84, BC86b, Bro82b, Cal83b, CPLA85, COR83a, Cou80, Day83b, Don86, Dou84, EH83, GPKBG89a, GPAL88, Gre82b, GWK87, GM87, Har84b, Hof83, Hok86b, Kha86, Koo86, KWG87, LIWK86, MVV89, Mur83, NA84, OWN89, PK80, Pug87a, Rei86a, Rei86b, Rei86c, Sch88, SCS88, Smi84, Sza83, TIK89a, TC87, Tho82, Tys83b, TH85, TD86, WS82, Wis80, Yok81]. size-some [PK80]. sizes [Hoe80, Sea80]. skeletal [Fra84]. skeleton [KM87, MK88, SME86, WB89]. skewness [Gat82]. skin [Ale87, HM85b, Nag86, RW82, Sai82, Sax83, SHBS89, dGV80]. skylight [Bri80]. sleep [KFSW82]. sleep-wake [KFSW82]. sliding [CG81a, CG81b, SS87]. slime [Arm84, BS88b, GS86, RSE81]. slope [Bri84]. slopes [BWM80]. Sloppy [TD86]. slot [OK87c]. Slow [AT85]. slowly [BR86, MM80]. slug [PO86]. small [CRB89, CK87, Day83b, Fra87, Fuj82, Pin83b, PL87, SK88b, Win89b, Yor81]. smell [NM81]. Smith [Vic88]. Smoluchowski [HB86a]. snakes [dCB84]. sneaking [dBH86]. Social [Yok84, Yok85, BB81d, BSM82, Bul81, Bul83, Cra83a, Hug86, Iwa81, Lum82, MC89, Moy82a, SI81, Sta85, Tro84, Wad82, Wad85]. Socioinformatic [HH85b]. sodium [BA87, BW86b, Cha82a, JCK80, Mat84, MUI85, MTM86a, MTM86b, Ste88, Str85c, TIK89a]. soil [LS81]. solar [Her83a, Her83b]. solid [Hsu88, SN85, SSN86]. solid-phase [SN85, SSN86]. solids [Bir81, MM86b]. solubilization [Gro82]. soluble [GK80]. solute [FM88, NSJW84, TL88, WWN82]. solute-coupled [FM88]. solutes [And81, BW89]. Solution [Tro83, Ash85, Ber85, KL80, Lan82b, New83, NK88, Pel84, Pys80, Shi80, LL82, LL84]. Solutions [HL82, Bla84, EG87a, GA80, HS89, HS81, LSH88, New80, RCA89]. solvable [BP80]. Solvent [Goo84, KFPB85, Sch84a, War86]. solvents [KLK^{+}88]. solving [FGDS87]. somatic [HRG82, HK81, Nes84, Poz87]. Some [BWW80, BvdHB84, ES88, FQ81, GG84a, JCK80, McG81, Mot89, Moy82a, New80, Nus82a, RS87, RR83a, RR83b, Sau80, Sil83, SP84, Tho83, VFS80, Wel85, Wil81c, Wil84, dCJM86, Aki86, ABS82, Blu83, Dob86, HSd^{+}88, IO83, Inf85, Jam89c, JF85, KN83, Kon88, MSA^{+}81, McG80, McG84, MM83c, Moy82b, PK80, PS80a, Ped88, SRO84, Suc83, Suc85a, Suc85b, TN80, Uma82, VP86, WS81, Wei88, Hoe80, KK87a]. sometimes [Cla86]. song [CJ82, Fle88]. sorting [LL88, MF80, PO86, SCP84]. soup [Kin86]. source [Hic89, Lap80a]. Sources [vZvdSdL81, Bar82a, Gra80, Wal89b, WB82]. Soviet [Ros84b]. Space [BW89, Cal82, LL82, LSH88, Mor80, Mur82, Peu86, RR88b, SS81, SR89]. Space-filling [BW89]. spacer [WS85a]. spaces [INMM83]. spacing [BMR89, MK83, MM82c]. span [ZL84]. spanning [DRP^{+}87, DRL88]. Spatial [BB88b, CDC88, FS87, MN85, Moy82a, NM80, PP80, Pra84, TCY85, Vic89, Bel84, BT84, BPP88, FDC85, Hej82, KAM85, Miy86, NMM87, Nag88, Nam80, Pin84, RC89, Ren82, Ros85, SK86, Sho85, Sta81, TS83]. spatially [Bro81a, CB81, TW80b]. Spatio [RSSK87, Tak87]. Spatio-temporal [RSSK87, Tak87]. Special [DKW89, LB87, Sim85, SF83]. speciation [Fia88, Gla87, LK88, PL85c]. species [Arm88b, Arm89, BBHM85, BP80, Bri88, BA80, CW84, GP82c, Hol85, Ipi87, IAHA80, IM85, Kat81, Ker89, MJN88, MB85, Rei85b, Rei86c, Rid89, Rou80, SV84b, Syv84, Syv85, Tre81a, Tre82, VP86, Wat80, Wel86]. species-abundance [Rou80]. Specific [Kur83, Lam84a, Abu88, Bar81e, Cas82, Cas84, Fuj82, KO82, KK83a, Kur84b, LL87b, McG81, Ond80, SL81, SM88b, Tho85, Var83]. Specific-mate [Lam84a]. specificity [CB84, Kor82, SSF80]. specified [MH84]. specify [Cla86]. spectra [Cla82c, Mei82]. spectral [Zac81]. spectroscopy [Mik84b]. spectrum [FY81, Sir86, Whi82]. speculation [RSR88]. Speculations [Maw85, RS87]. speed [HJS84, Spe86]. sperm [BH83, Par82, Rik82, Rik85, Par82]. spermatozoa [ST80]. Sphere [Lan84a, GS80]. spheres [HE80a]. spherical [DGST82, KL80]. spheroid [Reg84b]. sphingomyelin [MSS81]. spider [COA85, FB82b]. spin [KC83b, Poh81]. spindle [CTS86, FBS85, Hun81, Kin83, PHTCS86]. spindle-free [Hun81]. spines [KT83, KP85]. spiral [BHH83]. spiral-type [BHH83]. spirilla [MS89b]. Spirulina [VFS80]. spleen [Jøn85c]. splenic [WG81]. splice [BBH84, Iid88]. Splicing [Nus82b, Bur82b, HB86b, Hic89, NDB80, Nus80b, Rea84, vHS82]. Split [CG85a, Gra86a]. Splitting [GMM89, RB85]. Spontaneous [Fra87]. spore [GS86]. spores [LA82]. spots [Cla82a]. spp [VFS80]. spread [ASTT80, KAM85, Kel87, Mac87, TATS81, WGP^{+}88, WWG^{+}88]. square [Rei84, Sta88]. square-lattice [Sta88]. squares [Bro80b]. squatting [MM86a]. squid [AM82, Mat84, MUI85, dANG83]. squirrel [SK87]. Stability [ADG80, Aug85, DZJ85, GW88, GS85, Hyv80, McN87, McN89, dMP85, Boy89, BF85, CLFL86, Cus80, Duf85, Esh83, Gal84a, Gal84b, Har87a, Has81, Has82, Kat83, KM83, KL81, Kra82, LS88, LS89, Mah84b, Mal81a, MKS^{+}86, Mor85a, MC89, PL85a, RKL83, Rog89, RHH85, SK85b, SCH84b, SW83, Tay89b, Wag89, Wel83, ZL84, ZG85, CG81a]. stabilization [Pen80, PA86, Sim81b]. stabilizing [BPM82, Has88]. Stable [KSL89, Ros85, BHMR85, CB86a, CN85, CHM80, Com82, CD87, GNG80, Har81, Har83, Hin80, HB84, Hou83, Ish84, KR82, KK86b, Kra82, Llo83, LA86, Lun87, Mat85, MN89a, Met82, Mot88, NST80, RS84, Sch88, Sel80, Smi88, Ste84a, TC80a, Tre82, VC87, VC88a, Wag89, vRV80]. stacking [ANOM82, Aid88, HO84]. Stage [HINT85, HN82, HKK82, HT88, PO86, vS85]. stage-structured [vS85]. stages [Cla82b, FS80, KB88a, PC83]. stained [Bas80c]. stalk [GS86]. stalled [Var83]. stand [Har84a]. standard [HG87, Pro84]. staphylococcal [vHU87]. staphylococci [SP87]. Staphylococcus [FB82a]. Starfish [AK84b, AK86, BHMR85]. starfish/coral [AK86]. START [Shc89b]. startle [And88]. starvation [Gro88]. Stasis [Bar84, WRW83]. State [MH89, AH86, Ash86, Bar80a, Bru85b, Bun80, EP86, FGDS87, FLR83, Ger83a, Gro80, Hir83, Hir84a, KB89a, KKNM80, KK87b, KWG87, KL80, KK80, KCM81, Lap81a, LL87a, LSH88, Lum88, Mor85a, MB88a, NFP88, Ola86, Peu83, PA86, Rap84, Rot82b, SA81, SM88a, SS87, Ste87, Toz89, Wag87, Wag89, WBKW82]. State-dependent [MH89]. state-space [LSH88]. states [BW86a, DW81a, EB87, HR80b, Kra82, MAL87, Pen80, Pok82, RN84c, Sch83a, Ven82]. states-Antagonism [RN84c]. Static [PL85b]. Stationary [Ham82, HKK82, BW86a, NFP88, vS85]. Statistical [Coo82, KK87a, KK87b, KK83b, SV84b, All85a, AZE83, Boo84a, Els83, LM89, MKL81, MH80b, Mic86, SSW81, SG89a, Ste87]. statistics [HCKB83]. status [Har87b, Jam85]. Steady [SA81, Toz89, AH86, Bar80a, Bun80, FGDS87, FLR83, HR80b, KKKK80, KK87b, KWG87, KCM81, LL87a, Lum88, MB88a, Ola86, Peu83, SM88a, Ven82, Wag87, WBKW82]. Steady-state [SA81, Bun80, FGDS87, KK87b, KWG87, Lum88, Ola86, Wag87]. stearothermophilus [BK81]. steered [BC86b]. Stein [Lán84b]. Steiner [Fou84]. stem [OWN89, Pro80]. stems [WJS89]. step [BP81a, CF87b, FEC87b, För88, Hsu87b, Hsu87c, Mag86, Tor80]. steps [GVC81b, Mel82]. Stereochemical [Gra83a, BHM^{+}85, HAG82]. Stereochemistry [GD81]. Stereospecific [LE84b]. Steric [CU88, BP81b]. sterile [Gri82a]. Sterility [MS83, WL87]. Stern [DG80]. Stern-like [DG80]. Steroid [Ros84c, Bog87, Mit88]. steroids [BHM^{+}85]. stimulated [Rat87, SSF80]. stimulation [DLN85, Der87, MUK88, NM81]. stimuli [MB85, Pin83a, Zac81]. Stimulus [MM83b, Sto84a, Sto89c]. Stimulus-response [MM83b, Sto84a, Sto89c]. Stochastic [Ash86, Fer88, Hsu87b, Hsu87d, Hsu87c, Sie81, TW80a, AY85, BP87, GTSNO^{+}88, GW83, Har84b, HS81, Hol89, HM87, HW86, HW87, IP83, JLS80, JS82, KK87a, KK87b, Man81b, Mat87, Rot82a, Rot87, STN^{+}84, SM80a, Sve81, SH82b, WdlT84]. stochastic-deterministic [GW83]. stochastic-mechanical [STN^{+}84]. stochastically [RT88]. stockhastic [Bre86]. stocking [BS82]. stoichiometry [Cha82a]. Stomatal [CF82, SD83, URC83]. stone [Cal83c]. STOP [Shc89b, EKSS88, LJ88]. Storage [Rob85b, SPST83, CBF87, HT83, ML86, Peu83, Rob81]. stragetic [BSS83]. straight [BT81]. strain [GT85a, HTG89, MM83d, RL84, Sto81]. strains [PH87a]. strand [Aid88, AS87, Cra80a, HT89b, McG89a, McG89b, TS80]. strand-exchange [TS80]. stranded [CCG87, GDHB87, Hop86, McG81, NMMO82, SW83]. strands [LBBL84]. strategic [Abu84]. Strategies [GP82a, SG86, Bre80a, Bul84, CHM80, Com82, CD87, Dus89b, Lun87, Mat85, Mot88, RS84, Row85a, Sch88, Sel80, Ste84a, Sza83]. strategy [Cra89, DK83, DPV83, Har81, Har83, HAC82, Hin80, HB84, Hin86, Hou83, Ish84, KR85a, Row88, SH82a, Smi88, TS86, TC80a, VC87, Vic87, vRV80]. streaming [Hay80a, Hay80b, YMS84]. strength [AWMK85, BN81, BB88a, EL83, SYM86]. strengths [Ale84]. stress [CDC88, Reg84b, TS88b]. Stresses [WJS89, PF89]. stretch [MM80]. stretched [BM89a]. string [LVN87, Oli85]. stringent [Eng89]. Strong [Nus80a, Ril80, SPS83]. Structural [LBBL84, NP84b, Pow88, Sin82, ACB81, Arg81, CMN81, CC82b, Coh84, DS84, LS82, MAGR87, Red88, RS86, Sab80, Was86a, Yag85]. structuralism [HL84]. structuralist [LH88]. structurally [Sin83]. Structure [Arm87a, KB89b, Tom82, TLT80, WS81, BB85, BN81, BRM82, BSB88a, BSB88b, BA80, CMM86, CC83a, CHT80, Coh84, Cou80, CM87, DD85a, Duf85, Eva80a, FBF85, GPAL88, GAK88, GRS82, HN82, Ham82, HKK82, Har84a, HU85, Hor89a, Kam80, Kat83, KHB83, KM85, KK86a, KO83, Kri85b, Lan82b, Lav84, Lev80a, LR88, Luc82, MW83a, MN89a, Mau87, McG84, McG89b, MHB83, MF85, NHT88, NA84, Nus80b, RMFM85, Ric80, RB83b, Ryt87, SR80a, SP87, SGH85, Sim81b, Sim85, TCY85, Wag82, Wic85b, YLBB81, YBBL82a, YBBL82b, YBBL82c, Yod81, Yod84]. structure-activity [KM85, KK86a]. structure-function [FBF85]. Structure-hepatotoxic [WS81]. structured [AM81a, Abu86, CC81, Cus80, Ebe87, KP89a, McN89, Mur83, SVN89, Tre83, Vla87, ZL84, vS85]. structures [AY85, BPP88, DB88, DRP^{+}87, DRL88, GRS82, Hun85a, Hyv80, Ji85, KS88, KSL89, LP88b, MM82b, McG81, MG80, Mil89, Oli85, Pap82, PDCD88, SF89, SBW83, SJK88, TS83, Tho85, Was82b, Wax81, dVR85]. studied [AH86, FEC87a, vGvGL88]. studies [Cra80b, DD85b, DW87, FGDS87, FCBK88, GPM89, KN82, KN83, KS81, Kot82, KK83c, Kot83, LBBL84, NDB80, SFR86, SRO85, Spe87, SC89, YLBB81, YBBL82a, YBBL82b, YBBL82c]. Study [QL82, ZBL80a, ZBL80b, ANOM82, Aid88, AP80, BB85, BW89, BLR83, BVDA87, BF81, CV89, DC87a, DD89, Fuk81, GPLC84, GL84, Has81, Hay80a, Hay80b, Hol89, HFO83, HCW89, HB89, ITK88, JCK80, KLR89, KT83, Kir86, KK83b, KB89b, Kol82, KS81, KK81, LO82, MW83b, MT84, MHB83, NA84, NU82, dANG83, OL86, Omh87b, PS81a, PS83, RSE81, SH82a, Shi81, SLFT86, Suc85a, TN80, TGP^{+}86, UN82, Von82, Wol80, YKS80, YVIG80, Zaj86, ZC89]. studying [Cur88, MCB^{+}86]. Sub [SG89b]. Sub-exponential [SG89b]. subdermal [SA81, Sax83]. subdivided [Fra86, Moo88, Yok85]. Subdivision [SPP84]. subendothelial [WGP^{+}88, WWG^{+}88]. Subfertility [Cra83b]. subject [Car81b, SPS83]. Subjective [Lib85, Wol89]. subjects [BCL83]. submerged [EH83]. submergence [Mar84]. subordinated [Hau82]. subsequence [LR85a]. subserving [CPLA85]. Substance [Bar86, Ham85]. substances [AS86, Bas80a]. substituents [ELM81]. substituted [YHJ89]. substitution [Cla83b, SR80b]. substitutions [Fil88, LVN87, SK84]. Substrate [TCR85, BW89, CA89, Der84, Fer87, GVC81a, HC87, HC89, Hsu88, JK81, KVW88, KL80, Lan82a, Law83, ME82, MM86a, RP84, RN84c, SM88a, WT87, Web81, WD81]. substrate-channelling [KVW88]. substrate-cycle [RP84]. substrate-mediated [BW89]. substrates [SAS85, WT87]. subtilis [BB81b, FLNF86]. subtraction [Smi84]. subtransitions [SG89a]. Subunit [RN84c, RN85, SA80, BF81]. Subunits [Vau86, MKL81]. Successes [Hol88]. succession [Fin82]. sucking [DK83, KR85a]. sucrose [SS81]. suction [MOV82]. Sufficient [Yok85]. sugar [OH85]. sugar-transport [OH85]. suggesting [Dou84]. suitable [Ola86]. sulphonamide [PR82]. sultones [RW82]. summary [Els81]. sun [PW88]. sunflower [BHH83]. supercoiling [For80]. supercritical [Gad81]. superfamily [BSB88a, BSB88b]. Superhelical [Fuk83, Fuk84]. supernumerary [TT87]. superorganism [Lum82, WS89]. superoxide [BK81, OO84a]. superpacking [NBZ^{+}83]. supply [Cze87]. supports [KL80]. Suppression [McM81, Bri82, FE85]. Supramolecular [KSM85]. Surface [Gro88, MOK87, PSL^{+}85, Wil82, BG84, Bir81, BK88, BF81, Cum85b, Cum89, DV83, DT80, Eco84, Frö84, Gal84a, Gal84b, GVP^{+}80, Gra81b, GWK87, HT89a, Joh84, KO82, KK83a, Koc88, Kwo84, Kwo89, Lan85, Ler83, Lev80b, MUC81, MCB^{+}86, MN89b, vdHN85, OK87b, Rab82, RAR83, RN84a, RN84b, SA80, SEN82, Ten80, Tod85, TY82, VA88a, Vro83, vHU87]. Surface-limited [Gro88]. surfaces [BO82, Cha89a, DV83, LE84b, ZR84, ZR85]. surfactant [Col85]. surfactants [KLK^{+}88]. surgery [MH81]. surname [BW81]. survey [HT87]. surveys [SP84]. survival [Abu85a, Alf83, Ker89, MB82, PWH^{+}88, PTD81, SH82a, SV89, Süh89, SH82b, WG81, Zno83]. survivor [Man81c]. survivorship [Ing89, Pea88b]. Susan [HLS89]. suspension [Koe83]. suspension-feeding [Koe83]. suspensions [BT85, HB86a]. sustained [Rob89]. SV40 [BBH84]. swarming [Bul83]. swarms [LGK84]. sweat [JD82, JK89]. sweet [MB85]. Swelling [Hol83b]. swimming [GBFH88, HJS84, MS89b]. switch [GH82, JC80, ND82]. switch-on [ND82]. switch-on/off [ND82]. switches [Mit89]. Switching [MKS^{+}86, Bar81d, Bar82b, JCZ84, Mat85, PH87b]. symbionts [Col82]. symbiotic [GJ88]. Symbolic [MH81, Hux86]. symbolism [WD87]. symmetric [Bri84]. symmetrical [Bia81, FE85]. symmetries [Ser87]. symmetry [DfP87, DK82a, DK82b, LO82, Shc88, Shc89a, Shc89c, Whi80c]. sympathetic [KT81]. sympatric [Fia88]. symptoms [Ewa80]. synapse [TO80, TG83]. synapsis [CG80, CG84]. Synaptic [LD83a, VM89, AdMM86, Poz87]. synchronization [Gro88, Hir84a, Koc86a]. synchronizations [YKS80]. synchronized [Gri82b, Gri82c]. synchronous [TH80]. synchrony [Agu85]. syndrome [Der84]. synergetics [ST86a]. Synthesis [Gra83b, Bar89, BJL85, Blo87, Coo88, Dou81, Dou85, Hun85b, JC80, Kuh82, LB87, MdMF83, Mil87, Nit86, Pet82, RQ81, Shu84, Spe84, Tys83a, WI80, Whi80a, vHBL87]. synthesize [Var83]. synthesized [WGHY82]. synthetases [Kum82]. syrinx [CG85b]. system [Bec85, Ble88, Ble89, Cha80d, CLFL86, CN85, DG85, DG87, DDBB83, FE85, FM88, FJFN89, GA80, GFA89, GH82, HH89, Hof86, Hof87, Hsu87d, Jaf80, KFSW82, KB89a, KK87b, Kol82, KS82, Kur84b, LS87, LSFT85, MKS^{+}86, MN85, MV88b, NM82, NM85, Nag86, Nit86, OKT^{+}80, OH85, Oon82, PWH^{+}88, Ras87, RB85, Row87, Row88, SBC80, Sim80a, Sim80b, Sim81a, Sin82, Str85a, Str85b, Str85c, Tan84, TIK89b, TG80, VC88a, War86, Web81, Wes83b, WI80, YMS84]. systematic [Ola86, Orl82]. systematics [Day83a, Fai85]. systemic [Jea80, SS89a]. systems [Abu84, ASK89, AH86, BB85, Bar80b, BC80, BS82, BS84b, Bro80b, CPPR87, CPPRS87, Coo84a, CB89b, CN84, CK85, DfP87, DW81a, Dem83, DZJ85, ER88, Fer85b, FSF89, FS80, GVC81a, GS81, Gar89, Get81, GEL80, GE82, GT88b, GS85, Gra85b, GH87, Hal80a, Her81b, Hun85a, Hyv80, Hyv84, IBJ^{+}89, JB80, Jur84, Kat81, KVW88, KK87a, KR82, KK82a, KC83b, Lam84a, LO82, LR80, Mal88, Mau87, Mor85a, NSJT81, OH83, OH84, OC81, Par85, PCTF83, Peu83, Peu85, Phi81a, Phi82, Pok82, RR88a, Ric80, RLS82, Rob87, Rob88, Sal84, SNW81, Ses82, SR80b, ST86a, Spa80, SB86, Tho82, TG88a, Wag82, WB85, WKV88, Wel83, Wel86, Zam86, vGvGL88]. systolic [Reg88].

T [Ano80j, BBH84, GAD80, Bec81, FE85, NMMO82, PH87a, Ses82, Tho85, dBH86, dBH89]. T-cell [Bec81, PH87a]. T-independent [GAD80]. T-lymphocytes [dBH86]. T/G [MF85]. TAA [Shc89b]. tactic [Kil80, MUK86]. tactics [WW84]. tadpole [Was86b]. TAG [Shc89b]. tail [Ser87]. takeover [LJ88]. tandemism [Gla82]. tank [Hsu87d]. Target [Dra81, GPKBG89b, Kar88b]. taste [Ven84]. taxa [Pea88a, Pea88b, Pea88c, Pea88d]. taxis [Gla80, KTS89]. taxon [Pea88a]. taxonomic [AM87]. taxonomy [MLB88]. Taylor [GKB86]. Tchuraev [Ano81i]. technique [KK86a, RTM81a, RTM81b, RTM82, RR88a, Rob85a, SR87b, SW84, Wei80]. techniques [Sir86]. teleology [O’G84]. Teleonomical [Lef83]. Telling [Car82b]. telomer [MM82b]. Temperature [Mon88, Sax83, AA84, FEC87b, FEC87a, HSV84, HLS85, KFSW82, KB86, Mul85, SSM81, Tho86]. temperature-dependent [SSM81]. temperatures [DY82, FEC87a, Sea80, TL88]. templates [SG89b]. Templeting [KsC85]. tempo [Mit89]. temporal [DS84, LG89b, RSSK87, SR89, Tak87]. Temporally [AM82, TC87]. temporary [BBH86]. Tension [GG84b, Gal84a, GM82a, Gro80, Sau80]. tension-displacement [Gro80]. tensor [AA85, HR84, HR86]. tensorial [Pel84]. teratogenesis [Shu84]. tergite [MM82a]. term [AA83, Gra88, Man82a, Pin83a, Pin83b]. terminal [NDB80, Poz87, RMFM85, Rei83b, Spe84]. terminals [LBW86, Maw85, Miy86, PS81a]. terminate [PS84]. termination [UK82]. terminator [OK82]. termini [Ber89]. termini-hypothesis [Ber89]. terminology [Wol85]. terminus [MDPF86]. terms [SGH85, Tak89c, WBKW82]. ternary [LH89, Mal88]. terrestrial [Dob86, Spe86]. Territorial [NDB83, Bro82b, ND82, NMB85]. territoriality [Sar85]. territories [TH80]. territory [PK80, TH80]. tertiary [CMM86, GRS82]. test [ASTT80, BM87, BMR89, Bel87, HH85a, Lum88, PJK^{+}83, PR82]. Testing [DS84, GP80, CF87b, Jam89a, SC87a]. Testosterone [Jam88, BCL83, Ber87, CH86, Jam89a]. Tests [Vau86, Rid89, Ton86, Wol85]. tetanus [SSF80, SSF87]. tetrad [Sch80]. tetrahedral [TRS84]. Tetrahymena [Lap84]. tetrameric [FJ84]. tetrapeptide [Bro83]. texana [ST80]. textured [Pin83a]. thalidomide [Shu84]. Their [PA82, Sch81, Ale87, ABS82, BSR^{+}88, BWM80, Ben89, BP81b, EG87b, Flo84, Hum81, Jak88, Kel87, KS88, KP85, KK80, LP83, MAGR87, MB88b, NSJT81, Rei85b, RW82, SCH84b, Smi82a, War85, WGP^{+}88, WWG^{+}88, Whi81b, ZC89, vZvdSdL81]. them [Whi80b]. theme [MF89]. theor [Ano80j, Phi82]. theorem [CB85, Gie88, Has82, Wag82]. Theorems [Wei85, BBH86, BM88]. theoretic [CK82, Gol83a, Sch83b]. Theoretical [Aki84, Bas80c, Bee80, BW89, Ber82, Blu84, BSB88a, BSB88b, CG85b, DD85b, DW87, Fel80, GM82a, GDHB87, Ger83b, GPLC84, Hay80b, Jaf80, KT83, KBR88, KN82, KN83, KHU^{+}83, KB88b, LJTH85, MUC81, MT84, Miy86, Mor86, MM80, Nus87, OH85, RS83, Sal84, Shi81, Sta86, Sto89a, VM82, vHBL87, Aki85, Bar81c, BKB80a, BKB80b, BP83, dCB84, CP85, CW88, CLF84, Cra81b, Cru80, Dug86, DD89, FQ81, GS85, Gra80, GWK87, HF85, HU85, Hoe80, Hof83, JCK80, JP85, KLK^{+}88, KK82a, KL83, LO81, LRKC83, LFP84, Mot89, dANG83, Nus82b, OL86, Omh87b, Ond80, PS80a, PS81a, Per82, PPC82, RAR83, RR83a, RR83b, TN80, TGP^{+}86, VFS80, Van82, VDH82, Von82, WR83, YKS80, JLS80, PHTCS86]. theoretically [Pen80]. Theories [HK83, DW82, Gre80, HK81, Hon84, Lib85, PLJ84]. Theory [BAB84, Bre83, Can87, Eld84, GM82b, Izu81, Kat83, Kat85, Man81b, MSS81, MS86b, NSJT81, PCTF83, Rot82b, Sch82b, Ton86, VD89, Wei80, Abu88, Alf83, Alm85, AOB85, AA85, AWP80, Ash85, Asp88, Bân83, Bar81a, Bec85, BN81, Bla80, Bot82, BS89, BvdHB84, Bul83, CG85a, CPPRS87, CL82, Coo81, CB89b, Cra80b, Cra81a, Cri82, Dal83, DC87b, DT80, Eld80, Els81, FM88, FFM82, FS83, Gin86, GP82b, Gla87, Gle84, Gra88, GG82, GG84a, Gri81a, Gri81b, Gri81c, Gri81d, Gri82b, Gri82c, Gri82d, Gri82e, Gri82a, Gri82f, Han86, Har80a, HH85a, Har87b, Hau82, Her81a, Hof87, Hof82, Hof85, HF88, Hor89a, Ibe83, JPL83, Jon85b, KKKK80, Kat81, KL87, KK87b, Kil80, KRB81, KM84, KVB83, Lew81a, Lib88, LPS89, Mäk86, McC84, MH80b, MP85b]. theory [O’G84, Obe81, PB81, PE82, PV83, Pic86, Pow88, PP80, RT85, Red88, Reg84a, Rot83, RT83b, Row85a, Row87, Row88, SK88a, SSM81, SC84a, See86, Shi87, SC87a, Sla80, SW81a, SS87, SN85, Ste81, SF83, SKF83, Stu86, Syv85, Tan80, Tre83, TLN85, Val85, Was86b, Whi80a, Whi89b, Wic80, Wic85b, Wil81b, Wol89, Woo80, Yoc81a, Yoc81b, GZWR80]. there [Abu85b, And80b, Bot82, Ere85, FJ84, Kin86, Len84, MB88b, Par82, SYM86, TDA86]. Thermal [Gre85, McA87, MM86b, Ten80, DY82, Dus88, KB86, SW83]. Thermodynamic [GEL80, BG84, BH88, CFLF88, CTS86, Dix83, FJ85, FM88, JF85, LP83, MTM86a, MTM86b, Pet82, Peu83, RT83b, Ste87, WK81, Wic80]. thermodynamically [Tai86]. Thermodynamics [MR89, Wic85b, AH86, Bry80, Cha80b, Di 89, Gla82, GE82, Ish81, LR88, McC80b, Wes83b]. Thermoelectric [SI81]. thermolysin [KB89b]. thermophila [Lap84]. thermophile [ANOM82]. Thermophiles [Mal81a, Dan86]. Thermoregulation [Omh87a]. thermostability [ANOM82]. Thermosynthesis [Mul85]. thermotaxis [MUK88]. thermotropic [ST86b]. thickness [HF88, HO84]. thinning [Giv86, Pah85a, Per84, Pic84, Whi81a]. third [SR80a, SM87]. third-order [SR80a]. thirty [Har87b]. thirty-four [Har87b]. Thomas [Hof87]. thoracic [RS86]. thorns [BHMR85]. thoughts [Wil84]. threat [And80b]. Three [BB81b, BP80, BP81a, Hun85a, Nus80b, Zam81, Ash86, Bru85b, CW84, CSH^{+}81, Got86, Ipi87, KFSW82, KO83, MKS^{+}86, Mel82, MTM86a, NHT88, NSJT81, Row85b, SP87, SW81b, VP86, WH83, KKNM80]. Three-dimensional [BB81b, Zam81, CSH^{+}81, KO83, NHT88, Row85b, SP87, SW81b]. three-oscillator [KFSW82]. Three-species [BP80, VP86]. three-state [Ash86, KKNM80]. Three-step [BP81a]. threshold [GTMM87, GTSNO^{+}88, SAS85, Sla80, SK84]. thresholds [Lib80, NS89a, Sta88]. throw [Cal83c]. thylakoid [RB85, ZG82]. thymic [SM88b]. thymus [SM88b]. thyroid [DO83, She80]. TIMA [Was82a, Was82b]. TIMA-based [Was82b]. Time [AWP80, CN80, Fer81a, Fer85b, KMSR84, LSH88, Phi81a, Phi82, SH87, SVN89, ASK89, ADG80, Bar80a, Blo85, Blo87, Bre86, Cal82, Coo84c, DR89, DY82, DKW89, DDJ83, FK83, Gad81, HKK82, HR80b, HW80a, Hon84, Hon86, Jam85, KM81, KSW^{+}85, LL84, Lew81a, LB87, Lin80, MMGM80, Mas86, Met82, New83, Oon82, Sen88, SF81, SM80a, SMGD80, Sny88, Tho86, TSRK88, Tys83a, TH85, Vla87, WGP^{+}88, WWG^{+}88, Whi89a, vS83, vZvdSdL81, FS83, SF83]. Time-constraint [LSH88]. Time-dependent [Fer81a, Fer85b, ASK89, HR80b, SMGD80, WGP^{+}88, FS83, SF83]. time-invariance [vS83]. time-persistent [Vla87]. time-variance [DKW89]. timer [Men82]. times [Bas82, HW80b, Law83, Pow85, SS88, Smi87, TC80b]. Timing [Cal83c, Fuk81, PC83]. tiny [Par82]. Tip [GT85a, BG88, PS82b]. tissue [Coo84b, CM86, Ede82, FJFN89, GPAL88, HL82, HG87, LL82, LL84, MB84b, NMM87, SE80, Sil89, TS83, TS88b, Wei83, Wis80]. tissues [AL81, Bar82a, KYU^{+}86, MKM80, SA81, Sax83, SR81, Suc83]. TMV [AVR82]. toads [LCDC84]. tobacco [Dun82]. together [Gri82f]. tolerance [PGKE87, PGKE88, SM88b, Sto81]. tonoplast [Tai86]. tool [AP80, MGL81]. Tooth [Hor89a]. TOP [Ham85, Leo84]. top-down [Leo84]. topical [SHBS89]. Topological [MP85a, Pys80, MF80, McG84, Tho83, Sim81a]. Topology [Str83, Kin82]. toponymic [Ham85]. torpor [PJ86]. torsion [Gre85]. total [LPS89, MV88a, MV88b, Pea88c]. toxic [Lil82]. toxicant [HH89, SC89]. toxicant-individual [HH89]. toxicants [HdL84]. toxicologic [BGB89]. Toxin [Str85c]. trace [Wal89a]. traceback [Got86]. Tracer [Tra85, BAB84, BBH86, FGDS87, FCBK88, Wei85]. Tracheal [MM80]. tracheary [Pic89]. tracings [BBS88]. tracking [TC87]. tract [AHFH81, MHII83]. tradeoff [Cal82]. trained [KMSR84]. traits [Rob85a]. trajectories [Gol88]. tramformations [McK83]. Trans [Wig81]. transacylations [Pen80]. transcarbamylase [MH80a]. transcribed [Bro83]. transcription [Bee80, BT87, Fra88, JSJS88, Mad88, MMGM80, Mil87, Nus86, PY88, San84, SRO85, Sek85]. transcription-translation [Mad88]. transcriptional [GDHB87, Lav84]. transducer [Dun82]. transducing [Lac84, ZG89]. transduction [LM82, Mal84a, Mal85, Mer87, ML84, MYS89]. transfection [Kle81]. transfer [Bar81d, Bar82b, EAL80, FS85, GBS89, Jur84, Kir80, KL81, Kri85a, KVB83, Lam84b, Lay87, MLS80, Mik81, Mik83, NU82, Ola82, PSG83, SR87a, SR80b, SLYF84, Syv85, SB86, TG88b, TD81]. transfer-Complete [TD81]. transferred [TH81]. transform [KT87b, LL82]. Transformation [Uma82, Alf83, BS84b, Gal84a, Gal84b, Hje86, Hyv80, Lan84a, Sco84b, Shc89a]. transformational [CV89]. transformational-grammar [CV89]. transformations [AA85, McG80, Pel84, Sla85, SOLG82, Wag82, Woo80]. Transient [Hsu87e, Poz87, Shi82, CLF84, CN80, HN82, LL82, LL84, MUC81, WGP^{+}88, WWG^{+}88, GVC81a, GV81, GVCC82]. transiently [SPG87]. transients [Bro81a]. transit [Bas82]. transition [Cla83a, DW82, FEC87b, Gra85a, Hor89a, HT88, JS80, JPGL86, LP82, MAC89, Pen80, PA86, RN84c, TL88]. Transitions [Hop83, Hao82, HC87, Kra82, Leu87b, MZM89, Sch85, ST86b, TH85]. Translation [LvH87, Bah80, Cra81b, HR80b, Lum88, Mad88]. translational [DCR82, EKSS88, MG84a]. translocated [SC84a]. translocation [Dou85, NMB85, Osu84]. transmembrane [EPH86, KFPB85]. Transmission [FB82b, BKMS87, BR83, Coo84a, FHL89, KRB81, Rég84c, Spe81, Vau86]. transmitter [Miy86, TS85]. transpeptidase [All85b, Osu84]. transpeptidase-mediated [Osu84]. transplant [TT85]. transplantable [SH83b]. transplantation [OWN89]. Transport [GM84, Rot83, AH86, AT85, AWP80, BBH86, BM88, Bha86, BVW83, BK81, BD80, CS82, DSD81, Dug82, DT80, ER84, Fer80, Fer81b, Fer81a, Fle80, FS83, FB86, Gai80, GK80, HL82, HN84, JCK80, JA88, KFPB85, KB89a, KC83b, MB84a, Mas84, MB84b, MF81, MTM86a, MTM86b, MRR80, Nus82b, OH85, Peu85, Sar82, Sch83a, SA86, SA82, SMGD80, SF83, SKF83, TG84, TGP^{+}86, Wig81]. transports [YMIH86]. transposable [Hic89, ND82]. transposon [Kin85]. transposon-based [Kin85]. trauma [GPAL88]. traumatic [TKB80]. travelling [Dob86]. traversal [HW80b]. treating [Coh84]. treatment [CG84, Ewa80, Ger83b, GG84b, IO83, KLK^{+}88, KN83, KP81]. Tree [ICL85, Lew82, Val85, DRP^{+}87, DRL88, Hyv80, Sch89, WJS89]. Tree-growth [Val85]. trees [AY85, Bri82, Day83b, HHP89, JLS83, KB81, LdJ86, Lew82, Mac84, Mäk86, McM81]. Treisman [Wil83]. tri [ZBL80b]. tri- [ZBL80b]. triad [ANS83]. triangular [MZM89]. triceratops [Yca82]. tridactyla [DC87b]. trigger [JS80]. triggering [Wri80, You85]. triggers [NSCH88]. triple [NMMO82]. triple-stranded [NMMO82]. triplet [BS83]. triply [HMF88]. trisomy [HCW80, Mar80a]. triver [Rob80]. tRNA [ANOM82, BS83, BS85, Ber89, Fel83, Jak88]. tRNA-like [Ber89]. tRNAs [MKM80, YI88]. trophic [Cou80, HBP89, Lev80a, MKS^{+}86]. tropomyosin [NMMO82]. tropomyosin-troponin [NMMO82]. troponin [BSB88a, BSB88b, NMMO82, RH80, Rei85a]. trp [CM82]. true [Ola82, Smi85b]. truncated [AS86]. truth [Car82b]. trypsin [NSKS85, UN82]. tryptophan [Man81b]. Tsukahara [KP85]. Tu [Jak88]. Tu-GTP [Jak88]. tube [EAL80, MHII83, Sil89]. tubes [PS82b]. Tubular [HE80a]. tubulin [MS86a]. Tumor [DH85, GB80, MMGL87, dBH86, KF87, MKM80, SSF80, SSF87]. tumor-tetanus [SSF80]. tumorigenesis [Kle81]. tumors [All85a, SH83b]. tumour [BM85, Dug82]. tumour-induced [BM85]. tumours [CCDF81]. turbulent [Li88]. Turing [HSV84, Har87b, Hun85a, Mur82]. turn [BC86b, Li83, Mur85]. turn-over [Li83]. Turning [Bor84]. Turnover [JB80, FCBK88, KD85, LM80, Law83, SR89, Tay88c, Wag89, Zaj86, vS85]. turns [Mur87]. turtle [Sim80b]. twice [Gho82]. twist [ZC89]. Two [Cas84, Cho81, Hao82, HH81, KS80b, Niv83, Tre81b, YMIH86, AL83, AM81b, Aok84, BP81b, Com82, DD85a, EN86, FJ84, Fer87, FEC87b, GVC81a, Ger83a, GPLC84, Gro80, Has81, Hoe80, Hof82, Hsu87b, Hsu87c, Ive88, Kat81, KK83b, KR85b, Koo85, KK86b, LRKC83, LH86, Les86, MHII83, MB88b, Nit86, NP84a, OKT^{+}80, Par82, SS88, Ser87, ST88, SS87, SBPF83, TT85, TS80, TS86, WT87, Web81, Wes83b, Wri84, YKS80, You81, YVIG80, KK80]. two-component [Hof82]. Two-dimensional [YMIH86, AL83, MHII83]. two-factor [OKT^{+}80]. two-locus [AM81b]. two-membrane [Nit86]. two-phenotype [Les86]. two-policy [Aok84]. two-state [Ger83a, Gro80, SS87, KK80]. two-step [FEC87b, Hsu87b, Hsu87c]. two-strategy [TS86]. two-substrate [Fer87, GVC81a]. Type [HBRN81, BJ84, BHH83, BLC86, JJ82a, JCZ84, MATGR87, MAGR87, NL86, PH87a]. types [BGB89, Hoe82, MB88b, SS88, Vil83, Whi80b, Whi81b, Wri84]. typhimurium [Cla82a, CM82].

U [MF85]. u.v [MPPF80, dGV80]. ultrastructural [Woo80]. ultravolet [MdMF83]. umbilical [FW81]. unbiased [JA88]. unbranched [AK82]. uncorrelated [Wei88]. uncoupled [Nan86]. underdominance [GZ86, ZG85, ZG85]. undergoing [NW85]. underlie [IT82]. Understanding [FBB89, Suk89, VLCK86]. undulating [Ray85, WS84]. unequal [HRG82, KDAT84, Kor89, Mot88]. unexpected [Orl81b]. unfavourable [VL86]. unfolding [HBP89]. Uni [GVCC82]. Uni-Bi [GVCC82]. unicellular [LO82, PL85c]. Unidirectional [KH80, Win81, Wag87]. unified [Emi82, MP85b, dL83]. uniform [BF83, Kaw84, Sch89, Whi80b]. unifying [HLS85, JGW83, MF89]. uniparental [Col82]. unipolar [MS89b]. unique [Hop86, vHU87]. uniqueness [Sar85]. unireactant [SM88a]. unit [Leu87b, MZM89, NS86, RTM81b]. unitary [MKL81]. units [BB81a, Cal87, Don86, Eil88, GJ88, GM87, KVB83, Leu87a, Rei85a, Rei86b, Tak89b]. universe [Fer82]. unknowns [Ols87]. unlikely [Bro82a]. unrelated [Tre81a]. unreliable [Wal89b]. unsaturated [Inf85]. Unsolved [Dan82, MDW80a, MDW80b, MDW80c]. Unstable [GGPM80, Tys83b, WS82, Dug86, Sin83, Top88]. unstirred [Gai80, HF88]. unswollen [SME86]. unusually [NST80]. upon [Bor84, BC86b, Bru85a, Her83a, Lav84, Men85, MHB83]. upper [Jac83, Law83]. Uptake [WMS81, Bas80a, Bru85a, VKV83, WP84]. Urbain [Hof87]. urchin [BH83, Rik85]. Urospasm [Smi85a]. usage [AS87, Bul88, Cla82a, LvH87, Wel85]. Use [Frö88, Gri81d, Gri82c, Gri82e, Gri82a, KM85, Oli85, BM87, CB83, Cur88, DG85, FW81, FM88, FDC^{+}89, Maw85, McG84, MH87, MB88a, Ola86, Rid89, Rog84, SPP84, SS83b, SK84, Gri81b, Gri81c, Gri82b, Gri82d, Gri82f]. used [Els83, Rei86b, Tay88b]. usefulness [Mac84]. using [AFHM86, AS86, BMR89, BB88b, DFS89, DRL88, HMF88, Ive88, Mac82, Pel84, RTM81a, RTM81b, RTM82, SPS83, SC83, Whi89b]. Ustilago [HL86]. usual [Mor85a]. uterine [GTH88]. uterus [CY84, GTH88]. utilizing [Tak89c]. UV [LBW86].

V [Ano81i, KK83c, Gri82b, KK82a, PL85b]. vaccines [Koo85]. valency [LP82]. valid [MB88a]. validity [Aal89, SF89]. value [CB85, CN84, EL87, Izu81, UN82]. values [Bar83a, KC83a, LdJ86, PDCD88, Sib83, Yos86]. vanishing [FS83]. Variability [SGJ86, ÅF80, AS86, BP85, GK85, HM85a, MATGR87, Pet83, RC89]. Variable [BA80, Cha82a, Fil88, Reg86, Sch88, SH82a, She83, Wis80]. variables [DCR82, Lap81a]. Variance [GR88, Sto84b, Bas82, Dom83, DKW89, HW80b, dCJM86]. Variance-covariance [Sto84b]. Variation [PH87a, Ber87, BR83, CG81b, CK82, EL87, HvL86, Jam89c, Miy86, PC83, SR89, SH85b, WR83, vZvdSdL81, NM81]. variations [BW85, Hej82, JL89, MdMF83, PV83, Sve82, Yor81]. variegation [Fra88]. Varieties [Hol89]. Variola [ASTT80, TATS81]. various [KWG87, Whi81b]. vary [Gul82]. varying [DY82, Lei81, SPG87, TC87]. vascular [AWP80, Bas82]. Vector [Hej82, Pel84, TT85, TT87]. vector-matrix [Pel84]. vee [BH81a]. vegetation [SP84]. velocities [SPS83]. velocity [Bar83a, BC83, Hou86, KTG86, KT87b, SM88a]. ventilation [CLF84]. ventricular [IYS83]. Verhulst [Ros85]. Verhulst-limited [Ros85]. verification [FM88, FP89]. verifying [AM87]. version [Mac82]. versus [Duf86, Hin86, Law83, PS83, Pea88d, Sin82, SLYF84, TACM82, Tro83]. vertebrate [CVNO87, Fil88, Hol83a, MDBRM87, Sim80a, Sim80b, Sim81a]. vertebrates [RL84]. Vertical [Rég84c]. vertically [IPS83]. very [Fra87, JJ80, Kac87, SPS83]. Vesicle [OCMP89, Las87, McK83, ZG89]. vesicles [Bre80b, Hol83b, Kat83, KM84, WMS81]. vesicular [AWP80, HN84]. vessel [CP82, Che83, Che84a, Wax81]. vestibulo [Mon88]. vestibulo-ocular [Mon88]. VFA [FGDS87]. VI. [Gri82c, Kot83, PG88]. via [AL87, Pin84]. viabilities [Has81]. viability [And86, Cre88a, Cre88b]. vibration [FB82b]. vibrational [De87]. vibrations [Hav89]. vicinity [Lap80a]. view [Kub85, McG85, SM87]. Vigilance [Lim87, Lim88, KLR89, MC89, Mot89, War85]. VII [Gri82d]. VII. [WS81]. VIII. [Gri82e]. Viking [LS81]. villosity [Win89b]. vinyl [Aki85]. violating [Tra86]. viral [Gar80, Mer87, RT83a, RT83b, WS85b]. viroids [Flo84]. virulence [BP83, Joh86, Kno89]. virus [Bec80, Dun82, FM86, Inc83b, LG89b, Luf82, MK81, PH87a, RMFM85, SV89, Süh89]. virus-infected [SV89, Süh89]. virus-receptor [Inc83b]. viruses [Bân83, Cha88a, EPH86, HHP89, Inc83a, Nee89c]. Viscoelastic [MAL87, DGST82, DZJ85]. viscosity [AA84, Bla84, GJG83, WSMP83]. Visual [LBW86, And88, Maw85, Pin83b, Sim80a, Sim80b, Sim81a, VD84]. visuomotor [LCDC84]. vitro [Ada80, BM89b, MB88a, Rap84, RQ81, Sýk84, VFS80]. viviparity [SG88]. vivo [FCBK88, Kle81, MK81, PHTCS86, PP85, Sýk84, WK81, dMP85]. vocalization [Ano80-35]. vocalizations [Mar80b, SK87]. voltage [BB84, Cha80c, Cla83a, Col84, Sch89, VG80]. voltage-sensitive [Cla83a]. Volterra [BP81a, GA80, Shi87, Tay88b, Tay88c]. volume [Ano84h, Ano84i, Ano84j, Ano84k, Ano84l, Ano85h, Ano85i, Ano85j, Ano85k, Ano85l, Ano85m, Ano86h, Ano86i, Ano86j, Ano86k, Ano86l, Ano86m, Ano87j, Ano87k, Ano87l, Ano87m, Ano87n, Ano87o, Ano88n, Ano88o, Ano88p, Ano88q, Ano88r, Ano89g, Ano89h, Ano89i, Ano89j, Ano89k, Ano89l, BS84a, DDJ83, FM88, GF81b, Mas84, Suc85b, Sve82]. volumes [Bon85, Tra85]. vortex [KSL89]. vs [Jan82, KR82, Mit89, Rak80, SKF83, WKV88].

wait [Jan82]. wake [KFSW82]. walk [BB88b, Lap81b, MJ88]. walking [Jac83]. walks [KL87]. wall [Bol80, EAL80, KD85, Koc88, LBBL84, iWSS88]. walls [Nev88, TGP^{+}86, VFS80]. war [HP82, NR85, Ril80]. Was [Kin86, FM83, TIK89b]. wasps [Gre82a, SI81]. wastage [ER88]. wasteful [Sza83]. Water [Hal80b, Bry80, CF82, Dam86, Gai80, HMF88, Har82a, HL82, JK89, Kat85, Lap80a, LP88b, Mor85a, NU82, Omh87b, RB85, Sch84a, Spe87, SR87b, SR89, SD83, Tat88, VGP^{+}80]. water-protein [Bry80]. Watson [Gad81, Ji88]. Watson-crick [Ji88]. wave [CBLM81, MUK86, Nag86, Whi82]. waveforms [KT87b, Rik85]. Wavelike [Nag89]. waves [CVNO87, Dob86, Fer83, Lan85, PY88, SH85a, TLS83, TG88b]. way [Gat86, Mic88, Nun81]. Ways [HL84, PS82a, Rat87, Shc89c, WJT87, You81]. weak [Coh84]. wearout [Mil89]. web [COA85, Leg87]. webs [FB82b, KL81, Lev80a]. weight [Pec88]. weighted [AP80]. weights [HT87]. were [Sta85]. western [YC89]. Wetting [FJFN89]. whale [KB88b]. Where [MM86a, Gho82, MG84b]. Whether [TIK89b]. which [Bre86, GTH88, Inc83b, Kat85, KBR88, PC86, Pic89, SS83b, Sil89, TLN85, Wee84b]. while [Lim87]. white [Str80, TW80b]. who [Har85]. whole [Amt86, CP85, MM86b]. whorl [GT85a, Hut89]. whorls [HSV84]. wide [Pin83a]. widespread [Pet83]. width [Fer83, Sho85]. wigner [Has82, KT87b]. wild [Lib88]. will [CN85, HB84, Hin87]. wind [Lew82]. window [Cal83c, SC87b]. windows [SC86b]. Wing [Cor83b, AL81, MRB83, Yat86]. wing-areas [Cor83b]. Wing-beat [Cor83b]. winter [OL86, Omh87a, Omh87b]. Wiring [Sim80a]. withdrawal [Der84]. within [ACB81, Bar82a, DV83, Fil88, FW85, GT85b, Hej82, IM85, KM85, Kor82, Mur87, ST80]. without [Lew81c, LS89, Peu83, Pys80]. wobble [BS83, GPLC84, Shc89c]. work [LH89, Wil86a]. Worker [Bul81, AM81b, Omh88b]. Worker-queen [Bul81]. workers [Cra82b, Sta85, WL87]. working [Klä83, WGHY82]. world [Col82, KC85]. worm [WB89]. worm-like [WB89]. worms [KF83]. worthwhile [SC86a]. wound [SSF87, WFL86]. Wright [HR87].

X [Gri82f, vHU87]. X174 [Dug81, OK82]. Xenopus [KDW88, SCH84b]. Xylidine [WS81]. xylidines [WS81].

years [Har87b]. yeast [HN82, HKK82, HINT85, JCZ84, WB88]. yield [Hol83b, PDC84, Van84]. yield-density [Van84]. young [Hou87].

Z [Hop83]. Z-DNA [Hop83]. zebra [Bar81b]. zero [Sav81]. zeros [Bri85]. zigzag [Str83]. zinc [NU82]. Zipper [Cla83a]. Zn [NU82, TIK89b]. zonal [GT89]. zone [SEN82]. zones [Bas81, Mal84a]. Zoom [KT87b]. zwitterions [BVDA87]. zymogen [VCC^{+}88].

References

Anninos:1983:MMD

[AA83] Photios Anninos and Pangs Argyrakis. A mathematical model for the decay of short-term memory with age. Journal of Theoretical Biology, 102(2):191–197, May 21, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

            URL http://www.sciencedirect.com/science/article/pii/
            0022519383903594.

Ayeni:1984:TFE

[AA84] R. O. Ayeni and E. A. Akinrelere. On temperature factor and the effect of viscosity on blood temperature in the arteries. Journal of Theoretical Biology, 109(4):479–487, August 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

            URL http://www.sciencedirect.com/science/article/pii/
            S0022519384801538.

Arbib:1985:SMT

[AA85] Michael A. Arbib and Shun-Ichi Amari. Sensori-motor transformations in the brain (with a critique of the tensor theory of cerebellum). Journal of Theoretical Biology, 112(1):123–155, January 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

            (electronic). URL http://www.sciencedirect.com/science/
            article/pii/S002251938580120X.

Anninos:1987:NNM

[AAAT87] P. Anninos, G. Anogianakis, M. Apostolakis, and G. Thomas. A neural net model for the release of prolactin during pseudopregnancy in the rat. Journal of Theoretical Biology, 129(4):359–368, December 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387800188.

Aalto:1989:MMV

[Aal89] Erkki Aalto. The Moran model and validity of the diffusion approximation in population genetics. Journal of Theoretical Biology, 140(3):317–326, October 9, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938980089X.

Aldana:1989:RCM

[ABB89] Daniel H. Aldana, J. Douglas Brailsford, and Brian S. Bull. Red cell membrane crenation: a macromodel of the echinocyte I. Journal of Theoretical Biology, 140(2):185–192, September 22, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519389801274.

Atlan:1986:ECP

[ABEFS^{+}86] Henri Atlan, Esther Ben-Ezra, Françoise Fogelman-Soulie, Didier Pellegrin, and Gerard Weisbuch. Emergence of classification procedures in automata networks as a model for functional self-organization. Journal of Theoretical Biology, 120(3):371– 380, June 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519386802089.

Armstrong:1982:MSA

[ABS82] D. R. Armstrong, R. J. Breckenridge, and C. J. Suckling. Muscimol and some aza analogues: molecular orbital calculations relating to their potency as GABA mimetics. Journal of Theoretical Biology, 97(2):267–276, July 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382901035.

Abugov:1981:NLB

[Abu81] Robert Abugov. Non-linear benefits and the evolution of eusociality in the hymenoptera. Journal of Theoretical Biology, 88(4):733–742, February 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519381902472.

Abugov:1984:GSM

[Abu84] Robert Abugov. Genetic and strategic models for the evolution of mating systems. Journal of Theoretical Biology, 108(2):285– 294, May 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384800715.

Abugov:1985:GDF

[Abu85a] Robert Abugov. Genetics of Darwinian fitness. II. Pleiotropic selection on fertility and survival. Journal of Theoretical Biology, 114(1):109–125, May 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519385802599.

Abugov:1985:TCM

[Abu85b] Robert Abugov. Is there a cost of meiosis in life history? Journal of Theoretical Biology, 116(4):613–623, October 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800916.

Abugov:1986:GDF

[Abu86] Robert Abugov. Genetics of Darwinian fitness. III. A generalized approach to age structured selection and life history. Journal of Theoretical Biology, 122(3):311–323, October 7, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801230.

Abugov:1988:SSQ

[Abu88] Robert Abugov. A sex-specific quantitative genetic theory for life history and development. Journal of Theoretical Biology, 132(4):437–447, June 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388800833.

Aon:1989:RPC

[AC89] Miguel Antonio Aon and Sonia Cortassa. The regulation of plant cell growth: a bio-electromechanochemical model. Journal of Theoretical Biology, 138(4):429–456, June 22, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389800438.

Amar-Costesec:1981:SBE

[ACB81] Alain Amar-Costesec and Henri Beaufay. A structural basis of enzymic heterogeneity within liver endoplasmic reticulum. Journal of Theoretical Biology, 89(2):217–230, March 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251938190309X.

Adam:1980:PDM

[Ada80] G. Adam. Population-dynamical model of cell-density dependent growth regulation and aging of fibroblasts in vitro. Journal of Theoretical Biology, 84(2):233–257, May 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519380800063.

Alejandre-Duran:1987:SEP

[ADdACP87] E. Alejandre-Duran, D. F. de Andres-Cara, and A. Porras. Simulation of the effect of population size on the evolution of the recombination fraction. Journal of Theoretical Biology, 128(4):399– 405, October 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387801868.

Adams:1980:SAT

[ADG80] Virginia Dale Adams, Donald L. DeAngelis, and Robert A. Goldstein. Stability analysis of the time delay in a host-parasitoid model. Journal of Theoretical Biology, 83(1):43–62, March 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380903719.

Alonso-deFlorida:1986:SMK

[AdMM86] F. Alonso-deFlorida, A. A. Minzoni, and M. A. Morales. A synaptic model for the kindling effect. Journal of Theoretical Biology, 120(3):285–302, June 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386802016.

Aagren:1980:EVL

[ÅF80] Göran I. Ågren and Torbjörn Fagerström. On environmental variability and limits to similarity. Journal of Theoretical Biology, 82(3):401–404, February 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380902453.

Ahmed:1986:MCH

[AFHM86] A. K. Ahmed, S. Y. Fakhouri, J. B. Harness, and A. J. Mearns. Modelling of the control of heart rate by breathing using a kernel method. Journal of Theoretical Biology, 119(1):67–79, March 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386800510.

Arditi:1989:CPP

[AG89] Roger Arditi and Lev R. Ginzburg. Coupling in predator– prey dynamics: Ratio-dependence. Journal of Theoretical Biology, 139(3):311–326, August 9, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389802115.

Agmon:1985:DMM

[Agm85] Noam Agmon. A diffusion Michaelis–Menten mechanism: Continuous conformational change in enzymatic kinetics. Journal of Theoretical Biology, 113(4):711–717, April 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801880.

Aagren:1985:LPP

[Ågr85] Göran I. Ågren. Limits to plant production. Journal of Theoretical Biology, 113(1):89–92, March 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800783.

Agur:1985:RSP

[Agu85] Zvia Agur. Randomness, synchrony and population persistence. Journal of Theoretical Biology, 112(4):677–693, February 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385800539.

Alonso:1986:ESS

[AH86] G. L. Alonso and J. P. Hecht. Equilibrium and steady state thermodynamics of active transport systems studied on simple models simulating Ca^{2+} transport through sarcoplasmic reticulum membranes. Journal of Theoretical Biology, 120(2):191– 203, May 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519386801734.

Amidon:1981:PAR

[AHFH81] G. E. Amidon, N. F. H. Ho, A. B. French, and W. I. Higuchi. Predicted absorption rates with simultaneous bulk fluid flow in the intestinal tract. Journal of Theoretical Biology, 89(2):195– 210, March 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381903064.

Ahlgren:1981:MBO

[Ahl81] J. D. Ahlgren. On the molecular basis of oncogenesis. Journal of Theoretical Biology, 93(1):215–223, November 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381900667.

Aida:1988:IMO

[Aid88] Misako Aida. An ab initio molecular orbital study on the sequence-dependency of DNA conformation: An evaluation of intra- and inter-strand stacking interaction energy. Journal of Theoretical Biology, 130(3):327–335, February 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388800328.

Apple:1980:WMB

[AK80] M. S. Apple and M. A. Korostyshevskiy. Why many biological parameters are connected by power dependence. Journal of Theoretical Biology, 85(4):569–573, August 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380902581.

Ainsworth:1982:IUE

[AK82] Stanley Ainsworth and Julian Kinderlerer. The identification of unbranched enzyme kinetic mechanisms by computer analysis of qualitative information: Logic of the method. Journal of Theoretical Biology, 96(2):253–279, May 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902247.

Anninos:1984:NNM

[AK84a] Photios Anninos and Michael Kokkinidis. A neural net model for multiple memory domains. Journal of Theoretical Biology, 109(1): 95–110, July 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384801137.

Antonelli:1984:SPG

[AK84b] P. L. Antonelli and N. D. Kazarinoff. Starfish predation of a growing coral reef community. Journal of Theoretical Biology, 107(4):667–684, April 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801381.

Antonelli:1986:CSC

[AK86] P. L. Antonelli and N. D. Kazarinoff. Comments on starfish/ coral cycles after R. Bradbury et al. Journal of Theoretical Biology, 119(4):501–502, April 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801989.

Albery:1987:EECa

[AK87a] W. John Albery and Jeremy R. Knowles. Energetics of enzyme catalysis. I. Isotopic experiments, enzyme interconversion, and oversaturation. Journal of Theoretical Biology, 124(2):137–171, January 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251938780259X.

Albery:1987:EECb

[AK87b] W. John Albery and Jeremy R. Knowles. Energetics of enzyme catalysis. II. Oversaturation, case diagrams, reversible and irreversible behaviour. Journal of Theoretical Biology, 124(2):173– 189, January 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387802606.

Akintonwa:1984:TAE

[Aki84] Dunstan A. A. Akintonwa. Theoretical aspects of enzyme induction and inhibition leading to the reversal of resistance to biocides. Journal of Theoretical Biology, 106(1):79–87, January 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519384900286.

Akintonwa:1985:CBT

[Aki85] Dunstan A. A. Akintonwa. The correlation between theoretical and experimental biotransformation of 2-chloro-1-(2’,4’ dichlorophenyl) vinyl diethylphosphate (chlorfenvinphos) in the dog and rat. Journal of Theoretical Biology, 114(1):103–108, May 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385802587.

Akintonwa:1986:MBS

[Aki86] Dunstan A. A. Akintonwa. Mechanistic biotransformations of some amphetamine drugs. Journal of Theoretical Biology, 120(3): 303–308, June 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519386802028.

Anninos:1984:NNN

[AKS84] Photios Anninos, Michael Kokkinidis, and Andreas Skouras. Noisy neural nets exhibiting memory domains. Journal of Theoretical Biology, 109(4):581–594, August 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801605.

Abbott:1981:MGC

[AL81] Lois A. Abbott and Aristid Lindenmayer. Models for growth of clones in hexagonal cell arrangements: Applications in drosophila wing disc epithelia and plant epidermal tissues. Journal of Theoretical Biology, 90(4):495–514, June 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381903027.

Abdulnur:1983:TDR

[AL83] S. F. Abdulnur and K. Laki. A two-dimensional representation of relative orientations of α-helix residues. Journal of Theoretical Biology, 104(4):599–603, October 21, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383902485.

Awerbuch:1987:MMD

[AL87] Tamara E. Awerbuch and Levi Lustman. A mathematical model for determining minimal inhibitory concentrations (MICs) via diffusion assays. Journal of Theoretical Biology, 129(2):219–230, November 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387800140.

Alexander:1984:OSB

[Ale84] R. McN. Alexander. Optimum strengths for bones liable to fatigue and accidental fracture. Journal of Theoretical Biology, 109(4):621–636, August 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801629.

Alexander:1987:BCA

[Ale87] R. McN. Alexander. Bending of cylindrical animals with helical fibres in their skin or cuticle. Journal of Theoretical Biology, 124(1):97–110, January 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387802552.

Alfano:1983:ITM

[Alf83] Francis D. Alfano. An information-theory model of cellular survival, mutation and transformation and the relevance of this model to chemical carcinogenesis. Journal of Theoretical Biology, 104(4):605–616, October 21, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383902497.

Allen:1985:MPT

[All85a] Terence B. Allen. Multiple parathyroid tumors: a statistical method for distinguishing multiple adenoma from hyperplasia. Journal of Theoretical Biology, 112(3):479–492, February 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385800163.

Allison:1985:COM

[All85b] R. Donald Allison. Concerning the oscillatory mechanism of osuji for γ-glutamyl transpeptidase and the γ-glutamyl cycle. Journal of Theoretical Biology, 116(3):469–473, October 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385802824.

Almagor:1983:MAD

[Alm83] Hagai Almagor. A Markov analysis of DNA sequences. Journal of Theoretical Biology, 104(4):633–645, October 21, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383902515.

Almagor:1985:NDR

[Alm85] Hagai Almagor. Nucleotide distribution and the recognition of coding regions in DNA sequences: An information theory approach. Journal of Theoretical Biology, 117(1):127–136, November 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385801685.

Arenas:1981:BMI

[ALST81] C. A. Arenas, N. Lacourly, A. Sepúlveda, and J. C. Tohá. Biological molecular information and its representation. Journal of Theoretical Biology, 89(3):335–339, April 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381903556.

Altschul:1989:GCM

[Alt89] Stephen F. Altschul. Gap costs for multiple sequence alignment. Journal of Theoretical Biology, 138(3):297–309, June 3, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389801961.

Abugov:1981:RFS

[AM81a] Robert Abugov and Richard E. Michod. On the relation of family structured models and inclusive fitness models for kin selection. Journal of Theoretical Biology, 88(4):743–754, February 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381902484.

Aoki:1981:OTL

[AM81b] Kenichi Aoki and Michael Moody. One- and two-locus models of the origin of worker behavior in hymenoptera. Journal of Theoretical Biology, 89(3):449–474, April 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381903623.

Aihara:1982:TCO

[AM82] Kazuyuki Aihara and Gen Matsumoto. Temporally coherent organization and instabilities in squid giant axons. Journal of Theoretical Biology, 95(4):697–720, April 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382903496.

Arques:1987:PPM

[AM87] Didier G. Arquès and Christian J. Michel. A purine-pyrimidine motif verifying an identical presence in almost all gene taxonomic groups. Journal of Theoretical Biology, 128(4):457–461, October 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938780190X.

Aihara:1984:PNP

[AMI84] Kazuyuki Aihara, Gen Matsumoto, and Yuhji Ikegaya. Periodic and non-periodic responses of a periodically forced Hodgkin– Huxley oscillator. Journal of Theoretical Biology, 109(2):249– 269, July 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384800053.

Alberghina:1980:AMC

[AMM80] L. Alberghina, L. Mariani, and E. Martegani. Analysis of a model of cell cycle in eukaryotes. Journal of Theoretical Biology, 87(1):171–188, November 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251938090226X.

Amthor:1986:EAW

[Amt86] Jeffrey S. Amthor. Evolution and applicability of a whole plant respiration model. Journal of Theoretical Biology, 122(4):473– 490, October 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519386801874.

Anderson:1980:DHP

[And80a] R. M. Anderson. Depression of host population abundance by direct life cycle macroparasites. Journal of Theoretical Biology, 82(2):283–311, January 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380901046.

Andersson:1980:WTM

[And80b] Malte Andersson. Why are there so many threat displays? Journal of Theoretical Biology, 86(4):773–781, October 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380903100.

Anderson:1981:CER

[And81] John L. Anderson. Configurational effect on the reflection coefficient for rigid solutes in capillary pores. Journal of Theoretical Biology, 90(3):405–426, June 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381903210.

Andersson:1986:SSI

[And86] Malte Andersson. Sexual selection and the importance of viability differences: a reply. Journal of Theoretical Biology, 120(2):251– 254, May 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519386801783.

Anderson:1988:NMV

[And88] James Jay Anderson. A neural model for visual activation of startle behavior in fish. Journal of Theoretical Biology, 131(3):279– 288, April 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519388802261.

Anonymous:1980:Ab

[Ano80a] Anonymous. Addendum. Journal of Theoretical Biology, 85(1): 96–97, July 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519380902830.

Anonymous:1980:Aa

[Ano80b] Anonymous. Announcement. Journal of Theoretical Biology, 82(3):545, February 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002251938090257X.

Anonymous:1980:CA

[Ano80c] Anonymous. Conference announcement. Journal of Theoretical Biology, 85(4):815, August 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251938090274X.

Anonymous:1980:EBa

[Ano80d] Anonymous. Editorial Board. Journal of Theoretical Biology, 82(1):ifc, January 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519380900867.

Anonymous:1980:EBb

[Ano80e] Anonymous. Editorial Board. Journal of Theoretical Biology, 83 (1):ifc, March 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519380903689.

Anonymous:1980:EBc

[Ano80f] Anonymous. Editorial Board. Journal of Theoretical Biology, 85 (1):ifc, July 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519380902751.

Anonymous:1980:EBd

[Ano80g] Anonymous. Editorial Board. Journal of Theoretical Biology, 86(1):ifc, September 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519380900624.

Anonymous:1980:EBe

[Ano80h] Anonymous. Editorial Board. Journal of Theoretical Biology, 87(1):ifc, November 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519380902143.

Anonymous:1980:E

[Ano80i] Anonymous. Errata. Journal of Theoretical Biology, 87(3):623, December 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519380902404.

Anonymous:1980:MEN

[Ano80j] Anonymous. On the mechanics and energetics of the nectar feeding in butterflies: J. G. Kingsolver and T. L. Daniel, J. theor. Biol. (1979) 76, 167–179. Journal of Theoretical Biology, 86(4): I, October 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519380903136.

Anonymous:1980:PJa

[Ano80k] Anonymous. Pages 1–169 (7 January 1980). Journal of Theoretical Biology, 82(1):??, January 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1980:PJe

[Ano80l] Anonymous. Pages 1–169 (7 July 1980). Journal of Theoretical Biology, 85(1):??, July 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1980:PMc

[Ano80m] Anonymous. Pages 1–179 (7 May 1980). Journal of Theoretical Biology, 84(1):??, May 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1980:PMa

[Ano80n] Anonymous. Pages 1–194 (7 March 1980). Journal of Theoretical Biology, 83(1):??, March 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1980:PSa

[Ano80o] Anonymous. Pages 1–201 (7 September 1980). Journal of Theoretical Biology, 86(1):??, September 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1980:PNa

[Ano80p] Anonymous. Pages 1–209 (7 November 1980). Journal of Theoretical Biology, 87(1):??, November 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1980:PJb

[Ano80q] Anonymous. Pages 171–345 (21 January 1980). Journal of Theoretical Biology, 82(2):??, January 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1980:PJf

[Ano80r] Anonymous. Pages 171–397 (21 July 1980). Journal of Theoretical Biology, 85(2):??, July 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1980:PMd

[Ano80s] Anonymous. Pages 181–400 (21 May 1980). Journal of Theoretical Biology, 84(2):??, May 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1980:PMb

[Ano80t] Anonymous. Pages 195–375 (21 March 1980). Journal of Theoretical Biology, 83(2):??, March 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1980:PSb

[Ano80u] Anonymous. Pages 203–408 (21 September 1980). Journal of Theoretical Biology, 86(2):??, September 21, 1980. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1980:PNb

[Ano80v] Anonymous. Pages 211–419 (21 November 1980). Journal of Theoretical Biology, 87(2):??, November 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1980:PFa

[Ano80w] Anonymous. Pages 347–545 (7 February 1980). Journal of Theoretical Biology, 82(3):??, February 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1980:PAa

[Ano80x] Anonymous. Pages 377–535 (7 April 1980). Journal of Theoretical Biology, 83(3):??, April 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1980:PAc

[Ano80y] Anonymous. Pages 399–567 (7 August 1980). Journal of Theoretical Biology, 85(3):??, August 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1980:PJc

[Ano80z] Anonymous. Pages 401–588 (7 June 1980). Journal of Theoretical Biology, 84(3):??, June 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1980:POa

[Ano80-27] Anonymous. Pages 409–606 (7 October 1980). Journal of Theoretical Biology, 86(3):??, October 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1980:PD

[Ano80-28] Anonymous. Pages 421–623 (7 December 1980). Journal of Theoretical Biology, 87(3):??, December 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1980:PAb

[Ano80-29] Anonymous. Pages 537–703 (21 April 1980). Journal of Theoretical Biology, 83(4):??, April 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1980:PFb

[Ano80-30] Anonymous. Pages 547–688 (21 February 1980). Journal of Theoretical Biology, 82(4):??, February 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1980:PAd

[Ano80-31] Anonymous. Pages 569–815 (21 August 1980). Journal of Theoretical Biology, 85(4):??, August 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1980:PJd

[Ano80-32] Anonymous. Pages 589–800 (21 June 1980). Journal of Theoretical Biology, 84(4):??, June 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1980:POb

[Ano80-33] Anonymous. Pages 607–815 (21 October 1980). Journal of Theoretical Biology, 86(4):??, October 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1980:P

[Ano80-34] Anonymous. Pages 623–821 (1980). Journal of Theoretical Biology, 87(4):??, ???? 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1980:RMM

[Ano80-35] Anonymous. Reply to Martindale — on the multivariate analysis of avian vocalization. Journal of Theoretical Biology, 83(1):193– 194, March 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519380903793.

Anonymous:1981:Aa

[Ano81a] Anonymous. Addendum. Journal of Theoretical Biology, 89(2):214–215, March 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519381903088.

Anonymous:1981:Ab

[Ano81b] Anonymous. Announcement. Journal of Theoretical Biology, 92(3):343, October 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519381902976.

Anonymous:1981:EBa

[Ano81c] Anonymous. Editorial Board. Journal of Theoretical Biology, 88(1):ifc, January 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519381903258.

Anonymous:1981:EBb

[Ano81d] Anonymous. Editorial Board. Journal of Theoretical Biology, 89 (1):ifc, March 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381901740.

Anonymous:1981:EBc

[Ano81e] Anonymous. Editorial Board. Journal of Theoretical Biology, 90 (1):ifc, May 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381901181.

Anonymous:1981:EBd

[Ano81f] Anonymous. Editorial Board. Journal of Theoretical Biology, 91 (1):ifc, July 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381903684.

Anonymous:1981:EBe

[Ano81g] Anonymous. Editorial Board. Journal of Theoretical Biology, 92(1):ifc, September 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519381901557.

Anonymous:1981:EBf

[Ano81h] Anonymous. Editorial Board. Journal of Theoretical Biology, 93(1):CO2, November 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519381900515.

Anonymous:1981:EML

[Ano81i] Anonymous. Enlarged model of lambda phage ontogenesis: G. Kh. Kananyan, V. A. Ratner and R. N. Tchuraev. Journal of Theoretical Biology, 90(2):299, May 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381900497.

Anonymous:1981:Ea

[Ano81j] Anonymous. Erratum. Journal of Theoretical Biology, 90(1): 159, May 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381901296.

Anonymous:1981:Ec

[Ano81k] Anonymous. Erratum. Journal of Theoretical Biology, 91(3): 525, August 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381902721.

Anonymous:1981:Ed

[Ano81l] Anonymous. Erratum. Journal of Theoretical Biology, 92(3):341, October 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381902964.

Anonymous:1981:PMc

[Ano81m] Anonymous. Pages 1–159 (7 May 1981). Journal of Theoretical Biology, 90(1):??, May 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1981:PMa

[Ano81n] Anonymous. Pages 1–194 (7 March 1981). Journal of Theoretical Biology, 89(1):??, March 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1981:PJa

[Ano81o] Anonymous. Pages 1–199 (7 January 1981). Journal of Theoretical Biology, 88(1):??, January 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1981:PJe

[Ano81p] Anonymous. Pages 1–239 (7 July 1981). Journal of Theoretical Biology, 91(1):??, July 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1981:PSa

[Ano81q] Anonymous. Pages 1–96 (7 September 1981). Journal of Theoretical Biology, 92(1):??, September 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1981:PMd

[Ano81r] Anonymous. Pages 161–315 (21 May 1981). Journal of Theoretical Biology, 90(2):??, May 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1981:POa

[Ano81s] Anonymous. Pages 185–343 (7 October 1981). Journal of Theoretical Biology, 92(3):??, October 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1981:PMb

[Ano81t] Anonymous. Pages 195–333 (21 March 1981). Journal of Theoretical Biology, 89(2):??, March 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1981:PJb

[Ano81u] Anonymous. Pages 201–392 (21 January 1981). Journal of Theoretical Biology, 88(2):??, January 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1981:PJf

[Ano81v] Anonymous. Pages 241–382 (21 July 1981). Journal of Theoretical Biology, 91(2):??, July 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1981:PN

[Ano81w] Anonymous. Pages 279–497 (21 November 1981). Journal of Theoretical Biology, 93(2):??, November 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1981:PJc

[Ano81x] Anonymous. Pages 317–443 (7 June 1981). Journal of Theoretical Biology, 90(3):??, June 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1981:PAa

[Ano81y] Anonymous. Pages 335–537 (7 April 1981). Journal of Theoretical Biology, 89(3):??, April 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1981:POb

[Ano81z] Anonymous. Pages 345–503 (21 October 1981). Journal of Theoretical Biology, 92(4):??, October 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1981:PAc

[Ano81-27] Anonymous. Pages 383–525 (7 August 1981). Journal of Theoretical Biology, 91(3):??, August 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1981:PFa

[Ano81-28] Anonymous. Pages 393–597 (7 February 1981). Journal of Theoretical Biology, 88(3):??, February 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1981:PJd

[Ano81-29] Anonymous. Pages 445–571 (21 June 1981). Journal of Theoretical Biology, 90(4):??, June 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1981:PDa

[Ano81-30] Anonymous. Pages 499–696 (7 December 1981). Journal of Theoretical Biology, 93(3):??, December 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1981:PAd

[Ano81-31] Anonymous. Pages 527–654 (21 August 1981). Journal of Theoretical Biology, 91(4):??, August 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1981:PAb

[Ano81-32] Anonymous. Pages 539–717 (21 April 1981). Journal of Theoretical Biology, 89(4):??, April 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1981:PFb

[Ano81-33] Anonymous. Pages 599–760 (21 February 1981). Journal of Theoretical Biology, 88(4):??, February 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1981:PDb

[Ano81-34] Anonymous. Pages 697–1045 (21 December 1981). Journal of Theoretical Biology, 93(4):??, December 21, 1981. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1981:PSb

[Ano81-35] Anonymous. Pages 97–184 (21 September 1981). Journal of Theoretical Biology, 92(2):??, September 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1981:PCN

[Ano81-36] Anonymous. Pages CO2, 1-277 (7 November 1981). Journal of Theoretical Biology, 93(1):??, November 7, 1981. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1981:PIB

[Ano81-37] Anonymous. Possible interactions between alkoxy amphetamines and brain serotonin receptors. Journal of Theoretical Biology, 91(1):237–239, July 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519381903842.

Anonymous:1982:A

[Ano82a] Anonymous. Announcement. Journal of Theoretical Biology, 96 (1):i–iii, May 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382901618.

Anonymous:1982:EBa

[Ano82b] Anonymous. Editorial Board. Journal of Theoretical Biology, 94(1):ifc, January 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519382903253.

Anonymous:1982:EBb

[Ano82c] Anonymous. Editorial Board. Journal of Theoretical Biology, 95 (1):ifc, March 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251938290282X.

Anonymous:1982:EBc

[Ano82d] Anonymous. Editorial Board. Journal of Theoretical Biology, 96 (1):ifc, May 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382901497.

Anonymous:1982:EBd

[Ano82e] Anonymous. Editorial Board. Journal of Theoretical Biology, 97 (1):ifc, July 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382902703.

Anonymous:1982:EBe

[Ano82f] Anonymous. Editorial Board. Journal of Theoretical Biology, 98(1):CO2, September 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519382900510.

Anonymous:1982:EBf

[Ano82g] Anonymous. Editorial Board. Journal of Theoretical Biology, 99(1):ifc, November 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519382903836.

Anonymous:1982:E

[Ano82h] Anonymous. Erratum. Journal of Theoretical Biology, 98(4):707, October 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382901485.

Anonymous:1982:PMc

[Ano82i] Anonymous. Pages 1–106 (7 May 1982). Journal of Theoretical Biology, 96(1):??, May 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1982:PJe

[Ano82j] Anonymous. Pages 1–127 (7 July 1982). Journal of Theoretical Biology, 97(1):??, July 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1982:PSa

[Ano82k] Anonymous. Pages 1–169 (7 September 1982). Journal of Theoretical Biology, 98(1):??, September 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1982:PNa

[Ano82l] Anonymous. Pages 1–201 (7 November 1982). Journal of Theoretical Biology, 99(1):??, November 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1982:PMa

[Ano82m] Anonymous. Pages 1–223 (7 March 1982). Journal of Theoretical Biology, 95(1):??, March 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1982:PJa

[Ano82n] Anonymous. Pages 1–224 (7 January 1982). Journal of Theoretical Biology, 94(1):??, January 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1982:PMd

[Ano82o] Anonymous. Pages 107–326 (21 May 1982). Journal of Theoretical Biology, 96(2):??, May 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1982:PJf

[Ano82p] Anonymous. Pages 129–349 (21 July 1982). Journal of Theoretical Biology, 97(2):??, July 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1982:PSb

[Ano82q] Anonymous. Pages 171–368 (21 September 1982). Journal of Theoretical Biology, 98(2):??, September 21, 1982. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1982:PNb

[Ano82r] Anonymous. Pages 203–405 (21 November 1982). Journal of Theoretical Biology, 99(2):??, November 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1982:PMb

[Ano82s] Anonymous. Pages 225–414 (21 March 1982). Journal of Theoretical Biology, 95(2):??, March 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1982:PJb

[Ano82t] Anonymous. Pages 225–518 (21 January 1982). Journal of Theoretical Biology, 94(2):??, January 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1982:PJc

[Ano82u] Anonymous. Pages 327–516 (7 June 1982). Journal of Theoretical Biology, 96(3):??, June 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1982:PAc

[Ano82v] Anonymous. Pages 351–542 (7 August 1982). Journal of Theoretical Biology, 97(3):??, August 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1982:POa

[Ano82w] Anonymous. Pages 369–548 (7 October 1982). Journal of Theoretical Biology, 98(3):??, October 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1982:PDa

[Ano82x] Anonymous. Pages 407–628 (7 December 1982). Journal of Theoretical Biology, 99(3):??, December 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1982:PAa

[Ano82y] Anonymous. Pages 415–606 (7 April 1982). Journal of Theoretical Biology, 95(3):??, April 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1982:PJd

[Ano82z] Anonymous. Pages 517–757 (21 June 1982). Journal of Theoretical Biology, 96(4):??, June 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1982:PFa

[Ano82-27] Anonymous. Pages 519–741 (7 February 1982). Journal of Theoretical Biology, 94(3):??, February 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1982:PAd

[Ano82-28] Anonymous. Pages 543–716 (21 August 1982). Journal of Theoretical Biology, 97(4):??, August 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1982:POb

[Ano82-29] Anonymous. Pages 549–707 (21 October 1982). Journal of Theoretical Biology, 98(4):??, October 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1982:PAb

[Ano82-30] Anonymous. Pages 607–828 (21 April 1982). Journal of Theoretical Biology, 95(4):??, April 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1982:PDb

[Ano82-31] Anonymous. Pages 629–838 (21 December 1982). Journal of Theoretical Biology, 99(4):??, December 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1982:PFb

[Ano82-32] Anonymous. Pages 743–997 (21 February 1982). Journal of Theoretical Biology, 94(4):??, February 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1983:EBa

[Ano83a] Anonymous. Editorial Board. Journal of Theoretical Biology, 100(1):ifc, January 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519383900899.

Anonymous:1983:EBb

[Ano83b] Anonymous. Editorial Board. Journal of Theoretical Biology, 101 (1):ifc, March 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383902692.

Anonymous:1983:EBc

[Ano83c] Anonymous. Editorial Board. Journal of Theoretical Biology, 102 (1):ifc, May 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383902588.

Anonymous:1983:EBd

[Ano83d] Anonymous. Editorial Board. Journal of Theoretical Biology, 103 (1):ifc, July 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383901935.

Anonymous:1983:EBe

[Ano83e] Anonymous. Editorial Board. Journal of Theoretical Biology, 104(1):ifc, September 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002251938390396X.

Anonymous:1983:EBf

[Ano83f] Anonymous. Editorial Board. Journal of Theoretical Biology, 105(1):ifc, November 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519383904198.

Anonymous:1983:Ea

[Ano83g] Anonymous. Erratum. Journal of Theoretical Biology, 101(2): 325, March 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383903454.

Anonymous:1983:Ec

[Ano83h] Anonymous. Erratum. Journal of Theoretical Biology, 101(4): 661, April 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251938390022X.

Anonymous:1983:Ed

[Ano83i] Anonymous. Erratum. Journal of Theoretical Biology, 104 (1):159, September 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519383904071.

Anonymous:1983:Ee

[Ano83j] Anonymous. Erratum. Journal of Theoretical Biology, 104 (2):319, September 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519383904186.

Anonymous:1983:PMa

[Ano83k] Anonymous. Pages 1–150 (7 March 1983). Journal of Theoretical Biology, 101(1):??, March 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1983:PSa

[Ano83l] Anonymous. Pages 1–159 (7 September 1983). Journal of Theoretical Biology, 104(1):??, September 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1983:PJe

[Ano83m] Anonymous. Pages 1–172 (7 July 1983). Journal of Theoretical Biology, 103(1):??, July 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1983:PMc

[Ano83n] Anonymous. Pages 1–173 (7 May 1983). Journal of Theoretical Biology, 102(1):??, May 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1983:PN

[Ano83o] Anonymous. Pages 1–181 (7 November 1983). Journal of Theoretical Biology, 105(1):??, November 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1983:PJa

[Ano83p] Anonymous. Pages 1–184 (7 January 1983). Journal of Theoretical Biology, 100(1):??, January 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1983:PMb

[Ano83q] Anonymous. Pages 151–325 (21 March 1983). Journal of Theoretical Biology, 101(2):??, March 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1983:PSb

[Ano83r] Anonymous. Pages 161–319 (21 September 1983). Journal of Theoretical Biology, 104(2):??, September 21, 1983. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1983:PJf

[Ano83s] Anonymous. Pages 173–332 (21 July 1983). Journal of Theoretical Biology, 103(2):??, July 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1983:PMd

[Ano83t] Anonymous. Pages 175–340 (21 May 1983). Journal of Theoretical Biology, 102(2):??, May 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1983:PJb

[Ano83u] Anonymous. Pages 185–357 (21 January 1983). Journal of Theoretical Biology, 100(2):??, January 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1983:P

[Ano83v] Anonymous. Pages 185–378 (1983). Journal of Theoretical Biology, 105(2):??, ???? 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1983:POa

[Ano83w] Anonymous. Pages 321–471 (7 October 1983). Journal of Theoretical Biology, 104(3):??, October 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1983:PAa

[Ano83x] Anonymous. Pages 327–502 (7 April 1983). Journal of Theoretical Biology, 101(3):??, April 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1983:PAc

[Ano83y] Anonymous. Pages 333–480 (7 August 1983). Journal of Theoretical Biology, 103(3):??, August 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1983:PJc

[Ano83z] Anonymous. Pages 339–467 (7 June 1983). Journal of Theoretical Biology, 102(3):??, June 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1983:PFa

[Ano83-27] Anonymous. Pages 359–550 (7 February 1983). Journal of Theoretical Biology, 100(3):??, February 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1983:PDa

[Ano83-28] Anonymous. Pages 379–543 (7 December 1983). Journal of Theoretical Biology, 105(3):??, December 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1983:PJd

[Ano83-29] Anonymous. Pages 469–628 (21 June 1983). Journal of Theoretical Biology, 102(4):??, June 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1983:POb

[Ano83-30] Anonymous. Pages 473–702 (21 October 1983). Journal of Theoretical Biology, 104(4):??, October 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1983:PAd

[Ano83-31] Anonymous. Pages 481–652 (21 August 1983). Journal of Theoretical Biology, 103(4):??, August 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1983:PAb

[Ano83-32] Anonymous. Pages 503–661 (21 April 1983). Journal of Theoretical Biology, 101(4):??, April 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1983:PDb

[Ano83-33] Anonymous. Pages 545–721 (21 December 1983). Journal of Theoretical Biology, 105(4):??, December 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1983:PFb

[Ano83-34] Anonymous. Pages 551–707 (21 February 1983). Journal of Theoretical Biology, 100(4):??, February 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1984:A

[Ano84a] Anonymous. Addendum. Journal of Theoretical Biology, 110(2):312, September 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519384800612.

Anonymous:1984:EBa

[Ano84b] Anonymous. Editorial Board. Journal of Theoretical Biology, 106(1):ifc, January 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519384900225.

Anonymous:1984:EBb

[Ano84c] Anonymous. Editorial Board. Journal of Theoretical Biology, 107 (1):i, March 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384801174.

Anonymous:1984:EBc

[Ano84d] Anonymous. Editorial Board. Journal of Theoretical Biology, 108 (1):i, May 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384801630.

Anonymous:1984:EBd

[Ano84e] Anonymous. Editorial Board. Journal of Theoretical Biology, 109 (1):i, July 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384801058.

Anonymous:1984:EBe

[Ano84f] Anonymous. Editorial Board. Journal of Theoretical Biology, 110(1):i, September 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519384800089.

Anonymous:1984:EBf

[Ano84g] Anonymous. Editorial Board. Journal of Theoretical Biology, 111(1):i, November 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519384801903.

Anonymous:1984:IVa

[Ano84h] Anonymous. Index to volume 107. Journal of Theoretical Biology, 107(1):vii–xiv, March 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801186.

Anonymous:1984:IVb

[Ano84i] Anonymous. Index to volume 108. Journal of Theoretical Biology, 108(1):vii–xiii, May 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519384801642.

Anonymous:1984:IVc

[Ano84j] Anonymous. Index to volume 109. Journal of Theoretical Biology, 109(1):vi–xii, July 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002251938480106X.

Anonymous:1984:IVd

[Ano84k] Anonymous. Index to volume 110. Journal of Theoretical Biology, 110(1):vii–ixv, September 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384800090.

Anonymous:1984:IVe

[Ano84l] Anonymous. Index to volume 111. Journal of Theoretical Biology, 111(1):vii–xvi, November 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801915.

Anonymous:1984:PSa

[Ano84m] Anonymous. Pages 1–133 (7 September 1984). Journal of Theoretical Biology, 110(1):??, September 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1984:PJe

[Ano84n] Anonymous. Pages 1–156 (7 July 1984). Journal of Theoretical Biology, 109(1):??, July 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1984:PMc

[Ano84o] Anonymous. Pages 1–162 (7 May 1984). Journal of Theoretical Biology, 108(1):??, May 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1984:PMa

[Ano84p] Anonymous. Pages 1–177 (7 March 1984). Journal of Theoretical Biology, 107(1):??, March 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1984:PNa

[Ano84q] Anonymous. Pages 1–204 (7 November 1984). Journal of Theoretical Biology, 111(1):??, November 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1984:PJa

[Ano84r] Anonymous. Pages 1–87 (7 January 1984). Journal of Theoretical Biology, 106(1):??, January 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1984:PSb

[Ano84s] Anonymous. Pages 135–314 (21 September 1984). Journal of Theoretical Biology, 110(2):??, September 21, 1984. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1984:PJf

[Ano84t] Anonymous. Pages 157–298 (21 July 1984). Journal of Theoretical Biology, 109(2):??, July 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1984:PMd

[Ano84u] Anonymous. Pages 163–318 (21 May 1984). Journal of Theoretical Biology, 108(2):??, May 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1984:PMb

[Ano84v] Anonymous. Pages 179–343 (21 March 1984). Journal of Theoretical Biology, 107(2):??, March 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1984:PNb

[Ano84w] Anonymous. Pages 205–423 (21 November 1984). Journal of Theoretical Biology, 111(2):??, November 21, 1984. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1984:PFa

[Ano84x] Anonymous. Pages 245–439 (7 February 1984). Journal of Theoretical Biology, 106(3):??, February 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1984:PAc

[Ano84y] Anonymous. Pages 299–474 (7 August 1984). Journal of Theoretical Biology, 109(3):??, August 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1984:POa

[Ano84z] Anonymous. Pages 315–503 (7 October 1984). Journal of Theoretical Biology, 110(3):??, October 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1984:PJc

[Ano84-27] Anonymous. Pages 319–479 (7 June 1984). Journal of Theoretical Biology, 108(3):??, June 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1984:PAa

[Ano84-28] Anonymous. Pages 345–520 (7 April 1984). Journal of Theoretical Biology, 107(3):??, April 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1984:PDa

[Ano84-29] Anonymous. Pages 425–613 (7 December 1984). Journal of Theoretical Biology, 111(3):??, December 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1984:PFb

[Ano84-30] Anonymous. Pages 441–615 (21 February 1984). Journal of Theoretical Biology, 106(4):??, February 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1984:PAd

[Ano84-31] Anonymous. Pages 475–636 (21 August 1984). Journal of Theoretical Biology, 109(4):??, August 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1984:PJd

[Ano84-32] Anonymous. Pages 481–674 (21 June 1984). Journal of Theoretical Biology, 108(4):??, June 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1984:POb

[Ano84-33] Anonymous. Pages 505–690 (21 October 1984). Journal of Theoretical Biology, 110(4):??, October 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1984:PAb

[Ano84-34] Anonymous. Pages 521–704 (21 April 1984). Journal of Theoretical Biology, 107(4):??, April 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1984:PDb

[Ano84-35] Anonymous. Pages 615–819 (21 December 1984). Journal of Theoretical Biology, 111(4):??, December 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1984:PJb

[Ano84-36] Anonymous. Pages 89–243 (21 January 1984). Journal of Theoretical Biology, 106(2):??, January 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1985:EBa

[Ano85a] Anonymous. Editorial Board. Journal of Theoretical Biology, 112 (1):i, January 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385801120.

Anonymous:1985:EBb

[Ano85b] Anonymous. Editorial Board. Journal of Theoretical Biology, 113 (1):i, March 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385800709.

Anonymous:1985:EBc

[Ano85c] Anonymous. Editorial Board. Journal of Theoretical Biology, 114 (1):i, May 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385802484.

Anonymous:1985:EBd

[Ano85d] Anonymous. Editorial Board. Journal of Theoretical Biology, 115 (1):i, July 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385800011.

Anonymous:1985:EBe

[Ano85e] Anonymous. Editorial Board. Journal of Theoretical Biology, 116(1):i, September 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519385801259.

Anonymous:1985:EBf

[Ano85f] Anonymous. Editorial Board. Journal of Theoretical Biology, 117(1):i, November 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519385801594.

Anonymous:1985:E

[Ano85g] Anonymous. Erratum. Journal of Theoretical Biology, 112 (4):755, February 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519385800588.

Anonymous:1985:IVa

[Ano85h] Anonymous. Index to volume 112. Journal of Theoretical Biology, 112(1):viii–xvi, January 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801132.

Anonymous:1985:IVb

[Ano85i] Anonymous. Index to volume 113. Journal of Theoretical Biology, 113(1):vii–xiv, March 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800710.

Anonymous:1985:IVc

[Ano85j] Anonymous. Index to volume 114. Journal of Theoretical Biology, 114(1):vii–xv, May 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519385802496.

Anonymous:1985:IVd

[Ano85k] Anonymous. Index to volume 115. Journal of Theoretical Biology, 115(1):vii–xiii, July 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519385800023.

Anonymous:1985:IVe

[Ano85l] Anonymous. Index to volume 116. Journal of Theoretical Biology, 116(1):vii–xiv, September 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801260.

Anonymous:1985:IVf

[Ano85m] Anonymous. Index to volume 117. Journal of Theoretical Biology, 117(1):vi–xiii, November 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801600.

Anonymous:1985:PSa

[Ano85n] Anonymous. Pages 1–159 (7 September 1985). Journal of Theoretical Biology, 116(1):??, September 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1985:PJe

[Ano85o] Anonymous. Pages 1–160 (7 July 1985). Journal of Theoretical Biology, 115(1):??, July 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1985:PNa

[Ano85p] Anonymous. Pages 1–160 (7 November 1985). Journal of Theoretical Biology, 117(1):??, November 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1985:PMc

[Ano85q] Anonymous. Pages 1–192 (7 May 1985). Journal of Theoretical Biology, 114(1):??, May 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1985:PJa

[Ano85r] Anonymous. Pages 1–220 (7 January 1985). Journal of Theoretical Biology, 112(1):??, January 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1985:PMa

[Ano85s] Anonymous. Pages 1–223 (7 March 1985). Journal of Theoretical Biology, 113(1):??, March 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1985:PJf

[Ano85t] Anonymous. Pages 161–317 (21 July 1985). Journal of Theoretical Biology, 115(2):??, July 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1985:PNb

[Ano85u] Anonymous. Pages 161–317 (21 November 1985). Journal of Theoretical Biology, 117(2):??, November 21, 1985. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1985:PSb

[Ano85v] Anonymous. Pages 161–320 (21 September 1985). Journal of Theoretical Biology, 116(2):??, September 21, 1985. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1985:PMd

[Ano85w] Anonymous. Pages 193–352 (21 May 1985). Journal of Theoretical Biology, 114(2):??, May 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1985:PJb

[Ano85x] Anonymous. Pages 221–444 (21 January 1985). Journal of Theoretical Biology, 112(2):??, January 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1985:PMb

[Ano85y] Anonymous. Pages 225–405 (21 March 1985). Journal of Theoretical Biology, 113(2):??, March 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1985:PAc

[Ano85z] Anonymous. Pages 319–476 (7 August 1985). Journal of Theoretical Biology, 115(3):??, August 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1985:PDa

[Ano85-27] Anonymous. Pages 319–508 (7 December 1985). Journal of Theoretical Biology, 117(3):??, December 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1985:POa

[Ano85-28] Anonymous. Pages 321–478 (7 October 1985). Journal of Theoretical Biology, 116(3):??, October 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1985:PJc

[Ano85-29] Anonymous. Pages 353–510 (7 June 1985). Journal of Theoretical Biology, 114(3):??, June 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1985:PAa

[Ano85-30] Anonymous. Pages 407–597 (7 April 1985). Journal of Theoretical Biology, 113(3):??, April 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1985:PFa

[Ano85-31] Anonymous. Pages 445–666 (7 February 1985). Journal of Theoretical Biology, 112(3):??, February 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1985:PAd

[Ano85-32] Anonymous. Pages 477–634 (21 August 1985). Journal of Theoretical Biology, 115(4):??, August 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1985:POb

[Ano85-33] Anonymous. Pages 479–636 (21 October 1985). Journal of Theoretical Biology, 116(4):??, October 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1985:PDb

[Ano85-34] Anonymous. Pages 509–699 (21 December 1985). Journal of Theoretical Biology, 117(4):??, December 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1985:PJd

[Ano85-35] Anonymous. Pages 511–667 (21 June 1985). Journal of Theoretical Biology, 114(4):??, June 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1985:PAb

[Ano85-36] Anonymous. Pages 599–790 (21 April 1985). Journal of Theoretical Biology, 113(4):??, April 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1985:PFb

[Ano85-37] Anonymous. Pages 667–881 (21 February 1985). Journal of Theoretical Biology, 112(4):??, February 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1986:A

[Ano86a] Anonymous. Announcement. Journal of Theoretical Biology, 118(4):i–ii, February 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519386801680.

Anonymous:1986:EBa

[Ano86b] Anonymous. Editorial Board. Journal of Theoretical Biology, 118 (1):i, January 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519386800017.

Anonymous:1986:EBb

[Ano86c] Anonymous. Editorial Board. Journal of Theoretical Biology, 119 (1):i, March 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519386800455.

Anonymous:1986:EBc

[Ano86d] Anonymous. Editorial Board. Journal of Theoretical Biology, 120 (1):i, May 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519386800133.

Anonymous:1986:EBd

[Ano86e] Anonymous. Editorial Board. Journal of Theoretical Biology, 121 (1):i, July 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519386800236.

Anonymous:1986:EBe

[Ano86f] Anonymous. Editorial Board. Journal of Theoretical Biology, 122(1):i, September 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519386802181.

Anonymous:1986:EBf

[Ano86g] Anonymous. Editorial Board. Journal of Theoretical Biology, 123(1):i, November 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519386802296.

Anonymous:1986:IVa

[Ano86h] Anonymous. Index to volume 118. Journal of Theoretical Biology, 118(1):vi–xi, January 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386800029.

Anonymous:1986:IVb

[Ano86i] Anonymous. Index to volume 119. Journal of Theoretical Biology, 119(1):vi–xi, March 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519386800467.

Anonymous:1986:IVc

[Ano86j] Anonymous. Index to volume 120. Journal of Theoretical Biology, 120(1):vi–xi, May 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519386800145.

Anonymous:1986:IVd

[Ano86k] Anonymous. Index to volume 121. Journal of Theoretical Biology, 121(1):vi–xii, July 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519386800248.

Anonymous:1986:IVe

[Ano86l] Anonymous. Index to volume 122. Journal of Theoretical Biology, 122(1):vi–xi, September 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386802193.

Anonymous:1986:IVf

[Ano86m] Anonymous. Index to volume 123. Journal of Theoretical Biology, 123(1):vi–xi, November 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386802302.

Anonymous:1986:PMa

[Ano86n] Anonymous. Pages 1–124 (7 March 1986). Journal of Theoretical Biology, 119(1):??, March 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1986:PSa

[Ano86o] Anonymous. Pages 1–124 (7 September 1986). Journal of Theoretical Biology, 122(1):??, September 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1986:PJa

[Ano86p] Anonymous. Pages 1–125 (7 January 1986). Journal of Theoretical Biology, 118(1):??, January 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1986:PNa

[Ano86q] Anonymous. Pages 1–126 (7 November 1986). Journal of Theoretical Biology, 123(1):??, November 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1986:PJe

[Ano86r] Anonymous. Pages 1–127 (7 July 1986). Journal of Theoretical Biology, 121(1):??, July 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1986:PMc

[Ano86s] Anonymous. Pages 1–127 (7 May 1986). Journal of Theoretical Biology, 120(1):??, May 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1986:PMb

[Ano86t] Anonymous. Pages 125–249 (21 March 1986). Journal of Theoretical Biology, 119(2):??, March 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1986:PSb

[Ano86u] Anonymous. Pages 125–249 (21 September 1986). Journal of Theoretical Biology, 122(2):??, September 21, 1986. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1986:PJb

[Ano86v] Anonymous. Pages 127–250 (21 January 1986). Journal of Theoretical Biology, 118(2):??, January 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1986:PNb

[Ano86w] Anonymous. Pages 127–250 (21 November 1986). Journal of Theoretical Biology, 123(2):??, November 21, 1986. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1986:PJf

[Ano86x] Anonymous. Pages 129–254 (21 July 1986). Journal of Theoretical Biology, 121(2):??, July 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1986:PMd

[Ano86y] Anonymous. Pages 129–254 (21 May 1986). Journal of Theoretical Biology, 120(2):??, May 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1986:POa

[Ano86z] Anonymous. Pages 251–374 (7 October 1986). Journal of Theoretical Biology, 122(3):??, October 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1986:PDa

[Ano86-27] Anonymous. Pages 251–375 (7 December 1986). Journal of Theoretical Biology, 123(3):??, December 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1986:PFa

[Ano86-28] Anonymous. Pages 251–375 (7 February 1986). Journal of Theoretical Biology, 118(3):??, February 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1986:PAa

[Ano86-29] Anonymous. Pages 251–378 (7 April 1986). Journal of Theoretical Biology, 119(3):??, April 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1986:PAc

[Ano86-30] Anonymous. Pages 255–379 (7 August 1986). Journal of Theoretical Biology, 121(3):??, August 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1986:PJc

[Ano86-31] Anonymous. Pages 255–380 (7 June 1986). Journal of Theoretical Biology, 120(3):??, June 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1986:POb

[Ano86-32] Anonymous. Pages 375–498 (21 October 1986). Journal of Theoretical Biology, 122(4):??, October 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1986:PFb

[Ano86-33] Anonymous. Pages 377–499 (21 February 1986). Journal of Theoretical Biology, 118(4):??, February 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1986:PDb

[Ano86-34] Anonymous. Pages 377–502 (21 December 1986). Journal of Theoretical Biology, 123(4):??, December 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1986:PAb

[Ano86-35] Anonymous. Pages 379–502 (21 April 1986). Journal of Theoretical Biology, 119(4):??, April 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1986:PJd

[Ano86-36] Anonymous. Pages 381–505 (21 June 1986). Journal of Theoretical Biology, 120(4):??, June 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1986:PAd

[Ano86-37] Anonymous. Pages 381–508 (21 August 1986). Journal of Theoretical Biology, 121(4):??, August 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1987:A

[Ano87a] Anonymous. Announcement. Journal of Theoretical Biology, 128(4):525, October 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519387801960.

Anonymous:1987:EBa

[Ano87b] Anonymous. Editorial Board. Journal of Theoretical Biology, 124 (1):i, January 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387802461.

Anonymous:1987:EBb

[Ano87c] Anonymous. Editorial Board. Journal of Theoretical Biology, 125 (1):i, March 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251938780173X.

Anonymous:1987:EBc

[Ano87d] Anonymous. Editorial Board. Journal of Theoretical Biology, 126 (1):i, May 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387800942.

Anonymous:1987:EBd

[Ano87e] Anonymous. Editorial Board. Journal of Theoretical Biology, 127 (1):i, July 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251938780156X.

Anonymous:1987:EBe

[Ano87f] Anonymous. Editorial Board. Journal of Theoretical Biology, 128(1):i, September 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519387800267.

Anonymous:1987:EBf

[Ano87g] Anonymous. Editorial Board. Journal of Theoretical Biology, 129(1):i, November 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519387801984.

Anonymous:1987:Ea

[Ano87h] Anonymous. Errata. Journal of Theoretical Biology, 126(3): 322, June 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387802382.

Anonymous:1987:Ec

[Ano87i] Anonymous. Errata. Journal of Theoretical Biology, 128(4):527– 528, October 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387801972.

Anonymous:1987:IVa

[Ano87j] Anonymous. Index to volume 124. Journal of Theoretical Biology, 124(1):vi–x, January 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387802473.

Anonymous:1987:IVb

[Ano87k] Anonymous. Index to volume 125. Journal of Theoretical Biology, 125(1):vi–xii, March 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387801741.

Anonymous:1987:IVc

[Ano87l] Anonymous. Index to volume 126. Journal of Theoretical Biology, 126(1):vi–xii, May 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519387800954.

Anonymous:1987:IVd

[Ano87m] Anonymous. Index to volume 127. Journal of Theoretical Biology, 127(1):vi–xi, July 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519387801571.

Anonymous:1987:IVe

[Ano87n] Anonymous. Index to volume 128. Journal of Theoretical Biology, 128(1):vi–xi, September 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387800279.

Anonymous:1987:IVf

[Ano87o] Anonymous. Index to volume 129. Journal of Theoretical Biology, 129(4):vi–x, December 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387800255.

Anonymous:1987:PMa

[Ano87p] Anonymous. Pages 1–115 (7 March 1987). Journal of Theoretical Biology, 125(1):??, March 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1987:PJa

[Ano87q] Anonymous. Pages 1–125 (7 January 1987). Journal of Theoretical Biology, 124(1):??, January 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1987:PJe

[Ano87r] Anonymous. Pages 1–125 (7 July 1987). Journal of Theoretical Biology, 127(1):??, July 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1987:PSa

[Ano87s] Anonymous. Pages 1–125 (7 September 1987). Journal of Theoretical Biology, 128(1):??, September 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1987:PMc

[Ano87t] Anonymous. Pages 1–126 (7 May 1987). Journal of Theoretical Biology, 126(1):??, May 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1987:PNa

[Ano87u] Anonymous. Pages 1–140 (7 November 1987). Journal of Theoretical Biology, 129(1):??, November 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1987:PMb

[Ano87v] Anonymous. Pages 125–246 (21 March 1987). Journal of Theoretical Biology, 125(2):??, March 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1987:PJb

[Ano87w] Anonymous. Pages 127–250 (21 January 1987). Journal of Theoretical Biology, 124(2):??, January 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1987:PJf

[Ano87x] Anonymous. Pages 127–252 (21 July 1987). Journal of Theoretical Biology, 127(2):??, July 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1987:PMd

[Ano87y] Anonymous. Pages 127–252 (21 May 1987). Journal of Theoretical Biology, 126(2):??, May 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1987:PSb

[Ano87z] Anonymous. Pages 127–252 (21 September 1987). Journal of Theoretical Biology, 128(2):??, September 21, 1987. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1987:PNb

[Ano87-27] Anonymous. Pages 141–261 (21 November 1987). Journal of Theoretical Biology, 129(2):??, November 21, 1987. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1987:PAa

[Ano87-28] Anonymous. Pages 249–368 (7 April 1987). Journal of Theoretical Biology, 125(3):??, April 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1987:PFa

[Ano87-29] Anonymous. Pages 251–376 (7 February 1987). Journal of Theoretical Biology, 124(3):??, February 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1987:PAc

[Ano87-30] Anonymous. Pages 253–379 (7 August 1987). Journal of Theoretical Biology, 127(3):??, August 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1987:PJc

[Ano87-31] Anonymous. Pages 253–380 (7 June 1987). Journal of Theoretical Biology, 126(3):??, June 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1987:POa

[Ano87-32] Anonymous. Pages 253–398 (7 October 1987). Journal of Theoretical Biology, 128(3):??, October 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1987:PDa

[Ano87-33] Anonymous. Pages 263–357 (7 December 1987). Journal of Theoretical Biology, 129(3):??, December 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1987:PDb

[Ano87-34] Anonymous. Pages 359–455 (21 December 1987). Journal of Theoretical Biology, 129(4):??, December 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1987:PAb

[Ano87-35] Anonymous. Pages 369–496 (21 April 1987). Journal of Theoretical Biology, 125(4):??, April 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1987:PFb

[Ano87-36] Anonymous. Pages 377–504 (21 February 1987). Journal of Theoretical Biology, 124(4):??, February 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1987:PAd

[Ano87-37] Anonymous. Pages 381–506 (21 August 1987). Journal of Theoretical Biology, 127(4):??, August 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1987:PJd

[Ano87-38] Anonymous. Pages 381–506 (21 June 1987). Journal of Theoretical Biology, 126(4):??, June 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1987:POb

[Ano87-39] Anonymous. Pages 399–528 (21 October 1987). Journal of Theoretical Biology, 128(4):??, October 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1988:A

[Ano88a] Anonymous. Announcements. Journal of Theoretical Biology, 134(1):131–132, September 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519388803084.

Anonymous:1988:EBa

[Ano88b] Anonymous. Editorial Board. Journal of Theoretical Biology, 130 (1):i, January 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519388801590.

Anonymous:1988:EBb

[Ano88c] Anonymous. Editorial Board. Journal of Theoretical Biology, 131 (1):i, March 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519388801152.

Anonymous:1988:EBc

[Ano88d] Anonymous. Editorial Board. Journal of Theoretical Biology, 132 (1):i, May 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251938880184X.

Anonymous:1988:EBd

[Ano88e] Anonymous. Editorial Board. Journal of Theoretical Biology, 133 (1):i, July 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519388800183.

Anonymous:1988:EBe

[Ano88f] Anonymous. Editorial Board. Journal of Theoretical Biology, 134(1):i, September 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519388802972.

Anonymous:1988:EBf

[Ano88g] Anonymous. Editorial Board. Journal of Theoretical Biology, 135(1):i, November 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519388801693.

Anonymous:1988:Ec

[Ano88h] Anonymous. Errata. Journal of Theoretical Biology, 131(4): 515, April 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519388800468.

Anonymous:1988:Ee

[Ano88i] Anonymous. Errata. Journal of Theoretical Biology, 134(1):133, September 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519388803096.

Anonymous:1988:Ea

[Ano88j] Anonymous. Erratum. Journal of Theoretical Biology, 131(2): 173, March 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519388802340.

Anonymous:1988:Ed

[Ano88k] Anonymous. Erratum. Journal of Theoretical Biology, 132(3): 379, June 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519388802236.

Anonymous:1988:Ef

[Ano88l] Anonymous. Erratum. Journal of Theoretical Biology, 135 (1):137, November 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519388801838.

Anonymous:1988:Eg

[Ano88m] Anonymous. Erratum. Journal of Theoretical Biology, 135 (2):263, November 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002251938880078X.

Anonymous:1988:IVa

[Ano88n] Anonymous. Index to volume 131. Journal of Theoretical Biology, 131(4):vi–xii, April 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938880047X.

Anonymous:1988:IVb

[Ano88o] Anonymous. Index to volume 132. Journal of Theoretical Biology, 132(4):vi–xi, June 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519388800882.

Anonymous:1988:IVc

[Ano88p] Anonymous. Index to volume 133. Journal of Theoretical Biology, 133(1):vi–xi, July 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519388800195.

Anonymous:1988:IVd

[Ano88q] Anonymous. Index to volume 134. Journal of Theoretical Biology, 134(1):vi–xi, September 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388802984.

Anonymous:1988:IVe

[Ano88r] Anonymous. Index to volume 135. Journal of Theoretical Biology, 135(1):vi–xii, November 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938880170X.

Anonymous:1988:PJa

[Ano88s] Anonymous. Pages 1–126 (7 January 1988). Journal of Theoretical Biology, 130(1):??, January 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1988:PMc

[Ano88t] Anonymous. Pages 1–126 (7 May 1988). Journal of Theoretical Biology, 132(1):??, May 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1988:PJe

[Ano88u] Anonymous. Pages 1–131 (7 July 1988). Journal of Theoretical Biology, 133(1):??, July 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1988:PSa

[Ano88v] Anonymous. Pages 1–133 (7 September 1988). Journal of Theoretical Biology, 134(1):??, September 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1988:PMa

[Ano88w] Anonymous. Pages 1–136 (7 March 1988). Journal of Theoretical Biology, 131(1):??, March 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1988:PNa

[Ano88x] Anonymous. Pages 1–137 (7 November 1988). Journal of Theoretical Biology, 135(1):??, November 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1988:PMd

[Ano88y] Anonymous. Pages 127–250 (21 May 1988). Journal of Theoretical Biology, 132(2):??, May 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1988:PJb

[Ano88z] Anonymous. Pages 129–254 (21 January 1988). Journal of Theoretical Biology, 130(2):??, January 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1988:PJf

[Ano88-27] Anonymous. Pages 133–260 (21 July 1988). Journal of Theoretical Biology, 133(2):??, July 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1988:PSb

[Ano88-28] Anonymous. Pages 135–272 (21 September 1988). Journal of Theoretical Biology, 134(2):??, September 21, 1988. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1988:PMb

[Ano88-29] Anonymous. Pages 137–262 (21 March 1988). Journal of Theoretical Biology, 131(2):??, March 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1988:PNb

[Ano88-30] Anonymous. Pages 139–263 (21 November 1988). Journal of Theoretical Biology, 135(2):??, November 21, 1988. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1988:PJc

[Ano88-31] Anonymous. Pages 253–379 (7 June 1988). Journal of Theoretical Biology, 132(3):??, June 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1988:PFa

[Ano88-32] Anonymous. Pages 255–378 (7 February 1988). Journal of Theoretical Biology, 130(3):??, February 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1988:PAc

[Ano88-33] Anonymous. Pages 261–396 (7 August 1988). Journal of Theoretical Biology, 133(3):??, August 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1988:PAa

[Ano88-34] Anonymous. Pages 263–385 (7 April 1988). Journal of Theoretical Biology, 131(3):??, April 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1988:PDa

[Ano88-35] Anonymous. Pages 265–429 (7 December 1988). Journal of Theoretical Biology, 135(3):??, December 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1988:POa

[Ano88-36] Anonymous. Pages 273–415 (7 October 1988). Journal of Theoretical Biology, 134(3):??, October 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1988:PFb

[Ano88-37] Anonymous. Pages 379–515 (21 February 1988). Journal of Theoretical Biology, 130(4):??, February 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1988:PJd

[Ano88-38] Anonymous. Pages 381–510 (21 June 1988). Journal of Theoretical Biology, 132(4):??, June 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1988:PAb

[Ano88-39] Anonymous. Pages 389–515 (21 April 1988). Journal of Theoretical Biology, 131(4):??, April 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1988:PAd

[Ano88-40] Anonymous. Pages 397–528 (21 August 1988). Journal of Theoretical Biology, 133(4):??, August 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1988:POb

[Ano88-41] Anonymous. Pages 419–570 (21 October 1988). Journal of Theoretical Biology, 134(4):??, October 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1988:PDb

[Ano88-42] Anonymous. Pages 433–588 (21 December 1988). Journal of Theoretical Biology, 135(4):??, December 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1989:EBa

[Ano89a] Anonymous. Editorial Board. Journal of Theoretical Biology, 136 (1):i, January 9, 1989. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519389801821.

Anonymous:1989:EBb

[Ano89b] Anonymous. Editorial Board. Journal of Theoretical Biology, 137 (1):i, March 7, 1989. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519389801432.

Anonymous:1989:EBc

[Ano89c] Anonymous. Editorial Board. Journal of Theoretical Biology, 138 (1):i, May 7, 1989. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519389801729.

Anonymous:1989:EBd

[Ano89d] Anonymous. Editorial Board. Journal of Theoretical Biology, 139 (1):i, July 10, 1989. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519389800529.

Anonymous:1989:EBe

[Ano89e] Anonymous. Editorial Board. Journal of Theoretical Biology, 140(1):i, September 11, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519389800232.

Anonymous:1989:EBf

[Ano89f] Anonymous. Editorial Board. Journal of Theoretical Biology, 141(1):i, November 8, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519389800013.

Anonymous:1989:IVa

[Ano89g] Anonymous. Index to volume 136. Journal of Theoretical Biology, 136(1):vi–x, January 9, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389801833.

Anonymous:1989:IVb

[Ano89h] Anonymous. Index to volume 137. Journal of Theoretical Biology, 137(1):v–viii, March 7, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389801444.

Anonymous:1989:IVc

[Ano89i] Anonymous. Index to volume 137. Journal of Theoretical Biology, 138(1):vi–x, May 7, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519389801730.

Anonymous:1989:IVd

[Ano89j] Anonymous. Index to volume 139. Journal of Theoretical Biology, 139(1):vi–x, July 10, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519389800530.

Anonymous:1989:IVe

[Ano89k] Anonymous. Index to volume 140. Journal of Theoretical Biology, 140(1):vi–xii, September 11, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389800244.

Anonymous:1989:IVf

[Ano89l] Anonymous. Index to volume 141. Journal of Theoretical Biology, 141(1):vi–x, November 8, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389800025.

Anonymous:1989:PJa

[Ano89m] Anonymous. Pages 1–123 (9 January 1989). Journal of Theoretical Biology, 136(1):??, January 9, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1989:PMa

[Ano89n] Anonymous. Pages 1–125 (7 March 1989). Journal of Theoretical Biology, 137(1):??, March 7, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1989:PMc

[Ano89o] Anonymous. Pages 1–128 (7 May 1989). Journal of Theoretical Biology, 138(1):??, May 7, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1989:PJe

[Ano89p] Anonymous. Pages 1–139 (10 July 1989). Journal of Theoretical Biology, 139(1):??, July 10, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1989:PNa

[Ano89q] Anonymous. Pages 1–141 (8 November 1989). Journal of Theoretical Biology, 141(1):??, November 8, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1989:PSa

[Ano89r] Anonymous. Pages 1–143 (11 September 1989). Journal of Theoretical Biology, 140(1):??, September 11, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1989:PMb

[Ano89s] Anonymous. Pages 127–243 (21 March 1989). Journal of Theoretical Biology, 137(2):??, March 21, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1989:PJb

[Ano89t] Anonymous. Pages 127–243 (23 January 1989). Journal of Theoretical Biology, 136(2):??, January 23, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1989:PMd

[Ano89u] Anonymous. Pages 129–270 (22 May 1989). Journal of Theoretical Biology, 138(2):??, May 22, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1989:PNb

[Ano89v] Anonymous. Pages 143–283 (21 November 1989). Journal of Theoretical Biology, 141(2):??, November 21, 1989. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1989:PJf

[Ano89w] Anonymous. Pages 143–286 (21 July 1989). Journal of Theoretical Biology, 139(2):??, July 21, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1989:PSb

[Ano89x] Anonymous. Pages 145–287 (22 September 1989). Journal of Theoretical Biology, 140(2):??, September 22, 1989. CODEN JT-BIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1989:PAa

[Ano89y] Anonymous. Pages 245–363 (6 April 1989). Journal of Theoretical Biology, 137(3):??, April 6, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1989:PFa

[Ano89z] Anonymous. Pages 245–364 (8 February 1989). Journal of Theoretical Biology, 136(3):??, February 8, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1989:PJc

[Ano89-27] Anonymous. Pages 271–412 (3 June 1989). Journal of Theoretical Biology, 138(3):??, June 3, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1989:PDa

[Ano89-28] Anonymous. Pages 285–427 (7 December 1989). Journal of Theoretical Biology, 141(3):??, December 7, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1989:PAc

[Ano89-29] Anonymous. Pages 287–430 (9 August 1989). Journal of Theoretical Biology, 139(3):??, August 9, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1989:POa

[Ano89-30] Anonymous. Pages 289–430 (9 October 1989). Journal of Theoretical Biology, 140(3):??, October 9, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1989:PAb

[Ano89-31] Anonymous. Pages 365–488 (20 April 1989). Journal of Theoretical Biology, 137(4):??, April 20, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1989:PFb

[Ano89-32] Anonymous. Pages 365–492 (22 February 1989). Journal of Theoretical Biology, 136(4):??, February 22, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1989:PJd

[Ano89-33] Anonymous. Pages 413–554 (22 June 1989). Journal of Theoretical Biology, 138(4):??, June 22, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1989:PDb

[Ano89-34] Anonymous. Pages 429–571 (19 December 1989). Journal of Theoretical Biology, 141(4):??, December 19, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1989:PAd

[Ano89-35] Anonymous. Pages 431–574 (22 August 1989). Journal of Theoretical Biology, 139(4):??, August 22, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Anonymous:1989:POb

[Ano89-36] Anonymous. Pages 431–574 (23 October 1989). Journal of Theoretical Biology, 140(4):??, October 23, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

Aida:1982:IMO

[ANOM82] Misako Aida, Chikayoshi Nagata, Iwao Ohmine, and Keiji Morokuma. Ab initio molecular orbital study on the thermostability of the extreme thermophile tRNA: Role of the base stacking. Journal of Theoretical Biology, 99(3):599–608, December 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382902132.

Angyan:1983:CPE

[ANS83] János Angyan and Gábor Náray-Szabó. Comparison of protein electrostatic potential along the catalytic triad of serine proteinases. Journal of Theoretical Biology, 103(3):349–356, August 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383902916.

Azuma:1987:FSA

[AO87] Akira Azuma and Yoshinori Okuno. Flight of a samara, alsomitra macrocarpa. Journal of Theoretical Biology, 129(3):263–274, December 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387800012.

Antillon:1985:GTA

[AOB85] A. Antillón and I. Ortega-Blake. A group theory analysis of the ambiguities in the genetic code: on the existence of a generalized genetic code. Journal of Theoretical Biology, 112(4):757–769, February 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251938580059X.

Aoki:1984:QGM

[Aok84] Kenichi Aoki. A quantitative genetic model of two-policy games between relatives. Journal of Theoretical Biology, 109(1):111– 126, July 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384801149.

Aoki:1989:EFE

[Aok89] Ichiro Aoki. Entropy flow and entropy production in the human body in basal conditions. Journal of Theoretical Biology, 141(1):11–21, November 8, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389800049.

Armont:1980:WRP

[AP80] Carmel Armont and Yoram Palti. The weighted reversal potential — a tool to study refractoriness and regenerativity in nerves. Journal of Theoretical Biology, 83(3):505–515, April 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380900569.

Argos:1981:PSM

[Arg81] Patrick Argos. Protein structural models for nucleic acid interactions. Journal of Theoretical Biology, 93(3):609–626, December 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381902241.

Armstrong:1984:WDC

[Arm84] Doug P. Armstrong. Why don’t cellular slime molds cheat? Journal of Theoretical Biology, 109(2):271–283, July 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384800065.

Armaleo:1987:SEP

[Arm87a] Daniele Armaleo. Structure and evolution of prokaryotic and eukaryotic RNA polymerases: a model. Journal of Theoretical Biology, 127(3):301–314, August 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938780108X.

Armstrong:1987:PMM

[Arm87b] Robert A. Armstrong. A patch model of mutualism. Journal of Theoretical Biology, 125(2):243–246, March 21, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387800450.

Armaleo:1988:NRP

[Arm88a] Daniele Armaleo. A note on RNA polymerase phylogeny. Journal of Theoretical Biology, 131(1):135–136, March 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388801267.

Armstrong:1988:EDP

[Arm88b] Robert A. Armstrong. The effects of disturbance patch size on species coexistence. Journal of Theoretical Biology, 133(2):169– 184, July 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519388800031.

Armstrong:1989:CSP

[Arm89] Robert A. Armstrong. Competition, seed predation, and species coexistence. Journal of Theoretical Biology, 141(2):191–195, November 21, 1989. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519389800177.

Aron:1984:SPD

[AS84] Joan L. Aron and Ira B. Schwartz. Seasonality and perioddoubling bifurcations in an epidemic model. Journal of Theoretical Biology, 110(4):665–679, October 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801502.

Anderson:1986:DBU

[AS86] James Jay Anderson and Patrick J. Sullivan. A dynamic basis for using a truncated normal distribution to describe variability of chemical substances in aquatic environments. Journal of Theoretical Biology, 123(2):213–220, November 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801552.

Alff-Steinberger:1987:CUH

[AS87] C. Alff-Steinberger. Codon usage in homo sapiens: Evidence for a coding pattern on the non-coding strand and evolutionary implications of dinucleotide discrimination. Journal of Theoretical Biology, 124(1):89–95, January 7, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387802540.

Ashida:1985:GSC

[Ash85] H. Ashida. General solution of cable theory with both ends sealed. Journal of Theoretical Biology, 112(4):727–740, February 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385800564.

Ashida:1986:SDT

[Ash86] Hiroshi Ashida. Stochastic description of the three-state model of the na channel. Journal of Theoretical Biology, 121(1):45– 57, July 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519386800273.

Acerenza:1989:CAT

[ASK89] Luis Acerenza, Herbert M. Sauro, and Henrik Kacser. Control analysis of time-dependent metabolic systems. Journal of Theoretical Biology, 137(4):423–444, April 20, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389800384.

Asmussen:1983:DDS

[Asm83] Marjorie A. Asmussen. Density dependent selection incorporating intraspecific competition 1. A haploid model. Journal of Theoretical Biology, 101(1):113–127, March 7, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383902758.

Aspden:1988:TFR

[Asp88] R. M. Aspden. The theory of fibre-reinforced composite materials applied to changes in the mechanical properties of the cervix during pregnancy. Journal of Theoretical Biology, 130(2):213–221, January 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519388800961.

Angulo:1980:PDM

[ASTT80] Juan J. Angulo, Truett L. Smith, Janice O. Tsokos, and Chris P. Tsokos. Population-dynamics models and a sequential test in the analysis of the influence of household setting on the spread of variola minor. Journal of Theoretical Biology, 82(1):91–103, January 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519380900910.

Alvarez:1985:SAT

[AT85] Jaime Alvarez and Juan Carlos Torres. Slow axoplasmic transport: a fiction? Journal of Theoretical Biology, 112(3):627–651, February 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251938580028X.

Aarnaes:1981:MAC

[ATC81] E. Aarnæs, E. Thorud, and O. P. Clausen. Model analysis of circadian rhythms in mouse epidermal basal cell proliferation. Journal of Theoretical Biology, 88(2):355–370, January 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381900813.

Auger:1985:SIP

[Aug85] P. Auger. Stability of interacting populations with age-class distributions. Journal of Theoretical Biology, 112(3):595–605, February 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385800254.

Aruja:1982:DPP

[AVR82] A. Aruja, R. Vilu, and E. Raukas. Detection of periodic patterns in RNA sequences: the first encapsidated region of the TMV RNA. Journal of Theoretical Biology, 94(2):457–470, January 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382903216.

Allen:1983:HMC

[AW83] T. F. H. Allen and E. Paul Wyleto. A hierarchical model for the complexity of plant communities. Journal of Theoretical Biology, 101(4):529–540, April 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383900140.

Adler:1985:MMA

[AWM85] David Adler, Alan Y. K. Wong, and Yona Mahler. Model of mechanical alternans in the mammalian myocardium. Journal of Theoretical Biology, 117(4):563–577, December 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385802381.

Adler:1985:MCM

[AWMK85] David Adler, Alan Y. K. Wong, Y. Mahler, and G. A. Klassen. Model of calcium movements in the mammalian myocardium: Interval-strength relationship. Journal of Theoretical Biology, 113(2):379–394, March 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385802332.

Arminskit:1980:TDT

[AWP80] Keslie Arminskit, Sheldon Weinbaum, and Robert Pfeffer. Time dependent theory for vesicular transport across vascular endothelium. Journal of Theoretical Biology, 85(1):13–43, July 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380902787.

Allaerts:1988:DAM

[AWV^{+}88] Wilfried Allaerts, Alain Wouters, Dirk Van der Massen, André Persoons, and Carl Denef. A diffusion-adsorption model for the computation of the amount of hormone around a secreting cell, detected by the reverse hemolytic plaque assay. Journal of Theoretical Biology, 131(4):441–459, April 21, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388800407.

Agu:1985:SDB

[AY85] Masahiro Agu and Yota Yokoi. A stochastic description of branching structures of trees. Journal of Theoretical Biology, 112(4):667–676, February 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519385800527.

Azuma:1989:FPR

[AY89] Akira Azuma and Kunio Yasuda. Flight performance of rotary seeds. Journal of Theoretical Biology, 138(1):23–53, May 7, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519389801766.

Anninos:1983:ANN

[AZE83] P. Anninos, S. Zenone, and R. Elul. Artificial neural nets: Dependence of the EEG amplitude’s probability distribution on statistical parameters. Journal of Theoretical Biology, 103(3):339– 348, August 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383902904.

Brown:1980:VOG

[BA80] A. H. D. Brown and L. Albrecht. Variable outcrossing and the genetic structure of predominantly self-pollinated species. Journal of Theoretical Biology, 82(4):591–606, February 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380901800.

Babitch:1987:GCB

[BA87] Joseph A. Babitch and Frank A. Anthony. Grasping for calcium binding sites in sodium channels with an EF hand. Journal of Theoretical Biology, 127(4):451–459, August 21, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938780142X.

Bass:1984:AFT

[BAB84] Ludvik Bass, Janet Aisbett, and Anthony J. Bracken. Asymptotic forms of tracer clearance curves: Theory and applications of improved extrapolations. Journal of Theoretical Biology, 111(4):755–785, December 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384802660.

Babitch:1988:CBC

[Bab88] Joseph A. Babitch. On calcium binding to channel proteins. Journal of Theoretical Biology, 133(4):525–528, August 21, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388803394.

Bahramian:1980:HBR

[Bah80] M. B. Bahramian. How bacterial ribosomes select translation initiation sites. Journal of Theoretical Biology, 84(1):103–118, May 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519380810393.

Bandea:1983:NTO

[Bân83] Claudiu I. Bândea. A new theory on the origin and the nature of viruses. Journal of Theoretical Biology, 105(4):591–602, December 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383902217.

Bartha:1980:EEQ

[Bar80a] F. Bartha. Effect of error of the quasi-steady-state approximation on the estimation of K_{m} and V_{m} from a single time curve. Journal of Theoretical Biology, 86(1):105–115, September 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380900685.

Bartha:1980:CIP

[Bar80b] F. Bartha. On a criterion of instability for positive feedback control systems. Journal of Theoretical Biology, 82(1):141–147, January 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519380900934.

Barbieri:1981:RTO

[Bar81a] Marcello Barbieri. The ribotype theory on the origin of life. Journal of Theoretical Biology, 91(4):545–601, August 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902113.

Bard:1981:MGA

[Bar81b] Jonathan B. L. Bard. A model for generating aspects of zebra and other mammalian coat patterns. Journal of Theoretical Biology, 93(2):363–385, November 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381901090.

Barrett:1981:TED

[Bar81c] A. N. Barrett. A theoretical explanation of directional hearing characteristics in amphibia. Journal of Theoretical Biology, 93(3):591–596, December 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902228.

Barrett:1981:ETM

[Bar81d] T. W. Barrett. Energy transfer and molecular switching. I. The nerve action potential. Journal of Theoretical Biology, 92(3):185– 207, October 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381902873.

Bartosz:1981:NSR

[Bar81e] G. Bartosz. Non-specific reactions: Molecular basis for ageing. Journal of Theoretical Biology, 91(1):233–235, July 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381903830.

Bard:1982:MCM

[Bar82a] Jonathan B. L. Bard. A membrane-carrier mechanism for generating new sources and sinks within gradient-controlled tissues — its application to regeneration in oncopeltus. Journal of Theoretical Biology, 94(3):689–708, February 7, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382903071.

Barrett:1982:ETM

[Bar82b] T. W. Barrett. Energy transfer and molecular switching II. Muscle contraction and enzymatic reactions. Journal of Theoretical Biology, 99(2):293–307, November 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382900066.

Bardsley:1983:SEK

[Bar83a] William G. Bardsley. Simple enzyme kinetic mechanisms that can give all possible velocity profiles with chemically reasonable rate constant values. Journal of Theoretical Biology, 104(4):485–491, October 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383902412.

Barrai:1983:MD

[Bar83b] Italo Barrai. A model for differentiation. Journal of Theoretical Biology, 105(3):379–389, December 7, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383901820.

Barnard:1984:SCM

[Bar84] C. J. Barnard. Stasis: a coevolutionary model. Journal of Theoretical Biology, 110(1):27–34, September 7, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384800132.

Barker:1986:SPP

[Bar86] Roger Barker. Substance P and Parkinson’s disease: a causal relationship? Journal of Theoretical Biology, 120(3):353–362, June 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386802065.

Barach:1987:EOR

[Bar87] John Paul Barach. The effect of ohmic return currents on biomagnetic fields. Journal of Theoretical Biology, 125(2):187–191, March 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387800401.

Barker:1988:HDB

[Bar88] Roger Barker. How does the brain control its own activity? A new function for the basal ganglia. Journal of Theoretical Biology, 131(4):497–507, April 21, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388800444.

Baranowitz:1989:RNC

[Bar89] Steven A. Baranowitz. Regeneration, neural crest derivatives and retinoids: A new synthesis. Journal of Theoretical Biology, 140(2):231–242, September 22, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389801316.

Bass:1980:FDF

[Bas80a] L. Bass. Flow dependence of first-order uptake of substances by heterogeneous perfused organs. Journal of Theoretical Biology, 86(2):365–376, September 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380900120.

Bass:1980:LCF

[Bas80b] L. Bass. On the location of cellular functions in perfused organs. Journal of Theoretical Biology, 82(3):347–351, February 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380902416.

Basu:1980:TEC

[Bas80c] Samarendra Basu. Theoretical estimates of cathodoluminescence of giemsa stained chromosomes. Journal of Theoretical Biology, 85(1):125–151, July 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380902854.

Bass:1981:FZR

[Bas81] L. Bass. Functional zones in rat liver: The degree of overlap. Journal of Theoretical Biology, 89(2):303–319, March 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381903131.

Bass:1982:VTT

[Bas82] L. Bass. The variance of transit times of indicators through a vascular bed. Journal of Theoretical Biology, 95(1):81–87, March 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382902880.

Bass:1983:EIC

[Bas83] L. Bass. Estimates and implications of co-operativity for enzyme kinetics in the intact liver: Method of flow invariants. Journal of Theoretical Biology, 100(1):113–121, January 7, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383900966.

Barnard:1981:IAU

[BB81a] C. J. Barnard and Theodore Burke. Individuals as assessment units — reply to Breed and Bekoff. Journal of Theoretical Biology, 88(3):595–597, February 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902861.

Barrett:1981:TDM

[BB81b] A. N. Barrett and I. D. J. Burdett. A three-dimensional model reconstruction of pole assembly in bacillus subtilis. Journal of Theoretical Biology, 92(2):127–139, September 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251938190388X.

Borgia:1981:SCE

[BB81c] Gerald Borgia and James Blick. Sexual competition and the evolution of hermaphroditism. Journal of Theoretical Biology, 89(3):523–532, April 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519381903660.

Breed:1981:IRS

[BB81d] Michael D. Breed and Marc Bekoff. Individual recognition and social relationships. Journal of Theoretical Biology, 88(3):589– 593, February 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251938190285X.

Bowman:1984:AGH

[BB84] Charles L. Bowman and Adriano Baglioni. Application of the Goldman–Hodgkin–Katz current equation to membrane current– voltage data. Journal of Theoretical Biology, 108(1):1–29, May 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384801654.

Balanovski:1985:ODB

[BB85] Eduardo Balanovski and Peter Beaconsfield. Order and disorder in biophysical systems: a study of the correlation between structure and function of DNA. Journal of Theoretical Biology, 114(1): 21–33, May 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385802538.

Baake:1986:CGR

[BB86] M. Baake and E. Buff. From cell growth to root growth. Journal of Theoretical Biology, 123(3):347–365, December 7, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386802508.

Bertram:1988:BCS

[BB88a] John E. A. Bertram and Andrew A. Biewener. Bone curvature: Sacrificing strength for load predictability? Journal of Theoretical Biology, 131(1):75–92, March 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938880122X.

Bovet:1988:SAA

[BB88b] Pierre Bovet and Simon Benhamou. Spatial analysis of animals’ movements using a correlated random walk model. Journal of Theoretical Biology, 131(4):419–433, April 21, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388800389.

Burnett:1984:CRB

[BBH84] Leslie Burnett, Antony Basten, and William J. Hensley. Complementary relationship between repetitive sequences in the SV40 early gene pre-promoter enhancing region and big-t antigen splice junction exons. Journal of Theoretical Biology, 110(4):533–540, October 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384801435.

Bass:1986:FRT

[BBH86] L. Bass, A. J. Bracken, and J. Hilden. Flux ratio theorems for nonstationary membrane transport with temporary capture of tracer. Journal of Theoretical Biology, 118(3):327–338, February 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386800637.

Bernstein:1984:OS

[BBHM84] Harris Bernstein, Henry C. Byerly, Frederic A. Hopf, and Richard E. Michod. Origin of sex. Journal of Theoretical Biology, 110(3):323–351, October 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801782.

Bernstein:1985:SES

[BBHM85] Harris Bernstein, Henry C. Byerly, Frederic A. Hopf, and Richard E. Michod. Sex and the emergence of species. Journal of Theoretical Biology, 117(4):665–690, December 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385802460.

Bates:1987:DFD

[BBM87] Jason H. T. Bates, Diana A. Bates, and William Mackillop. On the difficulties of fitting the double Michaelis–Menten equation to kinetic data. Journal of Theoretical Biology, 125(2):237–241, March 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387800449.

Bonifazzi:1988:KAI

[BBS88] C. Bonifazzi, O. Belluzzi, and O. Sacchi. Kinetic analysis of incomplete current tracings according to the Hodgkin–Huxley model. Journal of Theoretical Biology, 130(2):183–190, January 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388800936.

Belintsev:1987:MPF

[BBZ87] B. N. Belintsev, L. V. Beloussov, and A. G. Zaraisky. Model of pattern formation in epithelial morphogenesis. Journal of Theoretical Biology, 129(4):369–394, December 21, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938780019X.

Boeynaems:1980:CBK

[BC80] J. M. Boeynaems and F. R. L. Cantraine. Comparison between the kinetics of heterogeneous and sequentially co-operative binding systems. Journal of Theoretical Biology, 83(3):447–456, April 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251938090051X.

Bremer:1981:CCE

[BC81a] H. Bremer and Lily Chuang. The cell cycle in Escherichia coli B/r. Journal of Theoretical Biology, 88(1):47–81, January 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381903283.

Bremer:1981:CDA

[BC81b] H. Bremer and Lily Chuang. Cell division after inhibition of chromosome replication in Escherichia coli. Journal of Theoretical Biology, 93(4):909–926, December 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381903477.

Bremer:1983:CDE

[BC83] H. Bremer and L. Chuang. Cell division in Escherichia coli after changes in the velocity of DNA replication. Journal of Theoretical Biology, 102(1):101–120, May 7, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383902643.

Bishop:1986:OCR

[BC86a] D. T. Bishop and C. Cannings. Ordinal conflicts with random rewards. Journal of Theoretical Biology, 122(2):225–230, September 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386800832.

Bornbusch:1986:ESS

[BC86b] Alan H. Bornbusch and William E. Conner. Effects of self-steered turn size and turn bias upon simulated chemoklinotactic behavior. Journal of Theoretical Biology, 122(1):7–18, September 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386802211.

Barnett:1983:EMT

[BCL83] Gene Barnett, Chia-Whei N. Chiang, and Vojtech Licko. Effects of marijuana on testosterone in male subjects. Journal of Theoretical Biology, 104(4):685–692, October 21, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383902552.

Burres:1980:MTT

[BD80] Neal Burres and A. Keith Dunker. Membrane transport through α-heltical bundles. IV. preliminary model building investigation of helix-helix interactions. Journal of Theoretical Biology, 87(4): 723–736, ???? 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519380901137.

Baumann:1981:MPC

[BE81a] Gilbert Baumann and George S. Easton. Markov process characterization of a single membrane gating site. Journal of Theoretical Biology, 93(4):785–804, December 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381903404.

Blomberg:1981:ECK

[BE81b] Clas Blomberg and Måns Ehrenberg. Energy considerations for kinetic proofreading in biosynthesis. Journal of Theoretical Biology, 88(4):631–670, February 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902423.

Baumann:1982:CIL

[BE82] Gilbert Baumann and George S. Easton. Charge immobilization linked to inactivation in the aggregation model of channel gating. Journal of Theoretical Biology, 99(2):249–261, November 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382900030.

Bezeau:1986:DEP

[BE86] Mary Bezeau and Laszlo Endrenyi. Design of experiments for the precise estimation of dose-response parameters: the Hill equation. Journal of Theoretical Biology, 123(4):415–430, December 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386802119.

Becker:1980:HBV

[Bec80] Yechiel Becker. Is hepatitis B virus DNA in the dane particles a partial virus genome resembling escaping genes? Journal of Theoretical Biology, 86(2):255–257, September 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380900053.

Beck:1981:MMC

[Bec81] Karen Beck. A mathematical model of T-cell effects in the humoral immune response. Journal of Theoretical Biology, 89(4):593–610, April 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002251938190031X.

Beck:1985:DST

[Bec85] Stanley D. Beck. Dual system theory of the biological clock. Journal of Theoretical Biology, 113(1):93–115, March 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800795.

Beck:1988:IHR

[Bec88] James S. Beck. On internalization of hormone-receptor complex and receptor recycling. Journal of Theoretical Biology, 132(3): 263–276, June 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519388802157.

Beebee:1980:QRT

[Bee80] Trevor J. C. Beebee. Quantitative regulation of transcription in eurkaryotes: Theoretical considerations of RNA polymerase involvement. Journal of Theoretical Biology, 86(4):803–815, October 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519380903124.

Bellett:1980:MRL

[Bel80] A. J. D. Bellett. Models for the replication of linear DNA must account for the completion of both molecular ends. Journal of Theoretical Biology, 85(3):549–551, August 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380903276.

Belintsev:1984:SFS

[Bel84] B. N. Belintsev. The spatial form self-organization in the development of Dictyostelium discoideum. Journal of Theoretical Biology, 108(1):123–129, May 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801733.

Bell:1986:EEF

[Bel86] Graham Bell. The evolution of empty flowers. Journal of Theoretical Biology, 118(3):253–258, February 7, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386800571.

Belovsky:1987:FOB

[Bel87] Gary E. Belovsky. Foraging and optimal body size: An overview, new data and a test of alternative models. Journal of Theoretical Biology, 129(3):275–287, December 7, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387800024.

Benhamou:1989:OOM

[Ben89] Simon Benhamou. An olfactory orientation model for mammals’ movements in their home ranges. Journal of Theoretical Biology, 139(3):379–388, August 9, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389802164.

Bernard:1982:TAC

[Ber82] Christian Bernard. Theoretical aspects of cupula deflection in semicircular canal. Journal of Theoretical Biology, 98(4):637–643, October 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382901424.

Bergman:1983:MMC

[Ber83] Mark O. Bergman. Mathematical model for contact inhibited cell division. Journal of Theoretical Biology, 102(3):375–386, June 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383903752.

Berenbaum:1985:EEC

[Ber85] Morris C. Berenbaum. The expected effect of a combination of agents: the general solution. Journal of Theoretical Biology, 114(3):413–431, June 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801764.

Bernstein:1987:DVT

[Ber87] Marianne E. Bernstein. Does variation in the testosterone level of the seminal plasma affect the primary sex ratio? Journal of Theoretical Biology, 126(3):377–378, June 7, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387802448.

Bertling:1989:FLL

[Ber89] Wolf M. Bertling. Full length L1 retroposons contain tRNA-like sequences near the 5’ termini-hypothesis on the replication mechanism of retroposons. Journal of Theoretical Biology, 138(2):185– 194, May 22, 1989. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519389801389.

Bardakjian:1984:ICN

[BESD84] B. L. Bardakjian, T. Y. El-Sharkawy, and N. E. Diamant. Interaction of coupled nonlinear oscillators having different intrinsic resting levels. Journal of Theoretical Biology, 106(1):9–23, January 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519384900249.

Bustin:1981:ICS

[BF81] Michael Bustin and Richard J. Feldmann. Interactive computer surface graphic study of the binding of an antibody to the chromatin subunit. Journal of Theoretical Biology, 92(2):97–102, September 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381903854.

Bywater:1983:PGC

[BF83] R. P. Bywater and J. L. Finney. Particle geometry and the chromatographic separation of cells: The simplest case of uniform particles. Journal of Theoretical Biology, 105(2):333–343, ???? 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383800125.

Bremermann:1985:SPH

[BF85] H. J. Bremermann and B. Fiedler. On the stability of polymorphic host-pathogen populations. Journal of Theoretical Biology, 117(4):621–631, December 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385802435.

Boyett:1988:CSE

[BF88] M. R. Boyett and D. Fedida. A computer simulation of the effect of heart rate on ion concentrations in the heart. Journal of Theoretical Biology, 132(1):15–27, May 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388801875.

Beretta:1984:TIF

[BG84] Edoardo Beretta and Franco Gambale. A thermodynamic interpretation to formation of clusters at the cell surface. Journal of Theoretical Biology, 108(1):85–109, May 7, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938480171X.

Benigni:1986:NIC

[BG86] R. Benigni and A. Giuliani. A new insight into chemical mutagenesis by multivariate data analysis. Journal of Theoretical Biology, 121(4):477–486, August 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801035.

Benigni:1987:CFM

[BG87] R. Benigni and A. Giuliani. The conceptual framework of monitoring mutagens/carcinogens: a critical appraisal. Journal of Theoretical Biology, 125(4):465–473, April 21, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938780214X.

Briere:1988:GDT

[BG88] Christian Brière and Brian Goodwin. Geometry and dynamics of tip morphogenesis in acetabularia. Journal of Theoretical Biology, 131(4):461–475, April 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388800419.

Bass:1989:GMD

[BGB89] L. Bass, H. S. Green, and H. Boxenbaum. Gompertzian mortality derived from competition between cell-types: Congenital, toxicologic and biometric determinants of longevity. Journal of Theoretical Biology, 140(2):263–278, September 22, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938980133X.

Bruni:1984:APC

[BGG84] Carlo Bruni, Alberto Gandolfi, and Alfredo Germani. Analysis of the parameter constraints for a proposed antibody affinity distribution. Journal of Theoretical Biology, 109(1):71–76, July 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384801113.

Bertuzzi:1983:GES

[BGGV83] A. Bertuzzi, A. Gandolfi, A. Germani, and R. Vitelli. A general expression for sequential DNA-fluorescence histograms. Journal of Theoretical Biology, 102(1):55–67, May 7, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251938390262X.

Badgerow:1981:EST

[BH81a] John P. Badgerow and F. Reed Hainsworth. Energy savings through formation flight? A re-examination of the vee formation. Journal of Theoretical Biology, 93(1):41–52, November 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381900552.

Bell:1981:PMR

[BH81b] R. Charles Bell and John E. Hunter. Periodic majority recoding and reverberatory memory. Journal of Theoretical Biology, 88(2):323–332, January 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381900795.

Bosserman:1981:DCG

[BH81c] Robert W. Bosserman and Frank Harary. Demiarcs, creaons and genons. Journal of Theoretical Biology, 92(3):241–254, October 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381902903.

Blum:1983:DCQ

[BH83] J. J. Blum and M. H. Hines. Diffusion of Ca^{2+} and the quiescent response of sea urchin sperm flagella. Journal of Theoretical Biology, 100(3):511–523, February 7, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383904435.

Buxbaum:1988:TMF

[BH88] R. E. Buxbaum and S. R. Heidemann. A thermodynamic model for force integration and microtubule assembly during axonal elongation. Journal of Theoretical Biology, 134(3):379–390, October 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388800687.

Bhatt:1986:CMO

[Bha86] B. S. Bhatt. Comments on a model of ovum transport. Journal of Theoretical Biology, 121(2):249–254, July 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386800972.

Berding:1983:PPF

[BHH83] C. Berding, T. Harbich, and H. Haken. A pre-pattern formation mechanism for the spiral-type patterns of the sunflower head. Journal of Theoretical Biology, 104(1):53–70, September 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383904010.

Bransome:1985:ASC

[BHM^{+}85] E. D. Bransome, L. B. Hendry, T. G. Muldoon, V. B. Mahesh, M. S. Hutson, and L. K. Campbell. Apparent stereochemical complementarity of estrogens and helical cavities between DNA base pairs: Implications for the mechanism of action of steroids. Journal of Theoretical Biology, 112(1):97–108, January 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801181.

Bradbury:1985:CRC

[BHMR85] R. H. Bradbury, L. S. Hammond, P. J. Moran, and R. E. Reichelt. Coral reef communities and the crown-of-thorns starfish: Evidence for qualitatively stable cycles. Journal of Theoretical Biology, 113(1):69–80, March 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938580076X.

Bissell:1982:HDE

[BHP82] Mina J. Bissell, H. Glenn Hall, and Gordon Parry. How does the extracellular matrix direct gene expression? Journal of Theoretical Biology, 99(1):31–68, November 7, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382903885.

Bianchi:1981:MMP

[Bia81] N. O. Bianchi. Molecular mechanisms of production of symmetrical and asymmetrical chromosome exchanges. Journal of Theoretical Biology, 93(3):681–689, December 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902290.

Birdi:1981:CAS

[Bir81] K. S. Birdi. Cell adhesion on solids and the role of surface forces. Journal of Theoretical Biology, 93(1):1–5, November 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381900527.

Bloch:1980:MSD

[BIS^{+}80] Ralph Bloch, David Ingram, George D. Sweeney, Khursheed Ahmed, and C. John Dickinson. MacDope: a simulation of drug disposition in the human body. Mathematical considerations. Journal of Theoretical Biology, 87(2):211–236, November 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380903574.

Beckwith:1982:AMA

[BJ82] A. C. Beckwith and P. A. Jursinic. An alternative mathematical approach to the analysis of photosynthetic oxygen evolutiont. Journal of Theoretical Biology, 97(2):251–265, July 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382901023.

Bak:1984:RPI

[BJ84] A. Leth Bak and A. Lund Jørgensen. RNA polymerase III control regions in retrovirus LTR, alu-type repetitive DNA, and papovavirus. Journal of Theoretical Biology, 108(3):339–348, June 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384800387.

Bhimarao:1989:AAB

[BJ89] Chandrika Bhimarao and Rajani R. Joshi. On antigen-antibody binding distribution. Journal of Theoretical Biology, 141 (3):285–302, December 7, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002251938980116X.

Bodnar:1989:DME

[BJE89] John W. Bodnar, Gwilym S. Jones, and Charles H. Ellis. The domain model for eukaryotic DNA organization 2: A molecular basis for constraints on development and evolution. Journal of Theoretical Biology, 137(3):281–320, April 6, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389800748.

Blomberg:1985:EPC

[BJL85] C. Blomberg, J. Johansson, and H. Liljenström. Error propagation in E. coli protein synthesis. Journal of Theoretical Biology, 113(3):407–423, April 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800308.

Bugrii:1981:EET

[BK81] G. V. Bugrii and V. V. Kukhtin. The effect of electron transport on the kinetics of the manganese-containing superoxide dismutase from bacillus stearothermophilus. Journal of Theoretical Biology, 90(2):161–167, May 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381900394.

Burdett:1983:HDB

[BK83] I. D. J. Burdett and T. B. L. Kirkwood. How does a bacterium grow during its cell cycle? Journal of Theoretical Biology, 103(1): 11–20, July 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383901959.

Bulliman:1988:SES

[BK88] Brian T. Bulliman and Philip W. Kuchel. A series expression for the surface area of an ellipsoid and its application to the computation of the surface area of avian erythrocytes. Journal of Theoretical Biology, 134(1):113–123, September 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388803060.

Brailsford:1980:CCRa

[BKB80a] J. Douglas Brailsford, Ralph A. Korpman, and Brian S. Bull. Crenation and cupping of the red cell: a new theoretical approach. Part I. Crenation. Journal of Theoretical Biology, 86(3):513–529, October 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519380903501.

Brailsford:1980:CCRb

[BKB80b] J. Douglas Brailsford, Ralph A. Korpman, and Brian S. Bull. Crenation and cupping of the red cell: a new theoretical approach. Part II. Cupping. Journal of Theoretical Biology, 86(3):531–546, October 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519380903513.

Bunow:1980:PFR

[BKJT80] Barry Bunow, Jean-Pierre Kernevez, Gislaine Joly, and Daniel Thomas. Pattern formation by reaction–diffusion instabilities: Application to morphogenesis in drosophila. Journal of Theoretical Biology, 84(4):629–649, June 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519380800245.

Berding:1986:PDA

[BKMS86] C. Berding, A. E. Keymer, J. D. Murray, and A. F. G. Slater. The population dynamics of acquired immunity to helminth infection. Journal of Theoretical Biology, 122(4):459–471, October 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386801862.

Berding:1987:PDA

[BKMS87] C. Berding, A. E. Keymer, J. D. Murray, and A. F. G. Slater. The population dynamics of acquired immunity to helminth infection: Experimental and natural transmission. Journal of Theoretical Biology, 126(2):167–182, May 21, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387802278.

Bookstaber:1985:OCB

[BL85] Richard Bookstaber and Joseph Langsam. On the optimality of coarse behavior rules. Journal of Theoretical Biology, 116(2):161–193, September 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385802629.

Beauchamp:1988:FFC

[BL88] Guy Beauchamp and Louis Lefebvre. Food finding in colonially nesting birds. Journal of Theoretical Biology, 132(3):357– 368, June 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519388802200.

Blake:1980:CCT

[Bla80] Iván Ortega Blake. A.C. cable theory and its implications on the measurements of the membrane electric parameters. Journal of Theoretical Biology, 83(4):595–621, April 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380901915.

Blake:1982:FPA

[BLA82] J. R. Blake, N. Liron, and G. K. Aldis. Flow patterns around ciliated microorganisms and in ciliated ducts. Journal of Theoretical Biology, 98(1):127–141, September 7, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382900625.

Blank:1984:MAV

[Bla84] Martin Blank. Molecular association and the viscosity of hemoglobin solutions. Journal of Theoretical Biology, 108(1):55– 64, May 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251938480168X.

Brooks:1984:IES

[BLC84] Daniel R. Brooks, Paul H. Leblond, and D. David Cumming. Information and entropy in a simple evolution model. Journal of Theoretical Biology, 109(1):77–93, July 7, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801125.

Bremer:1986:APC

[BLC86] H. Bremer and S. Lin-Chao. Analysis of the physiological control of replication of ColE1-type plasmids. Journal of Theoretical Biology, 123(4):453–470, December 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386802132.

Bleecken:1988:MFC

[Ble88] Stefan Bleecken. Model for the feedback control system of bacterial growth I. Growth in discontinuous culture. Journal of Theoretical Biology, 133(1):37–65, July 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388800237.

Bleecken:1989:MFC

[Ble89] Stefan Bleecken. Model for the feedback control system of bacterial growth II. Growth in continuous culture. Journal of Theoretical Biology, 141(3):325–362, December 7, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389801183.

Blomberg:1985:OEC

[Blo85] Clas Blomberg. Optimization of error-correction processes with respect to time. Comparison to free energy aspects. Journal of Theoretical Biology, 115(2):241–268, July 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800990.

Blomberg:1987:FET

[Blo87] Clas Blomberg. Free energy and time economy for the mutual selection of monomers in biosynthesis, primarily protein synthesis. Journal of Theoretical Biology, 128(1):87–107, September 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519387800334.

Blake:1983:IAA

[BLR83] I. Ortega Blake, A. Leś, and S. Rybak. The interaction of γaminobutiric acid with hydrated Ca^{2+} and Mg^{2+}. A pseudopotential ab initio study. Journal of Theoretical Biology, 104(4):571– 590, October 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383902461.

Blute:1983:SDD

[Blu83] Marion Blute. “selfish” DNA and differential parental investment: some implications for sex chromosomes. Journal of Theoretical Biology, 102(4):603–610, June 21, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383903934.

Blum:1984:HRC

[Blu84] Jacob J. Blum. Hormone receptor-coupling factor-adenylate cyclase interaction: Theoretical considerations. Journal of Theoretical Biology, 111(3):589–608, December 7, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384802398.

Blumenthal:1985:PPO

[Blu85] Steven J. Blumenthal. Palindromic polynucleotides and the origin of biological chirality. Journal of Theoretical Biology, 114(4):665– 667, June 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385800515.

Balding:1985:MMT

[BM85] D. Balding and D. L. S. McElwain. A mathematical model of tumour-induced capillary growth. Journal of Theoretical Biology, 114(1):53–73, May 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519385802551.

Bardsley:1987:UNL

[BM87] William G. Bardsley and Philip B. McGinlay. The use of nonlinear regression analysis and the F test for model discrimination with dose-response curves and ligand binding data. Journal of Theoretical Biology, 126(2):183–201, May 21, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938780228X.

Bass:1988:FRT

[BM88] L. Bass and A. McNabb. Flux ratio theorems for nonlinear membrane transport under nonstationary conditions. Journal of Theoretical Biology, 133(2):185–191, July 21, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388800043.

Bachouchi:1989:BCS

[BM89a] Nacima Bachouchi and Jean E. Morel. Behaviour of the crossbridges in stretched or compressed muscle fibres. Journal of Theoretical Biology, 141(2):143–157, November 21, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389800141.

Bachouchi:1989:MCV

[BM89b] Nacima Bachouchi and Jean E. Morel. Muscle contraction and in vitro movements. Journal of Theoretical Biology, 141(3):425–427, December 7, 1989. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519389801249.

Bardsley:1989:ODM

[BMR89] W. G. Bardsley, P. B. McGinlay, and M. G. Roig. Optimal design for model discrimination using the F -test with non-linear biochemical models. Criteria for choosing the number and spacing of experimental points. Journal of Theoretical Biology, 139(1): 85–102, July 10, 1989. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519389800591.

Bernhardt:1980:AGF

[BN80] J. Bernhardt and E. Neumann. Analysis of gated flux from or into sealed membrane fragments. Journal of Theoretical Biology, 86(4):649–661, October 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380903045.

Belmont:1981:PTC

[BN81] A. Belmont and C. Nicolini. Polyelectrolyte theory and chromatin-DNA quaternary structure: Role of ionic strength and hl histone. Journal of Theoretical Biology, 90(2):169–179, May 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381900400.

Bresch:1980:HPP

[BNH80] C. Bresch, U. Niesert, and D. Harnasch. Hypercycles, parasites and packages. Journal of Theoretical Biology, 85(3):399– 405, August 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519380903148.

Berlin:1982:MBL

[BO82] R. D. Berlin and J. M. Oliver. The movement of bound ligands over cell surfaces. Journal of Theoretical Biology, 99(1):69–80, November 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382903897.

Bodnar:1988:DME

[Bod88] John W. Bodnar. A domain model for eukaryotic DNA organization: a molecular basis for cell differentiation and chromosome evolution. Journal of Theoretical Biology, 132(4):479–507, June 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388800869.

Bogart:1987:SDH

[Bog87] Mark H. Bogart. Sex determination: a hypothesis based on steroid ratios. Journal of Theoretical Biology, 128(3):349–357, October 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387800772.

Boldrini:1980:FAW

[Bol80] P. Boldrini. Ferroelectricity in the arterial wall: a new physical component of atherosclerosis. Journal of Theoretical Biology, 87(2):263–273, November 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380903604.

Boldrini:1982:IMC

[Bol82a] P. Boldrini. Intracellular magnesium crystallized in the arteries and myocardium: New chemical-physical hypothesis. Journal of Theoretical Biology, 95(4):641–664, April 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382903460.

Boldrini:1982:NCI

[Bol82b] P. Boldrini. New chemical interpretation of what calcifies in the artieries. Journal of Theoretical Biology, 95(2):325–344, March 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251938290248X.

Bont:1982:MLP

[Bon82] W. S. Bont. A model for the lymphocyte production pathway in bone marrow. Journal of Theoretical Biology, 96(4):707– 722, June 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382902387.

Bont:1985:GPN

[Bon85] Willy S. Bont. The geometric progression of nuclear volumes. Journal of Theoretical Biology, 112(3):505–512, February 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385800187.

Bookstein:1984:SMB

[Boo84a] Fred L. Bookstein. A statistical method for biological shape comparisons. Journal of Theoretical Biology, 107(3):475–520, April 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384801046.

Boon:1984:MAC

[Boo84b] Jean-Pierre Boon. Model for approaching colonies. Journal of Theoretical Biology, 110(3):501–503, October 7, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801897.

Boroujerdi:1982:CPC

[Bor82] Mehdi Boroujerdi. The comparability of pharmacokinetics of creatinine in rabbit and man: a mathematical approach. Journal of Theoretical Biology, 95(2):369–380, March 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251938290251X.

Bornbusch:1984:TFS

[Bor84] Alan H. Bornbusch. Turning field size and its effects upon computer-simulated klinotactic orientation. Journal of Theoretical Biology, 107(1):151–163, March 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801265.

Bormann:1988:EBM

[Bor88] E. J. Bormann. On the entropy balance of microbial product formation. Journal of Theoretical Biology, 133(2):215–221, July 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388800067.

Botkin:1982:CTT

[Bot82] Daniel B. Botkin. Can there be a theory of global ecology? Journal of Theoretical Biology, 96(1):95–98, May 7, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251938290159X.

Bourne:1984:HCR

[Bou84] I. H. J. Bourne. An hypothesis concerning the role of the Golgi apparatus in the phenomenon of memory. Journal of Theoretical Biology, 110(4):541–557, October 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801447.

Boy:1983:CCK

[Boy83] Andrew W. Boy. Cell cycle kinetic data can be simulated by a simple chemical kinetic model. Journal of Theoretical Biology, 101(3):355–372, April 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383901443.

Boyle:1988:MHB

[Boy88] Joseph Boyle. Microcirculatory hematocrit and blood flow. Journal of Theoretical Biology, 131(2):223–229, March 21, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388802388.

Boyd:1989:MAE

[Boy89] Robert Boyd. Mistakes allow evolutionary stability in the repeated prisoner’s dilemma game. Journal of Theoretical Biology, 136(1):47–56, January 9, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519389801882.

Bhat:1980:TSE

[BP80] Nalini Bhat and L. K. Pande. Three-species ecosystems in a solvable model. Journal of Theoretical Biology, 83(2):321–344, March 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380902969.

Bhat:1981:TSF

[BP81a] Nalini Bhat and L. K. Pande. Three-step food chains in Gompertz and Lotka–Volterra models. Journal of Theoretical Biology, 91(3):429–435, August 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902666.

Bivas:1981:FSI

[BP81b] Isak Bivas and Alexander G. Petrov. Flexoelectric and steric interactions between two bilayer lipid membranes resulting from their curvature fluctuations. Journal of Theoretical Biology, 88(3):459–483, February 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902770.

Bremermann:1983:GTM

[BP83] Hans J. Bremermann and John Pickering. A game-theoretical model of parasite virulence. Journal of Theoretical Biology, 100(3):411–426, February 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519383904381.

Brown:1985:OHR

[BP85] Jerram L. Brown and Stuart L. Pimm. The origin of helping: The role of variability in reproductive potential. Journal of Theoretical Biology, 112(3):465–477, February 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800151.

Barton:1986:SCA

[BP86] N. H. Barton and R. J. Post. Sibling competition and the advantage of mixed families. Journal of Theoretical Biology, 120(4):381–387, June 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519386800339.

Brendel:1987:NSM

[BP87] Volker Brendel and Alan S. Perelson. A note on stochastic models for bacterial adhesion. Journal of Theoretical Biology, 126(2):247–249, May 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519387802321.

Borstnik:1987:PMO

[BPH87] B. Borstnik, D. Pumpernik, and G. L. Hofacker. Point mutations as an optimal search process in biological evolution. Journal of Theoretical Biology, 125(3):249–268, April 7, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387800589.

Bliss:1982:RFI

[BPM82] Richard D. Bliss, Page R. Painter, and Allen G. Marr. Role of feedback inhibition in stabilizing the classical operon. Journal of Theoretical Biology, 97(2):177–193, July 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382900984.

Balbi:1986:SMD

[BPP^{+}86] C. Balbi, M. Pala, S. Parodi, G. Figari, B. Cavazza, V. Trefiletti, and E. Patrone. A simple model for DNA elution from filters. Journal of Theoretical Biology, 118(2):183–198, January 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386801333.

Budriene:1988:MMI

[BPP88] E. O. Budriené, A. A. Polezhaev, and M. O. Ptitsyn. Mathematical modelling of intercellular regulation causing the formation of spatial structures in bacterial colonies. Journal of Theoretical Biology, 135(3):323–341, December 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388802480.

Boyd:1983:CTA

[BR83] Robert Boyd and Peter J. Richerson. The cultural transmission of acquired variation: Effects on genetic fitness. Journal of Theoretical Biology, 100(4):567–596, February 21, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383903247.

Brown:1986:HSS

[BR86] Joel S. Brown and Michael L. Rosenzweig. Habitat selection in slowly regenerating environments. Journal of Theoretical Biology, 123(2):151–171, November 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801503.

Barton:1987:FSB

[BR87] N. H. Barton and S. Rouhani. The frequency of shifts between alternative equilibria. Journal of Theoretical Biology, 125(4):397– 418, April 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387802102.

Boyd:1988:ERS

[BR88] Robert Boyd and Peter J. Richerson. The evolution of reciprocity in sizable groups. Journal of Theoretical Biology, 132(3):337– 356, June 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519388802194.

Bremermann:1980:SPS

[Bre80a] Hans J. Bremermann. Sex and polymorphism as strategies in host-pathogen interactions. Journal of Theoretical Biology, 87(4): 671–702, ???? 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519380901113.

Brewer:1980:EDM

[Bre80b] Gregory J. Brewer. Exocytosis and directed movement of cytoplasmic vesicles toward the plasma membrane. Journal of Theoretical Biology, 85(1):75–82, July 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380902817.

Bremer:1981:SBG

[Bre81] H. Bremer. Simulated bacterial growth. Journal of Theoretical Biology, 92(1):23–38, September 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381901582.

Bremermann:1982:RPC

[Bre82] Hans J. Bremermann. Reliability of proliferation controls. The Hayflick limit and its breakdown in cancer. Journal of Theoretical Biology, 97(4):641–662, August 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382903642.

Bremermann:1983:TCD

[Bre83] Hans J. Bremermann. Theory of catastrophic diseases of cultivated plants. Journal of Theoretical Biology, 100(2):255–274, January 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383903508.

Bremer:1986:SPD

[Bre86] H. Bremer. A stockhastic process determines the time at which cell division begins in Escherichia coli. Journal of Theoretical Biology, 118(3):351–365, February 7, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386800662.

Brentani:1988:BIC

[Bre88] Ricardo R. Brentani. Biological implications of complementary hydropathy of amino acids. Journal of Theoretical Biology, 135(4):495–499, December 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388802728.

Brines:1980:DPS

[Bri80] Michael L. Brines. Dynamic patterns of skylight polarization as clock and compass. Journal of Theoretical Biology, 86(3):507– 512, October 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519380903495.

Britton:1982:MSD

[Bri82] N. F. Britton. A model for suppression and dominance in trees. Journal of Theoretical Biology, 97(4):691–698, August 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382903678.

Briggs:1984:CEH

[Bri84] W. E. Briggs. Cooperativity and extrema of the Hill slope for symmetric protein-ligand binding polynomials. Journal of Theoretical Biology, 108(1):77–83, May 7, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801708.

Briggs:1985:RBZ

[Bri85] W. E. Briggs. The relationship between zeros and factors of binding polynomials and cooperativity in protein-ligand binding. Journal of Theoretical Biology, 114(4):605–614, June 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385800473.

Briggs:1988:PLB

[Bri88] W. E. Briggs. Protein-ligand binding with a missing species. Journal of Theoretical Biology, 133(4):429–436, August 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388803321.

Britton:1989:ACE

[Bri89] N. F. Britton. Aggregation and the competitive exclusion principle. Journal of Theoretical Biology, 136(1):57–66, January 9, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519389801894.

Benegas:1982:DSS

[BRM82] Julio C. Benegas, Daniel R. Ripoll, and Ezio Marchi. Denaturation of β-sheet structure. Journal of Theoretical Biology, 94(1):111–118, January 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519382903332.

Brown:1980:ERO

[Bro80a] A. P. Brown. Evidence for remnants of an ordered codon sequence and a restricted codon composition in selected proteins. Journal of Theoretical Biology, 83(4):537–560, April 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380901885.

Brown:1980:ANL

[Bro80b] Rodney Jay Brown. Application of non-linear least squares fitting to analysis of inhibition in biphasic kinetic systems. Journal of Theoretical Biology, 83(3):527–531, April 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380900600.

Brockman:1981:LNP

[Bro81a] W. H. Brockman. Ladder network prediction of transients in linear spatially inhomogeneous cables. Journal of Theoretical Biology, 92(4):469–478, October 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902599.

Brown:1981:IAC

[Bro81b] Charles Eric Brown. Interactions among carnosine, anserine, ophidine and copper in biochemical adaptation. Journal of Theoretical Biology, 88(2):245–256, January 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381900734.

Brookfield:1982:IRD

[Bro82a] J. F. Y. Brookfield. Interspersed repetitive DNA sequences are unlikely to be parasitic. Journal of Theoretical Biology, 94(2):281–299, January 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519382903137.

Brown:1982:OGS

[Bro82b] Jerram L. Brown. Optimal group size in territorial animals. Journal of Theoretical Biology, 95(4):793–810, April 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251938290354X.

Brown:1983:CLA

[Bro83] A. P. Brown. Codon-level analysis of histone primary sequence: Evidence of a repeat tetrapeptide origin and later inclusion of transcribed sequence. Journal of Theoretical Biology, 104(3):401– 416, October 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383901145.

Bass:1987:RBE

[BRR87] L. Bass, M. S. Roberts, and P. J. Robinson. On the relation between extended forms of the sinusoidal perfusion and of the convection-dispersion models of hepatic elimination. Journal of Theoretical Biology, 126(4):457–482, June 21, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387801522.

Bajaj:1987:MMI

[BRRK87] J. S. Bajaj, Gita Subba Rao, J. Subba Rao, and R. Khardori. A mathematical model for insulin kinetics and its application to protein-deficient (malnutrition-related) diabetes mellitus (PDDM). Journal of Theoretical Biology, 126(4):491–503, June 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519387801546.

Brugge:1985:MMG

[Bru85a] R. Brugge. A mechanistic model of grass root growth and development dependent upon photosynthesis and nitrogen uptake. Journal of Theoretical Biology, 116(3):443–467, October 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385802812.

Bruner:1985:TSM

[Bru85b] L. J. Bruner. A three state model for alamethicin conductance in bilayer membranes. Journal of Theoretical Biology, 117(2):265–276, November 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519385802216.

Barnard:1989:EMI

[BRW89] Ross Barnard, S. W. Rowlinson, and M. J. Waters. An electrostatic model for the interaction between growth hormone and its receptor involving chelation of Ca^{2+} to the human growth hormone molecule. Journal of Theoretical Biology, 140(3):355–367, October 9, 1989. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251938980092X.

Bryan:1980:TWP

[Bry80] William P. Bryan. The thermodynamics of water-protein interactions. Journal of Theoretical Biology, 87(4):639–661, ???? 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380901095.

Bade:1981:CCK

[BS81] M. L. Bade and A. Stinson. Chitin and chitinase: a kinetic model. Journal of Theoretical Biology, 93(4):697–700, December 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381903350.

Brauer:1982:CEH

[BS82] F. Brauer and A. C. Soudack. On constant effort harvesting and stocking in a class of predator–prey systems. Journal of Theoretical Biology, 95(2):247–252, March 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902429.

Balasubramanian:1983:CRW

[BS83] R. Balasubramanian and P. Seetharamulu. A conformational rationale for the wobble behaviour of the first base of the anticodon triplet in tRNA. Journal of Theoretical Biology, 101(1): 77–86, March 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383902734.

Birky:1984:RPC

[BS84a] C. William Birky and Russell V. Skavaril. Random partitioning of cytoplasmic organelles at cell division: The effect of organelle and cell volume. Journal of Theoretical Biology, 106(4):441–447, February 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519384900018.

Briedis:1984:ETE

[BS84b] D. Briedis and R. C. Seagrave. Energy transformation and entropy production in living systems I. Applications to embryonic growth. Journal of Theoretical Biology, 110(2):173–193, September 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251938480051X.

Balasubramanian:1985:OLC

[BS85] R. Balasubramanian and P. Seetharamulu. Origins of life: Conformational energy calculations on primitive tRNA nestling an amino acid. Journal of Theoretical Biology, 113(1):15–28, March 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385800734.

Brendel:1987:MRR

[BS87] Volker Brendel and Lee A. Segel. On modes of recombination, replication, and segregation of the higher plant mitochondrial genome. Journal of Theoretical Biology, 125(2):163–176, March 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519387800383.

Bains:1988:NMN

[BS88a] William Bains and Geoff C. Smith. A novel method for nucleic acid sequence determination. Journal of Theoretical Biology, 135(3):303–307, December 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388802467.

Barchilon:1988:AOR

[BS88b] Mercedes Barchilon and Lee A. Segel. Adaptation, oscillations and relay in a model for cAMP secretion in cellular slime molds. Journal of Theoretical Biology, 133(4):437–446, August 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388803333.

Britton:1989:MMG

[BS89] N. F. Britton and Suzanne M. Skevington. A mathematical model of the gate control theory of pain. Journal of Theoretical Biology, 137(1):91–105, March 7, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389801511.

Boguta:1988:TECa

[BSB88a] Grzegorz Boguta, Dariusz Stepkowski, and Andrzej Bierzyński. Theoretical estimation of the calcium-binding constants for proteins from the troponin C superfamily based on a secondary structure prediction method I. Estimation procedure. Journal of Theoretical Biology, 135(1):41–61, November 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388801735.

Boguta:1988:TECb

[BSB88b] Grzegorz Boguta, Dariusz Stepkowski, and Andrzej Bierzyński. Theoretical estimation of the calcium-binding constants for proteins from the troponin C superfamily based on a secondary structure prediction method: II. Applications. Journal of Theoretical Biology, 135(1):63–73, November 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388801747.

Brown:1982:ESB

[BSM82] Joel S. Brown, Michael J. Sanderson, and Richard E. Michod. Evolution of social behavior by reciprocation. Journal of Theoretical Biology, 99(2):319–339, November 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251938290008X.

Balaz:1988:KDA

[BSR^{+}88] S. Baláz, E. Sturdík, M. Rosenberg, J. Augustín, and B. Skára. Kinetics of drug activities as influenced by their physicochemical properties: Antibacterial effects of alkylating 2furylethylenes. Journal of Theoretical Biology, 131(1):115–134, March 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519388801255.

Bomze:1983:RMG

[BSS83] Immanuel M. Bomze, Peter Schuster, and Karl Sigmund. The role of Mendelian genetics in stragetic models on animal behaviour. Journal of Theoretical Biology, 101(1):19–38, March 7, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383902710.

Bulmer:1980:SRU

[BT80] M. G. Bulmer and P. D. Taylor. Sex ratio under the haystack model. Journal of Theoretical Biology, 86(1):83–89, September 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380900661.

Busby:1981:ASL

[BT81] Robert C. Busby and Ronald J. Tallarida. On the analysis of straight line data in pharmacology and biochemistry. Journal of Theoretical Biology, 93(4):867–879, December 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381903453.

Berkowitz:1984:PAS

[BT84] Gary C. Berkowitz and Robert D. Tschirgi. On the phylogeny of asymmetry and spatial discrimination. Journal of Theoretical Biology, 106(4):495–528, February 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519384900043.

Bergersen:1985:FAA

[BT85] Fraser J. Bergersen and Jean-Charles Trinchant. Flux of O_{2} affects apparent optimal concentrations of O_{2} in suspensions of bacteria respiring microaerobically. Journal of Theoretical Biology, 115(1):93–102, July 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800084.

Bucher:1987:NCE

[BT87] Philipp Bucher and Edward N. Trifonov. On Nussinov’s compilation of eukaryotic transcription initiation sites. Journal of Theoretical Biology, 126(3):373–375, June 7, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387802436.

Bulmer:1980:SCM

[Bul80] M. G. Bulmer. The sib competition model for the maintenance of sex and recombination. Journal of Theoretical Biology, 82(2):335– 345, January 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519380901071.

Bulmer:1981:WQC

[Bul81] M. G. Bulmer. Worker-queen conflict in annual social hymenoptera. Journal of Theoretical Biology, 93(1):239–251, November 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381900680.

Bulmer:1982:CPC

[Bul82] M. G. Bulmer. Cyclical parthenogenesis and the cost of sex. Journal of Theoretical Biology, 94(1):197–207, January 7, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382903411.

Bulmer:1983:SRT

[Bul83] M. G. Bulmer. Sex ratio theory in social insects with swarming. Journal of Theoretical Biology, 100(2):329–339, January 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383903569.

Bulmer:1984:RAN

[Bul84] M. G. Bulmer. Risk avoidance and nesting strategies. Journal of Theoretical Biology, 106(4):529–535, February 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519384900055.

Bulmer:1988:CUI

[Bul88] Michael Bulmer. Codon usage and intragenic position. Journal of Theoretical Biology, 133(1):67–71, July 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388800249.

Bunow:1980:CES

[Bun80] Barry Bunow. Cellular enzymology: The steady-state kinetics of compartmentalized enzymes. Journal of Theoretical Biology, 84(4):611–627, June 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519380800233.

Burgess:1982:PEC

[Bur82a] Antony W. Burgess. Prebiotic evolution: Conformational perspectives. Journal of Theoretical Biology, 96(1):21–38, May 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382901539.

Burnett:1982:MMC

[Bur82b] Leslie Burnett. A model for the mechanism and control of eukaryote gene splicing. Journal of Theoretical Biology, 97(3):351– 366, August 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382903708.

Burger:1983:EDM

[Bur83] Reinhard Burger. On the evolution of dominance modifiers I. A nonlinear analysis. Journal of Theoretical Biology, 101(4):585– 598, April 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383900176.

Busetta:1982:IRR

[Bus82] Bernard Busetta. Improving the residual representation of proteins. Journal of Theoretical Biology, 98(4):621–635, October 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382901412.

Boulanger:1987:IZI

[BVDA87] Thierry Boulanger, Daniel P. Vercauteren, Francois Durant, and Jean-Marie Andre. 3- and 5-isoxazolol zwitterions: an ab initio molecular orbital study relating to GABA agonism and antagonism. Journal of Theoretical Biology, 127(4):479–489, August 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519387801443.

Brugman:1984:SAR

[BvdHB84] C. J. M. Brugman, H. van den Hooff, and T. Blangé. Some aspects of the role of quantum mechanics in the theory of muscle contraction. Journal of Theoretical Biology, 107(1):173–177, March 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384801290.

Blake:1983:MOT

[BVW83] J. R. Blake, P. G. Vann, and H. Winet. A model of ovum transport. Journal of Theoretical Biology, 102(1):145–166, May 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383902679.

Bhatia:1981:AGD

[BW81] Kuldeep Bhatia and Susan R. Wilson. The application of gene diversity analyses to surname diversity data. Journal of Theoretical Biology, 88(1):121–133, January 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381903313.

Buse:1985:HBH

[BW85] M. Buse and J. Werner. Heat balance of the human body: Influence of variations of locally distributed parameters. Journal of Theoretical Biology, 114(1):34–51, May 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938580254X.

Bermudez:1986:EPM

[BW86a] J. Bermudez and J. Wagensberg. On the entropy production in microbiological stationary states. Journal of Theoretical Biology, 122(3):347–358, October 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801266.

Binggeli:1986:MPS

[BW86b] Richard Binggeli and Roy C. Weinstein. Membrane potentials and sodium channels: Hypotheses for growth regulation and cancer formation based on changes in sodium channels and gap junctions. Journal of Theoretical Biology, 123(4):377–401, December 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386802090.

Britton:1987:MMG

[BW87] N. F. Britton and A. E. Wheals. Mathematical models for a G_{0} phase in Saccharomyces cerevisiae. Journal of Theoretical Biology, 125(3):269–281, April 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387800590.

Bergman:1989:SFE

[BW89] Douglas A. Bergman and Donald J. Winzor. Space-filling effects of inert solutes as probes for the detection and study of substratemediated conformational changes by enzyme kinetics: Theoretical considerations. Journal of Theoretical Biology, 137(2):171– 189, March 21, 1989. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519389802048.

Bardsley:1980:RBM

[BWM80] William G. Bardsley, Richard Woolfson, and J.-P. Mazat. Relationships between the magnitude of Hill plot slopes, apparent binding constants and factorability of binding polynomials and their hessians. Journal of Theoretical Biology, 85(2):247– 284, July 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251938090020X.

Britton:1982:MMC

[BWM82] N. F. Britton, N. A. Wright, and J. D. Murray. A mathematical model for cell population kinetics in the intestine. Journal of Theoretical Biology, 98(3):531–541, October 7, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382901357.

Bardsley:1980:SMR

[BWW80] William G. Bardsley, Richard Woolfson, and Reg M. W. Wood. Some mathematical results concerning Hessians of binding polynomials and co-operativity coefficients. Journal of Theoretical Biology, 85(1):45–51, July 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380902799.

Czerlinski:1989:MAA

[CA89] George H. Czerlinski and Byron Anderson. Metabolic amplification: Analysis of the kinetic detection of low substrate levels. Journal of Theoretical Biology, 139(2):187–200, July 21, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519389800992.

Calder:1982:TBS

[Cal82] William A. Calder. A tradeoff between space and time: Dimensional constants in mammalian ecology. Journal of Theoretical Biology, 98(3):393–400, October 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382901266.

Calder:1983:AAP

[Cal83a] William A. Calder. An allometric approach to population cycles of mammals. Journal of Theoretical Biology, 100(2):275–282, January 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251938390351X.

Calder:1983:BSM

[Cal83b] William A. Calder. Body size, mortality, and longevity. Journal of Theoretical Biology, 102(1):135–144, May 7, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383902667.

Calvin:1983:STL

[Cal83c] William H. Calvin. A stone’s throw and its launch window: Timing precision and its implications for language and hominid brains. Journal of Theoretical Biology, 104(1):121–135, September 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383904058.

Calder:1987:MAB

[Cal87] William A. Calder. Metabolic allometry: Basic correlations are independent of units when properly converted. Journal of Theoretical Biology, 128(4):523–524, October 21, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387801959.

Cann:1987:TEH

[Can87] John R. Cann. Theory of electrophoresis of hybridizing enzymes with kinetic control: Implications for population genetics of electrophoretic markers. Journal of Theoretical Biology, 127(4):461– 477, August 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387801431.

Carlin:1980:PNE

[Car80] Richard K. Carlin. Poly(a): a new evolutionary probe. Journal of Theoretical Biology, 82(3):353–362, February 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380902428.

Carpenter:1981:DHD

[Car81a] Stephen R. Carpenter. Decay of heterogeneous detritus: a general model. Journal of Theoretical Biology, 89(4):539–547, April 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381900266.

Carpenter:1981:ECM

[Car81b] Stephen R. Carpenter. Effect of control measures on pest populations subject to regulation by parasites and pathogens. Journal of Theoretical Biology, 92(2):181–184, September 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251938190391X.

Carroll:1982:E

[Car82a] James P. Carroll. On emmetropization. Journal of Theoretical Biology, 95(1):135–144, March 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902922.

Caryl:1982:TTA

[Car82b] P. G. Caryl. Telling the truth about intentions. Journal of Theoretical Biology, 97(4):679–689, August 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382903666.

Carpenter:1983:LGI

[Car83] Stephen R. Carpenter. Lake geometry: Implications for production and sediment accretion rates. Journal of Theoretical Biology, 105(2):273–286, ???? 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519383800083.

Caro:1985:ISC

[Car85] T. M. Caro. Intersexual selection and cooption. Journal of Theoretical Biology, 112(2):275–277, January 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385802873.

Casolari:1981:MDM

[Cas81] Antonio Casolari. A model describing microbial inactivation and growth kinetics. Journal of Theoretical Biology, 88(1):1– 34, January 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251938190326X.

Caswell:1982:OLH

[Cas82] Hal Caswell. Optimal life histories and the age-specific costs of reproduction. Journal of Theoretical Biology, 98(3):519–529, October 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382901345.

Caswell:1984:OLH

[Cas84] Hal Caswell. Optimal life histories and age-specific costs of reproduction: Two extensions. Journal of Theoretical Biology, 107(1):169–172, March 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519384801289.

Case:1981:RAF

[CB81] Ted J. Case and Edward A. Bender. Is recombination advantageous in fluctuating and spatially heterogeneous environments? Journal of Theoretical Biology, 90(2):181–190, May 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381900412.

Cornish-Bowden:1983:PMC

[CB83] Athel Cornish-Bowden. Phenetic methods of classification use information that is disregarded by minimum-length methods. Journal of Theoretical Biology, 101(2):317–319, March 21, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383903430.

Cornish-Bowden:1984:ESM

[CB84] Athel Cornish-Bowden. Enzyme specificity: Its meaning in the general case. Journal of Theoretical Biology, 108(3):451– 457, June 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384800454.

Charnov:1985:SAP

[CB85] E. L. Charnov and J. J. Bull. Sex allocation in a patchy environment: a marginal value theorem. Journal of Theoretical Biology, 115(4):619–624, August 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801442.

Caraco:1986:GBC

[CB86a] Thomas Caraco and Jerram L. Brown. A game between communal breeders: When is food-sharing stable? Journal of Theoretical Biology, 118(4):379–393, February 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801606.

Charnov:1986:SAP

[CB86b] Eric L. Charnov and James J. Bull. Sex allocation, pollinator attraction and fruit dispersal in cosexual plants. Journal of Theoretical Biology, 118(3):321–325, February 7, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386800625.

Cote:1986:CRD

[CB86c] Gary G. Coté and Stuart Brody. Circadian rhythms in Drosophila melanogaster: Analysis of period as a function of gene dosage at the per (period) locus. Journal of Theoretical Biology, 121(4):487–503, August 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801047.

Chen:1987:QEM

[CB87a] Isaac I. H. Chen and Harold D. Battarbee. The quantitative estimation of microvessels in microvascular networks. Journal of Theoretical Biology, 127(3):315–320, August 7, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387801091.

Cornish-Bowden:1987:DI

[CB87b] Athel Cornish-Bowden. Dominance is not inevitable. Journal of Theoretical Biology, 125(3):333–338, April 7, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387800656.

Cornish-Bowden:1988:SFN

[CB88a] Athel Cornish-Bowden. Saturation functions as a nested set. Journal of Theoretical Biology, 130(1):125–126, January 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388801681.

Cornish-Bowden:1988:SPP

[CB88b] Athel Cornish-Bowden. Significance of the purine-pyrimidine motif present in most gene groups. Journal of Theoretical Biology, 134(1):1–7, September 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388802996.

Charnov:1989:PSR

[CB89a] Eric L. Charnov and James J. Bull. The primary sex ratio under environmental sex determination. Journal of Theoretical Biology, 139(4):431–436, August 22, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389800633.

Cornish-Bowden:1989:MCT

[CB89b] Athel Cornish-Bowden. Metabolic control theory and biochemical systems theory: Different objectives, different assumptions, different results. Journal of Theoretical Biology, 136(4):365–377, February 22, 1989. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519389801547.

Cornish-Bowden:1987:COM

[CBC87] Athel Cornish-Bowden and María Luz Cárdenas. Co-operativity in monomeric enzymes. Journal of Theoretical Biology, 124(1):1– 23, January 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387802485.

Colacino:1987:MOS

[CBF87] J. M. Colacino, J. R. Brannan, and Mary Anne Fields. A model for oxygen storage by hemoglobin. Journal of Theoretical Biology, 126(4):437–447, June 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387801509.

Crowley:1981:PDM

[CBLM81] Philip H. Crowley, Craig J. Benham, Suzanne M. Lenhart, and Janet L. Morgan. Peripheral doublet microtubules and wave generation in eukaryotic flagella. Journal of Theoretical Biology, 93(4):769–784, December 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381903398.

Churchward:1982:MCB

[CBY82] Gordon Churchward, Hans Bremer, and Ry Young. Macromolecular composition of bacteria. Journal of Theoretical Biology, 94(3):651–670, February 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382903058.

Charlesworth:1981:KSA

[CC81] Brian Charlesworth and Eric L. Charnov. Kin selection in age-structured populations. Journal of Theoretical Biology, 88(1):103–119, January 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519381903301.

Casper:1982:SAH

[CC82a] Brenda B. Casper and Eric L. Charnov. Sex allocation in heterostylous plants. Journal of Theoretical Biology, 96(2):143– 149, May 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251938290217X.

Charton:1982:SDA

[CC82b] Marvin Charton and Barbara I. Charton. The structural dependence of amino acid hydrophobicity parameters. Journal of Theoretical Biology, 99(4):629–644, December 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382901916.

Charton:1983:DCF

[CC83a] Marvin Charton and Barbara I. Charton. The dependence of the Chou-fasman parameters on amino acid side chain structure. Journal of Theoretical Biology, 102(1):121–134, May 7, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383902655.

Coulombe:1983:SPA

[CC83b] Alain Coulombe and Edouard Coraboeuf. Simulation of potassium accumulation in clefts of Purkinje fibers: Effect on membrane electrical activity. Journal of Theoretical Biology, 104(2):211–229, September 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383904113.

Cohen:1981:IHT

[CCDF81] Patricia Cohen, Barbara Connetta, Douglas Dix, and John Flannery. The incidence of hematologic tumours: a cellular model for the age dependence. Journal of Theoretical Biology, 90(3):427– 436, June 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381903222.

Calvo:1982:ACC

[CCG^{+}82] A. Calvo, J. L. Cruz, C. Gutiérrez, G. García-Herdugo, and … J. F. López-Sáez. Analysis of cell cycle in root meristems. Journal of Theoretical Biology, 96(2):295–308, May 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902260.

Cowan:1987:BDS

[CCG87] Richard Cowan, Christina M. Collis, and Geoffrey W. Grigg. Breakage of double-stranded DNA due to single-stranded nicking. Journal of Theoretical Biology, 127(2):229–245, July 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519387801339.

Chen:1984:MMC

[CD84] I. I. H. Chen and J. N. Diana. Mathematical modeling of capillary density. Journal of Theoretical Biology, 108(2):221–225, May 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384800685.

Cressman:1987:DDE

[CD87] R. Cressman and A. T. Dash. Density dependence and evolutionary stable strategies. Journal of Theoretical Biology, 126(4):393– 406, June 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387801479.

Chatterjee:1988:SDS

[CDC88] Avik P. Chatterjee, Anjan Dasgupta, and Anadi N. Chatterjee. Spatial dependence of stress distribution for rod-shaped bacteria. Journal of Theoretical Biology, 135(3):309–321, December 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388802479.

Comins:1982:SRW

[CF82] H. N. Comins and G. D. Farquhar. Stomatal regulation and water economy in crassulacean acid metabolism plants: an optimization model. Journal of Theoretical Biology, 99(2):263–284, November 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382900042.

Chela-Flores:1985:ECP

[CF85] Julián Chela-Flores. Evolution as a collective phenomenon. Journal of Theoretical Biology, 117(1):107–118, November 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385801661.

Chela-Flores:1987:TCB

[CF87a] Julian Chela-Flores. Towards a collective biology of the gene. Journal of Theoretical Biology, 126(2):127–136, May 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519387802242.

Culpepper:1987:AMD

[CF87b] Michael I. Culpepper and Kennon T. Francis. An anatomical model to determine step height in step testing for estimating aerobic capacity. Journal of Theoretical Biology, 129(1):1–8, November 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519387801996.

Chela-Flores:1988:KTP

[CFLF88] Julian Chela-Flores, Alfonso M. Liquori, and Arnaldo Florio. A kinetic thermodynamic phenomenological approach to genetic expression of heat-shock proteins. Journal of Theoretical Biology, 134(3):319–325, October 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388800626.

Callow:1980:CSH

[CG80] R. S. Callow and I. Gladwell. Chromosome synapsis in hexaploids. Journal of Theoretical Biology, 87(4):703–722, ???? 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380901125.

Cheung:1981:SFMa

[CG81a] A. S. Cheung and B. F. Gray. The sliding filament model of muscle contraction. III. Stability analysis and sinusoidal perturbations. Journal of Theoretical Biology, 91(3):493–506, August 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381902691.

Cheung:1981:SFMb

[CG81b] A. S. Cheung and B. F. Gray. The sliding filament model of muscle contraction. IV. the effect of variation of ATP concentration. Journal of Theoretical Biology, 91(3):507–513, August 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381902708.

Callow:1984:GTC

[CG84] R. S. Callow and I. Gladwell. A general treatment of chromosome synapsis in even-numbered polyploids. Journal of Theoretical Biology, 106(4):455–494, February 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519384900031.

Carpenter:1985:NTC

[CG85a] Gail A. Carpenter and Stephen Grossberg. A neural theory of circadian rhythms: Split rhythms, after-effects and motivational interactions. Journal of Theoretical Biology, 113(1):163–223, March 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385800837.

Casey:1985:TMA

[CG85b] Richard M. Casey and Abbot S. Gaunt. Theoretical models of the avian syrinx. Journal of Theoretical Biology, 116(1):45–64, September 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385801302.

Chandler:1980:AGP

[CH80] H. D. Chandler and H. R. Hepburn. The analysis of globular proteins as defect containing cylindrical crystals. Journal of Theoretical Biology, 84(2):311–334, May 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519380800099.

Coutelle:1982:ISP

[CH82] R. Coutelle and G. L. Hofacker. Influence of selective processes on the amino acid compositions of proteins: Collagen, cytochrome c, ferredoxin and α-crystallin. Journal of Theoretical Biology, 95(4):615–639, April 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382903459.

Conrad:1985:SCE

[CH85] Michael Conrad and Harold M. Hastings. Scale change and the emergence of information processing primitives. Journal of Theoretical Biology, 112(4):741–755, February 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800576.

Cartwright:1986:MCT

[CH86] Mark Cartwright and Masud Husain. A model for the control of testosterone secretion. Journal of Theoretical Biology, 123(2):239–250, November 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519386801588.

Chalker:1980:MLS

[Cha80a] Bruce E. Chalker. Modeling light saturation curves for photosynthesis: An exponential function. Journal of Theoretical Biology, 84(2):205–215, May 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938080004X.

Chapman:1980:CBT

[Cha80b] J. Brian Chapman. Consistency between thermodynamics and the kinetics of n, m, and h in the Hodgkin–Huxley equations. Journal of Theoretical Biology, 85(3):487–495, August 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380903227.

Chapman:1980:CVR

[Cha80c] J. Brian Chapman. On current-voltage relations of ionic channels. Journal of Theoretical Biology, 85(1):165–169, July 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380902878.

Charlesworth:1980:CSR

[Cha80d] B. Charlesworth. The cost of sex in relation to mating system. Journal of Theoretical Biology, 84(4):655–671, June 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519380800269.

Charlesworth:1980:CMA

[Cha80e] Brian Charlesworth. The cost of meiosis with alternation of sexual and asexual generations. Journal of Theoretical Biology, 87(3):517–528, December 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380902337.

Charton:1981:PFG

[Cha81] M. Charton. Protein folding and the genetic code: An alternative quantitative model. Journal of Theoretical Biology, 91(1):115– 123, July 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381903775.

Chapman:1982:KIV

[Cha82a] J. Brian Chapman. A kinetic interpretation of “variable” stoichiometry for an electrogenic sodium pump obeying chemiosmotic principles. Journal of Theoretical Biology, 95(4):665–678, April 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382903472.

Charlesworth:1982:RBC

[Cha82b] Brian Charlesworth. Reply to Bulmer on the cost of sex with cyclical parthenogenesis. Journal of Theoretical Biology, 94(1):223–224, January 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519382903435.

Charnov:1986:SAM

[Cha86] Eric L. Charnov. Size advantage may not always favor sex change. Journal of Theoretical Biology, 119(3):283–285, April 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386801412.

Chao:1988:ESR

[Cha88a] Lin Chao. Evolution of sex in RNA viruses. Journal of Theoretical Biology, 133(1):99–112, July 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388800274.

Charlesworth:1988:EMC

[Cha88b] Brian Charlesworth. The evolution of mate choice in a fluctuating environment. Journal of Theoretical Biology, 130(2):191–204, January 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519388800948.

Chay:1988:KMC

[Cha88c] Teresa Ree Chay. Kinetic modeling for the channel gating process from single channel patch clamp data. Journal of Theoretical Biology, 132(4):449–468, June 21, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388800845.

Chang:1989:ECE

[Cha89a] You-Im Chang. The effect of cationic electrolytes on the electrostatic behavior of cellular surfaces with ionizable groups. Journal of Theoretical Biology, 139(4):561–571, August 22, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389800700.

Chappell:1989:FBL

[Cha89b] Rick Chappell. Fitting bent lines to data, with applications to allometry. Journal of Theoretical Biology, 138(2):235–256, May 22, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519389801419.

Charlesworth:1989:AMF

[Cha89c] D. Charlesworth. Allocation to male and female function in hermaphrodites, in sexually polymorphic populations. Journal of Theoretical Biology, 139(3):327–342, August 9, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389802127.

Chen:1983:MRV

[Che83] I. I. H. Chen. A mathematical representation for vessel networks II. Journal of Theoretical Biology, 104(4):647–654, October 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383902527.

Chen:1984:MRV

[Che84a] I. I. H. Chen. A mathematical representation for vessel network III. Journal of Theoretical Biology, 111(1):115–121, November 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938480199X.

Cheverud:1984:QGD

[Che84b] James M. Cheverud. Quantitative genetics and developmental constraints on evolution by selection. Journal of Theoretical Biology, 110(2):155–171, September 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384800508.

Chichilnisky:1986:MMP

[Chi86] Eduardo J. Chichilnisky. A mathematical model of pattern formation. Journal of Theoretical Biology, 123(1):81–101, November 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386802375.

Comins:1980:ESD

[CHM80] Hugh N. Comins, William D. Hamilton, and Robert M. May. Evolutionarily stable dispersal strategies. Journal of Theoretical Biology, 82(2):205–230, January 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380900995.

Chou:1981:TNS

[Cho81] Kuo-Chen Chou. Two new schematic rules for rate laws of enzyme-catalysed reactions. Journal of Theoretical Biology, 89(4):581–592, April 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519381900308.

Cheetham:1980:BSA

[CHT80] Alan H. Cheetham, Lee-Ann C. Hayek, and Erik Thomsen. Branching structure in arborescent animals: Models of relative growth. Journal of Theoretical Biology, 85(2):335–369, July 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380900259.

Craig:1982:ECB

[CJ82] John L. Craig and Peter F. Jenkins. The evolution of complexity in broadcast song of passerines. Journal of Theoretical Biology, 95(3):415–422, April 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382900285.

Colding-Jorgensen:1983:MFP

[CJ83] Morten Colding-Jørgensen. A model for the firing pattern of a paced nerve cell. Journal of Theoretical Biology, 101(4):541– 568, April 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383900152.

Childress:1980:LG

[CK80] Stephen Childress and Joseph B. Keller. Lichen growth. Journal of Theoretical Biology, 82(1):157–165, January 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380900958.

Cooper:1982:ACF

[CK82] William S. Cooper and Robert H. Kaplan. Adaptive “coinflipping”: a decision-theoretic examination of natural selection for random individual variation. Journal of Theoretical Biology, 94(1):135–151, January 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382903368.

Csanyi:1985:AER

[CK85] V. Csányi and G. Kampis. Autogenesis: The evolution of replicative systems. Journal of Theoretical Biology, 114(2):303–321, May 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385801089.

Cheer:1987:PRF

[CK87] A. Y. L. Cheer and M. A. R. Koehl. Paddles and rakes: Fluid flow through bristled appendages of small organisms. Journal of Theoretical Biology, 129(1):17–39, November 7, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387802011.

Conrad:1982:MCL

[CL82] Michael Conrad and E. A. Liberman. Molecular computing as a link between biological and physical theory. Journal of Theoretical Biology, 98(2):239–252, September 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902624.

Crispe:1984:CDE

[CL84] I. N. Crispe and N. B. LaThangue. Cancer is not due to the expression of introns. Journal of Theoretical Biology, 107(2):339– 340, March 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384800338.

Clapp:1981:EFC

[Cla81] Roger E. Clapp. Entropy flow in cyclic biological processes. Journal of Theoretical Biology, 92(1):15–21, September 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381901570.

Clarke:1982:EIP

[Cla82a] Colin H. Clarke. Evolutionary implications of potential frameshift hot-spots, methylation sites, and codon usage differences between genes and between organisms in Escherichia coli and salmonella typhimurium. Journal of Theoretical Biology, 96(3): 461–471, June 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382901217.

Clarke:1982:HSM

[Cla82b] Colin H. Clarke. Hypothetical stages in the mutational evolution of the present-day genetic code from its RNY, comma-free ancestor. Journal of Theoretical Biology, 99(2):397–403, November 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382900133.

Clarke:1982:IDB

[Cla82c] Colin H. Clarke. The influence of DNA base ratios on intragenic mutational spectra in prokaryotes. Journal of Theoretical Biology, 94(3):671–687, February 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251938290306X.

Clarke:1982:IBD

[Cla82d] Colin H. Clarke. Influences of bacterial DNA base-ratios and amino acid composition of the gene product on the consequences of frameshift mutations. Journal of Theoretical Biology, 98(4):661–674, October 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382901448.

Clarke:1982:MEA

[Cla82e] Colin H. Clarke. Mutational and evolutionary analysis of an archaebacterial gene. Journal of Theoretical Biology, 98(4):609–620, October 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382901400.

Clarke:1982:PSF

[Cla82f] Colin H. Clarke. Preferred sites of frameshift mutagenesis in AT- and GC-rich microbial DNA. Journal of Theoretical Biology, 94(3):519–534, February 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902983.

Clapp:1983:ZTA

[Cla83a] Roger E. Clapp. Zipper transition in an alpha-helix: a mechanism for gating of voltage-sensitive ion channels in a biological membrane. Journal of Theoretical Biology, 104(1):137–158, September 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251938390406X.

Clarke:1983:CBP

[Cla83b] Colin H. Clarke. The consequences of base-pair substitution mutations in AT- and GC-rich bacteria. Journal of Theoretical Biology, 105(1):117–131, November 7, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383904289.

Clark:1986:EES

[Cla86] Stephen J. Clark. ESS equations sometimes do not specify an ESS. Journal of Theoretical Biology, 123(3):325–331, December 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938680248X.

Cleave:1984:CVT

[CLF84] John P. Cleave, Michael R. Levine, and Peter J. Fleming. The control of ventilation: a theoretical analysis of the response to transient disturbances. Journal of Theoretical Biology, 108(2):261–283, May 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384800703.

Cleave:1986:HBS

[CLFL86] J. P. Cleave, M. R. Levine, P. J. Fleming, and A. M. Long. Hopf bifurcations and the stability of the respiratory control system. Journal of Theoretical Biology, 119(3):299–318, April 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386801436.

Clarke:1982:CFM

[CM82] C. H. Clarke and P. G. G. Miller. Consequences of frameshift mutations in the trp A, trp B and lac I genes of Escherichia coli and in Salmonella typhimurium. Journal of Theoretical Biology, 96(3):367–379, June 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382901163.

Cowan:1986:CPD

[CM86] Richard Cowan and Valerie B. Morris. Cell population dynamics during the differentiative phase of tissue development. Journal of Theoretical Biology, 122(2):205–224, September 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386800820.

Crawley:1987:PDP

[CM87] M. J. Crawley and R. M. May. Population dynamics and plant community structure: Competition between annuals and perrenials. Journal of Theoretical Biology, 125(4):475–489, April 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519387802151.

Cowan:1988:DRP

[CM88] Richard Cowan and Valerie B. Morris. Division rules for polygonal cells. Journal of Theoretical Biology, 131(1):33–42, March 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388801188.

Carlson:1986:CAP

[CMM86] Gail M. Carlson, Raymond J. MacDonald, and Edgar F. Meyer. Computer aided prediction and evaluation of the tertiary structure for rat elastase II. Journal of Theoretical Biology, 119(1):107–124, March 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386800558.

Cecchini:1981:SPI

[CMN81] Albert B. P. Cecchini, Julius Melbin, and Abraham Noordergraaf. Set-point: Is it a distinct structural entity in biological control? Journal of Theoretical Biology, 93(2):387–394, November 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381901107.

Cunningham:1980:TLC

[CN80] A. Cunningham and R. M. Nisbet. Time lag and co-operativity in the transient growth dynamics of microalgae. Journal of Theoretical Biology, 84(2):189–203, May 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519380800038.

Costanza:1984:EII

[CN84] Robert Costanza and Christopher Neill. Energy intensities, interdependence, and value in ecological systems: a linear programming approach. Journal of Theoretical Biology, 106(1):41–57, January 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519384900262.

Cohen:1985:WWL

[CN85] Joel E. Cohen and Charles M. Newman. When will a large complex system be stable? Journal of Theoretical Biology, 113(1):153–156, March 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519385800813.

Craig:1985:ESO

[COA85] Catherine L. Craig, Akira Okubo, and Viggo Andreasen. Effect of spider orb-web and insect oscillations on prey interception. Journal of Theoretical Biology, 115(2):201–211, July 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800965.

Cale:1983:AEN

[COG83] W. G. Cale, R. V. O’Neill, and R. H. Gardner. Aggregation error in nonlinear ecological models. Journal of Theoretical Biology, 100(3):539–550, February 7, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383904460.

Cohen:1984:NCC

[Coh84] E. B. Cohen. New concepts of chemical and biological structure: Consequences of consistently treating weak bonds as chemical structural determinants. Journal of Theoretical Biology, 108(3):

369–376, June 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384800405.

Cohen:1985:WDB

[Coh85] Stephen R. Cohen. Why does brain make lactate? Journal of Theoretical Biology, 112(2):429–432, January 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385802952.

Coleman:1982:SDW

[Col82] Annette W. Coleman. Sex is dangerous in a world of potential symbionts or the basis of selection for uniparental inheritance. Journal of Theoretical Biology, 97(3):367–369, August 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251938290371X.

Colombini:1984:NMV

[Col84] Marco Colombini. A novel mechanism for voltage control of channel conductance. Journal of Theoretical Biology, 110(4):559–567, October 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384801459.

Colacicco:1985:AAA

[Col85] Giuseppe Colacicco. Arguments against and alternatives for an extracellular surfactant layer in the alveoli of mammalian lung. Journal of Theoretical Biology, 114(4):641–656, June 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385800497.

Comins:1982:ESS

[Com82] Hugh N. Comins. Evolutionarily stable strategies for localized dispersal in two dimensions. Journal of Theoretical Biology, 94(3):579–606, February 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382903022.

Conway:1983:PMC

[Con83] Francis J. Conway. A proposed model of cancer as the inappropriate expression of non-body introns. Journal of Theoretical Biology, 100(1):1–24, January 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383900905.

Cooper:1981:NDT

[Coo81] W. S. Cooper. Natural decision theory: a general formalism for the analysis of evolved characteristics. Journal of Theoretical Biology, 92(4):401–415, October 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902563.

Cooper:1982:CMS

[Coo82] Stephen Cooper. The continuum model: Statistical implications. Journal of Theoretical Biology, 94(4):783–800, February 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382900789.

Cook:1984:TIN

[Coo84a] Norman D. Cook. The transmission of information in natural systems. Journal of Theoretical Biology, 108(3):349–367, June 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384800399.

Cooper:1984:GJI

[Coo84b] Mark S. Cooper. Gap junctions increase the sensitivity of tissue cells to exogenous electric fields. Journal of Theoretical Biology, 111(1):123–130, November 7, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384802003.

Cooper:1984:ETE

[Coo84c] William S. Cooper. Expected time to extinction and the concept of fundamental fitness. Journal of Theoretical Biology, 107(4):603–629, April 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519384801356.

Cooper:1987:CCC

[Coo87] Stephen Cooper. Cell cycle controls in eukaryotic cells: a reply. Journal of Theoretical Biology, 127(2):247–249, July 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519387801340.

Cooper:1988:CMC

[Coo88] Stephen Cooper. The continuum model and c-myc synthesis during the division cycle. Journal of Theoretical Biology, 135(3):393–400, December 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519388802534.

Corrsin:1982:MEE

[Cor82] S. Corrsin. A more explicit estimate for the “implications of athlete’s bradycardia on lifespan”. Journal of Theoretical Biology, 96(4):683–688, June 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902363.

Collier:1983:MCP

[COR83a] Ivan E. Collier, William H. Olson, and James D. Regan. A mechanism for the control of patch size in mammalian cell DNA excision repair. Journal of Theoretical Biology, 102(4):487–499, June 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383903855.

Corben:1983:WBF

[Cor83b] Herbert C. Corben. Wing-beat frequencies, wing-areas and masses of flying insects and hummingbirds. Journal of Theoretical Biology, 102(4):611–623, June 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383903946.

Cousins:1980:TCD

[Cou80] S. H. Cousins. A trophic continuum derived from plant structure, animal size and a detritus cascade. Journal of Theoretical Biology, 82(4):607–618, February 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380901812.

Chen:1982:MRV

[CP82] I. I. H. Chen and R. L. Prewitt. A mathematical representation for vessel network. Journal of Theoretical Biology, 98(2):211–219, September 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382902594.

Chang:1985:TMI

[CP85] Ho Nam Chang and Tai Hyun Park. A theoretical model for immobilized whole cell enzyme. Journal of Theoretical Biology, 116(1):9–20, September 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801284.

Cervantes-Perez:1985:NMI

[CPLA85] Francisco Cervantes-Pérez, Rolando Lara, and Michael Arbib. A neural model of interactions subserving prey-predator discrimination and size preference in anuran amphibia. Journal of Theoretical Biology, 113(1):117–152, March 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800801.

Cocho:1987:DSCa

[CPPR87] G. Cocho, R. Pérez-Pascual, and J. L. Rius. Discrete systems, cell-cell interactions and color pattern of animals. I. Conflicting dynamics and pattern formation. Journal of Theoretical Biology, 125(4):419–435, April 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387802114.

Cocho:1987:DSCb

[CPPRS87] G. Cocho, R. Pérez-Pascual, J. L. Rius, and F. Soto. Discrete systems, cell-cell interactions and color pattern of animals II. Clonal theory and cellular automata. Journal of Theoretical Biology, 125(4):437–447, April 21, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387802126.

Craig:1980:FCA

[Cra80a] A. G. Craig. The formation of chromosome aberrations from single-strand damage in irradiated DNA. Journal of Theoretical Biology, 82(4):633–642, February 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380901836.

Craig:1980:SRC

[Cra80b] Robin Craig. Sex ratio changes and the evolution of eusociality in the hymenoptera: Simulation and games theory studies. Journal of Theoretical Biology, 87(1):55–70, November 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380902192.

Craig:1981:PMM

[Cra81a] A. G. Craig. A proposed modification to the microdosimetric theory of RBE. Journal of Theoretical Biology, 91(1):47–61, July 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381903738.

Craig:1981:TPR

[Cra81b] Robin Craig. The theoretical possibility of reverse translation of proteins into genes. Journal of Theoretical Biology, 88(4):757–760, February 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519381902502.

Craig:1982:EEK

[Cra82a] Robin Craig. Evolution of eusociality by kin selection: the effect of inbreeding between siblings. Journal of Theoretical Biology, 94(1):119–128, January 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382903344.

Craig:1982:EMW

[Cra82b] Robin Craig. Evolution of male workers in the hymenoptera. Journal of Theoretical Biology, 94(1):95–105, January 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382903319.

Craig:1983:FSE

[Cra83a] Robin Craig. Familial selection and the evolution of social behaviour re-examined. Journal of Theoretical Biology, 103(2):287– 294, July 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383900292.

Craig:1983:SEE

[Cra83b] Robin Craig. Subfertility and the evolution of eusociality by kin selection. Journal of Theoretical Biology, 100(3):379–397, February 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383904368.

Crawford:1989:LMS

[Cra89] Vincent P. Crawford. Learning and mixed-strategy equilibria in evolutionary games. Journal of Theoretical Biology, 140(4):537– 550, October 23, 1989. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519389801134.

Canals:1989:EGH

[CRB89] Mauricio Canals, Mario Rosenmann, and Francisco Bozinovic. Energetics and geometry of huddling in small mammals. Journal of Theoretical Biology, 141(2):181–189, November 21, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389800165.

Cressman:1988:CDB

[Cre88a] R. Cressman. Complex dynamical behaviour of frequencydependent viability selection: An example. Journal of Theoretical Biology, 130(2):167–173, January 21, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388800912.

Cressman:1988:FDV

[Cre88b] R. Cressman. Frequency-dependent viability selection (a single-locus, multi-phenotype model). Journal of Theoretical Biology, 130(2):147–165, January 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388800900.

Crippa:1981:EEP

[Cri81] P. R. Crippa. Electronic effects in the pigmented cuticle of insects. Journal of Theoretical Biology, 92(4):369–371, October 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251938190254X.

Crisci:1982:PET

[Cri82] Jorge V. Crisci. Parsimony in evolutionary theory: Law or methodological prescription? Journal of Theoretical Biology, 97(1):35–41, July 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519382902740.

Crumplin:1980:DTR

[Cru80] G. C. Crumplin. The derivation of theoretical resistance mechanisms to antibacterial agents. Journal of Theoretical Biology, 85(4):597–621, August 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251938090260X.

Chapais:1980:EMF

[CS80] Bernard Chapais and Steven R. Schulman. An evolutionary model of female dominance relations in primates. Journal of Theoretical Biology, 82(1):47–89, January 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380900909.

Clawson:1982:MNT

[CS82] Gary A. Clawson and Edward A. Smuckler. A model for nucleocytoplasmic transport of ribonucleoprotein particles. Journal of Theoretical Biology, 95(4):607–613, April 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382903447.

Clay:1984:AEC

[CS84] John R. Clay and Michael F. Shlesinger. Analysis of the effects of cesium ions on potassium channel currents in biological membranes. Journal of Theoretical Biology, 107(2):189–201, March 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384800211.

Cavalier-Smith:1987:BDS

[CS87] T. Cavalier-Smith. Bacterial DNA segregation: Its motors and positional control. Journal of Theoretical Biology, 127(3):361– 372, August 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387801121.

Crowley:1980:MPB

[CSC^{+}80] Philip H. Crowley, Susan C. Straley, Robert J. Craig, Joseph D. Culin, and … Joseph P. Straley. A model of prey bacteria, predator bacteria, and bacteriophage in continuous culture. Journal of Theoretical Biology, 86(2):377–400, September 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380900132.

Coutre:1981:PTD

[CSH^{+}81] S. E. Coutre, J. M. Stanford, J. G. Hovis, P. W. Stevens, and T. T. Wu. Possible three-dimensional backbone folding around antibody combining site of immunoglobulin MOPC167. Journal of Theoretical Biology, 92(4):417–434, October 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902575.

Caple:1986:BIN

[CSL^{+}86] Gerald Caple, David C. Sands, R. Gary Layton, William V. Zucker, and J. R. Snider. Biogenic ice nucleation: Could it be metabolically initiated? Journal of Theoretical Biology, 119(1): 37–45, March 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519386800492.

Chiang:1981:ADC

[CT81a] Anthony S. T. Chiang and Robert W. Thompson. The age distribution from continuous biochemical reactors with cell reproduction by mitosis. Journal of Theoretical Biology, 89(2):321– 333, March 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381903143.

Cosmides:1981:CII

[CT81b] Leda Murlas Cosmides and John Tooby. Cytoplasmic inheritance and intragenomic conflict. Journal of Theoretical Biology, 89(1):83–129, March 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519381901818.

Cohn:1986:MDS

[CTS86] Stanley A. Cohn, David H. Tippit, and Timothy P. Spurck. Microtubule dynamics in the spindle. II. A thermodynamic and kinetic description. Journal of Theoretical Biology, 122(3):277–301, October 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519386801217.

Cowan:1988:SEA

[CU88] Richard Cowan and P. Anne Underwood. Steric effects in antibody reactions with polyvalent antigen. Journal of Theoretical Biology, 132(3):319–335, June 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388802182.

Cummings:1985:MPF

[Cum85a] F. W. Cummings. A model of pattern and form in development. Journal of Theoretical Biology, 112(4):707–726, February 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385800552.

Cummings:1985:PSI

[Cum85b] F. W. Cummings. A pattern-surface interactive model of morphogenesis. Journal of Theoretical Biology, 116(2):243–273, September 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385802678.

Cummings:1989:SGC

[Cum89] F. W. Cummings. On surface geometry coupled to morphogen. Journal of Theoretical Biology, 137(2):215–219, March 21, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519389802073.

Curnow:1988:UMC

[Cur88] R. N. Curnow. The use of Markov chain models in studying the evolution of the proteins. Journal of Theoretical Biology, 134(1):51–57, September 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388803011.

Cushing:1980:MSI

[Cus80] J. M. Cushing. Model stability and instability in age structured populations. Journal of Theoretical Biology, 86(4):709–730, October 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519380903070.

Conrad:1981:RAA

[CV81] Michael Conrad and M. V. Volkenstein. Replaceability of amino acids and the self-facilitation of evolution. Journal of Theoretical Biology, 92(3):293–299, October 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902939.

Cannings:1988:PEI

[CV88] C. Cannings and G. T. Vickers. Patterns of ESS’s II. Journal of Theoretical Biology, 132(4):409–420, June 21, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938880081X.

Collado-Vides:1989:TGA

[CV89] Julio Collado-Vides. A transformational-grammar approach to the study of the regulation of gene expression. Journal of Theoretical Biology, 136(4):403–425, February 22, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389801560.

Cheer:1987:CAV

[CVNO87] Angela Cheer, Jean-Paul Vincent, Richard Nuccitelli, and George Oster. Cortical activity in vertebrate eggs i: The activation waves. Journal of Theoretical Biology, 124(4):377–404, February 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519387802175.

Caraco:1984:IEC

[CW84] Thomas Caraco and Thomas S. Whittam. Immigrationextinction competition on islands: Associations among three species. Journal of Theoretical Biology, 110(2):241–252, September 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384800569.

Chen:1988:TAP

[CW88] Jia-Lin Chen and Qi Wang. A theoretical analysis of the potential productivity of ryegrass under grazing. Journal of Theoretical Biology, 133(3):371–383, August 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938880328X.

Cummings:1984:DUP

[CY84] A. M. Cummings and J. M. Yochim. Differentiation of the uterus in preparation for gestation: a model for the action of progesterone. Journal of Theoretical Biology, 106(3):353–374, February 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519384900353.

Czerkawski:1987:RCP

[Cze87] J. W. Czerkawski. Reassessment of the contribution of protozoa to the microbial protein supply to the host ruminant animal. Journal of Theoretical Biology, 126(3):335–341, June 7, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387802400.

Calleja:1980:AFS

[CZJY80] G. B. Calleja, M. Zuker, Byron F. Johnson, and B. Y. Yoo. Analyses of fission scars as permanent records of cell division in schizosaccharomyces pombe. Journal of Theoretical Biology, 84(3):523–544, June 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938080018X.

Dall:1983:TMP

[Dal83] D. J. Dall. A theory for the mechanism of polar filament extrusion in the microspora. Journal of Theoretical Biology, 105(4):647–659, December 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383902254.

Damstra:1986:KAE

[Dam86] K. St J. Damstra. A kinetic analysis of evaporative water loss barriers. Journal of Theoretical Biology, 121(4):439–448, August 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386801011.

Danielli:1982:UP

[Dan82] James F. Danielli. Unsolved problems. Journal of Theoretical Biology, 96(1):1, May 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519382901503.

Daniel:1986:IRP

[Dan86] R. M. Daniel. The immune response to proteins from extreme thermophiles. Journal of Theoretical Biology, 120(1):125– 127, May 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519386800224.

Nogueira:1983:TSH

[dANG83] R. de A. Nogueira and E. A. Conde Garcia. A theoretical study on heat production in squid giant axon. Journal of Theoretical Biology, 104(1):43–52, September 7, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383904009.

Davison:1984:SME

[Dav84] John A. Davison. Semi-meiosis as an evolutionary mechanism. Journal of Theoretical Biology, 111(4):725–735, December 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384802647.

Davison:1987:SME

[Dav87] John A. Davison. Semi-meiosis and evolution: a response. Journal of Theoretical Biology, 126(3):379–380, June 7, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938780245X.

Day:1983:CDP

[Day83a] William H. E. Day. Computationally difficult parsimony problems in phylogenetic systematics. Journal of Theoretical Biology, 103(3):429–438, August 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383902965.

Day:1983:PNN

[Day83b] William H. E. Day. Properties of the nearest neighbor interchange metric for trees of small size. Journal of Theoretical Biology, 101(2):275–288, March 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383903417.

Dufton:1988:MDC

[DB88] M. J. Dufton and P. Bladon. A method for detecting centres of natural selection in protein structures: Potential for predicting the location of functional areas. Journal of Theoretical Biology, 134(3):331–339, October 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938880064X.

Deranleau:1980:ESF

[DBB80] David A. Deranleau, Th. Binkert, and Peter Bally. Estimation of the saturation fraction of bound fluorescent ligands from the intensity and anisotropy of the emission. Journal of Theoretical Biology, 86(3):477–485, October 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251938090346X.

deBoer:1986:IBM

[dBH86] Rob J. de Boer and Pauline Hogeweg. Interactions between macrophages and T-lymphocytes: Tumor sneaking through intrinsic to helper T cell dynamics. Journal of Theoretical Biology, 120(3):331–351, June 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386802053.

deBoer:1987:IDB

[dBH87] Rob J. de Boer and Pauline Hogeweg. Immunological discrimination between self and non-self by precursor depletion and memory accumulation. Journal of Theoretical Biology, 124(3):343–369, February 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387801212.

deBoer:1989:INI

[dBH89] Rob J. de Boer and Pauline Hogeweg. Idiotypic networks incorporating T–B cell co-operation. The conditions for percolation. Journal of Theoretical Biology, 139(1):17–38, July 10, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389800554.

DiGiulio:1987:AMC

[DC87a] Massimo Di Giulio and Franco Caldararo. Absorbent Markov chains as a model for the study of the evolution of proteins. Journal of Theoretical Biology, 124(4):485–494, February 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519387802229.

Dickins:1987:GTS

[DC87b] David W. Dickins and Roger A. Clark. Games theory and siblicide in the kittiwake gull, rissa tridactyla. Journal of Theoretical Biology, 125(3):301–305, April 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387800619.

Buning:1984:TAH

[dCB84] Tjard de Cock Buning. A theoretical approach to the heat sensitive pit organs of snakes. Journal of Theoretical Biology, 111(3):509–529, December 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384802374.

Dix:1980:RAC

[DCF80] Douglas Dix, Patricia Cohen, and John Flannery. On the role of aging in cancer incidence. Journal of Theoretical Biology, 83(1):163–173, March 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251938090377X.

delCastillo:1986:SCV

[dCJM86] Francisco del Castillo, Juan Jimenez, and J. R. Medina. Some comments on the variance of heterozygosity in finite populations. Journal of Theoretical Biology, 119(1):103–106, March 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386800546.

Drescher:1982:PRC

[DCR82] K. Drescher, G. Cornelius, and L. Rensing. Phase response curves obtained by perturbing different variables of a 24 hr model oscillator based on translational control. Journal of Theoretical Biology, 94(2):345–353, January 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382903150.

Danziger:1985:SFC

[DD85a] D. J. Danziger and P. M. Dean. The search for functional correspondences in molecular structure between two dissimilar molecules. Journal of Theoretical Biology, 116(2):215–224, September 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385802654.

Doucet:1985:TSA

[DD85b] Paul G. Doucet and Nicolien J. Drost. Theoretical studies on animal orientation: II. Directional displacement in kineses. Journal of Theoretical Biology, 117(3):337–361, December 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938580148X.

Duncan:1989:MCM

[DD89] Richard H. Duncan and G. S. Davies. Mechanistic clues to the mutagenicity of alkylated DNA bases: a theoretical study. Journal of Theoretical Biology, 140(3):345–354, October 9, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389800918.

DeFuria:1985:CSP

[DDA85] Robert R. DeFuria, Mary K. Dygert, and Ghiath M. Alachi. Computer simulation of the P^{31} NMR equations governing the creatine kinase reaction. Journal of Theoretical Biology, 114(1): 75–91, May 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385802563.

Dinh:1983:SBO

[DDBB83] T. Pham Dinh, J. Demongeot, P. Baconnier, and G. Benchetrit. Simulation of a biological oscillator: The respiratory system. Journal of Theoretical Biology, 103(1):113–132, July 7, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383902023.

Dvoaak:1983:ATH

[DDJ83] I. Dvoaak, J. Dainty, and K. Janácek. Analysis of time hierarchy in plant cell volume changes. Journal of Theoretical Biology, 101(1):87–112, March 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383902746.

Douglas:1982:QMC

[DE82] Michael Edward Douglas and John A. Endler. Quantitative matrix comparisons in ecological and evolutionary investigations. Journal of Theoretical Biology, 99(4):777–795, December 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382901977.

De:1987:VME

[De87] Sasadhar De. On the vibrational mode of an eukaryotic cell during the cell cycle. Journal of Theoretical Biology, 125(1):117– 120, March 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387801844.

Deak:1980:MLS

[Dea80] Ilan I. Deak. A model linking segmentation, compartmentalization and regeneration in drosophila development. Journal of Theoretical Biology, 84(3):477–504, June 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519380800154.

Demetrius:1981:MPE

[Dem81] Lloyd Demetrius. The Malthusian parameter and the effective rate of increase. Journal of Theoretical Biology, 92(2):141–161, September 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381903891.

Demetrius:1983:SEM

[Dem83] Lloyd Demetrius. Selection and evolution in macromolecular systems. Journal of Theoretical Biology, 103(4):619–643, August 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383902862.

Derr:1984:EWS

[Der84] Robert F. Derr. The ethanol withdrawal syndrome: a consequence of lack of substrate for a cerebral Krebs-cycle. Journal of Theoretical Biology, 106(3):375–381, February 7, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519384900365.

Deranleau:1987:KMC

[Der87] David A. Deranleau. Kinetic models of cell shape changes and other optically-detected responses to stimulation. Journal of Theoretical Biology, 128(4):487–498, October 21, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387801923.

Da-fu:1987:MCS

[DfP87] Ding Da-fu and I. Prigogine. Morphological and chiral symmetry breaking in reaction–diffusion systems. Journal of Theoretical Biology, 128(2):135–157, September 21, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387801662.

Dhanoa:1989:UDE

[DFS89] M. S. Dhanoa, J. France, and R. C. Siddons. On using a double-exponential model for describing faecal marker concentration curves. Journal of Theoretical Biology, 141(2):247–257, November 21, 1989. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519389800207.

Dolowy:1980:CEC

[DG80] Krzysztof Dolowy and ZdzisLaw Godlewski. Computation of the erythrocyte cell membrane parameters from electrophoretical and biochemical data: Stern-like electrochemical model of the cell membrane. Journal of Theoretical Biology, 84(4):709–723, June 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519380800294.

Decroly:1985:SBM

[DG85] O. Decroly and A. Goldbeter. Selection between multiple periodic regimes in a biochemical system: Complex dynamic behaviour resolved by use of one-dimensional maps. Journal of Theoretical Biology, 113(4):649–671, April 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801855.

Decroly:1987:SCO

[DG87] Olivier Decroly and Albert Goldbeter. From simple to complex oscillatory behaviour: Analysis of bursting in a multiply regulated biochemical system. Journal of Theoretical Biology, 124(2):219– 250, January 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387802643.

Dimitrov:1982:DVS

[DGST82] Dimiter S. Dimitrov, Georgi A. Georgiev, Natalia G. Stoicheva, and Trayko T. Traykov. Dynamics of viscoelastic spherical membranes-the balloon model of the alveolus. Journal of Theoretical Biology, 96(4):517–532, June 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902284.

deGruul:1980:DRM

[dGV80] F. R. de Gruul and J. C. Van Der Leun. A dose-response model for skin cancer induction by chronic u.v. exposure of a human population. Journal of Theoretical Biology, 83(3):487–504, April 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380900557.

DeBoer:1985:TEI

[DH85] R. J. De Boer and P. Hogeweg. Tumor escape from immune elimination: Simplified precursor bound cytotoxicity models. Journal of Theoretical Biology, 113(4):719–736, April 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801892.

DiCera:1988:LRI

[Di 88] Enrico Di Cera. Linkage relations for individual-site binding data. Journal of Theoretical Biology, 135(1):123–124, November 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388801796.

DiCera:1989:LTI

[Di 89] Enrico Di Cera. Linkage thermodynamics of individual site binding phenomena. Journal of Theoretical Biology, 136(4):467–474, February 22, 1989. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519389801584.

Dilman:1986:OMA

[Dil86] V. M. Dilman. Ontogenetic model of ageing and disease formation and mechanisms of natural selection. Journal of Theoretical Biology, 118(1):73–81, January 7, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386800091.

Dix:1983:TDL

[Dix83] Douglas Dix. Toward a definition of life: Semantic thermodynamic considerations. Journal of Theoretical Biology, 102(2):337– 340, May 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383903715.

Davenport:1982:OBSa

[DK82a] Demorest Davenport and Youssef Kohanzadeh. Orchids, bilateral symmetry and insect perception. Journal of Theoretical Biology, 94(2):241–252, January 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382903113.

Davenport:1982:OBSb

[DK82b] Demorest Davenport and Youssef Kohanzadeh. Orchids, bilateral symmetry and insect perception. Journal of Theoretical Biology, 97(3):541–542, August 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382903824.

Daniel:1983:FSM

[DK83] Thomas L. Daniel and Joel G. Kingsolver. Feeding strategy and the mechanics of blood sucking in insects. Journal of Theoretical Biology, 105(4):661–677, December 21, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383902266.

Drzewiecki:1989:PBM

[DKW89] G. M. Drzewiecki, E. Karam, and W. Welkowitz. Physiological basis for mechanical time-variance in the heart: Special consideration of non-linear function. Journal of Theoretical Biology, 139(4):465–486, August 22, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389800669.

deLong:1983:UCM

[dL83] R. de Long. A unified concept for microbial genetics. Journal of Theoretical Biology, 103(1):163–165, July 7, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383902059.

Davenport:1985:IAO

[DL85] Demorest Davenport and Hua Lee. Image analysis in the orchidaceae. Journal of Theoretical Biology, 114(2):199–222, May 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385801041.

Delville:1985:CCP

[DLN85] A. Delville, P. Laszlo, and D. J. Nelson. Calmodulin: Calcium, potassium, and magnesium ion multiple equilibria and kinetics for interconversion, including the effect of repeated stimulation. Journal of Theoretical Biology, 112(1):157–175, January 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385801211.

Day:1980:MCD

[DML80] R. Day, F. Martino, and J. Ladik. Model calculation on DNAprotein interaction. Journal of Theoretical Biology, 84(4):651– 653, June 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519380800257.

deMaertelaer:1985:SCC

[dMP85] V. de Maertelaer and D. Pottier. Stability conditions of cell populations in vivo. Journal of Theoretical Biology, 113(2):299–310, March 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385802290.

Davis:1983:BTH

[DO83] Paul J. Davis and Yoshio Osawa. Binding of thyroid hormone by heme proteins may signal protein age. Journal of Theoretical Biology, 103(1):133–136, July 7, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383902035.

Dobrolyubov:1986:MLS

[Dob86] A. I. Dobrolyubov. The mechanism of locomotion of some terrestrial animals by travelling waves of deformation. Journal of Theoretical Biology, 119(4):457–466, April 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801953.

Dominey:1983:SSA

[Dom83] Wallace J. Dominey. Sexual selection, additive genetic variance and the “phenotypic handicap”. Journal of Theoretical Biology, 101(3):495–502, April 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383901534.

Donhoffer:1986:BSM

[Don86] Sz. Donhoffer. Body size and metabolic rate: Exponent and coefficient of the allometric equation. The role of units. Journal of Theoretical Biology, 119(2):125–137, March 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386800686.

Dounce:1981:OLP

[Dou81] Alexander L. Dounce. Origin of life-proposed mechanisms for primeval polynucleotide and peptide chain synthesis. Journal of Theoretical Biology, 90(1):63–79, May 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381901223.

Dounce:1983:PMC

[Dou83] Alexander L. Dounce. A proposed mechanism for the catalatic ation of catalase. Journal of Theoretical Biology, 105(4):553–567, December 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383902199.

Douglas:1984:HED

[Dou84] J. Douglas. Hypothetical entropy-driven mechanism for selfregulation of the size and division of primitive cells, suggesting the origin and nature of mesosomes. Journal of Theoretical Biology, 109(4):475–478, August 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801526.

Dounce:1985:HEM

[Dou85] Alexander L. Dounce. An hypothesis for explaining the mechanism of translocation in peptide chain synthesis. Journal of Theoretical Biology, 113(1):157–162, March 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800825.

Deneubourg:1983:PBA

[DPV83] J. L. Deneubourg, J. M. Pasteels, and J. C. Verhaeghe. Probabilistic behaviour in ants: a strategy of errors? Journal of Theoretical Biology, 105(2):259–271, ???? 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519383800071.

Dowse:1989:SHP

[DR89] Harold B. Dowse and John M. Ringo. The search for hidden periodicities in biological time series revisited. Journal of Theoretical Biology, 139(4):487–515, August 22, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389800670.

Drake:1981:TDE

[Dra81] V. A. Drake. Target density estimation in radar biology. Journal of Theoretical Biology, 90(4):545–571, June 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381903052.

Dowse:1986:MDK

[DRB86] Harold B. Dowse, John M. Ringo, and Karen M. Barton. A model describing the kinetics of mating in drosophila. Journal of Theoretical Biology, 121(2):173–183, July 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386800911.

Dussert:1988:QDP

[DRL88] C. Dussert, G. Rasigni, and A. Llebaria. Quantization of directional properties in biological structures using the minimal spanning tree. Journal of Theoretical Biology, 135(3):295–302, December 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519388802455.

Dusser:1987:MST

[DRP^{+}87] C. Dusser, M. Rasigni, J. Palmari, G. Rasigni, A. Llebaria, and F. Marty. Minimal spanning tree analysis of biological structures. Journal of Theoretical Biology, 125(3):317–323, April 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519387800632.

Domach:1984:TPM

[DS84] M. M. Domach and M. L. Shuler. Testing of a potential mechanism for E. coli temporal cycle imprecision with a structural model. Journal of Theoretical Biology, 106(4):577–585, February 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519384900080.

Day:1987:CCI

[DS87] William H. E. Day and David Sankoff. Computational complexity of inferring phylogenies from chromosome inversion data. Journal of Theoretical Biology, 124(2):213–218, January 21, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387802631.

Dasgupta:1981:MTI

[DSD81] Anjan Dasgupta, Dillip Sen, and Jyotirmoy Das. On membrane transport and internal entropy production. Journal of Theoretical Biology, 88(3):569–574, February 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902836.

Dexter:1989:SDA

[DSR89] Franklin Dexter, Gerald M. Saidel, and Yoram Rudy. Simulation of the diffusion of acetylcholine in the neuroeffector junctions of the sinus node. Journal of Theoretical Biology, 141(4):505–514, December 19, 1989. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519389802322.

Thomas:1983:ARC

[dSTSM83] D. des S. Thomas, K. Suryanarayana, and E. K. Manavathu. Asexual reproduction coupled with heterothallism; possible consequences for fungi. Journal of Theoretical Biology, 105(2):373– 378, ???? 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519383800162.

Damjanovich:1983:PFE

[DSW83] S. Damjanovich, B. Somogyi, and G. R. Welch. Protein fluctuation and enzyme activity. Journal of Theoretical Biology, 105(1):25–33, November 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383904228.

Duniec:1980:TCS

[DT80] J. T. Duniec and S. W. Thorne. A theory of charge separation, ion, electron and proton transport in photosynthetic membranes based on asymmetry of surface charges. Journal of Theoretical Biology, 85(4):691–711, August 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380902660.

Dufton:1983:SRG

[Duf83] M. J. Dufton. The significance of redundancy in the genetic code. Journal of Theoretical Biology, 102(4):521–526, June 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383903880.

Dufton:1985:GCR

[Duf85] M. J. Dufton. Genetic code redundancy and the evolutionary stability of protein secondary structure. Journal of Theoretical Biology, 116(3):343–348, October 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385802721.

Dufton:1986:GCR

[Duf86] M. J. Dufton. Genetic code redundancy and its differential influence on the evolution of protein interiors versus exteriors. Journal of Theoretical Biology, 122(2):231–236, September 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386800844.

Duggleby:1981:PP

[Dug81] Ronald G. Duggleby. A paucity of palindromes in ϕX174. Journal of Theoretical Biology, 93(1):143–155, November 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381900618.

Duggleby:1982:ANL

[Dug82] Ronald G. Duggleby. Application of non-linear regression to the analysis of histidine transport by s37 ascites tumour cells. Journal of Theoretical Biology, 95(4):823–828, April 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382903563.

Duggleby:1986:PCR

[Dug86] Ronald G. Duggleby. Progress curves of reactions catalyzed by unstable enzymes. A theoretical approach. Journal of Theoretical Biology, 123(1):67–80, November 7, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386802363.

Duggleby:1988:FDM

[Dug88] Ronald G. Duggleby. Fitting the double Michaelis–Menten equation to kinetic data. Journal of Theoretical Biology, 130(1):123– 124, January 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251938880167X.

Dunker:1982:PMF

[Dun82] A. Keith Dunker. A proton motive force transducer and its role in proton pumps, proton engines, tobacco mosaic virus assembly and hemoglobin allosterism. Journal of Theoretical Biology, 97(1):95– 127, July 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382902818.

Durup:1982:RBE

[Dur82] Jean Durup. On the relations between error rates in DNA replication and elementary chemical rate constants. Journal of Theoretical Biology, 94(3):607–632, February 7, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382903034.

Dusenbery:1988:LTS

[Dus88] David B. Dusenbery. Limits of thermal sensation. Journal of Theoretical Biology, 131(3):263–271, April 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388802248.

Dusenbery:1989:ERA

[Dus89a] David B. Dusenbery. Efficiency and the role of adaptation in klinokinesis. Journal of Theoretical Biology, 136(3):281–293, February 8, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938980164X.

Dusenbery:1989:RS

[Dus89b] David B. Dusenbery. Ranging strategies. Journal of Theoretical Biology, 136(3):309–316, February 8, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389801663.

Donath:1983:CDW

[DV83] Edwin Donath and Andreas Voigt. Charge distribution within cell surface coats of single and interacting surfaces — a minimum free electrostatic energy approach. Conclusions for electrophoretic mobility measurements. Journal of Theoretical Biology, 101(4):569–584, April 21, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383900164.

deVienne:1985:BGP

[dVR85] Dominique de Vienne and François Rodolphe. Biochemical and genetic properties of oligomeric structures: a general approach. Journal of Theoretical Biology, 116(4):527–568, October 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385800874.

Day:1981:DSC

[DW81a] J. Day and C. R. Willis. Determination of the single chain density of states for membrane systems. Journal of Theoretical Biology, 88(4):693–718, February 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902459.

deWachter:1981:NRE

[dW81b] Rupert de Wachter. The number of repeats expected in random nucleic acid sequences and found in genes. Journal of Theoretical Biology, 91(1):71–98, July 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381903751.

Day:1982:AMP

[DW82] J. Day and C. R. Willis. An analysis of membrane phase transition theories based on kink and jog defects. Journal of Theoretical Biology, 94(2):367–390, January 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382903174.

Delisi:1983:MFP

[DW83] Charles Delisi and Frederik W. Wiegel. Membrane fluidity and the probability of complement fixation. Journal of Theoretical Biology, 102(2):307–322, May 21, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383903685.

Doucet:1987:TSA

[DW87] Paul G. Doucet and Annita N. Wilschut. Theoretical studies on animal orientation. III. A model for kinesis. Journal of Theoretical Biology, 127(1):111–125, July 7, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387801649.

Drane:1982:MTT

[DY82] C. R. Drane and M. J. Yerbury. Measuring thermal time constants of reptiles under conditions of linearly-varying ambient temperatures. Journal of Theoretical Biology, 95(3):531–542, April 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382900339.

Dimitrov:1985:SMS

[DZJ85] D. S. Dimitrov, D. V. Zhflev, and R. K. Jain. Stability of membrane systems modeled as multilayered viscoelastic films. Journal of Theoretical Biology, 113(2):353–377, March 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385802320.

Elliott:1980:CMT

[EAL80] R. L. Elliott, G. L. Amidon, and E. N. Lightfoot. A convective mass transfer model for determining intestinal wall permeabilities: laminar flow in a circular tube. Journal of Theoretical Biology, 87(4):757–771, ???? 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380901150.

Ellington:1987:FED

[EB87] Andrew D. Ellington and Steven A. Benner. Free energy differences between enzyme bound states. Journal of Theoretical Biology, 127(4):491–506, August 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387801455.

Ebenman:1987:NDB

[Ebe87] Bo Ebenman. Niche differences between age classes and intraspecific competition in age-structured populations. Journal of Theoretical Biology, 124(1):25–33, January 7, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387802497.

Ebenman:1988:CBA

[Ebe88] Bo Ebenman. Competition between age classes and population dynamics. Journal of Theoretical Biology, 131(4):389–400, April 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388800365.

Endrenyi:1981:ODE

[EC81] Laszlo Endrenyi and Fung-Yee Chan. Optimal design of experiments for the estimation of precise hyperbolic kinetic and binding parameters. Journal of Theoretical Biology, 90(2):241–263, May 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251938190045X.

Economos:1981:LLM

[Eco81] Angelos C. Economos. The largest land mammal. Journal of Theoretical Biology, 89(2):211–214, March 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381903076.

Economos:1982:OBS

[Eco82] Angelos C. Economos. On the origin of biological similarity. Journal of Theoretical Biology, 94(1):25–60, January 7, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382903289.

Economos:1983:EGS

[Eco83] A. C. Economos. Elastic and/or geometric similarity in mammalian design? Journal of Theoretical Biology, 103(1):167– 172, July 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383902060.

Economos:1984:SIE

[Eco84] Angelos C. Economos. The surface illusion and the elastic/geometric similarity paradox, encore. Journal of Theoretical Biology, 109(3):463–470, August 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384800946.

Edelstein:1982:PFC

[Ede82] Leah Edelstein. The propagation of fungal colonies: a model for tissue growth. Journal of Theoretical Biology, 98(4):679–701, October 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382901461.

Elliott:1987:AIR

[EG87a] Robert J. Elliott and Julia M. Goodfellow. Ammonium ion representation in Monte Carlo simulations of biomolecular solutions. Journal of Theoretical Biology, 128(1):121–125, September 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519387800358.

Elliott:1987:MCS

[EG87b] Robert J. Elliott and Julia M. Goodfellow. Monte Carlo simulations of nucleotide crystal hydrates and their counter-ions. Journal of Theoretical Biology, 127(4):403–412, August 21, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387801388.

Eigen:1980:HCC

[EGS80] M. Eigen, W. C. Gardiner, and P. Schuster. Hypercycles and compartments: Compartments assists — but do not replace — hypercyclic organization of early genetic information. Journal of Theoretical Biology, 85(3):407–411, August 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251938090315X.

Edelstein:1983:MPS

[EH83] Leah Edelstein and Yitzhak Hadar. A model for pellet size distributions in submerged mycelial cultures. Journal of Theoretical Biology, 105(3):427–452, December 7, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383901856.

Eilbert:1988:ASB

[Eil88] James L. Eilbert. Axon shape as a basis for multinode functional units in a hierarchical neural model. Journal of Theoretical Biology, 130(1):95–122, January 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388801668.

Engelberg-Kulka:1988:SEP

[EKSS88] Hanna Engelberg-Kulka and Rachel Schoulaker-Schwarz. Stop is not the end: Physiological implications of translational readthrough. Journal of Theoretical Biology, 131(4):477–485, April 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519388800420.

Enquist:1983:EFB

[EL83] Magnus Enquist and Olof Leimar. Evolution of fighting behaviour: Decision rules and assessment of relative strength. Journal of Theoretical Biology, 102(3):387–410, June 7, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383903764.

Enquist:1987:EFB

[EL87] Magnus Enquist and Olof Leimar. Evolution of fighting behaviour: The effect of variation in resource value. Journal of Theoretical Biology, 127(2):187–205, July 21, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387801303.

Eldridge:1980:TBC

[Eld80] Charles A. Eldridge. A theory of Brownian coagulation in a bounded planar region. Journal of Theoretical Biology, 87 (2):385–400, November 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519380903665.

Elder:1983:EA

[Eld83] D. Elder. The evolution of altruism. Journal of Theoretical Biology, 105(2):371–372, ???? 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519383800150.

Elder:1984:TEC

[Eld84] D. Elder. Theory of epigenetic coding. Journal of Theoretical Biology, 108(3):327–332, June 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384800363.

Ellner:1986:GDP

[Ell86] Stephen Ellner. Germination dimorphisms and parent-offspring conflict in seed germination. Journal of Theoretical Biology, 123(2):173–185, November 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519386801515.

Eriksson:1981:MPO

[ELM81] Anders Eriksson, Peter Lindner, and Olle Mårtensson. Molecular properties and odour. Effects of substituents on pyridine and pyridine odour. Journal of Theoretical Biology, 90(4):477–486, June 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381903003.

Elsasser:1981:PNB

[Els81] Walter M. Elsasser. Principles of a new biological theory: a summary. Journal of Theoretical Biology, 89(1):131–150, March 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251938190182X.

Elsasser:1982:OSM

[Els82] Walter M. Elsasser. The other side of molecular biology. Journal of Theoretical Biology, 96(1):67–76, May 7, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382901564.

Elsasser:1983:BAS

[Els83] Walter M. Elsasser. Biological application of the statistical concepts used in the second law. Journal of Theoretical Biology, 105(1):103–116, November 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383904277.

Enos:1981:EDC

[EM81] Bruce E. Enos and Donald A. McQuarrie. The effect of discrete charges on the electrical properties of membranes. II. Journal of Theoretical Biology, 93(3):499–522, December 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902162.

Emiliani:1982:EEE

[Emi82] Cesare Emiliani. Extinctive evolution: Extinctive and competitive evolution combine into a unified model of evolution. Journal of Theoretical Biology, 97(1):13–33, July 7, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902739.

Elwood:1986:ACI

[EN86] Robert W. Elwood and Susan J. Neil. Asymmetric contests involving two resources. Journal of Theoretical Biology, 120(2):237– 249, May 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519386801771.

Engstrom:1989:MRB

[Eng89] K. G. Engström. Mechanics of the red blood cell corpuscle: Section modulus, shape and stringent profile-curvature equation. Journal of Theoretical Biology, 138(4):529–544, June 22, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519389800499.

Epifanova:1986:CCC

[EP86] Olga I. Epifanova and V. A. Polunovsky. Cell cycle controls in higher eukaryotic cells: Resting state or a prolonged G_{1} period? Journal of Theoretical Biology, 120(4):467–477, June 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386800406.

Elfassi:1986:SAP

[EPH86] Emile Elfassi, Roberto Patarca, and William A. Haseltine. Similarities among the pre-S regions of hepatitis B viruses: Analogy with retroviral transmembrane proteins. Journal of Theoretical Biology, 121(3):371–374, August 7, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801163.

Epstein:1986:YIR

[Eps86] R. J. Epstein. Is your initiator really necessary? Journal of Theoretical Biology, 122(3):359–374, October 7, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801278.

El-Rifai:1984:ATM

[ER84] M. A. El-Rifai. Active transport membrane concentration amplifiers. Journal of Theoretical Biology, 106(4):559–576, February 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519384900079.

Edmonds:1988:RWE

[ER88] Diane E. Edmonds and Michael R. Rose. Reproductive wastage and the evolution of genetic systems. Journal of Theoretical Biology, 132(2):247–250, May 21, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388801589.

Eremia:1985:TCF

[Ere85] Dan Eremia. Is there a continuity of filaments in the sarcomere? Journal of Theoretical Biology, 112(1):177–181, January 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385801223.

Edelstein:1983:GMM

[ES83] L. Edelstein and L. A. Segel. Growth and metabolism in mycelial fungi. Journal of Theoretical Biology, 104(2):187–210, September 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383904101.

Engen:1985:ESS

[ES85] Steinar Engen and Bernt-Erik Sæther. The evolutionary significance of sexual selection. Journal of Theoretical Biology, 117(2):277–289, November 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519385802228.

Engen:1988:OCL

[ES88] Steiner Engen and Bernt-Erik Sæther. Optimalization of constrained life histories: Some generalizations. Journal of Theoretical Biology, 130(2):229–237, January 21, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388800985.

Eshel:1983:ECS

[Esh83] Ilan Eshel. Evolutionary and continuous stability. Journal of Theoretical Biology, 103(1):99–111, July 7, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383902011.

Eshel:1984:ICC

[Esh84a] Ilan Eshel. Are intragametic conflicts common in nature? Do they represent an important factor in evolution? Journal of Theoretical Biology, 108(1):159–162, May 7, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801769.

Eshel:1984:EIC

[Esh84b] Ilan Eshel. On the evolution of an inner conflict. Journal of Theoretical Biology, 108(1):65–76, May 7, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801691.

Evans:1980:DSC

[Eva80a] Geoffrey T. Evans. Diffusive structure: Counterexamples to any explanation? Journal of Theoretical Biology, 82(2):313–315, January 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519380901058.

Evans:1980:IIS

[Eva80b] Geoffrey T. Evans. Instability induced by shear diffusion. Journal of Theoretical Biology, 87(2):297–306, November 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380903628.

Ewald:1980:EBT

[Ewa80] Paul W. Ewald. Evolutionary biology and the treatment of signs and symptoms of infectious disease. Journal of Theoretical Biology, 86(1):169–176, September 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380900739.

Faith:1985:MID

[Fai85] Daniel P. Faith. A model of immunological distances in systematics. Journal of Theoretical Biology, 114(4):511–526, June 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385800412.

Faichney:1989:DDF

[Fai89] G. J. Faichney. Determination of digesta flow by continuous marker infusion. Journal of Theoretical Biology, 136(4):485–486, February 22, 1989. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519389801602.

Fajszi:1982:OOP

[Faj82] Cs. Fajszi. Origin of optical purity: the role of physical asymmetry. Journal of Theoretical Biology, 99(2):285–292, November 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382900054.

Freer:1982:PCD

[FB82a] J. H. Freer and T. H. Birkbeck. Possible conformation of delta-lysin, a membrane-damaging peptide of staphylococcus aureus. Journal of Theoretical Biology, 94(3):535–540, February 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382902995.

Frohlich:1982:TAV

[FB82b] Cliff Frohlich and Ruth E. Buskirk. Transmission and attenuation of vibration in orb spider webs. Journal of Theoretical Biology, 95(1):13–36, March 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902843.

Fulford:1986:MCT

[FB86] G. R. Fulford and J. R. Blake. Muco-ciliary transport in the lung. Journal of Theoretical Biology, 121(4):381–402, August 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386800984.

Frappell:1989:URC

[FBB89] P. B. Frappell, H. A. Blevin, and R. V. Baudinette. Understanding respirometry chambers: What goes in must come out. Journal of Theoretical Biology, 138(4):479–494, June 22, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389800463.

Fetisova:1985:ASF

[FBF85] Z. G. Fetisova, A. Yu. Borisov, and M. V. Fok. Analysis of structure-function correlations in light-harvesting photosynthetic antenna: Structure optimization parameters. Journal of Theoretical Biology, 112(1):41–75, January 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801168.

Fisher:1987:MAL

[FBP87] D. S. Fisher, J. C. Burns, and K. R. Pond. Modeling ad libitum dry matter intake by ruminants as regulated by distension and chemostatic feedbacks. Journal of Theoretical Biology, 126(4):407–418, June 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387801480.

Fuge:1985:MCM

[FBS85] Harald Fuge, Martin Bastmeyer, and Walter Steffen. A model for chromosome movement based on lateral interaction of spindle microtubules. Journal of Theoretical Biology, 115(3):391–399, August 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385801995.

Fajszi:1981:AAP

[FC81] Cs. Fajszi and J. Czégé. Amplification of asymmetry in polymerization. Journal of Theoretical Biology, 88(3):523–531, February 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381902812.

France:1988:ACM

[FCBK88] J. France, C. C. Calvert, R. L. Baldwin, and K. C. Klasing. On the application of compartmental models to radioactive tracer kinetic studies of in vivo protein turnover in animals. Journal of Theoretical Biology, 133(4):447–471, August 21, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388803345.

Focardi:1985:HSM

[FDC85] Stefano Focardi, Jean Louis Deneubourg, and Guido Chelazzi. How shore morphology and orientation mechanisms can affect the spatial organization of intertidal molluscs. Journal of Theoretical Biology, 112(4):771–782, February 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800606.

France:1989:URF

[FDC^{+}89] J. France, M. S. Dhanoa, S. B. Cammell, M. Gill, D. E. Beever, and J. H. M. Thornley. On the use of response functions in energy balance analysis. Journal of Theoretical Biology, 140(1):83–99, September 11, 1989. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519389800311.

Firestone:1983:ODN

[FDR83] M. P. Firestone, R. De Levie, and S. K. Rangarajan. On onedimensional nucleation and growth of “living” polymers I. Homogeneous nucleation. Journal of Theoretical Biology, 104(4):535– 552, October 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383902448.

France:1988:EPF

[FDS^{+}88] J. France, M. S. Dhanoa, R. C. Siddons, J. H. M. Thornley, and D. P. Poppi. Estimating the production of faeces by ruminants from faecal marker concentration curves. Journal of Theoretical Biology, 135(3):383–391, December 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388802522.

Fey:1985:CFS

[FE85] K. Fey and K. Eichmann. Cluster formation in a symmetrical network: a dynamical system for the description of the suppression among non-immune T lymphocytes and its application to the effects of immunization. Journal of Theoretical Biology, 114(4):615–640, June 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800485.

Fishman:1987:TDDb

[FEC87a] Svetlana Fishman, A. Erez, and G. A. Couvillon. The temperature dependence of dormancy breaking in plants: Computer simulation of processes studied under controlled temperatures. Journal of Theoretical Biology, 126(3):309–321, June 7, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387802370.

Fishman:1987:TDDa

[FEC87b] Svetlana Fishman, A. Erez, and G. A. Couvillon. The temperature dependence of dormancy breaking in plants: Mathematical analysis of a two-step model involving a cooperative transition. Journal of Theoretical Biology, 124(4):473–483, February 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519387802217.

Feder:1980:RSA

[Fed80] Jens Feder. Random sequential adsorption. Journal of Theoretical Biology, 87(2):237–254, November 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380903586.

Fell:1980:TAF

[Fel80] David A. Fell. Theoretical analyses of the functioning of the high- and low-K_{m} cyclic nucleotide phosphodiesterases in the regulation of the concentration of adenosine 3’,5’-cyclic monophosphate in animal cells. Journal of Theoretical Biology, 84(2):361– 385, May 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519380800117.

Felsenstein:1982:HCW

[Fel82] Joseph Felsenstein. How can we infer geography and history from gene frequencies? Journal of Theoretical Biology, 96(1): 9–20, May 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382901527.

Feldman:1983:CLM

[Fel83] Mark Y. Feldman. “cross-linkage” model for tRNA interactions with mRNA and ribosome. Journal of Theoretical Biology, 101(3):479–491, April 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383901510.

Ferrier:1980:EIT

[Fer80] J. M. Ferrier. Extracellular ion transport. Journal of Theoretical Biology, 85(4):739–743, August 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380902684.

Ferrier:1981:TDE

[Fer81a] J. M. Ferrier. Time-dependent extracellular ion transport. Journal of Theoretical Biology, 92(4):363–368, October 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902538.

Ferrier:1981:BCE

[Fer81b] Jack M. Ferrier. On a boundary condition of extracellular ion transport. Journal of Theoretical Biology, 93(2):495–496, November 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381901168.

Ferracin:1982:IUO

[Fer82] Alessandro Ferracin. The isotropy of the universe and the origin of chirality. Journal of Theoretical Biology, 94(2):517–518, January 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382903241.

Ferrier:1983:MRL

[Fer83] Jack Ferrier. A mechanism for the regulation of ligament width based on the resonance frequency of ion concentration waves. Journal of Theoretical Biology, 102(4):477–486, June 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383903843.

Ferrier:1984:OIF

[Fer84] Jack Ferrier. Osmosis and intermolecular force. Journal of Theoretical Biology, 106(4):449–453, February 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251938490002X.

Ferrandino:1985:MMB

[Fer85a] Francis J. Ferrandino. A mathematical model of biological development. Journal of Theoretical Biology, 112(3):609–626, February 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385800278.

Ferrier:1985:TDI

[Fer85b] Jack Ferrier. Time-dependent interfacial charging effects of electrical fields applied to biological systems. Journal of Theoretical Biology, 115(1):129–136, July 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800102.

Ferreira:1987:TSM

[Fer87] Ricardo Ferreira. A two-substrate Michaelis–Menten model for the growth of self-replicating polymers. Journal of Theoretical Biology, 128(3):289–295, October 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387800723.

Fernandez:1988:SDC

[Fer88] Ariel Fernández. Stochastic dynamical constraints in de novo RNA replication. Journal of Theoretical Biology, 134(4):419–430, October 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519388800481.

Forsdyke:1983:RFM

[FF83] D. R. Forsdyke and P. M. Ford. Rouleaux formation as a measure of the phase separating ability of plasma. Journal of Theoretical Biology, 103(3):467–472, August 7, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383902989.

Findley:1982:GCA

[FFM82] A. M. Findley, G. L. Findley, and S. P. McGlynn. Genetic coding: approaches to theory construction. Journal of Theoretical Biology, 97(2):299–318, July 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382901060.

France:1987:SFI

[FGDS87] J. France, M. Gill, M. S. Dhanoa, and R. C. Siddons. On solving the fully-interchanging N-compartment model in steady-state tracer kinetic studies with reference to VFA absorption from the rumen. Journal of Theoretical Biology, 125(2):193–211, March 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519387800413.

Fuhr:1989:DMN

[FH89] G. Fuhr and R. Hagedorn. Dielectric motors — a new hypothesis for the bacterial flagella. Journal of Theoretical Biology, 139(1): 39–58, July 10, 1989. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519389800566.

Findlay:1989:BEB

[FHL89] C. Scott Findlay, Roger I. C. Hansell, and Charles J. Lumsden. Behavioral evolution and biocultural games: Oblique and horizontal cultural transmission. Journal of Theoretical Biology, 137(3):245–269, April 6, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389800724.

Fialkowski:1988:LSS

[Fia88] Konrad R. Fialkowski. Lottery of sympatric speciation — a computer model. Journal of Theoretical Biology, 130(4):379–390, February 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519388802042.

Filipski:1988:WRS

[Fil88] Jan Filipski. Why the rate of silent codon substitutions is variable within a vertebrate’s genome. Journal of Theoretical Biology, 134(2):159–164, September 21, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388801991.

Finerty:1981:HC

[Fin81] Michael Finerty. The homology of collagens. Journal of Theoretical Biology, 93(2):279–301, November 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381901053.

Finn:1982:ESN

[Fin82] John T. Finn. Ecosystem succession, nutrient cycling and outputinput ratios. Journal of Theoretical Biology, 99(3):479–489, December 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382902077.

Fischer:1980:REH

[Fis80] Karen M. Fischer. Rapid evolution and histocompatibility antigen function. Journal of Theoretical Biology, 84(3):513–522, June 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519380800178.

Fischbach:1985:RGO

[Fis85] F. A. Fischbach. Radii of gyration of opaque biological microparticles from light scattering. Journal of Theoretical Biology, 116(3):435–442, October 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385802800.

Ferreira:1984:WTT

[FJ84] R. Ferreira and S. G. Jacchieri. Why are there two kinds of chain in tetrameric hemoglobins? Journal of Theoretical Biology, 108(2):191–201, May 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938480065X.

Ferrer:1985:NLT

[FJ85] F. Ferrer and D. Jou. Non-local thermodynamic effects and efficiency of oxidative phosphorylation. Journal of Theoretical Biology, 115(1):153–160, July 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800126.

Forgacs:1989:WPM

[FJFN89] Gabor Forgacs, Navdeep S. Jaikaria, H. L. Frisch, and Stuart A. Newman. Wetting, percolation and morphogenesis in a model tissue system. Journal of Theoretical Biology, 140(3):417–422, IN1, 423–430, October 9, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389800967.

Fritzsch:1983:ABT

[FK83] Günter Fritzsch and Hermann Koepsell. An analysis of biphasic time courses: the inactivation of (Na + +K + )- ATPase and Ca^{2+}-ATPase by ATP-analogs. Journal of Theoretical Biology, 102(4):469–476, June 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383903831.

Fletcher:1980:RBA

[Fle80] C. R. Fletcher. The relationship between active transport and the exchange diffusion effect. Journal of Theoretical Biology, 82(4):643–661, February 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380901848.

Fletcher:1988:BSQ

[Fle88] N. H. Fletcher. Bird song — a quantitative acoustic model. Journal of Theoretical Biology, 135(4):455–481, December 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388802704.

Fortnagel:1986:SHG

[FLNF86] Peter Fortnagel, Keith A. Lampel, Klaus-Dieter Neitzke, and Ernst Freese. Sequence homologies of glucose-dehydrogenases of bacillus megaterium and bacillus subtilis. Journal of Theoretical Biology, 120(4):489–497, June 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938680042X.

Flores:1984:CVI

[Flo84] Richardo Flores. Is the conformation of viroids involved in their pathogenicity? Journal of Theoretical Biology, 108(4):519–527, June 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384800776.

Frere:1983:LBH

[FLR83] Jean-Marie Frère, Bernard Leyh, and André Renard. Lineweaver– Burk, Hanes, Eadie–Hofstee and Dixon plots in non-steady-state situations. Journal of Theoretical Biology, 101(3):387–400, April 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383901467.

Fox:1983:SOP

[FM83] Sidney W. Fox and Koichiro Matsuno. Self-organization of the protocell was a forward process. Journal of Theoretical Biology, 101(2):321–323, March 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383903442.

Fisher:1986:EDV

[FM86] Ronald E. Fisher and Heather D. Mayor. Evolution of a defective virus from a cellular defense mechanism. Journal of Theoretical Biology, 118(4):395–404, February 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801618.

Fidelman:1988:NTA

[FM88] Mark L. Fidelman and Donald C. Mikulecky. Network thermodynamic analysis and simulation of isotonic solute-coupled volume flow in leaky epithelia: An example of the use of network theory to provide the qualitative aspects of a complex system and its verification by simulation. Journal of Theoretical Biology, 130(1):73– 93, January 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519388801656.

Forterre:1980:MSR

[For80] Patrick Forterre. Model for the supercoiling reaction catalyzed by DNA gyrase. Journal of Theoretical Biology, 82(2):255–269, January 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519380901022.

Forsdyke:1981:ISE

[For81] Donald R. Forsdyke. Are introns in-series error-detecting sequences? Journal of Theoretical Biology, 93(4):861–866, December 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381903441.

Forsdyke:1985:HSP

[For85] D. R. Forsdyke. Heat shock proteins defend against intracellular pathogens: a non-immunological basis for self/non-self discrimination? Journal of Theoretical Biology, 115(3):471–473, August 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385802058.

Forster:1988:CAC

[För88] Malte E. C. Förster. Citric acid cycle as a “one-step” reaction. Journal of Theoretical Biology, 133(1):1–11, July 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388800201.

Foulds:1984:MSS

[Fou84] L. R. Foulds. Maximum savings in the Steiner problem in phylogeny. Journal of Theoretical Biology, 107(3):471–474, April 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384801034.

Fox:1981:PCE

[Fox81] Denis L. Fox. Presumed carbaminoprotein equilibria and free energy exchanges in reversible carbon dioxide narcosis of protoplasm. Journal of Theoretical Biology, 90(3):441–443, June 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381903246.

Fikus:1989:BMC

[FP89] Magdalena Fikus and Piotr Pawlowski. Bioelectrorheological model of the cell. 2. Analysis of creep and its experimental verification. Journal of Theoretical Biology, 137(4):365–373, April 20, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519389800359.

Fields:1981:STC

[FQ81] Jeremy H. A. Fields and James F. Quinn. Some theoretical considerations on cytosolic redox balance during anaerobiosis in marine invertebrates. Journal of Theoretical Biology, 88(1):35–45, January 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381903271.

Frankel:1982:GPS

[Fra82] Joseph Frankel. Global patterning in single cells. Journal of Theoretical Biology, 99(1):119–134, November 7, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382903940.

Frasier:1984:ERB

[Fra84] Charles C. Frasier. An explanation of the relationships between mass, metabolic rate and characteristic skeletal length for birds and mammals. Journal of Theoretical Biology, 109(3):331–371, August 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384800867.

Frank:1986:DPS

[Fra86] Steven A. Frank. Dispersal polymorphisms in subdivided populations. Journal of Theoretical Biology, 122(3):303–309, October 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386801229.

Franciolini:1987:SFM

[Fra87] Fabio Franciolini. Spontaneous firing and myelination of very small axons. Journal of Theoretical Biology, 128(2):127–134, September 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387801650.

Frankham:1988:MHP

[Fra88] R. Frankham. Molecular hypotheses for position-effect variegation: Anti-sense transcription and promoter occlusion. Journal of Theoretical Biology, 135(1):85–107, November 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388801760.

Friedhoff:1983:EIC

[Fri83] Lawrence T. Friedhoff. The effect of ion channels on cell membrane motion. Journal of Theoretical Biology, 101(2):205–210, March 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251938390334X.

Fritzsch:1985:KAE

[Fri85] Günter K. Fritzsch. A kinetic analysis of enzyme inactivation as applied to the covalent modification of Na^{+} + K^{+}-ATPase and Ca^{2+}-ATPase. Journal of Theoretical Biology, 117(3):397–415, December 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251938580151X.

Frommel:1984:ASA

[Frö84] Cornelius Frömmel. The apolar surface area of amino acids and its empirical correlation with hydrophobic free energy. Journal of Theoretical Biology, 111(2):247–260, November 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938480209X.

Frommel:1988:UAM

[Frö88] Cornelius Frömmel. Use of the averaged mutation rate in pieces of protein sequences to predict the location of antigenic determinants. Journal of Theoretical Biology, 132(2):171–177, May 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388801553.

Freter:1980:PSM

[FS80] Rolf R. Freter and Michael A. Savageau. Proofreading systems of multiple stages for improved accuracy of biological discrimination. Journal of Theoretical Biology, 85(1):99–123, July 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380902842.

Frehland:1983:RTM

[FS83] E. Frehland and W. Stephan. Rate theory models for ion transport through rigid pores. I. Time-dependent analysis in the case of vanishing interactions. Journal of Theoretical Biology, 103(1): 77–97, July 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251938390200X.

Fisher:1985:ACI

[FS85] Rod R. Fisher and Richard C. Seagrave. Analysis of chemically induced growth with one-dimensional mass transfer. Journal of Theoretical Biology, 115(2):221–240, July 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800989.

France:1986:DDF

[FS86] J. France and R. C. Siddons. Determination of digesta flow by continuous market infusion. Journal of Theoretical Biology, 121(1):105–119, July 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519386800315.

Fukshansky:1987:SFP

[FS87] L. Fukshansky and A. R. Steinhardt. Spatial factors in phycomyces phototropism: Analysis of balanced responses. Journal of Theoretical Biology, 129(3):301–323, December 7, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387800048.

Frank:1988:SRU

[FS88] Steven A. Frank and Ian R. Swingland. Sex ratio under conditional sex expression. Journal of Theoretical Biology, 135 (3):415–418, December 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002251938880256X.

Fetisova:1989:BEO

[FSF89] Z. G. Fetisova, L. V. Shibaeva, and M. V. Fok. Biological expedience of oligomerization of chlorophyllous pigments in natural photosynthetic systems. Journal of Theoretical Biology, 140(2):167–184, September 22, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389801262.

Fares:1980:PDF

[FSM80] Youhanna Fares, Peter J. H. Sharpe, and Charles E. Magnuson. Pheromone dispersion in forests. Journal of Theoretical Biology, 84(2):335–359, May 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519380800105.

Ferreira:1985:RCB

[FT85] Ricardo Ferreira and Constantino Tsallis. On the role of complementarity in biogenesis: a critical phenomenon approach. Journal of Theoretical Biology, 117(2):303–317, November 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385802241.

Feldman:1987:BDC

[FT87] Marcus W. Feldman and Ewart A. C. Thomas. Behaviordependent contexts for repeated plays of the Prisoner’s dilemma II: Dynamical aspects of the evolution of cooperation. Journal of Theoretical Biology, 128(3):297–315, October 7, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387800735.

France:1985:MDF

[FTDS85] J. France, J. H. M. Thornley, M. S. Dhanoa, and R. C. Siddons. On the mathematics of digesta flow kinetics. Journal of Theoretical Biology, 113(4):743–758, April 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801910.

Fuerst:1980:BSF

[Fue80] J. A. Fuerst. Bacterial sheathed flagella and the rotary motor model for the mechanism of bacterial motility. Journal of Theoretical Biology, 84(4):761–774, June 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519380800324.

Fujita:1982:BLG

[Fuj82] Michiya Fujita. Bioassembly lines: a general mechanism for maintaining the specific architecture of a cell, with an implication for the function of a small RNA. Journal of Theoretical Biology, 99(1):9–13, November 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251938290385X.

Fukshansky:1981:QST

[Fuk81] Leonid Fukshansky. A quantitative study of timing in plant photoperiodism. Journal of Theoretical Biology, 93(1):63–91, November 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381900576.

Fukada:1983:SMD

[Fuk83] Masako Fukada. Superhelical model of DNA for regulation of gene expression. Journal of Theoretical Biology, 100(2):239–254, January 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383903491.

Fukada:1984:ASM

[Fuk84] Masako Fukada. Applicability of superhelical model for regulation of gene expression. Journal of Theoretical Biology, 109(3):431– 452, August 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384800910.

Fulda:1981:LEP

[Ful81] Joseph S. Fulda. The logistic equation and population decline. Journal of Theoretical Biology, 91(2):255–259, July 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902320.

Futcher:1986:CNA

[Fut86] A. Bruce Futcher. Copy number amplification of the 2μ m circle plasmid of Saccharomyces cerevisiae. Journal of Theoretical Biology, 119(2):197–204, March 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386800741.

Faber:1981:UCA

[FW81] J. Job Faber and Lori L. Woods. The use of chemical analyses of umbilical cord blood for the computation of instantaneous fetal growth. Journal of Theoretical Biology, 89(3):433–448, April 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381903611.

Firbank:1985:MIW

[FW85] L. G. Firbank and A. R. Watkinson. A model of interference within plant monocultures. Journal of Theoretical Biology, 116(2):291–311, September 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385802691.

Fedyanin:1981:ESA

[FY81] V. K. Fedyanin and L. V. Yakushevich. The exciton spectrum of the alpha-helical protein molecule model. Journal of Theoretical Biology, 91(1):1–11, July 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381903696.

Gopalsamy:1980:NEP

[GA80] K. Gopalsamy and B. D. Aggarwala. On the non-existence of periodic solutions of the reactive–diffusive Volterra system of equations. Journal of Theoretical Biology, 82(3):537–540, February 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380902556.

Grossman:1980:DAS

[GAD80] Zvi Grossman, Richard Asofsky, and Charles DeLisi. The dynamics of antibody secreting cell production: Regulation of growth and oscillations in the response to t-independent antigens. Journal of Theoretical Biology, 84(1):49–92, May 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938081037X.

Gadag:1981:FPT

[Gad81] V. G. Gadag. First passage time for a supercritical Galton– Watson process restricted to the non-extinction set. Journal of Theoretical Biology, 93(3):585–589, December 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902216.

Gains:1980:DKP

[Gai80] Nigel Gains. The determination of the kinetic parameters of a carrier mediated transport process in the presence of an unstirred water layer. Journal of Theoretical Biology, 87(3):559–568, December 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519380902350.

Ginzburg:1988:ECS

[GAK88] Lev R. Ginzburg, H. Resit Akçakaya, and Junhyong Kim. Evolution of community structure: Competition. Journal of Theoretical Biology, 133(4):513–523, August 21, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388803382.

Gallant:1983:COM

[Gal83] John K. Gallant. The complexity of the overlap method for sequencing biopolymers. Journal of Theoretical Biology, 101(1): 1–17, March 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383902709.

Gallez:1984:CMAa

[Gal84a] Dominique Gallez. Cell membranes after malignant transformation part i: Dynamic stability at low surface tension. Journal of Theoretical Biology, 111(2):323–340, November 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384802131.

Gallez:1984:CMAb

[Gal84b] Dominique Gallez. Cell membranes after malignant transformation part II: Dynamic stability with surface chemical reactions. Journal of Theoretical Biology, 111(2):341–354, November 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384802143.

Garden:1980:MAV

[Gar80] Peter W. Garden. Markov analysis of viral DNA/RNA sequences. Journal of Theoretical Biology, 82(4):679–684, February 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380901861.

Garfinkel:1982:HCW

[Gar82] David Garfinkel. How can we effectively handle the large information content and complexity of biology? Journal of Theoretical Biology, 96(1):3–8, May 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382901515.

Garte:1987:AMO

[Gar87] Seymour J. Garte. Activation of multiple oncogene pathways: a model for experimental carcinogenesis. Journal of Theoretical Biology, 129(2):177–188, November 21, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387800115.

Garfinkel:1989:CMM

[Gar89] David Garfinkel. Constraint matching as a means of designing biochemical experiments in multi-enzyme systems. Journal of Theoretical Biology, 137(2):221–234, March 21, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389802085.

Gates:1982:CSP

[Gat82] D. J. Gates. Competition and skewness in plantations. Journal of Theoretical Biology, 94(4):909–922, February 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382900868.

Gates:1986:SWL

[Gat86] Michael A. Gates. A simple way to look at DNA. Journal of Theoretical Biology, 119(3):319–328, April 7, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801448.

Grossman:1980:TEI

[GB80] Zvi Grossman and Gideon Berke. Tumor escape from immune elimination. Journal of Theoretical Biology, 83(2):267–296, March 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519380902933.

Green:1987:PPC

[GB87] Paul B. Green and D. Ray Baxter. Phyllotactic patterns: Characterization by geometrical activity at the formative region. Journal of Theoretical Biology, 128(3):387–395, October 7, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387800802.

Guell:1988:HFB

[GBFH88] D. C. Guell, H. Brenner, R. B. Frankel, and H. Hartman. Hydrodynamic forces and band formation in swimming magnetotactic bacteria. Journal of Theoretical Biology, 135(4):525–542, December 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519388802741.

Gololobov:1989:AGT

[GBS89] Mikhail Yu. Gololobov, Ivan L. Borisov, and Vytas K. Svedas. Acyl group transfer by proteases forming an acylenzyme intermediate: Kinetic model analysis (including hydrolysis of acylenzyme-nucleophile complex). Journal of Theoretical Biology, 140(2):193–204, September 22, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389801286.

Ghim:1983:DFK

[GC83] Young Sung Ghim and Ho Nam Chang. Diffusional falsification of kinetic constants on Lineweaver–Burk plots. Journal of Theoretical Biology, 105(1):91–102, November 7, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383904265.

Geber:1986:SAH

[GC86] Monica A. Geber and Eric L. Charnov. Sex allocation in hermaphrodites with partial overlap in male/female resource inputs. Journal of Theoretical Biology, 118(1):33–43, January 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386800066.

Gust:1981:SOP

[GD81] Devens Gust and Gary Dirks. Stereochemistry of oligomeric proteins. Journal of Theoretical Biology, 92(1):39–55, September 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381901594.

Gariglio:1987:TMP

[GDHB87] Patricio Gariglio, Rosamaria Del Angel, Alfredo Herrera, and Manuel Bonilla. Theoretical model for the post-transcriptional regulation of the human c-myc gene expression, involving doublestranded RNA processing. Journal of Theoretical Biology, 125(1): 83–92, March 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387801819.

Gladyshev:1982:PTB

[GE82] G. P. Gladyshev and Yu. A. Ershov. Principles of the thermodynamics of biological systems. Journal of Theoretical Biology, 94(2):301–343, January 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382903149.

Gladyshev:1980:TPB

[GEL80] G. P. Gladyshev, Yu. A. Ershov, and V. I. Loshchilov. Thermodynamic principles of the behavior of biological systems. Journal of Theoretical Biology, 83(1):17–42, March 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380903707.

Genoni:1988:SDI

[Gen88] Giulio P. Genoni. Sensory discrimination and the incipient advantage of mutations. Journal of Theoretical Biology, 134(3):327– 329, October 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519388800638.

Gero:1983:DTS

[Ger83a] Alexander Gero. Desensitization, two-state receptors and pharmacological parameters. Journal of Theoretical Biology, 103(1): 137–161, July 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383902047.

Gero:1983:TTE

[Ger83b] Alexander Gero. Theoretical treatment of events between the drug-receptor association and the effect. Journal of Theoretical Biology, 104(2):249–259, September 21, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383904137.

Gero:1985:DCR

[Ger85] Alexander Gero. Desensitization and coupled receptors: a model of drug dependence. Journal of Theoretical Biology, 115(4):603– 617, August 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385801430.

Getz:1981:GBK

[Get81] Wayne M. Getz. Genetically based kin recognition systems. Journal of Theoretical Biology, 92(3):209–226, October 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902885.

Getz:1982:ALK

[Get82] Wayne M. Getz. An analysis of learned kin recognition in hymenoptera. Journal of Theoretical Biology, 99(3):585–597, December 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382902120.

Getz:1984:PDP

[Get84] Wayne M. Getz. Population dynamics: a per capita resource approach. Journal of Theoretical Biology, 108(4):623–643, June 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938480082X.

Gailit:1981:PPD

[GF81a] James O. Gailit and Dixie W. Frederiksen. Possible palindromic DNA. Journal of Theoretical Biology, 88(2):241–243, January 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381900722.

Gotham:1981:SMC

[GF81b] Ivan J. Gotham and Harry L. Frisch. A simple model for cell volume and developmental compartments in nutrient limited cyclostat cultures of algae. Journal of Theoretical Biology, 92(4):435– 467, October 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381902587.

Guderley:1989:PCA

[GFA89] Helga Guderley, Paul Fournier, and Jean-Christian Auclair. Phylogeny congruence analysis and isozyme classification: The pyruvate kinase system. Journal of Theoretical Biology, 140 (2):205–220, September 22, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519389801298.

Gray:1982:PTE

[GG82] B. F. Gray and I. Gonda. Physical theory of enzyme catalysis. Journal of Theoretical Biology, 94(2):513–515, January 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251938290323X.

Gray:1984:CSA

[GG84a] B. F. Gray and I. Gonda. A comment on “Some aspects of the role of quantum mechanics in the theory of muscle contraction”. Journal of Theoretical Biology, 110(1):131–133, September 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384800193.

Groome:1984:THM

[GG84b] Lynn J. Groome and Sharon D. Groome. Tension at the heme: a mathematical treatment of hemoglobin cooperativity. Journal of Theoretical Biology, 109(2):157–172, July 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384800016.

Grossman:1986:PIL

[GGC86] Zvi Grossman, Charles L. Greenblatt, and Irun R. Cohen. Parasite immunology and lymphocyte population dynamics. Journal of Theoretical Biology, 121(2):129–139, July 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386800881.

Glass:1980:UDP

[GGPM80] Leon Glass, Carl Graves, Gino A. Petrillo, and Michael C. Mackey. Unstable dynamics of a periodically driven oscillator in the presence of noise. Journal of Theoretical Biology, 86(3):455– 475, October 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519380903458.

Gunther:1982:QDN

[GH82] Nicholas Gunther and Geoffrey W. Hoffmann. Qualitative dynamics of a network model of regulation of the immune system: a rationale for the IgM to IgG switch. Journal of Theoretical Biology, 94(4):815–855, February 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382900807.

Gates:1983:DC

[GH83] Michael A. Gates and Roger I. C. Hansell. On the distinctness of clusters. Journal of Theoretical Biology, 101(2):263–273, March 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383903405.

Gregorius:1987:CMS

[GH87] H.-R. Gregorius and H. H. Hattemer. Characterization of mating systems. Journal of Theoretical Biology, 128(4):407–422, October 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938780187X.

Ghosh:1981:PSP

[Gho81] Robin Ghosh. On the physiological significance of positive and negative co-operativity. Journal of Theoretical Biology, 93(2):395–401, November 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519381901119.

Ghosh:1982:KAD

[Gho82] Robin Ghosh. Kinetic analysis of dye-linked dehydrogenases and related enzymes where the electron acceptor binds twice during the catalytic sequence. Journal of Theoretical Biology, 94(4):857–867, February 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382900819.

Giersch:1988:CAB

[Gie88] Christoph Giersch. Control analysis of biochemical pathways: a novel procedure for calculating control coefficients, and an additional theorem for branched pathways. Journal of Theoretical Biology, 134(4):451–462, October 21, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388800511.

Ginzburg:1986:TPD

[Gin86] Lev R. Ginzburg. The theory of population dynamics: I. Back to first principles. Journal of Theoretical Biology, 122(4):385–399, October 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519386801801.

Givnish:1986:BCS

[Giv86] Thomas J. Givnish. Biomechanical constraints on self-thinning in plant populations. Journal of Theoretical Biology, 119(2):139– 146, March 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519386800698.

Gattuso:1988:CMQ

[GJ88] Jean-Pierre Gattuso and Jean Jaubert. Computation of metabolic quotients in plant-animal symbiotic units. Journal of Theoretical Biology, 130(2):205–212, January 21, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938880095X.

Gadgil:1983:MPV

[GJG83] Madhav Gadgil, N. V. Joshi, and Sulochana Gadgil. On the moulding of population viscosity by natural selection. Journal of Theoretical Biology, 104(1):21–42, September 7, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383903995.

Grasbeck:1980:POF

[GK80] Ralph Gräsbeck and Ilkka Kouvonen. Pseudo-oligomer formation between related soluble and membrane-bound proteins. Implications on the attachment of proteins to receptors and membranes and the evolution of transport proteins. Journal of Theoretical Biology, 84(3):505–511, June 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519380800166.

Gladyshev:1981:OAE

[GK81] G. P. Gladyshev and M. M. Khasanov. Optical activity and evolution. Journal of Theoretical Biology, 90(2):191–198, May 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381900424.

Giddens:1985:NHR

[GK85] Don P. Giddens and Richard I. Kitney. Neonatal heart rate variability and its relation to respiration. Journal of Theoretical Biology, 113(4):759–780, April 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801922.

Gillis:1986:TPL

[GKB86] Darren M. Gillis, Donald L. Kramer, and Graham Bell. Taylor’s power law as a consequence of Fretwell’s ideal free distribution. Journal of Theoretical Biology, 123(3):281–287, December 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386802430.

Goodwin:1984:FSH

[GL84] B. C. Goodwin and Norbert H. J. Lacroix. A further study of the holoblastic cleavage field. Journal of Theoretical Biology, 109(1): 41–58, July 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384801095.

Glagolev:1980:REL

[Gla80] A. N. Glagolev. Reception of the energy level in bacterial taxis. Journal of Theoretical Biology, 82(2):171–185, January 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380900971.

Gladyshev:1982:CTT

[Gla82] Georgy P. Gladyshev. Classical thermodynamics, tandemism and biological evolution. Journal of Theoretical Biology, 94(2):225– 239, January 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382903101.

Glazier:1987:TPT

[Gla87] Douglas S. Glazier. Toward a predictive theory of speciation: The ecology of isolate selection. Journal of Theoretical Biology, 126(3): 323–333, June 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387802394.

Gleeson:1984:MTS

[Gle84] Scott K. Gleeson. Medawar’s theory of senescence. Journal of Theoretical Biology, 108(3):475–479, June 7, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938480048X.

Glick:1980:MMP

[Gli80] Bernard R. Glick. A molecular model of peptide chain propagation. Journal of Theoretical Biology, 82(1):149–155, January 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380900946.

Gandevia:1982:TRI

[GM82a] S. C. Gandevia and C. K. Mahutte. Theoretical requirements for the interpretation of signals of intramuscular tension. Journal of Theoretical Biology, 97(2):141–153, July 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382900959.

Gunther:1982:TBS

[GM82b] B. Günther and E. Morgado. Theory of biological similarity revisited. Journal of Theoretical Biology, 96(4):543–559, June 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382902302.

Gunther:1984:TSD

[GM84] B. Günther and E. Morgado. Transport similarity: Dimensional analysis of diffusion at cellular level. Journal of Theoretical Biology, 108(3):437–449, June 7, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384800442.

Gunther:1987:BSM

[GM87] B. Günther and E. Morgado. Body size and metabolic rate: the role of units. Journal of Theoretical Biology, 128(3):397–398, October 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387800814.

Gardner:1989:EBA

[GM89] Roy Gardner and Molly R. Morris. The evolution of bluffing in animal contests: an ESS approach. Journal of Theoretical Biology, 137(2):235–243, March 21, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389802097.

Goldwasser:1989:SCC

[GMM89] Lloyd Goldwasser, P. K. Maini, and J. D. Murray. Splitting of cell clusters and bifurcation of bryozoan branches. Journal of Theoretical Biology, 137(3):271–279, April 6, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389800736.

Goltsov:1980:RED

[GMTG80] Victor A. Goltsov, Mikhail A. Mazo, Vyacheslav Z. Tarantul, and Karlen G. Gasaryan. Reassociation of eukaryotic DNA fragments containing interspersed repeats. Journal of Theoretical Biology, 83(3):389–403, April 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380900466.

Green:1986:CPF

[GN86] Richard F. Green and Adeline Taylor Nunez. Centralplace foraging in a patchy environment. Journal of Theoretical Biology, 123(1):35–43, November 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386802338.

Gadgil:1980:ESC

[GNG80] Sulochana Gadgil, Vidyanand Nanjundiah, and Madhav Gadgil. On evolutionarily stable compositions of populations of interacting genotypes. Journal of Theoretical Biology, 84(4):737–759, June 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519380800312.

Goldstein:1983:ACE

[Gol83a] B. N. Goldstein. Analysis of cyclic enzyme reaction schemes by the graph-theoretic method. Journal of Theoretical Biology, 103(2):247–264, July 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519383900279.

Goldsworthy:1983:EPA

[Gol83b] A. Goldsworthy. The evolution of plant action potentials. Journal of Theoretical Biology, 103(4):645–648, August 21, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383902874.

Goldwasser:1988:BPG

[Gol88] Lloyd Goldwasser. Branching patterns, generating rules, and astogenetic trajectories in arborescent bryozoans. Journal of Theoretical Biology, 132(2):179–201, May 21, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388801565.

Goodwin:1982:DE

[Goo82] B. C. Goodwin. Development and evolution. Journal of Theoretical Biology, 97(1):43–55, July 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902752.

Goodfellow:1984:SIN

[Goo84] Julia M. Goodfellow. Solvent interactions in nucleic acid crystal hydrates. Journal of Theoretical Biology, 107(2):261–274, March 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384800272.

Gordon:1987:RRA

[Gor87] Richard Gordon. A retaliatory role for algal projectiles, with implications for the mechanochemistry of diatom gliding motility. Journal of Theoretical Biology, 126(4):419–436, June 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519387801492.

Gotoh:1986:ATB

[Got86] Osamu Gotoh. Alignment of three biological sequences with an efficient traceback procedure. Journal of Theoretical Biology, 121(3):327–337, August 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801126.

Goyns:1982:RPA

[Goy82] M. H. Goyns. The role of polyamines in animal cell physiology. Journal of Theoretical Biology, 97(4):577–589, August 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382903599.

Gallant:1980:TME

[GP80] Jonathan A. Gallant and John Prothero. Testing models of error propagation. Journal of Theoretical Biology, 83(4):561–578, April 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380901897.

Gaubatz:1982:SCC

[GP82a] Jim Gaubatz and Gary V. Paddock. Strategies for constructing complementary DNA for cloning. Journal of Theoretical Biology, 95(4):679–696, April 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382903484.

Glasser:1982:NTN

[GP82b] John W. Glasser and Holly J. Price. Niche theory: New insights from an old paradigm. Journal of Theoretical Biology, 99(3):437–460, December 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519382902053.

Gunasekaran:1982:LND

[GP82c] N. Gunasekaran and L. K. Pande. Lon normal distribution for the intrinsic abundance of species from the Gompertz model. Journal of Theoretical Biology, 98(2):301–305, September 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902661.

Gaylor:1988:SCS

[GPAL88] D. C. Gaylor, K. Prakah-Asante, and R. C. Lee. Significance of cell size and tissue structure in electrical trauma. Journal of Theoretical Biology, 133(2):223–237, July 21, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388800079.

Garcia-Penarrubia:1989:ECS

[GPKBG89a] Pilar Garcia-Peñarrubia, Frederick T. Koster, Arthur D. Bankhurst, and Jesus Galvez. Effect of conjugate size on the kinetics of cell-mediated cytotoxicity at the population level. Journal of Theoretical Biology, 138(1):93–115, May 7, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389801808.

Garcia-Penarrubia:1989:MPD

[GPKBG89b] Pilar Garcia-Peñarrubia, Frederick T. Koster, Arthur D. Bankhurst, and Jesus Galvez. Model for population distributions of lymphocyte-target cell conjugates. Journal of Theoretical Biology, 138(1):77–92, May 7, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389801791.

Giessner-Prettre:1984:TSG

[GPLC84] C. Giessner-Prettre, J. Langlet, and F. Caron. Theoretical study of the geometrical arrangement of GT and GA wobble pairs in two short duplexes. Proton NMR chemical shifts and interaction energy calculations. Journal of Theoretical Biology, 107(2):211– 228, March 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384800235.

Grassi:1989:MDD

[GPM89] Antonio Grassi, Giuseppe C. Pappalardo, and Agata Marletta. Molecular determinants for drug-receptor interactions. 12. Molecular orbital studies on opioid analgesics N-allyl-Normetazocine and nalorphine: Conformations and electrostatic molecular potentials. Journal of Theoretical Biology, 140(4):551–570, October 23, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519389801146.

Granero-Porati:1980:IPE

[GPPZ80] M. I. Granero-Porati, A. Porati, and L. Zani. Informational parameters of an exact DNA base sequence. Journal of Theoretical Biology, 86(2):401–403, September 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380900144.

Grafen:1983:MMG

[GR83] Alan Grafen and Mark Ridley. A model of mate guarding. Journal of Theoretical Biology, 102(4):549–567, June 21, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383903909.

Galen:1988:VPC

[GR88] Candace Galen and John T. Rotenberry. Variance in pollen carryover in animal-pollinated plants: Implications for mate choice. Journal of Theoretical Biology, 135(3):419–429, December 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388802571.

Gray:1980:DHS

[Gra80] B. F. Gray. The distribution of heat sources in the human head-theoretical considerations. Journal of Theoretical Biology, 82(3):473–476, February 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380902507.

Grant:1981:SNH

[Gra81a] Bruce Grant. The safe-neighborhood hypothesis of junk DNA. Journal of Theoretical Biology, 90(1):149–150, May 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381901272.

Gray:1981:SLB

[Gra81b] B. F. Gray. On the “surface law” and basal metabolic rate. Journal of Theoretical Biology, 93(4):757–767, December 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381903386.

Grafstein:1983:SOG

[Gra83a] Daniel Grafstein. Stereochemical origins of the genetic code. Journal of Theoretical Biology, 105(1):157–174, November 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383904319.

Graham:1983:CES

[Gra83b] James Graham. Cancer and evolution: Synthesis. Journal of Theoretical Biology, 101(4):657–659, April 21, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383900218.

Graham:1984:CEA

[Gra84] James Graham. Cancer and evolution: Amplification. Journal of Theoretical Biology, 107(2):341–343, March 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938480034X.

Grafstein:1985:MRO

[Gra85a] Daniel Grafstein. Meso and racemic organisms and the transition to asymmetry. Journal of Theoretical Biology, 115(3):415– 427, August 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385802010.

Grafstein:1985:RME

[Gra85b] Daniel Grafstein. A role for metabolite enantiomers in the evolution and integration of enzyme driven systems. Journal of Theoretical Biology, 114(1):11–20, May 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385802526.

Grafen:1986:SSR

[Gra86a] Alan Grafen. Split sex ratios and the evolutionary origins of eusociality. Journal of Theoretical Biology, 122(1):95–121, September 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386802272.

Graur:1986:EEM

[Gra86b] Dan Graur. The evolution of electrophoretic mobility of proteins. Journal of Theoretical Biology, 118(4):443–469, February 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386801643.

Grafen:1988:CTS

[Gra88] Alan Grafen. A centrosomal theory of the short term evolutionary maintenance of sexual reproduction. Journal of Theoretical Biology, 131(2):163–173, March 21, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388802339.

Gready:1980:EMM

[Gre80] Jill E. Gready. Elaboration of microscopic and macroscopic theories of enzyme action: with implications for enzyme evolution, reversibility and kinetics. Journal of Theoretical Biology, 83(1):175– 191, March 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519380903781.

Green:1982:OFS

[Gre82a] Richard F. Green. Optimal foraging and sex ratio in parasitic wasps. Journal of Theoretical Biology, 95(1):43–48, March 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382902867.

Gregorius:1982:SPP

[Gre82b] Hans-Rolf Gregorius. Selection in plant populations of effectively infinite size II. Protectedness of a biallelic polymorphism. Journal of Theoretical Biology, 96(4):689–705, June 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902375.

Gregorius:1984:APF

[Gre84a] Hans-Rolf Gregorius. Allele protectedness in frequency dependent biallelic selection models with separated generations. Journal of Theoretical Biology, 111(3):425–446, December 7, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384802313.

Gregorius:1984:FFE

[Gre84b] Hans-Rolf Gregorius. Fractional fitnesses in exclusively sexually reproducing populations. Journal of Theoretical Biology, 111(2):205–229, November 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519384802064.

Greene:1985:TTF

[Gre85] Peter R. Greene. Thermal torsion of f-actin. Journal of Theoretical Biology, 117(3):489–492, December 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801569.

Green:1989:EEG

[Gre89] Michael E. Green. Electrorheological effects and gating of membrane channels. Journal of Theoretical Biology, 138(4):413–428, June 22, 1989. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519389800426.

Griffing:1981:TNSa

[Gri81a] Bruce Griffing. A theory of natural selection incorporating interaction among individuals. I. The modeling process. Journal of Theoretical Biology, 89(4):635–658, April 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381900333.

Griffing:1981:TNSb

[Gri81b] Bruce Griffing. A theory of natural selection incorporating interaction among individuals. II. Use of related groups. Journal of Theoretical Biology, 89(4):659–677, April 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381900345.

Griffing:1981:TNSc

[Gri81c] Bruce Griffing. A theory of natural selection incorporating interaction among individuals. III. Use of random groups of inbred individuals. Journal of Theoretical Biology, 89(4):679–690, April 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381900357.

Griffing:1981:TNSd

[Gri81d] Bruce Griffing. A theory of natural selection incorporating interaction among individuals. IV. use of related groups of inbred individuals. Journal of Theoretical Biology, 89(4):691–710, April 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381900369.

Griffing:1982:TNSe

[Gri82a] Bruce Griffing. A theory of natural selection incorporating interaction among individuals. IX. use of groups consisting of a mating pair and sterile diploid caste members. Journal of Theoretical Biology, 95(1):181–198, March 7, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251938290296X.

Griffing:1982:TNSa

[Gri82b] Bruce Griffing. A theory of natural selection incorporating interaction among individuals. V. Use of random synchronized groups. Journal of Theoretical Biology, 94(3):709–728, February 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382903083.

Griffing:1982:TNSb

[Gri82c] Bruce Griffing. A theory of natural selection incorporating interaction among individuals. VI. use of non-random synchronized groups. Journal of Theoretical Biology, 94(3):729–741, February 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382903095.

Griffing:1982:TNSc

[Gri82d] Bruce Griffing. A theory of natural selection incorporating interaction among individuals. VII. Use of groups consisting of one sire and one dam. Journal of Theoretical Biology, 94(4):951–966, February 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382900893.

Griffing:1982:TNSd

[Gri82e] Bruce Griffing. A theory of natural selection incorporating interaction among individuals. VIII. use of groups consisting of a sire and several dams. Journal of Theoretical Biology, 94(4):967–983, February 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251938290090X.

Griffing:1982:TNSf

[Gri82f] Bruce Griffing. A theory of natural selection incorporating interaction among individuate. X. Use of groups consisting of a mating pair together with haploid and diploid caste members. Journal of Theoretical Biology, 95(1):199–223, March 7, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902971.

Griffin:1983:HME

[Gri83] Malcolm Griffin. How much evolutionary advantage does sex confer? Journal of Theoretical Biology, 102(3):447–458, June 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383903806.

Gunning:1981:CEP

[GRJ81] B. E. S. Gunning, R. J. Robins, and B. E. Juniper. Considerations of the estimation of plasmodesmatal frequencies: Emendations to a previous paper by robins and juniper. Journal of Theoretical Biology, 89(4):711–713, April 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381900370.

Groome:1980:TST

[Gro80] Lynn J. Groome. A two-state tension-displacement model for hemoglobin-ligand binding. Journal of Theoretical Biology, 83(3):477–486, April 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519380900545.

Grosovksy:1982:WSA

[Gro82] Betsey Dexter-Dyer Grosovksy. What are the selective advantages of metal deposition and solubilization by microbes? Journal of Theoretical Biology, 97(1):83–86, July 7, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251938290279X.

Grover:1988:SLG

[Gro88] N. B. Grover. Surface-limited growth: a model for the synchronization of a growing bacterial culture through periodic starvation. Journal of Theoretical Biology, 134(1):77–87, September 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388803035.

Goel:1982:CTS

[GRS82] Narendra S. Goel, Bijan Rouyanian, and Mohammad Sanati. On the computation of the tertiary structure of globular proteins. III. Inter-residue distances and computed structures. Journal of Theoretical Biology, 99(4):705–757, December 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382901953.

Ginsborg:1980:DCS

[GS80] B. L. Ginsborg and J. W. Searl. On the determination of the concentration over a sphere with a barrier layer. Journal of Theoretical Biology, 82(3):521–524, February 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380902532.

Gandolfi:1981:ABC

[GS81] Alberto Gandolfi and Roberto Strom. Analysis of binding curves in multivalent antigen-heterogeneous antibody systems. Journal of Theoretical Biology, 92(1):57–84, September 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381901600.

Goldstein:1985:SMS

[GS85] B. N. Goldstein and E. L. Shevelev. Stability of multienzyme systems with feedback regulation: a graph theoretical approach. Journal of Theoretical Biology, 112(3):493–503, February 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385800175.

Grinfeld:1986:IEM

[GS86] Michael Grinfeld and Lee A. Segel. Implementation and extension of MacWilliams model for pre-stalk pre-spore regulation in cellular slime mold. Journal of Theoretical Biology, 121(1):23– 44, July 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519386800261.

Goodwin:1980:FDC

[GT80] B. C. Goodwin and L. E. H. Trainor. A field description of the cleavage process in embryogenesis. Journal of Theoretical Biology, 85(4):757–770, August 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380902702.

Goodwin:1985:TWM

[GT85a] B. C. Goodwin and L. E. H. Trainor. Tip and whorl morphogenesis in acetabularia by calcium-regulated strain fields. Journal of Theoretical Biology, 117(1):79–106, November 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938580165X.

Green:1985:EFW

[GT85b] H. S. Green and T. Triffet. Extracellular fields within the cortex. Journal of Theoretical Biology, 115(1):43–64, July 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800060.

Gilpin:1988:CHR

[GT88a] Michael E. Gilpin and Barbara L. Taylor. Comment on Harpending’s and Roger’s model of intergroup selection. Journal of Theoretical Biology, 135(1):131–135, November 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388801826.

Goel:1988:MFA

[GT88b] Narendra S. Goel and Richard L. Thompson. Movable finite automata (MFA) models for biological systems II: Protein biosynthesis. Journal of Theoretical Biology, 134(1):9–49, September 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938880300X.

Green:1989:ZMC

[GT89] H. S. Green and T. Triffet. A zonal model of cortical functions. Journal of Theoretical Biology, 136(1):87–116, January 9, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519389801924.

Gee:1988:MGI

[GTH88] H. Gee, E. W. Taylor, and R. Hancox. A model for the generation of intra-uterine pressure in the human parturient uterus which demonstrates the critical role of the cervix. Journal of Theoretical Biology, 133(3):281–291, August 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388803229.

Garcia-Tejedor:1987:IHO

[GTMM87] Alvaro García-Tejedor, Federico Morán, and Francisco Montero. Influence of the hypercyclic organization on the error threshold. Journal of Theoretical Biology, 127(4):393–402, August 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519387801376.

Garcia-Tejedor:1988:IHE

[GTSNO^{+}88] Alvaro García-Tejedor, Juan Carlos Sanz-Nuño, José Olarrea, Francisco Javier de la Rubia, and Francisco Montero. Influence of the hypercycle on the error threshold: a stochastic approach. Journal of Theoretical Biology, 134(4):431–443, October 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388800493.

Gulland:1982:WDF

[Gul82]  J. A. Gulland. Why do fish numbers vary?  Journal

of Theoretical Biology, 97(1):69–75, July 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902776.

Gunji:1989:BGR

[Gun89] Yukio Gunji. Branching growth of ribbing patterns and the significance of concealed signals. Journal of Theoretical Biology, 139(2):251–270, July 21, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389801031.

Galvez:1981:TPK

[GV81] J. Galvez and R. Varon. I. Transient phase kinetics of enzyme reactions. Journal of Theoretical Biology, 89(1):1–17, March 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381901752.

Galvez:1981:ITP

[GVC81a] J. Galvez, R. Varon, and F. Garcia Canovas. II. Transient phase of two-substrate enzyme systems. Journal of Theoretical Biology, 89(1):19–35, March 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381901764.

Galvez:1981:IKE

[GVC81b] J. Galvez, R. Varon, and F. G. Carmona. III. Kinetics of enzyme reactions with inactivation steps. Journal of Theoretical Biology, 89(1):37–44, March 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381901776.

Galvez:1982:ITP

[GVCC82] J. Galvez, R. Varon, F. Garcia Canovas, and F. Garcia Carmona. IV. Transient phase of the Uni-Bi mechanisms. Journal of Theoretical Biology, 94(2):413–420, January 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382903198.

Genest:1980:HPC

[GVP^{+}80] M. Genest, F. Vovelle, M. Ptak, B. Maigret, and S. Premilat. Hydration of peptides. I. Calculation of accessible surface areas for several conformations of a cyclic dipeptide. Journal of Theoretical Biology, 87(1):71–84, November 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380902209.

Glasser:1983:RCB

[GW83] John W. Glasser and Richard G. Wiegert. On the role of chance in biology: The stochastic-deterministic continuum. Journal of Theoretical Biology, 102(3):463–467, June 7, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251938390382X.

Gates:1988:SPM

[GW88] D. J. Gates and M. Westcott. Stability of plant mixtures in the absence of infection or predation. Journal of Theoretical Biology, 131(1):15–31, March 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388801176.

Grover:1987:ESE

[GWK87] N. B. Grover, C. L. Woldringh, and L. J. Koppes. Elongation and surface extension of individual cells of Escherichia coli B/r: Comparision of theoretical and experimental size distributions. Journal of Theoretical Biology, 129(3):337–348, December 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519387800061.

Gregorius:1986:SUM

[GZ86] H.-R. Gregorius and Martin Ziehe. The significance of over- and underdominance for the maintenance of genetic polymorphisms II. Overdominance and instability with random mating. Journal of Theoretical Biology, 118(1):115–125, January 7, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386800121.

Grover:1980:DRR

[GZWR80] N. B. Grover, A. Zaritsky, C. L. Woldringh, and R. F. Rosenberger. Dimensional rearrangement of rod-shaped bacteria following nutritional shift-up. I. Theory. Journal of Theoretical Biology, 86(3):421–439, October 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380903434.

Hausfater:1982:IAM

[HAC82] Glenn Hausfater, Susanne Aref, and Sara J. Cairns. Infanticide as an alternative male reproductive strategy in langurs: a mathematical model. Journal of Theoretical Biology, 94(2):391–412, January 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382903186.

Hanauske-Abel:1982:SCC

[HAG82] H. M. Hanauske-Abel and V. Günzler. A stereochemical concept for the catalytic mechanism of prolylhydroxylase: Applicability to classification and design of inhibitors. Journal of Theoretical Biology, 94(2):421–455, January 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382903204.

Haig:1986:CAM

[Hai86] David Haig. Conflicts among megaspores. Journal of Theoretical Biology, 123(4):471–480, December 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386802144.

Hallam:1980:ECC

[Hal80a] Thomas G. Hallam. Effects of cooperation on competitive systems. Journal of Theoretical Biology, 82(3):415–423, February 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380902477.

Halsey:1980:WIM

[Hal80b] Yadviga Dowmont Halsey. Water and its ions:mass action control of reaction rates in lipid-protein membranes. Journal of Theoretical Biology, 84(2):185–188, May 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519380800026.

Hamada:1982:SSC

[Ham82] T. Hamada. Stationary scar-class structure of populations of schizosaccharomyces pombe. Journal of Theoretical Biology, 99(4):835–838, December 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519382902028.

Hammond:1985:OOD

[Ham85] B. J. Hammond. On the origin of the observed distribution of toponymic substance (TOP) in avian retina. Journal of Theoretical Biology, 117(3):453–469, December 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801545.

Hannon:1985:LDE

[Han85] Bruce Hannon. Linear dynamic ecosystems. Journal of Theoretical Biology, 116(1):89–110, September 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801326.

Hannon:1986:ECT

[Han86] Bruce Hannon. Ecosystem control theory. Journal of Theoretical Biology, 121(4):417–437, August 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938680100X.

Hao:1982:TKE

[Hao82] Bai-Lin Hao. Two kinds of entrainment-beating transitions in a driven limit-cycle oscillator. Journal of Theoretical Biology, 98(1):9–14, September 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382900534.

Hardt:1980:MCM

[Har80a] Shoshana L. Hardt. A measurable consequence of the Mitchell theory of phosphorylation. Journal of Theoretical Biology, 87(1):1–7, November 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519380902155.

Hartung:1980:POC

[Har80b] John Hartung. Parent-offspring conflict — a retraction. Journal of Theoretical Biology, 87(4):815–817, ???? 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380901198.

Harley:1981:LES

[Har81] Calvin B. Harley. Learning the evolutionarily stable strategy. Journal of Theoretical Biology, 89(4):611–633, April 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381900321.

Harts:1982:MLW

[Har82a] John Harts. Modeling lake-water mineralization processes. Journal of Theoretical Biology, 99(3):553–569, December 7, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902107.

Hartung:1982:CPH

[Har82b] John Hartung. Carcinogenesis and post hoc conclusions from cybrids. Journal of Theoretical Biology, 97(4):715–716, August 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382903691.

Hartung:1982:MCF

[Har82c] John Hartung. Might cancer be a failed response to renegade mitochondria? Journal of Theoretical Biology, 94(1):173–178, January 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382903381.

Harvey:1982:RA

[Har82d] Paul H. Harvey. On rethinking allometry. Journal of Theoretical Biology, 95(1):37–41, March 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902855.

Harley:1983:WDA

[Har83] Calvin B. Harley. When do animals learn the evolutionarily stable strategy? Journal of Theoretical Biology, 105(1):179–181, November 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383904332.

Hara:1984:DSS

[Har84a] Toshihiko Hara. Dynamics of stand structure in plant monocultures. Journal of Theoretical Biology, 110(2):223–239, September 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384800557.

Hara:1984:SMM

[Har84b] Toshihiko Hara. A stochastic model and the moment dynamics of the growth and size distribution in plant populations. Journal of Theoretical Biology, 109(2):173–190, July 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384800028.

Hartung:1985:ECD

[Har85] John Hartung. Eusocial coprophagous diploids-who is reproducing? Journal of Theoretical Biology, 114(1):9–10, May 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385802514.

Harley:1987:LRO

[Har87a] Calvin B. Harley. Learning rules, optimal behaviour, and evolutionary stability. Journal of Theoretical Biology, 127(3):377– 379, August 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387801145.

Harrison:1987:WSR

[Har87b] Lionel G. Harrison. What is the status of reaction–diffusion theory thirty-four years after Turing? Journal of Theoretical Biology, 125(4):369–384, April 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387802084.

Hastings:1981:SSM

[Has81] Alan Hastings. Simultaneous stability of D = 0 and D ≠ 0 for multiplicative viabilities at two loci: An analytical study. Journal of Theoretical Biology, 89(1):69–81, March 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381901806.

Hastings:1982:MWS

[Has82] Harold M. Hastings. The may-wigner stability theorem. Journal of Theoretical Biology, 97(2):155–166, July 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382900960.

Hastings:1986:IQ

[Has86] Alan Hastings. The invasion question. Journal of Theoretical Biology, 121(2):211–220, July 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386800947.

Hastings:1988:DEH

[Has88] Alan Hastings. Dependence of expected heterozygosity on locus number with stabilizing selection and drift. Journal of Theoretical Biology, 134(1):103–112, September 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388803059.

Haukioja:1982:IRS

[Hau82] Erkki Haukioja. Are individuals really subordinated to genes? A theory of living entities. Journal of Theoretical Biology, 99(2):357–375, November 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382900108.

Havsteen:1989:NPE

[Hav89] Bent H. Havsteen. A new principle of enzyme catalysis: Coupled vibrations facilitate conformational changes. Journal of Theoretical Biology, 140(1):101–127, September 11, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389800323.

Hayashi:1980:FDS

[Hay80a] Yoshio Hayashi. Fluid-dynamical study of protoplasmic streaming in a plant cell. Journal of Theoretical Biology, 85(3):451– 467, August 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519380903197.

Hayashi:1980:TSM

[Hay80b] Yoshio Hayashi. Theoretical study of motive force of protoplasmic streaming in a plant cell. Journal of Theoretical Biology, 85(3):469–480, August 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380903203.

Hines:1984:CWS

[HB84] W. G. S. Hines and D. T. Bishop. Can and will a sexual diploid population attain an evolutionary stable strategy? Journal of Theoretical Biology, 111(4):667–686, December 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384802611.

Hardy:1986:CCS

[HB86a] C. D. Hardy and J. S. Beck. Coagulation in cell suspensions: Extensions of the von Smoluchowski model. Journal of Theoretical Biology, 120(2):181–189, May 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801722.

Hickey:1986:IRR

[HB86b] Donal A. Hickey and Bernhard Benkel. Introns as relict retrotransposons: Implications for the evolutionary origin of eukaryotic mRNA splicing mechanisms. Journal of Theoretical Biology, 121(3):283–291, August 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801084.

Harper:1988:EPP

[HB88] David G. Harper and Robert W. Blake. Energetics of piscivorous predator–prey interactions. Journal of Theoretical Biology, 134(1):59–76, September 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388803023.

Hunter:1989:MMS

[HB89] Graeme K. Hunter and Shawn M. Bader. A mathematical modelling study of epiphyseal cartilage calcification. Journal of Theoretical Biology, 138(2):195–211, May 22, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389801390.

Hickey:1989:GME

[HBA89] Donal A. Hickey, Bernhard F. Benkel, and Sumaia M. Abukashawa. A general model for the evolution of nuclear premRNA introns. Journal of Theoretical Biology, 137(1):41–53, March 7, 1989. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519389801481.

Horovitz:1983:MKC

[HBH83] Amnon Horovitz and Tamir Ben-Hur. A model for the kinetics of crime. Journal of Theoretical Biology, 103(4):609–617, August 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383902850.

Higashi:1989:FNU

[HBP89] Masahiko Higashi, Thomas P. Burns, and Bernard C. Patten. Food network unfolding: An extension of trophic dynamics for application to natural ecosystems. Journal of Theoretical Biology, 140(2):243–261, September 22, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389801328.

Hafner:1981:PEH

[HBRN81] D. Hafner, U. Borchard, O. Richter, and M. Neugebauer. Parameter estimation in Hodgkin–Huxley-type equations for membrane action potentials in nerve and heart muscle. Journal of Theoretical Biology, 91(2):321–345, July 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902368.

Hagan:1981:DIM

[HC81a] P. S. Hagan and M. S. Cohen. Diffusion-induced morphogenesis in the development of dictyostelium. Journal of Theoretical Biology, 93(4):881–908, December 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381903465.

Hallam:1981:NAL

[HC81b] T. G. Hallam and C. E. Clark. Non-autonomous logistic equations as models of populations in a deteriorating environment. Journal of Theoretical Biology, 93(2):303–311, November 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381901065.

Hervagault:1987:BIT

[HC87] Jean-François Hervagault and Stéphane Canu. Bistability and irreversible transitions in a simple substrate cycle. Journal of Theoretical Biology, 127(4):439–449, August 21, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387801418.

Hervagault:1989:DBO

[HC89] Jean-François Hervagault and Alain Cimino. Dynamic behaviors of an open substrate cycle: a graphical approach. Journal of Theoretical Biology, 140(3):399–416, October 9, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389800955.

Havel:1983:CDGb

[HCKB83] Timothy F. Havel, Gordon M. Crippen, Irwin D. Kuntz, and Jeffrey M. Blaney. The combinatorial distance geometry method for the calculation of molecular conformation II. Sample problems and computational statistics. Journal of Theoretical Biology, 104(3):383–400, October 7, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383901133.

Hejblum:1988:CCM

[HCVM88] Gilles Hejblum, Dominique Costagliola, Alain-Jacques Valleron, and Jean-Yves Mary. Cell cycle models and mother-daughter correlation. Journal of Theoretical Biology, 131(2):255–262, March 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938880242X.

Hayden:1980:MRI

[HCW80] Thomas L. Hayden, Philip H. Crowley, and Emery Wilson. A model relating the incidence of meiotic trisomy to maternal age. Journal of Theoretical Biology, 86(1):123–136, September 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380900703.

Hsu:1989:KSG

[HCW89] J. P. Hsu, T. H. Chen, and H. H. Wang. A kinetic study of the growth of a rhizopus colony. Journal of Theoretical Biology, 140(4):445–451, October 23, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389801080.

Hunt:1984:MPR

[HD84] Roderick Hunt and C. J. Doyle. Modelling the partitioning of research effort in ecology. Journal of Theoretical Biology, 111(3):451–461, December 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384802337.

Hallam:1984:ETP

[HdL84] T. G. Hallam and J. T. de Luna. Effects of toxicants on populations: a qualitative: Approach III. Environmental and food chain pathways. Journal of Theoretical Biology, 109(3):411–429, August 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384800909.

Harris:1980:TAS

[HE80a] William F. Harris and Ralph O. Erickson. Tubular arrays of spheres: Geometry, continuous and discontinuous contraction, and the role of moving dislocations. Journal of Theoretical Biology, 83(2):215–246, March 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380902908.

Honda:1980:HMD

[HE80b] Hisao Honda and Goro Eguchi. How much does the cell boundary contract in a monolayered cell sheet? Journal of Theoretical Biology, 84(3):575–588, June 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938080021X.

Hinegardner:1983:BC

[HE83] Ralph Hinegardner and Joseph Engelberg. Biological complexity. Journal of Theoretical Biology, 104(1):7–20, September 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383903983.

Haefner:1984:CIC

[HE84] James W. Haefner and Judith L. Edson. Community invasion by complex life cycles. Journal of Theoretical Biology, 108(3):377– 404, June 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384800417.

Hedges:1985:IMP

[Hed85] R. W. Hedges. Inheritance of magnetosome polarity in magnetotropic bacteria. Journal of Theoretical Biology, 112(3):607–608, February 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385800266.

Hejnowicz:1982:VSF

[Hej82] Z. Hejnowicz. Vector and scalar fields in modeling of spatial variations of growth rates within plant organs. Journal of Theoretical Biology, 96(2):161–173, May 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902193.

Heller:1980:FPH

[Hel80] Rolf Heller. Foraging on potentially harmful prey. Journal of Theoretical Biology, 85(4):807–813, August 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380902738.

Herbers:1981:RTF

[Her81a] Joan M. Herbers. Reliability theory and foraging by ants. Journal of Theoretical Biology, 89(1):175–189, March 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381901843.

Herendeen:1981:EIE

[Her81b] Robert A. Herendeen. Energy intensities in ecological and economic systems. Journal of Theoretical Biology, 91(4):607–620, August 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381902137.

Herbert:1983:IAR

[Her83a] Thomas J. Herbert. The influence of axial rotation upon interception of solar radiation by plant leaves. Journal of Theoretical Biology, 105(4):603–618, December 21, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383902229.

Herbert:1983:RBI

[Her83b] Thomas J. Herbert. On the relationship between interception of direct solar radiation by cactus bodies and plant leaves. Journal of Theoretical Biology, 105(4):619–629, December 21, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383902230.

Heideger:1985:TMC

[HF85] William J. Heideger and Mark E. Ferguson. A theoretical model for calcium absorption from the intestinal lumen. Journal of Theoretical Biology, 114(4):657–664, June 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800503.

Homble:1988:ADT

[HF88] Fabrice Homble and Jack M. Ferrier. Analysis of the diffusion theory of negative capacitance: the role of K^{+} and the unstirred layer thickness. Journal of Theoretical Biology, 131(2):183–197, March 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388802364.

Hernandez:1989:KME

[HFL89] Julio Hernandez, Jorge Fischbarg, and Larry S. Liebovitch. Kinetic model of the effects of electrogenic enzymes on the membrane potential. Journal of Theoretical Biology, 137(1):113–125, March 7, 1989. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519389801535.

Hosoya:1983:MOS

[HFO83] Toichiro Hosoya, Takashi Fujii, and Satoshi Ogawa. A molecular orbital study on the oxidation of hydrogen donor molecules by peroxidase compound II. Journal of Theoretical Biology, 100(2):283–292, January 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383903521.

Harguindey:1980:ECP

[HG80] S. Harguindey and M. Gillis. Evolution and cancer: Possible relationships to changes in environmental hydrogen ion concentration. Journal of Theoretical Biology, 86(3):487–492, October 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380903471.

Hopwood:1987:BIT

[HG87] P. R. Hopwood and D. A. Griffiths. Biochemical investigation of tissue growth: Towards definition of “standard” tissue samples. Journal of Theoretical Biology, 128(4):513–521, October 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519387801947.

Heinrich:1982:EPP

[HGG82] Reinhart Heinrich, Matthias Gaestel, and Roland Glaser. The electric potential profile across the erythrocyte membrane. Journal of Theoretical Biology, 96(2):211–231, May 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902223.

Hogeweg:1981:TPO

[HH81] P. Hogeweg and B. Hesper. Two predators and one prey in a patchy environment: An application of MICMAC modelling. Journal of Theoretical Biology, 93(2):411–432, November 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381901132.

Harrison:1985:QCA

[HH85a] Lionel G. Harrison and Naomi A. Hillier. Quantitative control of acetabularia morphogenesis by extracellular calcium: a test of kinetic theory. Journal of Theoretical Biology, 114(1):177– 192, May 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385802617.

Hogeweg:1985:SPM

[HH85b] P. Hogeweg and B. Hesper. Socioinformatic processes: MIR-ROR modelling methodology. Journal of Theoretical Biology, 113(2):311–330, March 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519385802307.

Hallam:1989:DTI

[HH89] Thomas G. Hallam and Yongnian Huang. On the dynamics of a toxicant-individual system. Journal of Theoretical Biology, 141(1):65–72, November 8, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389800086.

Henderson:1989:IVC

[HHP89] I. M. Henderson, Michael D. Hendy, and David Penny. Influenza viruses, comets and the science of evolutionary trees. Journal of Theoretical Biology, 140(3):289–303, October 9, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389800876.

Hickson:1989:SSI

[Hic89] R. E. Hickson. Self-splicing introns as a source for transposable genetic elements. Journal of Theoretical Biology, 141(1):1–10, November 8, 1989. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519389800037.

Higazi:1985:EEB

[Hig85] A. Higazi. The exchange of energy between the medium and the active site. Journal of Theoretical Biology, 117(4):609–619, December 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385802423.

Hines:1980:ESS

[Hin80] W. G. S. Hines. An evolutionarily stable strategy model for randomly mating diploid populations. Journal of Theoretical Biology, 87(2):379–384, November 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380903653.

Hines:1986:CMS

[Hin86] W. G. S. Hines. Constancy of mean strategy versus constancy of gene frequencies in a sexual diploid population. Journal of Theoretical Biology, 121(3):367–369, August 7, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801151.

Hines:1987:CWS

[Hin87] W. G. S. Hines. Can and will a sexual diploid population evolve to an ESS: The multi-locus linkage equilibrium case. Journal of Theoretical Biology, 126(1):1–5, May 7, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387800966.

Hamada:1985:SDD

[HINT85] T. Hamada, S. Iguti, N. Nishimura, and K. Tainaka. Stage dependent density effect on yeast populations. Journal of Theoretical Biology, 113(4):737–742, April 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801909.

Hirsch:1980:ESI

[Hir80] Henry R. Hirsch. Evolution of senescence: Influence of agedependent death rates on the natural increase of a hypothetical population. Journal of Theoretical Biology, 86(1):149–168, September 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519380900727.

Hirsch:1982:ESN

[Hir82] Henry R. Hirsch. Evolution of senescence: Natural increase of populations displaying Gompertz- or power-law death rates and constant or age-dependent maternity rates. Journal of Theoretical Biology, 98(2):321–346, September 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902685.

Hirsch:1983:IER

[Hir83] Henry R. Hirsch. Influence of the existence of a resting state on the probability of cell division in culture. Journal of Theoretical Biology, 100(3):399–410, February 7, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251938390437X.

Hirsch:1984:IER

[Hir84a] Henry R. Hirsch. Influence of the existence of a resting state on the decay of synchronization in cell culture. Journal of Theoretical Biology, 111(1):61–79, November 7, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801964.

Hirsh:1984:PPM

[Hir84b] Allen G. Hirsh. On the possibility of the production of many rare proteins by higher eukaryotes. Journal of Theoretical Biology, 110(3):399–410, October 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801824.

Hjelm:1986:NGI

[Hje86] K. K. Hjelm. Is non-genic inheritance involved in carcinogenesis? A cytotactic model of transformation. Journal of Theoretical Biology, 119(1):89–101, March 7, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386800534.

Hoekstra:1984:EGMa

[HJS84] Rolf F. Hoekstra, Robert F. Janz, and A. J. Schilstra. Evolution of gamete motility differences I. Relation between swimming speed and pheromonal attraction. Journal of Theoretical Biology, 107(1):57–70, March 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801216.

Hersh:1980:GND

[HK80] R. T. Hersh and P. A. Kitos. Is G_{1} normally distributed? Journal of Theoretical Biology, 86(1):117–122, September 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380900697.

Holliday:1981:PSM

[HK81] R. Holliday and T. B. L. Kirkwood. Predictions of the somatic mutation and mortalization theories of cellular ageing are contrary to experimental observations. Journal of Theoretical Biology, 93(3):627–642, December 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902253.

Holliday:1983:TCA

[HK83] R. Holliday and T. B. L. Kirkwood. Theories of cell ageing: a case of mistaken identity. Journal of Theoretical Biology, 103(2):329– 330, July 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383900322.

Harrison:1988:CBR

[HK88] Lionel G. Harrison and Miroslav Kolár. Coupling between reaction–diffusion prepattern and expressed morphogenesis, applied to desmids and dasyclads. Journal of Theoretical Biology, 130(4):493–515, February 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388802133.

Havel:1983:CDGa

[HKC83] Timothy F. Havel, Irwin D. Kuntz, and Gordon M. Crippen. The combinatorial distance geometry method for the calculation of molecular conformation. I. A new approach to an old problem. Journal of Theoretical Biology, 104(3):359–381, October 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383901121.

Hamada:1982:SSS

[HKK82] T. Hamada, S. Kanno, and E. Kano. Stationary stage structure of yeast population with stage dependent generation time. Journal of Theoretical Biology, 97(3):393–414, August 7, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382903733.

Hill:1982:SWT

[HL82] James M. Hill and Alexander I. Lee. Solutions for water transport through plant tissue. Journal of Theoretical Biology, 95(2):409– 411, March 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382902557.

Hughes:1984:FSW

[HL84] A. J. Hughes and D. M. Lambert. Functionalism, structuralism, and “ways of seeing”. Journal of Theoretical Biology, 111(4):787–800, December 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519384802672.

Hippe:1986:SPM

[HL86] Sigrun Hippe and H. Lüth. A simple physical model for fungicide induced hexagonal clustering of intramembrane particles in the plasmalemma of ustilago avenae. Journal of Theoretical Biology, 121(3):351–366, August 7, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938680114X.

Horovitz:1989:KBP

[HL89] Amnon Horovitz and R. D. Levine. Knowledge-based predictions of binding constants for engineered enzymes. Journal of Theoretical Biology, 141(4):557–562, December 19, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938980236X.

Hilbert:1985:UHT

[HLS85] D. W. Hilbert, J. A. Logan, and D. M. Swift. A unifying hypothesis of temperature effects on egg development and diapause of the migratory grasshopper, melanoplus sanguinipes (orthoptera: Acrididae). Journal of Theoretical Biology, 112(4):827–838, February 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385800643.

Hilbert:1989:CME

[HLS89] David W. Hilbert, Jesse A. Logan, and David M. Swift. Comments on a model of egg development and diapause: a reply to Susan Scott. Journal of Theoretical Biology, 136(1):121–123, January 9, 1989. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519389801948.

Horiike:1980:ELC

[HM80] Kihachiro Horiike and Donald B. Mccormick. Effect of ligand on chemical modification of proteins. Graphical determinations of dissociation constant and number of essential residues affected by ligand binding. Journal of Theoretical Biology, 84(4):691– 708, June 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519380800282.

Houston:1985:VBC

[HM85a] A. I. Houston and J. M. McNamara. The variability of behaviour and constrained optimization. Journal of Theoretical Biology, 112(2):265–273, January 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385802861.

Huf:1985:CAF

[HM85b] Ernst G. Huf and Donald C. Mikulecky. Compartmental analysis of the Na + flux ratio with application to data on frog skin epidermis. Journal of Theoretical Biology, 112(1):193–220, January 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385801247.

Houston:1987:SAM

[HM87] Alasdair I. Houston and John M. McNamara. Singing to attract a mate: a stochastic dynamic game. Journal of Theoretical Biology, 129(1):57–68, November 7, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387802035.

Haggarty:1988:MFW

[HMF88] P. Haggarty, B. A. McGaw, and M. F. Franklin. Measurement of fractionated water loss and CO 2 production using triply labelled water. Journal of Theoretical Biology, 134(3):291–308, October 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388800602.

Hamada:1982:OTS

[HN82] T. Hamada and Y. Nakamura. On the oscillatory transient stage structure of yeast population. Journal of Theoretical Biology, 99(4):797–805, December 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382901989.

Hiram:1984:MAV

[HN84] Yael Hiram and Avinoam Nir. On mechanical aspects of vesicular transport. Journal of Theoretical Biology, 111(1):91–113, November 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384801988.

Hagedorn:1985:EPE

[HN85] Rolf Hagedorn and Ekkehart Neef. The efficiency of photosynthetic energy conversion in continuous and intermittent light. Journal of Theoretical Biology, 114(1):93–101, May 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385802575.

Honda:1984:CSC

[HO84] Hisao Honda and Setsuko Oshibe. A computer simulation of cell stacking for even thickness in mammalian epidermis. Journal of Theoretical Biology, 111(4):625–633, December 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384802581.

Hoekstra:1980:WDO

[Hoe80] Rolf F. Hoekstra. Why do organisms produce gametes of only two different sizes? Some theoretical aspects of the evolution of anisogamy. Journal of Theoretical Biology, 87(4):785–793, ????

1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380901174.

Hoekstra:1982:ASE

[Hoe82] Rolf F. Hoekstra. On the asymmetry of sex: Evolution of mating types in isogamous populations. Journal of Theoretical Biology, 98(3):427–451, October 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382901291.

Hoekstra:1984:EGMb

[Hoe84] Rolf F. Hoekstra. Evolution of gamete motility differences II. Interaction with the evolution of anisogamy. Journal of Theoretical Biology, 107(1):71–83, March 7, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801228.

Hofman:1982:TCT

[Hof82] Michel A. Hofman. A two-component theory of encephalization in mammals. Journal of Theoretical Biology, 99(3):571–584, December 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382902119.

Hofman:1983:EBS

[Hof83] Michel A. Hofman. Evolution of brain size in neonatal and adult placental mammals: a theoretical approach. Journal of Theoretical Biology, 105(2):317–332, ???? 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519383800113.

Hofman:1985:NCC

[Hof85] Michel A. Hofman. Neuronal correlates of corticalization in mammals: a theory. Journal of Theoretical Biology, 112(1):77–95, January 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251938580117X.

Hoffmann:1986:NNM

[Hof86] Geoffrey W. Hoffmann. A neural network model based on the analogy with the immune system. Journal of Theoretical Biology, 122(1):33–67, September 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386802247.

Hoffmann:1987:CKU

[Hof87] Geoffrey W. Hoffmann. A critique of the Kaufman-urbain-Thomas immune system network theory. Journal of Theoretical Biology, 129(3):355–357, December 7, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387800085.

Hokkanen:1986:NCE

[Hok86a] J. E. I. Hokkanen. Notes concerning elastic similarity. Journal of Theoretical Biology, 120(4):499–501, June 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386800431.

Hokkanen:1986:SLL

[Hok86b] J. E. I. Hokkanen. The size of the largest land animal. Journal of Theoretical Biology, 118(4):491–499, February 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801679.

Holder:1983:DCE

[Hol83a] Nigel Holder. Developmental constraints and the evolution of vertebrate digit patterns. Journal of Theoretical Biology, 104(3):451– 471, October 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383901170.

Holt:1983:SGV

[Hol83b] C. Holt. Swelling of Golgi vesicles in mammary secretory cells and its relation to the yield and quantitative composition of milk. Journal of Theoretical Biology, 101(2):247–261, March 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383903399.

Holt:1985:DIM

[Hol85] Robert D. Holt. Density-independent mortality, non-linear competitive interactions, and species coexistence. Journal of Theoretical Biology, 116(4):479–493, October 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800849.

Holliday:1988:SLM

[Hol88] Robin Holliday. Successes and limitations of molecular biology. Journal of Theoretical Biology, 132(3):253–262, June 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388802145.

Holgate:1989:VSM

[Hol89] Philip Holgate. Varieties of stochastic model: a comparative study of the Gompertz effect. Journal of Theoretical Biology, 139(3):369–377, August 9, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389802152.

Honderich:1984:TCS

[Hon84] Ted Honderich. The time of a conscious sensory experience and mind-brain theories. Journal of Theoretical Biology, 110 (1):115–129, September 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519384800181.

Honderich:1986:MBT

[Hon86] Ted Honderich. Mind, brain and time: Rejoinder to libet. Journal of Theoretical Biology, 118(3):367–375, February 7, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386800674.

Hopkins:1983:TBB

[Hop83] Robert C. Hopkins. Transitions between b-DNA and z-DNA: a dilemma. Journal of Theoretical Biology, 101(3):327–333, April 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383901418.

Hopkins:1986:UFS

[Hop86] Robert C. Hopkins. A unique four-stranded model of a homologous recombination intermediate. Journal of Theoretical Biology, 120(2):215–222, May 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801758.

Horsfield:1980:DLB

[Hor80] Keith Horsfield. Are diameter, length and branching ratios meaningful in the lung? Journal of Theoretical Biology, 87(4):773–784, ???? 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380901162.

Horii:1989:TEP

[Hor89a] Zene Ichiro Horii. Tooth eruption: The phase transition theory on biological formation of an orderly structure. Journal of Theoretical Biology, 139(4):449–464, August 22, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389800657.

Horvath:1989:GOO

[Hor89b]  Gábor Horváth.  Geometric optical optimization of the

corneal lens of notonecta glauca. Journal of Theoretical Biology, 139(3):389–404, August 9, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389802176.

Houston:1983:CLE

[Hou83] Alasdair Houston. Comments on “learning the evolutionarily stable strategy”. Journal of Theoretical Biology, 105(1):175–178, November 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383904320.

Houston:1986:OFV

[Hou86] Alasdair I. Houston. The optimal flight velocity for a bird exploiting patches of food. Journal of Theoretical Biology, 119(3):345– 362, April 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519386801461.

Houston:1987:OFP

[Hou87] Alasdair I. Houston. Optimal foraging by parent birds feeding dependent young. Journal of Theoretical Biology, 124(3):251– 274, February 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387801157.

Hammerstein:1982:AWA

[HP82] Peter Hammerstein and Geoffrey A. Parker. The asymmetric war of attrition. Journal of Theoretical Biology, 96(4):647–682, June 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382902351.

Haigh:1980:EGA

[HR80a] John Haigh and Michael R. Rose. Evolutionary game auctions. Journal of Theoretical Biology, 85(2):381–397, July 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380900272.

Heinrich:1980:MMT

[HR80b] Reinhart Heinrich and Tom A. Rapoport. Mathematical modelling of translation of mRNA in eucaryotes; steady states, timedependent processes and application to reticulocytest. Journal of Theoretical Biology, 86(2):279–313, September 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380900089.

Hejnowicz:1984:GTP

[HR84] Z. Hejnowicz and John A. Romberger. Growth tensor of plant organs. Journal of Theoretical Biology, 110(1):93–114, September 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938480017X.

Horovitz:1985:PPI

[HR85] Amnon Horovitz and Meir Rigbi. Protein-protein interactions: Additivity of the free energies of association of amino acid residues. Journal of Theoretical Biology, 116(1):149–159, September 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385801351.

Hejnowicz:1986:CRC

[HR86] Z. Hejnowicz and John A. Romberger. Calculation of rates of cell division in shoot meristems with reference to the growth tensor. Journal of Theoretical Biology, 121(1):59–71, July 7, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386800285.

Harpending:1987:WMI

[HR87] Henry Harpending and Alan Rogers. On Wright’s mechanism for intergroup selection. Journal of Theoretical Biology, 127(1): 51–61, July 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387801601.

Harley:1982:LRD

[HRG82] C. B. Harley, R. J. Shmookler Reis, and S. Goldstein. Loss of repetitious DNA in proliferating somatic cells may be due to unequal recombination. Journal of Theoretical Biology, 94(1):1–12, January 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382903265.

Heinrich:1981:ASE

[HS81] Reinhart Heinrich and Ingrid Sonntag. Analysis of the selection equations for a multivariable population model. Deterministic and stochastic solutions and discussion of the approach for populations of self-reproducing biochemical networks. Journal of Theoretical Biology, 93(2):325–361, November 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381901089.

Houston:1987:LRM

[HS87] Alasdair I. Houston and Brian H. Sumida. Learning rules, matching and frequency dependence. Journal of Theoretical Biology, 126(3):289–308, June 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387802369.

Hassard:1989:IPS

[HS89] Brian D. Hassard and Lie June Shiau. Isolated periodic solutions of the Hodgkin–Huxley equations. Journal of Theoretical Biology, 136(3):267–279, February 8, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389801638.

Hu:1988:MPP

[HSd^{+}88] M. Hu, P. J. Sinko, A. L. J. deMeere, D. A. Johnson, and G. L. Amidon. Membrane permeability parameters for some amino acids and β-lactam antibiotics: Application of the boundary layer approach. Journal of Theoretical Biology, 131(1):107–114, March 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388801243.

Hsu:1987:MCS

[Hsu87a] Jyh-Ping Hsu. A Monte Carlo simulation of bacterial adhesion. Journal of Theoretical Biology, 126(1):7–13, May 7, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387800978.

Hsu:1987:SMBa

[Hsu87b] Jyh-Ping Hsu. Stochastic modeling of bacterial adhesion: a two-step mechanism with linear adhesion rate. Journal of Theoretical Biology, 124(4):495–504, February 21, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387802230.

Hsu:1987:SMBc

[Hsu87c] Jyh-Ping Hsu. Stochastic modeling of bacterial adhesion: a two-step mechanism with nonlinear adhesion rate. Journal of Theoretical Biology, 129(1):9–15, November 7, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938780200X.

Hsu:1987:SMBb

[Hsu87d] Jyh-Ping Hsu. Stochastic modeling of bacterial adhesion in an open flow system I. Single mixed tank. Journal of Theoretical Biology, 128(3):279–288, October 7, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387800711.

Hsu:1987:TAB

[Hsu87e] Jyh-Ping Hsu. Transient analysis of bacterial adhesion. Journal of Theoretical Biology, 126(4):449–455, June 21, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387801510.

Hsu:1988:BSS

[Hsu88] Jyh-Ping Hsu. Biodigestion of solid substrate: Effect of random initial condition. Journal of Theoretical Biology, 132(4):381– 386, June 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519388800791.

Harrison:1984:TMP

[HSV84] Lionel G. Harrison, J. Snell, and R. Verdi. Turing’s model and pattern adjustment after temperature shock, with application to acetabularia whorls. Journal of Theoretical Biology, 106(1):59– 78, January 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519384900274.

Hearon:1980:SDF

[HT80] John Z. Hearon and Yoshihiko Tsukamoto. Simple derivation of the function of genetic information and correction of the proof. Journal of Theoretical Biology, 83(3):525–526, April 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380900594.

Hoffman:1983:ESP

[HT83] Arnold Hoffman and Edward L. Tepper. Energy storage in photosynthetic photophosphorylation: Lessons from the physical chemistry of the photographic latent image. Journal of Theoretical Biology, 103(1):43–62, July 7, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383901984.

Hutchinson:1984:PAS

[HT84] G. Evelyn Hutchinson and Nils Tongring. The possible adaptive significance of the Brooks-dyar rule. Journal of Theoretical Biology, 106(3):437–439, February 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519384900407.

Horie:1987:MCM

[HT87] Shigeo Horie and Naosada Takizawa. Molecular characteristics of many hemoproteins: a survey of molecular weights, sedimentation coefficients, other molecular parameters and amino acid compositions. Journal of Theoretical Biology, 129(1):117–137, November 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387802060.

Hsu:1988:KAM

[HT88] Jyh-Ping Hsu and Shiojenn Tseng. Kinetic analysis of mating behavior in flies: a transition stage model. Journal of Theoretical Biology, 130(2):175–182, January 21, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388800924.

Hart:1989:GAS

[HT89a] T. N. Hart and L. E. H. Trainor. Geometrical aspects of surface morphogenesis. Journal of Theoretical Biology, 138(3):271– 296, June 3, 1989. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251938980195X.

Hsu:1989:SKM

[HT89b] J. P. Hsu and Shiojenn Tseng. A simple kinetic model for the repair of radiation-induced DNA double-strand breaks. Journal of Theoretical Biology, 136(3):357–359, February 8, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389801705.

Hsu:1989:EAP

[HT89c] Jyh-Ping Hsu and Shiojenn Tseng. Estimating the adjustable parameters of the double Michaelis–Menten model. Journal of Theoretical Biology, 137(3):355–363, April 6, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389800773.

Hart:1989:DEC

[HTG89] Trevor N. Hart, L. E. H. Trainor, and B. C. Goodwin. Diffusion effects in calcium-regulated strain fields. Journal of Theoretical Biology, 136(3):327–336, February 8, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389801687.

Hirata:1985:ITA

[HU85] Hironori Hirata and Robert E. Ulanowicz. Information theoretical analysis of the aggregation and hierarchical structure of ecological networks. Journal of Theoretical Biology, 116(3):321–341, October 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251938580271X.

Huang:1980:CRD

[Hua80a] Henry H. Huang. Chemical reaction of DNA with cross-linking agents: Influences of base permutation on chemical reactivity. Journal of Theoretical Biology, 87(3):585–596, December 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380902374.

Huang:1980:DDC

[Hua80b] Henry H. Huang. Detection and differentiation of crosslink formation in an eucaryotic DNA population. Journal of Theoretical Biology, 87(3):597–608, December 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380902386.

Hughes:1986:KCS

[Hug86] Austin L. Hughes. Kin coalitions and social dominance. Journal of Theoretical Biology, 123(1):55–66, November 7, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386802351.

Humphries:1981:CSB

[Hum81] R. L. Humphries. A computer simulation of the binding between flexible ligands and their polymeric receptor binding sites. Journal of Theoretical Biology, 91(3):477–491, August 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251938190268X.

Hummel:1983:AAF

[Hum83] D. Hummel. Aerodynamic aspects of formation flight in birds. Journal of Theoretical Biology, 104(3):321–347, October 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383901108.

Hunding:1981:PPG

[Hun81] Axel Hunding. Possible prepatterns governing mitosis: The mechanism of spindle-free chromosome movement in aulacantha scolymantha. Journal of Theoretical Biology, 89(3):353–385, April 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251938190357X.

Hunding:1985:MPD

[Hun85a] Axel Hunding. Morphogen prepatterns during mitosis and cytokinesis in flattened cells: Three dimensional Turing structures of reaction–diffusion systems in cylindrical coordinates. Journal of Theoretical Biology, 114(4):571–588, June 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800448.

Hunt:1985:PLS

[Hun85b] Fern Hunt. Patterns of LPS synthesis in Gram negative bacteria. Journal of Theoretical Biology, 115(2):213–219, July 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385800977.

Hutchinson:1989:CGS

[Hut89] J. M. C. Hutchinson. Control of gastropod shell shape; the role of the preceding whorl. Journal of Theoretical Biology, 140(4):431– 444, October 23, 1989. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519389801079.

Huxley:1986:PES

[Hux86] Janet R. Huxley. A possible explanation for the symbolic nature of language. Journal of Theoretical Biology, 123(3):321–324, December 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519386802478.

Hoekstra:1986:CSH

[HvL86] Rolf F. Hoekstra and E. N. van Loo. The cost of sex in hermaphrodite populations with variation in functional sex. Journal of Theoretical Biology, 122(4):441–452, October 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801849.

Holmes:1980:REP

[HW80a] M. G. Holmes and E. Wagner. A re-evaluation of phytochrome involvement in time measurement in plants. Journal of Theoretical Biology, 83(2):255–265, March 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380902921.

Homer:1980:VDT

[HW80b] Louis D. Homer and Paul K. Weathersby. The variance of the distribution of traversal times in a capillary bed. Journal of Theoretical Biology, 87(2):349–377, November 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380903641.

Hameroff:1982:IPM

[HW82] Stuart R. Hameroff and Richard C. Watt. Information processing in microtubules. Journal of Theoretical Biology, 98(4):549–561, October 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382901370.

Hsu:1986:SAB

[HW86] Jyh-Ping Hsu and Hsi-Hua Wang. A stochastic analysis of bacterial adhesion. Journal of Theoretical Biology, 119(4):435–444, April 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251938680193X.

Hsu:1987:KBA

[HW87] Jyh-Ping Hsu and Hsi-Hua Wang. Kinetics of bacterial adhesion — a stochastic analysis. Journal of Theoretical Biology, 124(4):405–413, February 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387802187.

Haig:1988:MOH

[HW88] David Haig and Mark Westoby. A model for the origin of heterospory. Journal of Theoretical Biology, 134(2):257–272, September 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519388802030.

Hwang:1985:LM

[Hwa85] Robert H. Hwang. A lignification mechanism. Journal of Theoretical Biology, 116(1):21–44, September 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801296.

Honda:1984:CSG

[HYDS84] H. Honda, H. Yamanaka, and M. Dan-Sohkawa. A computer simulation of geometrical configurations during cell division. Journal of Theoretical Biology, 106(3):423–435, February 7, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519384900390.

Hyver:1980:STS

[Hyv80] C. Hyver. Stability of transformation systems representable by tree graphs: Extension to structures of biological importance. Journal of Theoretical Biology, 82(2):187–204, January 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380900983.

Hyver:1984:LGL

[Hyv84] C. Hyver. Local and global limit cycles in biochemical systems. Journal of Theoretical Biology, 107(2):203–209, March 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384800223.

Ishay:1980:PCG

[IAHA80] Jacob S. Ishay, Revital Attia, Yosef Hochberg, and Mina Goldstein (Asanta). Photoresistivity of the cuticle in the gaster segments in the ant species cataglyphis bicolor, messor arenarius and messor semirufus. Journal of Theoretical Biology, 85(2):371– 379, July 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519380900260.

Iberall:1983:WLC

[Ibe83] A. S. Iberall. What is “language” that can facilitate the flow of information? A contribution to a fundamental theory of language and communication. Journal of Theoretical Biology, 102(3):347– 359, June 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383903739.

Iannaccone:1989:PDS

[IBJ^{+}89] P. M. Iannaccone, L. Berkwits, J. Joglar, J. Lindsay, and A. Lunde. Probabilistic division systems modeling the generation of mosaic fields. Journal of Theoretical Biology, 141(3):363– 366, IN1–IN2, 367–374, IN3, 375–377, December 7, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389801195.

Iwasa:1985:THC

[ICL85] Yoh Iwasa, Dan Cohen, and Jesus Alberto Leon. Tree height and crown shape, as results of competitive games. Journal of Theoretical Biology, 112(2):279–297, January 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385802885.

Innis:1980:NMC

[IH80] George S. Innis and James W. Haefner. A neutral model of community organization. Journal of Theoretical Biology, 87(3):529–558, December 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519380902349.

Infante:1980:NKM

[IHK80] Juan P. Infante, George E. Houghton, and John E. Kinsella. A novel kinetic mechanism explaining the non-hyperbolic behavior of metal activated enzymes. Case of choline kinase from rat liver. Journal of Theoretical Biology, 86(1):177–201, September 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380900740.

Ishida:1988:IPS

[IHTN88] Tetsuo Ishida, Kihachiro Horiike, Hiromasa Tojo, and Mitsuhiro Nozaki. Interaction of protein with a self-associating ligand. Deviation from a hyperbolic binding curve and the appearance of apparent co-operativity in the scatchard plot. Journal of Theoretical Biology, 130(1):49–66, January 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388801632.

Iida:1988:CDA

[Iid88] Yôichi Iida. Categorical discriminant analysis of 3’-splice site signals of mRNA precursors in higher eukaryote genes. Journal of Theoretical Biology, 135(1):109–118, November 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388801772.

Ives:1985:CWB

[IM85] Anthony R. Ives and Robert M. May. Competition within and between species in a patchy environment: Relations between microscopic and macroscopic models. Journal of Theoretical Biology, 115(1):65–92, July 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800072.

Incardona:1983:BVL

[Inc83a] N. L. Incardona. Binding of viruses and ligands to cells: Effect of receptor density. Journal of Theoretical Biology, 104(4):693–699, October 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383902564.

Incardona:1983:KMV

[Inc83b] N. L. Incardona. A kinetic model for virus binding which involves release of cell-bound virus-receptor complexes. Journal of Theoretical Biology, 105(4):631–645, December 21, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383902242.

Infante:1985:IBH

[Inf85] Juan P. Infante. Impaired biosynthesis of highly unsaturated phosphatidylcholines: a hypothesis on the molecular etiology of some muscular dystrophies. Journal of Theoretical Biology, 116(1):65–88, September 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801314.

Inghe:1989:GRS

[Ing89] Ola Inghe. Genet and ramet survivorship under different mortality regimes — a cellular automata model. Journal of Theoretical Biology, 138(2):257–270, May 22, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389801420.

Isogai:1983:DFM

[INMM83] Y. Isogai, H. Nose, K. Miki, and T. Morimoto. Dynamics of fluid movement between intravascular and interstitial spaces. Journal of Theoretical Biology, 100(2):305–317, January 21, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383903545.

Imamura:1983:MOT

[IO83] Akira Imamura and Masaru Ohsaku. A molecular orbital treatment on the reactivity of some polycyclic aromatic hydrocarbon diol-epoxides. Journal of Theoretical Biology, 102(4):511– 519, June 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383903879.

Izquierdo:1983:SAI

[IP83] Jose M. Izquierdo and Consuelo Perez. A stochastic approach for the interpretation of single pulse experiments in morphological multicompartments of renewing and exponentially growing cell populations. Journal of Theoretical Biology, 101(1):39–75, March 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383902722.

Ipina:1987:MPA

[Ipi87] Santiago L. Ipiña. A multivariate phenetic approach to neuronal nuclei resemblances across species, with examples from three regions of cerebral cortex of a mammal, a bird and a reptile. Journal of Theoretical Biology, 126(1):105–124, May 7, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387801042.

Ishay:1983:DFH

[IPS83] Jacob S. Ishay, Vered Paniry, and Dror Sadeh. Direction finding by hornets in a vertically rotating centrifuge. Journal of Theoretical Biology, 102(2):269–276, May 21, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383903648.

Ishida:1981:NET

[Ish81] Kenzi Ishida. Non-equilibrium thermodynamics of the selection of biological macromolecules. Journal of Theoretical Biology, 88(2):257–273, January 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381900746.

Ishida:1984:GIE

[Ish84] Kenzi Ishida. Gain of information and evolutionarily stable strategy in hypercyclic organization. Journal of Theoretical Biology, 108(3):469–474, June 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384800478.

Ishida:1986:FMH

[Ish86] Kenzi Ishida. Fluctuations and mutation in hypercyclic organization. Journal of Theoretical Biology, 118(1):3–13, January 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386800042.

Iwasa:1980:CLH

[IT80] Yoh Iwasa and Ei Teramoto. A criterion of life history evolution based on density dependent selection. Journal of Theoretical Biology, 84(3):545–566, June 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519380800191.

Iwasa:1982:MEU

[IT82] K. Iwasa and I. Tasaki. Molecular events that underlie membrane excitation. Journal of Theoretical Biology, 99(1):87–99, November 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382903915.

Imamura:1988:MOS

[ITK88] Akira Imamura, Seiji Tani, and Katsuya Kanda. Molecular orbital study on the metabolic pathway through the diol epoxide form of carcinogenic benzene in comparison with benzo[ a ]pyrene. Journal of Theoretical Biology, 135(2):215–218, November 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388800754.

Ito:1982:QMC

[Ito82] Masao Ito. Questions in modeling the cerebellum. Journal of Theoretical Biology, 99(1):81–86, November 7, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382903903.

Ives:1988:CCP

[Ive88] Anthony R. Ives. Covariance, coexistence and the population dynamics of two competitors using a patchy resource. Journal of Theoretical Biology, 133(3):345–361, August 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388803266.

Iwasa:1981:RSR

[Iwa81] Yoh Iwasa. Role of sex ratio in the evolution of eusociality in haplodiploid social insects. Journal of Theoretical Biology, 93(1):125–142, November 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381900606.

Iwasa:1988:FFA

[Iwa88] Yoh Iwasa. Free fitness that always increases in evolution. Journal of Theoretical Biology, 135(3):265–281, December 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388802431.

Wu:1988:PCW

[iWSS88] Hsin i Wu, Richard D. Spence, and Peter J. H. Sharpe. Plant cell wall elasticity II: Polymer elastic properties of the microfibrils. Journal of Theoretical Biology, 133(2):239–253, July 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388800080.

Ikeda:1983:DEM

[IYS83] N. Ikeda, S. Yoshizawa, and T. Sato. Difference equation model of ventricular parasystole as an interaction between cardiac pacemakers based on the phase response curve. Journal of Theoretical Biology, 103(3):439–465, August 7, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383902977.

Izutsu:1981:TMB

[Izu81] Kenneth T. Izutsu. Theory and measurement of the buffer value of bicarbonate in saliva. Journal of Theoretical Biology, 90(3): 397–403, June 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381903209.

Johnson:1988:DIM

[JA88] Donald A. Johnson and Gordon L. Amidon. Determination of intrinsic membrane transport parameters from perfused intestine experiments: a boundary layer approach to estimating the aqueous and unbiased membrane permeabilities. Journal of Theoretical Biology, 131(1):93–106, March 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388801231.

Jackson:1982:HMI

[Jac82] Robert R. Jackson. Habituation as a mechanism of intersexual selection. Journal of Theoretical Biology, 97(2):333–335, July 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382901096.

Jackson:1983:WUL

[Jac83] K. M. Jackson. Why the upper limbs move during human walking. Journal of Theoretical Biology, 105(2):311–315, ???? 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519383800101.

Jaffe:1980:TAC

[Jaf80] Klaus Jaffe. Theoretical analysis of the communication system for chemical mass recruitment in ants. Journal of Theoretical Biology, 84(4):589–609, June 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519380800221.

Jaffe:1984:NEC

[Jaf84] Klaus Jaffe. Negentropy and the evolution of chemical recruitment in ants (Hymenoptera: Formicidae). Journal of Theoretical Biology, 106(4):587–604, February 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519384900092.

Jakubowski:1988:NCB

[Jak88] Hieronim Jakubowski. Negative correlation between the abundance of Escherichia coli aminoacyl-tRNA families and their affinities for elongation factor Tu-GTP. Journal of Theoretical Biology, 133(3):363–370, August 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388803278.

James:1980:LRR

[Jam80] Alan T. James. Liver redox resistance: a dynamic model of gluconeogenic lactate metabolism. Journal of Theoretical Biology, 83(4):623–646, April 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380901927.

James:1982:NLR

[Jam82] Alan T. James. A note on the linear relation between lactate redox potential and the hydrogen shuttle flux. Journal of Theoretical Biology, 94(1):129–133, January 7, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382903356.

James:1985:SRD

[Jam85] William H. James. Sex ratio, dominance status and maternal hormone levels at the time of conception. Journal of Theoretical Biology, 114(3):505–510, June 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801818.

James:1986:HPS

[Jam86a] William H. James. Handedness of parents and sex of progeny. Journal of Theoretical Biology, 122(2):243–245, September 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386800868.

James:1986:HCS

[Jam86b] William H. James. Hormonal control of sex ratio. Journal of Theoretical Biology, 118(4):427–441, February 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801631.

James:1987:HLP

[Jam87] W. H. James. Hormone levels of parents and sex ratios of offspring. Journal of Theoretical Biology, 129(1):139–140, November 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519387802072.

James:1988:TLH

[Jam88] William H. James. Testosterone levels, handedness and sex ratio at birth. Journal of Theoretical Biology, 133(3):261–266, August 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388803205.

James:1989:FTL

[Jam89a] William H. James. Foetal testosterone levels, homosexuality and handedness: a research proposal for jointly testing Geschwind’s and Dörner’s hypotheses. Journal of Theoretical Biology, 136(2):177–180, January 23, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389802243.

James:1989:PHLa

[Jam89b] William H. James. Parental hormone levels and mammalian sex ratios at birth. Journal of Theoretical Biology, 139(1):59–67, July 10, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519389800578.

James:1989:PHLb

[Jam89c] William H. James. Parental hormone levels and the possibility of establishing that some mammalian sex ratio variation is adaptive. Journal of Theoretical Biology, 140(1):39–40, September 11, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519389800281.

Janetos:1982:AFV

[Jan82] Anthony C. Janetos. Active foragers vs. sit-and-wait predators: a simple model. Journal of Theoretical Biology, 95(2):381–385, March 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382902521.

Jung:1980:TMS

[JB80] A. Jung and P. Bartholdi. Turnover of multicompartment systems. Journal of Theoretical Biology, 85(4):657–663, August 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380902635.

Jagadeeswaran:1980:GMC

[JC80] P. Jagadeeswaran and Joseph D. Cherayil. A general model for the conformational switch in 5s RNA during protein synthesis. Journal of Theoretical Biology, 83(2):369–375, March 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251938090301X.

Johnson:1980:SEC

[JCK80] Edward A. Johnson, J. Brian Chapman, and J. Mailen Kootsey. Some electrophysiological consequences of electrogenic sodium and potassium transport in cardiac muscle: a theoretical study. Journal of Theoretical Biology, 87(4):737–756, ???? 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380901149.

Johnson:1984:MTG

[JCZ84] Byron F. Johnson, G. B. Calleja, and M. Zuker. Matingtype gene switching in a homothallic fission yeast. Journal of Theoretical Biology, 110(2):299–312, September 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384800600.

Jackson:1982:MAP

[JD82] Robert J. Jackson and W. B. Davis. A mathematical approach to the prediction of the rate of penetration of ions in the sweat duct. Journal of Theoretical Biology, 97(3):481–489, August 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382903769.

Jean:1980:SMG

[Jea80] Roger V. Jean. A systemic model of growth in botanometry. Journal of Theoretical Biology, 87(3):569–584, December 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380902362.

Jean:1985:GRP

[Jea85] Roger V. Jean. A geometric representation of α-polypeptide chains revisited. Journal of Theoretical Biology, 112(4):819–825, February 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385800631.

Jean:1987:MMM

[Jea87] Roger V. Jean. A mathematical model and a method for the practical assessment of the phyllotactic patterns. Journal of Theoretical Biology, 129(1):69–90, November 7, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387802047.

Jou:1985:SNT

[JF85] David Jou and Fernando Ferrer. A simple nonequilibrium thermodynamic description of some inhibitors of oxidative phosphorylation. Journal of Theoretical Biology, 117(3):471–488, December 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385801557.

Jones:1983:UCC

[JGW83] T. D. Jones, G. D. Griffin, and P. J. Walsh. A unifying concept for carcinogenic risk assessments. Journal of Theoretical Biology, 105(1):35–61, November 7, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251938390423X.

Ji:1985:BMM

[Ji85] Sungchul Ji. The bhopalator: a molecular model of the living cell based on the concepts of conformons and dissipative structures. Journal of Theoretical Biology, 116(3):399–426, October 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385802782.

Ji:1988:WCP

[Ji88] Sungchul Ji. Watson-crick and prigoginian forms of genetic information. Journal of Theoretical Biology, 130(2):239–245, January 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388800997.

Jeng:1980:NMV

[JJ80] Yunhua Jeng and Ingming Jeng. A new model for very low density lipoprotein metabolism — nomenclature for very low density lipoprotein derivatives. Journal of Theoretical Biology, 86(2):237–245, September 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251938090003X.

Jablonka:1982:NRS

[JJ82a] Paul Jablonka and Eileen Jablonka. Non-random sister chromatid segregation by cell type. Journal of Theoretical Biology, 99(3):427–436, December 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902041.

Jablonka:1982:CHP

[JJ82b] Paul Jablonka and Eileen Jablonka. On the control of hnRNA production. Journal of Theoretical Biology, 99(3):407–425, December 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251938290203X.

Jackson:1986:GND

[JJN86] Ernest R. Jackson, Diana Johnson, and William G. Nash. Gene networks in development. Journal of Theoretical Biology, 119(4):379–396, April 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519386801904.

Johansen:1981:FME

[JK81] Søren Johansen and Sussane Keiding. A family of models for the elimination of substrate in the liver. Journal of Theoretical Biology, 89(4):549–556, April 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381900278.

Jamaluddin:1987:REB

[JK87] Moideen P. Jamaluddin and Lissy K. Krishnan. A rate equation for blood platelet aggregation. Journal of Theoretical Biology, 129(2):257–261, November 21, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387800176.

Janovy:1988:MEB

[JK88] J. Janovy and G. W. Kutish. A model of encounters between host and parasite populations. Journal of Theoretical Biology, 134(3):391–401, October 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388800699.

Jeje:1989:ARR

[JK89] Ayodeji Jeje and Darren Koon. An analysis on the rates and regulation of insensible water loss through the eccrine sweat glands. Journal of Theoretical Biology, 141(3):303–324, December 7, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519389801171.

Jablonka:1988:MPC

[JL88] Eva Jablonka and Marion J. Lamb. Meiotic pairing constraints and the activity of sex chromosomes. Journal of Theoretical Biology, 133(1):23–36, July 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388800225.

Jablonka:1989:IAE

[JL89] Eva Jablonka and Marion J. Lamb. The inheritance of acquired epigenetic variations. Journal of Theoretical Biology, 139(1):69– 83, July 10, 1989. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251938980058X.

Jones:1980:SMC

[JLS80] Richard B. Jones, Charles K. Lumpkin, and James R. Smith. A stochastic model for cellular senescence. Part I. Theoretical considerations. Journal of Theoretical Biology, 86(3):581–592, October 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519380903549.

Jarvis:1983:CCS

[JLS83] J. P. Jarvis, J. K. Luedeman, and D. R. Shier. Comments on computing the similarity of binary trees. Journal of Theoretical Biology, 100(3):427–433, February 7, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383904393.

John:1984:MRC

[Joh84] Philip John. Mitochondrial regulation of cell surface components in relation to carcinogenesis. Journal of Theoretical Biology, 110(3):377–381, October 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801800.

Johnson:1986:HDE

[Joh86] Richard B. Johnson. Human disease and the evolution of pathogen virulence. Journal of Theoretical Biology, 122(1):19–24, September 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519386802223.

Jolliffe:1988:EEC

[Jol88] Peter A. Jolliffe. Evaluating the effects of competitive interference on plant performance. Journal of Theoretical Biology, 130(4):447–459, February 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388802091.

Jolicoeur:1989:SMB

[Jol89] Pierre Jolicoeur. A simplified model for bivariate complex allometry. Journal of Theoretical Biology, 140(1):41–49, September 11, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519389800293.

Jones:1985:NFC

[Jon85a] Karen Gay Cronquist Jones. Nitrogen fixation as a control in the nitrogen cycle. Journal of Theoretical Biology, 112(2):315–332, January 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385802903.

Jongeling:1985:AET

[Jon85b] T. B. Jongeling. On an axiomatization of evolutionary theory. Journal of Theoretical Biology, 117(4):529–543, December 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385802368.

Jonsson:1985:CDS

[Jøn85c] Viggo Jønsson. Comparison and definition of spleen and lymph node: a phylogenetic analysis. Journal of Theoretical Biology, 117(4):691–699, December 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385802472.

Johnson:1985:TAG

[JP85] I. R. Johnson and A. J. Parsons. A theoretical analysis of grass growth under grazing. Journal of Theoretical Biology, 112(2):345– 367, January 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385802927.

Jolicoeur:1989:PGD

[JP89a] Pierre Jolicoeur and Jacques Pontier. Population growth and decline: a four-parameter generalization of the logistic curve. Journal of Theoretical Biology, 141(4):563–571, December 19, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519389802371.

Joshi:1989:MDHa

[JP89b] Abhay Joshi and Bernhard O. Palsson. Metabolic dynamics in the human red cell: Part I — a comprehensive kinetic model. Journal of Theoretical Biology, 141(4):515–528, December 19, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389802334.

Joshi:1989:MDHb

[JP89c] Abhay Joshi and Bernhard O. Palsson. Metabolic dynamics in the human red cell: Part II-interactions with the environment. Journal of Theoretical Biology, 141(4):529–545, December 19, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519389802346.

Jou:1986:BFR

[JPGL86] D. Jou, C. Perez-Garcia, and J. E. Llebot. Bacterial flagellar rotation as a nonequilibrium phase transition. Journal of Theoretical Biology, 122(4):453–458, October 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801850.

Johnston:1983:NTC

[JPL83] Victor S. Johnston, Derek Partridge, and Patricia D. Lopez. A neural theory of cognitive development. Journal of Theoretical Biology, 100(3):485–509, February 7, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383904423.

Jungerman:1980:MIS

[JR80] Roger L. Jungerman and Bruce Rosenblum. Magnetic induction for the sensing of magnetic fields by animals — an analysis. Journal of Theoretical Biology, 87(1):25–32, November 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380902179.

Jerzmanowski:1980:MTC

[JS80] Andrzej Jerzmanowski and Krzysztof Staron. Mg^{2+} as a trigger of condenstion-decondensation transition of chromatin during mitosis. Journal of Theoretical Biology, 82(1):41–46, January 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380900892.

Jerzmanowski:1981:CPH

[JS81] Andrezej Jerzmanowski and Krzysztof Staron. Can phosphorylation of histone H1 be responsible for chromatin condensation in mitosis? Journal of Theoretical Biology, 89(1):191–194, March 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381901855.

Jones:1982:SMC

[JS82] Richard B. Jones and James R. Smith. A stochastic model of cellular senescence. II. Concordance with experimental data. Journal of Theoretical Biology, 96(3):443–460, June 7, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382901205.

Job:1988:PMH

[JSJS88] Dominique Job, Jean Michel Soulié, Claudette Job, and David Shire. Potential memory and hysteretic effects in transcription. Journal of Theoretical Biology, 134(3):273–289, October 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388800596.

Johnson:1984:MID

[JT84] I. R. Johnson and J. H. M. Thornley. A model of instantaneous and daily canopy photosynthesis. Journal of Theoretical Biology, 107(4):531–545, April 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801319.

Juretic:1984:EFE

[Jur84] Davor Juretić. Efficiency of free energy transfer and entropy production in photosynthetic systems. Journal of Theoretical Biology, 106(3):315–327, February 7, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251938490033X.

Kokkinidis:1985:NNN

[KA85] M. Kokkinidis and P. Anninos. Noisy neural nets exhibiting epileptic features. Journal of Theoretical Biology, 113(3):559– 588, April 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385800394.

Khan:1989:DAP

[KA89] Azhar Zaman Khan and Leon Aarons. Design and analysis of protein binding experiments. Journal of Theoretical Biology, 140(2):145–166, September 22, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389801250.

Kacser:1987:DIV

[Kac87] Henrik Kacser. Dominance not inevitable but very likely. Journal of Theoretical Biology, 126(4):505–506, June 21, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387801558.

Kam:1980:RSE

[Kam80] Zvi Kam. The reconstruction of structure from electron micrographs of randomly oriented particles. Journal of Theoretical Biology, 82(1):15–39, January 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380900880.

Kallen:1985:SMS

[KAM85] A. Källén, P. Arcuri, and J. D. Murray. A simple model for the spatial spread and control of rabies. Journal of Theoretical Biology, 116(3):377–393, October 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385802769.

Kaplan:1985:REC

[Kap85] J. I. Kaplan. Re-evaluation of the corrected signal from fluoroscopic probe measurements. Journal of Theoretical Biology, 115(4):633–634, August 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519385801466.

Karlin:1980:MEC

[Kar80] Samuel Karlin. Measures of enzyme co-operativity. Journal of Theoretical Biology, 87(1):33–54, November 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380902180.

Karlovsky:1988:CIC

[Kar88a] Petr Karlovský. Calculation of individual cleavage rates from partial digests in restriction endonuclease kinetics. Journal of Theoretical Biology, 132(1):7–14, May 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388801863.

Karlovsky:1988:MCI

[Kar88b] Petr Karlovsky. Mutual competitive inhibition between target sites during the restriction endonuclease digestion of DNA. Journal of Theoretical Biology, 132(1):1–6, May 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388801851.

Kastrikin:1980:FGA

[Kas80] N. F. Kastrikin. Force generation and ATP hydrolysis in muscle contraction. Journal of Theoretical Biology, 84(2):387–400, May 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519380800129.

Katsura:1981:ASM

[Kat81] Isao Katsura. Assembly systems in molecular biology: a graph theory of genetic complementation between two related species. Journal of Theoretical Biology, 88(3):503–512, February 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381902794.

Katsura:1983:TSS

[Kat83] Isao Katsura. Theory on the structure and stability of coated vesicles. Journal of Theoretical Biology, 103(1):63–75, July 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383901996.

Katz:1985:NFW

[Kat85] Murray A. Katz. New formulation of water and macromolecular flux which corrects for non-ideality: Theory and derivation, predictions, and experimental results. Journal of Theoretical Biology, 112(2):369–401, January 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385802939.

Kauffman:1986:ASP

[Kau86] Stuart A. Kauffman. Autocatalytic sets of proteins. Journal of Theoretical Biology, 119(1):1–24, March 7, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386800479.

Kawato:1984:CPN

[Kaw84] Mitsuo Kawato. Cable properties of a neuron model with non-uniform membrane resistivity. Journal of Theoretical Biology, 111(1):149–169, November 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384802027.

Klotz:1981:PMC

[KB81] Lynn C. Klotz and Roger L. Blanken. A practical method for calculating evolutionary trees from sequence data. Journal of Theoretical Biology, 91(2):261–272, July 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902332.

Kleijn:1983:IHI

[KB83] W. B. Kleijn and L. J. Bruner. The influence of hydrophobic ions and dipolar molecules on the electrostatic barrier in biomembranes. Journal of Theoretical Biology, 100(1):139–152, January 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251938390098X.

Khanin:1986:OLC

[KB86] M. A. Khanin and O. G. Bat. Optimal lower critical temperature and thermal insulation of homeotherms. Journal of Theoretical Biology, 123(3):251–260, December 7, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386802417.

Kirkwood:1988:EGP

[KB88a] T. B. L. Kirkwood and I. D. J. Burdett. Estimating the growth pattern of micro-organisms in distinct stages of the cell cycle. Journal of Theoretical Biology, 130(3):255–273, February 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388800298.

Kshatriya:1988:TMM

[KB88b] M. Kshatriya and R. W. Blake. Theoretical model of migration energetics in the blue whale, balaenoptera musculus. Journal of Theoretical Biology, 133(4):479–498, August 21, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388803369.

Khanin:1989:MME

[KB89a] M. A. Khanin and I. B. Bukharov. A mathematical model of the exercise functional state of the oxygen transport system. Journal of Theoretical Biology, 137(2):191–201, March 21, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938980205X.

Klopman:1989:CAS

[KB89b] Gilles Klopman and Rajiv D. Bendale. Computer automated structure evaluation (CASE): a study of inhibitors of the thermolysin enzyme. Journal of Theoretical Biology, 136(1):67–77, January 9, 1989. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519389801900.

Kemp:1988:TII

[KBR88] G. J. Kemp, A. Bevington, and R. G. G. Russell. Theoretical interpretation of isotope labelling experiments in cells in which the label is chemically incorporated: The example of orthophosphate. Journal of Theoretical Biology, 134(3):351–364, October 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388800663.

Keller:1981:LG

[KC81] J. B. Keller and S. Childress. Lichen growth. Journal of Theoretical Biology, 93(2):497, November 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251938190117X.

Kohn:1982:MNS

[KC82] Michael C. Kohn and Ernest Chiang. Metabolic network sensitivity analysis. Journal of Theoretical Biology, 98(1):109–126, September 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382900613.

Kohn:1983:SVR

[KC83a] Michael C. Kohn and Ernest Chiang. Sensitivity to values of the rate constants in a neurochemical metabolic model. Journal of Theoretical Biology, 100(4):551–565, February 21, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383903235.

Kuchel:1983:NSE

[KC83b] P. W. Kuchel and B. E. Chapman. NMR spin exchange kinetics at equilibrium in membrane transport and enzyme systems. Journal of Theoretical Biology, 105(4):569–589, December 21, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383902205.

Kampis:1985:SMD

[KC85] G. Kampis and V. Csányi. Simple models do not eliminate complexity from the real world. Journal of Theoretical Biology, 115(3):467–469, August 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385802046.

Kampis:1987:NOC

[KC87] G. Kampis and V. Csányi. Notes on order and complexity. Journal of Theoretical Biology, 124(1):111–121, January 7, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387802564.

Kuo-Chen:1981:GRN

[KCM81] Chou Kuo-Chen and Liu Wei Min. Graphical rules for non-steady state enzyme kinetic. Journal of Theoretical Biology, 91(4):637– 654, August 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381902150.

Kuchel:1980:CSK

[KD80] Philip W. Kuchel and Keith Dalziel. Computer simulations of the kinetics of reversible ligand binding to a multivalent macromolecule. Consequences of negative binding co-operativity associated with changes of the association and dissociation rate constants. Journal of Theoretical Biology, 85(3):497–505, August 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380903239.

Koch:1985:IOG

[KD85] Arthur L. Koch and Ronald J. Doyle. Inside-to-outside growth and turnover of the wall of gram-positive rods. Journal of Theoretical Biology, 117(1):137–157, November 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801697.

Kurland:1988:MML

[KD88] Irwin J. Kurland and David Z. D’Argenio. A minimal model of liver glycogen metabolism; feasibility for predicting flux rates.

Journal of Theoretical Biology, 135(3):343–358, December 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388802492.

Kimmel:1984:ACC

[KDAT84] Marek Kimmel, Zbigniew Darzynkiewicz, Ovide Arino, and Frank Traganos. Analysis of a cell cycle model based on unequal division of metabolic constituents to daughter cells during cytokinesis. Journal of Theoretical Biology, 110(4):637–664, October 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384801496.

Kell:1985:EEA

[KDG85] Michael J. Kell, Louis J. DeFelice, and W. Goolsby. Equilibrium energy analysis of freeze-fracture planes in membranes. Journal of Theoretical Biology, 115(1):103–127, July 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800096.

Koster:1988:KHG

[KDW88] Johanna G. Koster, Olivier H. J. Destrée, and Hans V. Westerhoff. Kinetics of histone gene expression during early development of xenopus laevis. Journal of Theoretical Biology, 135(2):139–167, November 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519388800717.

Kelly:1987:MER

[Kel87] Christopher Kelly. A model to explore the rate of spread of mimicry and rejection in hypothetical populations of cuckoos and their hosts. Journal of Theoretical Biology, 125(3):283–299, April 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519387800607.

Kerszberg:1989:DCC

[Ker89] Michel Kerszberg. Developmental canalization can enhance species survival. Journal of Theoretical Biology, 139(3):287–309, August 9, 1989. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519389802103.

Keller:1983:CW

[KF83] Joseph B. Keller and Meira S. Falkovitz. Crawling of worms. Journal of Theoretical Biology, 104(3):417–442, October 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383901157.

Klawansky:1984:GRD

[KF84] Sidney Klawansky and Maurice S. Fox. A growth rate distribution model for the age dependence of human cancer incidence: a proposed role for promotion in cancer of the lung and breast. Journal of Theoretical Biology, 111(3):531–587, December 7, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384802386.

Kendal:1987:CGI

[KF87] Wayne S. Kendal and Philip Frost. Constancy of genomic instability in tumor progression. Journal of Theoretical Biology, 126(3):369–371, June 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387802424.

Kappas:1985:SEI

[KFPB85] U. Kappas, W. Fischer, E. E. Polymeropoulos, and J. Brickmann. Solvent effects in ionic transport through transmembrane protein channels. Journal of Theoretical Biology, 112(3):459–464, February 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251938580014X.

Kawato:1982:TOM

[KFSW82] Mitsuo Kawato, Koji Fujita, Ryoji Suzuki, and Arthur T. Winfree. A three-oscillator model of the human circadian system controlling the core temperature rhythm and the sleep-wake cycle. Journal of Theoretical Biology, 98(3):369–392, October 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382901254.

Kat:1982:DMN

[KG82] Michael J. Kat and Ulf Grenander. Developmental matching and the numerical matching hypothesis for neuronal cell death. Journal of Theoretical Biology, 98(3):501–517, October 7, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382901333.

Kohler:1980:UFS

[KH80] Hans-Helmut Kohler and Klaus Heckmann. Unidirectional fluxes in saturated single-file pores of biological and artificial membranes. II. Asymptotic behavior at high degrees of saturation. Journal of Theoretical Biology, 85(4):575–595, August 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380902593.

Khait:1986:HMS

[Kha86] Alexander Khait. Hormonal mechanisms for size measurement in living organisms in the context of maturing juvenile plants. Journal of Theoretical Biology, 118(4):471–483, February 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386801655.

Katz:1983:PAM

[KHB83] Yehuda Katz, Milo E. Hoffman, and Robert Blumenthal. Parametric analysis of membrane characteristics and membrane structure. Journal of Theoretical Biology, 105(3):493–510, December 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383901893.

Kholodenko:1984:CMO

[Kho84] Boris N. Kholodenko. Control of mitochondrial oxidative phosphorylation. Journal of Theoretical Biology, 107(2):179–188, March 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251938480020X.

Kitazoe:1983:TAR

[KHU^{+}83] Yasuhiro Kitazoe, Nobuaki Hiraoka, Hisashi Ueta, Hisakazu Ogura, and … Hideo Saito. Theoretical analysis on relationship between the neural activity and the EEG. Journal of Theoretical Biology, 104(4):667–683, October 21, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383902540.

Kiltie:1980:AST

[Kil80] Richard A. Kiltie. Application of search theory to the analysis of prey aggregation as an antipredation tactic. Journal of Theoretical Biology, 87(1):201–206, November 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380902283.

King:1982:FTC

[Kin82] R. Bruce King. The flow topology of chemical reaction networks. Journal of Theoretical Biology, 98(2):347–368, September 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382902697.

King:1983:RMS

[Kin83] Stephen M. King. A regulatory model for spindle function during mitosis. Journal of Theoretical Biology, 102(4):501–510, June 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383903867.

King:1985:MTB

[Kin85] C. C. King. A model for transposon-based eucaryote regulatory evolution. Journal of Theoretical Biology, 114(3):447–462, June 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385801788.

King:1986:WTP

[Kin86] G. A. M. King. Was there a prebiotic soup? Journal of Theoretical Biology, 123(4):493–498, December 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386802168.

Kobuchi:1988:MRP

[KIO88] Youichi Kobuchi, Tadanao Ito, and Atsushi Ogiwara. A model for rouleaux pattern formation of red blood cells. Journal of Theoretical Biology, 130(2):129–145, January 21, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388800894.

Kirkwood:1980:EPI

[Kir80] T. B. L. Kirkwood. Error propagation in intracellular information transfer. Journal of Theoretical Biology, 82(3):363–382, February 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251938090243X.

Kirkpatrick:1986:SSC

[Kir86] Mark Kirkpatrick. Sexual selection and cycling parasites: a simulation study of Hamilton’s hypothesis. Journal of Theoretical Biology, 119(3):263–271, April 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801394.

Kuima:1980:CRC

[KK80] Hiromasa Kuima and Shoko Kuima. Cooperative response of chemically excitable membrane II. Two-state models and their limitations. Journal of Theoretical Biology, 82(3):425–463, February 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380902489.

Kothekar:1981:BSMb

[KK81] V. Kothekar and S. Kailash. Biophysical study of molecular mechanism of abortificient action of prostaglandins II. The study of the long and short range interaction between different fragments of PGF_{2α}, PGF_{1β} and PGA_{1}. Journal of Theoretical Biology, 93(1):25–40, November 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381900540.

Klinov:1982:TAK

[KK82a] S. V. Klinov and B. I. Kurganov. The theoretical analysis of kinetic behaviour of “hysteretic” allosteric enzymes V. Relaxation kinetics of dissociating enzyme systems. Journal of Theoretical Biology, 98(1):73–90, September 7, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382900595.

Korzuhin:1982:GDP

[KK82b] M. D. Korzuhin and O. Z. Kaganova. Genotype dynamics in populations with breeding pairs. Journal of Theoretical Biology, 96(2):201–209, May 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902211.

Katsurai:1983:GPH

[KK83a] Tominosuke Katsurai and Toshitaka Kajino. Growth process of the human body considered in connection with basal metabolism and specific surface area. Journal of Theoretical Biology, 101(3):493–494, April 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383901522.

Kitatsuji:1983:SSR

[KK83b] Yasutsugu Kitatsuji and Tsukasa Kuroda. Statistical study on reverberation between two mutually excitatory neuron groups. Journal of Theoretical Biology, 100(1):25–55, January 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383900917.

Kothekar:1983:BSMa

[KK83c] V. Kothekar and S. Kailash. Biophysical studies on molecular mechanisms of abortificient action of prostaglandins v. CNDO/2 estimation of the relative affinities of the cations Na +, Mg^{2+} and Ca^{2+} to the carboxylic group. Journal of Theoretical Biology, 101(2):225–231, March 21, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383903363.

Kharkats:1985:MRE

[KK85] Yu. I. Kharkats and L. I. Krishtalik. Medium reorganization energy and enzymatic reaction activation energy. Journal of Theoretical Biology, 112(2):221–249, January 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385802848.

Klopman:1986:QSA

[KK86a] Gilles Klopman and Alexander N. Kalos. Quantitative structureactivity relationships of beta-adrenergic agents. Application of the computer automated structure evaluation (CASE) technique of molecular fragment recognition. Journal of Theoretical Biology, 118(2):199–214, January 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801345.

Koth:1986:OLT

[KK86b] Maria Koth and Fritz Kemler. A one locus-two allele selection model admitting stable limit cycles. Journal of Theoretical Biology, 122(3):263–267, October 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801199.

Kijima:1987:SACa

[KK87a] Shôko Kijima and Hiromasa Kijima. Statistical analysis of channel current from a membrane patch I. Some stochastic properties of ion channels or molecular systems in equilibrium. Journal of Theoretical Biology, 128(4):423–434, October 21, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387801881.

Kijima:1987:SACb

[KK87b] Shôko Kijima and Hiromasa Kijima. Statistical analysis of channel current from a membrane patch II. A stochastic theory of a multi-channel system in the steady-state. Journal of Theoretical Biology, 128(4):435–455, October 21, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387801893.

Krishnan:1981:EKR

[KKJ81] Leela Krishnan, Engikolai C. Krishnan, and William R. Jewell. Energy kinetics and role of nutrition in cancer patients. Journal of Theoretical Biology, 91(2):303–320, July 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902356.

Kamo:1980:TDS

[KKKK80] Naoki Kamo, Tadashi Kashiwagura, Kenzo Kurihara, and Yonosuke Kobatake. A theory of dynamic and steady responses in chemoreception. Journal of Theoretical Biology, 83(1):111–130, March 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519380903756.

Kijima:1980:COR

[KKNM80] H. Kijima, S. Kijima, M. Noguchi, and H. Morita. Cooperative response of chemically excitable membrane III. Threestate model. Journal of Theoretical Biology, 83(1):131–161, March 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519380903768.

Kuhnau:1980:ASE

[KL80] Reiner Kühnau and Jürgen Lasch. An analytical solution to the evolution of presteady state substrate and product concentration profiles in spherical enzyme supports. Journal of Theoretical Biology, 84(2):217–232, May 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519380800051.

Kirkwood:1981:EBT

[KL81] R. S. M. Kirkwood and J. H. Lawton. Efficiency of biomass transfer and the stability of model food-webs. Journal of Theoretical Biology, 93(1):225–237, November 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381900679.

Kohn:1983:GTA

[KL83] Michael C. Kohn and William J. Letzkus. A graph-theoretical analysis of metabolic regulation. Journal of Theoretical Biology, 100(2):293–304, January 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519383903533.

Kauffman:1987:TGT

[KL87] Stuart Kauffman and Simon Levin. Towards a general theory of adaptive walks on rugged landscapes. Journal of Theoretical Biology, 128(1):11–45, September 7, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387800292.

Klambt:1983:OPM

[Klä83] Dieter Klämbt. Oligopeptides and plant morphogenesis: a working hypothesis. Journal of Theoretical Biology, 100(3):435–441, February 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251938390440X.

Klein:1981:CTR

[Kle81] Benjamin Y. Klein. Can tumorigenesis result from in vivo “transfection” of a normal cell with DNA escaped from disintegrating nontransformed cells? Journal of Theoretical Biology, 90(2):199– 211, May 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381900436.

Kabanov:1988:EER

[KLK^{+}88] A. V. Kabanov, A. V. Levashov, N. L. Klyachko, S. N. Namyotkin, A. V. Pshezhetsky, and Karel Martinek. Enzymes entrapped in reversed micelles of surfactants in organic solvents: a theoretical treatment of the catalytic activity regulation. Journal of Theoretical Biology, 133(3):327–343, August 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388803254.

Kaitala:1989:FVR

[KLR89] Veijo Kaitala, Kai Lindström, and Esa Ranta. Foraging, vigilance and risk of predation in birds — a dynamic game study of ESS.

Journal of Theoretical Biology, 138(3):329–345, June 3, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389801985.

Kaina:1981:TDF

[KM81] Bernd Kaina and Armin Meister. On time dependency of frequency and distribution of chromatid aberrations induced by mutagens with delayed effects. An interpretation. Journal of Theoretical Biology, 88(3):485–501, February 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902782.

Kidd:1983:EER

[KM83] N. A. C. Kidd and A. D. Mayer. The effect of escape responses on the stability of insect host-parasite models. Journal of Theoretical Biology, 104(2):275–287, September 21, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383904150.

Koslov:1984:TOL

[KM84] M. M. Koslov and V. S. Markin. A theory of osmotic lysis of lipid vesicles. Journal of Theoretical Biology, 109(1):17–39, July 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384801083.

Klopman:1985:UCA

[KM85] Gilles Klopman and Orest T. Macina. Use of the computer automated structure evaluation program in determining quantitative structure-activity relationships within hallucinogenic phenylalkylamines. Journal of Theoretical Biology, 113(4):637–648, April 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385801843.

Kozlov:1987:MRB

[KM87] M. M. Kozlov and V. S. Markin. Model of red blood cell membrane skeleton: Electrical and mechanical properties. Journal of Theoretical Biology, 129(4):439–452, December 21, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387800231.

Kypr:1989:RFH

[KM89] Jaroslav Kypr and Jan Mrázek. Reading frames of HIV genes. Journal of Theoretical Biology, 141(3):423–424, December 7, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519389801237.

Kitney:1984:TSA

[KMSR84] R. I. Kitney, R. C. Miall, J. F. Stein, and P. M. Riddell. Time series analysis of neuronal signals recorded in the cerebellum of trained monkeys. Journal of Theoretical Biology, 107(3):367– 385, April 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384800971.

Kikuchi:1982:TSH

[KN82] Takeshi Kikuchi and Kichisuke Nishimoto. Theoretical studies of hemoproteins. II. Dynamical behaviour of ligand binding of allosteric proteins. Journal of Theoretical Biology, 97(3):491– 504, August 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382903770.

Kikuchi:1983:TSH

[KN83] Takeshi Kikuchi and Kichisuke Nishimoto. Theoretical studies of hemoproteins. III. Dynamical treatment of some abnormal hemoglobins. Journal of Theoretical Biology, 102(1):167– 173, May 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383902680.

Knopf:1982:HPC

[Kno82] U. C. Knopf. The herbicide for the plant concept as opposed to the plant for the herbicide concept. Journal of Theoretical Biology, 94(4):985–987, February 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382900911.

Knolle:1989:HDE

[Kno89] Helmut Knolle. Host density and the evolution of parasite virulence. Journal of Theoretical Biology, 136(2):199–207, January 23, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519389802267.

Katsurai:1982:GPI

[KO82] Tominosuke Katsurai and Shigeru Okamoto. Growth process and interfacial process considered in connection with specific surface area. Journal of Theoretical Biology, 96(3):511–516, June 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382901242.

Korn:1983:MTD

[KO83] Alex P. Korn and F. P. Ottensmeyer. A model for the three-dimensional structure of glucagon. Journal of Theoretical Biology, 105(3):403–425, December 7, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383901844.

Koch:1982:MKC

[Koc82] Arthur L. Koch. Multistep kinetics: Choice of models for the growth of bacteria. Journal of Theoretical Biology, 98(3):401– 417, October 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382901278.

Koch:1986:BSR

[Koc86a] Arthur L. Koch. The basis of synchronization by repetitive dilution of a growing culture. Journal of Theoretical Biology, 123(3):333–346, December 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386802491.

Koch:1986:PNM

[Koc86b] Arthur L. Koch. The pH in the neighborhood of membranes generating a protonmotive force. Journal of Theoretical Biology, 120(1):73–84, May 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519386800182.

Koch:1988:PAB

[Koc88] Arthur L. Koch. Partition of autolysins between the medium, the internal part of the wall, and the surface of the wall of grampositive rods. Journal of Theoretical Biology, 134(4):463–472, October 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519388800523.

Koch:1989:ORO

[Koc89] Arthur L. Koch. The origin of the rotation of one end of a cell relative to the other end during growth of Gram-positive rods. Journal of Theoretical Biology, 141(3):391–402, December 7, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519389801213.

Koehl:1983:MPS

[Koe83] M. A. R. Koehl. The morphology and performance of suspensionfeeding appendages. Journal of Theoretical Biology, 105(1):1–11, November 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383904204.

Kolber:1982:MDC

[Kol82] Michael A. Kolber. Mono- and divalent competitive adsorption to a charged membrane in a closed system: a comparative study. Journal of Theoretical Biology, 94(3):633–649, February 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382903046.

Kondrashov:1984:REC

[Kon84a] Alexey S. Kondrashov. Rate of evolution in a changing environment. Journal of Theoretical Biology, 107(2):249–260, March 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384800260.

Konopka:1984:ICD

[Kon84b] Andrzej Konopka. Is the information content of DNA evolutionarily significant? Journal of Theoretical Biology, 107(4):697– 704, April 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251938480140X.

Kondrashov:1988:DME

[Kon88] Alexey S. Kondrashov. Deleterious mutations as an evolutionary factor: III. Mating preference and some general remarks. Journal of Theoretical Biology, 131(4):487–496, April 21, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388800432.

Koopman:1985:AJE

[Koo85] James S. Koopman. Analyzing the joint effects of two antibodies and the design of molecularly engineered vaccines. Journal of Theoretical Biology, 116(4):569–585, October 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800886.

Kooijman:1986:EBC

[Koo86] S. A. L. M. Kooijman. Energy budgets can explain body size relations. Journal of Theoretical Biology, 121(3):269–282, August 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386801072.

Korn:1982:PSW

[Kor82] Robert W. Korn. Positional specificity within plant cells. Journal of Theoretical Biology, 95(3):543–568, April 7, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382900340.

Korona:1989:IFD

[Kor89] Ryszard Korona. Ideal free distribution of unequal competitors can be determined by the form of competition. Journal of Theoretical Biology, 138(3):347–352, June 3, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389801997.

Kothekar:1982:BSM

[Kot82] V. Kothekar. Biophysical studies on molecular mechanism of abortificient action of prostaglandins-IV. conformation energy calculations on PGA 1, PGB 1 and PGE 1. Journal of Theoretical Biology, 94(4):943–949, February 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382900881.

Kothekar:1983:BSMb

[Kot83] V. Kothekar. Biophysical studies on molecular mechanism of abortificient action of prostaglands VI. conformation energy calculation on PGE 2, PGF 2α and 15-(s)-methyl PGF 2α. Journal of Theoretical Biology, 101(2):233–240, March 21, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383903375.

Kristofferson:1981:QTR

[KP81] David Kristofferson and Daniel L. Purich. A quantitative treatment of reversible monomer-polymer exchange reactions with microtubules and other biopolymers. Journal of Theoretical Biology, 92(1):85–96, September 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381901612.

Koch:1985:BPS

[KP85] Christof Koch and Tomaso Poggio. The biophysical properties of spines as a basis for their electrical function: a comment on kawato and tsukahara (1983). Journal of Theoretical Biology, 113(2):225–229, March 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385802253.

Keasling:1989:CPR

[KP89a] J. D. Keasling and B. O. Palsson. ColE1 plasmid replication: a simple kinetic description from a structured model. Journal of Theoretical Biology, 141(4):447–461, December 19, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389802309.

Keasling:1989:KPR

[KP89b] J. D. Keasling and B. O. Palsson. On the kinetics of plasmid replication. Journal of Theoretical Biology, 136(4):487–492, February 22, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519389801614.

Kindlmann:1982:CVD

[KR82] P. Kindlmann and M. Rejmék. Continuous vs. discrete models of multispecies systems: How much less stable are the latter ones? Journal of Theoretical Biology, 94(4):989–993, February 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382900923.

Kesavan:1985:FSM

[KR85a] Sunil K. Kesavan and Narender P. Reddy. On the feeding strategy and the mechanics of blood sucking in insects. Journal of Theoretical Biology, 113(4):781–783, April 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801934.

Knoppien:1985:PTM

[KR85b] P. Knoppien and J. Reddingius. Predators with two modes of searching: a mathematical model. Journal of Theoretical Biology, 114(2):273–301, May 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801077.

Krieber:1986:MPM

[KR86] Michel Krieber and Michael R. Rose. Males, parthenogenesis, and the maintenance of anisogamous sex. Journal of Theoretical Biology, 122(4):421–440, October 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801837.

Kramer:1982:AST

[Kra82] Lorenz Kramer. Absolute stability and transitions in ecosystems with a multiplicity of stable states and dispersal. Journal of Theoretical Biology, 98(1):91–108, September 7, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382900601.

King:1981:CTD

[KRB81] Roy King, Joachim D. Raese, and Jack D. Barchas. Catastrophe theory of dopaminergic transmission: a revised dopamine hypothesis of schizophrenia. Journal of Theoretical Biology, 92(4):373– 400, October 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381902551.

Kremen:1982:IAG

[Kre82] Alexandr Kremen. Informational aspects of Gibbs function output from nucleotide pools and of the adenylate kinase reaction. Journal of Theoretical Biology, 96(3):425–441, June 7, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382901199.

Kremen:1984:LEL

[Kre84] Alexandr Kremen. Logical elements in living cells. Journal of Theoretical Biology, 111(1):47–60, November 7, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801952.

Krishtalik:1980:CAR

[Kri80] L. I. Krishtalik. Catalytic acceleration of reactions by enzymes. Effect of screening of a polar medium by a protein globule. Journal of Theoretical Biology, 86(4):757–771, October 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380903094.

Krishtalik:1985:CMP

[Kri85a] L. I. Krishtalik. Conformational mobility of proteins and activation energy of enzymatic charge transfer reactions:

An analysis in dielectric formalism. Journal of Theoretical Biology, 116(2):201–214, September 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385802642.

Krishtalik:1985:EAE

[Kri85b] L. I. Krishtalik. Effective activation energy of enzymatic and nonenzymatic reactions. Evolution-imposed requirements to enzyme structure. Journal of Theoretical Biology, 112(2):251–264, January 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251938580285X.

Krishtalik:1989:DCC

[Kri89] L. I. Krishtalik. Dielectric constant in calculations of the electrostatics of biopolymers. Journal of Theoretical Biology, 139(2):143–154, July 21, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519389800979.

Kananyan:1981:EMLa

[KRT81a] G. Kh. Kananyan, V. A. Ratner, and R. N. Tchuraev. Enlarged model of lambda phage ontogenesis. Journal of Theoretical Biology, 88(3):393–407, February 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902733.

Kananyan:1981:EMLb

[KRT81b] G. Kh. Kananyan, V. A. Ratner, and R. N. Tchuraev. Enlarged model of lambda phage ontogenesis. Journal of Theoretical Biology, 90(2):301–315, May 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381900503.

Katz:1980:MIP

[KS80a] Sam Katz and William Squire. Metal ion-protein complex formation: Buffer competitive effects. Journal of Theoretical Biology, 82(3):405–414, February 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380902465.

Kawato:1980:TCN

[KS80b] Mitsuo Kawato and Ryoji Suzuki. Two coupled neural oscillators as a model of the circadian pacemaker. Journal of Theoretical Biology, 86(3):547–575, October 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380903525.

Kothekar:1981:BSMa

[KS81] V. Kothekar and A. K. Srivastava. Biophysical studies on molecular mechanism of abortificient action of prostaglandins I. The study of molecular electrostatic potential distribution of PGF_{2α}, PGF_{1β} and PGA_{1}. Journal of Theoretical Biology, 93(1):7–23, November 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381900539.

Koushanpour:1982:MSB

[KS82] Esmail Koushanpour and Gordon K. Stipp. Mathematical simulation of the body fluid regulating system in dog. Journal of Theoretical Biology, 99(2):203–235, November 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382900017.

Khait:1984:MEP

[KS84] A. Khait and L. A. Segel. A model for the establishment of pattern by positional differentiation with memory. Journal of Theoretical Biology, 110(2):135–153, September 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384800491.

Klein:1988:NMD

[KS88] Peter Klein and R. L. Somorjai. Nonlinear methods for discrimination and their application to classification of protein structures. Journal of Theoretical Biology, 130(4):461–468, February 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388802108.

Khanin:1989:MMK

[KS89] M. A. Khanin and V. V. Semenov. A mathematical model of the kinetics of blood coagulation. Journal of Theoretical Biology, 136(2):127–134, January 23, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389802206.

Katz:1985:TSA

[KsC85] Michael J. Katz and Yun shyong Chow. Templeting and selfassembly. Journal of Theoretical Biology, 113(1):1–13, March 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385800722.

Kretschmer:1987:ACI

[KSH87] Matthias Kretschmer, Wolfgang Schellenberger, and Eberhard Hofmann. The appearance of a critical input concentration of fructose 6-phosphate in the 6-phosphofructo-2-kinase/fructose 2,6-bisphosphatase cycle. Journal of Theoretical Biology, 127(2):181–186, July 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387801297.

Kliss:1989:SVS

[KSL89] M. Kliss, C. Somps, and M. W. Luttges. Stable vortex structures: a flat plate model of dragonfly hovering. Journal of Theoretical Biology, 136(2):209–228, January 23, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389802279.

Kurganov:1985:SOG

[KSM85] B. I. Kurganov, N. P. Sugrobova, and L. S. Mil’man. Supramolecular organization of glycolytic enzymes. Journal of Theoretical Biology, 116(4):509–526, October 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800862.

Kemeny:1981:SFI

[KSS81] Gabor Kemeny, Joseph P. Straley, and Katherine A. Strong. Single file ion diffusion and the schmoluchowski equation. Journal of Theoretical Biology, 89(2):231–247, March 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381903106.

Kirber:1985:PFD

[KSW^{+}85] Michael T. Kirber, Joshua J. Singer, John V. Walsh, Marc S. Fuller, and Robert A. Peura. Possible forms for dwell-time histograms from single-channel current records. Journal of Theoretical Biology, 116(1):111–126, September 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801338.

Kuba:1981:SIC

[KT81] Kenji Kuba and Setsuko Takeshita. Simulation of intracellular Ca^{2+} oscillation in a sympathetic neurone. Journal of Theoretical Biology, 93(4):1009–1031, December 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381903520.

Kawato:1983:TSE

[KT83] Mitsuo Kawato and Nakaakira Tsukahara. Theoretical study on electrical properties of dendritic spines. Journal of Theoretical Biology, 103(4):507–522, August 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383902801.

Kaufman:1987:MAB

[KT87a] M. Kaufman and R. Thomas. Model analysis of the bases of multistationarity in the humoral immune response. Journal of Theoretical Biology, 129(2):141–162, November 21, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387800097.

Kitney:1987:ZWT

[KT87b] R. I. Kitney and H. Talhami. The Zoom Wigner transform and its application to the analysis of blood velocity waveforms. Journal of Theoretical Biology, 129(4):395–409, December 21, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387800206.

Kitney:1986:ABV

[KTG86] R. I. Kitney, H. Talhami, and D. P. Giddens. The analysis of blood velocity measurements by autoregressive modelling. Journal of Theoretical Biology, 120(4):419–442, June 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386800364.

Kubota:1981:HPS

[KTNO81] Yasushi Kubota, Sho Takahashi, Ken Nishikawa, and Tatsuo Ooi. Homology in protein sequences expressed by correlation coefficients. Journal of Theoretical Biology, 91(2):347–361, July 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251938190237X.

Kawakubo:1989:PDT

[KTS89] Tatsuyuki Kawakubo, Yoshimi Tsuchiya, and Hitoshi Suzuki. Population-dependent taxis of protozoa. Journal of Theoretical Biology, 139(3):405–411, August 9, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389802188.

Kubo:1985:VBA

[Kub85] Kazumi Kubo. A view on the break in the Arrhenius plots. Journal of Theoretical Biology, 115(4):551–569, August 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801405.

Kuehn:1981:IAB

[Kue81] L. A. Kuehn. Implications of athlete’s bradycardia on lifespan. Journal of Theoretical Biology, 88(2):279–286, January 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251938190076X.

Kuhn:1982:PSI

[Kuh82] Ekkehard Kuhn. Programmed synthesis of informational molecules monitored by linear and circular genetic control chains.

Journal of Theoretical Biology, 95(3):423–464, April 7, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382900297.

Kumar:1982:MMA

[Kum82] Suriender Kumar. A model for the mechanism of action of animal fatty acid synthetases. Journal of Theoretical Biology, 95(2):263– 283, March 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382902442.

Kurganov:1983:SLI

[Kur83] B. I. Kurganov. Specific ligand-induced association of an enzyme. A new model of dissociating allosteric enzyme. Journal of Theoretical Biology, 103(2):227–245, July 21, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383900267.

Kurganov:1984:APE

[Kur84a] B. I. Kurganov. Adsorption of peripheral enzymes to membrane anchor proteins. Journal of Theoretical Biology, 111(4):707–723, December 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384802635.

Kurganov:1984:BSL

[Kur84b] B. I. Kurganov. Binding of specific ligand by linearly associating enzyme system. Journal of Theoretical Biology, 109(1):59– 70, July 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384801101.

Kurganov:1986:RMC

[Kur86] B. I. Kurganov. The role of multienzyme complexes in integration of cellular metabolism. Journal of Theoretical Biology, 119(4):445–455, April 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519386801941.

Kaufman:1985:TLA

[KUT85] M. Kaufman, J. Urbain, and R. Thomas. Towards a logical analysis of the immune response. Journal of Theoretical Biology, 114(4):527–561, June 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800424.

Kudzmauskas:1983:TET

[KVB83] S. Kudzmauskas, L. Valkunas, and A. Y. Borisov. A theory of excitation transfer in photosynthetic units. Journal of Theoretical Biology, 105(1):13–23, November 7, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383904216.

Konings:1987:MEF

[KvDVH87] D. A. M. Konings, L. P. van Duijn, H. O. Voorma, and P. Hogeweg. Minimal energy foldings of eukaryotic mRNAs form a separate leader domain. Journal of Theoretical Biology, 127(1): 63–78, July 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387801613.

Keleti:1988:PSC

[KVW88] Tamás Keleti, Beáta Vértessy, and G. Rickey Welch. The perfection of substrate-channelling in interacting enzyme systems: Energetics and evolution. Journal of Theoretical Biology, 135(1):75– 83, November 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519388801759.

Kauffman:1986:NEa

[KW86a] S. A. Kauffman and L. Wolpert. Note from the editors. Journal of Theoretical Biology, 118(1):1, January 7, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386800030.

Kauffman:1986:NEb

[KW86b] S. A. Kauffman and L. Wolpert. Note from the editors. Journal of Theoretical Biology, 118(2):127, January 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938680128X.

Kauffman:1986:NEc

[KW86c] S. A. Kauffman and L. Wolpert. Note from the editors. Journal of Theoretical Biology, 118(3):251, February 7, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938680056X.

Kauffman:1986:NEd

[KW86d] S. A. Kauffman and L. Wolpert. Note from the editors. Journal of Theoretical Biology, 118(4):377, February 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938680159X.

Kauffman:1989:NMR

[KW89] Stuart A. Kauffman and Edward D. Weinberger. The NK model of rugged fitness landscapes and its application to maturation of the immune response. Journal of Theoretical Biology, 141(2):211–245, November 21, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389800190.

Koppes:1987:PSS

[KWG87] L. J. Koppes, C. L. Woldringh, and N. B. Grover. Predicted steady-state cell size distributions for various growth models. Journal of Theoretical Biology, 129(3):325–335, December 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938780005X.

Kwok:1984:CAA

[Kwo84] L. Stephen Kwok. Calculation and application of the anterior surface area of a model human cornea. Journal of Theoretical Biology, 108(2):295–313, May 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384800727.

Kwok:1989:SAE

[Kwo89] L. Stephen Kwok. The surface area of an ellipsoid revisited. Journal of Theoretical Biology, 139(4):573–574, August 22, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389800712.

Kypr:1988:PRP

[Kyp88] Jaroslav Kypr. Possible reason for the preferential insertion of adenine opposite abasic lesions in DNA. Journal of Theoretical Biology, 135(1):125–126, November 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388801802.

Kawato:1986:SAE

[KYU^{+}86] Mitsuo Kawato, Atsushi Yamanaka, Seiichi Urushibara, Osamu Nagata, Hiroshi Irisawa, and Ryoji Suzuki. Simulation analysis of excitation conduction in the heart: Propagation of excitation in different tissues. Journal of Theoretical Biology, 120(4):389– 409, June 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519386800340.

Lefebvre:1982:KGB

[LA82] Gerald M. Lefebvre and Adel F. Antippa. The kinetics of germination in bacterial spores. Journal of Theoretical Biology, 95(3):489–515, April 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382900315.

Lande:1985:EMP

[LA85] Russell Lande and Stevan J. Arnold. Evolution of mating preference and sexual dimorphism. Journal of Theoretical Biology, 117(4):651–664, December 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385802459.

Lundberg:1986:BMP

[LA86] Sigfrid Lundberg and Thomas Alerstam. Bird migration patterns: Conditions for stable geographical population segregation. Journal of Theoretical Biology, 123(4):403–414, December 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386802107.

Lacy:1984:PEC

[Lac84] Mark E. Lacy. Phonon-electron coupling as a possible transducing mechanism in bioelectronic processes involving neuromelanin. Journal of Theoretical Biology, 111(1):201–204, November 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384802052.

Lambert:1984:SMR

[Lam84a] D. M. Lambert. Specific-mate recognition systems, phylogenies and asymmetrical evolution. Journal of Theoretical Biology, 109(1):147–156, July 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801162.

Lambert:1984:EEM

[Lam84b] Grant R. Lambert. Enzymic editing mechanisms and the origin of biological information transfer. Journal of Theoretical Biology, 107(3):387–403, April 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384800983.

Lam:1988:CEM

[Lam88] Gabriel K. Y. Lam. The combined effects of mixtures of ionizing radiations. Journal of Theoretical Biology, 134(4):531–546, October 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519388800559.

Lane:1982:DFS

[Lan82a] Andrew N. Lane. The dependence of flux, sensitivity and response of the flux on the concentrations of substrate and modifiers for cooperative enzymes. Journal of Theoretical Biology, 99(3):491–508, December 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382902089.

Lane:1982:GMP

[Lan82b] Andrew N. Lane. Generation of models of protein structure from solution scattering experiments. Journal of Theoretical Biology, 97(3):511–527, August 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382903794.

Landman:1984:CMR

[Lan84a] Kerry A. Landman. A continuum model for a red blood cell transformation: Sphere to crenated sphere. Journal of Theoretical Biology, 106(3):329–351, February 7, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519384900341.

Lansky:1984:ASN

[Lán84b] Petr Lánský. On approximations of Stein’s neuronal model. Journal of Theoretical Biology, 107(4):631–647, April 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801368.

Landman:1985:CMS

[Lan85] Kerry A. Landman. A continuum model for surface contraction waves. Journal of Theoretical Biology, 115(1):9–25, July 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385800047.

Lapidus:1980:MCF

[Lap80a] I. Richard Lapidus. Microbial chemotaxis in flowing water in the vicinity of a source of attractant or repellent. Journal of Theoretical Biology, 85(3):543–547, August 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380903264.

Lapidus:1980:PMO

[Lap80b] I. Richard Lapidus. “pseudochemotaxis” by micro-organisms in an attractant gradient. Journal of Theoretical Biology, 86 (1):91–103, September 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519380900673.

Lapidus:1981:BMI

[Lap81a] I. Richard Lapidus. Behavior of microorganisms and internal state variables. Journal of Theoretical Biology, 92(4):359–362, October 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381902526.

Lapidus:1981:CRM

[Lap81b] I. Richard Lapidus. Chemosensory responses of microorganisms: Asymmetric diffusion as the limit of a biased random walk. Journal of Theoretical Biology, 92(4):345–358, October 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902514.

Lapidus:1984:GDK

[Lap84] I. Richard Lapidus. Growth and division kinetics of asymmetrically dividing tetrahymena thermophila. Journal of Theoretical Biology, 106(2):135–140, January 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519384900146.

Lasic:1987:GMV

[Las87] Danilo D. Lasic. A general model of vesicle formation. Journal of Theoretical Biology, 124(1):35–41, January 7, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387802503.

Latimer:1983:PAC

[Lat83] Paul Latimer. Photometric assays of cell shrinkage-the resolution of a conflict. Journal of Theoretical Biology, 102(2):249–259, May 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383903624.

Lauder:1982:HBP

[Lau82] George V. Lauder. Historical biology and the problem of design. Journal of Theoretical Biology, 97(1):57–67, July 7, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902764.

Lavett:1984:MTR

[Lav84] Diane K. Lavett. A model for transcriptional regulation based upon homology between introns and promoter regions and secondary structure in the promoter regions: The human globins. Journal of Theoretical Biology, 107(1):1–36, March 7, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801198.

Laws:1983:PTT

[Law83] Edward A. Laws. Plots of turnover times versus added substrate concentrations provide only upper bounds to in situ substrate concentrations. Journal of Theoretical Biology, 101(1):147– 150, March 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383902783.

Laws:1984:IEP

[Law84] Edward A. Laws. Improved estimates of phytoplankton carbon based on ^{14}C incorporation into chlorophyll a. Journal of Theoretical Biology, 110(3):425–434, October 7, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801848.

Layton:1987:EAC

[Lay87] H. E. Layton. Energy advantage of counter-current oxygen transfer in fish gills. Journal of Theoretical Biology, 125(3):307– 316, April 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387800620.

Liljenstrom:1987:SDT

[LB87] Hans Liljenström and Clas Blomberg. Site dependent time optimization of protein synthesis with special regard to accuracy. Journal of Theoretical Biology, 129(1):41–56, November 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519387802023.

Leps:1984:SSB

[LBBL84] Bernd Leps, Gerhard Barnickel, Hans Bradaczek, and Harald Labischinski. Structural studies on the bacterial cell wall peptidoglycan pseudomurein. I. Conformational energy calculations on the glycan strands in C1 conformation and comparison with murein. Journal of Theoretical Biology, 107(1):85–114, March 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938480123X.

Liden:1986:VDT

[LBW86] Carola Lidén, Ulf Bergqvist, and Göran Wennersten. Visual display terminals do not emit UV — a radiation of clinical relevance. Journal of Theoretical Biology, 122(4):491–492, October 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386801886.

Lavietes:1980:RLM

[LC80] Beverly B. Lavietes and Peter S. Coleman. The role of lipid metabolism in neoplastic differentiation. Journal of Theoretical Biology, 85(3):523–542, August 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380903252.

Lenarz:1982:RGB

[LC82] Mark S. Lenarz and Walt Conley. Reproductive gambling in bighorn sheep (ovis): a simulation. Journal of Theoretical Biology, 98(1):1–7, September 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382900522.

Lara:1984:GMN

[LCDC84] Rolando Lara, Manuel Carmona, Federico Daza, and Armando Cruz. A global model of the neural mechanisms responsible for visuomotor coordination in toads. Journal of Theoretical Biology, 110(4):587–618, October 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801472.

Lara:1983:SCD

[LD83a] Rolando Lara and Gonzalo Viana Di Prisco. Synaptic competition in developmental binocular plasticity. Journal of Theoretical Biology, 101(4):503–527, April 21, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383900139.

Leclerco:1983:BAC

[LD83b] J. Leclerco and J. E. Dumont. Boolean analysis of cell regulation networks. Journal of Theoretical Biology, 104(4):507–534, October 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383902436.

Leunissen:1986:PTC

[LdJ86] Jack A. M. Leunissen and Wilfried W. de Jong. Phylogenetic trees constructed from hydrophobicity values of protein sequences. Journal of Theoretical Biology, 119(2):189–196, March 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938680073X.

Leimar:1984:EAO

[LE84a] Olof Leimar and Magnus Enquist. Effects of asymmetries in owner-intruder conflicts. Journal of Theoretical Biology, 111 (3):475–491, December 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519384802350.

Lundstrom:1984:SBC

[LE84b] Ingemar Lundström and Hans Elwing. Stereospecific binding capacity of proteins on surfaces — simple mathematical models. Journal of Theoretical Biology, 110(2):195–204, September 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384800521.

Lefevre:1983:TOF

[Lef83] Jacques Lefèvre. Teleonomical optimization of a fractal model of the pulmonary arterial bed. Journal of Theoretical Biology, 102(2):225–248, May 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383903612.

Leff:1986:PEA

[Lef86] P. Leff. Potential errors in agonist dissociation constant estimation caused by desensitization. Journal of Theoretical Biology, 121(2):221–232, July 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386800959.

Legovic:1987:DPF

[Leg87] T. Legović. Determination of parameters for food web models from field observations. Journal of Theoretical Biology, 129 (2):211–218, November 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519387800139.

Leigh:1981:ALP

[Lei81] Egbert Giles Leigh. The average lifetime of a population in a varying environment. Journal of Theoretical Biology, 90(2):213– 239, May 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381900448.

Lenski:1984:CBP

[Len84] Richard E. Lenski. Coevolution of bacteria and phage: Are there endless cycles of bacterial defenses and phage counterdefenses? Journal of Theoretical Biology, 108(3):319–325, June 7, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384800351.

Leonard:1984:TAA

[Leo84] Janet L. Leonard. A top-down approach to the analysis of behavioural organization. Journal of Theoretical Biology, 107(3):457–470, April 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519384801022.

Lerche:1983:BMI

[Ler83] Dietmar Lerche. A biophysical model for interaction of cells with a surface coat (glycocalyx). I. Electrostatic interaction profile. Journal of Theoretical Biology, 104(2):231–248, September 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383904125.

Lessard:1986:EPG

[Les86] Sabin Lessard. Evolutionary principles for general frequencydependent two-phenotype models in sexual populations. Journal of Theoretical Biology, 119(3):329–344, April 7, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938680145X.

Leuchtag:1987:IEF

[Leu87a] H. Richard Leuchtag. Indications of the existence of ferroelectric units in excitable-membrane channels. Journal of Theoretical Biology, 127(3):321–340, August 7, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387801108.

Leuchtag:1987:PTI

[Leu87b] H. Richard Leuchtag. Phase transitions and ion currents in a model ferroelectric channel unit. Journal of Theoretical Biology, 127(3):341–359, August 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938780111X.

Levine:1980:SMT

[Lev80a] Stephen Levine. Several measures of trophic structure applicable to complex food webs. Journal of Theoretical Biology, 83(2):195– 207, March 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251938090288X.

Levitan:1980:EIC

[Lev80b] E. Levitan. Estimation of ionic channel surface density from relaxational ionic and gating currents. Journal of Theoretical Biology, 85(3):423–427, August 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380903173.

Lewis:1981:CPHa

[Lew81a] Jerry W. Lewis. On the coevolution of pathogen and host: I. General theory of discrete time coevolution. Journal of Theoretical Biology, 93(4):927–951, December 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381903489.

Lewis:1981:CPHb

[Lew81b] Jerry W. Lewis. On the coevolution of pathogen and host: II. Selfing hosts and haploid pathogens. Journal of Theoretical Biology, 93(4):953–985, December 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381903490.

Lewis:1981:SRE

[Lew81c] Julian Lewis. Simpler rules for epimorphic regeneration: The polar-coordinate model without polar coordinates. Journal of Theoretical Biology, 88(2):371–392, January 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381900825.

Lewis:1982:TDP

[Lew82] Allen R. Lewis. Tree density, pollen-ovule ratio, and sex ratio in dioecious wind pollinated trees. Journal of Theoretical Biology, 97(2):289–298, July 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382901059.

Liebovitch:1986:MPC

[LF86] Larry S. Liebovitch and Jorge Fischbarg. Membrane pores: a computer simulation of interacting pores analyzed by g_{1}(τ) and g_{2}(τ) correlation functions. Journal of Theoretical Biology, 119(3):287–297, April 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519386801424.

Lisanti:1984:CLR

[LFP84] Michael P. Lisanti, Mark Flanagan, and Saul Puszkin. Clathrin lattice reorganization: theoretical considerations. Journal of Theoretical Biology, 108(1):143–157, May 7, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801757.

Lucas:1985:PPC

[LG85] Jeffrey R. Lucas and Alan Grafen. Partial prey consumption by ambush predators. Journal of Theoretical Biology, 113(3):455– 473, April 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385800333.

Li:1989:OIS

[LG89a] Yue-Xian Li and Albert Goldbeter. Oscillatory isozymes as the simplest model for coupled biochemical oscillators. Journal of Theoretical Biology, 138(2):149–174, May 22, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389801365.

Liquori:1989:TEA

[LG89b] Alfonso M. Liquori and Silvio Giove. The temporal evolution of AIDS manifestation in a group of HIV virus infected individuals. Journal of Theoretical Biology, 140(4):571–574, October 23, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519389801158.

Lauffenburger:1984:HAS

[LGK84] Douglas Lauffenburger, Mary Grady, and Kenneth H. Keller. An hypothesis for approaching swarms of myxobacteria. Journal of Theoretical Biology, 110(2):257–274, September 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384800582.

Lyko:1980:ACN

[LH80] Henry C. Lyko and James X. Hartmann. Ascorbate, cyclic nucleotides, citrus and a model for preventing large bowel cancer. Journal of Theoretical Biology, 83(4):675–686, April 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380901952.

Lumsden:1983:RCI

[LH83] Charles J. Lumsden and Bert Hölldobler. Ritualized combat and intercolony communication in ants. Journal of Theoretical Biology, 100(1):81–98, January 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383900930.

Lenski:1986:CTC

[LH86] Richard E. Lenski and Susan E. Hattingh. Coexistence of two competitors on one resource and one inhibitor: a chemostat model based on bacteria and antibiotics. Journal of Theoretical Biology, 122(1):83–93, September 7, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386802260.

Lambert:1988:KCS

[LH88] D. M. Lambert and A. J. Hughes. Keywords and concepts in structuralist and functionalist biology. Journal of Theoretical Biology, 133(2):133–145, July 21, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388800018.

Leff:1989:DPM

[LH89] P. Leff and D. Harper. Do pharmacological methods for the quantification of agonists work when the ternary complex mechanism operates? Journal of Theoretical Biology, 140(3):381–397, October 9, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519389800943.

Li:1983:HSM

[Li83] John K-J. Li. Hemodynamic significance of metabolic turn-over rate. Journal of Theoretical Biology, 103(3):333–338, August 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383902898.

Li:1986:CCM

[Li86] John K-J. Li. Comparative cardiac mechanics: Laplace’s law. Journal of Theoretical Biology, 118(3):339–343, February 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386800649.

Li:1988:LTF

[Li88] John K-J. Li. Laminar and turbulent flow in the mammalian aorta: Reynolds number. Journal of Theoretical Biology, 135(3):409–414, December 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388802558.

Liboff:1980:NT

[Lib80] Richard L. Liboff. Neuromagnetic thresholds. Journal of Theoretical Biology, 83(3):427–436, April 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380900491.

Libet:1985:SAS

[Lib85] Benjamin Libet. Subjective antedating of a sensory experience and mind-brain theories: Reply to honderich (1984). Journal of Theoretical Biology, 114(4):563–570, June 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800436.

Libertini:1988:ATI

[Lib88] Giacinto Libertini. An adaptive theory of the increasing mortality with increasing chronological age in populations in the wild. Journal of Theoretical Biology, 132(2):145–162, May 21, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938880153X.

Lillehoj:1982:EBE

[Lil82] E. B. Lillehoj. Evolutionary basis and ecological role of toxic microbial secondary metabolites. Journal of Theoretical Biology, 97(2):325–332, July 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382901084.

Lima:1987:VWF

[Lim87] Steven L. Lima. Vigilance while feeding and its relation to the risk of predation. Journal of Theoretical Biology, 124(3):303–316, February 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387801182.

Lima:1988:VDS

[Lim88] Steven L. Lima. Vigilance and diet selection: The classical diet model reconsidered. Journal of Theoretical Biology, 132(2):127– 143, May 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519388801528.

Linhart:1980:FTC

[Lin80] J. G. Linhart. Forgetting time in coupled, parallel channels. Journal of Theoretical Biology, 86(2):405–408, September 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380900156.

Lavigne:1986:MRB

[LIWK86] D. M. Lavigne, S. Innes, G. A. J. Worthy, and K. M. Kovacs. Metabolic rate — body size relations in marine mammals. Journal of Theoretical Biology, 122(1):123–124, September 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386802284.

Lehman:1988:GCD

[LJ88] Niles Lehman and Thomas H. Jukes. Genetic code development by stop codon takeover. Journal of Theoretical Biology, 135(2):203–214, November 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388800742.

Laws:1985:MRM

[LJTH85] Edward A. Laws, David R. Jones, Kenneth L. Terry, and Janice A. Hirata. Modifications in recent models of phytoplankton growth: Theoretical developments and experimental examination of predictions. Journal of Theoretical Biology, 114(2):323– 341, May 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385801090.

Linkens:1983:IRE

[LK83] D. A. Linkens and R. I. Kitney. Integer-ratio entrainment in mutually-coupled non-linear oscillators. Journal of Theoretical Biology, 101(4):599–617, April 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383900188.

Lande:1988:ESS

[LK88] Russell Lande and Mark Kirkpatrick. Ecological speciation by sexual selection. Journal of Theoretical Biology, 133(1):85– 98, July 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519388800262.

Leikin:1987:MFO

[LKC^{+}87] Sergey L. Leikin, Michael M. Kozlov, Leonid V. Chernomordik, Vladislav S. Markin, and Yuri A. Chizmadzhev. Membrane fusion: Overcoming of the hydration barrier and local restructuring. Journal of Theoretical Biology, 129(4):411–425, December 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519387800218.

Lenhoff:1982:EAD

[LL82] A. M. Lenhoff and E. N. Lightfoot. The effects of axial diffusion and permeability barriers on the transient response of tissue cylinders. I. Solution in transform space. Journal of Theoretical Biology, 97(4):663–677, August 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382903654.

Lenhoff:1984:EAD

[LL84] A. M. Lenhoff and E. N. Lightfoot. The effects of axial diffusion and permeability barriers on the transient response of tissue cylinders. II. Solution in time domain. Journal of Theoretical Biology, 106(2):207–238, January 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519384900201.

Liao:1987:EQS

[LL87a] J. C. Liao and E. N. Lightfoot. Extending the quasi-steady state concept to analysis of metabolic networks. Journal of Theoretical Biology, 126(3):253–273, June 7, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387802345.

Lundberg:1987:ISC

[LL87b] Sigfrid Lundberg and Christer Löfstedt. Intra-specific competition in the sex communication channel: a selective force in the evolution of moth pheromones? Journal of Theoretical Biology, 125(1):15–24, March 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387801765.

Lyster:1987:CSR

[LL87c] W. R. Lyster and M. M. Lloyd. Contrasting sex ratios in siblings of male and female left-handed persons. Journal of Theoretical Biology, 126(1):125–126, May 7, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387801054.

Linderman:1988:AIR

[LL88] Jennifer J. Linderman and Douglas A. Lauffenburger. Analysis of intracellular receptor/ligand sorting in endosomes. Journal of Theoretical Biology, 132(2):203–245, May 21, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388801577.

Lloyd:1983:ESS

[Llo83] David G. Lloyd. Evolutionarily stable sex ratios and sex allocations. Journal of Theoretical Biology, 105(3):525–539, December 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383901911.

Lappin:1980:PTM

[LM80] Sylvia A. Lappin and F. H. Malpress. Proton turnover at mitochondrial coupling sites. Journal of Theoretical Biology, 84(2):181–184, May 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519380800014.

Lundstrom:1982:EMM

[LM82] Ingemar Lundström and Douglas H. McQueen. An electrostatic model for mechano-electrical transduction, especially in the ear. Journal of Theoretical Biology, 97(4):611–623, August 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382903629.

Lacey:1985:GCC

[LM85] James C. Lacey and Dail W. Mullins. Genetic coding catalysis. Journal of Theoretical Biology, 115(4):595–601, August 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385801429.

Le:1989:MAS

[LM89] Shu-Yun Le and Jacob V. Maizel. A method for assessing the statistical significance of RNA folding. Journal of Theoretical Biology, 138(4):495–510, June 22, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389800475.

Lundstrom:1981:PMR

[LN81] Ingemar Lundstrom and Class Nylander. A possible model for receptors in excitable membranes. Journal of Theoretical Biology, 88(4):671–683, February 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902435.

Lennon:1985:EOF

[LN85] Gregory G. Lennon and Ruth Nussinov. Eukaryotic oligomer frequencies are correlated with certain DNA helical parameters. Journal of Theoretical Biology, 116(3):427–433, October 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385802794.

Larter:1981:TBS

[LO81] Raima Larter and Peter Ortoleva. A theoretical basis for selfelectrophoresis. Journal of Theoretical Biology, 88(4):599–630, February 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381902411.

Larter:1982:SIE

[LO82] Raima Larter and Peter Ortoleva. A study of instability to electrical symmetry breaking in unicellular systems. Journal of Theoretical Biology, 96(2):175–200, May 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251938290220X.

Lockard:1982:BBB

[Loc82] Isabel Lockard. The blood-brain barrier: an alternative hypothesis. Journal of Theoretical Biology, 97(2):167–176, July 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382900972.

Levitan:1982:EVT

[LP82] E. Levitan and Y. Palti. The equivalent valency and the transition dipole moment in the interpretation of the n, m and h parameters in the Hodgkin–Huxley model. Journal of Theoretical Biology, 95(2):413–414, March 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902569.

Levitan:1983:TCM

[LP83] E. Levitan and Y. Palti. On the thermodynamic characterization of membrane gating particles by their Boltzmann distribution. Journal of Theoretical Biology, 100(1):107–111, January 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383900954.

Lokajicek:1986:MDS

[LP86] M. Lokajícek and I. Pridal. Modelling of DNA sedimentation profiles from irradiated eukaryotic cells. Journal of Theoretical Biology, 121(1):121–127, July 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386800327.

Laredo:1988:MPM

[LP88a] C. Laredo and J. Pernes. Models for pearl millet domestication as an example of cereal domestication. I. A one locus asymmetrical model. Journal of Theoretical Biology, 131(3):289–305, April 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388802273.

Lawaczeck:1988:PWT

[LP88b] Rüdiger Lawaczeck and Horst Pitterich. Pathways of water through erythrocyte membranes. Routes along defect structures. Journal of Theoretical Biology, 135(3):401–407, December 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388802546.

Lustig:1986:QHN

[LPS86] Cornel Lustig, Hanna Parnas, and Lee A. Segel. On the quantal hypothesis of neurotransmitter release: an explanation for the calcium dependence of the binomial parameters. Journal of Theoretical Biology, 120(2):205–213, May 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801746.

Lustig:1989:NRD

[LPS89] Cornel Lustig, Hanna Parnas, and Lee A. Segel. Neurotransmitter release: Development of a theory for total release based on kinetics. Journal of Theoretical Biology, 136(2):151–170, January 23, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938980222X.

Lopez-Quintela:1989:RME

[LQC89] M. Arturo López-Quintela and Julio Casado. Revision of the methodology in enzyme kinetics: a fractal approach. Journal of Theoretical Biology, 139(1):129–139, July 10, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389800621.

Lee:1980:EDE

[LR80] Sun Bok Lee and Dewey D. Y. Ryu. External diffusion effects on the kinetic constants of immobilized enzyme systems. Journal of Theoretical Biology, 84(2):259–279, May 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519380800075.

Lowry:1983:NMP

[LR83] R. K. Lowry and D. A. Ratkowsky. A note on models for poikilotherm development. Journal of Theoretical Biology, 105(3):453–459, December 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383901868.

Lawrence:1985:MLE

[LR85a] Charles E. Lawrence and Andrew A. Reilly. Maximum likelihood estimation of subsequence conservation. Journal of Theoretical Biology, 113(3):425–439, April 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938580031X.

Licko:1985:RSB

[LR85b] Vojtech Licko and Hershel Raff. Rate sensitivity of blood pressure to hypoxia. Journal of Theoretical Biology, 112(4):839–845, February 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385800655.

Louie:1988:MSI

[LR88] A. H. Louie and I. W. Richardson. The metric structure of irreversible thermodynamics. Journal of Theoretical Biology, 132(1): 125–126, May 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519388801978.

Laws:1983:TEE

[LRKC83] Edward A. Laws, Donald G. Redalje, David M. Karl, and Michael S. Chalup. A theoretical and experimental examination of the predictions of two recent models of phytoplankton growth. Journal of Theoretical Biology, 105(3):469–491, December 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383901881.

Louie:1982:PCR

[LRS82] A. H. Louie, I. W. Richardson, and S. Swaminathan. A phenomenological calculus for recognition processes. Journal of Theoretical Biology, 94(1):77–93, January 7, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382903307.

Levin:1981:ASN

[LS81] Gilbert V. Levin and Patricia A. Straat. Antarctic soil no. 726 and implications for the Viking labeled release experiment. Journal of Theoretical Biology, 91(1):41–45, July 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381903726.

Louie:1982:DGP

[LS82] A. H. Louie and R. L. Somorjai. Differential geometry of proteins: a structural and dynamical representation of patterns. Journal of Theoretical Biology, 98(2):189–209, September 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902582.

Lewis:1987:DCO

[LS87] R. D. Lewis and D. S. Saunders. A damped circadian oscillator model of an insect photoperiodic clock. I. Description of the model based on a feedback control system. Journal of Theoretical Biology, 128(1):47–59, September 7, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387800309.

Lomnicki:1988:RPP

[LS88] Adam Lomnicki and Stanisław Sedziwy. Resource partitioning and population stability under exploitation competition. Journal of Theoretical Biology, 132(1):119–120, May 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388801954.

Lomnicki:1989:DID

[LS89] Adam Lomnicki and Stanisław Sedziwy. Do individual differences in resource intakes without monopolization cause population stability and persistence? Journal of Theoretical Biology, 136(3):317–326, February 8, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389801675.

Longo:1985:CMS

[LSFT85] E. Longo, F. M. L. G. Stamato, R. Ferreira, and O. Tapia. The catalytic mechanism of serine proteases II: The effect of the protein environment in the α-chymotrypsin proton relay system. Journal of Theoretical Biology, 112(4):783–798, February 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385800618.

Lucas:1988:DCP

[LSH88] Jeffrey R. Lucas and Paul Schmid-Hempel. Diet choice in patches: Time-constraint and state-space solutions. Journal of Theoretical Biology, 131(3):307–332, April 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388802285.

Levy:1985:MMD

[LTM85] Bernard I. Levy, Alain Tedgui, and Jean Baptiste Michel. A mechanical model of the dynamics of the coronary circulation in dog. Journal of Theoretical Biology, 116(2):225–242, September 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385802666.

Luo:1988:IPN

[LTZ88] Liao Fu Luo, Lu Tsai, and Yu Min Zhou. Informational parameters of nucleic acid and molecular evolution. Journal of Theoretical Biology, 130(3):351–361, February 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388800341.

Luchnik:1982:ECS

[Luc82] Andrey N. Luchnik. Eukaryote chromosome structure: Foundation of the bineme model. Journal of Theoretical Biology, 99(1):135–157, November 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519382903952.

Luftig:1982:DCP

[Luf82] Ronald B. Luftig. Does the cytoskeleton play a significant role in animal virus replication? Journal of Theoretical Biology, 99(1):173–191, November 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382903976.

Luke:1984:GPS

[Luk84] Robert L. Luke. Glomerular permselection: Shape and flow. Journal of Theoretical Biology, 106(2):141–156, January 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519384900158.

Lumsden:1982:SRP

[Lum82] Charles J. Lumsden. The social regulation of physical caste: the superorganism revived. Journal of Theoretical Biology, 95(4):749– 781, April 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382903526.

Lumsden:1984:DIH

[Lum84] Charles J. Lumsden. Dual inheritance in haploid organisms: a model of magnetotactic bacteria. Journal of Theoretical Biology, 111(1):1–16, November 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801927.

Lumsden:1988:GCC

[Lum88] Charles J. Lumsden. Gene-culture coevolution: a test of the steady-state hypothesis for gene-culture translation. Journal of Theoretical Biology, 130(4):391–406, February 21, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388802054.

Lundberg:1987:PBM

[Lun87] Per Lundberg. Partial bird migration and evolutionarily stable strategies. Journal of Theoretical Biology, 125(3):351–360, April 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938780067X.

Liljenstrom:1987:TRM

[LvH87] Hans Liljenström and Gunnar von Heijne. Translation rate modification by preferential codon usage: Intragenic position effects. Journal of Theoretical Biology, 124(1):43–55, January 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519387802515.

Landau:1987:ESM

[LVN87] Gad M. Landau, Uzi Vishkin, and Ruth Nussinov. An efficient string matching algorithm with K substitutions for nucleotide and amino acid sequences. Journal of Theoretical Biology, 126(4):483–490, June 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387801534.

Milgram:1983:PAM

[MA83] M. Milgram and H. Atlan. Probabilistic automata as a model for epigenesis of cellular networks. Journal of Theoretical Biology, 103(4):523–547, August 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383902813.

MacWilliams:1982:NSH

[Mac82] Harry K. MacWilliams. Numerical simulations of hydra head regeneration using a proportion-regulating version of the Gierer– Meinhardt model. Journal of Theoretical Biology, 99(4):681–703, December 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382901941.

Morrison:1983:PMD

[MAC^{+}83] P. F. Morrison, J. Aroesty, S. P. Creekmore, P. E. Barker, and T. L. Lincoln. A preliminary model of doable-minate-mediated gene amplification. Journal of Theoretical Biology, 104(1):71–91, September 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383904022.

MacDonald:1984:UGM

[Mac84] N. MacDonald. The usefulness of growth models for trees. Journal of Theoretical Biology, 111(2):419–423, November 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384802209.

Macnair:1987:MSG

[Mac87] Mark R. Macnair. A model for the spread of genes for postmating reproductive isolation by natural selection. Journal of Theoretical Biology, 125(1):105–115, March 7, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387801832.

McGilchrist:1989:MTM

[MAC89] C. A. McGilchrist, C. W. Aisbett, and S. Cooper. A Markov transition model in the analysis of the immune response. Journal of Theoretical Biology, 138(1):17–21, May 7, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389801754.

Mader:1988:TTA

[Mad88] A. Mader. A transcription-translation activation feedback circuit as a function of protein degradation, with the quality of protein mass adaptation related to the average functional load. Journal of Theoretical Biology, 134(2):135–157, September 21, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938880198X.

Maes:1984:PLR

[Mae84] F. W. Maes. Problems with linear regression analysis of pharmacological and biochemical data. Journal of Theoretical Biology, 111(4):817–819, December 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384802696.

Magner:1982:ISP

[Mag82] James A. Magner. Information in the signal peptide? Journal of Theoretical Biology, 99(4):831–833, December 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902016.

Maguire:1984:MMH

[Mag84] Marjorie P. Maguire. The mechanism of meiotic homologue pairing. Journal of Theoretical Biology, 106(4):605–615, February 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519384900109.

Maguire:1986:SMF

[Mag86] Marjorie P. Maguire. “semi-meiosis” as the first step in the evolution of meiosis. Journal of Theoretical Biology, 120(4):503– 505, June 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519386800443.

Maguire:1988:CSD

[Mag88] M. P. Maguire. Crossover site determination and interference. Journal of Theoretical Biology, 134(4):565–570, October 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388800584.

Menendez-Arias:1987:SCT

[MAGR87] Luis Menéndez-Arias, José G. Gavilanes, and Rosalía Rodriguez. A structural comparison of type c lysozymes based on their hydropathic profiles. Journal of Theoretical Biology, 127(2):221– 228, July 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387801327.

Mahaffy:1984:CCMa

[Mah84a] J. M. Mahaffy. Cellular control models with linked positive and negative feedback and delays. I. The models. Journal of Theoretical Biology, 106(2):89–102, January 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519384900110.

Mahaffy:1984:CCMb

[Mah84b] J. M. Mahaffy. Cellular control models with linked positive and negative feedback and delays. II. Linear analysis and local stability. Journal of Theoretical Biology, 106(2):103–118, January 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519384900122.

Makela:1986:IPM

[Mäk86] Annikki Mäkelä. Implications of the pipe model theory on dry matter partitioning and height growth in trees. Journal of Theoretical Biology, 123(1):103–120, November 7, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386802387.

Malcolm:1981:TRS

[Mal81a] Alan D. B. Malcolm. Thermophiles and RNA stability. Journal of Theoretical Biology, 92(4):479–481, October 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902605.

Malpress:1981:CHN

[Mal81b] F. H. Malpress. On the chemiosmotic hypothesis and the nature of the mitochondrial protonmotive force. Journal of Theoretical Biology, 92(3):255–265, October 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902915.

Malpress:1984:CHMb

[Mal84a] F. H. Malpress. The Coulombic hypothesis of mitochondrial energy transduction: an attempt to quantify relationships in the energized zones. Journal of Theoretical Biology, 111(2):397–405, November 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384802180.

Malpress:1984:CHMa

[Mal84b] F. H. Malpress. A Coulombic hypothesis of mitochondrial oxidative phosphorylation. Journal of Theoretical Biology, 109(4):501– 521, August 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384801551.

Malpress:1984:LEP

[Mal84c] F. H. Malpress. Local energized proton hypotheses of mitochondrial oxidative phosphorylation. Journal of Theoretical Biology, 109(4):523–532, August 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801563.

Malpress:1985:CHM

[Mal85] F. H. Malpress. The Coulombic hypothesis of mitochondrial energy transduction: a quantitative emendation. Journal of Theoretical Biology, 114(2):351–352, May 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801119.

Mizraji:1987:VME

[MAL87] E. Mizraji, L. Acerenza, and J. Lin. Viscoelastic models for enzymes with multiple conformational states. Journal of Theoretical Biology, 129(2):163–175, November 21, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387800103.

Malchow:1988:DPF

[Mal88] Horst Malchow. Dissipative pattern formation in ternary non-linear reaction-electrodiffusion systems with concentrationdependent diffusivities. Journal of Theoretical Biology, 135 (3):371–381, December 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519388802510.

Manabe:1981:MER

[Man81a] Tomohiro Manabe. Mathematical expression for the rate of initiation and certain prediction on the non-competitive binding operon: the Escherichia coli lactose operon. Journal of Theoretical Biology, 89(2):271–302, March 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251938190312X.

Manabe:1981:TRA

[Man81b] Tomohiro Manabe. Theory of regulation by the attenuation mechanism: stochastic model for the attenuation of the Escherichia coli tryptophan operon. Journal of Theoretical Biology, 91(4):527–544, August 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902101.

Manning:1981:SEE

[Man81c] J. T. Manning. The “survivor effect” and the evolution of parthenogenesis and self-fertilization. Journal of Theoretical Biology, 93(2):491–493, November 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381901156.

Manning:1982:HHS

[Man82a] J. T. Manning. Hitch-hiking and a short-term advantage for sex. Journal of Theoretical Biology, 98(4):703–705, October 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382901473.

Manning:1982:SFS

[Man82b] J. T. Manning. Sex and the fixation of single favourable mutations. Journal of Theoretical Biology, 94(4):905–908, February 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382900856.

Manning:1984:MAS

[Man84] J. T. Manning. Males and the advantage of sex. Journal of Theoretical Biology, 108(2):215–220, May 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384800673.

Manning:1985:CFC

[Man85] J. T. Manning. Choosy females and correlates of male age. Journal of Theoretical Biology, 116(3):349–354, October 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385802733.

Maroun:1980:IAC

[Mar80a] Leonard E. Maroun. Interferon action and chromosome 21 trisomy. Journal of Theoretical Biology, 86(3):603–606, October 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380903562.

Martindale:1980:MAA

[Mar80b] Steven Martindale. On the multivariate analysis of avian vocalizations. Journal of Theoretical Biology, 83(1):107–110, March 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380903744.

Maroudas:1984:CME

[Mar84] N. G. Maroudas. Cell membrane entropy as regulator of receptor submergence. Journal of Theoretical Biology, 107(2):243–248, March 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384800259.

Massaldi:1984:PCMb

[Mas84] Hugo A. Massaldi. The physico-chemical mechanism of mediated transport: II. Osmotic and isosmotic volume flow. Journal of Theoretical Biology, 110(1):35–57, September 7, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384800144.

Massing:1986:BTR

[Mas86] Walter Massing. The brain, time reversal and Libet’s concept of antedating. Journal of Theoretical Biology, 120(4):443–446, June 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386800376.

Matthews:1980:CCM

[Mat80] Harry R. Matthews. Chromosome condensation in mitosis. Journal of Theoretical Biology, 83(2):367–368, March 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380903008.

Mattson:1981:NHM

[Mat81] Peter A. H. Mattson. A new hypothesis for molecular behavior. Journal of Theoretical Biology, 93(4):829–853, December 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381903428.

Matsuno:1983:ECI

[Mat83] Koichiro Matsuno. Evolutionary changes in the information content of polypeptides. Journal of Theoretical Biology, 105(2):185– 199, ???? 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519383800010.

Matsumoto:1984:PMM

[Mat84] Gen Matsumoto. A proposed membrane model for generation of sodium currents in squid giant axons. Journal of Theoretical Biology, 107(4):649–666, April 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938480137X.

Matsuda:1985:ESS

[Mat85] Hiroyuki Matsuda. Evolutionarily stable strategies for predator switching. Journal of Theoretical Biology, 115(3):351–366, August 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385801971.

Matis:1987:CSM

[Mat87] J. H. Matis. The case for stochastic models of digesta flow. Journal of Theoretical Biology, 124(3):371–376, February 7, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387801224.

Menendez-Arias:1987:RBH

[MATGR87] Luis Menéndez-Arias, Javier Turnay, José G. Gavilanes, and Rosalía Rodríguez. Relationship between hydropathic variability and functional properties of α-lactalbumins and type c lysozymes. Journal of Theoretical Biology, 126(1):91–100, May 7, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387801029.

Maurer:1987:SBC

[Mau87] Brian A. Maurer. Scaling of biological community structure: a systems approach to community complexity. Journal of Theoretical Biology, 127(1):97–110, July 7, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387801637.

Mawson:1985:SOI

[Maw85] Anthony R. Mawson. Speculations on the origin of the illeffects associated with the use of visual display terminals. Journal of Theoretical Biology, 114(2):223–241, May 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801053.

Majkowski:1980:SBG

[MB80] J. Majkowski and L. Bramall. Sensitivity of bioenergetic growth models of animals to changes in the energy balance parameters. Journal of Theoretical Biology, 85(4):643–655, August 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380902623.

Magnani:1982:CSG

[MB82] M. Magnani and E. Beretta. Cell survival and glutathione reductase. Journal of Theoretical Biology, 94(4):995–997, February 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382900935.

Massaldi:1984:PCMa

[MB84a] H. A. Massaldi and C. H. Borzi. The physico-chemical mechanism of mediated transport I. Facilitated diffusion. Journal of Theoretical Biology, 106(4):537–557, February 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519384900067.

McCoy:1984:PCM

[MB84b] E. L. McCoy and L. Boersma. The principles of continuum mechanics applied to transport processes and deformation in plant tissue. Journal of Theoretical Biology, 111(4):687–705, December 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384802623.

Munton:1985:ASS

[MB85] S. L. Munton and G. G. Birch. Accession of sweet stimuli to receptors. I. Absolute dominance of one molecular species in binary mixtures. Journal of Theoretical Biology, 112(3):539–551, February 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385800217.

McCoy:1986:BAP

[MB86] E. L. McCoy and L. Boersma. The bioenergetics of axial plant growth I. Concept and model formulation. Journal of Theoretical Biology, 118(1):83–100, January 7, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386800108.

Morel:1988:CMJ

[MB88a] Jean E. Morel and Nacima Bachouchi. Comments on the molecular jet process for the in vitro movements: Is it valid to use steady state kinetics to analyse actin-myosin based motility? Journal of Theoretical Biology, 135(1):119–121, November 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388801784.

Morel:1988:MCM

[MB88b] Jean E. Morel and Nacima Bachouchi. Muscle contraction and movement of cellular organelles: Are there two different types of mechanisms for their generation? Journal of Theoretical Biology, 132(1):83–96, May 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388801929.

Mendz:1989:MPF

[MBJK89] George L. Mendz, Brian T. Bulliman, Natalie L. James, and Philip W. Kuchel. Magnetic potential and field gradients of a model cell. Journal of Theoretical Biology, 137(1):55–69, March 7, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519389801493.

Motro:1989:NVB

[MC89] Uzi Motro and Dan Cohen. A note on vigilance behavior and stability against recognizable social parasites. Journal of Theoretical Biology, 136(1):21–25, January 9, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389801869.

McArthur:1987:TIB

[McA87] A. J. McArthur. Thermal interaction between animal and microclimate: a comprehensive model. Journal of Theoretical Biology, 126(2):203–238, May 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387802291.

Mege:1986:QCS

[MCB^{+}86] Jean-Louis Mege, Christian Capo, Anne-Marie Benoliel, Collette Foa, Rémy Galindo, and Pierre Bongrand. Quantification of cell surface roughness; a method for studying cell mechanical and adhesive properties. Journal of Theoretical Biology, 119(2):147– 160, March 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519386800704.

McClendon:1980:ECI

[McC80a] John H. McClendon. The evolution of the chemical isotopes as an analog of biological evolution. Journal of Theoretical Biology, 87(1):113–128, November 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380902234.

McClendon:1980:NDB

[McC80b] John H. McClendon. The necessary distinction between metabolism and evolution: a comment on the thermodynamics of evolution. Journal of Theoretical Biology, 83(3):523–524, April 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380900582.

McConnaughey:1980:LE

[McC80c] William McConnaughey. Letter to the editor. Journal of Theoretical Biology, 87(1):207–209, November 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380902295.

McCabe:1984:SRT

[McC84] M. McCabe. Should reptation theory for migration of linear DNA through gels include screwing? Journal of Theoretical Biology, 107(1):165–167, March 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801277.

McCallum:1987:PRA

[McC87] H. I. McCallum. Predator regulation of acanthaster planci. Journal of Theoretical Biology, 127(2):207–220, July 21, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387801315.

McGavin:1980:RTB

[McG80] Stewart McGavin. Relationships and transformations between some nucleic acid models. Journal of Theoretical Biology, 85(4):665–672, August 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519380902647.

McGavin:1981:SFC

[McG81] Stewart McGavin. Some formal consequences of a specific association between double stranded nucleic acid molecules-properties of multistranded structures. Journal of Theoretical Biology, 91(1): 33–40, July 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381903714.

McGavin:1984:CUM

[McG84] Stewart McGavin. Construction and use of models designed to demonstrate some topological properties of multistranded nucleic acid structure and its possible bearing on recombination. Journal of Theoretical Biology, 107(1):37–56, March 7, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801204.

McGinness:1985:NVP

[McG85] J. McGinness. A new view of pigmented neurons. Journal of Theoretical Biology, 115(3):475–476, August 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938580206X.

McGavin:1989:FSR

[McG89a] Stewart McGavin. Four strand recombination models. Journal of Theoretical Biology, 136(2):135–150, January 23, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389802218.

McGavin:1989:FSS

[McG89b] Stewart McGavin. Four-strand structure, kinks and cruciforms in DNA. Journal of Theoretical Biology, 138(1):117–128, May 7, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938980181X.

McKinley:1983:MTC

[McK83] Dana Noel McKinley. Model for tramformations of the clathrin lattice in the coated vesicle pathway. Journal of Theoretical Biology, 103(3):405–419, August 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383902941.

McMurtrie:1981:SDT

[McM81] R. McMurtrie. Suppression and dominance of trees with overlapping crowns. Journal of Theoretical Biology, 89(1):151–174, March 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381901831.

McNair:1987:SEP

[McN87] James N. McNair. Stability effects of prey refuges with entry-exit dynamics. Journal of Theoretical Biology, 125(4):449–464, April 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519387802138.

McNair:1989:SEJ

[McN89] James N. McNair. Stability effects of a juvenile period in age-structured populations. Journal of Theoretical Biology, 137(4):397–422, April 20, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519389800372.

Manning:1986:ARP

[MD86] J. T. Manning and D. P. E. Dickson. Asexual reproduction, polyploidy and optimal mutation rates. Journal of Theoretical Biology, 118(4):485–489, February 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801667.

Mignotte:1987:SDP

[MDBRM87] Bernard Mignotte, Dominique Dunon-Bluteau, Claude Reiss, and Jean-Claude Mounolou. Sequence deduced physical properties in the d-loop region common to five vertebrate mitochondrial DNAs. Journal of Theoretical Biology, 124(1):57–69, January 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519387802527.

Meneghini:1983:RDS

[MdMF83] Rogerio Meneghini and Alberto C. de Mello Filho. Rate of DNA synthesis in mammalian cells irradiated with ultravolet light: a model based on the variations in the rate of movement of the replication fork and in the number of active replicons. Journal of Theoretical Biology, 100(3):359–372, February 7, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383904344.

Maroun:1986:EML

[MDPF86] Leonard E. Maroun, Marykay Degner, James W. Precup, and Phillip P. Franciskovich. Eukaryotic mRNA 5’-leader sequences have dual regions of complementarity to the 3’-terminus of 18s rRNA. Journal of Theoretical Biology, 120(1):85–98, May 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386800194.

Margulis:1980:SAUa

[MDW80a] L. Margulis, J. F. Danielli, and L. Wolpert. Short articles on unsolved problems. Journal of Theoretical Biology, 84(1):1–2, May 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519380810344.

Margulis:1980:SAUb

[MDW80b] L. Margulis, J. F. Danielli, and L. Wolpert. Short articles on unsolved problems. Journal of Theoretical Biology, 85(1):1–2, July 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380902763.

Margulis:1980:SAUc

[MDW80c] L. Margulis, J. F. Danielli, and L. Wolpert. Short articles on unsolved problems. Journal of Theoretical Biology, 86(1):1–2, September 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519380900636.

Marc:1982:ISP

[ME82] Annie Marc and Jean Marc Engasser. Influence of substrate and product diffusion on the heterogeneous kinetics of enzymic reversible reactions. Journal of Theoretical Biology, 94(1):179–189, January 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382903393.

Meakin:1986:NMB

[Mea86] Paul Meakin. A new model for biological pattern formation. Journal of Theoretical Biology, 118(1):101–113, January 7, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938680011X.

Meister:1982:LDS

[Mei82] Armin Meister. Limits for decomposition of spectra into single bands. Journal of Theoretical Biology, 94(3):541–553, February 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382903009.

Melicow:1982:TSC

[Mel82] Meyer M. Melicow. The three steps to cancer: a new concept of cancerigenesis. Journal of Theoretical Biology, 94(2):471–511, January 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382903228.

Mendelson:1982:HCP

[Men82] Neil H. Mendelson. The helix clock: a potential biomechanical cell cycle timer. Journal of Theoretical Biology, 94(1):209–222, January 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382903423.

Mendelson:1985:MBD

[Men85] Neil H. Mendelson. A model of bacterial DNA segregation based upon helical geometry. Journal of Theoretical Biology, 112(1):25– 39, January 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385801156.

Meredith:1982:D

[Mer82] A. Meredith. Devolution. Journal of Theoretical Biology, 96(1): 49–65, May 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382901552.

Mercola:1987:POC

[Mer87] Dan Mercola. The proto-oncogene c-fos encodes a potential regulatory site that is disrupted by viral transduction. Journal of Theoretical Biology, 126(2):243–246, May 21, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938780231X.

Metz:1982:STS

[Met82] J. A. J. Metz. The short time scale behaviour of an enzymatic reaction is always stable: Comment to a paper by seshadri and fritzsch. Journal of Theoretical Biology, 97(2):341–342, July 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382901114.

Matela:1980:CSC

[MF80] Raymond J. Matela and Robert J. Fletterick. Computer simulation of cellular self-sorting: a topological exchange model. Journal of Theoretical Biology, 84(4):673–690, June 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519380800270.

Mikulecky:1981:RBL

[MF81] D. C. Mikulecky and J. J. Feher. The relation between local and global transport parameters in an epithelial membrane having an asymmetric transport mechanism. Journal of Theoretical Biology, 88(3):575–587, February 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902848.

Muller:1985:BSG

[MF85] Uwe R. Müller and Walter M. Fitch. The biological significance of G-T/G-U mispairing in nucleic acid secondary structure. Journal of Theoretical Biology, 117(1):119–126, November 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801673.

Morel:1989:OPU

[MF89] Richard E. Morel and George Fleck. Onsager’s principle: a unifying bio-theme. Journal of Theoretical Biology, 136(2):171–175, January 23, 1989. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519389802231.

Meinhardt:1980:GRS

[MG80] Hans Meinhardt and Alfred Gierer. Generation and regeneration of sequence of structures during morphogenesis. Journal of Theoretical Biology, 85(3):429–450, August 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380903185.

Margineanu:1981:SMP

[MG81] Ilca Mărgineanu and Victor Ghetie. A selective model of plasma protein catabolism. Journal of Theoretical Biology, 90(1):101– 110, May 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381901247.

Mosteller:1984:MMA

[MG84a] Raymond D. Mosteller and Bruce E. Goldstein. A mathematical model that applies to protein degradation and posttranslational processing of proteins and to analogous processes for other molecules in non-growing and exponentially growing cells. Journal of Theoretical Biology, 108(4):597–621, June 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384800818.

Murray:1984:CNM

[MG84b] Martyn G. Murray and Russell Gerrard. Conflict in the neighbourhood: Models where close relatives are in direct competition. Journal of Theoretical Biology, 111(2):237–246, November 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384802088.

Murray:1985:PCR

[MG85] Martyn G. Murray and Russell Gerrard. Putting the challenge into resource exploitation: a model of contest competition. Journal of Theoretical Biology, 115(3):367–389, August 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801983.

Moore:1981:SRE

[MGH81] William S. Moore, W. Gordon, and S. Hines. Sex in random environments. Journal of Theoretical Biology, 92(3):301–316, October 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381902940.

Mazat:1981:CRE

[MGL81] Jean-Pierre Mazat, Pierre-Vincent Graves, and Bernard Labouesse. Coupling in relaxation experiments as a diagnostic tool to discriminate among different antico-operative models. Journal of Theoretical Biology, 89(1):53–67, March 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251938190179X.

Marchi:1980:ATK

[MH80a] Ezio Marchi and Jorge Horas. On aspartate transcarbamylase kinetics. Journal of Theoretical Biology, 85(3):413–421, August 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380903161.

McNamara:1980:ASD

[MH80b] John McNamara and Alasdair Houston. The application of statistical decision theory to animal behaviour. Journal of Theoretical Biology, 85(4):673–690, August 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380902659.

Mohammad:1980:CRM

[MH80c] S. Noor Mohammad and A. J. Hopfinger. Chemical reactivity of a methyldiazonium ion with nucleophilic centers of DNA bases. Journal of Theoretical Biology, 87(2):401–419, November 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380903677.

Moran:1980:LNQ

[MH80d] Nancy Moran and W. D. Hamilton. Low nutritive quality as defense against herbivores. Journal of Theoretical Biology, 86(2):247–254, September 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380900041.

Moore:1981:SLA

[MH81] G. William Moore and Grover M. Hutchins. Symbolic logic analysis of cause of dealth in humans: Application to 108 patients after coronary artery bypass graft surgery. Journal of Theoretical Biology, 92(3):267–291, October 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902927.

Muller-Herold:1983:WH

[MH83]  U. Müller-Herold. What is a hypercycle?  Journal of

Theoretical Biology, 102(4):569–584, June 21, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383903910.

Muller-Herold:1984:FSM

[MH84] U. Müller-Herold. A fully specified model of a non-linear hypercycle. Journal of Theoretical Biology, 110(1):21–26, September 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384800120.

McNamara:1985:OFL

[MH85] John M. McNamara and Alasdair I. Houston. Optimal foraging and learning. Journal of Theoretical Biology, 117(2):231–249, November 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385802198.

McNamara:1987:MEU

[MH87] John M. McNamara and Alasdair I. Houston. Memory and the efficient use of information. Journal of Theoretical Biology, 125(4):385–395, April 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387802096.

McNamara:1989:SDC

[MH89] John M. McNamara and Alasdair I. Houston. State-dependent contests for food. Journal of Theoretical Biology, 137(4):457– 479, April 20, 1989. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519389800402.

Mohammad:1983:IMP

[MHB83] S. Noor Mohammad, A. J. Hopfinger, and D. R. Bickers. Intrinsic mutagenicity of polycyclic aromatic hydrocarbons: a quantitative structure activity study based upon molecular shape analysis. Journal of Theoretical Biology, 102(2):323–331, May 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383903697.

Murase:1989:PED

[MHB89] M. Murase, M. Hines, and J. J. Blum. Properties of an excitable dynein model for bend propagation in cilia and flagella. Journal of Theoretical Biology, 139(3):413–430, August 9, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938980219X.

Melendez-Hevia:1985:GPP

[MHI85] Enrique Meléndez-Hevia and Angel Isidoro. The game of the pentose phosphate cycle. Journal of Theoretical Biology, 117 (2):251–263, November 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519385802204.

Miyamoto:1983:CPI

[MHII83] Yusei Miyamoto, Manabu Hanano, Tatsuji Iga, and Masanobu Ishikawa. Concentration profile in the intestinal tract and drug absorption model: two-dimensional laminar flow in a circular porous tube. Journal of Theoretical Biology, 102(4):585–601, June 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383903922.

Muller-Herold:1987:QMC

[MHLK87] U. Müller-Herold, H. U. Lutz, and O. Kedem. Quantum mechanical coherence in red blood cells: No experimental evidence. Journal of Theoretical Biology, 126(2):251–252, May 21, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387802333.

Melendez-Hevia:1988:EDM

[MHT88] Enrique Meléndez-Hevia and Néstor V. Torres. Economy of design in metabolic pathways: Further remarks on the game of the pentose phosphate cycle. Journal of Theoretical Biology, 132(1): 97–111, May 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519388801930.

Maruyama:1981:ERM

[MI81] Takeo Maruyama and Hirotami T. Imai. Evolutionary rate of the mammalian karyotype. Journal of Theoretical Biology, 90(1): 111–121, May 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381901259.

Michel:1986:NSA

[Mic86] Christian J. Michel. New statistical approach to discriminate between protein coding and non-coding regions in DNA sequences and its evaluation. Journal of Theoretical Biology, 120(2):223– 236, May 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251938680176X.

Michaux:1988:OAW

[Mic88] B. Michaux. Organotaxism: an alternative “way of seeing” the fossil record. Journal of Theoretical Biology, 133(4):397–408, August 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519388803308.

Mikelsaar:1981:CBP

[Mik81] Raik Mikelsaar. C–A base pairs in transfer ribonucleic acids and codon-anticodon complexes. Journal of Theoretical Biology, 92(2):163–180, September 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381903908.

Mikelsaar:1983:HMG

[Mik83] Raik Mikelsaar. Human mitochondrial genome and the evolution of methionine transfer ribonucleic acids. Journal of Theoretical Biology, 105(2):221–232, ???? 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519383800046.

Mikhaltsev:1984:PN

[Mik84a] Igor E. Mikhaltsev. Parametric neuromorphology. Journal of Theoretical Biology, 109(3):461–462, August 7, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384800934.

Miklavc:1984:CBS

[Mik84b] Adolf Miklavc. Could binding sites of a receptor be “seen”? A great challenge to picosecond spectroscopy. Journal of Theoretical Biology, 111(2):231–236, November 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384802076.

Miller:1987:SRS

[Mil87] Mark J. Miller. Sensitivity of RNA synthesis to actinomycin D inhibition is dependent on the frequency of transcription: a mathematical model. Journal of Theoretical Biology, 129(3):289–299, December 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387800036.

Miller:1989:DWE

[Mil89] Arnold R. Miller. The distribution of wearout over evolved reliability structures. Journal of Theoretical Biology, 136(1):27–46, January 9, 1989. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519389801870.

Minkevich:1982:PCP

[Min82] I. G. Minkevich. Physico-chemical properties of organic compounds and the energetics of metabolism. Journal of Theoretical Biology, 95(3):569–590, April 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382900352.

Mittwoch:1988:SDS

[Mit88] Ursula Mittwoch. Sex determination and steroid ratios. Journal of Theoretical Biology, 133(2):259–260, July 21, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388800109.

Mittwoch:1989:SDM

[Mit89] Ursula Mittwoch. Sex differentiation in mammals and tempo of growth: Probabilities vs. switches. Journal of Theoretical Biology, 137(4):445–455, April 20, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389800396.

Miyamoto:1986:PQT

[Miy86] Michael D. Miyamoto. Probability of quantal transmitter release from nerve terminals: Theoretical considerations in the determination of spatial variation. Journal of Theoretical Biology, 123(3):289–304, December 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386802442.

Manning:1980:BAE

[MJ80] J. T. Manning and J. Jenkins. The “balance” argument and the evolution of sex. Journal of Theoretical Biology, 86(3):593–601, October 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519380903550.

Marsh:1988:FCR

[MJ88] L. M. Marsh and R. E. Jones. The form and consequences of random walk movement models. Journal of Theoretical Biology, 133(1):113–131, July 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388800286.

Mark:1988:BIA

[MJN88] Alan E. Mark, Peter D. Jeffrey, and Lawrence W. Nichol. The binding of an indefinitely associating ligand to acceptor: Consideration of monovalent ligand species binding to a multivalent acceptor. Journal of Theoretical Biology, 131(2):137–149, March 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388802315.

MacCarthy:1981:DMV

[MK81] John E. MacCarthy and John J. Kozak. A dynamic model for in vivo virus replication. Journal of Theoretical Biology, 90(2):265– 281, May 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381900461.

Marzec:1983:PMS

[MK83] C. Marzec and J. Kappraff. Properties of maximal spacing on a circle related to phyllotaxis and to the golden mean. Journal of Theoretical Biology, 103(2):201–226, July 21, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383900255.

Meinhardt:1987:MPF

[MK87] Hans Meinhardt and Martin Klingler. A model for pattern formation on the shells of molluscs. Journal of Theoretical Biology, 126(1):63–89, May 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519387801017.

Markin:1988:MPR

[MK88] V. S. Markin and M. M. Kozlov. Mechanical properties of the red cell membrane skeleton: Analysis of axisymmetric deformations. Journal of Theoretical Biology, 133(2):147–167, July 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938880002X.

Matteson:1981:SMI

[MKL81] D. R. Matteson, M. E. Kriebel, and F. Llados. A statistical model indicates that miniature end-plate potentials and unitary evoked end-plate potentials are composed of subunits. Journal of Theoretical Biology, 90(3):337–363, June 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381903167.

Mushinski:1980:CAA

[MKM80] J. Frederic Mushinski, James A. Koziol, and Marina Marini. Cluster analysis of aminoacyl-tRNAs from mouse plasmacytomas correlates chromatographic profiles with myeloma protein similarity, clonal origin of tumor lines, and the neoplastic nature of the tissues. Journal of Theoretical Biology, 85(3):507–521, August 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380903240.

Matsuda:1986:SES

[MKS^{+}86] Hiroyuki Matsuda, Kohkichi Kawasaki, Nanako Shigesada, Ei Teramoto, and Luigi M. Ricciardi. Switching effect on the stability of the prey-predator system with three trophic levels. Journal of Theoretical Biology, 122(3):251–262, October 7, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801187.

Mizraji:1982:MEL

[ML82] E. Mizraji and J. Lin. Migration of enzymes on lidear polymers. Journal of Theoretical Biology, 96(4):561–569, June 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902314.

Mizraji:1984:CTP

[ML84] E. Mizraji and J. Lin. Chemicomechanical transduction performed by enzymes activated by polymers. Journal of Theoretical Biology, 108(2):173–189, May 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384800648.

Mizraji:1986:ESD

[ML86] E. Mizraji and J. Lin. Energy storage during DNA girase activity. Journal of Theoretical Biology, 120(1):63–71, May 7, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386800170.

MacFie:1988:NTC

[MLB88] Halliday J. H. MacFie, Nicholas D. Light, and Allen J. Bailey. Natural taxonomy of collagen based on amino acid composition. Journal of Theoretical Biology, 131(4):401–418, April 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388800377.

Milton:1989:CDB

[MLB^{+}89] John G. Milton, André Longtin, Anne Beuter, Michael C. Mackey, and Leon Glass. Complex dynamics and bifurcations in neurology. Journal of Theoretical Biology, 138(2):129–147, May 22, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519389801353.

Maroti:1980:DDL

[MLS80] P. Maróti, G. Laczkó, and L. Szalay. Determination of the distance law of the transfer of electronic excitation energy. Journal of Theoretical Biology, 86(4):663–671, October 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380903057.

Mortola:1980:TSA

[MM80] J. P. Mortola and S. A. Mortola. Tracheal slowly adapting stretch receptors: Theoretical models. Journal of Theoretical Biology, 83(2):313–320, March 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380902957.

Madhavan:1982:PRT

[MM82a] Mekkara Mandaravally Madhavan and Kornath Madhavan. Pattern regulation in tergite of drosophila: a model. Journal of Theoretical Biology, 95(4):731–748, April 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382903514.

McFadden:1982:DCS

[MM82b] Grant McFadden and A. Richard Morgan. DNA cruciform structures: Implications for telomer replication in eukaryotes and instability of long palindromic DNA sequences in prokaryotes. Journal of Theoretical Biology, 97(2):343–349, July 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382901126.

Moore:1982:ASR

[MM82c] Gordon P. Moore and Arnold R. Moore. The average spacing of restriction enzyme recognition sites in DNA. Journal of Theoretical Biology, 98(1):165–169, September 7, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382900649.

Mankin:1983:PMP

[MM83a] R. W. Mankin and M. S. Mayer. A phenomenological model of the perceived intensity of single odorants. Journal of Theoretical Biology, 100(1):123–138, January 7, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383900978.

Mankin:1983:SRR

[MM83b] R. W. Mankin and M. S. Mayer. Stimulus-response relationships of insect olfaction: Correlations among neurophysiological and behavioral measures of response. Journal of Theoretical Biology, 100(4):613–630, February 21, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383903260.

Millan:1983:DSG

[MM83c] Leopoldo Millán and Ezio Marchi. Denaturation of some general DNA sequences. Journal of Theoretical Biology, 105(2):245–258, ???? 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938380006X.

Mittenthal:1983:MSG

[MM83d] Jay E. Mittenthal and Robert M. Mazo. A model for shape generation by strain and cell-cell adhesion in the epithelium of an arthropod leg segment. Journal of Theoretical Biology, 100(3):443–483, February 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383904411.

McLean:1985:MCM

[MM85] K. R. McLean and J. T. Manning. Mate choice, monogamy and mutational load. Journal of Theoretical Biology, 116(3):369–376, October 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385802757.

Mazat:1986:DSE

[MM86a] Jean Pierre Mazat and Françoise Mazat. Double site enzyme and squatting: Where one regulatory ligand is also a substrate of the reaction. Journal of Theoretical Biology, 121(1):89–103, July 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386800303.

Morozov:1986:TMW

[MM86b] V. N. Morozov and T. Ya. Morozova. Thermal motion of whole protein molecules in protein solids. Journal of Theoretical Biology, 121(1):73–88, July 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386800297.

Moore:1984:FMS

[MMG84] Gordon P. Moore, Arnold R. Moore, and Lawrence I. Grossman. The frequency of matching sequences in DNA. Journal of Theoretical Biology, 108(1):111–122, May 7, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801721.

Michelson:1987:TME

[MMGL87] S. Michelson, B. E. Miller, A. S. Glicksman, and J. T. Leith. Tumor micro-ecology and competitive interactions. Journal of Theoretical Biology, 128(2):233–246, September 21, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387801716.

Mahon:1980:TCT

[MMGM80] Garry A. T. Mahon, Peter McWilliam, Robert L. Gordon, and David J. McConnell. The time course of transcription. Journal of Theoretical Biology, 87(3):483–515, December 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380902325.

Mooney:1985:SPP

[MN85] J. R. Mooney and B. N. Nagorcka. Spatial patterns produced by a reaction–diffusion system in primary hair follicles. Journal of Theoretical Biology, 115(2):299–317, July 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801028.

Matsuda:1989:CES

[MN89a] Hiroyuki Matsuda and Toshiyuki Namba. Co-evolutionarily stable community structure in a patchy environment. Journal of Theoretical Biology, 136(2):229–243, January 23, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389802280.

Miklavic:1989:CSB

[MN89b] S. J. Miklavic and B. W. Ninham. Conformation of surface bound polyelectrolytes: I. Implications for cell electrophoresis. Journal of Theoretical Biology, 137(1):71–89, March 7, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938980150X.

Mogie:1986:RBA

[Mog86] Michael Mogie. On the relationship between asexual reproduction and polyploidy. Journal of Theoretical Biology, 122(4):493–498, October 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519386801898.

Makino:1987:SPI

[MOK87] Kimiko Makino, Hiroyuki Ohshima, and Tamotsu Kondo. Surface potential of an ion-penetrable charged membrane. Journal of Theoretical Biology, 125(3):367–368, April 7, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387800693.

Monoi:1983:IIA

[Mon83] Hiroshi Monoi. Ionic interactions and anion binding in the gramicidin channel. An electrostatic calculation. Journal of Theoretical Biology, 102(1):69–99, May 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383902631.

Montgomery:1988:TCV

[Mon88] John C. Montgomery. Temperature compensation in the vestibulo-ocular reflex: a novel hypothesis of cerebellar function. Journal of Theoretical Biology, 132(2):163–170, May 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388801541.

Moody:1988:EMS

[Moo88] Michael E. Moody. The evolution of migration in subdivided populations. I. Haploids. Journal of Theoretical Biology, 131(1): 1–14, March 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519388801164.

Morowitz:1980:DNS

[Mor80] Harold J. Morowitz. The dimensionality of niche space. Journal of Theoretical Biology, 86(2):259–263, September 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380900065.

Morley:1982:ARD

[Mor82] Alexander A. Morley. Is ageing the result of dominant and codominant mutations? Journal of Theoretical Biology, 98(3):469– 474, October 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251938290131X.

Moroz:1984:CFG

[Mor84] Pavel E. Moroz. The cell in the field of gravity and the centrifugal field. Journal of Theoretical Biology, 107(2):303–320, March 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384800302.

Morel:1985:DSW

[Mor85a] Jean E. Morel. Discussion on the state of water in the myofilament lattice and other biological systems, based on the fact that the usual concepts of colloid stability can not explain the stability of the myofilament lattice. Journal of Theoretical Biology, 112(4):847–858, February 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800667.

Morpurgo:1985:DIE

[Mor85b] Giorgio Morpurgo. Deregulation of immunoglobulin expression: Is this the main cause of Burkitt’s lymphoma? Journal of Theoretical Biology, 116(3):395–398, October 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385802770.

Moreno:1986:TAA

[Mor86] Cosme M. Moreno. Theoretical approaches to d-amino acid oxidase. Journal of Theoretical Biology, 119(3):369–378, April 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386801485.

Motro:1982:CHP

[Mot82] Uzi Motro. The courtship handicap-phenotypic effect. Journal of Theoretical Biology, 97(2):319–324, July 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382901072.

Motro:1988:ESS

[Mot88] Uzi Motro. Evolutionarily stable strategies of mutual help between relatives having unequal fertilities. Journal of Theoretical Biology, 135(1):31–39, November 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388801723.

Motro:1989:SPE

[Mot89] Uzi Motro. Should a parasite expose itself? (some theoretical aspects of begging and vigilance behavior). Journal of Theoretical Biology, 140(2):279–287, September 22, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389801341.

Muller:1982:QHM

[MOV82] M. Muller, J. W. M. Osse, and J. H. G. Verhagen. A quantitative hydrodynamical model of suction feeding in fish. Journal of Theoretical Biology, 95(1):49–79, March 7, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902879.

Moynihan:1982:SPI

[Moy82a] M. Moynihan. Spatial parameters and interspecific social relations: Some differences between birds and fishes. Journal of Theoretical Biology, 95(2):253–262, March 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902430.

Moynihan:1982:WLA

[Moy82b] M. Moynihan. Why is lying about intentions rare during some kinds of contests? Journal of Theoretical Biology, 97(1):7– 12, July 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382902727.

Maresin:1985:TAE

[MP85a] V. M. Maresin and E. V. Presnov. Topological approach to embryogenesis. Journal of Theoretical Biology, 114(3):387–398, June 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385801740.

Moskowitz:1985:UTP

[MP85b] Nathan Moskowitz and Saul Puszkin. A unified theory of presynaptic chemical neurotransmission. Journal of Theoretical Biology, 112(3):513–534, February 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800199.

McKinney:1987:DRL

[MP87] James D. McKinney and Lee G. Pedersen. Do residue levels of polychlorinated biphenyls (PCBs) in human blood produce mild hypothyroidism? Journal of Theoretical Biology, 129(2):231–241, November 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387800152.

Morpurgo:1980:RUV

[MPPF80] Giorgio Morpurgo, Marcella Nazzaro Porro, Siro Passi, and Corrado Fanelli. The role of u.v. light in the control of melanogenesis. Journal of Theoretical Biology, 83(2):247–254, March 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251938090291X.

Maurer:1986:DHC

[MR86] Calvin R. Maurer and David M. Regen. Dependence of heart chamber dimensions and dynamics on chamber demands and myocardial properties. Journal of Theoretical Biology, 120(1): 1–29, May 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519386800157.

Malev:1989:TME

[MR89] Valery V. Malev and Anatoly I. Rusanov. Thermodynamics, mechanical and electrical properties of biomembranes. Journal of Theoretical Biology, 136(3):295–307, February 8, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389801651.

Matela:1983:CSC

[MRB83] Raymond J. Matela, Robert Ransomt, and M. A. Bowles. Computer simulation of compartment maintenance in the drosophila wing imaginal disc. Journal of Theoretical Biology, 103(3):357– 378, August 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383902928.

Margalit:1984:IDR

[MRG84] H. Margalit, R. F. Rosenberger, and N. B. Grover. Initiation of DNA replication in bacteria: Analysis of an autorepressor control model. Journal of Theoretical Biology, 111(1):183–199, November 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384802040.

Murray:1980:CCA

[MRR80] W. J. Murray, K. W. Reed, and E. B. Roche. Conformations of carnitine and acetylcarnitine and the relationship to mitochondrial transport of fatty acids. Journal of Theoretical Biology, 82(4):559–572, February 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380901782.

Machlup:1980:DOL

[MS80] S. Machlup and T. J. Sluckin. Driven oscillations of a limit-cycle oscillator. Journal of Theoretical Biology, 84(1):119–134, May 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938081040X.

Maruani:1983:SFE

[MS83] P. Maruani and D. Schwartz. Sterility and fecundability estimation. Journal of Theoretical Biology, 105(2):211–219, ???? 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383800034.

Murase:1986:MFM

[MS86a] M. Murase and H. Shimizu. A model of flagellar movement based on cooperative dynamics of dynein-tubulin cross-bridges. Journal of Theoretical Biology, 119(4):409–433, April 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801928.

Murray:1986:TRC

[MS86b] J. D. Murray and D. A. Smith. Theory of the rotational contribution to facilitated diffusion. Journal of Theoretical Biology, 118(2):231–246, January 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801369.

Markevich:1989:PRA

[MS89a] N. I. Markevich and E. E. Sel’kov. Parametric resonance and amplification in excitable membranes. The Hodgkin–Huxley model. Journal of Theoretical Biology, 140(1):27–38, September 11, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938980027X.

Myerscough:1989:MSU

[MS89b] M. R. Myerscough and M. A. Swan. A model for swimming unipolar spirilla. Journal of Theoretical Biology, 139(2):201–218, July 21, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519389801006.

MacMahon:1981:OCA

[MSA^{+}81] James A. MacMahon, David J. Schimpf, Douglas C. Andersen, Kimberly G. Smith, and Robert L. Bayn. An organism-centered approach to some community and ecosystem concepts. Journal of Theoretical Biology, 88(2):287–307, January 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381900771.

Moldawer:1984:RAL

[MSBB84] Lyle L. Moldawer, Javier Sobrado, George L. Blackburn, and Bruce R. Bistrian. A rationale for administering leukocyte endogenous mediator to protein malnourished, hospitalized patients. Journal of Theoretical Biology, 106(2):119–133, January 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519384900134.

Merajver:1981:TCO

[MSS81] Sofia D. Merajver, James P. Sheridan, and Eduardo N. Siguel. Theory of the conformational order of single component sphingomyelin bilayers. Journal of Theoretical Biology, 93(4):737–755, December 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381903374.

Meredith:1981:SSA

[MT81] David Meredith and Jack Tomlinson. Simultaneous sensory adaptation to light. Journal of Theoretical Biology, 91(3):397–415, August 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381902642.

Mierts:1984:TSA

[MT84] Stanislav Mierts and Magdaléna Trebatická. Theoretical study on the alkylation and protonation of the nucleic acid bases. Journal of Theoretical Biology, 108(4):509–517, June 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384800764.

Mintz:1986:EASa

[MTM86a] E. Mintz, S. R. Thomas, and D. C. Mikulecky. Exploration of apical sodium transport mechanisms in an epithelial model by network thermodynamic simulation of the effect of mucosal sodium depletion: I. Comparison of three different apical sodium permeability expressions. Journal of Theoretical Biology, 123(1):1–19, November 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519386802314.

Mintz:1986:EASb

[MTM86b] E. Mintz, S. R. Thomas, and D. C. Mikulecky. Exploration of apical sodium transport mechanisms in an epithelial model by network thermodynamic simulation of the effect of mucosal sodium depletion: II. An apical sodium channel and amiloride blocking. Journal of Theoretical Biology, 123(1):21–34, November 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386802326.

Massaldi:1981:TBI

[MUC81] H. A. Massaldi, G. Urrutia, and E. G. Cerrella. Theoretical basis for the interpretation of surface permeability data by a transient method. Journal of Theoretical Biology, 93(2):313–323, November 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381901077.

Matsumoto:1985:MHH

[MUI85] Gen Matsumoto, Masashi Urayama, and Michinori Ichikawa. Modified Hodgkin–Huxley gating kinetics of sodium activation in giant axons of squid (Doryteuthis bleekeri). Journal of Theoretical Biology, 112(4):695–705, February 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800540.

Matsumoto:1986:PPW

[MUK86] Kenji Matsumoto, Tetsuo Ueda, and Yonosuke Kobatake. Propagation of phase wave in relation to tactic responses by the plasmodium of Physarum polycephalum. Journal of Theoretical Biology, 122(3):339–345, October 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801254.

Matsumoto:1988:RTO

[MUK88] Kenji Matsumoto, Tetsuo Ueda, and Yonosuke Kobatake. Reversal of thermotaxis with oscillatory stimulation in the plasmodium of Physarum polycephalum. Journal of Theoretical Biology, 131(2):175–182, March 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388802352.

Muller:1985:TBE

[Mul85] Anthonie W. J. Muller. Thermosynthesis by biomembranes: Energy gain from cyclic temperature changes. Journal of Theoretical Biology, 115(3):429–453, August 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385802022.

Muller:1987:OPA

[Mul87] M. Muller. Optimization principles applied to the mechanism of neurocranium levation and mouth bottom depression in bony fishes (halecostomi). Journal of Theoretical Biology, 126(3):343– 368, June 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387802412.

Murray:1981:PPF

[Mur81] J. D. Murray. A pre-pattern formation mechanism for animal coat markings. Journal of Theoretical Biology, 88(1):161–199, January 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381903349.

Murray:1982:PST

[Mur82] J. D. Murray. Parameter space for Turing instability in reaction diffusion mechanisms: a comparison of models. Journal of Theoretical Biology, 98(1):143–163, September 7, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382900637.

Murphy:1983:NGM

[Mur83] Lea F. Murphy. A nonlinear growth mechanism in size structured population dynamics. Journal of Theoretical Biology, 104(4):493– 506, October 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383902424.

Murakami:1985:MAA

[Mur85] Makoto Murakami. Mutations affecting the 12th and 61st amino acids of p21 protein result in decreased probability of β-turn occurrence around the mutation positions: a prediction. Journal of Theoretical Biology, 114(2):193–198, May 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938580103X.

Murakami:1987:CAA

[Mur87] Makoto Murakami. Critical amino acids of p21 protein are located within β-turns: Further evaluation. Journal of Theoretical Biology, 128(3):339–347, October 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387800760.

Muller:1988:MAE

[MV88a] M. Muller and J. H. G. Verhagen. A mathematical approach enabling the calculation of the total endolymph flow in the semicircular ducts. Journal of Theoretical Biology, 134(4):503–529, October 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519388800547.

Muller:1988:NQM

[MV88b] M. Muller and J. H. G. Verhagen. A new quantitative model of total endolymph flow in the system of semicircular ducts. Journal of Theoretical Biology, 134(4):473–501, October 21, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388800535.

Metz:1989:ASC

[MVV89] J. A. J. Metz, J. Verboom, and F. Van den Bosch. The average size of chain-like planktonic algal colonies. Journal of Theoretical Biology, 140(1):19–26, September 11, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389800268.

MacDonald:1981:MAP

[MW81] N. MacDonald and A. R. Watkinson. Models of an annual plant population with a seedbank. Journal of Theoretical Biology, 93(3):643–653, December 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902265.

Markowsky:1983:MII

[MW83a] George Markowsky and Andrew Wohlgemuth. Mathematical immunogenetics II antibody incidence structure. Journal of Theoretical Biology, 102(3):425–437, June 7, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383903788.

McGraw:1983:SPG

[MW83b] James B. McGraw and R. D. Wulff. The study of plant growth: a link between the physiological ecology and population biology of plants. Journal of Theoretical Biology, 103(1):21–28, July 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383901960.

Murata:1989:MBF

[MYS89] T. Murata, M. Yano, and H. Shimizu. A model for bacterial flagellar motor: Free energy transduction and self-organization of rotational motion. Journal of Theoretical Biology, 139(4):531– 559, August 22, 1989. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519389800694.

Mahaffy:1989:MIR

[MZ89] Joseph M. Mahaffy and Judith W. Zyskind. A model for the initiation of replication in Escherichia coli. Journal of Theoretical Biology, 140(4):453–477, October 23, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389801092.

Meicenheimer:1989:CGP

[MZM89] Roger D. Meicenheimer and Beata Zagorska-Marek. Consideration of the geometry of the phyllotaxic triangular unit and discontinuous phyllotactic transitions. Journal of Theoretical Biology, 139(3):359–368, August 9, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389802140.

Nagata:1984:IMO

[NA84] Chikayoshi Nagata and Misako Aida. Ab initio molecular orbital study of the interaction of Li^{+}, Na^{+} and K^{+} with the pore components of ion channels: Consideration of the size, structure and selectivity of the pore of the channels. Journal of Theoretical Biology, 110(4):569–585, October 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801460.

Nagasawa:1981:ASM

[Nag81] Masao Nagasawa. An application of the segregation model for septation of Escherichia coli. Journal of Theoretical Biology, 90(4):445–455, June 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519381902988.

Nagylaki:1982:GIP

[Nag82] Thomas Nagylaki. Geographical invariance in population genetics. Journal of Theoretical Biology, 99(1):159–172, November 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382903964.

Nagorcka:1986:RRD

[Nag86] B. N. Nagorcka. The role of a reaction–diffusion system in the initiation of skin organ primordia. I. The first wave of initiation. Journal of Theoretical Biology, 121(4):449–475, August 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386801023.

Nagorcka:1988:PFM

[Nag88] B. N. Nagorcka. A pattern formation mechanism to control spatial organization in the embryo of Drosophila melanogaster. Journal of Theoretical Biology, 132(3):277–306, June 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388802169.

Nagorcka:1989:WIP

[Nag89] B. N. Nagorcka. Wavelike isomorphic prepatterns in development. Journal of Theoretical Biology, 137(2):127–162, March 21, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519389802024.

Namba:1980:DDD

[Nam80] Toshiyuki Namba. Density-dependent dispersal and spatial distribution of a population. Journal of Theoretical Biology, 86(2):351–363, September 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380900119.

Namba:1984:CCE

[Nam84] Toshiyuki Namba. Competitive co-existence in a seasonally fluctuating environment. Journal of Theoretical Biology, 111 (2):369–386, November 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519384802167.

Nanjundiah:1986:HRD

[Nan86] Vidyanand Nanjundiah. How rapidly do uncoupled oscillators desynchronize? Journal of Theoretical Biology, 121(3):375–379, August 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519386801175.

Nicolini:1983:PCM

[NBZ^{+}83] C. Nicolini, A. Belmont, S. Zietz, A. Maura, and … G. Brambilla. Physico-chemical model for DNA alkaline elution: New experimental evidence and differential role of DNA length, chain flexibility and superpacking. Journal of Theoretical Biology, 100(2):341–357, January 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383903570.

Newman:1989:CON

[NC89] Jonathan A. Newman and Thomas Caraco. Co-operative and non-co-operative bases of food-calling. Journal of Theoretical Biology, 141(2):197–209, November 21, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389800189.

Naora:1982:IEE

[ND82] H. Naora and N. J. Deacon. Implication of the effect of extragenic territorial DNA sequences on a mechanism involving switch-on/ off of neighbouring genes by transposable elements in eukaryotes. Journal of Theoretical Biology, 95(3):601–606, April 7, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382900376.

Naora:1980:FSS

[NDB80] H. Naora, N. J. Deacon, and D. Buckle. Further studies on similarity of terminal nucleotide sequences of eukaryote mRNA — a possible mechanism of RNA splicing. Journal of Theoretical Biology, 85(1):83–96, July 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380902829.

Naora:1983:TCA

[NDB83] H. Naora, N. J. Deacon, and A. W. Braithwaite. Territorial confrontation among genes and activation of a cellular oncogene. Journal of Theoretical Biology, 105(4):715–721, December 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383902291.

Nee:1989:ACE

[Nee89a] Sean Nee. Antagonistic co-evolution and the evolution of genotypic randomization. Journal of Theoretical Biology, 140(4):499– 518, October 23, 1989. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519389801110.

Nee:1989:DHR

[Nee89b] Sean Nee. Does Hamilton’s rule describe the evolution of reciprocal altruism? Journal of Theoretical Biology, 141(1):81–91, November 8, 1989. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519389800104.

Nee:1989:ESR

[Nee89c] Sean Nee. On the evolution of sex in RNA viruses. Journal of Theoretical Biology, 138(3):407–412, June 3, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389802012.

Nesterenko:1984:CIR

[Nes84] V. G. Nesterenko. A concept of immune regulation of somatic cell differentiation. Journal of Theoretical Biology, 107(3):443– 456, April 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384801010.

Neville:1988:PCM

[Nev88] A. C. Neville. A pipe-cleaner molecular model for morphogenesis of helicoidal plant cell walls based on hemicellulose complexity. Journal of Theoretical Biology, 131(2):243–254, March 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388802418.

Newman:1980:SES

[New80] William I. Newman. Some exact solutions to a non-linear diffusion problem in population genetics and combustion. Journal of Theoretical Biology, 85(2):325–334, July 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380900247.

Newman:1983:LTB

[New83] William I. Newman. The long-time behavior of the solution to a non-linear diffusion problem in population genetics and combustion. Journal of Theoretical Biology, 104(4):473–484, October 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383902400.

Newman:1988:SSA

[NFP88] Stuart A. Newman, H. L. Frisch, and J. K. Percus. On the stationary state analysis of reaction–diffusion mechanisms for biological pattern formation. Journal of Theoretical Biology, 134(2):183–197, September 21, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388802017.

Nur:1984:PPH

[NH84] Nadav Nur and Oren Hasson. Phenotypic plasticity and the handicap principle. Journal of Theoretical Biology, 110(2):275–297, September 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384800594.

Nagano:1988:PTD

[NHT88] Kozo Nagano, Michal Harel, and Mikio Takezawa. Prediction of three-dimensional structure of Escherichia coli ribosomal RNA. Journal of Theoretical Biology, 134(2):199–256, September 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388802029.

Nickerson:1984:HRP

[Nic84] Kenneth W. Nickerson. A hypothesis on the role of pressure in the origin of life. Journal of Theoretical Biology, 110(3):487–499, October 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384801885.

Ninio:1983:BG

[Nin83] Jacoues Ninio. The brain as a geometer. Journal of Theoretical Biology, 103(1):1–9, July 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383901947.

Nitschmann:1986:CCC

[Nit86] Wolfgang H. Nitschmann. Chemiosmotic coupling in cyanobacteria: ATP synthesis in a two-membrane system. Journal of Theoretical Biology, 122(4):409–419, October 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801825.

Niven:1983:TDA

[Niv83] B. S. Niven. Two different animals may not have the same environment. Journal of Theoretical Biology, 105(2):369–370, ???? 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383800149.

Nunogaki:1988:HOF

[NK88] Kazuki Nunogaki and Michiki Kasai. The H + OH^{−} flux localizes around the channel mouth in buffered solution. Journal of Theoretical Biology, 134(3):403–415, October 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388800705.

Novak:1986:LCT

[NL86] B. Novák and E. László. A limit cycle type oscillation model for the cell cycle regulated by carbon dioxide. Journal of Theoretical Biology, 120(3):309–320, June 7, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938680203X.

Namba:1980:SDC

[NM80] Toshiyuki Namba and Masayasu Mimura. Spatial distribution of competing populations. Journal of Theoretical Biology, 87(4): 795–814, ???? 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519380901186.

Nachbar:1981:GCG

[NM81] Robert B. Nachbar and Thomas Hellman Morton. A gas chromatographic (GLPC) model for the sense of smell. Variation of olfactory sensitivity with conditions of stimulation. Journal of Theoretical Biology, 89(3):387–407, April 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381903581.

Nagorcka:1982:RRD

[NM82] B. N. Nagorcka and J. R. Mooney. The role of a reaction– diffusion system in the formation of hair fibres. Journal of Theoretical Biology, 98(4):575–607, October 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382901394.

Nagorcka:1985:RRD

[NM85] B. N. Nagorcka and J. R. Mooney. The role of a reaction– diffusion system in the initiation of primary hair follicles. Journal of Theoretical Biology, 114(2):243–272, May 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801065.

Naora:1985:FIT

[NMB85] H. Naora, K. Miyahara, and A. W. Braithwaite. Further implications of the territorial effects of genes: chromosomal translocation and oncogene activation. Journal of Theoretical Biology, 116(2):313–320, September 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385802708.

Nakamoto:1984:RAA

[NMM84] Tetsuo Nakamoto, Mark McCroskey, and Henry M. Mallek. The role of ascorbic acid deficiency in human gingivitis — a new hypothesis. Journal of Theoretical Biology, 108(2):163–171, May 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384800636.

Nagorcka:1987:CSP

[NMM87] B. N. Nagorcka, V. S. Manoranjan, and J. D. Murray. Complex spatial patterns from tissue interactions — an illustrative model. Journal of Theoretical Biology, 128(3):359–374, October 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519387800784.

Nagano:1982:PTS

[NMMO82] Kozo Nagano, Shuichi Miyamoto, Makiko Matsumura, and Iwao Ohtsuki. Prediction of a triple-stranded coiled-coil region in tropomyosin-troponin T complex. Journal of Theoretical Biology, 94(4):743–782, February 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382900777.

Noda:1982:PLR

[Nod82] Haruhiko Noda. Probability of life. Rareness of realization in evolution. Journal of Theoretical Biology, 95(1):145–150, March 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382902934.

Norwich:1982:CDC

[Nor82] Kenneth H. Norwich. Can diffusion coefficients be estimated from plasma clearance curves of intact animals? Journal of Theoretical Biology, 95(2):399–407, March 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902545.

Norwich:1983:PDR

[Nor83] Kenneth H. Norwich. To perceive is to doubt: The relativity of perception. Journal of Theoretical Biology, 102(2):175–190, May 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383903582.

Norris:1989:PFC

[Nor89] Vic Norris. Phospholipid flip-out controls the cell cycle of Escherichia coli. Journal of Theoretical Biology, 139(1):117–128, July 10, 1989. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251938980061X.

Neal:1981:CAI

[NP81] J. V. Neal and C. S. Potten. A critical assessment of the intestinal proliferon hypothesis. Journal of Theoretical Biology, 91(1):63– 70, July 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251938190374X.

Neal:1982:IPH

[NP82] J. V. Neal and C. S. Porren. The intestinal proliferon hypothesis. Journal of Theoretical Biology, 97(3):539–540, August 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382903812.

Nussinov:1984:FTL

[NP84a] Ruth Nussinov and George Pieczenik. Folding of two large nucleotide chains. Journal of Theoretical Biology, 106(3):261–273, February 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519384900304.

Nussinov:1984:SCC

[NP84b] Ruth Nussinov and George Pieczenik. Structural and combinatorial constraints on base pairing in large nucleotide sequences. Journal of Theoretical Biology, 106(3):245–259, February 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519384900298.

Nalebuff:1985:AEW

[NR85] Barry Nalebuff and John Riley. Asymmetric equilibria in the war of attrition. Journal of Theoretical Biology, 113(3):517–527, April 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385800369.

Norwich:1982:PFP

[NS82] Kenneth H. Norwich and Sidney Siu. Power functions in physiology and pharmacology. Journal of Theoretical Biology, 95(2):387– 398, March 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382902533.

Nedbal:1986:HLD

[NS86] L. Nedbal and V. Szöcs. How long does excitonic motion in the photosynthetic unit remain coherent? Journal of Theoretical Biology, 120(4):411–418, June 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386800352.

Nowak:1989:ETR

[NS89a] Martin Nowak and Peter Schuster. Error thresholds of replication in finite populations mutation frequencies and the onset of Muller’s ratchet. Journal of Theoretical Biology, 137(4):375– 395, April 20, 1989. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519389800360.

Nowak:1989:OER

[NS89b] Martin Nowak and Karl Sigmund. Oscillations in the evolution of reciprocity. Journal of Theoretical Biology, 137(1):21–26, March 7, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519389801468.

Norris:1988:SCF

[NSCH88] V. Norris, S. J. Seror, S. Casaregola, and I. B. Holland. A single calcium flux triggers chromosome replication, segregation and septation in bacteria: a model. Journal of Theoretical Biology, 134(3):341–350, October 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388800651.

Nichol:1981:CLI

[NSJT81] L. W. Nichol, M. J. Sculley, P. D. Jeffrey, and R. D. Teasdale. Cross-linking interactions in systems of three reactants: Theory pertaining to their experimental elucidation. Journal of Theoretical Biology, 91(1):215–230, July 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381903817.

Nichol:1984:ISA

[NSJW84] L. W. Nichol, M. J. Sculley, P. D. Jeffrey, and D. J. Winzor. Indefinite self-association of a solute in linear and branched arrays. Journal of Theoretical Biology, 109(2):285–298, July 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384800077.

Naray-Szabo:1985:EET

[NSKS85] G. Náray-Szabó, G. Kramer, and I. Simon. Electrostatic effect of trypsin binding on the hydrogen exchange rate of bovine pancreatic trypsin inhibitor β-sheet NH’s. Journal of Theoretical Biology, 117(3):505–508, December 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801582.

Nussinov:1980:MRP

[NST80] R. Nussinov, J. L. Sussman, and E. N. Trifonov. MS2 RNA has a potential to form an unusually large number of stable hairpins. Journal of Theoretical Biology, 85(3):481–486, August 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380903215.

Nichol:1982:CDE

[NSW82] L. W. Nichol, M. J. Sculley, and D. J. Winzor. The composition-dependence of the extent of reaction between a multivalent acceptor and a bivalent ligand: Implications of self-interaction of the ligand in precipitin effects. Journal of Theoretical Biology, 96(4):723–740, June 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902399.

Nakagawa:1982:IMO

[NU82] Setsuko Nakagawa and Hideaki Umeyama. Ab initio molecular orbital study on the effects of zinc ion, its ligands and ionic amino acid residues for proton transfer energetics between glu 270 and Zn co-ordinated water molecule in carboxypeptidase a. Journal of Theoretical Biology, 96(3):473–493, June 7, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382901229.

Nunney:1980:DCI

[Nun80] Len Nunney. Density compensation, isocline shape and singlelevel competition models. Journal of Theoretical Biology, 86 (2):323–349, September 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519380900107.

Nunes:1981:SCA

[Nun81] Marlies Vaz Nunes. A ‘simple clock’ approach of circadian rhythms: An easy way to predict the clock’s singularity. Journal of Theoretical Biology, 92(3):227–239, October 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902897.

Nussinov:1980:SAC

[Nus80a] Ruth Nussinov. Strong adenine clustering in nucleotide sequences. Journal of Theoretical Biology, 85(2):285–291, July 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380900211.

Nussinov:1980:TDS

[Nus80b] Ruth Nussinov. Three dimensional structure and sequence homology determine splicing sites in eukaryotic precursor RNA. Journal of Theoretical Biology, 83(4):647–662, April 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380901939.

Nussinov:1982:SII

[Nus82a] Ruth Nussinov. Some indications for inverse DNA duplication. Journal of Theoretical Biology, 95(4):783–791, April 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382903538.

Nussinov:1982:STE

[Nus82b] Ruth Nussinov. Splicing and transport of eukaryotic mRNAs: a theoretical model. Journal of Theoretical Biology, 97(2):129– 140, July 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382900947.

Nussinov:1983:ECS

[Nus83] Ruth Nussinov. An efficient code searching for sequence homology and DNA duplication. Journal of Theoretical Biology, 100(2):319–328, January 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519383903557.

Nussinov:1985:LHC

[Nus85] Ruth Nussinov. Large helical conformational deviations from ideal b-DNA and prokaryotic regulatory sites. Journal of Theoretical Biology, 115(2):179–189, July 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800941.

Nussinov:1986:CES

[Nus86] Ruth Nussinov. Compilation of eukaryotic sequences around transcription initiation sites. Journal of Theoretical Biology, 120(4):479–487, June 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519386800418.

Nussinov:1987:TMB

[Nus87] Ruth Nussinov. Theoretical molecular biology: Prospectives and perspectives. Journal of Theoretical Biology, 125(2):219–235, March 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387800437.

Nussinov:1988:CQN

[Nus88] Ruth Nussinov. Conserved quartets near 5’ intron junctions in primate nuclear pre-mRNA. Journal of Theoretical Biology, 133(1):73–84, July 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519388800250.

Nichol:1985:BLA

[NW85] Lawrence W. Nichol and Donald J. Winzor. The binding of a ligand to an acceptor undergoing indefinite self-association. Journal of Theoretical Biology, 117(4):597–608, December 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385802411.

Obendorf:1981:RMF

[Obe81] P. Obendorf. A rotating myosin filament theory of muscular contraction. Journal of Theoretical Biology, 93(3):667–680, December 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381902289.

Olson:1981:CMD

[OC81] W. H. Olson and R. B. Cumming. Chemical mutagens: Dosimetry, Haber’s rule and linear systems. Journal of Theoretical Biology, 91(3):383–395, August 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902630.

Oster:1989:VFG

[OCMP89] George F. Oster, Louis Y. Cheng, Hsiao-Ping H. Moore, and Alan S. Perelson. Vesicle formation in the Golgi apparatus. Journal of Theoretical Biology, 141(4):463–504, December 19, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519389802310.

OGrady:1984:ETT

[O’G84] Richard T. O’Grady. Evolutionary theory and teleology. Journal of Theoretical Biology, 107(4):563–578, April 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801332.

Okamoto:1983:DBC

[OH83] Masahiro Okamoto and Katsuya Hayashi. Dynamic behavior of cyclic enzyme systems. Journal of Theoretical Biology, 104(4):591–598, October 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383902473.

Okamoto:1984:FCM

[OH84] Masahiro Okamoto and Katsuya Hayashi. Frequency conversion mechanism in enzymatic feedback systems. Journal of Theoretical Biology, 108(4):529–537, June 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384800788.

Okamoto:1985:CMB

[OH85] Masahiro Okamoto and Katsuya Hayashi. Control mechanism for a bacterial sugar-transport system: Theoretical hypothesis. Journal of Theoretical Biology, 113(4):785–790, April 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801946.

Ohta:1987:MEA

[Oht87] Tomoko Ohta. A model of evolution for accumulating genetic information. Journal of Theoretical Biology, 124(2):199–211, January 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251938780262X.

Otsuka:1982:CBS

[OK82] Jinya Otsuka and Takashi Kunisawa. Characteristic base sequence patterns of promoter and terminator sites in ϕX174 and fd phage DNAs. Journal of Theoretical Biology, 97(3):415–436, August 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382903745.

Ohshima:1987:ERI

[OK87a] Hiroyuki Ohshima and Tamotsu Kondo. Electrostatic repulsion of ion penetrable charged membranes: Role of Donnan potential. Journal of Theoretical Biology, 128(2):187–194, September 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519387801686.

Ohshima:1987:PPD

[OK87b] Hiroyuki Ohshima and Tamotsu Kondo. Potential and pH distribution across a membrane with a surface layer of ionizable groups. Journal of Theoretical Biology, 124(2):191–198, January 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519387802618.

Opadia-Kadima:1987:HSM

[OK87c] G. Z. Opadia-Kadima. How the slot machine led biologists astray. Journal of Theoretical Biology, 124(2):127–135, January 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519387802588.

Okamoto:1980:DBE

[OKT^{+}80] Masahiro Okamoto, Atsuko Katsurayama, Michiko Tsukiji, Yoichi Aso, and Katsuya Hayashi. Dynamic behavior of enzymatic system realizing two-factor model. Journal of Theoretical Biology, 83(1):1–16, March 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380903690.

Omholt:1986:HPW

[OL86] Stig W. Omholt and Knut Lønvik. Heat production in the winter cluster of the honeybee, Apis mellifera. A theoretical study. Journal of Theoretical Biology, 120(4):447–456, June 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386800388.

Olavarria:1982:DCB

[Ola82] José Manuel Olavarría. Does the coupling between conformational fluctuation and enzyme catalysis involve a true phase transfer catalysis? Journal of Theoretical Biology, 99(1):21–30, November 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382903873.

Olavarria:1986:SMC

[Ola86] José M. Olavarría. A systematic method for the compilation of the King and Altman graphs of a steady-state enzyme model suitable for manual or computer use. Journal of Theoretical Biology, 122(3):269–275, October 7, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801205.

Olivier:1985:UKC

[Oli85] Thomas J. Olivier. Use of keyed character string data structures and operators in models of primate groups. Journal of Theoretical Biology, 115(4):539–549, August 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801399.

Olson:1987:PLC

[Ols87] Arthur David Olson. Picking lowest-cost measurements for identifying unknowns. Journal of Theoretical Biology, 128(1):1–9, September 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387800280.

Omholt:1987:TWC

[Omh87a] Stig W. Omholt. Thermoregulation in the winter cluster of the honeybee, Apis Mellifera. Journal of Theoretical Biology, 128(2):219–231, September 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387801704.

Omholt:1987:WHR

[Omh87b] Stig W. Omholt. Why honeybees rear brood in winter. A theoretical study of the water conditions in the winter cluster of the honeybee, Apis mellifera. Journal of Theoretical Biology, 128(3):329– 337, October 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387800759.

Omholt:1988:DPH

[Omh88a] Stig W. Omholt. Drone production in honeybee colonies: Controlled by a longlasting inhibitory pheromone from the drones? Journal of Theoretical Biology, 134(3):309–318, October 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388800614.

Omholt:1988:RBW

[Omh88b] Stig W. Omholt. Relationships between worker longevity and the intracolonial population dynamics of the honeybee. Journal of Theoretical Biology, 130(3):275–284, February 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388800304.

Omoto:1989:MCE

[Omo89] Charlotte K. Omoto. Mechanochemical coupling in eukaryotic flagella. Journal of Theoretical Biology, 137(2):163–169, March 21, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519389802036.

Onda:1980:TCF

[Ond80] Hiromasa Onda. A theoretical consideration of fundamental biological phenomena on cell-specific mitosis-inhibiting protein excretion hypothesis. Journal of Theoretical Biology, 85(4):771–787, August 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519380902714.

Oberley:1984:RSD

[OO84a] Larry W. Oberley and Terry D. Oberley. The role of superoxide dismutase and gene amplification in carcinogenesis. Journal of Theoretical Biology, 106(3):403–422, February 7, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519384900389.

Owen:1984:MPM

[OO84b] Robin E. Owen and A. R. G. Owen. Mathematical paradigms for mimicry: Recurrent sampling. Journal of Theoretical Biology, 109(2):217–247, July 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519384800041.

Oonuki:1982:MLT

[Oon82] Makoto Oonuki. Macroscopic law for time and activity in a neural system. Journal of Theoretical Biology, 94(1):191–196, January 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251938290340X.

Orgel:1986:RCO

[Org86] Leslie E. Orgel. RNA catalysis and the origins of life. Journal of Theoretical Biology, 123(2):127–149, November 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801497.

Orlove:1981:EEL

[Orl81a] Michael J. Orlove. The elevator effect or a lift in a lift: How a locus in neutral equilibrium can provide a free ride for a neutral allele at another locus. Journal of Theoretical Biology, 90(1):81– 100, May 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381901235.

Orlove:1981:MSR

[Orl81b] Michael J. Orlove. A model of the sex ratio in flying ants with an unexpected equilibrium and an expected ESS. Journal of Theoretical Biology, 93(3):523–532, December 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902174.

Orlove:1982:SBA

[Orl82] Michael J. Orlove. A systematic bias against altruism due to an inevitable hitchhiking process in the hymenoptera (a sequel to the elevator effect). Journal of Theoretical Biology, 96(2):107– 127, May 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382902156.

Osuji:1984:OMG

[Osu84] G. O. Osuji. An oscillatory mechanism for the γ-glutamyl transpeptidase-mediated translocation of amino acids across the cell membrane. Journal of Theoretical Biology, 109(1):1–15, July 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384801071.

Othmer:1986:NFM

[Oth86] H. G. Othmer. On the Newman-frisch model of limb chondrogenesis. Journal of Theoretical Biology, 121(4):505–508, August 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386801059.

OQuigley:1989:ESD

[OWN89] John O’Quigley, Patrick Ward, and Richard Nash. Estimating the size of the dividing stem cell pool after allogeneic bone marrow transplantation. Journal of Theoretical Biology, 141(1):73–79, November 8, 1989. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519389800098.

Ponnuswamy:1982:MNT

[PA82] P. K. Ponnuswamy and A. Anukanth. Modified nucleotides: Their conformational characteristics. Journal of Theoretical Biology, 96(2):233–251, May 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519382902235.

Polgar:1986:BDC

[PA86] L. Polgár and B. Asbóth. The basic difference in catalyses by serine and cysteine proteinases resides in charge stabilization in the transition state. Journal of Theoretical Biology, 121(3):323– 326, August 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519386801114.

Pahor:1985:PTL

[Pah85a] S. Pahor. On the −3/2 power thinning law in plant ecology. Journal of Theoretical Biology, 112(3):535–537, February 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800205.

Pahor:1985:PEM

[Pah85b] S. Pahor. Plant ecology models. Journal of Theoretical Biology, 117(1):159–160, November 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801703.

Paietta:1982:PEC

[Pai82] J. Paietta. Photooxidation and the evolution of circadian rhythmicity. Journal of Theoretical Biology, 97(1):77–82, July 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382902788.

Papentin:1980:OCG

[Pap80] Frank Papentin. On order and complexity. I. General considerations. Journal of Theoretical Biology, 87(3):421–456, December 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380902301.

Papentin:1982:OCI

[Pap82] Frank Papentin. On order and complexity. II. Application to chemical and biochemical structures. Journal of Theoretical Biology, 95(2):225–245, March 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902417.

Papageorgiou:1984:HPC

[Pap84] S. Papageorgiou. A hierarchical polar coordinate model for epimorphic regeneration. Journal of Theoretical Biology, 109(4):533– 554, August 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384801575.

Papageorgiou:1989:CPC

[Pap89] S. Papageorgiou. Cartesian or polar co-ordinates in pattern formation? Journal of Theoretical Biology, 141(2):281–283, November 21, 1989. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519389800220.

Parker:1982:WTM

[Par82] G. A. Parker. Why are there so many tiny sperm? Sperm competition and the maintenance of two sexes. Journal of Theoretical Biology, 96(2):281–294, May 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902259.

Parker:1983:ARE

[Par83a] G. A. Parker. Arms races in evolution — an ESS to the opponent-independent costs game. Journal of Theoretical Biology, 101(4):619–648, April 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251938390019X.

Parry:1983:ICG

[Par83b] G. D. Parry. The influence of the cost of growth on ectotherm metabolism. Journal of Theoretical Biology, 101(3):453–477, April 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383901509.

Park:1985:DHL

[Par85] D. Park. Does Horton’s law of branch length apply to open branching systems? Journal of Theoretical Biology, 112(2):299– 313, January 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385802897.

Pate:1984:ORP

[Pat84] Edward Pate. The organizer region and pattern regulation in amphibian embryos. Journal of Theoretical Biology, 111(2):387–396, November 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384802179.

Painter:1981:RTO

[PB81] Page R. Painter and Richard D. Bliss. Reconsideration of the theory of oscillatory repression. Journal of Theoretical Biology, 90(2):293–298, May 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381900485.

Parker:1983:SIA

[PC83] G. A. Parker and P. Courtney. Seasonal incidence: Adaptive variation in the timing of life history stages. Journal of Theoretical Biology, 105(1):147–155, November 7, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383904307.

Pate:1986:MIM

[PC86] Edward Pate and Roger Cooke. A model for the interaction of muscle cross-bridges with ligands which compete with ATP. Journal of Theoretical Biology, 118(2):215–230, January 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801357.

Paul:1983:TLR

[PCTF83] R. Paul, R. Chatterjee, J. A. Tuszyński, and O. G. Fritz. Theory of long-range coherence in biological systems. I. The anomalous behaviour of human erythrocytes. Journal of Theoretical Biology, 104(2):169–185, September 21, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383904095.

Paltridge:1984:DSY

[PDC84] G. W. Paltridge, J. V. Denholm, and D. J. Connor. Determinism, senescence and the yield of plants. Journal of Theoretical Biology, 110(3):383–398, October 7, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801812.

Pineda:1988:ESR

[PDCD88] F. D. Pineda, C. L. De Pablo, M. A. Casado, and J. M. De Miguel. Ecological structures recognized by means of entropy analysis: Assessment of differences between entropy values. Journal of Theoretical Biology, 135(3):283–293, December 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388802443.

Papadopulos-Eleopulos:1982:MT

[PE82] Eleni Papadopulos-Eleopulos. A mitotic theory. Journal of Theoretical Biology, 96(4):741–757, June 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902405.

Pease:1988:BPT

[Pea88a] Craig M. Pease. Biases in the per-taxon origination and extinction rates of fossil taxa. Journal of Theoretical Biology, 130(1): 9–30, January 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519388801619.

Pease:1988:BSC

[Pea88b] Craig M. Pease. Biases in the survivorship curves of fossil taxa. Journal of Theoretical Biology, 130(1):31–48, January 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388801620.

Pease:1988:BTE

[Pea88c] Craig M. Pease. Biases in the total extinction rates of fossil taxa. Journal of Theoretical Biology, 130(1):1–7, January 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388801607.

Pease:1988:CGD

[Pea88d] Craig M. Pease. On comparing the geologic durations of easily versus poorly fossilized taxa. Journal of Theoretical Biology, 133(2):255–257, July 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388800092.

Peczkis:1988:PBW

[Pec88] Jan Peczkis. Predicting body-weight distribution of mammalian genera in families and orders. Journal of Theoretical Biology, 132(4):509–510, June 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388800870.

Pedemonte:1988:KMI

[Ped88] Carlos Humberto Pedemonte. Kinetic mechanism of inhibition of the Na^{+}-pump and some of its partial reactions by external Na^{+} (Na_{o}^{+}). Journal of Theoretical Biology, 134(2):165–182, September 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519388802005.

Pellionisz:1984:CVM

[Pel84] András J. Pellionisz. Coordination: a vector-matrix description of transformations of overcomplete CNS coordinates and a tensorial solution using the Moore–Penrose generalized inverse. Journal of Theoretical Biology, 110(3):353–375, October 7, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801794.

Penninckx:1980:ESD

[Pen80] Michel Penninckx. Electrostatic stabilization of dipolar transition states in enzyme catalyzed transacylations. A theoretically plausible mechanism. Journal of Theoretical Biology, 82(3):541–543, February 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519380902568.

Penny:1982:TBC

[Pen82] David Penny. Towards a basis for classification: the incompleteness of distance measures, incompatibility analysis and phenetic classification. Journal of Theoretical Biology, 96(2):129–142, May 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382902168.

Percheron:1982:PMG

[Per82] G. Percheron. Principles and methods of the graph-theoretical analysis of natural binary arborescences. Journal of Theoretical Biology, 99(3):509–552, December 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902090.

Perry:1984:MPA

[Per84] David A. Perry. A model of physiological and allometric factors in the self-thinning curve. Journal of Theoretical Biology, 106(3):383–401, February 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519384900377.

Peterson:1980:LCI

[Pet80] Eric L. Peterson. A limit cycle interpretation of a mosquito circadian oscillator. Journal of Theoretical Biology, 84(2):281– 310, May 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519380800087.

Petkov:1982:EPS

[Pet82] D. D. Petkov. Enzyme peptide synthesis and semisynthesis: Kinetic and thermodynamic aspects. Journal of Theoretical Biology, 98(3):419–425, October 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251938290128X.

Peterson:1983:WNP

[Pet83] Jerry A. Peterson. The widespread nature of phenotypic variability in hepatomas and cell lines, in the form of a geometric series. Journal of Theoretical Biology, 102(1):41–53, May 7, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383902618.

Peusner:1983:HIE

[Peu83] Leonardo Peusner. Hierarchies of irreversible energy conversion systems: A network thermodynamic approach. I. Linear steady state without storage. Journal of Theoretical Biology, 102(1): 7–39, May 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383902606.

Peusner:1985:HIE

[Peu85] Leonardo Peusner. Hierarchies of irreversible energy conversion systems. II. Network derivation of linear transport equations. Journal of Theoretical Biology, 115(3):319–335, August 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385801958.

Peusner:1986:HEC

[Peu86] Leonardo Peusner. Hierarchies of energy conversion processes III. Why are Onsager equations reciprocal? The Euclidean geometry of fluctuation–dissipation space. Journal of Theoretical Biology, 122(2):125–155, September 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386800789.

Pawlowski:1989:BMC

[PF89] Piotr Pawlowski and Magdalena Fikus. Bioelectrorheological model of the cell 1. Analysis of stresses and deformations. Journal of Theoretical Biology, 137(3):321–337, April 6, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938980075X.

Palsson:1988:MMD

[PG88] Bernhard O. Palsson and Thomas M. Groshans. Mathematical modelling of dynamics and control in metabolic networks: VI. dynamic bifurcations in single biochemical control loops. Journal of Theoretical Biology, 131(1):43–53, March 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938880119X.

Peper:1987:AMT

[PGKE87] A. Peper, C. A. Grimbergen, J. W. Kraal, and J. H. Engelbart. An approach to the modeling of the tolerance mechanism in the drug effect. I: The drug effect as a disturbance of regulations. Journal of Theoretical Biology, 127(4):413–426, August 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938780139X.

Peper:1988:AMT

[PGKE88] A. Peper, C. A. Grimbergen, J. W. Kraal, and J. H. Engelbart. An approach to the modeling of the tolerance mechanism in the drug effect. II: On the implications of compensatory regulation.

Journal of Theoretical Biology, 132(1):29–41, May 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388801887.

Patarca:1987:VAH

[PH87a] Roberto Patarca and William A. Haseltine. Variation among the human T-cell lymphotropic virus type III (HTLV-III/LAV) strains. Journal of Theoretical Biology, 125(2):213–217, March 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519387800425.

Prajneshu:1987:PPM

[PH87b] Prajneshu and P. Holgate. A prey-predator model with switching effect. Journal of Theoretical Biology, 125(1):61–66, March 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519387801790.

Phillips:1989:CCA

[PH89] D. S. Phillips and S. F. Hubbard. Controlled covariance in the Atkinson-shorrocks model of multispecies competition. Journal of Theoretical Biology, 141(1):137–141, November 8, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938980013X.

Phillis:1981:TLC

[Phi81a] Yannis A. Phillis. Time limits in the control of pollution systems. Journal of Theoretical Biology, 93(3):541–546, December 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381902198.

Phipps:1981:ECP

[Phi81b] Michel Phipps. Entropy and community pattern analysis. Journal of Theoretical Biology, 93(1):253–273, November 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381900692.

Phillis:1982:TLC

[Phi82] Yannis A. Phillis. Time limits in the control of pollution systems: J. theor. Biol (1981) 93, 541–546. Journal of Theoretical Biology, 99(2):405, November 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382900145.

Pickett-Heaps:1986:MDS

[PHTCS86] Jeremy D. Pickett-Heaps, David H. Tippit, Stanley A. Cohn, and Timothy P. Spurck. Microtubule dynamics in the spindle. Theoretical aspects of assembly/disassembly reactions in vivo. Journal of Theoretical Biology, 118(2):153–169, January 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938680131X.

Pickard:1984:STR

[Pic84] William F. Pickard. The self-thinning rule. Journal of Theoretical Biology, 110(2):313–314, September 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384800624.

Pickard:1986:TSO

[Pic86] William F. Pickard. A theory of the saturable ohmic channel. Journal of Theoretical Biology, 122(1):1–6, September 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938680220X.

Pickard:1989:HMT

[Pic89] William F. Pickard. How might a tracheary element which is embolized by day be healed by night? Journal of Theoretical Biology, 141(2):259–279, November 21, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389800219.

Pinter:1983:EBL

[Pin83a] R. B. Pinter. The electrophysiological bases for linear and for nonlinear product term lateral inhibition and the consequences for wide field textured stimuli. Journal of Theoretical Biology, 105(2):

233–243, ???? 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519383800058.

Pinter:1983:PTN

[Pin83b] Robert B. Pinter. Product term nonlinear lateral inhibition enhances visual selectivity for small objects or edges. Journal of Theoretical Biology, 100(3):525–531, February 7, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383904447.

Pinter:1984:ARF

[Pin84] Robert B. Pinter. Adaptation of receptive field spatial organization via multiplicative lateral inhibition. Journal of Theoretical Biology, 110(3):435–444, October 7, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938480185X.

Prothero:1986:EMD

[PJ86] John Prothero and Klaus Dieter Jürgens. An energetic model of daily torpor in endotherms. Journal of Theoretical Biology, 121(4):403–415, August 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386800996.

Pokorny:1983:FER

[PJK^{+}83] J. Pokorný, A. Jandová, J. Kobilková, K. Heyberger, and T. Hraba. Fröhlich electromagnetic radiation from human leukocytes: Implications for leukocyte adherence inhibition test. Journal of Theoretical Biology, 102(2):295–305, May 21, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383903673.

Palsson:1984:MMDb

[PJL84] Bernhard O. Palsson, Robert Jamier, and Edwin N. Lightfoot. Mathematical modelling of dynamics and control in metabolic networks. II. Simple dimeric enzymes. Journal of Theoretical Biology, 111(2):303–321, November 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938480212X.

Parker:1980:ETS

[PK80] G. A. Parker and N. Knowlton. The evolution of territory sizesome ESS models. Journal of Theoretical Biology, 84(3):445– 476, June 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519380800142.

Perkkio:1982:EGA

[PK82] J. Perkkiö and R. Keskinen. Exponential growth and allometric equations. Journal of Theoretical Biology, 98(4):675–677, October 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251938290145X.

Perkkio:1985:RBG

[PK85] J. Perkkiö and R. Keskinen. The relationship between growth and allometry. Journal of Theoretical Biology, 113(1):81–87, March 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385800771.

Palsson:1984:MMDa

[PL84a] Bernhard O. Palsson and Edwin N. Lightfoot. Mathematical modelling of dynamics and control in metabolic networks. I. On Michaelis–Menten kinetics. Journal of Theoretical Biology, 111(2):273–302, November 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384802118.

Pridal:1984:MDR

[PL84b] I. Pridal and M. V. Lokajícek. A model of DSB-repair kinetics. Journal of Theoretical Biology, 111(1):81–90, November 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384801976.

Palsson:1985:MMDb

[PL85a] Bernhard O. Palsson and Edwin N. Lightfoot. Mathematical modelling of dynamics and control in metabolic networks. IV. local stability analysis of single biochemical control loops. Journal of Theoretical Biology, 113(2):261–277, March 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385802277.

Palsson:1985:MMDc

[PL85b] Bernhard O. Palsson and Edwin N. Lightfoot. Mathematical modelling of dynamics and control in metabolic networks. V. Static bifurcations in single biochemical control loops. Journal of Theoretical Biology, 113(2):279–298, March 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385802289.

Pohja:1985:SMS

[PL85c] Pasi Pohja and Juhani Lehtonen. Simulation of morphogenesis and speciation in the unicellular alga micrasterias (chlorophyta, conjugatophyceae). Journal of Theoretical Biology, 115(3):401– 414, August 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385802009.

Potten:1987:CMC

[PL87] C. S. Potten and M. Loeffler. A comprehensive model of the crypts of the small intestine of the mouse provides insight into the mechanisms of cell migration and the proliferation hierarchy. Journal of Theoretical Biology, 127(4):381–391, August 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519387801364.

Plant:1982:CMR

[Pla82] Richard E. Plant. A continuum model for root growth. Journal of Theoretical Biology, 98(1):45–59, September 7, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382900571.

Partridge:1984:CPT

[PLJ84] Derek Partridge, Patricia D. Lopez, and Victor S. Johnston. Computer programs as theories in biology. Journal of Theoretical Biology, 108(4):539–564, June 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938480079X.

Potten:1980:RIC

[PM80] Christopher S. Potten and D. Major. Repeated injection (continuous labelling) experiments in mouse epidermis. Journal of Theoretical Biology, 82(3):465–472, February 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380902490.

Pate:1981:CSC

[PO81] Edward F. Pate and Garrett M. Odell. A computer simulation of chemical signaling during the aggregation phase of Dictyostelium discoideum. Journal of Theoretical Biology, 88(2):201– 239, January 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381900710.

Pate:1986:DCS

[PO86] E. F. Pate and H. G. Othmer. Differentiation, cell sorting and proportion regulation in the slug stage of Dictyostelium discoideum. Journal of Theoretical Biology, 118(3):301–319, February 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386800613.

Pohl:1981:CSR

[Poh81] Herbert A. Pohl. Cellular spin resonance. Journal of Theoretical Biology, 93(1):207–213, November 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381900655.

Pokorny:1982:MFC

[Pok82] J. Pokorný. Multiple Fröhlich coherent states in biological systems: Computer simulation. Journal of Theoretical Biology, 98(1):21–27, September 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382900558.

Pokhariyal:1986:MGI

[Pok86] G. P. Pokhariyal. A model for the growth of infection. Journal of Theoretical Biology, 119(2):181–188, March 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386800728.

Pomiankowski:1987:CCS

[Pom87] Andrew Pomiankowski. The costs of choice in sexual selection. Journal of Theoretical Biology, 128(2):195–218, September 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519387801698.

Popel:1982:ODS

[Pop82] Aleksander S. Popel. Oxygen diffusive shunts under conditions of heterogeneous oxygen delivery. Journal of Theoretical Biology, 96(4):533–541, June 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902296.

Portelli:1981:ABM

[Por81] C. Portelli. The asymmetry of biological molecules. Journal of Theoretical Biology, 93(1):93–96, November 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381900588.

Portelli:1982:RBP

[Por82] G. Portelli. The relations between the precodons of overlapping genes. Journal of Theoretical Biology, 95(2):345–350, March 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382902491.

Portelli:1983:CDC

[Por83] C. Portelli. The l-cell and d-cell of mitosis. Journal of Theoretical Biology, 101(2):241–246, March 21, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383903387.

Powell:1985:RTC

[Pow85] Godfrey E. Powell. Relaxation times in chemostat culture. Journal of Theoretical Biology, 112(3):589–594, February 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800242.

Powell:1988:SIT

[Pow88] Godfrey E. Powell. Structural instability of the theory of simple competition. Journal of Theoretical Biology, 132(4):421– 435, June 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519388800821.

Poznanski:1987:TRS

[Poz87] Roman R. Poznanski. Transient response in a somatic shunt cable model for synaptic input activated at the terminal. Journal of Theoretical Biology, 127(1):31–50, July 7, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387801595.

Prabhakaran:1980:SAA

[PP80] M. Prabhakaran and P. K. Ponnuswamy. Spatial assignment of amino acid residues in globular proteins: An approach from information theory. Journal of Theoretical Biology, 87(4):623–637, ???? 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380901083.

Palmer:1984:DFE

[PP84a] A. R. Palmer and A. C. Pinder. The directionality of the frog ear described by a mechanical model. Journal of Theoretical Biology, 110(2):205–215, September 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384800533.

Parikh:1984:EMN

[PP84b] Jitendra C. Parikh and R. Pratap. An evolutionary model of a neural network. Journal of Theoretical Biology, 108(1):31– 38, May 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384801666.

Ponsin:1985:EAB

[PP85]  Gabriel Ponsin and Henry J. Pownall.  Equilibrium of

apoproteins between high density lipoprotein and the aqueous phase: Modelling of in vivo metabolism. Journal of Theoretical Biology, 112(1):183–192, January 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801235.

Pulliam:1982:SBJ

[PPC82] H. Ronald Pulliam, Graham H. Pyke, and Thomas Caraco. The scanning behavior of juncos: a game-theoretical approach. Journal of Theoretical Biology, 95(1):89–103, March 7, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902892.

Palsson:1985:MMDa

[PPL85] Bernhard O. Palsson, Hakon Palsson, and Edwin N. Lightfoot. Mathematical modelling of dynamics and control in metabolic networks. III. Linear reaction sequences. Journal of Theoretical Biology, 113(2):231–259, March 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385802265.

Parnas:1986:NMD

[PPS86] Hanna Parnas, Itzchak Parnas, and Lee Segel. A new method for determining co-operativity in neurotransmitter release. Journal of Theoretical Biology, 119(4):481–499, April 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801977.

Platt:1982:IRO

[PR82] E. Platt and B. Robson. Ab-initio refinement of an orbitalcentred force field for biomolecules. Test cases including peptides, a sulphonamide and modelling of DNA helices. Journal of Theoretical Biology, 96(3):381–399, June 7, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382901175.

Prabhakaran:1984:SCG

[Pra84] M. Prabhakaran. Spatial constraints and group behaviour in globular proteins. Journal of Theoretical Biology, 106(1):25–40, January 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519384900250.

Platt:1981:IRO

[PRH81] E. Platt, B. Robson, and I. H. Hillier. Ab initio refinement of an orbital-centred force field for biomolecules: Orbital localisation and parameterisation of the C–O–P(O_{2})–O–C fragment of nucleotides. Journal of Theoretical Biology, 88(2):333–353, January 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381900801.

Prothero:1980:CSC

[Pro80] John Prothero. Control of stem cell proliferation: a densitydependent commitment model. Journal of Theoretical Biology, 84(4):725–736, June 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519380800300.

Prothero:1984:SSE

[Pro84] John Prothero. Scaling of standard energy metabolism in mammals: I. Neglect of circadian rhythms. Journal of Theoretical Biology, 106(1):1–8, January 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519384900237.

Prothero:1986:MAS

[Pro86] John Prothero. Methodological aspects of scaling in biology. Journal of Theoretical Biology, 118(3):259–286, February 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386800583.

Parnas:1980:TES

[PS80a] Hanna Parnas and Lee A. Segel. A theoretical explanation for some effects of calcium on the facilitation of neurotransmitter release. Journal of Theoretical Biology, 84(1):3–29, May 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519380810356.

Peterson:1980:CER

[PS80b] Eric L. Peterson and David S. Saunders. The circadian eclosion rhythm in sarcophaga argyrostoma: a limit cycle representation of the pacemaker. Journal of Theoretical Biology, 86(2):265–277, September 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519380900077.

Parnas:1981:TSC

[PS81a] Hanna Parnas and Lee A. Segel. A theoretical study of calcium entry in nerve terminals, with application to neurotransmitter release. Journal of Theoretical Biology, 91(1):125–169, July 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381903787.

Platt:1981:EPA

[PS81b] Trevor Platt and William Silvert. Ecology, physiology, allometry and dimensionality. Journal of Theoretical Biology, 93 (4):855–860, December 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002251938190343X.

Parnas:1982:WDP

[PS82a] Hanna Parnas and Lee A. Segel. Ways to discern the presynaptic effect of drugs on neurotransmitter release. Journal of Theoretical Biology, 94(4):923–941, February 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251938290087X.

Picton:1982:MMT

[PS82b] Jill M. Picton and Martin W. Steer. A model for the mechanism of tip extension in pollen tubes. Journal of Theoretical Biology, 98(1):15–20, September 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382900546.

Parnas:1983:CSL

[PS83] Hanna Parnas and Lee A. Segel. A case study of linear versus non-linear modelling. Journal of Theoretical Biology, 103(4):549– 580, August 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383902825.

Parnas:1984:ECD

[PS84] Hanna Parnas and Lee A. Segel. Exhaustion of calcium does not terminate evoked neurotransmitter release. Journal of Theoretical Biology, 107(3):345–365, April 7, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938480096X.

Pellegrino:1983:MTP

[PSG83] J. J. Pellegrino, R. L. Sani, and R. I. Gamow. Mass transfer from phycomyces in the avoidance and anemotropic responses: Numerical modeling. Journal of Theoretical Biology, 105(1):77– 90, November 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383904253.

Pepe:1985:SEP

[PSL^{+}85] G. Pèpe, B. Serres, D. Laporte, G. Del Re, and C. Minichino. Surface electrostatic potentials on macromolecules in a monopole approximation: a computer program and an application to cytochromes. Journal of Theoretical Biology, 115(4):571–593, August 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385801417.

Powers:1981:RBC

[PTD81] Hilary J. Powers, D. I. Thurnam, and A. E. Dugdale. Is red blood cell survival limited by the activity of a critical enzyme? Journal of Theoretical Biology, 88(4):685–691, February 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902447.

Pugliese:1987:ORA

[Pug87a] Andrea Pugliese. Optimal resource allocation and optimal size in perennial herbs. Journal of Theoretical Biology, 126(1):33– 49, May 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387800991.

Pugsley:1987:ELB

[Pug87b] A. G. Pugsley. An elastic limit to branch extension. Journal of Theoretical Biology, 125(4):491–496, April 21, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387802163.

Papageorgiou:1983:RDT

[PV83] Spyros Papageorgiou and Dionyssis Venieratos. A reaction– diffusion theory of morphogenesis with inherent pattern invariance under scale variations. Journal of Theoretical Biology, 100(1):57–79, January 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519383900929.

Phillips:1988:CPL

[PW88] John B. Phillips and Jerry A. Waldvogel. Celestial polarized light patterns as a calibration reference for sun compass of homing pigeons. Journal of Theoretical Biology, 131(1):55–67, March 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388801206.

Parker:1988:AMF

[PWH^{+}88] Rodger Parker, Charles Waldren, Tom K. Hei, D. F. Wong, and T. L. Puck. Analysis of mutant frequency curves and survival curves applied to the A_{L} hybrid cell system. Journal of Theoretical Biology, 132(1):113–117, May 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388801942.

Polozov:1988:NWD

[PY88] R. V. Polozov and L. V. Yakushevich. Nonlinear waves in DNA and regulation of transcription. Journal of Theoretical Biology, 130(4):423–430, February 21, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388802078.

Papadopoulos:1985:SEE

[PYN85] Kyriakos D. Papadopoulos, Akira Yato, and Hung Nguyen. Shape and elasticity effects on erythrocyte electrostatic repulsion. Journal of Theoretical Biology, 113(3):545–557, April 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800382.

Pyshnov:1980:TSC

[Pys80] M. B. Pyshnov. Topological solution for cell proliferation in intestinal crypt. I. Elastic growth without cell loss. Journal of Theoretical Biology, 87(1):189–200, November 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380902271.

Qiwu:1982:SMS

[QL82] Cui Qiwu and G. J. Lawson. Study on models of single populations: an expansion of the logistic and exponential equations. Journal of Theoretical Biology, 98(4):645–659, October 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382901436.

Queller:1983:KSC

[Que83] David C. Queller. Kin selection and conflict in seed maturation. Journal of Theoretical Biology, 100(1):153–172, January 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383900991.

Rabinowitz:1982:EEF

[Rab82] James R. Rabinowitz. The effect of electric field induced perturbation of the distribution of ions near the cell surface on migration of charged membrane components. Journal of Theoretical Biology, 99(2):377–387, November 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251938290011X.

Rakitzis:1980:KIE

[Rak80] Emmanuel T. Rakitzis. Kinetics of irreversible enzyme inhibition: The interpretation of the fractional enzyme activity vs. extent of protein modification plot. Journal of Theoretical Biology, 85(3):553–560, August 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380903288.

Rakitzis:1986:KPM

[Rak86] Emmanuel T. Rakitzis. Kinetics of protein modification reactions: Protein conformational isomerism may imitate adsorptive complex formation. Journal of Theoretical Biology, 122(2):247–249, September 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251938680087X.

Rao:1982:CMM

[Rao82] K. Murali Krishna Rao. Capping and mitogenesis: a model implicating microfilaments in lymphocyte activation. Journal of Theoretical Biology, 98(1):61–71, September 7, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382900583.

Rappaport:1984:HRC

[Rap84] Catherine Rappaport. An hypothesis on the role of cellular colloid osmotic pressure in determining behavior of cells in vitro including anchorage dependency and maintenance of the differentiated state. Journal of Theoretical Biology, 111(4):801–816, December 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384802684.

Raudino:1983:TMS

[RAR83] A. Raudino, A. Asero, and V. Rizza. A theoretical model for surface polarizability in phospholipid membranes. Journal of Theoretical Biology, 103(1):29–41, July 7, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383901972.

Rasmussen:1987:NMD

[Ras87] Nicolas Rasmussen. A new model of developmental constraints as applied to the drosophila system. Journal of Theoretical Biology, 127(3):271–299, August 7, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387801078.

Rattay:1986:HFE

[Rat86] Frank Rattay. High frequency electrostimulation of excitable cells. Journal of Theoretical Biology, 123(1):45–54, November 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938680234X.

Rattay:1987:WAC

[Rat87] Frank Rattay. Ways to approximate current-distance relations for electrically stimulated fibers. Journal of Theoretical Biology, 125(3):339–349, April 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387800668.

Rayner:1985:BUF

[Ray85] J. M. V. Rayner. Bounding and undulating flight in birds. Journal of Theoretical Biology, 117(1):47–77, November 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801648.

Rao:1982:GCM

[RB82a] I. Narayana Rao and S. G. Bhagwat. Gene centromere mapping and detection of post-meiotic segregation in angiosperms by genotyping the polar nucleus. Journal of Theoretical Biology, 97(4):603–609, August 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382903617.

Root-Bernstein:1982:AAP

[RB82b] Robert Scott Root-Bernstein. Amino acid pairing. Journal of Theoretical Biology, 94(4):885–894, February 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382900832.

Root-Bernstein:1982:OGC

[RB82c] Robert Scott Root-Bernstein. On the origin of the genetic code. Journal of Theoretical Biology, 94(4):895–904, February 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382900844.

Root-Bernstein:1982:PP

[RB82d] Robert Scott Root-Bernstein. The problem of problems. Journal of Theoretical Biology, 99(1):193–201, November 7, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382903988.

Root-Bernstein:1982:SCO

[RB82e] Robert Scott Root-Bernstein. Selective chirility and the origins of life. Journal of Theoretical Biology, 99(1):101–109, November 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382903927.

Root-Bernstein:1983:PRA

[RB83a] Robert S. Root-Bernstein. Protein replication by amino acid pairing. Journal of Theoretical Biology, 100(1):99–106, January 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383900942.

Root-Bernstein:1983:SSL

[RB83b] Robert Scott Root-Bernstein. The structure of a serotonin and LSD binding site of myelin basic protein. Journal of Theoretical Biology, 100(3):373–378, February 7, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383904356.

Raval:1985:MWS

[RB85] M. K. Raval and U. C. Biswal. A model for the water splitting system of thylakoid membranes. Journal of Theoretical Biology, 115(1):137–151, July 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800114.

Rouhani:1987:PPS

[RB87] S. Rouhani and N. H. Barton. The probability of peak shifts in a founder population. Journal of Theoretical Biology, 126(1):

51–62, May 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387801005.

Rajkowski:1981:CEC

[RC81] K. M. Rajkowski and N. Cittanova. Corrected equations for the calculation of proteinligand binding results from fluorescence polarization data. Journal of Theoretical Biology, 93(3):691–696, December 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381902307.

Recer:1989:SEP

[RC89] Gregg M. Recer and Thomas Caraco. Sequential-encounter prey choice and effects of spatial resource variability. Journal of Theoretical Biology, 139(2):239–249, July 21, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938980102X.

Riveros:1989:LJP

[RCA89] O. J. Riveros, T. L. Croxton, and W. McD. Armstrong. Liquid junction potentials calculated from numerical solutions of the Nernst–Planck and Poisson equations. Journal of Theoretical Biology, 140(2):221–230, September 22, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389801304.

Rangarajan:1983:ODN

[RdL83] S. K. Rangarajan and R. de Levie. On one-dimensional nucleation and growth of “living” polymers II. Growth at constant monomer concentration. Journal of Theoretical Biology, 104(4):553–570, October 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251938390245X.

Reanney:1984:RSE

[Rea84] D. C. Reanney. RNA splicing as an error-screening mechanism. Journal of Theoretical Biology, 110(3):315–321, October 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384801770.

Reder:1988:MCT

[Red88] Christine Reder. Metabolic control theory: a structural approach. Journal of Theoretical Biology, 135(2):175–201, November 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388800730.

Reed:1980:NDC

[Ree80] Larry D. Reed. Neighbor distances in culture media. Journal of Theoretical Biology, 87(4):819–821, ???? 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380901204.

Regelmann:1984:RTR

[Reg84a] Klaus Regelmann. A remark on the theory of risk-sensitive foraging. Journal of Theoretical Biology, 110(2):217–222, September 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384800545.

Regen:1984:MSE

[Reg84b] David M. Regen. Myocardial stress equations: Fiberstresses of the prolate spheroid. Journal of Theoretical Biology, 109(2):191– 215, July 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251938480003X.

Regniere:1984:VTD

[Rég84c] Jacques Régnière. Vertical transmission of diseases and population dynamics of insects with discrete generations: a model. Journal of Theoretical Biology, 107(2):287–301, March 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384800296.

Regelmann:1986:LFV

[Reg86] Klaus Regelmann. Learning to forage in a variable environment. Journal of Theoretical Biology, 120(3):321–329, June 7, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386802041.

Regen:1988:IDS

[Reg88] David M. Regen. Independent determinants of systolic effectiveness: Growth ability, contractility and mobility. Journal of Theoretical Biology, 132(1):61–81, May 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388801917.

Reiner:1981:PDL

[Rei81] J. M. Reiner. Population distribution of labelled DNA and the effects of cell divisional. Journal of Theoretical Biology, 93(3):533–539, December 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902186.

Reid:1983:DIC

[Rei83a] R. E. Reid. Drug interaction with calmodulin: The binding site. Journal of Theoretical Biology, 105(1):63–76, November 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383904241.

Reiner:1983:DAT

[Rei83b] John M. Reiner. Differentiation, ageing, and terminal differentiation: A semantic analysis. Journal of Theoretical Biology, 105(4):545–552, December 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383902187.

Reiner:1984:CPG

[Rei84] Brian S. Reiner. Cyclic permutations in the genetic code, the 4 × 4 magic square and the antiparallelism of DNA. Journal of Theoretical Biology, 110(4):681–690, October 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801514.

Reid:1985:FNC

[Rei85a] R. E. Reid. The functional nature of calcium binding units in calmodulin, troponin C and parvalbumin. Journal of Theoretical Biology, 114(3):353–374, June 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801715.

Reiss:1985:ARW

[Rei85b] Michael J. Reiss. The allometry of reproduction: Why larger species invest relatively less in their offspring. Journal of Theoretical Biology, 113(3):529–544, April 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800370.

Reiss:1986:BMO

[Rei86a] Michael J. Reiss. Belovsky’s model of optimal moose size. Journal of Theoretical Biology, 122(2):237–242, September 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386800856.

Reiss:1986:BSM

[Rei86b] Michael J. Reiss. Body size and metabolic rate: Calculated exponents are independent of the units used. Journal of Theoretical Biology, 123(1):125–126, November 7, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386802405.

Reiss:1986:SDB

[Rei86c] Michael J. Reiss. Sexual dimorphism in body size: Are larger species more dimorphic? Journal of Theoretical Biology, 121(2):163–172, July 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002251938680090X.

Reiss:1987:ECC

[Rei87] Michael J. Reiss. Evolutionary conflict over the control of offspring sex ratio. Journal of Theoretical Biology, 125(1):25–39, March 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387801777.

Renshaw:1982:DSP

[Ren82] Eric Renshaw. The development of a spatial predator– prey process on interconnected sites. Journal of Theoretical Biology, 94(2):355–365, January 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382903162.

Ressler:1982:EAE

[Res82] N. Ressler. Electronic aspects of enzyme catalysis proton-electron density displacements. Journal of Theoretical Biology, 97(2):195– 225, July 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382900996.

Roman-Franco:1982:NEE

[RF82] A. A. Roman-Franco. Non-enzymatic extramicrosomal bioactivation of chemical carcinogens by phagocytes: a proposed new pathway. Journal of Theoretical Biology, 97(4):543–555, August 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382903575.

Robertson:1987:BRQ

[RG87] A. D. J. Robertson and James F. Grutsch. Biphasic responses, quantal signals and cellular behaviour. Journal of Theoretical Biology, 125(1):41–60, March 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387801789.

Rao:1982:SIC

[RGR82] J. Subba Rao, C. P. Geevan, and Gita Subba Rao. Significance of the information content of DNA in mutations and evolution. Journal of Theoretical Biology, 96(4):571–577, June 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902326.

Reid:1980:COC

[RH80] R. E. Reid and R. S. Hodges. Co-operativity and calcium/magnesium binding to troponin C and muscle calcium binding parvalbumin: An hypothesis. Journal of Theoretical Biology, 84(3):401–444, June 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519380800130.

Rowe:1985:ICF

[RH85] Glenn W. Rowe and Ian F. Harvey. Information content in finite sequences: Communication between dragonfly larvae. Journal of Theoretical Biology, 116(2):275–290, September 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938580268X.

Rowe:1985:EPE

[RHH85] Glenn W. Rowe, Ian F. Harvey, and Stephen F. Hubbard. The essential properties of evolutionary stability. Journal of Theoretical Biology, 115(2):269–285, July 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801004.

Richardson:1980:MSA

[Ric80] I. W. Richardson. The metrical structure of aging (dissipative) systems. Journal of Theoretical Biology, 85(4):745–756, August 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380902696.

Ricard:1987:DME

[Ric87] Jacques Ricard. Dynamics of multi-enzyme reactions, cell growth and perception of ionic signals from the external milieu. Journal of Theoretical Biology, 128(3):253–278, October 7, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938780070X.

Ridley:1989:WUS

[Rid89] Mark Ridley. Why not to use species in comparative tests. Journal of Theoretical Biology, 136(3):361–364, February 8, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389801717.

Rikmenspoel:1982:CCM

[Rik82] Robert Rikmenspoel. Ciliary contractile model applied to sperm flagellar motion. Journal of Theoretical Biology, 96(4):617–645, June 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251938290234X.

Rikmenspoel:1985:AEW

[Rik85] Robert Rikmenspoel. Algebraic expressions for the waveforms of sea urchin sperm flagella. Journal of Theoretical Biology, 116(1):127–147, September 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938580134X.

Riley:1980:SEE

[Ril80] John G. Riley. Strong evolutionary equilibrium and the war of attrition. Journal of Theoretical Biology, 82(3):383–400, February 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380902441.

Robins:1980:CEP

[RJ80] R. J. Robins and B. E. Juniper. Considerations of the estimation of plasmodesmatal frequencies. Journal of Theoretical Biology, 83(3):405–409, April 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380900478.

Rejmanek:1983:ISC

[RKL83] M. Rejmánek, P. Kindlmann, and J. Leps. Increase of stability with connectance in model competition communities. Journal of Theoretical Biology, 101(4):649–656, April 21, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383900206.

Robinson:1982:FCE

[RL82] James V. Robinson and G. S. Ladde. Feasibility constraints on the elastic expansion of ecosystem models. Journal of Theoretical Biology, 97(2):277–287, July 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382901047.

Richardson:1983:PRP

[RL83] I. W. Richardson and A. H. Louie. Projections as representations of phenomena. Journal of Theoretical Biology, 102(2):199– 223, May 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383903600.

Rubin:1984:DSS

[RL84] Clinton T. Rubin and Lance E. Lanyon. Dynamic strain similarity in vertebrates; an alternative to allometric limb bone scaling. Journal of Theoretical Biology, 107(2):321–327, March 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384800314.

Rajasekar:1988:PDR

[RL88] S. Rajasekar and M. Lakshmanan. Period doubling route to chaos for a BVP oscillator with periodic external force. Journal of Theoretical Biology, 133(4):473–477, August 21, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388803357.

Rldley:1982:CSC

[Rld82] J. N. Rldley. Computer simulation of contact pressure in capitula. Journal of Theoretical Biology, 95(1):1–11, March 7, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902831.

Richardson:1982:PCC

[RLS82] I. W. Richardson, A. H. Louie, and S. Swaminathan. A phenomenological calculus for complex systems. Journal of Theoretical Biology, 94(1):61–76, January 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382903290.

Regen:1983:DCD

[RM83] David M. Regen and Calvin R. Maurer. The dependence of chamber dynamics on chamber dimensions. Journal of Theoretical Biology, 105(4):679–705, December 21, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383902278.

Regen:1986:EMP

[RM86] David M. Regen and Calvin R. Maurer. Evaluation of myocardial properties from image/pressure data: Chronic conditions. Journal of Theoretical Biology, 120(1):31–61, May 7, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386800169.

Ripoll:1981:NAM

[RMB81] Daniel Ripoll, Ezio Marchi, and Julio Benegas. A new approach to microtubule depolymerization. Journal of Theoretical Biology, 93(1):53–62, November 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381900564.

Rayfield:1985:CDS

[RMFM85] Mark Rayfield, George S. Michaels, Richard Feldmann, and Nicholas Muzyczka. Comparison of the DNA sequence and secondary structure of the herpes simplex virus L/S junction and the adeno-associated virus terminal repeat. Journal of Theoretical Biology, 115(4):477–494, August 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801363.

Rapp:1981:FEB

[RMS81] P. E. Rapp, A. I. Mees, and C. T. Sparrow. Frequency encoded biochemical regulation is more accurate than amplitude dependent control. Journal of Theoretical Biology, 90(4):531–544, June 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381903040.

Ricard:1982:ERE

[RN82] Jacques Ricard and Georges Noat. Evolution of regulatory enzymes towards functional simplicity. Journal of Theoretical Biology, 96(3):347–365, June 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382901151.

Ricard:1984:ERSa

[RN84a] Jacques Ricard and Georges Noat. Enzyme reactions at the surface of living cells: I. Electric repulsion of charged ligands and recognition of signals from the external milieu. Journal of Theoretical Biology, 109(4):555–569, August 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801587.

Ricard:1984:ERSb

[RN84b] Jacques Ricard and Georges Noat. Enzyme reactions at the surface of living cells: II. Destabilization in the membranes and conduction of signals. Journal of Theoretical Biology, 109(4):571– 580, August 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384801599.

Ricard:1984:SIE

[RN84c] Jacques Ricard and Georges Noat. Subunit interactions in enzyme transition states-antagonism between substrate binding and reaction rate. Journal of Theoretical Biology, 111(4):737–753, December 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384802659.

Ricard:1985:SCK

[RN85] Jacques Ricard and Georges Noat. Subunit coupling and kinetic co-operativity of polymeric enzymes. Amplification, attenuation and inversion effects. Journal of Theoretical Biology, 117(4):633–649, December 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519385802447.

Ricard:1986:CEK

[RN86] Jacques Ricard and Georges Noat. Catalytic efficiency, kinetic co-operativity of oligomeric enzymes and evolution. Journal of Theoretical Biology, 123(4):431–451, December 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386802120.

Robinson:1980:NMT

[Rob80] James V. Robinson. A necessary modification of trivers’ parent-offspring conflict model. Journal of Theoretical Biology, 83(3):533–535, April 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519380900612.

Robinson:1981:MGS

[Rob81] Charles E. Robinson. The molecule-group schema of memory storage. Journal of Theoretical Biology, 91(1):231–232, July 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381903829.

Roberts:1984:CCP

[Rob84] D. W. Roberts. A chemical contact pressure model for phyllotaxis. Journal of Theoretical Biology, 108(4):481–490, June 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384800740.

Robertson:1985:PTI

[Rob85a] Donald S. Robertson. A possible technique for isolating genic DNA for quantitative traits in plants. Journal of Theoretical Biology, 117(1):1–10, November 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801612.

Robinson:1985:MGS

[Rob85b] Charles E. Robinson. The molecule-group schema of memory: Storage, recognition and retrieval. Journal of Theoretical Biology, 112(4):859–871, February 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800679.

Roberts:1987:CCP

[Rob87] D. W. Roberts. The chemical contact pressure model for phyllotaxis — application to phyllotaxis changes in seedlings and to anomalous phyllotaxis systems. Journal of Theoretical Biology, 125(2):141–161, March 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387800371.

Robinson:1988:ORB

[Rob88] David Robinson. Optimal relations between root length and nutrient inflow rate in plant root systems. Journal of Theoretical Biology, 135(3):359–370, December 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388802509.

Robinson:1989:CND

[Rob89] David Robinson. Can the nutrient demand of a plant be sustained by an increase in local inflow rate? Journal of Theoretical Biology, 138(4):551–554, June 22, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389800517.

Rogatko:1983:DMC

[Rog83] André Rogatko. Diallelic model combining selection and incomplete penetrance. Journal of Theoretical Biology, 102(1):1–6, May 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251938390259X.

Rogers:1984:UDB

[Rog84] Edward F. Rogers. The use and design of biligand enzyme inhibitors. Journal of Theoretical Biology, 109(3):471–474, August 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384800958.

Rogers:1986:PDU

[Rog86] Alan R. Rogers. Population dynamics under exploitation competition. Journal of Theoretical Biology, 119(3):363–368, April 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386801473.

Rogers:1989:RPS

[Rog89] Alan R. Rogers. Resource partitioning and the stability of population dynamics: a reply to lomnicki and sedziwy. Journal of Theoretical Biology, 138(4):545–549, June 22, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389800505.

Rosin:1980:HBD

[Ros80a] R. Rosin. The honey-bee “dance language” hypothesis and the foundations of biology and behavior. Journal of Theoretical Biology, 87(3):457–481, December 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380902313.

Rosin:1980:PHB

[Ros80b] R. Rosin. Paradoxes of the honey-bee “dance language” hypothesis. Journal of Theoretical Biology, 84(4):775–800, June 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519380800336.

Rose:1982:PBM

[Ros82] Michael R. Rose. A physiological barrier for the maintenance of anisogamous sex. Journal of Theoretical Biology, 94(4):801–813, February 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382900790.

Rose:1983:CMO

[Ros83] Michael R. Rose. The contagion mechanism for the origin of sex. Journal of Theoretical Biology, 101(1):137–146, March 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383902771.

Rosen:1984:OPE

[Ros84a] Gerald Rosen. Orthogonal projection experiments for manyparameter dependent phenomena: Optimization of hydroponic plant growth. Journal of Theoretical Biology, 108(2):315–318, May 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384800739.

Rosin:1984:FAH

[Ros84b] R. Rosin. Further analysis of the honey bee “dance language” controversy. I. Presumed proofs for the “dance language” hypothesis, by Soviet scientists. Journal of Theoretical Biology, 107(3):417– 442, April 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384801009.

Rossini:1984:SRR

[Ros84c] Gian Paolo Rossini. Steroid receptor recycling and its possible role in the modulation of steroid hormone action. Journal of Theoretical Biology, 108(1):39–53, May 7, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801678.

Rosen:1985:SHS

[Ros85] Gerald Rosen. Stable heterogeneous spatial distributions of oxygen and chemotactic E. coli with Verhulst-limited population growth. Journal of Theoretical Biology, 112(4):877–881, February 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385800692.

Rothstein:1980:RAK

[Rot80] Stephen I. Rothstein. Reciprocal altruism and kin selection are not clearly separable phenomena. Journal of Theoretical Biology, 87(2):255–261, November 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380903598.

Rotenberg:1982:DLG

[Rot82a] Manuel Rotenberg. Diffusive logistic growth in deterministic and stochastic environments. Journal of Theoretical Biology, 94(2):253–280, January 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382903125.

Rotenberg:1982:TDQ

[Rot82b] Manuel Rotenberg. Theory of distributed quiescent state in the cell cycle. Journal of Theoretical Biology, 96(3):495–509, June 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382901230.

Rotenberg:1983:TTG

[Rot83] Manuel Rotenberg. Transport theory for growing cell populations. Journal of Theoretical Biology, 103(2):181–199, July 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383900243.

Rotenberg:1987:ECS

[Rot87] Manuel Rotenberg. Effect of certain stochastic parameters on extinction and harvested populations. Journal of Theoretical Biology, 124(4):455–471, February 21, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387802205.

Routledge:1980:FSA

[Rou80] R. D. Routledge. The form of species-abundance distributions. Journal of Theoretical Biology, 82(4):547–558, February 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380901770.

Rowe:1985:MMS

[Row85a] Glenn W. Rowe. Mutations, mixed strategies and game theory. Journal of Theoretical Biology, 117(2):291–302, November 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938580223X.

Rowe:1985:TDR

[Row85b] Glenn W. Rowe. A three-dimensional representation for base composition of protein-coding DNA sequences. Journal of Theoretical Biology, 112(2):433–444, January 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385802964.

Rowe:1987:DGT

[Row87] Glenn W. Rowe. A dynamic game theory model of a diploid genetic system. Journal of Theoretical Biology, 129(2):243–255, November 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387800164.

Rowe:1988:EGS

[Row88] Glenn W. Rowe. To each genotype a separate strategy — a dynamic game theory model of a general diploid system. Journal of Theoretical Biology, 134(1):89–101, September 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388803047.

Regen:1984:SPR

[RP84] David M. Regen and Simon J. Pilkis. Sensitivity of pathway rate to activities of substrate-cycle enzymes: Application to gluconeogenesis and glycolysis. Journal of Theoretical Biology, 111(4):635–658, December 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384802593.

Robinson:1981:TCS

[RQ81] D. G. Robinson and H. Quader. Towards cellulose synthesis in vitro from plant extracts. Journal of Theoretical Biology, 92(4):483–495, October 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902617.

Rodin:1983:STAa

[RR83a] Sergey N. Rodin and Vadim A. Ratner. Some theoretical aspects of protein coevolution in the ecosystem “phage-bacteria” I. The problem. Journal of Theoretical Biology, 100(2):185–195, January 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383903466.

Rodin:1983:STAb

[RR83b] Sergey N. Rodin and Vadim A. Ratner. Some theoretical aspects of protein coevolution in the ecosystem “phage-bacteria” II. The deterministic model of microevolution. Journal of Theoretical Biology, 100(2):197–210, January 21, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383903478.

Ratner:1984:CDC

[RR84] Vadim A. Ratner and Sergey N. Rodin. Coevolution of directly contacting proteins in phage-bacterium ecosystem: Possibility or fiction? Journal of Theoretical Biology, 108(3):405–411, June 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384800429.

Regan:1988:FDT

[RR88a] M. P. Regan and D. Regan. A frequency domain technique for characterizing nonlinearities in biological systems. Journal of Theoretical Biology, 133(3):293–317, August 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388803230.

Ruppersberg:1988:ESC

[RR88b] J. Peter Ruppersberg and Reinhardt Rüdel. The effects of space charge on the ionic currents through biological membranes. Journal of Theoretical Biology, 130(4):431–445, February 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938880208X.

Ricard:1982:SOD

[RS82] Jacques Ricard and Jean-Michel Soulie. Self-organization and dynamics of an open futile cycle. Journal of Theoretical Biology, 95(1):105–121, March 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902909.

Rao:1983:TAR

[RS83] A. S. Rao and Randhir Singh. Theoretical approaches for reducing eioenergetic costs and enhancing plant productivity. Journal of Theoretical Biology, 104(1):113–120, September 7, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383904046.

Reed:1984:ESS

[RS84] Jon Reed and Nils C. Stenseth. On evolutionarily stable strategies. Journal of Theoretical Biology, 108(4):491–508, June 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384800752.

Ruggiero:1986:SMT

[RS86] C. Ruggiero and B. McA. Sayers. A structural model of the thoracic cage in the cat. Journal of Theoretical Biology, 119(2):161– 180, March 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519386800716.

Rao:1987:SSP

[RS87] A. S. Rao and Randhir Singh. Some speculations on prion protein replication and morphology. Journal of Theoretical Biology, 127(2):251–252, July 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387801352.

Rubinow:1981:MFS

[RSE81] S. I. Rubinow, Lee A. Segel, and Witta Ebel. A mathematical framework for the study of morphogenetic development in the slime mold. Journal of Theoretical Biology, 91(1):99–113, July 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381903763.

Ruffner:1980:AHH

[RSM80] James A. Ruffner, Nick Sperelakis, and James E. Mann. Application of the Hodgkin–Huxley equations to an electric field model for interaction between excitable cells. Journal of Theoretical Biology, 87(1):129–152, November 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380902246.

Ralph:1988:SNC

[RSR88] R. K. Ralph, E. Schneider, and S. Rattray. A speculation on the nature of chromosomal DNA. Journal of Theoretical Biology, 135(1):127–129, November 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388801814.

Ruggiero:1987:STC

[RSSK87] C. Ruggiero, B. McA. Sayers, T. A. Sears, and P. A. Kirkwood. Spatio-temporal characterization of intercostal activity during breathing in the cat. Journal of Theoretical Biology, 125(2):125– 140, March 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251938780036X.

Rowe:1983:ICV

[RT83a] Glenn W. Rowe and L. E. H. Trainor. On the informational content of viral DNA. Journal of Theoretical Biology, 101(2):151– 170, March 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383903326.

Rowe:1983:TTC

[RT83b] Glenn W. Rowe and L. E. H. Trainor. A thermodynamic theory of codon bias in viral genes. Journal of Theoretical Biology, 101(2):171–203, March 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383903338.

Raffa:1985:CCR

[RT85] Robert B. Raffa and Ronald J. Tallarida. The concept of a changing receptor concentration: Implications for the theory of drug action. Journal of Theoretical Biology, 115(4):625–632, August 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385801454.

Rabinovitch:1988:BPO

[RT88] A. Rabinovitch and R. Thieberger. Biological populations obeying a stochastically perturbed logistic difference equation. Journal of Theoretical Biology, 131(4):509–514, April 21, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388800456.

Rao:1981:CMEa

[RTM81a] Gita Subba Rao, R. S. Tyagi, and R. K. Mishra. Calculation of the minimum energy conformation of biomolecules using a global optimization technique I. Methodology and application to a model molecular fragment (normal pentane). Journal of Theoretical Biology, 90(3):377–389, June 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381903180.

Rao:1981:CMEb

[RTM81b] Gita Subba Rao, R. S. Tyagi, and R. K. Mishra. Calculation of the minimum energy conformation of biomolecules using a global optimization technique II. Conformation of Nacetylglycine N-methyl amide, a model dipeptide unit. Journal of Theoretical Biology, 90(3):391–395, June 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381903192.

Rao:1982:CME

[RTM82] Gita Subba Rao, R. S. Tyagi, and R. K. Mishra. Calculation of the minimum energy conformation of biomolecules using a global optimization technique. III. Conformation of acetylcholine using molecular fragments. Journal of Theoretical Biology, 98(3):543– 548, October 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382901369.

Rebeiz:1988:CBK

[RTM88] Constantin A. Rebeiz, Baishnab C. Tripathy, and Joseph M. Mayasich. Chloroplast biogenesis 61: Kinetic analysis of precursor-product relationships in complex biosynthetic pathways. Journal of Theoretical Biology, 133(3):319–326, August 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388803242.

Rutberg:1983:EMP

[Rut83] Allen T. Rutberg. The evolution of monogamy in primates. Journal of Theoretical Biology, 104(1):93–112, September 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383904034.

Roberts:1982:DQC

[RW82] D. W. Roberts and D. L. Williams. The derivation of quantitative correlations between skin sensitisation and physiochemical parameters for alkylating agents, and their application to experimental data for sultones. Journal of Theoretical Biology, 99(4):807–825, December 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382901990.

Rzeszowska-Wolny:1983:GCE

[RW83] J. Rzeszowska-Wolny. Is genetic code error-correcting? Journal of Theoretical Biology, 104(4):701–702, October 21, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383902576.

Ryan:1980:IEI

[Rya80] J. Philip Ryan. Information-entropy interfaces and different levels of biological organization. Journal of Theoretical Biology, 84(1): 31–48, May 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519380810368.

Ryti:1987:AER

[Ryt87] Randall T. Ryti. Allocation of energy to reproduction: Effects of competition and population structure. Journal of Theoretical Biology, 128(4):499–512, October 21, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387801935.

Rizzotti:1986:AGB

[RZ86] Martino Rizzotti and Alberto Zanardo. Axiomatization of genetics 1. Biological meaning. Journal of Theoretical Biology, 118(1):61–71, January 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002251938680008X.

Simon:1980:SCS

[SA80] I. Simon and B. Asbóth. Subunit contact surface — an additional argument in favour of continuous folding during biosynthesis of proteins. Journal of Theoretical Biology, 82(4):685–688, February 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380901873.

Saxena:1981:SSH

[SA81] V. P. Saxena and D. Arya. Steady-state heat distribution in epidermis, dermis and subdermal tissues. Journal of Theoretical Biology, 89(3):423–432, April 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251938190360X.

Severns:1982:RBK

[SA82] M. L. Severns and J. Milton Adams. The relation between krogh and compartmental transport models. Journal of Theoretical Biology, 97(2):239–249, July 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382901011.

Schultz:1986:RBD

[SA86] T. W. Schultz and H. Y. Ando. A relationship between diffusional transport in lipid membranes and a lipophilic-eutectic parameter. Journal of Theoretical Biology, 123(3):367–375, December 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938680251X.

Sabesan:1980:SSA

[Sab80] M. N. Sabesan. Secondary structural and “active site” momologies between nerve growth factor and insulin. Journal of Theoretical Biology, 83(3):469–476, April 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380900533.

Sachs:1988:ESA

[Sac88] T. Sachs. Epigenetic selection: An alternative mechanism of pattern formation. Journal of Theoretical Biology, 134(4):547–559, October 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519388800560.

Saito:1982:ARP

[Sai82] Nobuhiko Saitô. Asymptotic regular pattern of epidermal cells in mammalian skin. Journal of Theoretical Biology, 95(3):591– 599, April 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382900364.

Salu:1984:TMC

[Sal84] Yehuda Salu. Theoretical models and computer simulations of neural learning systems. Journal of Theoretical Biology, 111(1):31–46, November 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801940.

Sanders:1984:STR

[San84] Dennis A. Sanders. Self-transcription: RNA polymerase transcription of its own genes, and its role in cellular differentiation and cell cycling. Journal of Theoretical Biology, 106(2):171–182, January 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519384900171.

Sanford:1988:APP

[San88] John C. Sanford. Applying the PDR principle to AIDS. Journal of Theoretical Biology, 130(4):469–480, February 21, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938880211X.

Sarai:1982:EET

[Sar82] Akinori Sarai. Efficiency of electron transport processes. Journal of Theoretical Biology, 99(2):341–355, November 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382900091.

Saroff:1985:UPS

[Sar85] Harry A. Saroff. The uniqueness of protein sequences: Molecular territoriality. Journal of Theoretical Biology, 115(2):191– 199, July 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385800953.

Saroff:1987:ICI

[Sar87] Harry A. Saroff. The ionization of clusters II. Pairwise isotropic interactions and individual site binding data. Journal of Theoretical Biology, 129(4):427–438, December 21, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938780022X.

Sarai:1989:MRI

[Sar89] Akinori Sarai. Molecular recognition and information gain. Journal of Theoretical Biology, 140(1):137–143, September 11, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519389800347.

Schmidt:1985:PTC

[SAS85] Steven K. Schmidt, Martin Alexander, and Michael L. Shuler. Predicting threshold concentrations of organic substrates for bacterial growth. Journal of Theoretical Biology, 114(1):1–8, May 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385802502.

Saunders:1980:OAC

[Sau80] Kenneth B. Saunders. Oscillations of arterial CO 2 tension in a respiratory model: Some implications for the control of breathing in exercise. Journal of Theoretical Biology, 84(1):163–179, May 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519380810423.

Savageau:1981:APZ

[Sav81] Michael A. Savageau. Accuracy of proofreading with zero energy cost. Journal of Theoretical Biology, 93(1):179–195, November 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381900631.

Saxena:1983:TDH

[Sax83] V. P. Saxena. Temperature distribution in human skin and subdermal tissues. Journal of Theoretical Biology, 102(2):277– 286, May 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251938390365X.

Szocs:1986:DET

[SB86] Vojtech Szöcs and Ivan Barvík. Description of excitation transfer and damping in the primary processes of bio-(chemical) systems. Journal of Theoretical Biology, 122(2):179–186, September 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386800807.

Saunders:1980:BMR

[SBC80] Kenneth B. Saunders, Hari N. Bali, and Ewert R. Carson. A breathing model of the respiratory system: The controlled system. Journal of Theoretical Biology, 84(1):135–161, May 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519380810411.

Sibbald:1989:CHO

[SBM89] Peter R. Sibbald, Satindranath Banerjee, and Jack Maze. Calculating higher order DNA sequence information measures. Journal of Theoretical Biology, 136(4):475–483, February 22, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389801596.

Sorarrain:1983:AAM

[SBPF83] Oscar M. Sorarrain, Rafael R. Boggio, Jorge L. Pousa, and Ewald A. Favret. Application of absorbent Markov chains for a selling model of two independent logi. Journal of Theoretical Biology, 103(2):173–180, July 21, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383900231.

Soloway:1983:CRS

[SBW83] A. H. Soloway, R. J. Brumbaugh, and D. T. Witiak. Carbinolamines and related structures — potential alkylating metabolites of clinically active anticancer drugs. Journal of Theoretical Biology, 102(3):361–373, June 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383903740.

Sibly:1982:ARP

[SC82] Richard Sibly and Peter Calow. Asexual reproduction in protozoa and invertebrates. Journal of Theoretical Biology, 96(3):401– 424, June 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382901187.

Sibly:1983:IAL

[SC83] Richard Sibly and Peter Calow. An integrated approach to life-cycle evolution using selective landscapes. Journal of Theoretical Biology, 102(4):527–547, June 21, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383903892.

Schwabe:1984:PSC

[SC84a] W. W. Schwabe and A. G. Clewer. Phyllotaxis — a simple computer model based on the theory of a polarly-translocated inhibitor. Journal of Theoretical Biology, 109(4):595–619, August 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384801617.

Sibly:1984:DAC

[SC84b] Richard Sibly and Peter Calow. Direct and absorption costing in the evolution of life cycles. Journal of Theoretical Biology, 111(3):463–473, December 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384802349.

Sibly:1986:WBE

[SC86a] R. Sibly and P. Calow. Why breeding earlier is always worthwhile. Journal of Theoretical Biology, 123(3):311–319, December 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386802466.

Soucek:1986:BWF

[SC86b] Branko Soucek and Albert D. Carlson. Brain windows in firefly communication. Journal of Theoretical Biology, 119(1):47– 65, March 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519386800509.

Sibly:1987:ECC

[SC87a] Richard Sibly and Peter Calow. Ecological compensation — a complication for testing life-history theory. Journal of Theoretical Biology, 125(2):177–186, March 21, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387800395.

Soucek:1987:BWL

[SC87b] Branko Soucek and Albert D. Carlson. Brain window language of fireflies. Journal of Theoretical Biology, 125(1):93–103, March 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519387801820.

Srikanth:1989:LRA

[SC89] Nadadur S. Srikanth and V. Chandrasekharam. Linear regression analysis for enzyme kinetic studies in toxicant interactions. Journal of Theoretical Biology, 138(4):467–478, June 22, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389800451.

Singh:1981:ERC

[SCB81] S. N. Singh, K. C. Chakrabarty, and D. Barman. On estimating the risk of conception from observations on first conceptive delays. Journal of Theoretical Biology, 90(1):1–7, May 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381901193.

Schoen:1980:HTA

[Sch80] Daniel J. Schoen. Half tetrad analysis in angiosperms. Journal of Theoretical Biology, 86(2):315–322, September 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380900090.

Schroeder:1981:CGE

[Sch81] Lauren Alfred Schroeder. Consumer growth efficiencies: Their limits and relationships to ecological energetics. Journal of Theoretical Biology, 93(4):805–828, December 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381903416.

Schmitz:1982:MAI

[Sch82a] Kenneth S. Schmitz. A model for the association of intercalating ligands with mononucleosomes and chromatin. Journal of Theoretical Biology, 98(1):29–43, September 7, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251938290056X.

Schole:1982:TMR

[Sch82b] J. Schole. Theory of metabolic regulation including hormonal effects on the molecular level. Journal of Theoretical Biology, 96(4):579–615, June 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902338.

Scherer:1983:BFS

[Sch83a] Siegfried Scherer. Basic functional states in the evolution of light-driven cyclic electron transport. Journal of Theoretical Biology, 104(2):289–299, September 21, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383904162.

Schmidt:1983:PCD

[Sch83b] Willi Schmidt. The problem of cell debris in pulse-cytophotometry: a probability theoretic model. Journal of Theoretical Biology, 103(4):581–599, August 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519383902837.

Schlitter:1984:REE

[Sch84a] Jürgen Schlitter. Rapid energy exchange of enzymes with solvent water by dipolar interaction. Journal of Theoretical Biology, 106(3):303–313, February 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519384900328.

Shoaf:1984:ARD

[SCH84b] Sarah A. Shoaf, Kevin Conway, and R. Kevin Hunt. Application of reaction–diffusion models to cell patterning in xenopus retina. Initiation of patterns and their biological stability. Journal of Theoretical Biology, 109(3):299–329, August 7, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384800855.

Schwemmler:1985:PTM

[Sch85] Werner Schwemmler. Possible transitions from molecules to cells. Journal of Theoretical Biology, 117(2):187–208, November 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385802174.

Schaffer:1988:ESS

[Sch88] Mark E. Schaffer. Evolutionarily stable strategies for a finite population and a variable contest size. Journal of Theoretical Biology, 132(4):469–478, June 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388800857.

Schierwagen:1989:NUE

[Sch89] Andreas K. Schierwagen. A non-uniform equivalent cable model of membrane voltage changes in a passive dendritic tree. Journal of Theoretical Biology, 141(2):159–179, November 21, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389800153.

Sibly:1985:PGA

[SCN85] R. Sibly, P. Calow, and N. Nichols. Are patterns of growth adaptive? Journal of Theoretical Biology, 112(3):553–574, February 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385800229.

Scott:1982:BSA

[Sco82] James A. Scott. On the biochemical similarities of ascorbic acid and interferon. Journal of Theoretical Biology, 98(2):235–238, September 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382902612.

Scott:1984:PAP

[Sco84a] James A. Scott. Phospholipase activity and plasma membrane homeostasis. Journal of Theoretical Biology, 111(4):659–665, December 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251938480260X.

Scott:1984:RCI

[Sco84b] James A. Scott. The role of cytoskeletal integrity in cellular transformation. Journal of Theoretical Biology, 106(2):183–188, January 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519384900183.

Scott:1989:CME

[Sco89] Susan M. Scott. Comments on a model of egg development and diapause. Journal of Theoretical Biology, 136(1):117–120, January 9, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519389801936.

Savage:1983:DBH

[SCP83] John R. K. Savage, A. H. Cawood, and D. G. Papworth. The disparity between homologous chromosomes during DNA replication. Journal of Theoretical Biology, 100(4):631–643, February 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383903272.

Sulsky:1984:MCS

[SCP84] D. Sulsky, S. Childress, and J. K. Percus. A model of cell sorting. Journal of Theoretical Biology, 106(3):275–301, February 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519384900316.

Sibly:1988:OSS

[SCS88] R. Sibly, P. Calow, and R. H. Smith. Optimal size of seasonal breeders. Journal of Theoretical Biology, 133(1):13–21, July 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388800213.

Soberon:1981:DPP

[SD81] Jorge M. Soberon and Carlos Martinez Del Rio. The dynamics of a plant-pollinator interaction. Journal of Theoretical Biology, 91(2):363–378, July 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902381.

Stewart:1983:SRP

[SD83] D. W. Stewart and L. M. Dwyer. Stomatal response to plant water deficits. Journal of Theoretical Biology, 104(4):655–666, October 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383902539.

Szathmary:1987:GSE

[SD87] Eörs Szathmáry and László Demeter. Group selection of early replicators and the origin of life. Journal of Theoretical Biology, 128(4):463–486, October 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387801911.

Skalak:1982:ADG

[SDM^{+}82] R. Skalak, G. Dasgupta, M. Moss, E. Otten, and … H. Vilmann. Analytical description of growth. Journal of Theoretical Biology, 94(3):555–577, February 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382903010.

Silk:1980:LBR

[SE80] Wendy Kuhn Silk and Ralph O. Erickson. Local biosynthesis rates of cytoplasmic constituents in growing tissue. Journal of Theoretical Biology, 83(4):701–703, April 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380901976.

Searcy:1980:OBS

[Sea80] William A. Searcy. Optimum body sizes at different ambient temperatures: an energetics explanation of Bergmann’s rule. Journal of Theoretical Biology, 83(4):579–593, April 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380901903.

Seeman:1982:NAJ

[See82] Nadrian C. Seeman. Nucleic acid junctions and lattices. Journal of Theoretical Biology, 99(2):237–247, November 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382900029.

Seely:1986:EMF

[See86] Stephen Seely. An electrodynamic (moving field) theory of muscular contraction. Journal of Theoretical Biology, 121(2):233– 248, July 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519386800960.

Seger:1981:KC

[Seg81] Jon Seger. Kinship and covariance. Journal of Theoretical Biology, 91(1):191–213, July 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381903805.

Sekeris:1985:RHC

[Sek85] Constantine E. Sekeris. The role of HnRNA in the control of ribosomal gene transcription. Journal of Theoretical Biology, 114(4):601–604, June 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800461.

Selten:1980:NES

[Sel80] Reinhard Selten. A note on evolutionarily stable strategies in asymmetric animal conflicts. Journal of Theoretical Biology, 84(1):93–101, May 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519380810381.

Stenberg:1982:KRZ

[SEN82] Manne Stenberg, Hans Elwing, and Håkan Nygren. Kinetics of reaction zone formation with radial diffusion of ligands over a receptor-coated surface. Journal of Theoretical Biology, 98(2):307–320, September 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902673.

Sen:1988:TCR

[Sen88] Asok K. Sen. On the time course of the reversible Michaelis– Menten reaction. Journal of Theoretical Biology, 135(4):483–493, December 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519388802716.

Serwer:1987:MPT

[Ser87] Philip Serwer. The mechanism for producing two symmetries at the head-tail junction of bacteriophages: a hypothesis. Journal of Theoretical Biology, 127(2):155–161, July 21, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387801261.

Seshi:1982:DRS

[Ses82] B. Seshi. The dual recognition systems of T lymphocytes: a model. Journal of Theoretical Biology, 99(4):827–830, December 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382902004.

Seshadri:1981:TES

[SF81] M. S. Seshadri and G. Fritzsch. The time evolution of sequential enzyme reactions: a singular perturbation approach. Journal of Theoretical Biology, 93(1):197–205, November 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381900643.

Stephan:1983:RTMa

[SF83] W. Stephan and E. Frehland. Rate theory models for ion transport through rigid pores. II. Time-dependent analysis of the single-file model with special reference to oscillatory behaviour. Journal of Theoretical Biology, 103(4):481–505, August 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383902795.

Small:1989:MMM

[SF89] J. Rankin Small and David A. Fell. The matrix method of metabolic control analysis: its validity for complex pathway structures. Journal of Theoretical Biology, 136(2):181–197, January 23, 1989. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519389802255.

Sankoff:1986:ALC

[SFR86] David Sankoff, Vincent Ferretti, and Dominique Roby. Allele and locus classification in electrophoretic population studies. Journal of Theoretical Biology, 119(4):397–407, April 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801916.

Stein:1982:EBR

[SG82] Murray B. Stein and Richard Gordon. Epithelia as bubble rafts: a new method for analysis of cell shape and intercellular adhesion in embryonic and other epithelia. Journal of Theoretical Biology, 97(4):625–639, August 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382903630.

Sibly:1986:SRC

[SG86] R. M. Sibly and J. P. Grime. Strategies of resource capture by plants-evidence for adversity selection. Journal of Theoretical Biology, 118(2):247–250, January 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801370.

Shine:1988:EVR

[SG88] Richard Shine and Louis J. Guillette. The evolution of viviparity in reptiles: a physiological model and its ecological consequences. Journal of Theoretical Biology, 132(1):43–50, May 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388801899.

Smeller:1989:SMM

[SG89a] L. Smeller and S. Györgyi. A statistical mechanical model of the pre- and subtransitions of lecithin membranes. Journal of Theoretical Biology, 137(2):203–214, March 21, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389802061.

Szathmary:1989:SEG

[SG89b] Eörs Szathmáry and Irina Gladkih. Sub-exponential growth and coexistence of non-enzymatically replicating templates. Journal of Theoretical Biology, 138(1):55–58, May 7, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389801778.

Staudte:1984:AMD

[SGd84] R. G. Staudte, M. Guiguet, and M. Collyn d’Hooghe. Additive models for dependent cell populations. Journal of Theoretical Biology, 109(1):127–146, July 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801150.

Segel:1986:MES

[SGDK86] Lee A. Segel, Albert Goldbeter, Peter N. Devreotes, and Barry E. Knox. A mechanism for exact sensory adaptation based on receptor modification. Journal of Theoretical Biology, 120(2):151– 179, May 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519386801710.

Sernetz:1985:OBI

[SGH85] M. Sernetz, B. Gelléri, and J. Hofmann. The organism as bioreactor. Interpretation of the reduction law of metabolism in terms of heterogeneous catalysis and fractal structure. Journal of Theoretical Biology, 117(2):209–230, November 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385802186.

Smith:1986:VBS

[SGJ86] Richard J. Smith, Rebecca Z. German, and William L. Jungers. Variability of biological similarity criteria. Journal of Theoretical Biology, 118(3):287–293, February 7, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386800595.

Sabesan:1980:ARE

[SH80a] M. N. Sabesan and Edwin T. Harper. Are aromatic residues essential at the “active sites” of peptide hormones? Journal of Theoretical Biology, 83(3):457–467, April 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380900521.

Sharpe:1980:RKN

[SH80b] Peter J. H. Sharpe and Lorinda C. Hu. Reaction kinetics of nutrition dependent poikilotherm developmentt. Journal of Theoretical Biology, 82(2):317–333, January 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251938090106X.

Saunders:1981:ICE

[SH81] P. T. Saunders and M. W. Ho. On the increase in complexity in evolution II. The relativity of complexity and the principle of minimum increase. Journal of Theoretical Biology, 90(4):515– 530, June 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381903039.

Schom:1982:KSS

[SH82a] Charles B. Schom and Efraim Halfon. Kin selection as a salmonid survival strategy explaining mortality and variable growth in families: a modeling study. Journal of Theoretical Biology, 94(1):153– 172, January 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251938290337X.

Sy:1982:ASM

[SH82b] W. N-C. Sy and P. Han. Analysis of a stochastic model of cell survival. Journal of Theoretical Biology, 96(2):309–326, May 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382902272.

Shimizu:1983:COD

[SH83a] H. Shimizu and H. Haken. Co-operative dynamics in organelles. Journal of Theoretical Biology, 104(2):261–273, September 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383904149.

Sluyser:1983:CPH

[SH83b] Mels Sluyser and Guus Hart. Calculation of percentage of hormone independent transplantable cells in experimental mammary tumors. Journal of Theoretical Biology, 100(4):701–707, February 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383903314.

Saunders:1985:PSW

[SH85a] P. T. Saunders and M. W. Ho. Primary and secondary waves in prepattern formation. Journal of Theoretical Biology, 114(3):491– 504, June 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385801806.

Starmer:1985:MAA

[SH85b] C. Frank Starmer and Michael D. Hollett. Mechanisms of apparent affinity variation of guarded receptors. Journal of Theoretical Biology, 115(3):337–349, August 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938580196X.

Schuster:1987:THE

[SH87] Stefan Schuster and Reinhart Heinrich. Time hierarchy in enzymatic reaction chains resulting from optimality principles.

Journal of Theoretical Biology, 129(2):189–209, November 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519387800127.

Stehle:1989:LTD

[SHBS89] R. G. Stehle, N. F. H. Ho, C. L. Barsuhn, and K. J. Stefanski. Local topical delivery of drugs: a model incorporating simultaneous diffusion and metabolic interconversion between drug and a single metabolite in the skin. Journal of Theoretical Biology, 138(1): 1–15, May 7, 1989. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519389801742.

Shcherbak:1988:COS

[Shc88] V. I. Shcherbak. The co-operative symmetry of the genetic code. Journal of Theoretical Biology, 132(1):121–124, May 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388801966.

Shcherbak:1989:RRT

[Shc89a] V. I. Shcherbak. Rumer’s rule and transformation in the context of the co-operative symmetry of the genetic code. Journal of Theoretical Biology, 139(2):271–276, July 21, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389801043.

Shcherbak:1989:SSG

[Shc89b] V. I. Shcherbak. The “START” and “STOP” of the genetic code: Why exactly ATG and TAG, TAA? Journal of Theoretical Biology, 139(2):283–286, July 21, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389801067.

Shcherbak:1989:WWP

[Shc89c] V. I. Shcherbak. Ways of wobble pairing are formalized with the co-operative symmetry of the genetic code. Journal of Theoretical Biology, 139(2):277–281, July 21, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389801055.

Shears:1980:TGL

[She80] S. B. Shears. The thyroid gland and the liver mitochondrial protonic electrochemical potential difference: a novel hormone action? Journal of Theoretical Biology, 82(1):1–13, January 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380900879.

Shears:1981:MPE

[She81] S. B. Shears. The mitochondrial protonic electrochemical potential difference as a point of hormone action. II. New proposals for the activity of glucagon. Journal of Theoretical Biology, 91(1): 171–189, July 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381903799.

Sheehe:1983:VRC

[She83] Paul R. Sheehe. A variable risk clinical prognostic compartmental model. Journal of Theoretical Biology, 105(3):511–523, December 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251938390190X.

Shinagawa:1980:IHD

[Shi80] Y. Shinagawa. Invalidity of the Henderson diffusion equation shown by the exact solution of the Nernst–Planck equations. Journal of Theoretical Biology, 83(2):359–364, March 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380902982.

Shipman:1981:TSB

[Shi81] Lester L. Shipman. Theoretical study of the binding site and mode of action for photosystem II herbicides. Journal of Theoretical Biology, 90(1):123–148, May 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381901260.

Shiner:1982:TIF

[Shi82]  J. S. Shiner.  Transient intermediates, fast “equilibria”

and the rate of enzyme catalysis. Journal of Theoretical Biology, 99(2):309–318, November 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382900078.

Shipley:1987:RBD

[Shi87] Bill Shipley. The relationship between dynamic game theory and the Lotka–Volterra competition equations. Journal of Theoretical Biology, 125(1):121–123, March 7, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387801856.

Shore:1985:CSL

[Sho85] Harold Shore. Complementary spatial locations, width, and disparity. Journal of Theoretical Biology, 113(4):673–702, April 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385801867.

Shull:1984:DIP

[Shu84] Gary E. Shull. Differential inhibition of protein synthesis: a possible biochemical mechanism of thalidomide teratogenesis. Journal of Theoretical Biology, 110(3):461–486, October 7, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801873.

Shimony:1981:TSE

[SI81] Tamar B. Shimony and Jacob S. Ishay. Thermoelectric (seebeck) effect of the cuticle of social wasps. Journal of Theoretical Biology, 92(4):497–503, October 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902629.

Sibatani:1981:PCO

[Sib81] Atuhiro Sibatani. The polar co-ordinate model for pattern regulation in epimorphic fields: a critical appraisal. Journal of Theoretical Biology, 93(2):433–489, November 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381901144.

Sibatani:1983:PMI

[Sib83] Atuhiro Sibatani. A plausible molecular interpretation of the polar coordinate model with a centripetal degeneracy of distinctive field values. Journal of Theoretical Biology, 103(3):421–428, August 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383902953.

Siegel:1981:SAG

[Sie81] Armand Siegel. Stochastic aspects of the generation of the electroencephalogram. Journal of Theoretical Biology, 92(3):317–339, October 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381902952.

Silvester:1983:SHA

[Sil83] N. R. Silvester. Some hydrodynamic aspects of fitter feeding with rectangular-mesh nets. Journal of Theoretical Biology, 103(2):265–286, July 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519383900280.

Silk:1984:LRC

[Sil84] Joan B. Silk. Local resource competition and the evolution of male-biased sex ratios. Journal of Theoretical Biology, 108(2):203–213, May 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519384800661.

Silk:1989:GRP

[Sil89] Wendy Kuhn Silk. Growth rate patterns which maintain a helical tissue tube. Journal of Theoretical Biology, 138(3):311–328, June 3, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519389801973.

Siminoff:1980:MVVa

[Sim80a] Robert Siminoff. Modeling of the vertebrate visual system. I. Wiring diagram of the cone retina. Journal of Theoretical Biology, 86(4):673–708, October 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380903069.

Siminoff:1980:MVVb

[Sim80b] Robert Siminoff. Modeling of the vertebrate visual system. II. Application to the turtle cone retina. Journal of Theoretical Biology, 87(2):307–347, November 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251938090363X.

Siminoff:1981:MVV

[Sim81a] Robert Siminoff. Modelling of the vertebrate visual system 3. Topological analysis of the cone mosaic. Journal of Theoretical Biology, 91(3):437–476, August 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902678.

Simon:1981:PMD

[Sim81b] I. Simon. Possible mechanism for the dynamic stabilization of protein structure. Journal of Theoretical Biology, 90(4):487– 493, June 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381903015.

Simon:1985:IPR

[Sim85] I. Simon. Investigation of protein refolding: a special feature of native structure responsible for refolding ability. Journal of Theoretical Biology, 113(4):703–710, April 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801879.

Simon:1986:PGC

[Sim86] I. Simon. Proteins as general crystals. Journal of Theoretical Biology, 123(1):121–124, November 7, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386802399.

Singh:1982:OSO

[Sin82] U. N. Singh. On the origin of a self-organizing system: Structural versus functional organization. Journal of Theoretical Biology, 99(1):15–20, November 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382903861.

Sinha:1983:COS

[Sin83] Somdatta Sinha. Are circadian oscillators structurally unstable? Journal of Theoretical Biology, 101(3):401–414, April 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383901479.

Sirotkin:1986:AMT

[Sir86] Karl Sirotkin. Advantages to mutagenesis techniques generating populations containing the complete spectrum of single codon changes. Journal of Theoretical Biology, 123(3):261–279, December 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386802429.

Sizuo:1987:PNP

[Siz87] Yokota Sizuo. A proof of a neglected proof. Journal of Theoretical Biology, 129(4):453–455, December 21, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387800243.

Staron:1981:MMB

[SJ81] Krzysztof Staron and Andrzej Jerzmanowski. Model of mitosis based on antagonistic changes of intracellular concentration of free calcium and magnesium ions. Journal of Theoretical Biology, 89(3):501–511, April 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381903647.

Sanford:1985:CPD

[SJ85] J. C. Sanford and S. A. Johnston. The concept of parasite-derived resistance — deriving resistance genes from the parasite’s own genome. Journal of Theoretical Biology, 113(2):395–405, March 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385802344.

Sung:1989:CPP

[SJ89] Shen-Shu Sung and Peter C. Jordan. The channel properties of possible gramicidin dimers. Journal of Theoretical Biology, 140(3):369–380, October 9, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389800931.

Stephens:1988:MPC

[SJK88] Trent D. Stephens, Alfred W. Jackson, and Lawrence J. Kratz. A method for predicting the cranio-caudal position of secondary embryonic structures. Journal of Theoretical Biology, 135(4):433–444, December 21, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388802686.

Solovyov:1984:SMC

[SK84] Victor V. Solovyov and Nicolay A. Kolchanov. A simple method for the calculation of low energy packings of α-helices — a threshold approximation. I. The use of the method to estimate the effects of amino acid substitutions, deletions and insertions in globins. Journal of Theoretical Biology, 110(1):67–91, September 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384800168.

Schaffer:1985:NOD

[SK85a] William M. Schaffer and Mark Kot. Nearly one dimensional dynamics in an epidemic. Journal of Theoretical Biology, 112 (2):403–427, January 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519385802940.

Shindyalov:1985:AFI

[SK85b] Iljya N. Shindyalov and Nicolay A. Kolchanov. Analysis of the factors and implications of an empirical method for estimating the stability of mutant human haemoglobins. Journal of Theoretical Biology, 117(1):19–46, November 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801636.

Sekimura:1986:SPF

[SK86] Toshio Sekimura and Youichi Kobuchi. A spatial pattern formation model for Dictyostelium discoideum. Journal of Theoretical Biology, 122(3):325–338, October 7, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801242.

Snyder:1987:MAS

[SK87] Richard Snyder and Roy King. Mathematical analysis of separation-induced vocalizations in squirrel monkeys. Journal of Theoretical Biology, 125(3):361–366, April 7, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387800681.

Schoner:1988:DPT

[SK88a] G. Schöner and J. A. S. Kelso. A dynamic pattern theory of behavioral change. Journal of Theoretical Biology, 135(4):501–524, December 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251938880273X.

Skowlund:1988:CRC

[SK88b] Christopher T. Skowlund and Dale W. Kirmse. Comparison of the relative concentration of motile and non-motile bacteria in small pores. Journal of Theoretical Biology, 133(4):499–512, August 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388803370.

Stephan:1983:RTMb

[SKF83] W. Stephan, B. Kleutsch, and E. Frehland. Rate theory models for ion transport through rigid: Pores. III. Continuum vs discrete models in single file diffusion. Journal of Theoretical Biology, 105(2):287–310, ???? 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519383800095.

Saud:1984:DO

[SKK84] Ibne Saud, M. Z. Khan, and A. A. Khan. Diversity of order α. Journal of Theoretical Biology, 107(4):521–530, April 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384801307.

Shapiro:1980:GSR

[SL80a] Douglas Y. Shapiro and Roger Lubbock. Group sex ratio and sex reversal. Journal of Theoretical Biology, 83(3):411–426, April 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251938090048X.

Soloway:1980:PEM

[SL80b] A. H. Soloway and P. W. Lequesne. Potential endogenous mutagens/carcinogens. Journal of Theoretical Biology, 85(1):153– 163, July 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519380902866.

Savageau:1981:OKP

[SL81] Michael A. Savageau and David S. Lapointe. Optimization of kinetic proofreading: a general method for derivation of the constraint relations and an exploration of a specific case. Journal of Theoretical Biology, 93(1):157–177, November 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251938190062X.

Silverman:1986:MDP

[SL86] B. D. Silverman and R. Linsker. A measure of DNA periodicity. Journal of Theoretical Biology, 118(3):295–300, February 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386800601.

Saunders:1987:DCOa

[SL87a] D. S. Saunders and R. D. Lewis. A damped circadian oscillator model of an insect photoperiodic clock: II. Simulations of the shapes of the photoperiodic response curves. Journal of Theoretical Biology, 128(1):61–71, September 7, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387800310.

Saunders:1987:DCOb

[SL87b] D. S. Saunders and R. D. Lewis. A damped circadian oscillator model of an insect photoperiodic clock: III. Circadian and “hourglass” responses. Journal of Theoretical Biology, 128(1):73–85, September 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387800322.

Slack:1980:STT

[Sla80] J. M. W. Slack. A serial threshold theory of regeneration. Journal of Theoretical Biology, 82(1):105–140, January 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380900922.

Slack:1985:HTM

[Sla85] J. M. W. Slack. Homoeotic transformations in man: Implications for the mechanism of embryonic development and for the organization of epithelia. Journal of Theoretical Biology, 114(3):463– 490, June 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251938580179X.

Sladek:1986:HMM

[Sla86] Todd L. Sladek. A hypothesis for the mechanism of mycoplasma evolution. Journal of Theoretical Biology, 120(4):457–465, June 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938680039X.

Stamato:1986:CMS

[SLFT86] F. M. L. G. Stamato, E. Longo, R. Ferreira, and O. Tapia. The catalytic mechanism of serine proteases. III an indo-ISCRF study of the methylacetate docking in α-chymotrypsin. Journal of Theoretical Biology, 118(1):45–59, January 7, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386800078.

Smith:1984:IRP

[SLN84] S. Jay Smith, Dale N. Lawrence, and Gary R. Noble. An immune response profile model for immunogenicity quantitation. Journal of Theoretical Biology, 110(1):1–10, September 7, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384800107.

Stamato:1984:CMS

[SLYF84] F. M. L. G. Stamato, E. Longo, L. M. Yoshioka, and R. C. Ferreira. The catalytic mechanism of serine proteases: Single proton versus double proton transfer. Journal of Theoretical Biology, 107(2):329–338, March 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384800326.

Smith:1980:DDT

[SM80a] R. H. Smith and R. Mead. The dynamics of discrete-time stochastic models of population growth. Journal of Theoretical Biology, 86(4):607–627, October 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380903021.

Smith:1980:PES

[SM80b] R. H. Smith and R. Mead. On predicting extinction in simple population models II. Numerical approximations. Journal of Theoretical Biology, 82(3):525–535, February 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380902544.

Schoffeniels:1981:IED

[SM81a] E. Schoffeniels and D.-G. Mărgineanu. Intramembrane events during the nerve impulse. Journal of Theoretical Biology, 92(1):1–13, September 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519381901569.

Silvius:1981:NLA

[SM81b] John R. Silvius and Ronald N. McElhaney. Non-linear Arrhenius plots and the analysis of reaction and motional rates in biological membranes. Journal of Theoretical Biology, 88(1):135–152, January 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381903325.

Srivastava:1981:PGC

[SM81c] S. K. Srivastava and P. C. Mishra. Phototautomerism of the G–C base pair of DNA. Journal of Theoretical Biology, 93(3):655–666, December 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381902277.

Sugino:1987:AVR

[SM87] K. Sugino and H. Machemer. Axial-view recording: An approach to assess the third dimension of the ciliary cycle. Journal of Theoretical Biology, 125(1):67–82, March 7, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387801807.

Segel:1988:GMA

[SM88a] Irwin H. Segel and Robert L. Martin. The general modifier (“allosteric”) unireactant enzyme mechanism: Redundant conditions for reduction of the steady state velocity equation to one that is first degree in substrate and effector. Journal of Theoretical Biology, 135(4):445–453, December 21, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388802698.

Spach:1988:NTP

[SM88b] C. Spach and R. Motta. Neonatal thymus provides all the information required for tolerance induction against all the nonthymic organ-specific minor histocompatibility antigens. Journal of Theoretical Biology, 131(1):69–73, March 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388801218.

Stokke:1986:HEM

[SME86] Bjørn Torger Stokke, Arne Mikkelsen, and Arnljot Elgsaeter. The human erythrocyte membrane skeleton may be an ionic gel. III. Micropipette aspiration of unswollen erythrocytes. Journal of Theoretical Biology, 123(2):205–211, November 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801540.

Smith:1980:TDM

[SMGD80] K. C. Smith, C. E. Magnuson, J. D. Goeschl, and D. W. DeMichele. A time-dependent mathematical expression of the münch hypothesis of phloem transport. Journal of Theoretical Biology, 86(3):493–505, October 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380903483.

Smith:1980:RA

[Smi80] Richard J. Smith. Rethinking allometry. Journal of Theoretical Biology, 87(1):97–111, November 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380902222.

Smith:1982:DAC

[Smi82a] J. Maynard Smith. Do animals convey information about their intentions? Journal of Theoretical Biology, 97(1):1–5, July 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382902715.

Smith:1982:MRF

[Smi82b] Richard J. Smith. On the mechanical reduction of functional morphology. Journal of Theoretical Biology, 96(1):99–106, May 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382901606.

Smith:1982:MLL

[Smi82c] Stuart W. Smith. A model of labelled-ligand displacement assays resulting in logit-log relationships. Journal of Theoretical Biology, 98(3):475–499, October 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382901321.

Smith:1984:DRS

[Smi84] Richard J. Smith. Determination of relative size: The “criterion of subtraction” problem in allometry. Journal of Theoretical Biology, 108(1):131–142, May 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801745.

Smith:1985:U

[Smi85a] Howard Smith. Urospasm. Journal of Theoretical Biology, 112(4):873–875, February 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519385800680.

Smith:1985:SSH

[Smi85b] J. Maynard Smith. Sexual selection, handicaps and true fintess. Journal of Theoretical Biology, 115(1):1–8, July 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800035.

Smith:1987:EOF

[Smi87] J. Maynard Smith. On the equality of origin and fixation times in genetics. Journal of Theoretical Biology, 128(2):247–252, September 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387801728.

Smith:1988:CMS

[Smi88] J. Maynard Smith. Can a mixed strategy be stable in a finite population? Journal of Theoretical Biology, 130(2):247–251, January 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388801000.

Sharpe:1980:DMH

[SN80] Peter J. H. Sharpe and Alan W. Nordheim. Distribution model of human ovulatory cycles. Journal of Theoretical Biology, 83(4):663–673, April 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380901940.

Stenberg:1985:DLR

[SN85] Manne Stenberg and Håkan Nygren. A diffusion limited reaction theory for a solid-phase immunoassay. Journal of Theoretical Biology, 113(3):589–597, April 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800400.

Sculley:1981:IBM

[SNW81] M. J. Sculley, L. W. Nichol, and D. J. Winzor. Interactions between micellar ligand systems and acceptors: Forms of binding curves. Journal of Theoretical Biology, 90(3):365–376, June 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381903179.

Snyder:1988:TCP

[Sny88] Douglas M. Snyder. On the time of a conscious peripheral sensation. Journal of Theoretical Biology, 130(2):253–254, January 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388801012.

Svetina:1982:MBE

[SOLG82] S. Svetina, Angela Ottova-Leitmannová, and R. Glaser. Membrane bending energy in relation to bilayer couples concept of red blood cell shape transformations. Journal of Theoretical Biology, 94(1):13–23, January 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382903277.

Soumpasis:1980:GDS

[Sou80] Dikeos Soumpasis. Genetic distance: a simple quantitative measure of biological relatedness. Journal of Theoretical Biology, 86(1):137–147, September 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380900715.

Story:1984:SNA

[SP84] R. Story and Toni A. Paine. Some neglected aspects of vegetation classification in integrated surveys. Journal of Theoretical Biology, 108(4):663–674, June 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384800843.

Seligman:1987:MTD

[SP87] Stephen J. Seligman and Matthew R. Pincus. A model for the three-dimensional structure of peptidoglycan in staphylococci. Journal of Theoretical Biology, 124(3):275–292, February 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519387801169.

Savage:1988:MCD

[SP88] John R. K. Savage and D. G. Papworth. A method for comparing DNA replication programmes at the level of the chromosome bands. Journal of Theoretical Biology, 134(3):365–377, October 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388800675.

Sparrow:1980:BCB

[Spa80] C. Sparrow. Bifurcation and chaotic behaviour in simple feedback systems. Journal of Theoretical Biology, 83(1):93–105, March 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380903732.

Spertell:1981:TEP

[Spe81] Robert B. Spertell. On the transmission of external pressure fluctuations to the cerebrospinal fluid. Journal of Theoretical Biology, 88(2):309–321, January 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381900783.

Spencer:1984:ATA

[Spe84] N. Spencer. Amino terminal acetylation and protein synthesis in eukaryotes. Journal of Theoretical Biology, 107(3):405–415, April 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384800995.

Speakman:1986:OSS

[Spe86] J. R. Speakman. The optimum search speed of terrestrial predators when feeding on sedentary prey: a predictive model. Journal of Theoretical Biology, 122(4):401–407, October 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801813.

Speakman:1987:CCP

[Spe87] J. R. Speakman. Calculation of CO 2 production in doublylabelled water studies. Journal of Theoretical Biology, 126(1): 101–104, May 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387801030.

Starmer:1987:LBT

[SPG87] C. F. Starmer, D. L. Packer, and A. O. Grant. Ligand binding to transiently accessible sites: Mechanisms for varying apparent binding rates. Journal of Theoretical Biology, 124(3):335–341, February 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387801200.

Sherman:1989:CDO

[SPKJ89] Thomas F. Sherman, Aleksander S. Popel, Akos Koller, and Paul C. Johnson. The cost of departure from optimal radii in microvascular networks. Journal of Theoretical Biology, 136(3):245– 265, February 8, 1989. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519389801626.

Savage:1984:SPU

[SPP84] John R. K. Savage, R. Prasad, and D. G. Papworth. Subdivision of s-phase and its use for comparative purposes in cultured human cells. Journal of Theoretical Biology, 111(2):355–367, November 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384802155.

Schmidt:1983:KES

[SPS83] Norman D. Schmidt, Jacques J. Peschon, and Irwin H. Segel. Kinetics of enzymes subject to very strong product inhibition: Analysis using simplified integrated rate equations and average velocities. Journal of Theoretical Biology, 100(4):597–611, February 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383903259.

Sepulveda:1983:SRB

[SPST83] A. Sepúlveda, M. Pieber, M. A. Soto, and J. C. Tohà. Storage and retrieval of biomolecule sequences. Journal of Theoretical Biology, 103(2):331–332, July 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383900334.

Schmitz:1980:GTO

[SR80a] Kenneth S. Schmitz and Brinda Ramanathan. Generation of a third-order folded structure for chromatin. Journal of Theoretical Biology, 83(2):297–311, March 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380902945.

Shaitan:1980:EIS

[SR80b] K. V. Shaitan and A. B. Rubin. Effects of isotope substitution on intermolecular electron transfer in biological systems. Journal of Theoretical Biology, 86(2):203–221, September 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380900016.

Shirazi:1981:EMD

[SR81] M. A. Shirazi and R. T. Riley. Estimation of molecular diffusivity in isolated animal tissues. Journal of Theoretical Biology, 93(4):1033–1036, December 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381903532.

Salmon:1987:MTL

[SR87a] Peter M. Salmon and Channing R. Robertson. Mass transfer limitations in gel beads containing growing immobilized cells. Journal of Theoretical Biology, 125(3):325–332, April 7, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387800644.

Speakman:1987:ECO

[SR87b] J. R. Speakman and P. A. Racey. The equilibrium concentration of oxygen-18 in body water: Implications for the accuracy of the doubly-labelled water technique and a potential new method of measuring RQ in free-living animals. Journal of Theoretical Biology, 127(1):79–95, July 7, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387801625.

Sinha:1988:CBR

[SR88] Somdatta Sinha and Ramakrishna Ramaswamy. Complex behaviour of the repressible operon. Journal of Theoretical Biology, 132(3):307–318, June 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388802170.

Speakman:1989:DLW

[SR89] J. R. Speakman and P. A. Racey. Doubly labelled water: Errors in the evaluation of oxygen isotope turnover due to temporal variation in CO 2 production do not always covary with dilution space estimate. Journal of Theoretical Biology, 141(4):547–556, December 19, 1989. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519389802358.

Sanyal:1984:MBD

[SRO84] Nitish K. Sanyal, Mihir Roychoudhury, and Rajendra P. Ojha. Molecular basis of drug action of some antibiotics. Journal of Theoretical Biology, 110(4):505–521, October 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801411.

Sanyal:1985:IES

[SRO85] Nitish K. Sanyal, Mihir Roychoudhury, and Rajendra Pd Ojha. Interaction energy studies of an antimetabolite 8-azaguanine during transcription. Journal of Theoretical Biology, 116(2):195–199, September 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385802630.

Salant:1989:MBC

[SRPS89] Evan P. Salant, J. Richard, L. Pink, and Charles M. Steinberg. A model of bursal colonization. Journal of Theoretical Biology, 139(1):1–16, July 10, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519389800542.

Sitaramam:1981:DSS

[SS81] V. Sitaramam and M. K. Janardana Sarma. Does sucrose space exist? Journal of Theoretical Biology, 90(3):317–336, June 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381903155.

Schuster:1982:NES

[SS82] Peter Schuster and Karl Sigmund. A note on the evolution of sexual dimorphism. Journal of Theoretical Biology, 94(1):107– 110, January 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382903320.

Schuster:1983:RD

[SS83a] Peter Schuster and Karl Sigmund. Replicator dynamics. Journal of Theoretical Biology, 100(3):533–538, February 7, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383904459.

Seim:1983:RAW

[SS83b] Eva Seim and Bernt-Erik Sæther. On rethinking allometry: which regression model to use? Journal of Theoretical Biology, 104(2):161–168, September 21, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383904083.

Sivakami:1983:MM

[SS83c] S. Sivakami and V. Srinivasan. A model for memory. Journal of Theoretical Biology, 102(2):287–294, May 21, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383903661.

Sauve:1985:IFI

[SS85] Rémy Sauvé and Gabor Szabo. Interpretation of 1/F fluctuations in ion conducting membranes. Journal of Theoretical Biology, 113(3):501–516, April 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800357.

Smith:1987:TSF

[SS87] D. A. Smith and S. Sicilia. The theory of sliding filament models for muscle contraction. I. The two-state model. Journal of Theoretical Biology, 127(1):1–30, July 7, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387801583.

Sennerstam:1988:EIR

[SS88] Roland Sennerstam and Jan-Olof Strömberg. Evidence for an intraclonal random shift between two types of cell cycle times in an embryonal carcinoma cell line. Journal of Theoretical Biology, 131(2):151–162, March 21, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388802327.

Savageau:1989:CAM

[SS89a] Michael A. Savageau and Albert Sorribas. Constraints among molecular and systemic properties: Implications for physiological genetics. Journal of Theoretical Biology, 141(1):93–115, November 8, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519389800116.

Soberon:1989:PDR

[SS89b] Jorge Soberón and Edna Shapiro. Population dynamics of a Rhizobium-legume interaction. A mathematical model. Journal of Theoretical Biology, 140(3):305–316, October 9, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389800888.

Schmidt:1980:NPS

[SSF80] Willi Schmidt, Ulrich Schneeweiss, and Eva-Maria Fabricius. A new principle of specificity in anaerobic infections-mathematical model of mitosis-stimulated bacterial growth. The application of the tumor-tetanus phenomenon. Journal of Theoretical Biology, 86(4):783–802, October 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380903112.

Schneeweiss:1987:AIM

[SSF87] Ulrich Schneeweiss, Willi Schmidt, and Eva-Maria Fabricius. Antibacterial immunity and its modeling in experimental tumor tetanus and wound tetanus of the mouse. Journal of Theoretical Biology, 126(1):15–32, May 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938780098X.

Schoolfield:1981:NLR

[SSM81] R. M. Schoolfield, P. J. H. Sharpe, and C. E. Magnuson. Non-linear regression of biological temperature-dependent rate models based on absolute reaction-rate theory. Journal of Theoretical Biology, 88(4):719–731, February 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902460.

Stenberg:1986:EDS

[SSN86] Manne Stenberg, Lars Stiblert, and Håkan Nygren. External diffusion in solid-phase immunoassays. Journal of Theoretical Biology, 120(2):129–140, May 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801692.

Shulman:1981:CFN

[SSW81] Marc J. Shulman, Charles M. Steinberg, and Nicholas Westmoreland. The coding function of nucleotide sequences can be discerned by statistical analysis. Journal of Theoretical Biology, 88(3):409–420, February 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902745.

Sprackling:1980:DMB

[ST80] M. T. Sprackling and J. Twigg. Dislocations in microtubular bundles within spermatozoa of the coccid insect neosteingelia texana. Journal of Theoretical Biology, 85(3):561–567, August 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251938090329X.

Shirane:1986:CDA

[ST86a] Kotaro Shirane and Takayuki Tokimoto. Chemical dynamic approach to synergetics and instability in biological systems. Journal of Theoretical Biology, 123(2):187–192, November 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386801527.

Shirane:1986:DAT

[ST86b] Kotaro Shirane and Takayuki Tokimoto. Dynamic approach to thermotropic and ionotropic phase transitions. Journal of Theoretical Biology, 118(3):345–350, February 7, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386800650.

Shirane:1988:NFN

[ST88] Kotaro Shirane and Takayuki Tokimoto. Network formation in negative charged membranes by two divalent cations and the catastrophe. Journal of Theoretical Biology, 130(2):223–227, January 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519388800973.

Stauffer:1981:DFS

[Sta81] Howard B. Stauffer. Distributions for Fraser’s spatial indicators. Journal of Theoretical Biology, 89(3):513–522, April 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381903659.

Starr:1985:WIW

[Sta85] Christopher K. Starr. What if workers in social hymenoptera were males? Journal of Theoretical Biology, 117(1):11–18, November 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385801624.

Starmer:1986:TCI

[Sta86] C. Frank Starmer. Theoretical characterization of ion channel blockade: Ligand binding to periodically accessible receptors. Journal of Theoretical Biology, 119(2):235–249, March 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386800777.

Stauffer:1988:PTS

[Sta88] D. Stauffer. Percolation thresholds in square-lattice Kauffman model. Journal of Theoretical Biology, 135(2):255–261, November 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388800778.

Stein:1980:EAM

[Ste80] W. D. Stein. The epigenetic address: a model for embryological development. Journal of Theoretical Biology, 82(4):663–677, February 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251938090185X.

Stenseth:1981:CTF

[Ste81] Nils Chr. Stenseth. On Chitty’s theory for fluctuating populations: The importance of genetic polymorphism in the generation of regular density cycles. Journal of Theoretical Biology, 90(1): 9–36, May 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251938190120X.

Stein:1982:RRD

[Ste82] W. D. Stein. Replicating RNA duplexes and the epigenetic address. Journal of Theoretical Biology, 94(4):869–883, February 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382900820.

Stenseth:1984:ESS

[Ste84a] Nils Chr. Stenseth. Evolutionarily stable strategies in food selection models with fitness sets. Journal of Theoretical Biology, 109(4):489–499, August 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938480154X.

Stern:1984:SME

[Ste84b] Claudio D. Stern. A simple model for early morphogenesis. Journal of Theoretical Biology, 107(2):229–242, March 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384800247.

Stenseth:1985:EC

[Ste85] Nils Chr. Stenseth. On the evolution of cannibalism. Journal of Theoretical Biology, 115(2):161–177, July 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938580093X.

Steinberg:1987:RBS

[Ste87] Izchak Z. Steinberg. Relationship between statistical properties of single ionic channel recordings and the thermodynamic state of the channels. Journal of Theoretical Biology, 124(1):71–87, January 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387802539.

Steinberg:1988:CSE

[Ste88] I. Z. Steinberg. Computer simulations of the effect of noninactivating sodium channels on the electric behavior of excitable cells. Journal of Theoretical Biology, 133(2):193–214, July 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388800055.

Shinozuka:1984:SMM

[STN^{+}84] M. Shinozuka, A. Tsurui, T. Naganuma, M. Moss, and L. Moss-Salentijn. A stochastic-mechanical model of longitudinal long bone growth. Journal of Theoretical Biology, 108(3):413–436, June 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384800430.

Stout:1981:DSA

[Sto81] Darryl G. Stout. Dehydration strain avoidance and tolerance in plant cold hardiness. Journal of Theoretical Biology, 88(3):513– 521, February 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381902800.

Stockert:1983:NEI

[Sto83] Juan C. Stockert. A neighbour-exclusion intercalating model for poly(ADP-ribose) binding to DNA. Journal of Theoretical Biology, 105(3):461–467, December 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251938390187X.

Stone:1984:ISR

[Sto84a] M. Stone. Informativeness in stimulus-response modelling. Journal of Theoretical Biology, 111(2):261–272, November 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384802106.

Stone:1984:VCM

[Sto84b] M. Stone. Variance-covariance modelling with chromosome markers. Journal of Theoretical Biology, 107(2):275–286, March 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384800284.

Stockbridge:1989:TRB

[Sto89a] N. Stockbridge. Theoretical response of a bifurcating axon with a locally altered axial resistivity. Journal of Theoretical Biology, 137(3):339–354, April 6, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389800761.

Stockert:1989:IMG

[Sto89b] J. C. Stockert. An intercalative and minor groove binding model for the DNA cleavage reagent, copper(I) (1, 10-phenanthroline)

2. Journal of Theoretical Biology, 137(1):107–111, March 7, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519389801523.

Stone:1989:IDR

[Sto89c] M. Stone. Isobolic determination of “ratio-linear” parameters in pharmacological stimulus-response models. Journal of Theoretical Biology, 137(4):481–488, April 20, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389800414.

Strebel:1980:LGP

[Str80] Donald E. Strebel. Logistic growth in the presence of non-white environmental noise. Journal of Theoretical Biology, 85(4):713– 738, August 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519380902672.

Strogatz:1983:TZC

[Str83] Steven Strogatz. Topology of zigzag chromatin. Journal of Theoretical Biology, 103(4):601–607, August 21, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383902849.

Strandberg:1985:HPS

[Str85a] M. W. P. Strandberg. Homomorphism on a physical system of the Hodgkin–Huxley equations for ion conductance in nerve. Journal of Theoretical Biology, 117(4):509–527, December 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385802356.

Strandberg:1985:IPSa

[Str85b] M. W. P. Strandberg. Isomorphism on a physical system of the Hodgkin–Huxley equations for potassium conductance. Journal of Theoretical Biology, 117(2):161–185, November 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385802162.

Strandberg:1985:IPSb

[Str85c] M. W. P. Strandberg. Isomorphism on a physical system of the Hodgkin–Huxley equations for sodium conductance: Toxin allosteric effects and gating currents. Journal of Theoretical Biology, 117(3):385–396, December 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801508.

Stuart:1985:BIE

[Stu85a] C. I. J. M. Stuart. Bio-informational equivalence. Journal of Theoretical Biology, 113(4):611–636, April 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801831.

Stuart:1985:PMB

[Stu85b] C. I. J. M. Stuart. Physical models of biological information and adaptation. Journal of Theoretical Biology, 113(3):441–454, April 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385800321.

Stubblefield:1986:TDC

[Stu86] Elton Stubblefield. A theory for developmental control by a program encoded in the genome. Journal of Theoretical Biology, 118(2):129–143, January 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801291.

Suciu:1983:CDA

[Suc83] Dan Suciu. Cellular death by apoptosis in some radiosensitive and radioresistant mammalian tissues. Journal of Theoretical Biology, 105(3):391–401, December 7, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383901832.

Suciu:1985:MSI

[Suc85a] Dan Suciu. Morphometric study of the interphase nucleus in some radiosensitive and radioresistant mammalian cells. Journal of Theoretical Biology, 113(4):599–609, April 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938580182X.

Suciu:1985:NVC

[Suc85b] Dan Suciu. Nuclear volume and chromatin organization in some radiosensitive and radioresistant mammalian cells. Journal of Theoretical Biology, 117(4):587–596, December 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938580240X.

Suhnel:1989:MSEb

[Süh89] Jürgen Sühnel. Modeling survival of experimentally virus-infected laboratory animals-exploration of the survival diagram. Journal of Theoretical Biology, 141(1):53–63, November 8, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389800074.

Sukhodolets:1989:GCC

[Suk89] Vitaly V. Sukhodolets. The genetic code as a clue to understanding of molecular evolution. Journal of Theoretical Biology, 141(3):379–389, December 7, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389801201.

Sitaram:1984:ICO

[SV84a] B. R. Sitaram and V. S. Varma. Information content and one point mutations in DNA. Journal of Theoretical Biology, 110(4):523–532, October 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801423.

Sitaram:1984:SMG

[SV84b] B. R. Sitaram and V. S. Varma. Statistical mechanics of the Gompertz model of interacting species. Journal of Theoretical Biology, 110(2):253–256, September 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384800570.

Suhnel:1989:MSEa

[SV89] Jurgen Sühnel and Anneliese Veckenstedt. Modeling survival of experimentally virus-infected laboratory animals. Journal of Theoretical Biology, 137(1):27–39, March 7, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938980147X.

Svetina:1981:PIP

[Sve81] S. Svetina. Protein induction process and stochastic nature of cell commitment to proliferation and differentiation. Journal of Theoretical Biology, 90(1):151–158, May 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381901284.

Svetina:1982:RAV

[Sve82] S. Svetina. Relations among variations in human red cell volume, density, membrane area, hemoglobin content and cation content. Journal of Theoretical Biology, 95(1):123–134, March 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382902910.

Sulsky:1989:TDA

[SVN89] Deborah Sulsky, Richard R. Vance, and William I. Newman. Time delays in age-structured populations. Journal of Theoretical Biology, 141(3):403–422, December 7, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389801225.

Smith:1981:OGI

[SW81a] T. F. Smith and M. S. Waterman. Overlapping genes and information theory. Journal of Theoretical Biology, 91(2):379– 380, July 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381902393.

Stanford:1981:PMD

[SW81b] Janette M. Stanford and Tai Te Wu. A predictive method for determining possible three-dimensional foldings of immunoglobulin backbones around antibody combining sites. Journal of Theoretical Biology, 88(3):421–439, February 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902757.

Suyama:1983:CBT

[SW83] Akira Suyama and Akiyoshi Wada. Correlation between thermal stability maps and genetic maps of double-stranded DNAs.

Journal of Theoretical Biology, 105(1):133–145, November 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383904290.

Suh:1984:DTM

[SW84] Okhee Suh and Leonard Weiss. The development of a technique for the morphometric analysis of invasion in cancer. Journal of Theoretical Biology, 107(4):547–561, April 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801320.

Sibbald:1987:HPA

[SW87] Peter R. Sibbald and Michael J. White. How probable are antibody cross-reactions? Journal of Theoretical Biology, 127(2):163– 169, July 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387801273.

Swabb:1982:DFA

[Swa82] Edward A. Swabb. Dynamics of fluid accumulation in intestinal segments. Journal of Theoretical Biology, 95(3):517–530, April 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382900327.

Sykora:1984:DBV

[Sýk84] Peter Sýkora. The discrepancy between in vivo and in vitro experiments with polycyclic aromatic hydrocarbon(PAH) carcinogens: a hypothetical explanation. Journal of Theoretical Biology, 110(1):59–66, September 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384800156.

Sasaki:1986:TRB

[SYM86] Naoki Sasaki, Hidemi Yamamura, and Norio Matsushima. Is there a relation between bone strength and percolation? Journal of Theoretical Biology, 122(1):25–31, September 7, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386802235.

Syvanen:1984:CRM

[Syv84] Michael Syvanen. Conserved regions in mammalian β-globins: Could they arise by cross-species gene exchange? Journal of Theoretical Biology, 107(4):685–696, April 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801393.

Syvanen:1985:CSG

[Syv85] Michael Syvanen. Cross-species gene transfer; implications for a new theory of evolution. Journal of Theoretical Biology, 112(2):333–343, January 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519385802915.

Szarski:1983:CSC

[Sza83] Henryk Szarski. Cell size and the concept of wasteful and frugal evolutionary strategies. Journal of Theoretical Biology, 105(2): 201–209, ???? 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519383800022.

Szathmary:1988:HI

[Sza88] Eörs Szathmáry. A hypercyclic illusion. Journal of Theoretical Biology, 134(4):561–563, October 21, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388800572.

Szumiel:1983:CMA

[Szu83] Irena Szumiel. Ca^{2+}, Mg^{2+} and (adenosine Diphosphate ribose) n in cellular response to irradiation. Journal of Theoretical Biology, 101(3):441–451, April 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383901492.

Toro:1982:EVH

[TACM82] M. Toro, R. Abugov, B. Charlesworth, and R. E. Michod. Exact versus heuristic models of kin selection. Journal of Theoretical Biology, 97(4):699–713, August 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251938290368X.

Taiz:1986:BRP

[Tai86] Lincoln Taiz. Are biosynthetic reactions in plant cells thermodynamically coupled to glycolysis and the tonoplast proton motive force? Journal of Theoretical Biology, 123(2):231–238, November 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386801576.

Takahashi:1987:MST

[Tak87] Manabu Takahashi. A model for the spatio-temporal organization of DNA replication in mammalian cells. Journal of Theoretical Biology, 129(1):91–115, November 7, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387802059.

Takahashi:1989:FMC

[Tak89a] Manabu Takahashi. A fractal model of chromosomes and chromosomal DNA replication. Journal of Theoretical Biology, 141 (1):117–136, November 8, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519389800128.

Takahashi:1989:MCD

[Tak89b] Manabu Takahashi. A model of chromatin-dependent DNA replication sequences based on the decondensation units hypothesis. Journal of Theoretical Biology, 136(4):427–465, February 22, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519389801572.

Takenaka:1989:OLP

[Tak89c] Akio Takenaka. Optimal leaf photosynthetic capacity in terms of utilizing a natural light environment. Journal of Theoretical Biology, 139(4):517–529, August 22, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389800682.

Tamachi:1987:EAC

[Tam87] Nobuo Tamachi. The evolution of alarm calls: an altruism with nonlinear effect. Journal of Theoretical Biology, 127(2):141– 153, July 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251938780125X.

Tanaka:1980:AIT

[Tan80] Masahiro Tanaka. An application of information theory to biological evolution. Journal of Theoretical Biology, 85(4):789–806, August 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519380902726.

Tanaka:1984:PCB

[Tan84] Masahiro Tanaka. A physical characterization of biological information and communication system model of ecosystems. Journal of Theoretical Biology, 110(4):619–635, October 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801484.

Tatner:1988:MNA

[Tat88] Paul Tatner. A model of the natural abundance of oxygen-18 and deuterium in the body water of animals. Journal of Theoretical Biology, 133(3):267–280, August 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388803217.

Tsokos:1981:SAI

[TATS81] Janice O. Tsokos, Juan J. Angulo, Chris P. Tsokos, and Truett L. Smith. Sequential analysis of the influence of sex on the spread of the disease variola minor. Journal of Theoretical Biology, 89(3):341–351, April 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381903568.

Taylor:1981:FBA

[Tay81] Martin Taylor. Foraging behaviour of ants: a correction. Journal of Theoretical Biology, 93(1):275–277, November 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381900709.

Taylor:1985:EES

[Tay85a] Fritz Taylor. Estimating the ends of the sensitive period for photoperiodically induced diapause in arthropods. Journal of Theoretical Biology, 117(3):319–336, December 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801478.

Taylor:1985:GMM

[Tay85b] Peter D. Taylor. A general mathematical model for sex allocation. Journal of Theoretical Biology, 112(4):799–818, February 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938580062X.

Taylor:1986:CAA

[Tay86] William Ramsay Taylor. The classification of amino acid conservation. Journal of Theoretical Biology, 119(2):205–218, March 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386800753.

Taylor:1988:IFM

[Tay88a] Peter D. Taylor. An inclusive fitness model for dispersal of offspring. Journal of Theoretical Biology, 130(3):363–378, February 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388800353.

Taylor:1988:CSP

[Tay88b] Peter J. Taylor. Consistent scaling and parameter choice for linear and generalized Lotka–Volterra models used in community ecology. Journal of Theoretical Biology, 135(4):543–568, December 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388802753.

Taylor:1988:CTC

[Tay88c] Peter J. Taylor. The construction and turnover of complex community models having generalized Lotka–Volterra dynamics. Journal of Theoretical Biology, 135(4):569–588, December 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519388802765.

Taylor:1989:DIC

[Tay89a] Fritz Taylor. Diapause induction in changing photoperiods. Journal of Theoretical Biology, 139(1):103–116, July 10, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389800608.

Taylor:1989:LSS

[Tay89b] Peter D. Taylor. Local stability in a seed provisioning model. Journal of Theoretical Biology, 136(1):1–12, January 9, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389801845.

Taylor:1980:LMC

[TB80] P. D. Taylor and M. G. Bulmer. Local mate competition and the sex ratio. Journal of Theoretical Biology, 86(3):409–419, October 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380903422.

Treisman:1980:PFE

[TC80a] Michel Treisman and John S. Collins. Prior fitness and the evolutionarily stable strategy for intraspecific conflict. Journal of Theoretical Biology, 84(3):567–573, June 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519380800208.

Tuckwell:1980:ANI

[TC80b] Henry C. Tuckwell and Davis K. Cope. Accuracy of neuronal interspike times calculated from a diffusion approximation. Journal of Theoretical Biology, 83(3):377–387, April 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380900454.

Temme:1987:BSA

[TC87] David H. Temme and Eric L. Charnov. Brood size adjustment in birds: Economical tracking in a temporally varying environment. Journal of Theoretical Biology, 126(2):137–147, May 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519387802254.

Teasdale:1985:SAC

[TCR85] R. D. Teasdale, A. R. Carr, and R. S. D. Read. Substrate aggregation and cooperative enzyme kinetics: Consideration of enzyme access with large aggregates. Journal of Theoretical Biology, 114(3):375–382, June 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801727.

Toko:1985:DSC

[TCY85] Kiyoshi Toko, Hiroyuki Chosa, and Kaoru Yamafuji. Dissipative structure in the characea: Spatial pattern of proton flux as a dissipative structure in characean cells. Journal of Theoretical Biology, 114(1):125–175, May 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385802605.

Troitsky:1981:CCT

[TD81] N. A. Troitsky and S. E. Dromashko. Conjugational chromosome transfer-complete or partial? Kinetics of chromosome transfer in bacterial conjugation. Journal of Theoretical Biology, 90(2):283– 291, May 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381900473.

Tyson:1986:SSC

[TD86] John J. Tyson and Odo Diekmann. Sloppy size control of the cell division cycle. Journal of Theoretical Biology, 118(4):405–426, February 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251938680162X.

Taylor:1986:TEC

[TDA86] M. J. Taylor, C. J. Duggleby, and T. Atkinson. Is there evidence for a common amino acid sequence in proteins with membrane attaching ability? Journal of Theoretical Biology, 121(3):293– 306, August 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519386801096.

Tenforde:1980:TAE

[Ten80] Tom S. Tenforde. Thermal aspects of electromagnetic field interactions with bound calcium ions at the nerve cell surface. Journal of Theoretical Biology, 83(3):517–521, April 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380900570.

Triffet:1980:IEF

[TG80] T. Triffet and H. S. Green. Information and energy flow in a simple nervous system. Journal of Theoretical Biology, 86(1):3–44, September 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519380900648.

Triffet:1983:CDP

[TG83] T. Triffet and H. S. Green. Calcium dynamics at a plastic synapse in APLYSIA. Journal of Theoretical Biology, 100(4):645–674, February 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383903284.

Takenaka:1984:SAT

[TG84] Toshifumi Takenaka and Hideki Gotoh. Simulation of axoplasmic transport. Journal of Theoretical Biology, 107(4):579–601, April 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384801344.

Thompson:1988:MFA

[TG88a] Richard L. Thompson and Narendra S. Goel. Movable finite automata (MFA) models for biological systems i: Bacteriophage assembly and operation. Journal of Theoretical Biology, 131(3):351–385, April 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519388802303.

Triffet:1988:ITE

[TG88b] T. Triffet and H. S. Green. Information transfer by electromagnetic waves in cortex layers. Journal of Theoretical Biology, 131(2):199–221, March 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388802376.

Tzeghai:1986:TMS

[TGP^{+}86] G. Tzeghai, P. Ganatos, R. Pfeffer, S. Weinbaum, and A. Nir. A theoretical model to study the effect of convection and leaky junctions on macromolecule transport in artery walls. Journal of Theoretical Biology, 121(2):141–162, July 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386800893.

Tanemura:1980:GMT

[TH80] Masaharu Tanemura and Masami Hasegawa. Geometrical models of territory I. Models for synchronous and asynchronous settlement of territories. Journal of Theoretical Biology, 82(3):477–496, February 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519380902519.

Thornley:1981:DSS

[TH81] A. L. Thornley and A. Harington. Diploidy and sex as the selective advantages for retaining genes transferred from mitochondrial and plastid ancestors in the nuclear genome. Journal of Theoretical Biology, 91(3):515–523, August 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251938190271X.

Tyson:1985:DCS

[TH85] John J. Tyson and Kenneth B. Hannsgen. The distributions of cell size and generation time in a model of the cell cycle incorporating size control and random transitions. Journal of Theoretical Biology, 113(1):29–62, March 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800746.

Thompson:1982:CSD

[Tho82] Robert W. Thompson. Comments on size dispersion in living systems. Journal of Theoretical Biology, 96(1):87–94, May 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382901588.

Thor:1983:STC

[Tho83] Gautam Thor. Some topological considerations of the optic pathway. Journal of Theoretical Biology, 101(1):129–136, March 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251938390276X.

Thomas:1985:IFH

[Tho85] David W. Thomas. An induced fit hypothesis for antigen recognition by T lymphocytes: a role for specific antigen retention structures on antigen-presenting cells. Journal of Theoretical Biology, 117(3):417–429, December 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801521.

Thornley:1986:GMR

[Tho86] J. H. M. Thornley. A germination model: Responses to time and temperature. Journal of Theoretical Biology, 123(4):481–492, December 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519386802156.

Tani:1989:KAI

[TIK89a] Seiji Tani, Akira Imamura, and Katsuya Kanda. Kinetic approach with ab initio MO method on ionic selectivity and size in sodium channel. Journal of Theoretical Biology, 140(4):519–536, October 23, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519389801122.

Toh:1989:WZP

[TIK89b] Hiroyuki Toh, Akira Imamura, and Katsuya Kanda. Whether Zn 2+ for the polymerase system was selected inevitably or by historical accident. Journal of Theoretical Biology, 136(1):79–85, January 9, 1989. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519389801912.

Tory:1983:SRS

[TJP83] E. M. Tory, W. S. Jodrey, and D. K. Pickard. Simulation of random sequential adsorption: Efficient methods and resolution of conflicting results. Journal of Theoretical Biology, 102(3):439– 445, June 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251938390379X.

Trams:1980:PRE

[TKB80] Eberhard G. Trams, Heather Kaufmann, and Geoffrey Burnstock. A proposal for the role of ecto-enzymes and adenylates in traumatic shock. Journal of Theoretical Biology, 87(3):609–621, December 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519380902398.

Tao:1986:CPC

[TL86] Dali Tao and Paul H. Li. Classification of plant cell cryoprotectants. Journal of Theoretical Biology, 123(3):305–310, December 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386802454.

Tenforde:1988:MDP

[TL88] T. S. Tenforde and R. P. Liburdy. Magnetic deformation of phospholipid bilayers: Effects on liposome shape and solute permeability at prephase transition temperatures. Journal of Theoretical Biology, 133(3):385–396, August 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388803291.

Troquet:1985:EFG

[TLN85] J. Troquet, Ph. Lambin, and C. V. Nelson. Experimental foundation of the Gabor-Nelson theory applied to boundaries which are non-insulating. Journal of Theoretical Biology, 114(3):399– 411, June 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385801752.

Tedgui:1983:IBF

[TLS83] Alain S. Tedgui, Bernard I. Levy, and Simon R. Sebag. Instantaneous blood flow, impedance and elastic properties computed in man from aortic pulse waves. Journal of Theoretical Biology, 101(3):345–354, April 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383901431.

Totafurno:1980:SFB

[TLT80] John Totafurno, Charles J. Lumsden, and L. E. H. Trainor. Structure and function in biological hierarchies: An Ising model approach. Journal of Theoretical Biology, 85(2):171–198, July 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251938090017X.

Talekar:1980:NCE

[TMGB80] S. V. Talekar, D. C. Mikulecky, and C. M. Gary-Bobo. A new current expression for selective ion permeation across membranes. Journal of Theoretical Biology, 87(4):663–669, ???? 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380901101.

Thuraisingham:1980:TSC

[TN80] R. A. Thuraisingham and S. H. M. Nilar. A theoretical study of the carcinogenic nature of some aromatic amines. Journal of Theoretical Biology, 86(3):577–580, October 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380903537.

Takizawa:1980:ASD

[TO80] Naosada Takizawa and Makoto Oonuki. Analysis of synapse density in cerebral cortex. Journal of Theoretical Biology, 82 (4):573–590, February 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519380901794.

Tomlinson:1989:CEM

[TO89] Ian Tomlinson and Peter O’Donald. The co-evolution of multiple female mating preferences and preferred male characters: The “gene-for-gene” hypothesis of sexual selection. Journal of Theoretical Biology, 139(2):219–238, July 21, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389801018.

Todd:1982:GMC

[Tod82] P. H. Todd. A geometric model for the cortical folding pattern of simple folded brains. Journal of Theoretical Biology, 97(3):529– 538, August 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382903800.

Todd:1985:GCP

[Tod85] P. H. Todd. Gaussian curvature as a parameter of biological surface growth. Journal of Theoretical Biology, 113(1):63–68, March 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385800758.

Tomita:1982:SMC

[Tom82] Kazuhisa Tomita. Structure and macroscopic chaos in biology. Journal of Theoretical Biology, 99(1):111–118, November 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382903939.

Tonkyn:1986:PMM

[Ton86] David W. Tonkyn. Predator-mediated mutualism: Theory and tests in the homoptera. Journal of Theoretical Biology, 118(1):15– 31, January 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519386800054.

Tooby:1982:PPE

[Too82] John Tooby. Pathogens, polymorphism, and the evolution of sex. Journal of Theoretical Biology, 97(4):557–576, August 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382903587.

Topham:1988:ICA

[Top88] Christopher M. Topham. Ill-conditioning associated with the “end-point” method for the determination of kinetic parameters describing irreversible enzyme inactivation by an unstable inhibitor. Journal of Theoretical Biology, 135(2):169–173, November 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519388800729.

Tornheim:1980:COC

[Tor80] Keith Tornheim. Co-ordinate control of phosphofructokinase and pyruvate kinase by fructose diphosphate: a mechanism for amplification and step changes in the regulation of glycolysis in liver. Journal of Theoretical Biology, 85(2):199–222, July 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380900181.

Toro:1982:ASR

[Tor82] Miguel A. Toro. Altruism and sex ratio. Journal of Theoretical Biology, 95(2):305–311, March 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902466.

Tourenne:1985:MEF

[Tou85] Christian J. Tourenne. A model of the electric field of the brain at EEG and microwave frequencies. Journal of Theoretical Biology, 116(4):495–507, October 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800850.

Tozeren:1989:AIM

[Toz89] Aydin Tozeren. Adhesion induced by mobile cross-bridges: Steady state peeling of conjugated cell pairs. Journal of Theoretical Biology, 140(1):1–17, September 11, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389800256.

Trautman:1985:TMN

[Tra85] Edwin D. Trautman. Tracer measurements of non-equilibrium volumes of distribution. Journal of Theoretical Biology, 115 (4):515–538, August 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519385801387.

Tranter:1986:PVE

[Tra86] G. E. Tranter. Parity-violating energy differences and the origin of biomolecular homochirality. Journal of Theoretical Biology, 119(4):467–479, April 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801965.

Treisman:1981:ELF

[Tre81a] Michel Treisman. Evolutionary limits to the frequency of aggression between related or unrelated conspecifics in diploid species with simple Mendelian inheritance. Journal of Theoretical Biology, 93(1):97–124, November 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251938190059X.

Treisman:1981:SIR

[Tre81b] Michel Treisman. The significance of immunity restriction by the major histocompatibility complex, and of the occurrence of high polymorphism at MHC loci: Two hypotheses. Journal of Theoretical Biology, 89(3):409–421, April 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381903593.

Treisman:1982:ESL

[Tre82] Michel Treisman. Evolutionarily stable levels of altruism in haploid and diploid species. Journal of Theoretical Biology, 97(3):437–479, August 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519382903757.

Treisman:1983:ETS

[Tre83] Michel Treisman. Errors in the theory of the structured deme. Journal of Theoretical Biology, 102(3):339–346, June 7, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383903727.

Troouft:1983:SVD

[Tro83] J. Troouft. Solution versus diffusion at the alveolar lining in the lung. Journal of Theoretical Biology, 102(2):333–336, May 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383903703.

Troyer:1984:MHE

[Tro84] Katherine Troyer. Microbes, herbivory and the evolution of social behavior. Journal of Theoretical Biology, 106(2):157–169, January 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251938490016X.

Trainor:1984:TRP

[TRS84] L. E. H. Trainor, G. W. Rowe, and V. L. Szabo. A tetrahedral representation of poly-codon sequences and a possible origin of codon degeneracy. Journal of Theoretical Biology, 108(3):459– 468, June 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384800466.

Toompuu:1980:GRF

[TS80] Oleg G. Toompuu and Victor P. Shcherbakov. Genetic recombination: Formal implications of a crossed strand-exchange between two homologous DNA molecules. Journal of Theoretical Biology, 82(3):497–520, February 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380902520.

Tarumi:1983:OSS

[TS83] Kazuaki Tarumi and Helmut Schwegler. Oscillating spatial structures of a tissue. Journal of Theoretical Biology, 101(3):373– 386, April 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383901455.

Todd:1984:SBE

[TS84] P. H. Todd and I. H. M. Smart. The shape of birds’ eggs. Journal of Theoretical Biology, 106(2):239–243, January 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519384900213.

Taylor:1985:ANA

[TS85] Dermot B. Taylor and Charles E. Spivak. The activation of the nicotinic acetylcholine receptor by the transmitter. Journal of Theoretical Biology, 112(3):653–666, February 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800291.

Toro:1986:AET

[TS86] M. Toro and L. Silio. Assortment of encounters in the twostrategy game. Journal of Theoretical Biology, 123(2):193–204, November 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519386801539.

Tokimoto:1988:SOC

[TS88a] Takayuki Tokimoto and Kotaro Shirane. A self-organized chemical model and reaction cascade. Journal of Theoretical Biology, 130(1):67–72, January 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388801644.

Tozeren:1988:ISG

[TS88b] Aydin Tözeren and Richard Skalak. Interaction of stress and growth in a fibrous tissue. Journal of Theoretical Biology, 130(3):337–350, February 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002251938880033X.

Turvey:1988:TAC

[TSRK88] M. T. Turvey, R. C. Schmidt, Lawrence D. Rosenblum, and Peter N. Kugler. On the time allometry of co-ordinated rhythmic movements. Journal of Theoretical Biology, 130(3):285–325, February 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519388800316.

Tevlin:1985:TVF

[TT85] Patrick Tevlin and L. E. H. Trainor. A two vector field model of limb regeneration and transplant phenomena. Journal of Theoretical Biology, 115(4):495–513, August 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801375.

Totafurno:1987:NLV

[TT87] John Totafurno and L. E. H. Trainor. A non-linear vector field model of supernumerary limb production in salamanders. Journal of Theoretical Biology, 124(4):415–454, February 21, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387802199.

Tuckwell:1987:DAC

[Tuc87] Henry C. Tuckwell. Diffusion approximations to channel noise. Journal of Theoretical Biology, 127(4):427–438, August 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519387801406.

Tuszynski:1988:PMR

[Tus88] J. A. Tuszyński. On possible mechanisms of rouleau formation in human erythrocytes. Journal of Theoretical Biology, 132(3):369– 373, June 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519388802212.

Tognetti:1980:SGM

[TW80a] Keith Tognetti and Graham Winley. Stochastic growth models with logistic mean population. Journal of Theoretical Biology, 82(1):167–169, January 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251938090096X.

Tuckwell:1980:RNC

[TW80b] Henry C. Tuckwell and Frederick Y. M. Wan. The response of a nerve cylinder to spatially distributed white noise inputs. Journal of Theoretical Biology, 87(2):275–295, November 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380903616.

Toko:1982:RCF

[TY82] Kiyoshi Toko and Kaoru Yamafuji. Rote of Ca^{2+} in the formation of an excitable membrane on the surface of a protoplasmic droplet of nitella. Journal of Theoretical Biology, 99(3):461–478, December 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382902065.

Tyson:1983:PES

[Tys83a] John J. Tyson. Periodic enzyme synthesis and oscillatory repression: Why is the period of oscillation close to the cell cycle time? Journal of Theoretical Biology, 103(2):313–328, July 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383900310.

Tyson:1983:UAM

[Tys83b] John J. Tyson. Unstable activator models for size control of the cell cycle. Journal of Theoretical Biology, 104(4):617–631, October 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383902503.

Tyson:1987:SCC

[Tys87] John J. Tyson. Size control of cell division. Journal of Theoretical Biology, 126(4):381–391, June 21, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387801467.

Ueda:1982:IDT

[UK82] Tetsuo Ueda and Yonosuke Kobatake. Initiation, development and termination of contraction rhythm in plasmodia of myxomycete Physarum polycephalum. Journal of Theoretical Biology, 97(1):87–93, July 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519382902806.

Urushibara:1987:SAC

[UKNS87] Seiichi Urushibara, Mitsuo Kawato, Kazuo Nakazawa, and Ryoji Suzuki. Simulation analysis of conduction of excitation in the atrioventricular node. Journal of Theoretical Biology, 126(3):275– 288, June 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387802357.

Ulanowicz:1980:HDNa

[Ula80a] Robert E. Ulanowicz. An hypothesis on the development of natural communities. Journal of Theoretical Biology, 85(2):i, July 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380900168.

Ulanowicz:1980:HDNb

[Ula80b] Robert E. Ulanowicz. An hypothesis on the development of natural communities. Journal of Theoretical Biology, 85(2):223– 245, July 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519380900193.

Umansky:1982:GPC

[Uma82] S. R. Umansky. The genetic program of cell death. Hypothesis and some applications: Transformation, carcinogenesis, ageing. Journal of Theoretical Biology, 97(4):591–602, August 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382903605.

Umeyama:1982:PVH

[UN82] Hideaki Umeyama and Setsuko Nakagawa. The pK_{a} value of his 57-Asp 102 couple in the active site of bovine pancreatic β-trypsin: A molecular orbital study. Journal of Theoretical Biology, 99(4):759–775, December 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382901965.

Upadhyaya:1983:MME

[URC83] S. K. Upadhyaya, R. H. Rand, and J. R. Cooke. A mathematical model of the effects of Co 2 on stomatal dynamics. Journal of Theoretical Biology, 101(3):415–440, April 7, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383901480.

Uryvaeva:1981:BSL

[Ury81] Irina V. Uryvaeva. Biological significance of liver cell polyploidy: an hypothesis. Journal of Theoretical Biology, 89(4):557–571, April 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251938190028X.

Voit:1985:DFD

[VA85] Eberhard O. Voit and Hermann Josef Anton. Derivation of the frequency distributions of cycle durations from continuous labeling curves. Journal of Theoretical Biology, 112(3):575–588, February 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385800230.

Vila:1988:AIH

[VA88a] J. Vila and J. L. Alessandrini. Adsorption of an ionic homopolypeptide to a flat surface induced by pH change. Journal of Theoretical Biology, 134(4):445–449, October 21, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938880050X.

Voit:1988:ECC

[VA88b] Eberhard O. Voit and Hermann J. Anton. Estimation of cell cycle parameters from double labeling experiments. Journal of Theoretical Biology, 131(4):435–440, April 21, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388800390.

Valentine:1985:TGM

[Val85] Harry T. Valentine. Tree-growth models: Derivations employing the pipe-model theory. Journal of Theoretical Biology, 117 (4):579–585, December 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519385802393.

VanHelden:1982:TCL

[Van82] Paul D. Van Helden. A theoretical consideration of the lysinerich histones: H1 from a mammal and an echinoderm, h5 from erythrocytes. Journal of Theoretical Biology, 96(3):327–336, June 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382901138.

Vandermeer:1984:PCY

[Van84] John Vandermeer. Plant competition and the yield-density relationship. Journal of Theoretical Biology, 109(3):393–399, August 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384800880.

Varshavsky:1983:DSR

[Var83] Alexander Varshavsky. Do stalled replication forks synthesize a specific alarmone? Journal of Theoretical Biology, 105 (4):707–714, December 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/002251938390228X.

Vautrin:1986:SQT

[Vau86] Jean Vautrin. Subunits in quantal transmission at the mouse neuromuscular junction: Tests of peak intervals in amplitude distributions. Journal of Theoretical Biology, 120(3):363–370, June 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386802077.

Vickers:1987:DES

[VC87] G. T. Vickers and C. Cannings. On the definition of an evolutionarily stable strategy. Journal of Theoretical Biology, 129 (3):349–353, December 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519387800073.

Vickers:1988:NSE

[VC88a] G. T. Vickers and C. Cannings. On the number of stable equilibria in a one-locus, multi-allelic system. Journal of Theoretical Biology, 131(3):273–277, April 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938880225X.

Vickers:1988:PE

[VC88b] G. T. Vickers and C. Cannings. Patterns of ESS’s I. Journal of Theoretical Biology, 132(4):387–408, June 21, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388800808.

Varon:1988:KMZ

[VCC^{+}88] R. Varón, F. García Cánovas, F. García Carmona, J. Tudela, A. Román, and A. Vázquez. Kinetics of a model for zymogen activation: The case of high activating enzyme concentrations. Journal of Theoretical Biology, 132(1):51–59, May 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388801905.

Victorri:1984:MMV

[VD84] Bernard Victorri and Jean-Robert Derome. Mathematical model of visual perception. Journal of Theoretical Biology, 108(2):227– 260, May 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384800697.

Viret:1989:BIM

[VD89] J. Viret and D. Daveloose. Biophysical interpretation of membrane fluidity by catastrophe theory. Journal of Theoretical Biology, 140(1):51–82, September 11, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938980030X.

Voigt:1982:FEE

[VDH82] Andreas Voigt, Edwin Donath, and Reinhart Heinrich. Free energy of the electrostatic interaction of cells with adjacent charged glycoprotein layer: a theoretical approach. Journal of Theoretical Biology, 98(2):269–282, September 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902648.

vanderHoeven:1984:MMC

[vdH84] Nelly van der Hoeven. A mathematical model for the co-existence of incompatible, conjugative plasmids in individual bacteria of a bacterial population. Journal of Theoretical Biology, 110(3):411– 423, October 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384801836.

Nelly:1985:EBS

[vdHN85] van der Hoeven Nelly. Evolution of bacterial surface exclusion against incompatible plasmids. Journal of Theoretical Biology, 117(3):431–452, December 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801533.

Vertregt:1987:RMD

[VdV87] N. Vertregt and F. W. T. Penning de Vries. A rapid method for determining the efficiency of biosynthesis of plant biomass. Journal of Theoretical Biology, 128(1):109–119, September 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519387800346.

vanDamme:1986:MGL

[vDvD86] Jos M. M. van Damme and Ruud van Damme. On the maintenance of gynodioecy: Lewis’ result extended. Journal of Theoretical Biology, 121(3):339–350, August 7, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801138.

Venieratos:1982:RCC

[Ven82] Dionyssis Venieratos. The role of competition in creating multiple steady states and favouring evolution of prebiotic polymers. Journal of Theoretical Biology, 95(2):313–323, March 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902478.

Venanzi:1984:HPB

[Ven84] Thomas J. Venanzi. Hydrophobicity parameters and the bitter taste of l-amino acids. Journal of Theoretical Biology, 111(3):447–450, December 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519384802325.

VanEykelenburg:1980:STC

[VFS80] C. Van Eykelenburg, A. Fuchs, and G. H. Schmidt. Some theoretical considerations on the in vitro shape of the cross-walls in spirulina spp. Journal of Theoretical Biology, 82(2):271–282, January 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519380901034.

VanDriessche:1980:ACV

[VG80] Willy Van Driessche and Heinz Gögelein. Attenuation of current and voltage noise signals recorded from epithelia. Journal of Theoretical Biology, 86(4):629–648, October 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380903033.

Vovelle:1980:HPI

[VGP^{+}80] F. Vovelle, M. Genest, M. Ptak, B. Maigret, and S. Premilat. Hydration of peptides. II. Determination of the preferred sites of interactions of a cyclic dipeptide with water. Journal of Theoretical Biology, 87(1):85–95, November 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380902210.

vanGurp:1988:OPB

[vGvGL88] M. van Gurp, G. van Ginkel, and Y. K. Levine. Orientational properties of biological pigments in ordered systems studied with polarized light: Photosynthetic pigment-protein complexes in membranes. Journal of Theoretical Biology, 131(3):333– 349, April 7, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519388802297.

vonHeijne:1987:TMP

[vHBL87] Gunnar von Heijne, Clas Blomberg, and Hans Liljenström. Theoretical modelling of protein synthesis. Journal of Theoretical Biology, 125(1):1–14, March 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387801753.

vonHeijne:1982:RSA

[vHS82] Gunnar von Heijne and Michael A. Savageau. RNA splicing: Advantages of parallel processing. Journal of Theoretical Biology, 98(4):563–574, October 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382901382.

vonHeijne:1987:HRX

[vHU87] Gunnar von Heijne and Mathias Uhlén. Homology to region X from staphylococcal protein A is not unique to cell surface proteins. Journal of Theoretical Biology, 127(3):373–376, August 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519387801133.

Vorotyntsev:1980:PPP

[VI80] M. A. Vorotyntsev and E. M. Itskovich. Primary photosynthetic processes: The problem of rapid irreversible redistribution of electronic energy. Journal of Theoretical Biology, 86(2):223–236, September 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519380900028.

Vickery:1987:HCA

[Vic87] William L. Vickery. How to cheat against a simple mixed strategy ESS. Journal of Theoretical Biology, 127(2):133–139, July 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519387801248.

Vickery:1988:RMS

[Vic88] William L. Vickery. Reply to Maynard Smith. Journal of Theoretical Biology, 132(3):375–378, June 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388802224.

Vickers:1989:SPE

[Vic89] G. T. Vickers. Spatial patterns and ESS’s. Journal of Theoretical Biology, 140(1):129–135, September 11, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389800335.

Vilu:1983:CPR

[Vil83] Raivo Vilu. Calculation of probabilities of random occurrence of different types of palindromes. Journal of Theoretical Biology, 102(2):261–268, May 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383903636.

Viswanadhan:1987:PFE

[Vis87] V. N. Viswanadhan. On the prediction of functionally equivalent residues at given sites in proteins. Journal of Theoretical Biology, 124(1):123–125, January 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387802576.

Vilstrup:1983:KGU

[VKV83] Hendrik Vilstrup, Susanne Keiding, and Per B. Vendsborg. Kinetics of galactose uptake by perfused rat livers: Applicability of a family of models. Journal of Theoretical Biology, 101(3):335– 344, April 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251938390142X.

Varenne:1986:EDU

[VL86] Stanislas Varenne and Claude Lazdunski. Effect of distribution of unfavourable codons on the maximum rate of gene expression by an heterologous organism. Journal of Theoretical Biology, 120(1): 99–110, May 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519386800200.

Vlad:1987:OTP

[Vla87] Marcel Ovidiu Vlad. The occurrence of time-persistent distribution functions in structured population dynamics. Journal of Theoretical Biology, 126(2):239–242, May 21, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387802308.

Voon:1986:SAU

[VLCK86] F. C. T. Voon, P. W. Lucas, K. L. Chew, and D. A. Kuke. A simulation approach to understanding the masticatory process. Journal of Theoretical Biology, 119(3):251–262, April 7, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801382.

Vila:1982:TAH

[VM82] J. Vila and E. Marvhi. Theoretical approach for α-helix → random coil denaturation. Journal of Theoretical Biology, 98 (2):253–268, September 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519382902636.

Vautrin:1989:SCB

[VM89] Jean Vautrin and Jean Mambrini. Synaptic current between neuromuscular junction folds. Journal of Theoretical Biology, 140(4):479–498, October 23, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389801109.

Voit:1985:CCG

[Voi85] Eberhard O. Voit. Cell cycles and growth laws: the CCC model. Journal of Theoretical Biology, 114(4):589–599, June 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938580045X.

Volkenstein:1981:SPP

[Vol81] M. K. Volkenstein. Simple physical presentation of enzymatic catalysis. Journal of Theoretical Biology, 89(1):45–51, March 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381901788.

VonHeune:1982:TSA

[Von82] Gunnar Von Heune. A theoretical study of the attenuation control mechanism. Journal of Theoretical Biology, 97(2):227–238, July 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251938290100X.

Varma:1986:AST

[VP86] V. S. Varma and L. K. Pande. Asymptotics of some three-species models. Journal of Theoretical Biology, 119(3):273–282, April 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386801400.

Vroman:1983:PRI

[Vro83] Leo Vroman. Possible relationships and interactions among events in clotting, platelet adhesion, immune surface reactions and granulocyte adhesion. Journal of Theoretical Biology, 105(3):541–543, December 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383901923.

vanRhijn:1980:BHA

[vRV80] Johan G. van Rhijn and Ron Vodegel. Being honest about one’s intentions: An evolutionary stable strategy for animal conflicts. Journal of Theoretical Biology, 85(4):623–641, August 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380902611.

vanStraalen:1983:PTT

[vS83] Nico M. van Straalen. Physiological time and time-invariance. Journal of Theoretical Biology, 104(3):349–357, October 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251938390111X.

vanStraalen:1985:PBT

[vS85] Nico M. van Straalen. Production and biomass turnover in stationary stage-structured populations. Journal of Theoretical Biology, 113(2):331–352, March 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385802319.

vanTienderen:1986:SRU

[vTdJ86] P. H. van Tienderen and G. de Jong. Sex ratio under the haystack model: Polymorphism may occur. Journal of Theoretical Biology, 122(1):69–81, September 7, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386802259.

Vidaver:1985:ELI

[VTL85] George A. Vidaver, Alexander Ting, and Jean W. Lee. Evidence that lysolecithin is an important causal agent of atherosclerosis.

Journal of Theoretical Biology, 115(1):27–41, July 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800059.

VanHouten:1982:CSP

[VV82] Judith Van Houten and J. Van Houten. Computer simulation of Paramecium chemokinesis behavior. Journal of Theoretical Biology, 98(3):453–468, October 7, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382901308.

vonZglinicki:1987:MMH

[vZ87] Thomas von Zglinicki. A mitochondrial membrane hypothesis of aging. Journal of Theoretical Biology, 127(2):127–132, July 21, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519387801236.

Valiquette:1985:MSP

[VZR85] G. Valiquette, E. A. Zimmerman, and J. L. Roberts. mRNA sequence predictions from homologous protein sequences. Journal of Theoretical Biology, 112(3):445–458, February 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800138.

vanZoelen:1981:SVI

[vZvdSdL81] E. J. J. van Zoelen, P. T. van der Saag, and S. W. de Laat. Sources of variation in intermitotic time and their relation to cell cycle kinetics. Journal of Theoretical Biology, 93(4):1037–1045, December 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381903544.

Wade:1982:EMI

[Wad82] Michael J. Wade. The effect of multiple inseminations on the evolution of social behaviors in diploid and haplo-diploid organisms. Journal of Theoretical Biology, 95(2):351–368, March 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519382902508.

Wade:1985:IMI

[Wad85] Michael J. Wade. The influence of multiple inseminations and multiple foundresses on social evolution. Journal of Theoretical Biology, 112(1):109–121, January 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801193.

Wagner:1981:MNM

[Wag81] H.-J. Wagner. The minimum number of mutations in an evolutionary network. Journal of Theoretical Biology, 91(4):621–636, August 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381902149.

Wagner:1982:LSI

[Wag82] G. P. Wagner. The logical structure of irreversible systems transformations: a theorem concerning Dollo’s law and chaotic movement. Journal of Theoretical Biology, 96(3):337–346, June 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251938290114X.

Wagg:1987:MDS

[Wag87] Jonathan Wagg. A method for defining steady-state unidirectional fluxes through branched chemical, osmotic and chemiosmotic reactions. Journal of Theoretical Biology, 128(3):375–385, October 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387800796.

Wagner:1989:RBR

[Wag89] Aurel Popa Wagner. On the relationships between the rate of cytoskeletal stable assemblies turnover, stability of the differentiated state, development and aging. Journal of Theoretical Biology, 138(2):175–184, May 22, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389801377.

Waldman:1987:MKR

[Wal87] Bruce Waldman. Mechanisms of kin recognition. Journal of Theoretical Biology, 128(2):159–185, September 21, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387801674.

Wallraff:1989:SNBb

[Wal89a] Hans G. Wallraff. Simulated navigation based on assumed gradients of atmospheric trace gases (models on pigeon homing, part 2). Journal of Theoretical Biology, 138(4):511–528, June 22, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519389800487.

Wallraff:1989:SNBa

[Wal89b] Hans G. Wallraff. Simulated navigation based on unreliable sources of information (models on pigeon homing. Part 1). Journal of Theoretical Biology, 137(1):1–19, March 7, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389801456.

Ward:1985:WBF

[War85] Paul I. Ward. Why birds in flocks do not coordinate their vigilance periods. Journal of Theoretical Biology, 114(3):383– 385, June 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385801739.

Warwicker:1986:CDM

[War86] J. Warwicker. Continuum dielectric modelling of the proteinsolvent system, and calculation of the long-range electrostatic field of the enzyme phosphoglycerate mutase. Journal of Theoretical Biology, 121(2):199–210, July 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386800935.

Wassermann:1982:TPTa

[Was82a] G. D. Wassermann. TIMA part 1. TIMA as a paradigm for the evolution of molecular complementarities and macromolecules. Journal of Theoretical Biology, 96(1):77–86, May 7, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382901576.

Wassermann:1982:TPTb

[Was82b] G. D. Wassermann. TIMA part 2. TIMA-based instructive evolution of macromolecules and organs and structures. Journal of Theoretical Biology, 99(3):609–628, December 7, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902144.

Wassermann:1986:NAO

[Was86a] G. D. Wassermann. Note on an abnormality of the operator for the structural gene of the dopamine D_{1} receptor as a possible partial cause of schizophrenia. Journal of Theoretical Biology, 120(3): 277–283, June 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519386802004.

Wassersug:1986:HDT

[Was86b] Richard Wassersug. How does a tadpole know when to metamorphose? A theory linking environmental and hormonal cues. Journal of Theoretical Biology, 118(2):171–181, January 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801321.

Watkinson:1980:DDS

[Wat80] A. R. Watkinson. Density-dependence in single-species populations of plants. Journal of Theoretical Biology, 83(2):345–357, March 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519380902970.

Waterman:1984:ESA

[Wat84] Michael S. Waterman. Efficient sequence alignment algorithms. Journal of Theoretical Biology, 108(3):333–337, June 7, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384800375.

Watanabe:1985:DPE

[Wat85] Tsuneo Watanabe. Dissipative processes of energetic heavy charged particles in matter. Journal of Theoretical Biology, 115(2):287–298, July 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519385801016.

Waxman:1981:CAB

[Wax81] Allen M. Waxman. A continuum approach to blood vessel growth: Axisymmetric elastic structures. Journal of Theoretical Biology, 91(2):273–301, July 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902344.

Wikswo:1982:PSN

[WB82] John P. Wikswo and John P. Barach. Possible sources of new information in the magnetocardiogram. Journal of Theoretical Biology, 95(4):721–729, April 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382903502.

White:1983:MCA

[WB83] J. G. White and G. G. Borisy. On the mechanisms of cytokinesis in animal cells. Journal of Theoretical Biology, 101(2):289– 316, March 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383903429.

Wagner:1985:EDM

[WB85] Günter P. Wagner and Reinhard Bürger. On the evolution of dominance modifiers II: a non-equilibrium approach to the evolution of genetic systems. Journal of Theoretical Biology, 113(3):475–500, April 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800345.

Wittrup:1988:MMR

[WB88] Karl Dane Wittrup and James E. Bailey. A mathematical model of recombinational amplification of the 2μ plasmid in the yeast Saccharomyces cerevisiae. Journal of Theoretical Biology, 130(4):481–492, February 21, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388802121.

Wadepuhl:1989:CSH

[WB89] Martin Wadepuhl and Wolf-Jürgen Beyn. Computer simulation of the hydrostatic skeleton. The physical equivalent, mathematics and application to worm-like forms. Journal of Theoretical Biology, 136(4):379–402, February 22, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389801559.

Wardell:1982:POC

[WBKW82] Jean M. Wardell, William G. Bardsley, John P. Kavanagh, and Reg M. Wood. The probability of obtaining complex kinetic curves for enzyme mechanisms with cubic terms in the pseudo-steady state rate equations. Journal of Theoretical Biology, 95(3):465–487, April 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382900303.

Weinand:1988:MIR

[WC88] Richard G. Weinand and Michael Conrad. Maturation of the immune response: a computational model. Journal of Theoretical Biology, 133(4):409–428, August 21, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938880331X.

Weiss:1981:NCS

[WD81] George H. Weiss and Ivan G. Darvey. A note on the choice of substrate concentration in enzyme kinetic experiments. Journal of Theoretical Biology, 90(3):437–439, June 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381903234.

Weiss:1986:MAL

[WD86] Leonard Weiss and Dimiter S. Dimitrov. Mechanical aspects of the lungs as cancer cell-killing organs during hematogenous metastasis. Journal of Theoretical Biology, 121(3):307–321, August 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519386801102.

Wohlgemuth:1987:ISI

[WD87] Andrew Wohlgemuth and Devendra P. Dubey. The impact of symbolism on immunogenetics: An application to HLA. Journal of Theoretical Biology, 126(2):149–165, May 21, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387802266.

Witten:1984:BPO

[WdlT84] Matthew Witten and Deborah L. de la Torre. Biological populations obeying difference equations: The effects of stochastic perturbation. Journal of Theoretical Biology, 111(3):493–507, December 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384802362.

Webster:1981:DOI

[Web81] Ian A. Webster. The dynamics of an open, isothermal homogeneous, single enzyme, two substrate system. Journal of Theoretical Biology, 91(3):417–428, August 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902654.

Weber:1983:CRE

[Web83] Charles E. Weber. Corticosteroid regulation of electrolytes. Journal of Theoretical Biology, 104(3):443–449, October 7, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383901169.

Weekley:1981:NNR

[Wee81] L. Bruce Weekley. On the neurogenic-neuroendocrine regulation of the pancreas. Journal of Theoretical Biology, 88(2):275–278, January 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381900758.

Weekley:1984:CNR

[Wee84a] L. Bruce Weekley. Central neuroendocrine regulation of renin secretion rate: a hypothesis. Journal of Theoretical Biology, 111(3):609–613, December 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384802404.

Weekley:1984:MWP

[Wee84b] L. Bruce Weekley. A mechanism by which primary or secondary hypothalamic involvement results in the development of insulin-dependent diabetes mellitus (IDDM). Journal of Theoretical Biology, 111(1):171–182, November 7, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384802039.

Watterson:1982:CIP

[WEHM82] G. A. Watterson, W. J. Ewens, T. E. Hall, and A. Morgan. The chromosome inversion problem. Journal of Theoretical Biology, 99(1):1–7, November 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382903848.

Weinbaum:1980:TFI

[Wei80] Sheldon Weinbaum. Theory for the formation of intercellular junctions based on intramembranous particle patterns observed in the freeze fracture technique. Journal of Theoretical Biology, 83(1):63–68, IN1–IN2, 69–92, March 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380903720.

Weiss:1983:ITD

[Wei83] M. Weiss. Importance of tissue distribution in determining drug disposition curves. Journal of Theoretical Biology, 103(4):649– 652, August 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383902886.

Weiss:1985:TLC

[Wei85] Michael Weiss. Theorems on log-convex disposition curves in drug and tracer kinetics. Journal of Theoretical Biology, 116(3):355– 368, October 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385802745.

Weisbuch:1986:NAB

[Wei86] G. Weisbuch. Networks of automata and biological organization. Journal of Theoretical Biology, 121(3):255–267, August 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386801060.

Weinberger:1988:MRD

[Wei88] Ed Weinberger. A more rigorous derivation of some properties of uncorrelated fitness landscapes. Journal of Theoretical Biology, 134(1):125–129, September 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388803072.

Wells:1983:PES

[Wel83] Harrington Wells. Population equilibria and stability in plantanimal pollination systems. Journal of Theoretical Biology, 100(4):685–699, February 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519383903302.

Welch:1985:SPU

[Wel85] G. Rickey Welch. Some problems in the usage of Gibbs free energy in biochemistry. Journal of Theoretical Biology, 114(3): 433–446, June 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385801776.

Wells:1986:PAP

[Wel86] Harrington Wells. Plant-animal pollination systems. 2. Fewer pollinator than plant species. Journal of Theoretical Biology, 122(4):375–384, October 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519386801795.

Werren:1987:CAC

[Wer87] John H. Werren. The coevolution of autosomal and cytoplasmic sex ratio factors. Journal of Theoretical Biology, 124(3):317–334, February 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519387801194.

Westall:1983:CBP

[Wes83a] FreD C. Westall. Cancer basic protein: a cellular growth factor in cancer cells? Journal of Theoretical Biology, 103(3):473– 480, August 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519383902990.

Westhaus:1983:TTD

[Wes83b] Paul Westhaus. The thermodynamics of a two dimensional system of hard core electric dipoles. Journal of Theoretical Biology, 100(1):181–184, January 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383901017.

Wichmann:1982:IHC

[WF82] H. E. Wichmann and K. Fesser. Influence of 3 HTdR on the circadian rhythm — a model analysis for mouse epidermis. Journal of Theoretical Biology, 97(3):371–391, August 7, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382903721.

Weigel:1986:MRH

[WFL86] Paul H. Weigel, Gerald M. Fuller, and Robert D. LeBoeuf. A model for the role of hyaluronic acid and fibrin in the early events during the inflammatory response and wound healing. Journal of Theoretical Biology, 119(2):219–234, March 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386800765.

Wichmann:1981:PSC

[WG81] H. E. Wichmann and M. D. Gerhardts. Platelet survival curves in man considering the splenic pool. Journal of Theoretical Biology, 88(1):83–101, January 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381903295.

Wheatley:1982:SRD

[WGHY82] Denys N. Wheatley, Santiago Grisolía, and Jose Hernández-Yago. Significance of the rapid degradation of newly synthesized proteins in mammalian cells: a working hypothesis. Journal of Theoretical Biology, 98(2):283–300, September 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251938290265X.

Waser:1983:CMM

[WGKG83] Mark R. Waser, Lillian Garfinkel, Michael C. Kohn, and David Garfinkel. Computer modeling of muscle phosphofructokinase kinetics. Journal of Theoretical Biology, 103(2):295–312, July 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383900309.

Weinbaum:1988:TDD

[WGP^{+}88] S. Weinbaum, P. Ganatos, R. Pfeffer, G. B. Wen, M. Lee, and S. Chien. On the time-dependent diffusion of macromolecules through transient open junctions and their subendothelial spread: I. Short-time model for cleft exit region. Journal of Theoretical Biology, 135(1):1–30, November 7, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388801711.

Woldringh:1980:DRR

[WGRZ80] C. L. Woldringh, N. B. Grover, R. F. Rosenberger, and A. Zaritsky. Dimensional rearrangement of rod-shaped bacteria following nutritional shift-up. II. Experiments with Escherichia coli Br. Journal of Theoretical Biology, 86(3):441–454, October 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380903446.

Werren:1980:PEM

[WGS80] John H. Werren, Mart R. Gross, and Richard Shine. Paternity and the evolution of male parental care. Journal of Theoretical Biology, 82(4):619–631, February 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380901824.

Woldenberg:1983:FOL

[WH83] Michael J. Woldenberg and Keith Horsfield. Finding the optimal lengths for three branches at a junction. Journal of Theoretical Biology, 104(2):301–318, September 21, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383904174.

Woldenberg:1986:RBA

[WH86] Michael J. Woldenberg and Keith Horsfield. Relation of branching angles to optimality for four cost principles. Journal of Theoretical Biology, 122(2):187–204, September 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386800819.

Wheatley:1984:IPD

[Whe84] D. N. Wheatley. Intracellular protein degradation: Basis of a self-regulating mechanism for the proteolysis of endogenous proteins. Journal of Theoretical Biology, 107(1):127–149, March 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384801253.

White:1980:TED

[Whi80a] R. Allen White. A theory for the estimation of DNA synthesis rates by flow cytometry. Journal of Theoretical Biology, 85(1): 53–73, July 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519380902805.

Whitehead:1980:MUC

[Whi80b] Edward P. Whitehead. Mixed and uniform co-operativity of ligand binding to multisite proteins: the co-operativity types allowed by the adair equation and conditions for them. Journal of Theoretical Biology, 87(1):153–170, November 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380902258.

Whitehead:1980:PSC

[Whi80c] Edward P. Whitehead. Protein symmetry and the co-operative ligand-binding behaviour predicted by allosteric Koshland models. Journal of Theoretical Biology, 86(1):45–82, September 7, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251938090065X.

White:1981:AIS

[Whi81a] James White. The allometric interpretation of the self-thinning rule. Journal of Theoretical Biology, 89(3):475–500, April 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381903635.

Whitehead:1981:CVT

[Whi81b] Edward P. Whitehead. Conditions for the various types of cooperativity allowed by the adair, equation and their relation with the algebraic character of binding polynomials. Journal of Theoretical Biology, 93(3):547–583, December 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902204.

Whittenburg:1982:EMC

[Whi82] Scott L. Whittenburg. The effect of the membrane compression wave on the light scattering spectrum from a protein monomerdimer equilibrium. Journal of Theoretical Biology, 96(2):151– 160, May 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519382902181.

Whittenburg:1983:MEL

[Whi83] Scott L. Whittenburg. Microviscosity and enzyme-ligand dissociation. Journal of Theoretical Biology, 100(4):675–683, February 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383903296.

White:1989:CMC

[Whi89a] R. A. White. Computing multiple cell kinetic properties from a single time point. Journal of Theoretical Biology, 141(4):429–446, December 19, 1989. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519389802292.

White:1989:AEB

[Whi89b] Thomas D. White. An analysis of epipubic bone function in mammals using scaling theory. Journal of Theoretical Biology, 139(3):343–357, August 9, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389802139.

Wheatley:1980:IPS

[WI80] Denys N. Wheatley and Marget S. Inglis. An intracellular perfusion system linking pools and protein synthesis. Journal of Theoretical Biology, 83(3):437–445, April 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380900508.

Wicken:1980:TTE

[Wic80] Jeffrey S. Wicken. A thermodynamic theory of evolution. Journal of Theoretical Biology, 87(1):9–23, November 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380902167.

Wicken:1985:OCM

[Wic85a] Jeffrey S. Wicken. An organismic critique of molecular darwinism. Journal of Theoretical Biology, 117(4):545–561, December 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938580237X.

Wicken:1985:TCS

[Wic85b] Jeffrey S. Wicken. Thermodynamics and the conceptual structure of evolutionary theory. Journal of Theoretical Biology, 117(3):363–383, December 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0022519385801491.

Wiese:1981:EAI

[Wie81] L. Wiese. On the evolution of anisogamy from isogamous monoecy and on the origin of sex. Journal of Theoretical Biology, 89(4):573–580, April 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381900291.

Wigler:1981:TEC

[Wig81] Paul W. Wigler. Trans effects in cell membrane transport. Journal of Theoretical Biology, 93(4):997–1008, December 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381903519.

Wiggins:1982:PMC

[Wig82a] Philippa M. Wiggins. A possible mechanism for the Ca-ATPase of sarcoplasmic reticulum. Journal of Theoretical Biology, 99(4):645–664, December 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382901928.

Wiggins:1982:PMK

[Wig82b] Philippa M. Wiggins. A possible mechanism for the na,k-ATPase. Journal of Theoretical Biology, 99(4):665–679, December 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251938290193X.

Wilkins:1981:ECR

[Wil81a] Adam S. Wilkins. Eukaryotic chromosome replication and the radial loop model. Journal of Theoretical Biology, 89(4):715– 717, April 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381900382.

Williams:1981:AGT

[Wil81b] Scott M. Williams. On the applicability of game theory to evolution: a response. Journal of Theoretical Biology, 91(4):603–605, August 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381902125.

Willson:1981:SEF

[Wil81c] Mary F. Willson. Sex expression in fern gametophytes: Some evolutionary possibilities. Journal of Theoretical Biology, 93 (2):403–409, November 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519381901120.

Williams:1982:SMC

[Wil82] Alan F. Williams. Surface molecules and cell interactions. Journal of Theoretical Biology, 98(2):221–234, September 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902600.

Wilson:1983:RT

[Wil83] David Sloan Wilson. Reply to Treisman. Journal of Theoretical Biology, 102(3):459–462, June 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383903818.

Wilson:1984:STE

[Wil84] John E. Wilson. Some thoughts on the evolutionary basis for the prominent role of ATP and ADP in cellular energy metabolism. Journal of Theoretical Biology, 111(4):615–623, December 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938480257X.

Williams:1986:CBP

[Wil86a] R. J. P. Williams. The connections between particle flow and mechanical, electrical and chemical work. Journal of Theoretical Biology, 121(1):1–22, July 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938680025X.

Wills:1986:SRP

[Wil86b] Peter R. Wills. Scrapie, ribosomal proteins and biological information. Journal of Theoretical Biology, 122(2):157–178, September 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519386800790.

Watari:1980:GMD

[WIN80] Hiroshi Watari, Yoshinori Isogai, and Hiroyasu Nlshikawa. Graphic method for determination of allosteric constants. Journal of Theoretical Biology, 83(2):209–214, March 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380902891.

Winne:1981:UFB

[Win81] D. Winne. Unidirectional fluxes and blood flow. Journal of Theoretical Biology, 93(4):987–996, December 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381903507.

Winfree:1989:EIC

[Win89a] A. T. Winfree. Electrical instability in cardiac muscle: Phase singularities and rotors. Journal of Theoretical Biology, 138(3): 353–405, June 3, 1989. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519389802000.

Winne:1989:EVD

[Win89b] Dietrich Winne. Effect of villosity and distension on the absorptive and secretory flux in the small intestine. Journal of Theoretical Biology, 139(2):155–186, July 21, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389800980.

Wiseman:1980:PIE

[Wis80] Lawrence L. Wiseman. Phase inversion in embryonic tissue engulfment experiments: Tissue size is not the important variable. Journal of Theoretical Biology, 83(2):365–366, March 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519380902994.

West:1989:SST

[WJS89] P. W. West, D. R. Jackett, and S. J. Sykes. Stresses in, and the shape of, tree stems in forest monoculture. Journal of Theoretical Biology, 140(3):327–343, October 9, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519389800906.

Wang:1987:NWJ

[WJT87] Zhi-Xin Wang, Ming Ju, and Chen-Lu Tsou. Number of ways of joining SH groups to form multi-peptide chain proteins. Journal of Theoretical Biology, 124(3):293–301, February 7, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387801170.

Welch:1981:CEA

[WK81] G. Rickey Welch and Tamás Keleti. On the “cytosociology” of enzyme action in vivo: a novel thermodynamic correlate of biological evolution. Journal of Theoretical Biology, 93(4):701–735, December 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381903362.

Welch:1988:CCM

[WKV88] G. Rickey Welch, Tamás Keleti, and Beáta Vértessy. The control of cell metabolism for homogeneous vs. heterogeneous enzyme systems. Journal of Theoretical Biology, 130(4):407–422, February 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519388802066.

Woyciechowski:1987:MMQ

[WL87] Michal Woyciechowski and Adam Lomnicki. Multiple mating of queens and the sterility of workers among eusocial hymenoptera. Journal of Theoretical Biology, 128(3):317–327, October 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519387800747.

Wohlgemuth:1983:MIM

[WM83] Andrew Wohlgemuth and George Markowsky. Mathematical immunogenetics I mathematics as language. Journal of Theoretical Biology, 102(3):411–424, June 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383903776.

Wynn:1986:CCGa

[WMR86] C. H. Wynn, A. Marsden, and B. Robson. Calculation of the conformation of glycosphingolipids 1. β-D-glucosyl-N-(2-D-hydroxyoctadecanoyl)-D-dihydrosphingosine (glucosyl-ceramide) and α-D-N0acetylneuraminyl-(2 → 3)-β-D-galactose. Journal of Theoretical Biology, 119(1):81–87, March 7, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386800522.

Weiss:1981:UPR

[WMS81] Steven D. Weiss, Pamela D. McNamara, and Stanton Segal. Uptake of proline by renal brushborder vesicles: a mathematical analysis. Journal of Theoretical Biology, 93(3):597–608, December 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251938190223X.

Wojcierowski:1983:EGM

[Woj83] Jacek K. Wojcierowski. Eukaryotic genome: Model considerations. Journal of Theoretical Biology, 103(3):379–403, August 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251938390293X.

Wolf:1980:SSA

[Wol80] Matthew B. Wolf. A simulation study of the anomalous osmotic behavior of red cells. Journal of Theoretical Biology, 83(4):687– 700, April 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519380901964.

Wolman:1985:DMC

[Wol85] Moshe Wolman. Different modes of cell death and related phenomena: a proposal for a meaningful terminology in relation to tests of lethality. Journal of Theoretical Biology, 114(2):343– 350, May 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385801107.

Wolf:1989:QPT

[Wol89] Fred Alan Wolf. On the quantum physical theory of subjective antedating. Journal of Theoretical Biology, 136(1):13–19, January 9, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519389801857.

Wong:1981:MEC

[Won81] Alan Y. K. Wong. A model of excitation-contraction coupling of mammalian cardiac muscle. Journal of Theoretical Biology, 90(1): 37–61, May 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381901211.

Wookey:1980:TCU

[Woo80] P. J. Wookey. A theory for the coupling of ultrastructural transformations of mitochondria to the activity of glycolysis. Journal of Theoretical Biology, 85(1):3–11, July 7, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380902775.

Weir:1985:GCM

[WOT85] B. S. Weir, T. Ohta, and H. Tachida. Gene conversion models. Journal of Theoretical Biology, 116(1):1–8, September 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519385801272.

Wiegel:1981:CPA

[WP81] Frederik W. Wiegel and Alan S. Perelson. Configurations of a polymeric antigen adsorbed to a b-cell membrane. Journal of Theoretical Biology, 88(3):533–568, February 7, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902824.

West:1984:WOM

[WP84] Ian C. West and Malcolm G. P. Page. When is the outer membrane of Escherichia coli rate-limiting for uptake of galactosides? Journal of Theoretical Biology, 110(1):11–19, September 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384800119.

Wilbur:1983:TBL

[WR83] W. John Wilbur and John Rinzel. A theoretical basis for large coefficient of variation and bimodality in neuronal interspike interval distributions. Journal of Theoretical Biology, 105(2):345–368, ???? 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519383800137.

Wynn:1986:CCGb

[WR86] C. H. Wynn and B. Robson. Calculation of the conformation of glycosphingolipids 2. G_{M1}- and G_{M2}-gangliosides. Journal of Theoretical Biology, 123(2):221–230, November 21, 1986. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801564.

Wright:1980:IAN

[Wri80] M. Wright. Involvement of an “all-or-none event” in the triggering of chromosome segregation and of S phase in procaryotic and eucaryotic cells. Journal of Theoretical Biology, 82(2):249–253, January 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519380901010.

Wright:1984:CMC

[Wri84] Barbara E. Wright. Constraints on the models of carbohydrate metabolism in the two cell types of Dictyostelium discoideum. Journal of Theoretical Biology, 110(3):445–460, October 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384801861.

Wake:1983:PSO

[WRW83] David B. Wake, Gerhard Roth, and Marvalee H. Wake. On the problem of stasis in organismal evolution. Journal of Theoretical Biology, 101(2):211–224, March 21, 1983. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383903351.

Weinberg:1981:XIV

[WS81] Josette Weinberg and V. Em. Sahini. Xylidine isomers. VII. structure-hepatotoxic activity relationships of some xylidines. Journal of Theoretical Biology, 88(1):153–159, January 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381903337.

Wheals:1982:UAM

[WS82] Alan Wheals and Bernard Silverman. Unstable activator model for size control of the cell cycle. Journal of Theoretical Biology, 97(3):505–510, August 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382903782.

Ward-Smith:1984:AEC

[WS84] A. J. Ward-Smith. Aerodynamic and energetic considerations relating to undulating and bounding flight in birds. Journal of Theoretical Biology, 111(2):407–417, November 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384802192.

Williams:1985:SCE

[WS85a] Scott M. Williams and Curtis Strobeck. Sister chromatid exchange and the evolution of rDNA spacer length. Journal of Theoretical Biology, 116(4):625–636, October 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800928.

Witkiewicz:1985:MDA

[WS85b] H. Witkiewicz and M. Schweiger. A model of λ DNA arrangement in the viral particle. Journal of Theoretical Biology, 116(4):587– 605, October 21, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385800898.

Wilson:1989:RS

[WS89] David Sloan Wilson and Elliott Sober. Reviving the superorganism. Journal of Theoretical Biology, 136(3):337–356, February 8, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519389801699.

Welch:1983:EVE

[WSMP83] G. Rickey Welch, Béla Somogyi, János Matkó, and Sándor Papp. Effect of viscosity on enzyme-ligand dissociation II. Role of the microenvironment. Journal of Theoretical Biology, 100(2):211– 238, January 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/002251938390348X.

Wang:1987:KSR

[WT87] Zhi-Xin Wang and Chen-Lu Tsou. Kinetics of substrate reaction during irreversible modification of enzyme activity for enzymes involving two substrates. Journal of Theoretical Biology, 127(3):253–270, August 7, 1987. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387801066.

Wu:1988:MNS

[Wu88] Tai Te Wu. Matching nucleotide sequences of human antibodies with other known sequences. Journal of Theoretical Biology, 131(2):231–234, March 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938880239X.

Weihs:1984:OAE

[WW84] D. Weihs and P. W. Webb. Optimal avoidance and evasion tactics in predator–prey interactions. Journal of Theoretical Biology, 106(2):189–206, January 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519384900195.

Wen:1988:TDD

[WWG^{+}88] G. B. Wen, S. Weinbaum, P. Ganatos, R. Pfeffer, and S. Chien. On the time dependent diffusion of macromolecules through transient open junctions and their subendothelial spread. 2. Long time model for interaction between leakage sites. Journal of Theoretical Biology, 135(2):219–253, November 21, 1988. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519388800766.

Winzor:1982:QCC

[WWN82] D. J. Winzor, L. D. Ward, and L. W. Nichol. Quantitative considerations of the consequences of an interplay between ligand binding and reversible adsorption of a macromolecular solute. Journal of Theoretical Biology, 98(2):171–187, September 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902570.

Young:1989:MMP

[YA89] B. A. Young and M. A. Ashley. A mathematical model for predicting relative muscle force with a perturbation analysis of selected muscle parameters. Journal of Theoretical Biology, 138(2):213– 223, May 22, 1989. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519389801407.

Yagil:1985:SCS

[Yag85] Gad Yagil. On the structural complexity of simple biosystems. Journal of Theoretical Biology, 112(1):1–23, January 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801144.

Yamamura:1987:MCB

[Yam87] Norio Yamamura. A model on correlation between precopulatory guarding and short receptivity to copulation. Journal of Theoretical Biology, 127(2):171–180, July 21, 1987. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519387801285.

Yates:1986:OPA

[Yat86] George T. Yates. Optimum pitching axes in flapping wing propulsion. Journal of Theoretical Biology, 120(3):255–276, June 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386801990.

Yadav:1982:QCSa

[YBBL82a] Janardan Singh Yadav, Gerhard Barnickel, Hans Bradaczek, and Harald Labischinski. Quantum chemical studies on the conformational structure of bacterial peptidoglycan. II. PCILO calculations on the mono-saccharides. Journal of Theoretical Biology, 95(1):151–166, March 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902946.

Yadav:1982:QCSb

[YBBL82b] Janardan Singh Yadav, Gerhard Barnickel, Hans Bradaczek, and Harald Labischinski. Quantum chemical studies on the conformational structure of bacterial peptidoglycan. III. CNDO and INDO calculations on the n-acetyl-giucosamine. Journal of Theoretical Biology, 95(1):167–179, March 7, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902958.

Yadav:1982:QCSc

[YBBL82c] Janardan Singh Yadav, Gerhard Barnickel, Hans Bradaczek, and Harald Labischinski. Quantum chemical studies on the conformational structure of bacterial peptidoglycan. IV. PCILO calculation on the di-saccharide part of the glycan moiety. Journal of Theoretical Biology, 95(2):285–303, March 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382902454.

Ydenberg:1989:AAD

[YC89] R. C. Ydenberg and Colin W. Clark. Aerobiosis and anaerobiosis during diving by western grebes: An optimal foraging approach. Journal of Theoretical Biology, 139(4):437–447, August 22, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519389800645.

Ycas:1982:MT

[Yca82]  M. Ycas. A message from triceratops?  Journal of

Theoretical Biology, 96(1):39–47, May 7, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382901540.

Yokoyama:1982:ECI

[YE82] Shozo Yokoyama and Alan R. Empleton. Effect of cultural inheritance of reproductive compensation on the incidence of a sex-linked lethal disease. Journal of Theoretical Biology, 99(2):389–395, November 21, 1982. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382900121.

Yuan:1989:QMS

[YHJ89] C. J. Yuan, A. J. Hopfinger, and M. E. Johnson. QSAR and molecular shape analysis of aryl-substituted alanine analogs as antigelling agents. Journal of Theoretical Biology, 141(1):41–52, November 8, 1989. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519389800062.

Yavachev:1988:WDH

[YI88] Libert Yavachev and Ivan Ivanov. What does the homology between E. coli tRNAs and RNAs controlling ColE1 plasmid replication mean? Journal of Theoretical Biology, 131(2):235–241, March 21, 1988. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519388802406.

Yamaguchi:1980:TSS

[YKS80] Y. Yamaguchi, K. Kometani, and H. Shimizu. A theoretical study of synchronizations of two myocardial cells. Journal of Theoretical Biology, 82(2):231–248, January 21, 1980. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380901009.

Yadav:1981:QCS

[YLBB81] Janardan Singh Yadav, Harald Labischinski, Gerhard Barnickel, and Hans Bradaczek. Quantum chemical studies on the conformational structure of bacterial peptidoglycan. I. MNDO calculations on the glycan moiety. Journal of Theoretical Biology, 88(3):441– 457, February 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381902769.

Young:1989:PAC

[YM89] Richard A. Young and Richard T. Marrocco. Predictions about chromatic receptive fields assuming random cone connections. Journal of Theoretical Biology, 141(1):23–40, November 8, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519389800050.

Yuasa:1986:IAC

[YMIH86] Hiroaki Yuasa, Yusei Miyamoto, Tatsuji Iga, and Manabu Hanano. Intestinal absorption by carrier-mediated transports: Two-dimensional laminar flow model. Journal of Theoretical Biology, 119(1):25–36, March 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386800480.

Yamamoto:1984:KIM

[YMS84] Yoshio Yamamoto, Hiroyuki Mioh, and Hiroshi Shimizu. A kinetic Ising model for dynamical behavior of anisotropic system — its significance in streaming phenomena. Journal of Theoretical Biology, 109(3):373–391, August 7, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384800879.

Yockey:1981:ROG

[Yoc81a] H. P. Yockey. Rebuttal of “overlapping genes and information theory”. Journal of Theoretical Biology, 91(2):381–382, July 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/002251938190240X.

Yockey:1981:SOO

[Yoc81b] Hubert P. Yockey. Self organization origin of life scenarios and information theory. Journal of Theoretical Biology, 91(1):13– 31, July 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0022519381903702.

Yodzis:1981:SAC

[Yod81] P. Yodzis. The structure of assembled communities. Journal of Theoretical Biology, 92(2):103–117, September 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381903866.

Yodzis:1984:SAC

[Yod84] Peter Yodzis. The structure of assembled communities II. Journal of Theoretical Biology, 107(1):115–126, March 7, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384801241.

Yokoyama:1981:FSE

[Yok81] Shozo Yokoyama. Family size and evolution of Rh polymorphism. Journal of Theoretical Biology, 92(2):119–125, September 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381903878.

Yokoyama:1983:EFG

[Yok83] Shozo Yokoyama. Equilibrium frequencies of α-globin genes. Journal of Theoretical Biology, 100(1):173–180, January 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519383901005.

Yokoyama:1984:SSH

[Yok84] Shozo Yokoyama. Social selection in human populations: Differential modification of the fitness by the sex of an affected parent. Journal of Theoretical Biology, 109(3):453–459, August 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384800922.

Yokoyama:1985:SSH

[Yok85] Shozo Yokoyama. Social selection in human populations: Sufficient conditions for protection of deleterious alleles in a subdivided population. Journal of Theoretical Biology, 115(3):455– 465, August 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385802034.

Yorke:1981:SPS

[Yor81] Ellen D. Yorke. Sensitivity of pigeons to small magnetic field variations. Journal of Theoretical Biology, 89(3):533–537, April 7, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381903672.

Yoshino:1986:KMV

[Yos86] Masataka Yoshino. Kinetic meaning of I_{0}.5 and A_{0}.5 values in allosteric enzymes. Journal of Theoretical Biology, 123(4):499–502, December 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002251938680217X.

Young:1981:SCC

[You81] J. P. W. Young. Sib competition can favour sex in two ways. Journal of Theoretical Biology, 88(4):755–756, February 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519381902496.

Young:1985:RIE

[You85] John Ding-E Young. Role of ionic events in the triggering of phagocytosis. Journal of Theoretical Biology, 116(3):475–478, October 7, 1985. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519385802836.

Ypey:1980:MET

[YVIG80] D. L. Ypey, W. P. M. VanMeerwijk, C. Ince, and G. Groos. Mutual entrainment of two pacemaker cells. A study with an electronic parallel conductance model. Journal of Theoretical Biology, 86(4):731–755, October 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380903082.

Zaciu:1981:MSC

[Zac81] Cornelia Zaciu. Microscopical significance of the critical characteristics of stimuli in excitation processes, as revealed by spectral analysis. Journal of Theoretical Biology, 91(2):241–254, July 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381902319.

Zajicek:1982:IPH

[Zaj82] G. Zajicek. The intestinal proliferon hypothesis. Journal of Theoretical Biology, 97(2):337–340, July 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382901102.

Zajicek:1986:AKE

[Zaj86] G. Zajicek. The application of kinematic equations for the study of cell turnover. Journal of Theoretical Biology, 120(2):141– 149, May 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519386801709.

Zamir:1981:TDA

[Zam81] M. Zamir. Three-dimensional aspects of arterial branching. Journal of Theoretical Biology, 90(4):457–476, June 21, 1981. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
002251938190299X.

Zamir:1986:CAA

[Zam86] M. Zamir. Cost analysis of arterial branching in the cardiovascular systems of man and animals. Journal of Theoretical Biology, 120(1):111–123, May 7, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386800212.

Zamir:1984:CDO

[ZB84] M. Zamir and D. C. Bigelow. Cost of departure from optimality in arterial branching. Journal of Theoretical Biology, 109(3):401– 409, August 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384800892.

Zink:1980:SCAa

[ZBL80a] Sandra Zink, Charles L. Beckel, and Allen C. Larson. Study of conformations of the adenosine phosphates. I: Adenosine monophosphate (AMP). Journal of Theoretical Biology, 85(2):293–303, July 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519380900223.

Zink:1980:SCAb

[ZBL80b] Sandra Zink, Charles L. Beckel, and Allen C. Larson. Study of conformations of the adenosine phosphates. II: Adenosine tri- and diphosphate (ATP, ADP). Journal of Theoretical Biology, 85(2):305–323, July 21, 1980. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519380900235.

Zhang:1989:SDS

[ZC89] Chun-Ting Zhang and Yue-Qin Cen. The study of DNA sequences by their sequences of twist angles. Journal of Theoretical Biology, 138(4):457–465, June 22, 1989. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S002251938980044X.

Zeeman:1981:DEA

[Zee81] E. C. Zeeman. Dynamics of the evolution of animal conflicts. Journal of Theoretical Biology, 89(2):249–270, March 21, 1981. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/0022519381903118.

Zimanyi:1982:CLI

[ZG82] László Zimányi and Gyözö Garab. Configuration of the light induced electric field in thylakoid and its possible role in the kinetics of the 515 nm absorbance chance. Journal of Theoretical Biology, 95(4):811–821, April 21, 1982. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519382903551.

Ziehe:1985:SUM

[ZG85] Martin Ziehe and H.-R. Gregorius. The significance of overand underdominance for the maintenance of genetic polymorphisms: I. Underdominance and stability. Journal of Theoretical Biology, 117(3):493–504, December 7, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385801570.

Zimanyi:1989:CEF

[ZG89] László Zimányi and Győző Garab. Configuration of the electric field and distribution of ions in energy transducing biological membranes: Model calculations in a vesicle containing discrete charges. Journal of Theoretical Biology, 138(1):59–76, May 7, 1989. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S002251938980178X.

Zhen-gang:1984:HAC

[ZgQy84] Li Zhen-gang and Wu Qiu-ying. A hypothesis on the action of chromosomal genes. Journal of Theoretical Biology, 108(4):645– 661, June 21, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384800831.

Zhadin:1984:RPC

[Zha84] M. N. Zhadin. Rhythmic processes in the cerebral cortex. Journal of Theoretical Biology, 108(4):565–595, June 21, 1984. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519384800806.

Zachmann:1984:MSA

[ZL84] David W. Zachmann and Jesse A. Logan. Mathematical stability analysis for an age-structured population with finite life span. Journal of Theoretical Biology, 111(1):131–148, November 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519384802015.

Znojil:1983:IEM

[Zno83] V. Znojil. An ignored error in the mathematical formulation of erythrocyte survival curves. Journal of Theoretical Biology, 102(4):625–628, June 21, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0022519383903958.

Zolman:1983:PCO

[Zol83] Joseph C. Zolman. Positive co-operativity in the membrane receptor mediated response. Journal of Theoretical Biology, 104(1):1–6, September 7, 1983. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0022519383903971.

Zielenkiewicz:1984:PPR

[ZR84] Piotr Zielenkiewicz and Andrzej Rabczenko. Protein-protein recognition: Method of finding complementary surfaces of interacting proteins. Journal of Theoretical Biology, 111(1):17–30, November 7, 1984. CODEN JTBIAP. ISSN 0022-5193 (print),

1095-8541 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0022519384801939.

Zielenkiewicz:1985:SIS

[ZR85] Piotr Zielenkiewicz and Andrzej Rabczenko. Searching for interacting surfaces of proteins-the improved method. Journal of Theoretical Biology, 116(4):607–612, October 21, 1985. CO-DEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519385800904.

Zanardo:1986:AGF

[ZR86]  Alberto Zanardo and Martino Rizzotti.  Axiomatization

of genetics 2. Formal development. Journal of Theoretical Biology, 118(2):145–152, January 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0022519386801308.

Zuckerkandl:1986:QPP

[Zuc86] Emile Zuckerkandl. Quantitative protein profiling: Determining lexotypes. Journal of Theoretical Biology, 121(2):185–198, July 21, 1986. CODEN JTBIAP. ISSN 0022-5193 (print), 1095-8541

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0022519386800923.